A Complete Bibliography of Publications in Limnology and Oceanography: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.08

Title word cross-reference

$79.95 [Sta01]. $89.95 [Bel01]. > 110 [GSL^{+}07]. _{3}^{−} [BARS01]. ^{12} [VW08]. ^{13} [AHM^{+}08, ASS03, BCC^{+}04, BLM^{+}06, CCKP02, CWC^{+}07, GRB09, GP06a, GSH04, HFR08, KTG^{+}06, KCP^{+}04, LBM^{+}04, MBCD04, MdG08, MBB^{+}00, RRSC01, RB01a, RHM04, SRL^{+}09, SKC^{+}01, SSL05, TMYH02, TB05, VW08, VMS^{+}04, VR01, VvOBM06, VMS09, WIK08, dGMH^{+}08]. ^{14} [CEBMF05, FKCF00, MBCD04, RRSC01, RB01a, SL02, YPBL07]. ^{15} [AS00a, BBP^{+}06, CWC^{+}07, GRB09, GHB^{+}05, GVT^{+}09, HB05, HTS^{+}09, HVP^{+}07, HWL^{+}00, LBM^{+}04, MM05, MBCD04, MVW^{+}04, MHS^{+}09, RRSC01, RHM04, SRG00, SHBM01, SSL05, TMYH02, TMS^{+}00, TB00, UMY^{+}02, VR01, VEMM07]. ^{16} [MWH01]. ^{18} [RV04, SRL^{+}09, YPBL07, vdSvdSG^{+}00]. ^{210} [JBvHS08, NSBA01, SCARM00]. ^{222} [CDBC00]. ^{226} [BSA^{+}00]. ^{234} [MWE^{+}03, QSH^{+}02]. ^{238} [MWE^{+}03]. ^{239,240} [NSBA01]. ^{31} [ART^{+}06, HRS04]. 32 [BNK02]. ^{33} [BNK02, NKJM06]. ^{34} [AHM^{+}08]. ^{64} [CBvEK03]. ^{7} [JBvHS08]. _{12} [BSR^{+}07, TS08]. _{2} [BD04, BSM01, Bat02, BDS^{+}04, BMH^{+}08, BS03, BARS01, CPGC09, CTCT03, lLCChTO08, CPL^{+}04, CR06, CRF01, DLZ^{+}09, DJB^{+}07, EZA^{+}05, FH03, FPZ09, FMG^{+}08, GEJ^{+}02, GR01, GGH07, GTK^{+}01, GGRJ03, GBFJ07, GSB^{+}09, GAPR06, GZY^{+}08, HVP^{+}07, HFZ^{+}07, KLS^{+}06, KLS^{+}07, KHQ01, KDRR09, LGJ02, LFFC06, MBG^{+}05, MdG08, MVBM04, MB05, ORH01, PCD^{+}08, RBC00, RG04, SNPK^{+}09, SH05b, VTO^{+}05, WC04, WKH^{+}06, WG04]. _{3} [MB01]. _{3}^{−} [Alt01, HWL^{+}00]. _{4} [CDBC00, LFFC06]. _{org} [WIK08]. β [AH02, BGHA04, SWS^{+}03]. : [ALRM08, Ano04, AD01, BSMR08, Ber01, BBP^{+}06, ETCP03, FQC^{+}07, HW07a, KHKS05, LG04, LMT05, MBLP01, NHPHH08, RFM07, SAS08, SAE^{+}08, SF04, VW08, VPA02, WGKE04]. Δ [MBCD04, AHM^{+}08, ASS03, BCC^{+}04, BBP^{+}06, CWC^{+}07, GRB09, GP06a, KTG^{+}06, LBM^{+}04, MM05, MdG08, RHM04, SRL^{+}09, SSL05, TMYH02, TB00, TB05, UMY^{+}02, VR01, VMS09, WIK08]. f [BSM01]. μ [GSL^{+}07, HBH00]. ω6 [PWP07]. Q [ZB01].

-dimethylsulfoniopropionate [SWS^{+}03]. -factor [ZB01]. -fixation [HVP^{+}07]. -glucosidase [AH02]. -proteobacteria [BGHA04]. -stabilizing [WKSM08].

0 [Bel01, Boy01, Bri01, Caf01, Dow02, Elm02, Fou02, Jun01, Kne02, Pom01, Ris02, SG02, Sta01, Wro01]. 0-12-389845-5 [Sta01]. 0-471-29993-6 [Pom01]. 0-521-66184-6 [Fou02]. 0-521-78367-4 [Bri01]. 0-632-05566-9 [Wro01]. 0-691-00493-5 [Boy01]. 0-691-04929-7 [Jun01]. 0-8493-0008-8 [Kne02]. 0-8493-2026-7 [Ris02]. 0-8493-9062-1 [Elm02]. 0-8493-9063-X [Bel01]. 0-87590-271-5 [Caf01]. 0-87590-272-3 [Dow02]. 0-87893-057-4 [SG02]. 000- [HBH00]. 000-m [KCCE06]. 000-year [ES00]. 0214 [Wan02]. 0214-8358 [Wan02].

1 [Elm02, Mil02]. 1-57003-431-1 [Mil02]. 101 [Ano04, KSWC^{+}03]. 110m [AGL03]. 17-year [SRKK08]. 170^{o} [ZMK^{+}08]. 1974-99 [LIE06]. 1998 [TAB^{+}00]. 1^{o} [ZMK^{+}08].

2-week-old [AGKK08]. 2.6 [KT02]. 2005 [RFM07]. 20th [AS06, OKO^{+}01]. 21st [Wan02]. 234Th [RFW08]. 234Th-based [RFW08].

3 [Dow02]. 3-mercaptopropionic [HMB06].

4 [Bri01, Gou01, SG02]. 46^{o} [PLW^{+}08].

5 [Boy01, Caf01, Sta01]. 5-yr [Pil04].

6 [Fou02, Pom01]. 6-week-old [AGKK08]. 65 [Wan02].

7 [Jun01, Ris02].

8 [Kne02]. 8-year [SKBC01]. 8358 [Wan02].

9 [Wro01]. 90 [Gou01]. 90-804341-2-4 [Gou01].

abiotic [ESA06, HWYL07]. above [JPK^{+}05]. abrupt [Boy01, YYB^{+}05]. absolute [DKKA09, HH03]. Absorbance [HMKZ02, BK01b, BK01c, MPD^{+}06]. Absorption [HSR^{+}08, BS02a, BS04, BVK02, CLC02, ETCP03, EMIP05, Fin01, HMKZ02, LMPY00, LBD^{+}09, MV00, MN04, SSMM00, WV00, DdG09a]. Abundance [BK03a, CG09, CDD^{+}03, Kiø00, vdSvdSG^{+}00, BLS^{+}06, CMH^{+}08, CW02, DGV01b, DPC^{+}06, GCG^{+}02, GRB02, GZY^{+}08, Har02, HOFD01, HVP^{+}07, IHHH00, INU04, KDO09, KN07, LHB04, MRM07, ML05, NTDV06, NML^{+}05, PD02, RL07, TA03, Thi00, WSHW05]. abundances [MCC02, RRSC01, SF05]. abundant [KS04a]. abyssal [DPC^{+}08, HBH^{+}07, SKBC01, SRKK08, VRS09]. Academic [Sta01]. Acartia [BB04b, Møl07, XW02]. accelerating [SABA02]. accelerations [dlFSIN08]. accelerator [FKCF00]. acceptors [MED^{+}08]. Accessory [TCB^{+}00]. acclimated [BARS01]. Acclimation [SWN08, vdPVB07, MGW^{+}05]. acclimatization [PGC07]. acclimatized [NRB08]. accompanying [TKAW03]. Accounting [MFF07]. accretion [Cra07]. Accumulation [CSK^{+}00, GW03, HP03, LNJL09, AFH07, Ber03, BRB^{+}01, BSA^{+}00, BSL^{+}07, CSWD07, Ceb02, Cra07, CHT02, Dal03, KKKT07, PNPA04, PLK^{+}05, RKH03, SSK06, SWCL01, Shi04, SlBWC^{+}00, SF03b, SF03c, SH00, TCD^{+}03, TVL^{+}00, TMB07, VM06]. accurate [MWE^{+}03]. accurately [BEM^{+}02]. Acetate [HST^{+}02]. aciculiferum [RL01a]. acid [BKK^{+}06, BMNB^{+}06, BJD05, BDK^{+}08, EMWL06, FHDKR03, FGM^{+}03, FSBB05, HMB06, KF01, LBC^{+}04, LLC^{+}09, LMZB08, MHRW02, MGL02, PWP07, PTLJ01, PF01b, PAS^{+}05, QSH^{+}02, RB06, SR07, Tho09, TAB^{+}00, THH02, THL^{+}09, TGFPT08, VVSM09, WTC^{+}05]. acid-sensitive [FHDKR03]. acidic [KT02, PTKP00]. acidification [AG03b, FHDKR03, GRV^{+}09, RG08, VMKN03, WMRK06]. acidified [BRB^{+}01, JBH^{+}04, KHB^{+}00, KHK^{+}03]. acidity [PTKP00]. acidity-driven [PTKP00]. acids [AP00b, AHR^{+}01, BB05, GvRE^{+}02, KAM04, KF01, KK00, KLAF04, MGW^{+}05, PPH03, PBGV04, RBMN03, RS01, SO07, SAP^{+}06, Shi04, TLS01, TS06, TvAVH06, TTG^{+}09, VW01, WBvE02, WvE04, WV00, vE02]. acoustic [HM03b, KJH03]. acquisition [KDRR09, MGW^{+}05, PQP05, RRBS03, RRS06, SRNW06, SKM05, TXH05, Tor00, TRM00]. Acropora [OFKH01, SE06]. Across [SVPK^{+}09, BNM^{+}09, CSK^{+}00, CFO^{+}00, CFG07, DET04, DMEPA04, DJG^{+}06, HMR^{+}07, JBM^{+}07, KB07, LES^{+}05, LMMW03, LZ09, MCJ^{+}06, NCPd08, ORH01, RHJ02, SC00, SLHK02, TLS01]. across-shelf [CFG07]. acrylic [EMWL06]. acting [HMJ04]. active [Dix00, HMS^{+}04, RCL^{+}02, VMH^{+}01, vdPAJ^{+}06]. activities [EMPI07, FSS^{+}00, SHA08]. Activity [CJ08, Bat00, DP01, DA03, DR06, EPS00, GCG^{+}02, GGRJ03, GSL^{+}07, HWL^{+}00, KHP^{+}06, KED^{+}07, MT06, MGF03, NDCP06, OS05, PS06, RORR^{+}09, RRH^{+}03, RvBB02, RGRPD04, SXLM06, TXH05, TCR03, TPFS09, TLvAH06, WG04, WdMD^{+}00, WSHW05, dBSN^{+}06]. acuminata [SGG^{+}08]. Adaptation [CMCB^{+}09, HHM^{+}09, SW05a, VVD^{+}09]. Adaptive [DHKH05, LN04, Nor04, RL01a]. addition [CPCK00, CCKP02, EVW07, EGP08, HTS^{+}09, HMR^{+}07, KCP^{+}04, MVW^{+}04, MHS^{+}09, OAA^{+}06, TMS^{+}00, WWMB05]. additions [Car01, HB05, LTM01]. additive [SA09]. Addressing [AD02, CG08]. adiabatic [GBH^{+}02]. Adirondack [SDM^{+}09]. adjacent [THL^{+}09]. adjustment [HRA07]. Adriatic [BZH00]. adult [HL06, JL08, NA05]. adult-juvenile [HL06]. adults [AMS^{+}02]. advances [Sch06b]. advantages [AD07]. advecting [YM03]. advection [AKG^{+}05, DAL02, JFH^{+}05, JHW05, MRK02, PHK02, RB09, STK^{+}04]. Advective [CR06, HR00b, MGGM06, PH03, CWG^{+}07]. Aeration [GHWG04]. Aerobic [LBG^{+}09, PM08, GKS^{+}02, SF05, SGT^{+}06]. aeruginosa [GPAP^{+}04, HTWG08, KJ00, Lür03, RSWH04, SWH^{+}05]. affect [MM03, PSH^{+}07, RKK09, Shi04, TKM07]. affected [IR06, LLDS06, MTC^{+}06]. affecting [KRZ06]. affects [FH09, LFPH07, LB00, MACB07, SFC^{+}00]. affinity [JGSB07, TRT03]. Africa [LJ09, CWJ09, HHT01, MRK02, MVD05, Pil04, VBH^{+}05, WBB09]. African [JM08, Pil04]. after [HLK^{+}07, JPJS06, OFKH01, SK04]. after-effect [OFKH01]. against [BH02, BJL01, KSP07]. Agardh [WLP^{+}07]. age [EWE02, HH05, LBC06, MCLM02, RM06]. Agent [HKNG08]. Agent-based [HKNG08]. agents [BVWB07, SFC^{+}00]. ages [BW01, MBCD04]. Aggregate [KKT05, PG00, VOG^{+}08]. aggregates [All00, GP00, GP01, GSL^{+}07, GSJ^{+}09, HMT^{+}08, Kiø00, LWU^{+}05, PG00, Plo01, PHK02, PP07, WSHN00]. aggregating [DG06]. aggregation [Ham02]. Aging [BLJ09, CW01, GRdL^{+}08]. Aglaura [CCGH05]. agricultural [AD01, FL06, HBM08, SNFT07]. agriculture [QLD^{+}02]. AHOD [ME06]. Aiptasia [GCG05]. Air [BMH^{+}08, BSM01, Bat02, BMD^{+}04, KSS^{+}07, LD01, SLHK02]. air-sea [BSM01]. Airborne [APH^{+}03, DZL^{+}03, HTWG08]. airsheds [Sch09]. akashiwo [BGC04a, LCS02, SNM^{+}08, TDT01]. Aktun [vHRB^{+}09]. al. [AP00a, EW03]. Alabama [CF01]. alarm [LBT06]. Alaska [BLSQ07, BB07, CGD08, ML05, PDS07, SMO07]. Alaskan [BLS^{+}06, MFK^{+}09]. albedos [CLL^{+}03]. Alboran [GST^{+}01]. aldehydes [CWPL05, DKt08]. aldose [HY01]. Alexander [Caf01]. Alexandrium [AKG^{+}05, DA03, JTL01, MWM^{+}04, SCP08]. alga [BGR^{+}09, BGC04a, JM02, LCS02, Lür03, MHDH00, PSJ00, SH03, VVD^{+}09]. algae [AGL03, BS02b, Boe00, BLM^{+}06, Chi00, DVV^{+}07, EPS00, EMIP05, HFLD03, HFR08, HFR09, KHKS05, LLB07, QYEA06, SSM^{+}04, SBZ^{+}09, TSG^{+}01, TMA03, VZM07, VEMM07, VAS02, vE02]. Algal [KL01, LTT07, PCvRR00, AP00c, BAPMMB01, BKM05, BKK^{+}06, BMNSK00, BLSQ07, CR08, CHH07, CWJ09, DGM^{+}05, EJL^{+}06, FKS^{+}01, Fin04, FBW^{+}08, HTS^{+}03, LBW^{+}08, LU08, MSAC06, MSHE09, MSVAC04, MLSS05, PS06, PMK08, RC07, SSLE02, SPM04, SEC08, SDH02, SJW03, TMP02, VMS^{+}04, WL06, YCM^{+}01, ZHL06]. algal- [PS06, SPM04]. algal-bacterial [MLSS05, VMS^{+}04]. algorithm [Mor03, SM01]. alike [TBG01]. alive [Duf09]. Alkaline [NDCP06, DP01, DR06, RRH^{+}03, SXLM06, TXH05]. alkalinity [KLC06, KHK^{+}03, WKH^{+}06]. alkalinity-dissolved [WKH^{+}06]. Allelopathic [HH04, PMK08, SEL^{+}02]. allelopathy [KHNV05]. Aller [Mil02, Mil02]. Alley [Boy01]. Allocation [OCNH00, ACW02]. Allochthonous [Jan03, BLJ09, JMBJ01, KJMJ03, KJJ07, KAJ08, KCP^{+}04, MID^{+}07, VVSM09, WR00, WCB^{+}06, vWSBT08, vWBT09]. allometric [CMH^{+}08]. allow [GF02, NSBA01]. Almeria [GST^{+}01]. Almeria-Oran [GST^{+}01]. ALOHA [BK03b, CDK02, Gre09]. along [BMT^{+}04, BdG02, BFCLÁ08, CS06, CLS05, CW02, FF04, FTC^{+}06, FTY^{+}06, GCM^{+}09, GK03, HOFD01, JCBA08, KF01, KVR03, LFFC06, MH00, PDS07, PHMR^{+}07, SIH^{+}06, SGB04, SMA^{+}06, SMD^{+}04, SMP05, SNFT07, SWSJ02, StLHB^{+}07, SLL^{+}04, TN09, dGB02]. alongshore [GDK03]. Alpine [MLRMR09, CPGC09, DSW05, EPP^{+}08, KLWS08, LRMV04, MMCW09, PS06, SK05, SCV04, TS06, WW00]. alpinus [Hol00]. ALPS [DIMP08]. alter [GV09]. alteration [Ham05, VW08]. alternate [Fra01]. Alternative [BCSW07, BP03, Law03, SD05, TA03]. alters [LW01b, PSH08]. Altitude [LLK05, BSS^{+}01, TLS01]. Altitude-dependent [LLK05]. alum [RHW00]. aluminum [KMHN07, VMKN03]. always [HD03b]. Amazon [AHR^{+}01, FSM^{+}07, HMT^{+}00, RLC^{+}01, TSMB05, TB09]. ambient [BCDH09]. ambigua [HWA03]. ambiguity [HMKZ02]. ambush [HK06]. ambush-feeding [HK06]. America [LBEP06, RBJ^{+}06]. American [Caf01, Dow02, MVD05, SFB^{+}09]. americana [CF05, NA06, NA09, SJ08]. americanus [Ann05]. americium [FNLG00]. amino [AHR^{+}01, BK03a, GvRE^{+}02, KF01, KLAF04, MGW^{+}05, MGL02, PTLJ01, PPH03, PF01b, RS01, Shi04, SR07, TLS01, TS06, TvAVH06, WV00]. ammonia [RlBW01]. Ammonium [AW00a, AW00b, CRJ08, CT02, SH07, Ano04, Ano08a, EPD09, GMA^{+}06, GMRVFP02, HLW^{+}07, JKGT08, KD02, KSWC^{+}03, MPB^{+}08, TDJ^{+}06, WPD^{+}01, YTVR07]. ammonium-supported [Ano04, KSWC^{+}03]. amoebae [vHRB^{+}09]. among [AFB01, CT04, CGH^{+}04, CPCLS04, GA01, LVC^{+}00, LSK07, SD05, SAS^{+}03, SMA00, SH04, TJV01, TLS01, TS08, VK06, WGKE04, Wey04]. among- [LVC^{+}00]. Amount [RHW00]. amounts [BLM^{+}06]. amphibian [CGWK^{+}09]. Amphipod [CT06, SDC00]. Amphipoda [ZB03b]. amphipods [BL07]. Amphiura [VTR03]. Amphiuridae [VTR03]. amplification [Ano08b, CSG^{+}06, DMS^{+}03]. amplifies [MSVS^{+}07]. amurensis [CTCT03]. Amyot [SH10]. Anabaena [HKNG08]. anadromous [CPB^{+}08]. Anaerobic [EPD09, HLW^{+}07, RG04, TDJ^{+}06, BLWB08, GKS^{+}02, GFJ^{+}02, JKGT08, KB07, TKBJ05]. analyses [CCKC05, DPC^{+}08, SHS03a, TMYH02, WST06]. Analysis [TMS^{+}00, WKH^{+}06, AD00b, ACW02, BMNSK00, BDK^{+}06b, Cal01, DKRZ06, FBW^{+}08, FTC^{+}06, FTY^{+}06, GGP01, GBFJ07, GJS01, GSJ^{+}09, HBG00, HM03b, JBM^{+}07, JP03, KR08, KJH03, PDH06, PE07, RA06, STST07, STP^{+}07, SM05a, SM05b, WMW^{+}08, YJMT08, vODL^{+}09]. anammox [HLK^{+}07, JHW^{+}07, RGRPD04, TN09]. Anatomy [ESI01]. anchialine [vHRB^{+}09]. ancient [MSVS^{+}07, Gou01]. Andean [CMC^{+}09]. Andes [AMW^{+}07, TSMB05]. anemones [LLF^{+}09]. angles [ZMK^{+}08]. anglica [GK03]. anguillarum [GKS^{+}02]. anhydrase [EPS00, MT06]. Anna [Gou01]. annandalei [INU04]. Annual [ASI^{+}02, JCC02, SP09, Lee01, LLB04, MVC^{+}05, RZMB05, WCGBG07]. Anomalously [TA03]. anophagefferens [BRL03, GRB02, MGL02]. Anostraca [GSA^{+}05]. anoxia [MJ03]. anoxic [ART^{+}06, AVJ^{+}00, BPOT04, BNOK^{+}04, CGCPA08, CDD07b, GZY^{+}08, HLCG07, HPAM01, LSC^{+}08, PBM^{+}07, TVL^{+}00, Tun00, WBH03]. anoxygenic [Goe02a, SF05, SGT^{+}06]. Antagonistic [CBG01]. Antarctic [ASAC04, AW00a, AW00b, AMS^{+}02, ASH^{+}06, CDD^{+}03, DJB^{+}07, LPD^{+}04, LLB04, MAS^{+}02, PAS^{+}05, PCD^{+}08, ROP^{+}02, SAP^{+}06, SRZ04, SHS03a, SMDG04, TSH^{+}06, TS08, TKM07, VHL^{+}04, WCM09, vdPAJ^{+}06]. Antarctica [CRBF^{+}06, DP01, GMRD04, MFG^{+}09, SWD^{+}05, TPM01, TKAW03, Vin00, VH06, WSNH06, dVKJK07, dVKT^{+}09, KMvD^{+}09, SG08]. antennule [SCH07]. Anthropogenic [PVP^{+}06, BDK^{+}06b, CRF01, KHQ01, Sch01, SSS02]. antioxidant [BSMR08]. antipredator [SD05, VK06]. antipredatory [Sak00]. Apalachicola [CL02]. Aphanizomenon [Ber01, Plo08]. Apopka [BHVC05, BHVC06, Sch06a]. Appalachian [LVSW07]. Apparent [YS06, HMKZ02, Kim05, MMS^{+}02, PIF08]. appear [TDT01]. appendicularian [AK00, FA03, LUHS04, TFNT07]. applicability [FHB^{+}09]. applicable [RSY04]. Application [GD01b, KSM06, NDCP08, BK01b, DL00, Har01, HFE07, NSBA01]. Applications [Gai00, CLF^{+}08]. applied [BJBA03, MCL06]. approach [Bia01, BNM^{+}09, BPLH04, BK07, DB07, DHP03, GL07, LBC06, LRS^{+}03, MBCD04, MBC06, MKV01, Mul01, MVW^{+}04, Nor04, PED^{+}09, RLM^{+}08, RB01a, SAS^{+}03, TBH07, WIK08, WKFP06, Wro01]. approaches [Cur00, LSF^{+}08]. Aqaba [FPZ09, FWM^{+}05, MLC^{+}07, LAG^{+}03, SMGP06]. Aquatic [Kem00, BS02a, BB04a, BP03, BSJ02, BSJM06, BK07, CCH02, HBTM05, HSF^{+}08, HM03a, HA08, HR03, IKT03, JCK^{+}08, LLT05, LL05, MPJ08, MD04, MKV01, Mul01, NA09, OH08, PNPA04, RHJ04, STST07, SL00, Thi00, TBH07, VR01]. aquaticus [ZB03b]. Arabian [BBJ^{+}06, FTC^{+}06, FTY^{+}06, Goe02b, JHW^{+}07, NDT07, WGC00]. aragonite [CPL^{+}04]. araucanus [MBC^{+}05]. Archaea [KED^{+}07, TLvAH06, WCB^{+}06, CDD^{+}03, CCCC07]. Archaeal [TvAVH06, WKW09]. archipelago [CRY^{+}08]. architecture [WW00]. Arctic [BJG^{+}00, BGHA04, CTY06, CTWFP06, DB07, Dit04, GLP^{+}08, GGH07, HM08, HBF^{+}06, KSM06, KED^{+}07, LFPH07, LN02, MSGK06, MFK^{+}09, MHK^{+}07, MVA^{+}09, OBB^{+}04, RG04, RGRPD04, SSH^{+}09, TMH^{+}06, VGP^{+}08, WCB^{+}06]. area [AOR07, BRW07, ET08, KKKT07, VMC^{+}07, FKS^{+}01]. areal [ME06]. areas [DBWS06, MTC^{+}06, Reu02, SF04]. arenaria [HM02]. Argopecten [TG09]. arguments [RS09]. argus [BFCLÁ08, GB09]. arid [CIWH09]. arid-zone [CIWH09]. arising [HLH09]. armatus [VVD^{+}09]. armored [MA00]. aromaticity [MBW^{+}01]. arrangement [KTM01]. arsenate [HFLD03]. arsenic [LVSW07]. Artemia [GSA^{+}05]. artifacts [HBH^{+}07]. artificial [EDF^{+}03, KTM01]. Ascending [ASP04]. ascidians [RCAK05]. Asellus [ZB03b]. Asexuality [AMW^{+}07]. ash [MSBHG01]. ash-free [MSBHG01]. Asian [SWF04]. ASLO [NDCP08]. Asparagopsis [MHDH00, PGC07]. aspect [MCV^{+}05]. aspects [EDF^{+}03]. assays [GWH06]. assemblage [Dav00, HSCF06, KLS^{+}06, KLS^{+}07, LB04b, Rit09]. assemblages [ASI^{+}02, CCC^{+}03, CS05, DMEPA04, FH09, FB05, HMS^{+}04, HLGI08, HPT^{+}09, KCWC06, MT06, NKE08, RP06, RM06, RJS^{+}03, SNPK^{+}09, SMA00, TMC06, WPW^{+}09, WBR03b, ZSS^{+}08]. assembly [AR06, DVV^{+}07]. assessed [CEBMF05, RHJ05]. Assessing [BJW^{+}02, BW05, BZS07, CDH^{+}08, CGWK^{+}09, FHB^{+}09, Lan03, MBF^{+}03, MBCD04, VD08, WNP04, YJMT08]. Assessment [BTS^{+}06, CLC02, Mul01, SMC08, SKH^{+}01, AD02, BVH02, DBE08, HHB^{+}09, HPV^{+}00, HBM08, KNC^{+}09, Law03, LWB^{+}01, SB03, VHL^{+}04, WJMD07, WR03]. assimilating [MKK04]. Assimilation [BFS02, WLTF03, CBG01, GSNH04, MKVK05, RlBW02, TMC06, WDX01, XW02, YTVR07, YW02, dlM09]. Assimilatory [ATJH08]. associated [BVN^{+}02, BSC^{+}09, ED05, KD02, KLSW00, RSWH04, SDC00, TG03, WCGBG07, WWWY05, dVKJK07]. Associates [SG02]. Associations [WGKE04, NA05]. Asteroidea [MSBHG01]. Asymmetric [PF01c]. Atlantic [ÁS07, BSSL03, DJG^{+}06, DKA02, GA00, GA01, GJT^{+}08, GAM^{+}01, GP03, GHN^{+}02, HDMB04, KLAF04, LD01, LHB04, MH04b, MML^{+}08b, MVC^{+}05, PLW^{+}08, TSSWK^{+}06, TSDK^{+}09, Wey04, dCCL09, ACW^{+}03, ASGA^{+}07, ÁSGMP01, ÁÁS07, BAG^{+}09, BGC^{+}04b, BZS07, BST^{+}08, CPL^{+}04, CRJ08, DTM^{+}08, Dix00, DAA^{+}01, DWA^{+}02, EJL^{+}06, EMPI07, FSM^{+}07, GSG^{+}05, GvRE^{+}02, GSL^{+}07, HVH00, KBD^{+}00, LES^{+}05, LA06, MKK04, MKVK05, Mar05, MML^{+}08a, MIN^{+}05, MCB08, MCC02, MN04, MY05, ORH01, ODR^{+}08, OS01b, RLP^{+}06, RvAV^{+}06, RSH08, SIH^{+}06, SRF^{+}01, SGT^{+}06, SAPA^{+}02, SMC08, StLHB^{+}07, TPSF01, TvAVH06, TLvAH06, TN09, UM05, WJMD07, ZJC^{+}07, vODL^{+}09]. atlanticum [LJ09]. Atmosphere [BMH^{+}08]. Atmospheric [MOK^{+}03, MBPVR06, RBC00, BBJ05, BGC^{+}05, Caf01, CH05a, DT00, EBS07, LH08]. Atoll [APH^{+}03, DACG01]. ATP [LTT07]. attached [MWH01, NH06b, PIF08]. Attachment [JJP04, Car02]. attack [LLF^{+}09]. attenuance [STK^{+}04]. Attenuation [LVC^{+}00, ARK^{+}00, BJG^{+}00, BVK02, BB09, Kir03, MV00]. attractants [HMJ04]. Atyidae [PH07, CRY^{+}08]. augmented [BSA^{+}00]. Augusta [LAN08]. Aulacoseira [JGZ^{+}08, VOG^{+}08]. Aurelia [RG06]. Aureococcus [BRL03, GRB02, MGL02]. Aureoumbra [LB00]. austral [PTK^{+}07, SR07]. Australia [Chi00, CIWH09, DCMS03, EM02, HMS^{+}07, HBTM05, HLD^{+}03, NSH^{+}09, SMC08]. Australian [EF05]. australis [MHRW02]. Austrian [LD01]. Authors [Ano01]. autocatalyzed [BFBHF04]. Autochthonous [KCP^{+}04, VVSM09, JMBJ01, KAJ08, MGMW06, SM05b]. autonomous [BST^{+}08, CLF^{+}08, FDH^{+}01, HM08, JN08, NAH^{+}08, SGLT08, DIMP08]. Autosub [NAH^{+}08]. autotrophic [BMCvE06, BGC^{+}05, Nil03, TRT04]. autotrophs [RH06]. autumn [MFG^{+}09, TMS^{+}00, XBD^{+}09]. Availability [BGC^{+}04b, CSE^{+}08, ALRM08, ARO^{+}04, BFCS^{+}01, CFMH01, FF04, Har02, KTLS03, KHB^{+}00, LCPS05, MLC^{+}07, MED^{+}08, MPA^{+}07, MML^{+}08b, PCA^{+}00, RH06, SXLM06, SP04, TGdB^{+}01, TvdWdB04, TMB07]. available [BVK02]. averaging [OHCP02]. Avicennia [KA06]. avoidance [ARB^{+}04, Fra01, IJA^{+}04, JS03b]. Avoiding [PPL^{+}02]. away [YM03]. axenic [SBL06]. azoricus [Dix00].

B [Bel01, Boy01, Caf01, Sta01, BSR^{+}07, FGM^{+}03, KCE^{+}00, MBC^{+}05, TS08, WdMD^{+}00]. Ba [TBHS01]. Ba/Ca [TBHS01]. Bacillariophyceae [PD02, TSG^{+}01, VOG^{+}08]. back [SRKK08]. backscattering [LMBP07, SSMM00]. Bacteria [KED^{+}07, TLvAH06, AGMP04, BVN^{+}02, BRT^{+}03, BVB03, BBL^{+}03, CMSVA02, CGCPA08, CNGV09, CBK^{+}08, CTY06, DMJ09, DAV^{+}05, FSSP02, GM04b, GD00b, GD01a, GW03, GSLT08, GBM^{+}04, GFJ^{+}02, GHN^{+}02, GNU00, HH04, HPHT00, KJ01, KCP^{+}04, LLT05, LLB07, LB04b, LRMV04, MKK04, MdG08, MGF03, MWH01, ORFD^{+}09, PDR^{+}09, PIF08, SF05, SGT^{+}06, SSM^{+}04, SC08, TRT04, TS08, TCD^{+}03, VMH^{+}01, VEMM07, ZHSJ09, dGMH^{+}08, vNHM^{+}09, CDD^{+}03]. bacteria-containing [dGMH^{+}08]. bacteria-mediated [BBL^{+}03]. bacteria-phytoplankton [DAV^{+}05, SC08]. Bacterial [BA01, CMC^{+}09, GP00, KB06, KAW05, PTLJ01, PW01, PG00, PBGV04, SP04, SVB^{+}08, TPM01, ZB03b, ZWC^{+}04, ASGA^{+}07, AH02, AKP^{+}09, BLJ09, BMBA07, BOC01, BKM05, BdG02, BCDH09, CC01, CSSNA02, Car01, CKA03, Coc01, CD03, DKA02, GCG^{+}02, Goe02a, GvRE^{+}02, HTS^{+}03, JCK^{+}08, KB08a, KB08b, KHP^{+}06, KL00, KLC06, KMC^{+}00, KDMC05, LJ01, LSC^{+}08, LFAB06, LTT07, LCLP02, MBCD04, MML^{+}08a, MNM^{+}09, MLSS05, OBSD04, PS06, RORR^{+}09, RSH08, RlBW01, SKJ^{+}07, SJW03, SC08, Thi00, TB09, TS00, VDK04, VMS^{+}04, WSHW05, WKW09, YCM^{+}01, YNCK04, dGB02, vOMSM06a, vOMSM06b]. bacterial-based [BMBA07]. bacterio [CG01]. bacterio- [CG01]. Bacteriochlorophyll [Goe02a]. Bacteriochlorophyll-a [Goe02a]. bacteriophage [BVN^{+}02]. bacterioplankter [BR00]. Bacterioplankton [HD03a, KCWC06, NFFK00, AWLH04, CH05b, GAPR06, HDSC09, HSCF06, LB04b, MSAC06, MPACG01, MSVAC04, MSPAE06, MVC^{+}05, NSM09, PPH03, Too01, UMASG07, UVY^{+}01, WBH03, WAGH09]. bacterium [MHB01]. bacterivorous [AVA^{+}02, DMJ09]. bacterivory [VGP^{+}08].

Bahamas [DZL^{+}03, LRS^{+}03, ZVRL03]. Bahamian [BZ02]. Baikal [RJS^{+}03, ES00, JGZ^{+}08, KSS^{+}07, MBF^{+}03, MMW06, RKWG00, SDG^{+}07, WRG^{+}05]. balance [AD05a, Ano07, BJBA03, CZRD^{+}04, CR10, DGH^{+}07, DGV01b, DdG09b, GHTJ00, GAM^{+}01, MCM06, PBC02, RB01a, RAKE05, RB01b, SB03, SRF^{+}01, SK01, STST07, WPV00]. balances [KFW09]. Baldeggersee [TB00, TB05]. baleen [Sch00]. Ballast [PIF08, BDvO^{+}03, GJRM07]. ballasting [PIKB08]. balthica [PvAB^{+}03, RHM04]. Baltic [SSA^{+}04, TKBJ05, ART^{+}06, BEW^{+}00, BEM^{+}02, Dup00, GEJ^{+}02, GLVFH09, GZY^{+}08, HLK^{+}07, HVH00, HH02, KAW05, KRM09, KUP^{+}06, LHWE01, LSK^{+}03, LL05, Plo08, RSY04, VJ07, WSBL06]. banding [EKdF07, WCGBG07]. Bangdu [HLK05]. bank [GRCS06, JS03a, SK04, SDH02, ODR^{+}08, vODL^{+}09]. banks [DLDE04, FRO01, DZL^{+}03, DDZ03]. bar [RSC04]. Barbara [BVB03, MSSW^{+}07]. Barcelona [Wan02]. Barents [NML^{+}05, OS01b]. Barium [TCL^{+}09]. Barnacle [PDS05, PHMR^{+}07]. Barrier [Chi00, GEP08, KDS03, LLvW^{+}03, WKH^{+}06]. basal [HGBB08]. base [DSD02, WTM09]. base- [WTM09]. based [ASH^{+}06, BRT^{+}03, BMBA07, CHW^{+}01, GDL^{+}00, HKNG08, Leo08, LM02, MBF^{+}03, PLK^{+}05, PMPC00, RNR08, RPR^{+}02, RMS^{+}08, SPM04, SY07, Too01, GFK00, RFW08]. baseline [MM03, MdG08]. baselines [WAB^{+}08]. Basin [AIK04, KR08, YT09, dBWF^{+}05, AIS00, BNOK^{+}04, BRB^{+}01, lLCChTO08, GGIA06, HISW00, HCD^{+}09, LES^{+}05, LH08, RHCS08, RvAV^{+}06, SI08, TVL^{+}00, dlFSIN08, AHR^{+}01, BVB03, HTAS06, HST^{+}02, LSC^{+}08, MOK^{+}03, ME07, TIH^{+}01, TVL^{+}00]. Basin-scale [AIK04, AIS00, GGIA06, HISW00, LES^{+}05, SI08]. basins [KSM06]. bass [BFS02]. batch [BS03, GD00b]. bathymetric [SGLT08]. bathymetry [DZL^{+}03, LRS^{+}03]. Bathymodiolus [BFS^{+}04, DBS^{+}04, Dix00]. bathypelagic [HD03a, HSCF06, RvAV^{+}06, TvAVH06]. BATS [GHN^{+}02, MN04]. Bay [Ano08b, AP00c, BMU04, BCBS04, BW99, BJBA03, BDNF01, CL02, CCKC05, CSG^{+}06, DD09a, DAMM^{+}09, DC05, EM02, GGM03, GTS^{+}09, GSB^{+}09, GGDC04, GGS06, HHY^{+}05, Leh00b, MCPP06b, Ris02, RHN^{+}02, TKBJ05, WKFP06, WCB^{+}06, YYTH07, GGRJ03, GFM^{+}01, KFW09, PP00, RTT^{+}02, SLE04, TG09, BEM^{+}02, BMD^{+}04, CZRD^{+}04, CB04, CTCT03, CGL03, CG03, CGL^{+}04, CCH02, CDBC00, DJH05, DAV04, FHBW06, GN00, HLCG07, HLK05, JNKdV09, JRS03, KCWC06, KRZ06, KN07, LBC06, LRF07, LF09, MCL06, MC06, NDCP06, NSBA01, NCC00, PD05, RNR08, RZMB05, SAH^{+}05, SAHS06, SAS08, SSA^{+}04, SH06, SHA08, SP02, TRT03, TRT04, THWS00, TWS01, TS01, Too01, TNO^{+}08, WV00, YTVR07, ZC02]. Bay-Delta [CGL^{+}04]. bayou [HLCG07, EB01, HLCG07]. be [KGYU06, TDT01, JBvHS08]. Beach [dSYL^{+}08, CBH08]. beachface [HU07]. Bear [DFBA07]. bearing [LEML09, YSM^{+}03]. beat [Mor06]. Beaufort [WCB^{+}06, XBD^{+}09]. Bed [JWRW07, AH01a, SDJT03]. bedload [SPM04]. beds [CMD^{+}09, DRSL07, FT03, JPK^{+}05, PVD^{+}07, RKH03]. been [NH06a]. before [YM03]. Behavior [CWLH06, BRM^{+}07, BDC00, EDF^{+}03, FMS04, FNM07, GPAP^{+}04, JWRW07, KDO09, KJ01, LLS08, LB00, RKGC^{+}01, SD05, SGPK07, SFC^{+}00, SSS06, TGF00, WM07, vDBL05]. Behavioral [CMJ07, GB09, RH03]. behaviors [BGC04a, DG06, SJ08]. being [LM02]. Belgian [BRP^{+}06]. Belgium [Gou01, SMH^{+}06]. Belize [SGLT08, WLRT06]. below [AB07]. belt [JCBA08, PDS07]. beneath [NDP^{+}07, SFL05]. Bengal [BMD^{+}04, RNR08]. Benguela [BZS07, PN06]. Benitez [MWE^{+}04]. Benitez-Nelson [MWE^{+}04]. Benthic [ALH01, BLM^{+}06, CWG^{+}07, Dal03, Dup00, EF05, GK03, JHW05, KFW09, KT02, KKT05, NHTF01, SLL^{+}04, TMA03, WBB09, AG03a, AJ01, ARO^{+}04, AGP07, BB04a, BST^{+}05a, BS02b, CCS06, CVH^{+}05, Com02, CDD07a, CJB00, DTFL03, DS07b, DPC^{+}08, FDH^{+}01, GMR^{+}09, GGH07, GTK^{+}01, GGRJ03, GM04b, GBFJ07, GFK00, GSFB02, HMS^{+}04, HBF^{+}06, HFR08, HLK05, JFH^{+}05, KBL08, KDS03, LLC^{+}09, LRS^{+}03, MG04, MTM^{+}01, NH06a, NTM^{+}05, Nil03, NHN^{+}05, PDR^{+}09, QMS^{+}07, QLD^{+}02, RCVK09, RTT^{+}02, SSAZ01, SABA02, SKBC01, SF04, TVT04, TS01, VJV^{+}03, VKV07, VMJ08, VH06, WBP09, WG04, WR03, WBD02, WA01, WLTF03, YYTH07, YM03]. Benthic-pelagic [NHTF01, KBL08, TS01]. benthopelagic [DS07b, MCV^{+}05]. benthos [HM03a, JTAM09]. benthoscape [ZLP^{+}03]. Bering [BB07, MT06, Sch00, Sch01]. Bermuda [BSM01, GHN^{+}02, MVC^{+}05, MN04, ZXN^{+}08]. Bertness [SG02]. between [ARV^{+}07, AVJ^{+}00, Ano08a, Arn00, BAPMMB01, BGE04, BSS09, BJBA03, CVH^{+}05, CL02, CLL^{+}09, CLC02, CPCS01, CWC^{+}07, Cul06, DC05, EVW07, ERH06, FMG^{+}08, GL07, GJRM07, HDSC09, HBTM05, Ham05, HHY^{+}05, HD09, HSLV05, HH04, HWL^{+}00, JL08, JM02, LAG^{+}03, LJCM08, LLT05, LRMV04, LTB09, MW02, MRM07, MHC05, MPB^{+}08, MLSS05, NHPHH08, OB04, OV03, OHT06, PPH03, RC05a, RJS^{+}03, SWH^{+}05, Sar07, SRF^{+}01, SGPM09, SM01, TMK^{+}07, VMH^{+}01, VDK04, VZM07, Wey04, WA01, WBB09, WLTF03, ZB01, ZMK^{+}08, vNHM^{+}09]. Beyond [LSK^{+}03]. bias [RFW08]. biasing [KHQ01]. Bicarbonate [MT06, BTdB03]. Bidirectional [MZV03, VM05, ZVRL03]. Big [PBGJ09]. bigger [HD03b]. Bight [JHW05, GJT^{+}08, MY05]. bilateral [WRD01]. binding [AKH04, BRT^{+}03, DNB^{+}04, SRM05]. Bio [BVG^{+}07, OCSB05, BPC06, GGDC04, NCPd08, RL01b]. bio-fluid [RL01b]. bio-irrigation [GGDC04]. bio-oceanographic [BPC06]. Bio-optical [BVG^{+}07, OCSB05, NCPd08]. Bioaccumulation [AGL03, SF03a]. Bioavailability [ACW^{+}03, BK03b, CW01, SDC00, SSS02, WTM09, HHB^{+}09, HPHT00, KS04a, KSM^{+}05, PRKHW04, SGR08, SHL^{+}05, SLJ00, WS04]. bioavailable [VJ07]. Biochemical [AFL07, BvE07, MH00, BWK^{+}07, LG04, LBC06, SHSH^{+}07]. biocomplexity [LALE04]. Biodegradable [KKH06]. biodeposition [NCO02, NHTF01]. Biodiversity [CSWD07, LN04, Nor04, BK07, DJT06, Hei08, HP06, KMHE04, WMWV04, WBP09]. biodiversity-inspired [BK07]. biofilm [AGKK08, Bat00, DKA^{+}03, EW07, HBH^{+}06, LZ09, RGAS04]. biofilms [AGP07, HMJ04, UPC^{+}05, WNAW09]. Biogas [DJB^{+}07]. Biogeneration [ORFD^{+}09]. Biogenic [BMD^{+}04, HMT^{+}08, KFN00, LVB08, SSK06, SDV^{+}06, THWS00]. Biogeochemical [ÁÁS07, BSM01, CFMH01, Chr05, CFO^{+}00, GEP08, MBPVR06, PTK^{+}07, BZ03, CV05, CA08, FZL^{+}05, KFW09, LBM^{+}04, LMBP07, LLDS06, MC06, NCPd08, NIE^{+}00, OCSB05, PDM^{+}09, PWS09, PCD^{+}08, SH10, Thi00, TS00, VHL^{+}04]. biogeochemically [MPACG01]. Biogeochemistry [AJS07, BDSB07, McM00, WSNH06, COK^{+}08, HAA09, HB06, HFZ^{+}07, JCBA08, KA06, KC08, LRF07, NEC03, SBC^{+}08, WGLT05, WMWV04]. biogeographic [PGC07]. Biogeography [KDMC05, GBL^{+}04, HLGI08]. Bioirrigation [MGGM06, MCJ^{+}06, MKV01, SdB06, VPH^{+}07]. Biological [AKE04, BSA^{+}00, Gay00, MJW03, PAF^{+}04, SFC^{+}00, TT05, THH02, dVKT^{+}09, FZL^{+}05, Gou01, HHB^{+}09, KCCE06, KHK^{+}03, LLB04, MSZ00, NSH^{+}09, OH00, OB04, SDJT03, SCS^{+}08, WdMD^{+}00, WW02, WCL^{+}07, YYG02, ZZPP01]. biologically [MMS^{+}02, RCL^{+}02, VJ07]. biology [DORS08, Kil00, TS01]. Bioluminescent [vDBL05]. biomagnification [CLS05]. biomanipulation [Sos00]. Biomarker [BKM05, PMPC00]. biomarkers [PAS^{+}05, SHS03a]. biomass [ADA00a, ADA00b, AFH07, AH03, BDK^{+}06a, BBJ05, CD03, Dup00, FMWC00, Goe02b, HFR08, HK01, JDS08, KRZ06, KLSW00, LJ01, Leh00b, LCC^{+}09, LTB09, MMO06, MGRE^{+}07, New01, RZMB05, SRZ04, SL01, VWPP^{+}06]. biome [FAL01]. biomonitor [CHT02]. biooptical [Zim03]. biophysical [PC04]. Biophysicochemical [AGP07]. Bioreactivity [MBC06, AFB01]. Biosphere [FAL01]. Biostabilization [DRSL07]. biosynthesis [Shi04]. biota [Sch01, SMA^{+}06]. Biotic [HWYL07, MJ03, BDNF01, ESA06, OBB^{+}04]. bioturbated [KGP^{+}02]. Bioturbation [VKV07, VPH^{+}07, Shu01, SDH02, VRS09]. Birch [SMS09]. births [JKS04]. BIT [HSB^{+}06, WIK08]. Bivalve [RH01, JPK^{+}05, LWU^{+}05, NCO02, PvAB^{+}03, SO07, TCL^{+}09, WSNM00]. bivalves [Ben00, GJB^{+}04, GWR^{+}09, NHTF01, PMPC00, RL00, RHM04, SLD00]. Bivalvia [Dix00, MBDL01, WSBL06]. Biwa [MS05, OKO^{+}01, BVK02, GNU00, INU04, KGYU06, UGY^{+}00]. Black [MH04a, AJS07, DKRZ06, FDH^{+}01, JKGT08, KLF^{+}03]. black-water [AJS07]. blackwater [PCA^{+}00]. Blackwell [Wro01]. Blanes [DAV04]. Bleaching [ACHG07, Fab06, FMWC00, HGFS^{+}05, LRYS06, OFKH01, RG07, RRHPT^{+}06, WLRT06, WAB^{+}08, ZJ03]. blocks [OV03]. Bloom [CPHM04, TMH^{+}06, AGL03, AH02, BST^{+}05b, BSSL03, CR08, EGPA02, EZA^{+}05, GEJ^{+}02, GRB02, GSW03, GPHS08, LFS^{+}04, LCS02, LU08, LRF07, MWM^{+}04, MG04, MLSS05, PSLO07, PF01b, RLP^{+}06, RHN^{+}02, RRBS03, RNA06, Sar07, ST03, SKH^{+}01, SMG06, TWC^{+}09, VMS^{+}04, WAGH09, WMW^{+}08, YCM^{+}01]. bloom-forming [LCS02, RRBS03, Sar07]. blooms [AKG^{+}05, Ber01, BEW^{+}00, BRP^{+}06, BRS01, CC04, CHH07, CFG07, EJL^{+}06, FA03, GFM^{+}01, KNE08, KNC^{+}09, Kut04, LMT05, MCL06, MGVM^{+}06, NKE08, PSEL08, Plo08, PMK08, RC07, RTT^{+}02, SNM^{+}08, TKAW03, UM05, WW03, WL06]. blue [Car02, GF02, HFC^{+}04, JWRW07, MV00, REL06, SGPM09, HLD^{+}03]. bodied [DGV01a]. bodies [Caf01]. Body [PSH08, ASH^{+}06, HBH00, HB03]. Boeckella [ROP^{+}02]. Bolivian [HMT^{+}00]. bomb [BW01]. Bonneville [KR08]. Bonney [LPD^{+}04]. Book [Bel01, Boy01, Bri01, Caf01, Dow02, Elm02, Fou02, Gou01, Jun01, Kne02, Mil02, Pom01, Ris02, SG02, Sta01, Wan02, Wro01]. Borderland [ME07]. Boreal [BCSW07, BST^{+}05a, BP03, BLJ09, CA08, DTFL03, FKS^{+}01, GVT06, GV09, HLL^{+}06, HSB^{+}04, KHP^{+}06, KFR^{+}01, PAF^{+}04, SH10, SLJ00, SKC^{+}01, TXH05, WK09, vWSBT08]. boryana [STHG07]. Bosmina [ABJ^{+}06, SCH07]. bosminids [TIH02]. Bosporus [KLF^{+}03]. both [BF08a, dGP08]. bottlenecks [VAS02]. Bottom [CKA03, Arn00, BZ03, BMWW08, CCS06, CLL^{+}03, CV05, DDL^{+}09, FH09, LPW05, LWS01, MSL^{+}03, MMO06, PGP01, SCDR07, SHS03b, TG03, WPW^{+}09, ZB03a]. Bottom-up [CKA03, FH09, LWS01, MMO06, SCDR07, WPW^{+}09]. bottoms [MS03, MZV03]. Boudreau [SR04]. bound [CW01, XW02]. boundaries [SC00]. Boundary [RMK^{+}06, ÁÁS07, BMWW08, GBFJ07, JWRW07, JPK^{+}05, KSM06, LMMW03, LPW05, NA09, PN06, PSJ00, RM04, RHJ02, RHJ04, SSAZ01, SABA02, YM03]. boundary-layer [JWRW07, LMMW03, SSAZ01]. Bow [BSS^{+}01]. Brachionus [Gil03, MA00, SS03, SS06, VZM07]. brackish [KAW05]. branched [RKM^{+}06]. branching [JKLR08]. Branchipodopsis [HMD^{+}07]. Bransfield [SWD^{+}05]. Brazil [RLC^{+}01]. breakdown [SSM^{+}09]. Breathing [LMMW03]. breeding [HRA07]. breeding-system [HRA07]. Brest [MCL06]. brevis [KHNV05, KSP07, PMK08]. brief [Ack03]. brightwellii [RA00, RNA06]. brine [GSA^{+}05]. brines [PTK^{+}07]. British [NK05, Tun00, BBP^{+}06, LCS02]. brittle [VTR03]. brominated [KKH04]. bromine [HM09]. bromoform [MB01]. Bronk [SRG00]. brown [GRB02, JM02, KCE^{+}00]. brownwater [HFE07]. Brunnsviken [RMG^{+}04]. Brunswick [EF05]. bryozoan [FRO01]. bubble [ZLL^{+}02]. bubbles [BGJ01a, BGJ01b, Ost03]. budget [BMT^{+}04, CHB08, CBH08, D’A08, JKLR08, KHM04, MHB03, SSA^{+}04, WBR^{+}03a, YW02]. budgets [ÁÁS07, AMS^{+}02, EM02, HLD^{+}03, KM02, MED^{+}08, MAS^{+}02, PCvRR00]. Buenoa [HGB00]. building [BPC06]. Bulk [YSM^{+}03, LSK^{+}03]. buoyant [CRC^{+}03, OIH^{+}03]. burial [KNE08, SDKZ^{+}09]. burrow [MGGM06]. burrowing [EMPI07, MHC05]. bush [HGBB08]. Bypassing [DPC^{+}08]. Byrne [BK01b]. Bythotrephes [MD09, FS09b, HBM06].

C [Fou02, Gou01, Jun01, MBCD04, Mil02, VW08, WIK08, AHM^{+}08, ALRM08, ASS03, BCC^{+}04, BSMR08, BLM^{+}06, CEBMF05, CCKP02, CWC^{+}07, FKCF00, GRB09, GP06a, GSH04, GHN^{+}02, HFR08, IR06, KTG^{+}06, KCP^{+}04, LBM^{+}04, MH07, MBLP01, MdG08, MBB^{+}00, RRSC01, RB01a, RHM04, SRL^{+}09, SL02, SKC^{+}01, SSL05, TMYH02, TB05, VW08, VMS^{+}04, VR01, VvOBM06, VPA02, VMS09, WGKE04, YPBL07, dGMH^{+}08]. C-DIC [BCC^{+}04]. C-enriched [BLM^{+}06, dGMH^{+}08]. C-labeling [MBB^{+}00, VvOBM06]. C. [WLP^{+}07]. C5 [Has06]. Ca [TBHS01]. caddisfly [SrL05]. cadmium [CLS05, CCC^{+}03, CS05, Cul06, DA05, FQC^{+}07, LJCM08, MHC05, NRB08, OHT06, SDC00, TvG02, TLM04, XTSM07]. caerulea [dGMH^{+}08]. Calanoid [BS01, KSEÖ^{+}02, BHL07, LUHS04, SC07, SSH^{+}07, ZRL06]. Calanoides [IOG^{+}04]. Calanopia [CF05]. Calanus [BKAE01, EAO^{+}02, Has06, IHH^{+}00, IHHH00, KBD^{+}00, ODR^{+}08, PFL07, SvKM01, TKK02]. calceolata [DBGR09]. Calcification [Chi00, GRVF^{+}00, AS00b, GCM^{+}09, LGJ02, LAHdBJ05, MTC^{+}06, MAL08, SE06, SLDE09, TLrR07]. calcifying [GRV^{+}09]. calcite [Gre09, MWW06, PIKB08]. calcitrans [TSG^{+}01]. calcium [AY08, HR00a, LGN^{+}06, TW09, TLrR07]. Calculation [OZM^{+}01]. calculations [MY05]. Caledonia [DJT06]. Calibration [JFH^{+}05]. California [MBR08, AD05b, AO09, BNM^{+}09, BBG05a, BW99, BRS01, BJBA03, CJ08, CSM^{+}08, CLG08, CHB08, CBH08, DORS08, FL08, HP06, KMC^{+}00, MCJF09, MCPP06b, ME07, PHMR^{+}07, RA09, RCE^{+}02, SMA^{+}06, SMC03, TAB^{+}00, dSYL^{+}08]. Callinectes [HFC^{+}04, JWRW07]. calyciflorus [MA00, VZM07]. Calyptogena [SO07]. Cambridge [Bri01, Fou02]. camouflage [JS03b]. campechanus [BW01]. Can [MIDH05, Has04, KVdB^{+}04, LAJ^{+}04, RES^{+}08, TSG^{+}01, WM07]. Canada [BBP^{+}06, LCS02, PDS07, RKH03]. Canadian [WCB^{+}06, BSS^{+}01, HBM06, KCE^{+}00, QLD^{+}02, TXH05]. canopies [GDK03, Zim03]. canopy [Ano08a, ERH06, LSF^{+}08, MPB^{+}08]. canyon [GCG^{+}02]. canyon-shaped [GCG^{+}02]. capability [HD09]. capacity [BPLH04, JP03, KCP08, LSK07, Sch00]. Cape [CBA01, GCP^{+}02, HU07]. capillary [BK01b]. capture [Ben00, PNPA04, RL00, SLD00, VD08, WSNM00]. carbohydrate [SR07]. carbohydrates [HY01, SHA08]. Carbon [CGCPA08, CYN05, EM02, GDM02, GRdL^{+}08, HPBD01, KMC^{+}00, KDRR09, LZ09, MDL00, MBR08, MSZ00, RRBS03, RRS06, Sch01, SBZ^{+}09, SKC^{+}01, TMYH02, TTG^{+}09, VMS^{+}04, WC04, All00, ASGA^{+}07, ÁSGMP01, ÁS07, AP00b, AG03b, Ano04, ARK^{+}00, ASS03, AGKK08, BMF^{+}06, BMT^{+}04, BGV02, BNC04, BLJ09, BMD^{+}04, BKM05, Bou04, BTS^{+}06, BBP^{+}06, CWWS00, CGD08, CL04, CGH^{+}04, CLPB02, Ceb02, CCH02, CCKP02, CWG^{+}07, CRF01, CS05, DBS^{+}04, DB07, DH01, ET08, EPS00, EGPA02, EB01, FQH^{+}01, Fin04, FLWP09, FSSP02, FSBB05, FXL04, GR01, GM04b, GSW03, GD00b, GMMW03, GGDC04, GKJ04, GP01, GSH04, GSL^{+}07, HPE01, HZG07, HD03a, HBCK03, HW07a, HW08, HPAM01, HDVB01, HLD^{+}03, HA08, HM06a, HST^{+}02, HB06, HKD01, Jan03, JCW08, JMBJ01, KB08a, KB08b, KNE08, KJMJ03, KJJ07, KAJ08]. carbon [KF01, KB07, KFR^{+}01, KWS06, KNN06, KFW09, KHK^{+}03, KGP^{+}02, KA06, KCP^{+}04, KSWC^{+}03, LTM01, LSPC02, LMSE05, Law08, Lee01, LTGG03, LBB^{+}04, LBM^{+}04, LFFC06, Lon00, MMDW09, MPJ08, MH04a, MCF^{+}04, MTC^{+}06, MHK^{+}07, MED^{+}08, MdG08, ML02, MHB03, MBB^{+}00, Møl07, MF01, MF02, MWE^{+}03, MWE^{+}04, MC04, MLPS09, MC01, MB05, NH06a, NA03, NA06, NHN^{+}05, OH00, OB04, OV03, OCNH00, PC02, PCC^{+}07, PWX04, PKKB05, PLW^{+}08, PHF03, PG00, RBC00, RA06, RFM07, RFW08, RCVK09, RZR02, SPW^{+}05, STP^{+}07, SDKZ^{+}09, SWN08, SlBWC^{+}00, SAE^{+}08, SC08, SL00, SH04, TPM01, TG09, TIH^{+}01, TWS03, TPSF01, TB05, TTSC04, TVL^{+}00, Too01, Tor00, TRM00, TM02, TMC06, TPG^{+}08, TSMB05, TDC^{+}09, TB09, THJ^{+}09, TR03, Tun00, VSSTS00, VMH^{+}01, VDK04, VVSM09, VRS09]. carbon [VM06, VMS09, WGLT05, WG01, WKH^{+}06, WBD02, WK09, WS04, WWMB05, WLS00, WCL^{+}07, XLF^{+}03, XBD^{+}09, YSM^{+}03, ZXN^{+}08, dGvD07, dGMH^{+}08, dGP08, vOMSM06a, vOMSM06b, vODL^{+}09, vWSBT08, vWBT09, BMH^{+}08]. carbon-based [Too01]. Carbon-nitrogen [VMS^{+}04]. carbonate [BZ02, BZH08, CFMH01, DFBA07, FF04, HA01, JNKdV09, LMZB08, MTC^{+}06, SE06, SLRM03, SMH^{+}02, TMK^{+}07]. carbonic [EPS00, MT06]. carbonyl [CCF04]. carboxylic [PBGV04]. carcasses [LJ09]. Cariaco [HTAS06, HST^{+}02, LSC^{+}08, TIH^{+}01, TVL^{+}00]. Caribbean [BFCLÁ08, GB09, LLvW^{+}03, MG08, MLG^{+}03, PVD^{+}07, RRHPT^{+}06, ZJ03]. Caridina [Har01, PH07]. carinatus [IOG^{+}04]. carnivore [MCB08]. Carolina [Mil02, VWPP^{+}06, DAL02, GHS05]. carotenoid [ETCP03]. carotenoids [MGW^{+}05]. carrying [Sch00]. Carwash [vHRB^{+}09]. cascade [RH03, SK05]. cascades [ZHSJ09]. Cascading [MVA^{+}09, FLT02a, SVPK^{+}09]. Casco [GGDC04]. case [CCAÆ06, COK^{+}08, GSB^{+}09, HTWG08, LBSY02, MK00, NSH^{+}09, PPL^{+}02, RVP^{+}09, SCB03, SHS03a, WSL^{+}02]. cases [SrL05]. Castro [Caf01]. Catalina [CHB08, CBH08]. catastrophic [CGWK^{+}09]. catchment [BLG^{+}07, FL06, HAL^{+}09, SM05a]. Caulerpa [CM03]. Caullerya [BRE02]. cause [Cul06, LAJ^{+}04]. caused [BBL^{+}03, FDP^{+}08, TMK^{+}07]. Causes [Car02, CC04, HK06, SSvNH01, VC05]. causing [LCS02, MCL06]. cave [SMP05, vHRB^{+}09]. Caveats [AWH07, HR07, HW07b, OSDB07]. cavernosa [LMPY00]. cavities [dGvD07]. Cayuga [Kra05]. CCFZ [KCCE06]. Cd [CHT02, SH00]. CDOM [CLG08, SM01]. Cell [AS02, DWA^{+}02, AD00a, BGC04a, BCV^{+}08, BFBHF04, BBG^{+}05b, CLC02, DCMS03, DK06, FZL^{+}05, GvRE^{+}02, GHN^{+}02, GDL^{+}00, Hol00, KB06, LBC^{+}04, LA06, MC00, MCR^{+}07, MH07, MAL08, TS00, WS06, vDBL05, GGP01]. cell-level [BGC04a]. Cell-specific [DWA^{+}02, TS00]. celled [Hol00]. cells [DK06, GFPG09, HSN^{+}03, KLC06, LB00, MG04, OHUB00, SDH02, TBF04, YPBL07]. cellular [BFCS^{+}01, GGP01, JS08, JTL01, PSJ00, SH04]. cenotes [SPR02]. center [RMS^{+}08]. Centimeter [OD08]. Centimeter-scale [OD08]. Central [WJMD07, HLK^{+}07, HPT^{+}09, LBEP06, NDT07, PLG^{+}02, PHMR^{+}07, RHCS08, SAH^{+}05, SAHS06, WSNH06, PPL^{+}02, VMKN03]. centrarchids [PDH06]. centric [RA00]. Centropages [Goe08, Møl07]. century [AS06, AC08, OKO^{+}01, Wan02]. Cerastoderma [RHM04]. Ceratium [HHB05]. Cercopagis [BJW^{+}02, CHW^{+}01, GL07]. cesium [AGL03]. cesium-137 [AGL03]. cf. [TS01]. CFF [GDL^{+}00]. CFF-SPLITT [GDL^{+}00]. CH [CDBC00, LFFC06]. Chaetoceros [KTMM00, LG04, TSG^{+}01]. Chaetopterus [TS01]. chain [LBEP06, OHT06, SO07]. chains [NSM09]. chalk [TSA09]. chamber [JFH^{+}05]. Change [BD07, AOZ^{+}09, BST^{+}05a, BF08b, Boy01, CWJ09, ES00, EJL^{+}06, GCJ^{+}06, Ham05, JCC02, JCK^{+}08, LA09, MLRMR09, Sch09, SA09, SMH^{+}06, TRF^{+}06, VH09, WA09, WGD^{+}02, WSVS09]. Changes [AH02, CV05, Dav00, GRdL^{+}08, GP06a, GRG02, HM06a, JJT07, JCBA08, MPACG01, MKG04, VM06, WR00, ZJ03, AS06, ASI^{+}02, BVN^{+}02, BdG02, EPP^{+}08, ESA06, FRO01, FCJ^{+}06, GA00, GN00, GJS01, HVH00, HLD^{+}03, Hol00, HP06, INU04, KM02, Kim05, KDRR09, LES^{+}05, MH07, ME06, MED^{+}08, MSHE09, MPD^{+}06, MCB08, MSZ00, NH06b, PvAB^{+}03, RHM04, RMG^{+}04, RL07, SSH^{+}07, SSS08, TS06, UGY^{+}00, VGP^{+}08, VKV07, WKW09, YYB^{+}05, vdPVB07]. changing [FMG^{+}08, Lon00]. channel [CH05a, ED05, MSLH02, RC05a, IHH^{+}00, IHHH00, LMBP07, MSSW^{+}07, RV04]. Chao [DBWS06]. Chaoborus [OHT06, CHT02, DST08, LLS08]. characteristic [BRT^{+}03]. Characteristics [PN06, GHS05, Has06, MYM^{+}05, PWN07, RSC04, SJH03]. Characterization [CCCC07, CCC09, HSB^{+}06, NHS07, ST04, LBC06, LBD^{+}09, MBW^{+}01, RSS^{+}03, SHSH^{+}07, SHL^{+}05, YT03]. Characterizations [PED^{+}09]. characterized [LASR09]. Charette [MWE^{+}04]. Charlotte [TRF^{+}06]. Chasing [KW04]. Chasmagnathus [EMPI07]. Chatham [NDP^{+}07]. CHBr [MB01]. Chemical [CT06, HLCG07, Hes00, KKS^{+}01, KSP07, MYM^{+}05, PSH^{+}07, RAKE05, SWH^{+}03, SWS^{+}03, AFB01, BNM^{+}09, GST^{+}01, HHB^{+}09, KTM01, KNH00, KHK^{+}03, LBT06, MMW06, MCJ^{+}06, MJ03, MCL06, MJW03, NA05, OBB^{+}04, PBM^{+}07, PWN07, RBJ^{+}06, SD00, SHL^{+}05, SS03, WW01, WRD01, WH07, CR10]. Chemically [THR00, TNO^{+}08]. Chemistry [DJH05, BH02, IKT03, KSM^{+}05, MW02, SE06, Wey04]. chemoautotrophic [HPE01, RSS^{+}03, VTO^{+}05]. Chemoautotrophy [TIH^{+}01]. chemolithoautotrophic [GZY^{+}08]. Chemosensory [WW01, KJ01]. chemostat [CG08]. chemosynthesis [LM02]. chemosynthesis-based [LM02]. chemosynthetic [CMCB^{+}09]. chemotaxonomic [BCM01]. chemotype [WŠN^{+}07]. chemotypes [RES^{+}08]. Chena [CGD08]. Chesapeake [CB04, DJH05, FHBW06, GN00, JRS03, KCWC06, KRZ06, KN07, LBC06, NSBA01, PD05, RZMB05, SHA08, TS01, TNO^{+}08, ZC02]. Chidami [SH10]. Chihuahuan [GBL^{+}04]. childressi [BFS^{+}04, DBS^{+}04]. Chile [PPL^{+}02, TDJ^{+}06, WBB09]. China [lLCChTO08, CLL^{+}09, HWYL07, HWL^{+}00]. chironomid [GKJ04, RP06]. Chironomids [BPLH04]. Chironomus [MPH08]. Chlorella [JS08]. chloride [CR10, RAKE05]. chlorophyll [BVWB07, BP03, BGC^{+}05, DCC^{+}01, FB03, GRG02, HB03, HMR^{+}07, HB08, JP03, KHH^{+}03, Kut04, LTGG03, LKAB04, MMLH06, MPAS00, MGRE^{+}07, Mor03, OHUB00, PSEL08, PW03, QYEA06, SHGM06, SH04, SŻKL08, TCB^{+}00, VM05, YSO02]. chlorophyll-a [GRG02, HB03, LTGG03, MPAS00, Mor03, OHUB00, QYEA06, SHGM06, SŻKL08, TCB^{+}00]. Chlorophyta [CM03, MB01]. chloroplasts [MZL04]. cholesterol [Has04]. Choptank [FHBW06]. chromatic [HHM^{+}09]. chromophoric [CLG08, DD09b, DdG09a, FW07, HSR^{+}08, LBD^{+}09, ORFD^{+}09, OZM^{+}01, OOB09, SM01, WdMD^{+}00, YT09]. chronic [Cul06]. chronologies [DT00]. Chrysophyceae [PCDB09]. chytrid [Hol00]. cichlids [KK00]. Ciències [Wan02]. Ciliary [RL01b]. ciliate [MBC^{+}05, WSL^{+}02]. ciliates [DJT06, HMS^{+}04, JJP04, LN02, MZL04, MBCB08, MA00, SSAZ01, WS06, WNAW09, ZHSJ09]. Circadian [JS04]. circuiting [LAN08]. circular [dlFSIN08]. Circulation [RSC04, GP03, HLM08, LIHS03, PL09, TvG02]. circulations [MHLN01]. circumneutral [ESSS01]. Cladocera [GL07]. cladoceran [HRA07, LED06, VKvD05]. cladocerans [BS01, GD00a, MBK^{+}09]. Cladophora [MGH08]. clam [HM02, TBHS01]. Clams [CTCT03, TG09]. Clarion [KCCE06]. clarity [BDNF01]. class [VWPP^{+}06]. class-level [VWPP^{+}06]. classes [HPL^{+}06]. classical [VMC^{+}07]. classification [LRS^{+}03]. clavipes [VAS02]. clay [Ham02]. clay-sized [Ham02]. Clayoquot [NK05]. clear [DPFL09, JJS^{+}03, KCE^{+}00, KJMJ03, KAJ08, PCC^{+}07, SJ00, WDL07]. clear-water [DPFL09, KJMJ03, KAJ08, PCC^{+}07, WDL07]. Clearance [BB04a]. clearest [MGC^{+}07]. cleavage [EKD^{+}07]. Climate [AWE07, CPGC09, DPFL09, GA00, KBL08, MFK^{+}09, PvAB^{+}03, SBR^{+}06, VRS09, AOZ^{+}09, AE08, AH03, BST^{+}05a, BF08b, Boy01, CWJ09, DS07a, ES00, EJL^{+}06, FS09a, FKCF00, FAM^{+}06, GDB^{+}09, GN00, GHJ^{+}05, Had08, Ham05, HAL^{+}09, HP06, HAG08, KN07, LSK07, LFA^{+}09, Leh00b, LH08, LP07, LA09, LTHW04, MNA^{+}09, MD09, MIN^{+}05, MCB08, MMKN08, MVA^{+}09, OS01b, PLMPN08, PD05, QLD^{+}02, RCVK09, Sch09, SSH^{+}07, SA09, TDC^{+}09, VH09, WA09, WGD^{+}02, WSVS09, WM06, WSEL09]. climate-change [LA09]. Climate-driven [GA00]. Climate-induced [KBL08, GN00]. climate-linked [MD09]. Climate-related [AWE07, MFK^{+}09, PvAB^{+}03, AE08, Had08]. Climatic [FKS^{+}01, HVH00, JSJ03, LCS^{+}03, PBMSG04, SSvNH01, SSvNH03, SLC^{+}09, ABDS04, BDK^{+}06b, FCJ^{+}06, IKT03, LLK05, LW02, PVP^{+}06, Sal05, SFB^{+}09]. Climatological [YSO02]. Clipperton [KCCE06]. Clipperton-Clarion [KCCE06]. clockwork [WSEL09]. clonal [SBS07]. clones [WH07]. Close [CLL^{+}09]. closure [GDCG08]. clouds [ARO^{+}04]. Clues [WCB^{+}06]. Cnidaria [WCGBG07]. cnidarian [GF02]. CO [BSM01, BHL07, MML^{+}08a, SXLM06, BSM01, Bat02, BDS^{+}04, BMH^{+}08, BS03, BARS01, CPGC09, CTCT03, CPL^{+}04, CR06, CRF01, DJB^{+}07, EZA^{+}05, FH03, GR01, GAPR06, GZY^{+}08, HFZ^{+}07, KLS^{+}06, KLS^{+}07, KHQ01, LGJ02, LFFC06, MBG^{+}05, MdG08, RBC00, SNPK^{+}09, SH05b, TT05, WC04, WKH^{+}06]. co-limitation [MML^{+}08a, SXLM06]. co-occurring [BHL07]. Coalescence [HSS06]. Coast [LJ09, Agr05, BMD^{+}04, BFCLÁ08, KNC^{+}09, KF01, PDS07, PSEL08, PHMR^{+}07, SMC08, TvG02, THH02, TWC^{+}09, WSC^{+}05, WSC^{+}06, YSO02]. Coastal [CCAÆ06, Dow02, KLG09, WLS00, AOR07, ÁSGMP01, ARO^{+}04, AHF^{+}05, AH02, AWH07, BMFS^{+}03, BEM^{+}02, BGE04, BMU04, BRP^{+}06, BLS^{+}06, BRS01, BS03, BM03, Caf01, CG09, CT05, CC04, CLF^{+}08, CLG08, CWC^{+}06, CTY06, DD09b, DLZ^{+}09, DCMS03, DC05, DRM^{+}09, DR06, Elm02, EM02, Fis00, GLP^{+}08, GM04b, GJB^{+}04, HR07, HFE07, HM06b, KKKT07, KM02, KCV06, KTMM00, KLAF04, LFS^{+}04, LCS02, LN02, LSK^{+}03, LRS^{+}03, LLDS06, MAC^{+}06, MH04a, MCF^{+}04, MMLH06, MG04, MCL06, MW06b, NCÁSP06, OS05, OH00, OSDB07, OAA^{+}06, OLCT08, OOB09, PLK^{+}05, Par01, PD02, PTLJ01, PQP05, PWN07, PCvRR00, RCL^{+}02, Reu02, RlBW02, RV04, RTT^{+}02, SBC^{+}08, SLE04, SP09, SK01, Smi06, SON^{+}03, SMO07, SH04, SH05b, TCL^{+}09, TVSS01, TRM00, TAB^{+}00, TCCG05, UMASG07, VGP^{+}08, VMC^{+}07, VCKS08, WNL^{+}08]. coastal [WMWV04, WDX01, WG04, WHC^{+}06, WW07, YJMT08, ZHFD06, Zha00, ZB01, ZWC^{+}04, Dow02]. Cobalt [SMCW02, AGL03, SRM05, SG08, TS08, TSG^{+}01, XTSM07]. cobalt-60 [AGL03]. Coccolith [LGN^{+}06, FHB^{+}09]. Coccolithophore [RLP^{+}06, MAL08, XTSM07]. coccolithophorid [BPF^{+}08, EZA^{+}05, RZR02]. coccoliths [GD01b]. Cochlodinium [PLK^{+}05]. cod [KBD^{+}00, CBA01]. coefficient [BS04, CLC02, HMKZ02]. coefficients [MKM06, SSMM00]. coexistent [CC02]. coherence [LP07]. coherent [DJG^{+}06, LD01]. cohesion [HWA03]. cohesive [DRSL07]. cohort [AHF^{+}05]. coin [MBG^{+}05]. Cold [WRG^{+}05, BFS^{+}04, DDL^{+}09, FLT02a, HGBB08, HLGI08, KRM09, RC07, ST03, TBHS01, vODL^{+}09]. cold-water [DDL^{+}09, HGBB08, KRM09, ST03, vODL^{+}09]. coli [GKS^{+}02]. colimitation [BSR^{+}07, BTdB03, NGS^{+}07, SGR08, SG08]. collected [AGL03, BDvO^{+}03, GHN^{+}02, SSL05, TBF04]. collector [PNPA04]. colloid [CW01]. colloid-bound [CW01]. colloidal [BM00, CG09, CW01, CGL03, CG03, FW07, GSW03, NP01, SW05b, WG01]. colloids [GST^{+}01, GDL^{+}00, HA01]. Colne [DNS06]. colonies [ACW^{+}03, BLHC08, LMPY00, NT08, OGR08, SJRB09, SWCL01, WLRB03]. Colonization [DVV^{+}07, Kiø00, BMT^{+}04, CSSNA02, MKG04]. Colony [LV01, CPHM04, CEA^{+}09, Fab06, PSJ00, VMS09]. colony-forming [PSJ00]. color [BST^{+}08, DZL^{+}03, DKRZ06, LES^{+}05, MF03, PC02, TPFS09, WSC^{+}08, WBR03b]. Colorado [BMC07, EPP^{+}08]. coloration [JS03b, VK06]. Colored [BGV02, ACZ00, DGCMK00, DKRZ06, SJ00, TNO^{+}08, ZSS^{+}08]. Columbia [BBP^{+}06, LCS02, NK05, Tun00]. column [Agr05, Aki00, BEM^{+}02, EDF^{+}03, HST^{+}02, JDS08, LCLP02, MS05, SKR01, TMA03]. columns [KL01]. combinations [GCG05]. Combined [AMS^{+}08, BBG^{+}05b, MBK^{+}09, GL07, HHB^{+}09, LBC06, MBC06, TMA^{+}01]. come [WGN^{+}08]. Comment [Sch06a, AP00a, Ben00, BK01b, CP03, CR08, HMKZ02, MCLM02, MWE^{+}04, RL00, SH10, SRG00, SR04]. comments [JSJ03]. common [GH00, HMKZ02]. communities [AS02, AH01b, BST^{+}05a, BSJ02, BGC^{+}05, BdG02, Ceb02, CLC02, CSK^{+}07, CH05b, DTFL03, DdG09b, FAC05, FHDKR03, FF04, FSSH03, GHM^{+}09, GVT06, GV09, GHJ^{+}05, HDSC09, JCK^{+}08, KDS03, LSK^{+}03, LA06, LED06, LCLP02, LSF^{+}08, LZ09, MAC^{+}03, MGL02, NDCP06, NGS^{+}07, OBB^{+}04, QLD^{+}02, RV04, RJS^{+}03, SKJ^{+}07, SlBWC^{+}00, SLRM03, SCV04, SSS08, SMH^{+}02, TS08, TRM00, WDVA08, WŠN^{+}07, WNAW09, WA01, WM06, WKW09, WCL^{+}07, dGB02]. Community [AS00b, AR06, BMT^{+}04, BSJM06, DVV^{+}07, ERH06, BKM05, BBLB05b, BBLB05a, BLM^{+}06, CJ08, CGH^{+}04, Car01, CWJ09, CSK^{+}07, DGM^{+}05, DGD04, ES00, FAM04, FGM^{+}03, FWM^{+}05, FTC^{+}06, FTY^{+}06, GLP^{+}08, GCG^{+}02, Goe02b, GAM^{+}01, GSJ^{+}09, Har02, HRGM07, HBM06, JDS08, KCWC06, KR08, LJ01, LLT05, LFS^{+}04, Lee01, Leh00b, LFAB06, LBSY02, MLC^{+}07, MKG04, MBR08, MY05, NSM09, NEE08, NHTF01, NML^{+}05, PS06, PCvRR00, RORR^{+}09, SEC08, SABA02, SK01, SWML08, SG02, SSAH05, SL01, SMG06, TCC06, TCD^{+}03, UHB^{+}08, UVY^{+}01, VDK04, WBD02, WBB09, WWS^{+}00, vdSvdSG^{+}00, vODL^{+}09]. Comparative [ACD10, ACW02, SH10, Cal01, JCW08, LRS^{+}03, OAA^{+}06, WLRT06, ZRL06]. compared [ET08]. Comparing [DJT06, SBB00, WW02]. Comparison [AP00b, Ano04, ASAC04, FSBB05, KSWC^{+}03, MMLH06, BMU04, EPS00, FMG01, HMT^{+}00, HCD^{+}09, JS03a, LN02, PGM^{+}09, THA01, WRD01, ZMK^{+}08]. Comparisons [EVW07, WBB09, XLF^{+}03]. compatible [Har01]. competent [PPL^{+}02]. competing [QYEA06]. Competition [ARV^{+}07, OB04, OHT06, CPCS01, CPCLS04, HTS^{+}03, JM02, LKB04, LTT07, MTV^{+}02, vNHM^{+}09]. competitive [HZ00, TRT04]. competitors [PMK08]. Complementary [HBG^{+}09, SAP^{+}06]. complete [GHTJ00, RSS^{+}00]. Complex [LW02, SWH^{+}05, SSH^{+}07, SJH03, BvE07, CWWS00, Fly03, GBL^{+}04, HLGI08, HKNG08, KTM01, KS04b, KTMM00, LN04, MHK^{+}07, MSVAC04, Nor04, RS02, SBS07, SS06, DDL^{+}09]. complexation [BMC07, QSH^{+}02, TWS01]. complexed [CMQ^{+}00]. complexes [SRNW06]. complexing [CMB00, SWS^{+}06]. complexity [Gay00, MWH01]. component [OH00, PED^{+}09, WNP04]. components [CSK^{+}00, ZHSJ09]. Composition [GWL^{+}07, KIB01, RV04, AFB01, BBKC03, BMNB^{+}06, CJ08, CV05, CMSVA02, CC04, Car01, CPH06, CCH02, CSK^{+}07, CRF01, DSD02, DGM^{+}05, DKt08, EB01, FHBK06, GCG^{+}02, GR01, GCS06, HY01, HM03a, HRGM07, HSN^{+}03, HSCF06, HWYL07, HPHT00, JDS08, KB06, KF01, KR08, LLT05, LBB^{+}04, LG04, LB04b, LRMV04, LMBP07, MH00, MH07, MPACG01, MCPP06b, MG04, ME07, NSM09, NML^{+}05, PKKB05, PTK^{+}07, PWS09, PS06, PF01b, Pri05, RORR^{+}09, RKNA09, SCDR07, SL01, TSSWK^{+}06, TGFPT08, TTG^{+}09, VWPP^{+}06, VC05, VMS09, WR00, WKFP06, WŠN^{+}07, WR03]. Compositional [BdG02, MM03, GCJ^{+}06]. compositions [CYN05, LRYS06]. Compound [EPP^{+}08, ZWC^{+}04]. Compound-specific [EPP^{+}08, ZWC^{+}04]. compounds [ART^{+}06, BSS^{+}01, EPP^{+}08, HMJ04, HBG^{+}09, LPD^{+}04, MGW^{+}05, TMA03, VCKS08, ZHFD06]. Comprehensive [BDK^{+}06b]. compressibility [GBH^{+}02]. concealed [LWU^{+}05]. concentration [AY08, BMFS^{+}03, EZA^{+}05, GAPR06, KLS^{+}06, KLS^{+}07, MB05, NKS^{+}02, PKKB05, QYEA06, RPR^{+}02, RHJ04, SGPK07, SL01, SLS^{+}02, TLS01, TMA^{+}01, WSL^{+}02]. concentrations [All00, AY08, BDK^{+}08, BARS01, CMQ^{+}00, CPH06, CBD^{+}06, CMCB^{+}09, ESA06, GR01, GMRVFP02, HW07a, JK08, JJT07, KD02, LGN^{+}06, LH05, MPH08, NKJM06, PWH^{+}04, SHA08, TG09, TCB^{+}00, TLrR07, TKAW03, VMKN03, VM05, WK09]. concept [SGR08]. concepts [HMKZ02]. conceptual [LNB04]. Conch [SSV^{+}04, SWML08]. Concholepas [PPL^{+}02]. condition [DBS^{+}04, MACB07, RRSC01]. Conditional [PP00]. conditions [AS06, BFC07, BSS09, CRBF^{+}06, GKS^{+}02, HP06, JVL00, LLF^{+}09, LCC^{+}09, MB00, Rit09, RQB^{+}00, Sar07, SGPK07, WJA^{+}07, WTM09, WWS^{+}00, dlM09]. Confirmation [SO07]. conflicting [SrL05]. confounding [SA09]. congeneric [GRDCS09, LMPY00]. conifer [GVT06]. connected [RC05b, VVD^{+}09]. Connecticut [HMB06, RB01b]. Connecting [WCM09]. connection [GSG^{+}05]. connectivity [JGE^{+}09, MS06, Reu02, SGPM09]. Conochilus [Ham05]. Consequences [Dow02, HNB^{+}07, HLNM^{+}03, JLB^{+}06, MM05, BVG^{+}07, Car02, CK00, FZL^{+}05, JWRW07, LSK07, RRCP03]. conservation [GBL^{+}04, Bri01]. conservative [UGSMN01]. considerations [SCARM00]. conspicuousness [TKK02]. Constance [AIK04]. constancy [TKAW03]. constant [HMS^{+}04, SCARM00]. constants [TRT03, TvdWdB04]. constituents [BJG^{+}00]. constrain [LBG^{+}09, RC07]. Constraints [BLS^{+}06, LBM^{+}04, Aks07, Ben00, BKM05, FMD03, Law08, MB01, MCvEH08, RB06, VVSM09]. Consumer [ABJ^{+}06, BS01, WLS00]. consumers [AFH07, BH02, CVH^{+}05, KB05a, KKH04, MVFD04, NH06a]. consumption [ACZ00, Ceb02, Duf09, GSB^{+}09, HPAM01, HBN^{+}00, PDR^{+}09, SS01, SS04, dVKT^{+}09]. containing [CWPL05, PP07, dGMH^{+}08]. contaminants [FCJ^{+}06]. contaminated [BMU04]. contamination [LVSW07]. contemporary [CCKC05, JGE^{+}09]. content [AS00a, CG08, DPC^{+}08, GHN^{+}02, JP03, JTL01, MF01, MF02, RB08, RB06, SAP^{+}06, SKC^{+}01]. contiguous [FS09a]. continental [GSO03, GvRE^{+}02, HCC^{+}02, LSPC02, PGP01, TN09, UWA^{+}07, YSO02, vODL^{+}09]. continuity [Shi04]. Continuous [CFO^{+}00, GD00b, LB04a, MB05]. continuum [VWPP^{+}06]. Contrasting [CJWR05, KSM^{+}05, LRF07, NSM09, SFR^{+}08, BHMS06, CJ08, DLZ^{+}09, GDCG08, OH00, OAA^{+}06, OCNH00, Reu02, TLM04, KGYU06]. Contrasts [Arn00]. contribute [VNH^{+}08, WDL07]. contributes [BRW07]. contributing [MJEM06, MJ03]. Contribution [ÁS07, BVK02, CD03, FVD01, FLT02a, KLC06, MF03, ADA00a, ADA00b, All00, BK01a, CDBC00, GVT^{+}09, Ham01, HABF03, KB08a, KB08b, KTG^{+}06, LV01, MBR08, RH06, SSH^{+}09, WHHS06, ZB03b]. Contributions [GSO03, TSMB05, MPdVG07, TB09, VVSM09]. Control [BJD05, GFPG09, KJMJ03, KA06, ADA00a, ADA00b, ANSS04, Alt01, BA01, BZS07, BMC07, BB09, CCS06, Ceb02, DPFL09, DHTH05, ET08, FS00, FAM^{+}06, Goe02b, HMS^{+}07, HBAK09, HZ00, HVH00, HSB^{+}04, HFZ^{+}07, JTAM09, LvdB09, LCS^{+}03, LFAB06, LMMW03, LWS01, MSAC06, MW06a, MSVAC04, MNM^{+}09, Mul01, RCAK05, RB03, SC00, SVB^{+}08, SMO07, TKBJ05, TMA03, WCL^{+}07, YYG02]. controlled [KLS^{+}06, KLS^{+}07, LAHdBJ05, MGMW06, SCH07, SDKZ^{+}09]. controlling [BTS^{+}06, EF05, GCG^{+}02, GCS06, LFS^{+}02, LVC^{+}00, TB05, Thi00, vDBL05]. Controls [BCC^{+}04, DAV04, HTRM^{+}07, MLPS09, AGP07, BLWB08, BMD^{+}04, Bou04, CKA03, DD09b, DS07a, DACG01, ERH06, Fin04, FCJ^{+}06, FUG02, HT03, LW02, LLDS06, MMO06, ME07, SR02, SR04, SCDR07, WvE04]. convection [SLHK02]. convective [LAG^{+}03, LPW05, MVC^{+}05, OS01a]. Convectively [FLPL08]. Convergences [SMS07]. conversion [Shi04]. conveyer [JCBA08]. conveyor [PDS07]. Conwy [DNS06]. Cooling [TYK^{+}08, PFH^{+}03, WS01]. Cooper [HBTM05]. coordinates [RWFK00]. Copepod [DKt08, HP06, AD07, AAD08, BF08a, BB04b, CF05, Has04, HKF01, HW09, Kiø07, LH06, MB00, MGW^{+}05, PIKB08, ROP^{+}02, SSH^{+}07, SMDG04, TVT04, WSEL09, WGC00, XW02]. copepodites [Has06]. copepods [AP00b, BSSL05, BHL07, BS01, CK08, Goe08, HB03, KB05b, KSEÖ^{+}02, KKT05, LUHS04, Møl07, PD05, SCB03, SC07, SSH^{+}09, SAS^{+}03, SvKM01, SSS06, SF03a, WWWY05, ZRL06]. coping [MPGB08]. Copper [MAC^{+}06, MAMC02, PQP05, QRF06, AKH04, BMC07, CZRD^{+}04, CB04, Cro03, DGD04, DNB^{+}04, DA05, GW03, PFN^{+}05, RRD04, SWS^{+}06, TWS01, TSKA05, TMB07, WTC^{+}05]. copper-binding [DNB^{+}04]. copper-complexing [SWS^{+}06]. Copper-dependent [MAC^{+}06]. copper-limited [GW03]. Coral [HGFS^{+}05, dGvD07, AS00b, ACHG07, AP00c, BSM01, Bat02, BPC06, BF08b, BTKGA07, CEA^{+}09, CCC^{+}01, DDL^{+}09, DOE^{+}06, EMIP05, EGP08, Fab06, FDP^{+}08, FAL01, FAC05, FMWC00, FSSV08, FB05, GRVF^{+}00, GMR^{+}09, GEP08, GFPG09, GCM^{+}09, GMRVFP02, GMAFP03, HMSI06, HLM08, HG03a, HG03b, JM02, JLP^{+}08, JP03, LBW^{+}08, LGJ02, LSM03, LRYS06, LLF^{+}09, LSF^{+}08, MSL^{+}03, MKM06, OFKH01, PC04, PSH^{+}07, PSS^{+}06, PVD^{+}07, QYEA06, RMK^{+}06, RKK09, RCAK05, Ris02, RG07, RRHPT^{+}06, SE06, Shi04, SLKP05, SMH^{+}02, SSL05, TMK^{+}07, TGFPT08, TTG^{+}09, TGF00, UMY^{+}02, VNH^{+}08, WLRT06, WAB^{+}08, WLRB03, YYG02, YYB^{+}05, ZJ03, ZSS^{+}08, dGMH^{+}08, vODL^{+}09]. coral-algal [LBW^{+}08]. coral-reef [FSSV08, HG03a, HG03b]. coralline [Chi00]. Corallistes [EKdF07]. corals [ACW02, Fab06, FAL^{+}07, FMWC00, FBW^{+}08, GHWG04, GKLF01, GF02, JKLR08, LLvW^{+}03, LMPY00, MF03, MLG^{+}03, OFKH01, RKM^{+}06, SLDE09, WLRT06]. cordgrass [New01]. Core [HBH^{+}07, BK01c, SA09, TB00]. cores [Boy01, SDJT03, TN09]. correction [STST07]. correlates [VMS09]. correlation [BRW07, BMWW08, CW02]. correspondence [ES00]. cosmogenic [BNK02]. cosmopolitan [CCGH05]. cost [GBA00]. Costa [SRM05]. costatum [AMP^{+}00, BB04b, SNM^{+}08]. costs [AD07, SFR^{+}08]. cotolerance [NRB08]. Could [Cul06, HSF^{+}08, RFW08]. Coulometric [Too01]. Coulter [Gou01]. counteracting [ROP^{+}02]. Coupled [GSNH04, GDL^{+}00, MH07, Nil03, SAPA^{+}02, AJ01, BNM^{+}09, CCKP02, HD03a, JNKdV09, RB02]. Coupling [CV05, GVT06, SRZ04, ALH01, AGP07, CVH^{+}05, CL02, CLL^{+}09, DAV^{+}05, GM04a, HDSC09, HBF^{+}06, LLT05, MCV^{+}05, NFFK00, NCPd08, NHTF01, SR07, SKBC01, TS01, VMH^{+}01, VMS^{+}04]. cover [BSG^{+}08, LA09, MSL^{+}03, MVA^{+}09]. covered [PCD^{+}08, VGP^{+}08, WSNH06]. covers [FS09a]. crab [EMPI07, HFC^{+}04, HH05, PDS07]. crabs [JWRW07, KA06, REL06]. Crassostrea [KN07, TG09]. crater [SKTD04, UVY^{+}01]. Crayfish [SFC^{+}00, KTM01]. CRC [Bel01, Elm02, Kne02, Ris02]. Creek [AD05b, HBTM05]. creeks [BdBS^{+}08, GHS05]. Crepidula [MTC^{+}06]. crepuscular [YYB^{+}05]. Cretan [BLM^{+}06, SH05a]. criteria [OIH^{+}03, SCS^{+}08]. critical [LG04, VHL^{+}04]. Critique [EWE02]. Crocosphaera [FMG^{+}08]. crops [PCA^{+}00]. Cross [GCJ^{+}06, HCD^{+}09, GZL02, GDL^{+}00, LN02]. Cross-basin [HCD^{+}09]. cross-flow [GDL^{+}00]. cross-latitude [LN02]. cross-shelf [GZL02]. Cross-system [GCJ^{+}06]. Crossed [GRDCS09]. crowding [Gil03, SS06]. CRS [SCARM00]. Crustacea [DD09a, Har01]. crustacean [BFS02, RYS^{+}02, CRY^{+}08]. crustaceans [WHHS06]. crustose [Chi00]. Cruz [BBG05a]. Cryptic [JS03b, MVD05, DST08, PH07]. Cryptomonas [UGY^{+}00]. ctenophore [Ano08b, CSG^{+}06, GLVFH09]. ctenophores [PD05]. Cu [ALRM08, CMB00, CBvEK03, MHRW02, MJW03]. cucullata [DGV01b]. cue [NA05, SS03]. cues [DHKH05, KDO09, KRM09, LBT06, WW01]. Cultural [FHBW06, VJV^{+}03, CVH^{+}05, Gou01]. culture [DGL^{+}04, EKD^{+}07, GD00b, LB04a, SBL06]. cultured [MC01]. cultures [Ano04, BS03, DA03, GW03, GSNH04, KSWC^{+}03, MB05]. Cumbria [GHTJ00]. Current [AO09, DORS08, ÁÁS07, HP06, HHW^{+}02, IOG^{+}04, MBR08, NCPd08, SMA^{+}06, SLKP05]. currents [CFG07, DDL^{+}09, FNM07, GZL02, GDK03]. curve [BMSA08]. curvirostris [DPK^{+}00]. cuspidata [TWC^{+}09]. cyanate [KCP08]. Cyanobacteria [KCP08, WA09, ARV^{+}07, BRT^{+}03, BSC^{+}09, BD07, BFQF^{+}07, CPHM04, CK00, DGV01a, FSM^{+}07, JJT07, KSEÖ^{+}02, LB09, MAMC02, MCvEH08, MHKP02, MPRC02, NKE08, NFS^{+}07, RSY04, SW05a, Sar07, SEL^{+}02, TRT04, TWR04]. Cyanobacterial [BEW^{+}00, Plo08, Ber01, GEJ^{+}02, HMJ04, KNE08, Kut04, MPZ08, MGVM^{+}06, SPG05, SKH^{+}01, VW01, WST06, WH07]. cyanobacterium [BVG^{+}07, BFBHF04, CT06, CBvEK03, FMG^{+}08, HKNG08, IB05, MTV^{+}02, PM08, RSWH04, RES^{+}08, SHSH^{+}07, STHG07, VMS09]. cycle [AR06, AG03b, BBAW09, BBG^{+}05b, Cro03, HPE01, HKNG08, LB09, Lee01, Lon00, MCV^{+}05, MC06, MN04, MAL08, PTKP00, SGG^{+}08, SMDG04, WMRK06]. cycles [HMS^{+}07, JS04, JKS04, SLDE09, YYG02]. cyclic [GSB00, HD09]. Cycling [GSW03, MWW06, SMH^{+}02, BGV02, BNK02, BNOK^{+}04, BRL03, CL04, CT05, Chr05, CCF04, DD09a, Dal03, DGL^{+}04, ESSS01, EF05, FCGC^{+}09, GM04b, GTS^{+}09, GFM^{+}01, GK03, GSH04, HZG07, HST^{+}02, JBH^{+}04, KB07, KFW09, KIB01, LBM^{+}04, LF09, LZ09, MPJ08, MAM07, MA08, RB01a, RA06, RTT^{+}02, SAPA^{+}02, SdB06, SH00, TMS^{+}00, TWS03, TS08, TDC^{+}09, VDK04, WKFP06, WCL^{+}07, Wur07, XBD^{+}09, ZXN^{+}08, ZB00, vODL^{+}09]. cyclone [FDP^{+}08]. cyclones [SCM08]. cyclopoid [MB00]. Cyclotella [AGL03]. cylindrica [JM02]. cylindrical [PNPA04]. cylindrus [KMvD^{+}09]. Cymodocea [BMT^{+}04]. cyst [BZS07, KRM09]. cytoarchitecture [BR00]. cytometer [SON^{+}03].

D [AE08, Pom01, SMS09, SG02]. D. [JS03a]. D.C [Dow02]. Dagow [CCC09]. Daily [YYG02]. dam [AD05b]. damage [ROP^{+}02, RMS^{+}08]. damaging [GWK01]. Damariscotta [FW07]. dammed [SCY00]. damping [SI08]. Danish [CCAÆ06]. Daphnia [AFL07, AKE04, AMW^{+}07, AY08, BSMR08, BAPMMB01, BB03, BB05, BD07, BRE02, Boe00, BMNB^{+}06, BvE07, BJL01, CT04, CWPL05, DGV01a, DGV01b, DLDE04, DPK^{+}00, FXL04, GPAP^{+}04, HCH00, HW07a, HW08, HWA03, HD09, HM03b, HR00a, JS03a, JLH01, KVdB^{+}04, KS04b, KWTS07, KBE^{+}04, LBT06, LW01b, LV01, Lür03, MCvEH08, MVD05, MSJMN02, PvE00, RK02, RBMN03, RB06, RCD^{+}05, Sak00, SD05, SW05a, Sar07, SSLE02, SJR^{+}04, SBS07, SS01, SS04, TW09, VW01, VPA02, WGKE04, WH07, WKB^{+}04, YW02, vE02]. daphniids [TBG01]. dAPS [Tho09]. dark [BAG^{+}09, CGCPA08, GZY^{+}08, WKW09, AGDH09]. data [AHHH^{+}07, BJBA03, KSS^{+}07, RBW09, WMW^{+}08]. dataset [MGRE^{+}07]. dating [SCARM00]. davisae [Kiø07, SCB03]. dead [LLF^{+}09]. Death [GWR^{+}09, BFBHF04, LA06]. deaths [JKS04]. debris [SD00]. Decadal [GRB09, LES^{+}05, AO09, BDK^{+}06b]. decades [HM06b]. Decapoda [CRY^{+}08, Har01]. decay [LRF07]. decaying [KLSW00, New01]. Deciphering [SA09]. decline [KB06]. declines [CGWK^{+}09]. Declining [Sch00]. decomposing [DVV^{+}07]. decomposition [BLWB08, CA08, SH10, VW08, WG01, ZGP^{+}00]. decoupling [EGPA02]. Decrease [HYH04, NHPHH08]. decreases [BSMR08, GCM^{+}09, Jan03, VMKN03]. Decreasing [BF08b]. Deep [PFH^{+}03, VBH^{+}05, AHHH^{+}07, AVJ^{+}00, AELS04, BVWB07, BBJ^{+}06, Bou04, BLM^{+}06, CMC^{+}09, CMCB^{+}09, DDL^{+}09, DH01, Dit04, DS07b, DKKA09, EKdF07, EPD09, FLT02a, FLT02b, FH03, GSG^{+}05, GTS^{+}09, GSB^{+}09, GSFB02, GSL^{+}07, HPAH^{+}02, Jun01, KSM06, KHH^{+}03, LAG^{+}03, LSPC02, NFFK00, NEE08, NDP^{+}07, NHN^{+}05, PSLO07, PGM^{+}09, PW03, PMPC00, RS01, RLJdB08, SO07, SL02, SR02, SR04, SJH03, TLvAH06, TTSC04, VRC07, WBH03, WRG^{+}05, YWR^{+}07]. deep-ocean [NDP^{+}07]. deep-sea [Bou04, CMCB^{+}09, DS07b, EKdF07, EPD09, GSFB02, Jun01, LSPC02, NHN^{+}05, PMPC00, SO07, SL02, SR02, SR04]. Deep-water [PFH^{+}03, VBH^{+}05, AELS04, DDL^{+}09, DKKA09, GSG^{+}05, HPAH^{+}02, WRG^{+}05]. deepening [MRK02]. deepest [BL07]. deepwater [DJG^{+}06]. defend [KKH04]. defense [CF05, KSP07, SWH^{+}03, SWS^{+}03, WH07]. defenses [BH02, CT06, LBT06, SD05]. defensive [vDBL05]. deficiency [DGV01b, DP01, FGM^{+}03, LBC^{+}04, MH07]. deficient [DH01, dSFFTD07, HW08, PP04, TDJ^{+}06]. deficit [ME06]. Definition [CHH07]. definitions [SGR08]. degassing [JCW08, WC04]. degeneration [BII05]. deglaciated [RM06]. Degradation [ART^{+}06, BPOT04, BLG^{+}07, CS06, FPPH06, GWL^{+}07, GP01, HR00c, KKT05, MSC^{+}04, MMCW09, MSZ00, NMK03, SM05b, SŻKL08, TVL^{+}00]. degrading [GHM^{+}09]. degrees [SMA^{+}06]. del [Wan02]. Delaware [CD03, HU07, KDMC05, YNCK04]. delayed [FL06]. delivery [FQH^{+}01, MSSW^{+}07]. delta [BDSB07, CGL^{+}04, MSJMN02]. demand [ASGA^{+}07, ORH01, SKBC01, SH05b, TPM01]. demersal [TMYH02]. demise [BFBHF04]. democratica [LIE06]. demographic [SGPM09]. demographics [PF01a]. Demonstration [SLB^{+}00, WLTF03]. denaturing [DMEPA04]. Denitrification [FCGC^{+}09, GMRD04, MHS^{+}09, RGRPD04, AJ01, ATJH08, BHV04, DNS06, HLK^{+}07, JKGT08, KD02, MVW^{+}04, Nil03, PCvRR00, SLL^{+}04, TN09]. denitrifying [GW03, GSLT08]. Denmark [PKKB05]. densities [AH03, FMWC00, JJT07, SLC^{+}09]. Density [DD09a, Ber03, BZ02, GBH^{+}02, HPV^{+}00, HL06, HG03a, HG03b, PP00, QYEA06, SS03, WCGBG07]. Density-dependent [DD09a, HL06, PP00, SS03]. depend [GM03a]. dependence [BFD^{+}04, CPHM04, KJMJ03, LGN^{+}06, PP00]. dependent [DD09a, DJO00, De 02, FSBB05, GBA00, HL06, HBS04, LLK05, MAC^{+}06, PP00, RZR02, SS03, SAE^{+}08, WGD^{+}02]. depends [DK06]. depleted [WJA^{+}07]. depletion [HSF^{+}08, JLB^{+}06, JPK^{+}05, KCSM07, RM09a, TvdWdB04]. deployment [FDH^{+}01]. deposit [CMQ^{+}00, GM04a, KNE08, PW01, RB08, RHM04, SM02, SdB06]. deposit- [RHM04]. deposit-feeder [PW01, SM02]. deposit-feeding [GM04a, KNE08, RB08, SdB06]. deposited [JM08]. Deposition [NTM^{+}05, BBJ05, BBJ^{+}06, CH05a, DFBA07, DT00, EBS07, FT03, GDM02, HSB^{+}06, HAL^{+}09, LJ09, MOK^{+}03, MED^{+}08, MLRMR09, MTM^{+}01, NDP^{+}07, SEC08, TNM^{+}01, TSSWK^{+}06, Tun00]. deposits [GJB^{+}04]. depress [TMP02]. Depth [DS07b, TS00, BD07, BK01a, Bou04, CMJ07, DET04, FBW^{+}08, HKF01, HM03a, JDS08, KCCE06, LLS08, LW01b, MKM06, SD05, SR02, SR04, SHS03b, SLL^{+}04]. Depth-related [DS07b, TS00]. depth-selection [LLS08]. depths [BDK^{+}06a, PDB03, SBR^{+}06]. derived [BLSQ07, CBA01, PS06, SBC^{+}08, SAH^{+}04, SMD^{+}04, SBL06, SM01, TCP^{+}03, THJ^{+}09, UM05, UMY^{+}02, VW08]. describing [JPJS06, WSNM00]. Description [Har01]. desensitization [vDBL05]. Desert [GBL^{+}04, KR08, KBG^{+}08, Vin00]. desferrioxamine [MHB01]. design [LW01a]. designed [MPACG01]. Desmodesmus [VVD^{+}09]. detached [GD01b]. detailed [NSBA01]. detected [GDB^{+}09]. Detecting [JLH01]. Detection [YT03, APH^{+}03, DWA^{+}02, Kut04, YTVR07]. determinant [Bat00]. determinants [HAL^{+}09, XLF^{+}03]. Determination [HKD01, MMS^{+}02, SHS03b, vE02, DHKH05, HU07, MSL^{+}03, Mor03, TPFS09]. determine [MC04]. determined [BYF09, KJJ07, KKH06, Lan03, MWH01, RlBW01, RWFK00, SMS09, VB00, WBR^{+}03a]. determines [ATJH08, KDO09, vNHM^{+}09]. Determining [MHK^{+}07, LAG^{+}03, LCLP02, MMW00, PWN07]. detrimental [MSVS^{+}07]. detrital [CSSNA02, KB08a, KB08b, LLB07, SPM04, TB09, WNL^{+}08, vNHM^{+}09]. detrital-based [SPM04]. detritivores [CJWR05, EW07]. detritivorous [vNHM^{+}09]. Detritus [CCC^{+}01, BA01, MSHE09, RPR^{+}02, XW02]. detritus-based [RPR^{+}02]. detritus-bound [XW02]. developed [PWH^{+}04]. developing [GEJ^{+}02, MBF^{+}03, SCS^{+}08, vDBL05]. Development [DL00, KHH^{+}03, DS07a, HM06a, HFR08, JHGGP06, MG04, MFG^{+}09, PDTR06, PSLO07, PN06, RNA06, SNFT07, SMDG04]. developmental [MSBHG01, SvKM01]. developments [GFM^{+}01]. Deviating [TKK02]. DGT [FDH^{+}01]. DGT/DET [FDH^{+}01]. diagenesis [CDD07a, FSK^{+}04, HYH04, HRS04, KCSM07, LNJL09, SMD^{+}04, Shu01, TB09]. Diagenetic [ME07, AFB01, BGJ01a, BGJ01b, GRB09, HAES02, MGF03]. diagnosed [HDMB04]. diagrams [WKH^{+}06]. diameter [DCMS03]. diapause [CT04, GRCS06, IOG^{+}04]. diapausing [BDvO^{+}03, GS04, JKS04, KS04b]. Diaptomus [VAS02]. diapycnal [D’A08]. Diatom [IKT03, PAS^{+}05, AMP^{+}00, BA01, BB04b, BRP^{+}06, BRS01, BS03, DKt08, ES00, EM08, EH00, EGPA02, FHDKR03, GR01, GP01, HBH^{+}06, Has04, Hol00, IHHH00, JGZ^{+}08, JGSB07, KTMM00, LJCM08, LBC^{+}04, MBF^{+}03, MMLH06, MH07, MW06b, MF01, MF02, PFL07, PG00, PHK02, PP07, Pri05, RLM^{+}04, RA00, RNA06, RJS^{+}03, RBJ^{+}06, SNM^{+}08, SWN08, ST03, SF04, SSC02, SH04, SH05b, TCC06, TSKA05, VHL^{+}04, vdPAJ^{+}06]. diatom-based [MBF^{+}03]. diatom-dominated [HBH^{+}06]. diatom/Phaeocystis [BRP^{+}06]. diatomas [MVBM04]. Diatoms [HMS^{+}07, Alv07, ALRM08, BARS01, CWPL05, Fin01, Gai00, GM03b, HMS^{+}04, Ham02, HM09, KVdB^{+}04, KFN00, KSM^{+}05, LSK07, LES^{+}05, LFPH07, LLC^{+}09, MAC^{+}06, MG04, MDL00, MTM^{+}01, OLCT08, PP04, PQP05, TS08, TGdB^{+}01, TvdWdB04, WTC^{+}05, WW07, dlM09]. diazotroph [NTDV06]. Diazotrophic [LB09, NFS^{+}07, TWR04]. diazotrophs [KFH^{+}09]. DIC [BCC^{+}04]. Diego [CZRD^{+}04]. Diel [BF08a, BBG^{+}05b, BDC00, CBB^{+}02, DKKA09, KKT^{+}07, SS01, TWR04, UGY^{+}00, YYB^{+}05, ZXN^{+}08, BBAW09, GZL02, HKF01, HM03b, IJA^{+}04, LLS08, SLDE09, TCCG05, SS04]. diet [ABJ^{+}06, BMNB^{+}06, DPC^{+}08, EKdF07, GJS01, Has04, RRSC01, VVSM09]. Dietary [MGW^{+}05, AS00a, DGV01b, SDC00]. diets [BL07, CWPL05, CD02, dSFFTD07, RH01]. difference [GNU00, SSAH05]. Differences [GA01, OH00, DJH05, LAJ^{+}04, LLK05, MBDL01]. Different [AFH07, HPL^{+}06, SR07, BK01a, BLM^{+}06, BARS01, BS01, CSSNA02, CDD07a, EAO^{+}02, FAL01, GR01, GCG05, GWL^{+}07, HW07a, HK01, JJT07, JHW05, KFR^{+}01, LB04b, Lür03, MMDW09, MW06b, MBC^{+}05, MPRC02, MJW03, MWW06, OHCP02, OCNH00, OCSB05, PGC07, PH07, TW09, WWS^{+}00, ZRL06]. Differential [BB05, FSSP02, JS08, LTM01, PWX04, PS06, WDVA08, AAD08, GDL^{+}00, HRGM07, LBC^{+}09, PFH^{+}03, SMDG04, TFNT07]. differentially [PSH^{+}07]. differentiation [Ham05, PCDB09, SJR^{+}04, SBS07, SM01]. differs [BGE04, LSK07]. diffuse [BLG^{+}07]. Diffusion [PHK02, GGRJ03, MNSV03, RENA06, SL02, YPBL07]. Diffusive [NA09, PSJ00, GBFJ07, LMMW03, RM04, RHJ02, RHJ04]. diffusivity [PP07, PIF08, WdBD^{+}01, WPV00]. digestion [VMJ08, ZB03b]. digestive [SM02, ST03]. dimensional [FG06, LH05, PGM^{+}09, RWFK00, SF04, ZB03a]. Dimethyl [RA09, SSS06, EMWL06]. dimethylated [LPD^{+}04, VCKS08]. dimethylated-sulfur [LPD^{+}04]. Dimethylsulfide [ZHFD06, DJB^{+}07, KL00, SAPA^{+}02, SKNK09, TKK^{+}03, dVKT^{+}09]. Dimethylsulfoniopropionate [MKVK05, BS03, EKD^{+}07, MKK04, SAPA^{+}02, SWS^{+}03]. dimethylsulfoxide [SKNK09]. Diminished [BBL^{+}03]. diminishing [GSG^{+}05]. dinitrogen [Ano04, BSC^{+}09, KSWC^{+}03, PLW^{+}08, PM08]. dinitrogen- [Ano04, KSWC^{+}03]. dinocysts [MPdVG07]. dinoflagellate [AD07, AAD08, BZS07, CFG07, DCMS03, DA03, HHB05, JVL00, JTL01, KHNV05, KSP07, LASR09, MPdVG07, PMK08, SEL^{+}02, ZGP^{+}00, vDBL05]. dinoflagellates [AR06, FMWC00, KBBJ00, KRM09, LES^{+}05, LN02, LNB02, MDL00, WLRT06]. Dinophyceae [GGP01]. Dinophysis [GGP01, SGG^{+}08]. dinophysistoxin [KUP^{+}06]. dinophysistoxin-1 [KUP^{+}06]. Dinosterols [MPdVG07]. dioica [AK00, LUHS04, TFNT07]. dioxide [BMD^{+}04, CGCPA08, CS05, HPBD01, JCW08, KFR^{+}01, Sch01, SWN08, SKC^{+}01, TG09, WC04]. dipeptides [ML09]. Diptera [BPLH04]. Direct [EMPI07, GP00, HGB00, HU07, HMD^{+}07, JKS04, MdG08, NP01, PC04, PP07, PIKB08, RCL^{+}02, WRG^{+}05, EZA^{+}05, WH07]. directed [PVD^{+}07]. Directly [HD03a]. discharge [CB04, Cha07, CIWH09, GGM03, GMMW03, HU07, HDFM08, HLK05, KH03, KC08, PSS^{+}06, SSA^{+}04, SMGP06, THA01, TSDK^{+}09, dCCL09, dSYL^{+}08]. discharges [PBGJ09, TSSWK^{+}06]. Discontinuities [AB07]. discontinuous [HC01]. discovered [AD07]. Discovery [AP00c, GGS06]. discriminate [FSSV08]. discrimination [KMMS00, KDS03, MVBM04]. disease [Duf09]. disequilibrium [EPS00, MT06, MWE^{+}03]. dislodgment [CMD^{+}09]. dismutases [PP04]. disparity [PHF03]. Dispelling [TF00]. Dispersal [AHF^{+}05, GSA^{+}05, KR08, PHMR^{+}07, AG03a, BPC06, DGCMK00, DL00, HCH00, HS04, HMD^{+}07, HR03, JGE^{+}09, MVD05, REL06, WNAW09, WM07, ZRL06]. dispersion [JLP^{+}08, LN06, SI03, UGSMN01]. dispersive [FRO04]. disposed [FH03]. Disrupted [WSEL09]. dissimilar [RES^{+}08]. dissimilatory [GMA^{+}06, GSLT08]. dissociate [JGE^{+}09]. Dissolution [FH03, GJB^{+}04, LVB08, SM02, BM00, BBL^{+}03, BJBA03, BZ02, BZH08, FAC05, GWR^{+}09, Gre09, JNKdV09, PSS00, RBJ^{+}06]. Dissolved [BFC07, BI07, BLBB07, HBH^{+}07, HPAM01, HDVB01, LBC06, OH08, RSH08, SH05a, TPSF01, VMSB08, AW05, AVJ^{+}00, All00, ÁSGMP01, AFB01, ACZ00, ARK^{+}00, ASS03, AHR^{+}01, BFDC01, BFC05, BMF^{+}06, BPOT04, BJG^{+}00, BGV02, BK03a, BRL03, BT00, BK03b, BTS^{+}06, BW00, BMC07, BB07, BM03, CWWS00, CKPS04, CGH^{+}04, CGL03, CG03, CDK02, CLG08, CSK^{+}07, CSPW00, CS05, Cul06, DD09b, DGKA06, DB07, DGCMK00, DKRZ06, Dit04, DGD04, DdG09a, DR06, EB01, FAL^{+}07, FS09a, FQH^{+}01, FSSH03, FSSP02, FW07, FXL04, GBM^{+}04, GFJ^{+}02, HY01, HBCK03, HSR^{+}08, HKD01, HFDO05, HFE07, JR07, KB01, KS04a, KB06, KD02, KWS06, KL00, KKH06, KNN06, KHK^{+}03, KCV06, LFZ^{+}07, Lee01, LBM^{+}04, LBG^{+}09, LFFC06, LBD^{+}09, LTT07, MYM^{+}05, MPJ08, MH00, MH04a, MCF^{+}04, MHK^{+}07, MBC06, MBW^{+}01, MCPP06b]. dissolved [MHB03, MMCW09, MPD^{+}06, ML05, Møl07, MSZ00, MC04, NSM09, NA03, NCÁSP06, NA06, OH00, ORH01, OB04, OV03, ORFD^{+}09, OZM^{+}01, OOB09, PC02, PWX04, PWH^{+}04, PS06, PAF^{+}04, PBGV04, RKGC^{+}01, RBC00, RA06, RvBB02, RS01, STC03, SMD^{+}04, SHL^{+}05, SDvEB06, SR07, SRG00, STP^{+}07, SMH^{+}06, SlBWC^{+}00, SWML08, SM01, SM05a, SM05b, SHA08, SC08, TPM01, TvG02, TYK^{+}08, Tho09, TNO^{+}08, UWA^{+}07, VSSTS00, VJ07, VW08, WKH^{+}06, WW07, WK09, WV00, WdMD^{+}00, WS04, WTM09, XLF^{+}03, YT03, YJMT08, YT09, ZSS^{+}08, ZB00, ZWC^{+}04, dGvD07, dGMH^{+}08, dVKJK07, vWSBT08, vWBT09]. distance [HMD^{+}07, ZRL06]. distances [HR03]. distinct [BLBB07, FTY^{+}06, Fyf03, KWTS07]. distinctive [TNO^{+}08]. distinctness [HSLV05]. Distinguishing [DJH05]. distributed [STP^{+}07]. Distribution [BB04b, CSSNA02, CGL03, CG03, DMEPA04, GSB00, HHM^{+}09, HMB06, KL00, MGF03, RKGC^{+}01, RCE^{+}02, SLE04, SGT^{+}06, TLvAH06, UWA^{+}07, Ann05, BKAE01, BJW^{+}02, BSJ02, BZS07, BRM^{+}07, BBLB05b, BBLB05a, CDD^{+}03, CMJ07, CMC^{+}09, CW02, DPC^{+}06, GD00a, GBFJ07, GSB^{+}09, GGS06, GSL^{+}07, HD03a, HDWDD09, HOFD01, Hol00, HFDO05, JL08, KLWS08, KFH^{+}09, KL01, LSK07, LW01b, LCC^{+}09, LBD^{+}09, MNSV03, MZV03, NFFK00, Par01, SSA^{+}04, SMP05, SVC^{+}07, SMR^{+}01, TvG02, THWS00, Tho09, TKK02, VB00, VMS^{+}03, WLRT06, WS06, WBP^{+}06, WSBL06, YT09, ZW06, Zim03, vNHM^{+}09]. distributions [BGC04a, BST^{+}08, CSM^{+}08, CBK^{+}08, DG06, EWE02, FSM^{+}07, GZL02, HC01, HTAS06, KH03, MH04b, RHK^{+}09, SBB00, SMA00, ZJC^{+}07, vHRB^{+}09]. disturb [MHLN01]. Disturbance [FDP^{+}08, BS02b, CMD^{+}09, CCKC05, EBS07, HW09, MW02, WA01]. disturbances [HZ00]. disturbed [KCCE06]. Ditylum [RA00, RNA06].

Diurnal [DLZ^{+}09, DGL^{+}04, KDRR09, ORH01]. divergence [HWA03]. divergences [SMS07]. diverse [AVA^{+}02, LTB09, QRF06, SMA00]. diversion [DFBA07]. Diversity [MG08, AH02, CMH^{+}08, CMC^{+}09, FQC^{+}07, FPZ09, GLP^{+}08, GBL^{+}04, HGBB08, HZ00, LLvW^{+}03, LFS^{+}04, NTDV06, OBB^{+}04, RB02, RL07, RA00, Thi00, vdSvdSG^{+}00, Gou01]. division [MC00]. DMS [SWH^{+}03, TKK^{+}03, dVKT^{+}09]. DMS-producing [SWH^{+}03]. DMSO [dVKJK07]. DMSP [BCM01, EKD^{+}07, KL00, MKK04, MKVK05, SWS^{+}03]. DNA [CDD07a, CDD07b, DSD02, LTT07, RA00, VPA02, WGKE04]. DNRA [GMA^{+}06]. Do [BP03, BEM^{+}02, Fly03, HD03b, JKLR08, KNC^{+}09, KKH04, WDL07, WKB^{+}04, NA09, TDT01, WGN^{+}08]. DOC [All00, AG03b, HTS^{+}03, RB01a, YSM^{+}03, dGvD07]. docosapentaenoic [PWP07]. document [CCM07]. documentation [JKS04]. Documenting [DCC^{+}01]. Does [JJS^{+}03, KHNV05, MWE^{+}03, MTV^{+}02, PCC^{+}07, RC07, SSAH05, GM03a, GRVF^{+}00, Sak00, VW01]. DOM [BFC05, KS04a, MH00, MRNT^{+}07, ORH01, WR00, YJMT08, dGMH^{+}08]. DOM-feeding [dGMH^{+}08]. dome [SRM05]. Dominance [BOC01, RSWH04, BHMS06, ERH06, WA09]. dominant [CLC02, DVV^{+}07, GHS05, GGIA06, NDT07]. dominate [HZG07]. dominated [CL02, Cha07, DD09b, DT00, DRSL07, DBWS06, EMPI07, FAM04, HMS^{+}07, HBH^{+}06, JCW08, MGL02, WC04]. dominates [KB05a]. Domoic [TAB^{+}00, WTC^{+}05, BKK^{+}06, MHRW02, THH02, THL^{+}09]. DON [Ber01, BW99, PWH^{+}04, SSS02, SLJ00]. DON/TDN [PWH^{+}04]. don’t [KKH04]. DOP [CSSNA02]. dormant [DLDE04, DPK^{+}00]. Doughnut [KBG^{+}08]. Dover [Jun01]. down [CCS06, CBD^{+}06, DHTH05, Goe02b, KMHE04, LWS01, MMO06, SCDR07, SMO07, WPW^{+}09]. downside [MK00]. downward [SSH^{+}09]. Downwelling [DDL^{+}09, BTKGA07, SGG^{+}08]. drag [MKM06]. drainage [AB07, SKR01]. draining [HA01]. Drake [BTB^{+}08, HMR^{+}07]. drawdown [PLW^{+}08]. Dreissena [BFC05, BFC07, BDNF01, MGH08, MBDL01, SWH^{+}05, WBvE02, WvE04]. drift [HR03]. drifters [BMH^{+}08]. driven [FLPL08, GZL02, GA00, HLM08, LKAB04, PFH^{+}03, PTKP00, PH03, PFPH04, RKK09, REL06, RLJdB08, SP09, SL02]. drivers [SK05]. drives [BFS^{+}04, HRGM07]. driving [CSM^{+}08]. droebachiensis [GWH06]. Dronning [WSNH06]. drop [OFKH01]. drought [FHDKR03]. drought-induced [FHDKR03]. Dry [DP01, GMRD04, TPM01, Vin00, MSBHG01]. drying [CH05a]. dryland [HBTM05]. dual [WKFP06]. Duarte [Fou02]. due [MH07]. duebeni [DHKH05]. Dumai [AJS07]. dune [HB05]. Dungeness [PDS07]. Duplin [WC04]. duration [LA09, MD09]. during [AGL03, AS06, AMS^{+}08, AELS04, AH02, BD04, BVWB07, BJG^{+}00, BLSQ07, BSSL03, BJD05, BAP^{+}05, CJ08, Coc01, CPCK00, CCCC07, CCC09, ET08, GRdL^{+}08, GP06a, GM04a, GRB02, GSW03, GSNH04, GSLT08, HYH04, Har02, HFC^{+}04, HM06a, JBM^{+}07, JJS^{+}03, JGSB07, JHGGP06, KKKT07, KTLS03, KB06, Leo08, LRYS06, LKM03, LP07, LNJL09, LRF07, MSGK06, MBL^{+}01, MPACG01, MAC^{+}03, MSZ00, MLSS05, OKO^{+}01, OBSD04, PTK^{+}07, PWP07, PSH08, PVD^{+}07, PPL^{+}02, PD05, RSC04, RCVK09, SKTD04, STHG07, SLJ00, SLC^{+}09, SGG^{+}08, SMG06, SLS^{+}02, TvG02, TMS^{+}00, TBF04, VMH^{+}01, VMS^{+}04, WMRK06, WAGH09, WW03, WHC^{+}06, WTM09, ZJ03, dSYL^{+}08]. dust [HCK02, MLRMR09, RORR^{+}09, SWF04, TSSWK^{+}06]. Dutch [DGV01a, DGV01b]. Dynamic [YCM^{+}01, BKB02, LKB04, vdPVB07]. dynamical [KSS^{+}07]. Dynamics [FWM^{+}05, HBH^{+}06, MVBM04, RB02, SHA08, AD05b, Ano08b, AP00c, BJW^{+}02, Boo00, CSM^{+}08, CNGV09, lLCChTO08, CBH08, CDD07b, CSG^{+}06, CIWH09, CFG07, DGH^{+}07, DAC^{+}07, DJB^{+}07, DAMM^{+}09, DBGR09, Duf09, EWE02, EGPA02, FL08, FMD03, FG06, GMA^{+}06, GA00, GD00a, GGRJ03, GM04b, GBM^{+}04, GAPR06, HNB^{+}07, HSB00, HFR08, HH02, HTWG08, JS03a, JPJS06, JBvHS08, KRZ06, Kiø07, LAG^{+}03, LFS^{+}04, LMB05, LSC^{+}08, LLDS06, LWB^{+}01, MFK^{+}09, MLC^{+}07, MFF07, MLPS09, New01, NCO02, ORH01, PFPH04, RKNA09, RBW09, SSLE02, SSH^{+}07, SDM^{+}09, SKJ^{+}07, SAPA^{+}02, SNFT07, SSV^{+}04, SSJ07, SC08, SLC^{+}09, TCL^{+}09, TVSS01, THH02, TMH^{+}06, VKvD05, VCKS08, Vin00, WGLT05, WG04, WMW^{+}08, WSEL09, WNP04, YJMT08, dVKT^{+}09].

Early [CDD07a, SMD^{+}04, HYH04, HAES02, LWS01, PTK^{+}07, PWP07, REL06, SvKM01, SLS^{+}02, TMH^{+}06]. earth [Gai00]. East [FSS^{+}00, MRK02, SLS^{+}02, VBH^{+}05, WSNH06, EF05, CWJ09, HWL^{+}00, MVD05, Pil04]. Eastern [BLM^{+}06, ÁÁS07, BZS07, Coc01, DPC^{+}08, FCGC^{+}09, GST^{+}01, HB08, KN07, LBW^{+}08, LMBP07, MC00, MOK^{+}03, MMPA04, MWH01, RvAV^{+}06, SRF^{+}01, SKBC01, StLHB^{+}07, SH05a, TM02, VDK04, KHM04, TG09]. Echinodermata [MSBHG01, VTR03]. echinulata [VMS09]. Eckernförde [SSA^{+}04, TKBJ05]. Ecklonia [WLP^{+}07]. Ecological [BRE02, Boo00, HD09, Par01, QMS^{+}07, RCL^{+}02, SJR^{+}04, SFB^{+}09, BSSL05, BBCW09, BBKC03, BCBS04, CG08, CGWK^{+}09, HVH00, HA08, KAM04, LW02, OS01b, PvE00, PH03, RLM^{+}08, RES^{+}08, SMP^{+}01, Tor00, TWC^{+}09, WLRT06, WSKL06]. Ecology [WGC00, AWH07, BB04b, Bri01, CR08, Fis00, HR07, Ham01, JGZ^{+}08, Jun01, Kem00, KW04, KBL08, Kne02, LU08, MAMC02, OSDB07, Pom01, Smi00, WNP04, Wro01, Fou02, SG02]. ecophysiological [AKE04, PCDB09]. Ecophysiology [BH07, RCE^{+}02, WLP^{+}07, WSBL06]. ecoregion [Fis00]. Ecosystem [HT03, Vin00, ÁSGMP01, BK07, CC04, CNGV09, DD09a, EPP^{+}08, GLP^{+}08, GHS05, GHB^{+}05, GBM^{+}04, HBG00, HP06, JCC02, KHB^{+}00, LAJ^{+}04, LRMV04, LVSW07, LLDS06, MCF^{+}04, MdG08, MCL06, Nor04, RNC03, RKNA09, Ris02, RlBW02, SCS^{+}08, SCDR07, SFC^{+}00, TCCG05, WGN^{+}08, WNL^{+}08, WGLT05, YYTH07]. ecosystem-specific [SCS^{+}08]. ecosystems [BMU04, Cal01, CBRC01, CDD07a, GBA00, HTS^{+}09, HPV^{+}00, HA08, HM06b, LN02, MPJ08, MVA^{+}09, Mul01, MHS^{+}09, PVP^{+}06, PSS^{+}06, RH06, SJH06, TVSS01, WBB09, ZHL06, vdHPH^{+}09, Dow02]. ecotone [SMD^{+}04]. ecotype [ZJC^{+}07]. ecotypes [SFR^{+}08]. ectoenzyme [FSS^{+}00]. ectohydrolases [BA01]. Ed [Pom01]. eddies [SLKP05]. Eddy [BRW07, BMWW08, THL^{+}09]. Edge [BBCW09, BGR^{+}09, STG^{+}07]. Edmondson [Leh00a]. Eds [Caf01, Gou01, Mil02, SG02, Sta01, Wan02, Wro01, Dow02]. edule [RHM04]. eelgrass [FHBK06, GBP03, Reu02]. EEM [YJMT08]. EEM-PARAFAC [YJMT08]. Effect [AKH04, BGC^{+}05, BM03, BR00, CWG^{+}07, CCC^{+}03, GR01, GBFJ07, GM03b, GCG05, GMRVFP02, HHB05, HK01, IJA^{+}04, JVL00, JTL01, KS04a, KDO09, KTMM00, LASR09, LH05, LBC^{+}04, MBC^{+}05, MF01, MF02, PL09, QMS^{+}07, RORR^{+}09, RRS06, SH00, SH05b, SBM04, TXH05, TGFPT08, TTG^{+}09, TLrR07, VTR03, WS06, WSBL06, ZHSJ09, AS00a, AOR07, Bat02, Ber01, BZ03, BCDH09, BMNSK00, Cal01, CCKC05, CBG01, CRF01, DKA^{+}03, EZA^{+}05, FRO04, GRV^{+}09, GAPR06, HDSC09, HBS05, HMNL06, KCE^{+}00, KLS^{+}06, KLS^{+}07, KHK^{+}03, KHQ01, KDRR09, LLS08, LEML09, MHRW02, MB05, MS05, NA06, OS01a, OFKH01, PS06, PFL07, RB09, RSS^{+}03, SE06, SHSH^{+}07, SDvEB06, SA09, SBR^{+}06, ST03, TCR03, TR03, UMASG07, WSL^{+}02, YYTH07]. effective [HH03]. Effects [ARO^{+}04, ACHG07, ABDS04, AKP^{+}09, BST^{+}05a, BBJ05, BS02b, CWPL05, CG01, CS05, DGV01b, DDZ03, DA05, EDF^{+}03, Fab06, FAC05, FAL^{+}07, FHDKR03, FQH^{+}01, FS09b, FGM^{+}03, FNM07, GPAP^{+}04, GEP08, GVT06, GSL^{+}07, GSJ^{+}09, HPV^{+}00, HFR08, HG03a, HG03b, HH05, KCSM07, LFZ^{+}07, LCC^{+}09, LH06, MPA^{+}07, MSLH02, MKM06, MPD^{+}06, MS03, MZV03, MPZ08, NPE08, NH06b, PMK08, PSHM04, PBGV04, PD05, RCVK09, SCB03, Sal05, SWCL01, SPM04, SCP08, SCM08, SWN08, SDH02, SL01, SLRM03, SSC02, SCV04, TYK^{+}08, TSSWK^{+}06, VKvD05, WG04, WWMB05, WST06, WBR03b, WL06, WWS^{+}00, ZB00, ALRM08, Ann05, Ano08a, AMS^{+}08, AY08, AH01b, BVH02, BB03, BB05, BTKGA07, BBAW09, BRM^{+}07, BSG^{+}08, BLSQ07, BBG^{+}05b, BvE07, BDNF01, BS01, CLL^{+}03, CFMH01, CVH^{+}05, CD02, CSK^{+}07, CT02]. effects [DD09a, DAMM^{+}09, DSW05, EMPI07, EMK08, FS09a, FH09, FSK^{+}04, GRB09, GGH07, GGS06, GNU00, HGB00, HH03, HS04, HFR09, HWL^{+}00, JWRW07, JS08, JM02, KNC^{+}09, KBL08, KBD^{+}00, KSM^{+}05, LTM01, LCS^{+}03, LBEP06, LFA^{+}09, Les01, LLB04, LL05, MBK^{+}09, MSVS^{+}07, MHKP02, MPB^{+}08, MY05, OBB^{+}04, PLMPN08, QLD^{+}02, RKM^{+}06, RNR08, RCD^{+}05, RPR^{+}02, SSvNH03, Sch09, STC03, SJH03, TG09, TDT01, WBvE02, WA09, WMWV04, WDVA08, WH07, vdPAJ^{+}06]. Efficacy [GJRM07]. efficiencies [GP00, HK06, Too01, XTSM07]. Efficiency [Aki00, BMBA07, NH06b, ASGA^{+}07, BBL^{+}03, CEBMF05, KTM01, MRNT^{+}07, MNM^{+}09, OHUB00, PBGV04, SP04, SDKZ^{+}09, SA04, TCD^{+}03, UPC^{+}05]. effluent [SLE04]. efflux [CBvEK03, KHP^{+}06]. effusus [KLSW00]. egesta [PW01]. egg [AFL07, GRCS06, Has04, IHH^{+}00, IHHH00, JS03a, JKS04, KS04b, SK04, WvE04]. egg-bank [JS03a]. egg-production [Has04]. eggs [BDvO^{+}03, BSSL05, DPK^{+}00, GS04, KSEÖ^{+}02, YM03]. Ehrenb [KHH^{+}03]. Eighty [LD01]. Eilat [BTKGA07]. eject [OH08]. Elbe [DBE08]. Electron [MED^{+}08, PTKP00]. electrophoresis [DMEPA04]. electrospray [SHL^{+}05]. electrospray-ionization [SHL^{+}05]. Element [CG08, TBF04, SDvEB06, WA01, YSM^{+}03]. Elemental [BBKC03, GHN^{+}02, HSN^{+}03, BFML05, DL00, FQR^{+}06, FH09, GR01, GGP01, HABF03, HPHT00, HFZ^{+}07, LG04, MH07, Pri05, SL01, VC05, WSC^{+}05, WSKL06]. Elements [Thi00, AVJ^{+}00, BFS02, CLL^{+}03]. elephant [BTB^{+}08]. Elevated [LPD^{+}04, SNPK^{+}09, CMCB^{+}09, GCG05, MAC^{+}03, SWN08, TG09]. elevation [NSM09, SMC03]. elevational [TSMB05]. Elizabeth [CB04]. Elkhorn [BNM^{+}09]. elodes [GGH07]. Elopidae [MH04b]. Elops [MH04b]. Elucidating [FMD03, GVT^{+}09]. embayment [EM02, MRK02, RC05a, RC05b]. embedded [WSHN00]. embryonic [JHGGP06]. embryos [KBD^{+}00, LLB04]. emergence [GS04, IOG^{+}04, SJ08, TVT04, VKV07]. emergent [KLSW00, OS01a]. emersion [HBH^{+}06]. emigration [PP00]. Emiliania [BTdB03, DNB^{+}04, DA05, EZA^{+}05, EMWL06, EKD^{+}07, EW08, FHB^{+}09, GRV^{+}09, GD01b, LGN^{+}06, Leo08, LMT05, MAL08, RZR02, SXLM06, SWH^{+}03, TSG^{+}01, TLrR07, XTSM07, vdPVB07]. emission [YJMT08]. Empirical [BMNSK00, KTAL04, BSJ02, JPJS06, MBC01, WSC^{+}08]. Emulsions [VMJ08]. enantiomeric [PPH03, TvAVH06]. enantiomerization [PTLJ01]. encounter [HK06, KB05b, KKT05]. encystment [KRM09]. endemic [GCM^{+}09]. endopeptidases [OS05]. endosymbionts [RSS^{+}03, ZB03b]. Energetics [AI01, SI08, ACHG07]. Energy [RG07, ACW02, AMS^{+}02, DPFL09, Jan03, KB05a, LFA^{+}09, MAS^{+}02, RS09, SO07, SIH^{+}06, WBR^{+}03a, WPV00]. energy-budget [WBR^{+}03a]. engineering [DD09a]. engineers [SFC^{+}00]. England [GSO03, TCP^{+}03]. English [FUG02, IHH^{+}00, IHHH00, LMBP07, RV04]. enhance [Has04]. Enhanced [ACG07, TCD^{+}03, WAGH09, Fra01, FT03, HSF^{+}08, LBC^{+}09, MV03, dSYL^{+}08]. Enhancement [AJ01, FA03, CPL^{+}04, GB09]. enhances [EMIP05, MSAC06]. enrich [vE02]. enriched [BRP^{+}06, BLM^{+}06, CBRC01, NH06a, dGMH^{+}08, dSYL^{+}08]. Enrichment [KLAF04, AS00a, AS06, AWH07, BHMS06, CV05, Coc01, CJWR05, HR07, HPV^{+}00, LSK^{+}03, LLF^{+}09, MBL^{+}01, OSDB07, RH01, SSK06, SMC03, Smi06, TMK^{+}07, TMA^{+}01, WA01, YS06]. enrichments [CLP^{+}03, GSH04, LCPS05]. enslave [MZL04]. Entrainment [STK^{+}04, FLT02a]. entrance [KHH^{+}03]. enumeration [MKK04]. Environment [BFS^{+}04, BN01, BMD^{+}04, CFMH01, DFBA07, DVV^{+}07, DC05, DRSL07, Dup00, DST08, EKdF07, FF04, FBW^{+}08, GSB00, Hei08, Hes00, JTAM09, KAW05, KBD^{+}00, Les01, MPGB08, MB00, NH06b, RLP^{+}06, SRNW06, SŻKL08]. environment-indicating [DST08]. Environmental [SSLE02, TKBJ05, AS00b, AHHH^{+}07, BZS07, BBLB05b, BBLB05a, CSM^{+}08, CCKC05, Dav00, DHKH05, EPP^{+}08, Gai00, HAHS^{+}09, LALE04, LCC^{+}09, MSH^{+}06, RM06, RBW09, RMG^{+}04, SGPK07]. environments [DLZ^{+}09, HGBB08, LRS^{+}03, MNM^{+}09, OH08, OH00, SLRM03, SMA00, TW09, TCL^{+}09, YJMT08]. enzymatic [CS06]. enzyme [SHA08]. enzymes [OS05, STC03]. ephemeral [CIWH09]. ephippial [KBE^{+}04]. epibiont [BAPMMB01]. epibionts [DLDE04]. epidemics [DHTH05]. epifluorescence [SF05]. Epilimnetic [HBM06, KFR^{+}01]. epilimnion [HHS05, MBDL01, PBC02]. epilimnion/hypolimnion [MBDL01]. epilimnion/hypolimnion-related [MBDL01]. Epilithic [HHT01, LZ09]. epilithon [ML05]. epipelagic [SRZ04]. epiphyte [DDZ03]. epiphytes [CT02]. epiphytic [TCC06]. Episodic [HLM08, LSM03, SCM08, GM03b, NSBA01]. Eppley [BMSA08]. EPS [HBH^{+}06, LB00]. Equatorial [TM02, Alt01, BDK^{+}08, Coc01, MC00, vdSvdSG^{+}00]. Erie [ALH01, BLHC08, HBG00, HBS04, NGS^{+}07, RHCS08]. erosion [EMPI07, SFC^{+}00]. Erratum [Ano07, Ano08a, Ano08b, CR10, DdG09a, KB08a, KLS^{+}07, MBH^{+}06a, SS04]. error [Lan03]. errors [McA03]. erythraeum [PM08]. Escape [BHL07]. Escherichia [GKS^{+}02]. Essential [KAM04, AP00b, PWP07, RB06]. Establishment [LED06, RB03]. esterase [RvBB02]. estimate [ACD10, MPdVG07, SH10]. estimated [BZH08, KM02, LHWE01, Lee01]. estimates [BNC04, BLS^{+}06, CRJ08, DB07, EWE02, IJA^{+}04, MSLH02, MV03, Too01]. Estimating [CDBC00, LMSE05, WPV00]. Estimation [HR03, SSMM00, Har01, OHUB00]. estuaries [BDS^{+}04, BdG02, CHH07, CBD^{+}06, DBE08, DNS06, FHBW06, GMA^{+}06, HOFD01, JCW08, LTOS04, McM00, OCNH00, SMP^{+}01, THA01, TLM04]. Estuarine [Bia01, Dow02, ACG07, AJ01, BN01, BKB02, CG09, CV05, CK08, CCH02, DAC^{+}07, DVV^{+}07, DGD04, GSW03, HBH^{+}06, JTAM09, JCBA08, KLG09, MYM^{+}05, MH00, MH04a, MVFD04, MBCD04, MBC06, MA08, MGF03, MSZ00, Nil03, PVP^{+}06, Par01, RBW09, SDC00, SBL06, SSS02, SWS^{+}06, SMJ^{+}06, Smi06, YTVR07, Zha00, ZWC^{+}04, dGB02, CWWS00, VWPP^{+}06]. Estuary [FW07, GGDC04, How00, MWM^{+}04, TRF^{+}06, AJS07, AHM^{+}08, AR06, BKM05, BDSB07, BLBB07, BDK^{+}06b, CL02, CBA01, Cha07, CGL03, CG03, CD03, DBE08, DAC^{+}07, EF05, GSG^{+}05, HMS^{+}07, HDWDD09, HDVB01, KM02, Kim05, LF09, MA08, MPZ08, NA03, RNC03, RBC00, RB01a, REL06, RL07, SBC^{+}08, SMD^{+}04, SCM08, SMH^{+}06, SM05a, TWS03, Tho09, THA01, TCP^{+}03, TM05, VVSM09, WC04, YNCK04, YS06, KCSM07, KDMC05, Leh00b, LBG^{+}09, SMJ^{+}06, WKFP06]. ether [WIK08]. eucaryotic [CMB00]. Eudiaptomus [ZRL06]. Eugene [Dow02]. eukaryotic [DMEPA04, DR06, SKM05, SFR^{+}08, TSG^{+}01, WNP04]. Euphausia [ASAC04, AW00a, AW00b, AMS^{+}02, ASH^{+}06, JHGGP06, MAS^{+}02, MFG^{+}09, RQB^{+}00, SAP^{+}06, SHS03a, TSH^{+}06, TKM07]. euphausiids [GYBA04, JHGGP06]. euphotic [AO09, BK03b, BW99, LKAB04, NEE08, TCB^{+}00]. EURO90 [Wan02].

Europe [DJG^{+}06]. European [BDS^{+}04, JLB^{+}06, JGE^{+}09, Law08, OAA^{+}06, PLG^{+}02, UWA^{+}07, VMKN03, WJA^{+}07]. eurystoma [SE06]. eutrophic [BOC01, BDK^{+}06b, CCC09, GCG^{+}02, GFM^{+}01, GCS06, GP06b, HLNM^{+}03, HTWG08, JPJS06, KFW09, Law08, LBM^{+}04, MCF^{+}04, MGVM^{+}06, PWS09, RA09, RB02, SKJ^{+}07, Zha00]. Eutrophication [Dod06, MSVS^{+}07, SJH06, SQTA07, CCS06, CCAÆ06, CHH07, CVH^{+}05, CS06, CWC^{+}06, FHBW06, GCJ^{+}06, GP06a, HOFD01, HM06b, HLK05, IR06, LFS^{+}02, PVP^{+}06, PD02, PDTR06, Sch06b, THT^{+}09, VJV^{+}03, VM06, WSBL06, WL06, YYTH07, ZC02, Elm02]. Eutrophication-induced [SQTA07]. evaluate [HBI01, MBC01]. Evaluating [CCKC05, MC06, RENA06, SJW03, CBA01]. Evaluation [BCM01, LTGG03, LBSY02, HHB^{+}09, WSC^{+}08]. Evaporation [WBR^{+}03a]. evasion [HPBD01, RBC00]. event [AGL03, BZH00, LJ09, LRYS06, RRHPT^{+}06]. events [AR06, CTR^{+}02, GWH06, GV09, HLK^{+}07, HAHS^{+}09, NSBA01, PPL^{+}02, SCM08, SVPK^{+}09]. Everglades [Ris02, CBD^{+}06, GCJ^{+}06, IR06, NHS07, SIRO05]. Evidence [AP00a, Aks07, ÁÁS07, BAG^{+}09, BAPMMB01, BZ03, BNM^{+}09, CTY06, CLS05, DGV01a, DLDE04, Dit04, DAA^{+}01, ED05, FAC05, FL06, FRO04, GKS^{+}02, HMS^{+}04, HDFM08, KRZ06, LAG^{+}03, LvdB09, LA06, ME06, NGS^{+}07, PDS07, PLW^{+}08, RHM04, SMC03, SQTA07, TSMB05, THJ^{+}09, TGF00, VMH^{+}01, WKW09, Zha00, ZWC^{+}04, dGMH^{+}08, ARB^{+}04, Alt01, AE08, BBCW09, DPK^{+}00, GLVFH09, Had08, HVP^{+}07, HR00b, HPHT00, KTG^{+}06, LNB04, LFA^{+}09, LM02, MW02, MCJ^{+}06, MVFD04, PC04, PD02, PTKP00, PCD^{+}08, RCL^{+}02, Sch00, WRG^{+}05, RC07]. Evolution [DHP03, AIS00, AI03, BST^{+}05b, CRY^{+}08, GPHS08, TIH02, VRC07, dlFSIN08]. Evolutionary [SBS07, Tor00, HA08]. Evolving [HM06b]. ex [DDZ03]. exacerbates [GRV^{+}09]. Examination [CL02, MW06a, SIRO05]. examined [TMYH02]. Example [DAC^{+}07, CTCT03, Leo08]. exceeds [LH06]. Exceptional [HAHS^{+}09]. Excess [HA08, SL02]. excessive [vdPAJ^{+}06, vdPVB07]. Exchange [MHC05, RC05a, BSM01, Bat02, BMD^{+}04, CGL^{+}04, CWG^{+}07, FUG02, GJRM07, HM08, HLM08, HDP09, MCJ^{+}06, OV03, PH03, PFPH04, RSC04, RHJ05, THA01, UL05, VPH^{+}07]. exchanges [SH06]. excitation [SIK07, YJMT08]. exclusion [BGV02, DGV01a, HZ00, PL09]. excreted [HY01, SH03]. excretion [AW00a, AW00b, AH01b, MTC^{+}06, ZHL06]. excursions [KSM06, TB05]. exhibit [BP03]. existence [TF00]. exopeptidases [OS05]. exopolymer [ASP04, LB00, PP07]. exopolymeric [MBLP01]. Exotic [HBG00, BKB02, CPB^{+}08, DPK^{+}00, MGH08, NPE08, RSWH04, RL07, WM06]. Expansion [SCY00, JGE^{+}09]. expeditions [Smi00]. Experiment [BJD05, BAP^{+}05, KTLS03, OBSD04, TBF04, BMD06, BHV04, Coc01, EGP08, GCG^{+}02, HTS^{+}03, HPAM01, KLS^{+}06, KLS^{+}07, LTT07, MJW03, PBMSG04, PW03, TCD^{+}03, VMS^{+}04, vOMSM06a]. Experimental [BVH02, DAV^{+}05, KTG^{+}06, KNC^{+}09, KW04, MW02, RRH^{+}03, SPM04, SF03b, SF03c, AD05a, BGR^{+}09, BPLH04, CSWD07, ED05, HPV^{+}00, HD09, JPK^{+}05, LW01a, LNB04, LED06, MSAC06, MW06a, MSPAE06, OAA^{+}06, PSHM04, RL00, SCDR07, SFC^{+}00, SCV04, SHS03a, VMS^{+}04, WMRK06, FKS^{+}01]. Experiments [EMK08, SSS08, ZMJ^{+}07, AWH07, BLM^{+}06, EVW07, EZA^{+}05, HR07, HTS^{+}09, HR00b, KCP^{+}04, MT06, MHS^{+}09, NHN^{+}05, OS01a, OSDB07, TKAW03, WST06, BVG^{+}07]. explain [KR08, MTV^{+}02, TF00]. Explaining [MPH08]. explains [SPW^{+}05]. explanation [BWK^{+}07, Fra01, TA03]. explicit [Ano08a, MCPP06a, MPB^{+}08]. Exploring [WGN^{+}08]. export [ÁS07, BAP^{+}05, BB09, Ceb02, DBTK02, ET08, GSL^{+}07, GSJ^{+}09, HBAK09, HD03a, MWE^{+}03, Pil04, SSH^{+}09, TVSS01, SVPK^{+}09, WC04, WTM09]. exports [FL06, KCV06]. exposed [ABFS09, AAD08, BVH02, JJT07, OAA^{+}06]. exposure [EMWL06, GCG05, HH03, HD03b, LW01b, SDKZ^{+}09, WLP^{+}07, ZSS^{+}08]. Expression [LPAQ^{+}05, MPZ08, PP00, ZMJ^{+}07]. extend [VPH^{+}07]. extended [AP00b]. extends [HGBB08]. extensive [RA00]. extent [LFS^{+}02]. External [LFAB06, ESA06, JPJS06, SKR01]. Extracellular [KSM^{+}05, CDD07a, CDD07b, CMB00, DSD02, EPS00, HBH^{+}06, MT06, TSKA05, VOG^{+}08]. extraction [AD00b]. extraordinary [HGBB08]. extratropical [PLW^{+}08]. Extreme [BL07, DMS^{+}03, GV09, GSH04]. extremes [GDCG08]. exudate [LV01]. exudates [LMZB08]. exudation [DA05]. exuded [DNB^{+}04].

faceted [CLL^{+}03]. facilitate [SPZ03]. facilitated [GKS^{+}02]. facilitates [ABFS09, NA05]. facilitation [ARV^{+}07]. factor [CPCLS04, GWR^{+}09, MRNT^{+}07, YJMT08, ZB01]. Factors [GCS06, LFS^{+}02, LVC^{+}00, MJEM06, SNM^{+}08, TB05, BSM01, BFCLÁ08, EF05, GCG^{+}02, HVH00, HDWDD09, KRZ06, KSS^{+}07, LALE04, MJ03, MBCB08, MJW03, RBW09, SGPK07, SA09, WW02]. facultative [vNHM^{+}09]. fairy [HMD^{+}07]. Falkenbergia [MHDH00]. fall [AMS^{+}08, GJT^{+}08, GBM^{+}04, SR07]. fallout [NSBA01]. farcta [WSL^{+}02]. Fast [RSY04, RP06, CEBMF05, DJB^{+}07, GFK00, Lan03, SKH^{+}01, SQTA07]. fat [Has06]. Fate [VvOBM06, BMU04, CZRD^{+}04, DSD02, Duf09, Hes00, KGYU06, MPJ08, MBB^{+}00, SGG^{+}08, TMA^{+}01, WNL^{+}08, vOMSM06a]. fatty [AP00b, BB05, BMNB^{+}06, KAM04, KK00, PWP07, PAS^{+}05, RBMN03, RB06, SO07, SAP^{+}06, TGFPT08, TTG^{+}09, VVSM09, VW01, WBvE02, WvE04, vE02]. fauna [PWP07, VEMM07, WG04]. faunal [GGRJ03]. faveolata [LMPY00, QYEA06, RRHPT^{+}06, SSL05, ZJ03]. Fe [ALRM08, BRT^{+}03, BFQF^{+}07, CT05, FMG^{+}08, MP00, MHRW02, MKS08, RW02, SWCL01, SH00, WKSM08]. Fe-limited [MP00]. feature [SRM05]. features [BSG^{+}08, FSSV08, GRCS06, GS04, MS06, PE07, SGLT08]. fecal [PIF08, PIKB08, RFM07]. fed [BSS^{+}01, DMJ09, HLD^{+}03]. feed [KSEÖ^{+}02]. feedback [ME06]. feeder [BB04a, PW01, SM02]. feeders [CMQ^{+}00, Com02, WLTF03]. Feeding [AMS^{+}02, IHH^{+}00, MAS^{+}02, SWH^{+}03, TJV01, BB04b, BDC00, GPAP^{+}04, GM04a, GMRVFP02, HK06, HW09, JS04, KNE08, LH06, MK00, MIDH05, MA00, Møl07, MTM^{+}01, NHTF01, NH06b, PSH08, RB08, RL00, RL01b, RHM04, SC07, SGPK07, SSAZ01, SLD00, SdB06, SWS^{+}03, THR00, TKM07, TS01, TGFPT08, TTG^{+}09, WSNM00, YSM^{+}03, YWR^{+}07, dGMH^{+}08]. Feitsui [CWLH06]. FeL [SRNW06]. female [HFC^{+}04]. females [YM03]. ferric [SRNW06]. ferrichrome [KTMM00]. ferrisiderophore [KTMM00]. fertilization [AWLH04, FF04, GWH06, GRG02, HPL^{+}06, LDC^{+}01, OFKH01]. fertilized [GS04, YM03]. few [Kiø07]. fiddler [KA06]. Field [ARB^{+}04, CMD^{+}09, DGL^{+}04, HWL^{+}00, Ano04, AIS00, BM00, DA03, Fra01, Har01, JLH01, KSWC^{+}03, Leo08, McA03, MVW^{+}04, RQB^{+}00, RA00, VB00, VCCS05]. field-compatible [Har01]. field-flow [VB00]. fields [Dix00, KCCE06, ZZPP01]. Fifteen [SMA^{+}06]. filamentous [BD07, NKE08, RSY04, STHG07]. filaments [ÁS07, DGV01a]. filiformis [VTR03]. filter [BPC06, BD07, MTM^{+}01, RL00, RL01b]. filter-feeding [MTM^{+}01, RL00]. filter-screen [BD07]. filtration [GDL^{+}00, JPK^{+}05]. finding [KTM01, Kiø07]. Fine [UMY^{+}02, WdBD^{+}01, AHR^{+}01, CSM^{+}08, DJH05, ETCP03, FT03, HA01, NTM^{+}05, Tun00]. fine-particle [FT03]. Fine-scale [UMY^{+}02, WdBD^{+}01, CSM^{+}08, DJH05, Tun00]. fingerprinting [DL00, RA00]. Fingerprints [LALE04, BFML05, MWH01]. fingers [KLF^{+}03]. Finland [CSPW00, GL07, KHH^{+}03]. finmarchicus [Has06, KBD^{+}00, ODR^{+}08, SvKM01]. Finnish [SJR^{+}04]. Fiordland [WLP^{+}07]. first [BJG^{+}00]. first-year [BJG^{+}00]. Fish [CA08, SPZ03, SH10, ANSS04, Aks07, AH01b, BSSL05, BBLB05b, BBLB05a, BJL01, CBG01, CCM07, Dav00, FSSV08, FB05, GHWG04, GCP^{+}02, HNB^{+}07, HCC^{+}02, HH02, HG03a, HG03b, HW09, KLWS08, LBC^{+}09, LKM03, MK00, MACB07, MD09, MM05, MAC^{+}03, MIDH05, MID^{+}07, NH06b, OKO^{+}01, PC04, PvE00, RHK^{+}09, Rit09, RH03, SK04, SK05, SD00, SRL^{+}09, SLC^{+}09, TMYH02, VNH^{+}08, WSC^{+}05, WSC^{+}06, WM06, Wur07, ZW06, ZHL06]. fisheries [HH02, Wro01, Wro01]. fishes [BTKGA07, BKB02, CCC^{+}01, DS07b, JB00, PVD^{+}07, SP09, SLKP05]. fishless [SCV04]. fitness [LH06, WKB^{+}04]. five [WM06]. fixation [ÁÁS07, Ano04, BFCS^{+}01, BFQF^{+}07, CGCPA08, lLCChTO08, CM03, FPZ09, FMG^{+}08, GEJ^{+}02, GMA^{+}06, GGH07, GZY^{+}08, HHT01, HVP^{+}07, HFZ^{+}07, HCD^{+}09, KFH^{+}09, KSWC^{+}03, LHWE01, MW06a, MCC02, MB05, MBH^{+}06a, MBH^{+}06b, NFS^{+}07, PLW^{+}08, PGL06, PM08, StLHB^{+}07, SH04, ZMJ^{+}07]. fixed [MBH^{+}06a, MBH^{+}06b]. fixing [ARV^{+}07, CPHM04, CBK^{+}08]. Fjord [Cro03, AELS04, CTWFP06, ESA06, MLSS05, PBM^{+}07, SLL^{+}04, Tun00, WGLT05, YWR^{+}07, BDS^{+}04]. fjords [WIK08]. flagellate [BWK^{+}07]. flagellates [CMJ07, SSAZ01, UMASG07, WNAW09]. flat [FAM04, GBM^{+}04, WBP^{+}06]. flats [RLJdB08]. flavicans [LLS08]. Flexible [WSKL06]. flip [GP03]. flip-side [GP03]. float [BST^{+}08, RNR08]. float-based [RNR08]. floats [MJR08]. flocculation [vWSBT08, vWBT09]. flood [AMS^{+}08, BS02b, SLJ00]. Flooding [LVSW07]. floodplain [TMUW02]. Floods [LAJ^{+}04]. floor [Bea03, WCL^{+}07]. Florida [CL02, DGCMK00, DCC^{+}01, Ris02, SSV^{+}04, TRF^{+}06, BHVC05, BHVC06, BSA^{+}00, BSL^{+}07, CCKC05, CBD^{+}06, CDBC00, DC05, GCJ^{+}06, GDB^{+}09, HRGM07, JNKdV09, LSM03, LWP^{+}07, LDC^{+}01, MCJ^{+}06, PF01a, Sch06a, SWML08, SLKP05, SP02, SSL05, ZSS^{+}08]. Flow [CEA^{+}09, RM09b, SPW^{+}05, AGKK08, DMS^{+}03, Fab06, FAC05, FRO01, Gay00, GRR^{+}07, GDL^{+}00, HMSI06, HZG07, HBTM05, HG03a, HG03b, JVL00, KBBJ00, LASR09, LAN08, LSF^{+}08, MKM06, PTKP00, PP00, RVP^{+}09, RG06, RMK^{+}06, RKK09, RLJdB08, SSAZ01, SABA02, SON^{+}03, TBHS01, VB00, WBD02, vDBL05, CBH08]. flow-induced [JVL00]. flowing [THJ^{+}09]. flowpath [FQH^{+}01]. flows [NA09, PF01c]. fluctuating [DBGR09]. Fluctuations [OKO^{+}01, LSK07, LIE06, RHJ04, Sal05]. fluid [BHVC05, BHVC06, JTL01, RL01b, Sch06a, TBHS01]. fluids [CH01, SM02]. Flume [FT03, Bea03]. fluorescence [DCC^{+}01, HFE07, LLA05, MF03, MSL^{+}03, Mor03, ORH01, OHUB00, SM05a, SM05b, YJMT08, ZJ03]. fluorescens [CG08]. Fluorescent [CKPS04, MLG^{+}03, NCÁSP06]. fluorometer [GFK00, SKH^{+}01]. fluorometry [CEBMF05, Lan03, RSY04, SQTA07]. fluorosensor [HAES02]. flushed [TCP^{+}03]. flushing [MCLM02, OS01a]. fluvial [ASS03, FAM^{+}06]. fluviatilis [LB04a]. Flux [GCP^{+}02, Lon00, All00, BNC04, BRW07, Bou04, BMH^{+}08, BMWW08, BB09, BLG^{+}07, CH01, DAL02, DS07a, FLWP09, LSPC02, LMSE05, MSLH02, MWE^{+}03, NHTF01, RFW08, RHJ02, SSH^{+}09, SKTD04, SR02, SR04, SlBWC^{+}00, SVB^{+}08, SBM04, VRS09, VTR03, VTO^{+}05]. Fluxes [BSS^{+}01, SKR01, ACD10, BNM^{+}09, CGD08, HKD01, HSB^{+}04, JFH^{+}05, JHW05, KCSM07, Law08, MTC^{+}06, MdG08, MWE^{+}04, NFFK00, Plo01, Plo08, RlBW01, SI01, SBC^{+}08, SH10, STG^{+}07, SWML08, SLL^{+}04, TVL^{+}00, TMA03, dBSN^{+}06]. flyash [BRB^{+}01]. Focusing [NDP^{+}07]. following [AD05b, GFM^{+}01, HGFS^{+}05, HLL^{+}06, MKG04]. Food [HB05, Law08, SH03, WBvE02, WvE04, ABJ^{+}06, AFH07, AY08, ASH^{+}06, AGKK08, BFS02, BFD^{+}04, BFC07, BSMR08, BMBA07, BL07, BSG^{+}08, BMNSK00, BvE07, CSWD07, CV05, CSK^{+}00, CPCK00, CGWK^{+}09, CCC^{+}01, CLS05, DGH^{+}07, DDL^{+}09, DGV01a, DET04, DPC^{+}08, EW07, FH09, HMSI06, HMS^{+}04, Ham05, HSS06, HSB00, HTS^{+}03, HW07a, HH05, JLH01, JSJ03, KAM04, KTG^{+}06, KDO09, KHKS05, KTM01, LFFC06, LFPH07, MB00, MBK^{+}09, MRM07, MBCB08, MGVM^{+}06, MSJMN02, NH06a, NKE08, OHCP02, OAA^{+}06, OHT06, PCC^{+}07, PSHM04, QMS^{+}07, RBMN03, SO07, SGPK07, SSvNH01, SAP^{+}06, SPM04, SBS07, SAPA^{+}02, SKBC01, SBR^{+}06, SRKK08, STST07, SS01, SS04, SHS03a, TRT03, TJV01, THJ^{+}09, VJ07, VR01, VMC^{+}07, VZM07, VKvD05, VW01, VPA02, WSL^{+}02, WWMB05, WH07, WLTF03, WLS00, ZHSJ09, vODL^{+}09, vE02]. food-finding [KTM01]. food-web [vODL^{+}09]. foodwebs [HL02]. Foraging [BL07, WM07, DG06, JS03b, JJP04]. foraminifer [GSFB02]. Foraminifera [AG03a, LEML09, NHN^{+}05, Spe00, vHRB^{+}09]. foraminiferans [HGBB08]. forced [NSH^{+}09, WPV00]. forces [CSM^{+}08, GDK03, HD03b, OD08, RKK09]. forcing [Cha07, ESA06, LLK05, LP07, MMKN08, PBMSG04, SSLE02, SLC^{+}09]. forecasting [CLF^{+}08]. forest [DTFL03, FKS^{+}01, GRR^{+}07, KJJ07, TYK^{+}08, TMS^{+}00]. forested [HBL02]. forests [GVT06]. form [BHMS06, RSS^{+}03, RES^{+}08, SHSH^{+}07]. formation [BdBS^{+}08, CR08, CPHM04, DNS06, KRM09, LU08, LV01, MMS^{+}02, NDT07, NEC03, SMS07, SSM^{+}09, VOG^{+}08, ZWC^{+}04]. formed [Plo08]. Forming [LL06, LCS02, PSJ00, RRBS03, Sar07]. forms [FAC05]. fornicata [MTC^{+}06]. fossil [BLS^{+}06, CPL^{+}04]. fostering [MID^{+}07]. found [TP03]. founder [FRO04]. four [RLM^{+}08, TvdWdB04]. Fourqurean [EW03]. fraction [MPAS00]. fractional [Alt01]. fractionated [CEBMF05, HWYL07]. fractionation [AHR^{+}01, BSC^{+}09, CCCC07, CCC09, GR01, GSNH04, GSLT08, GDL^{+}00, JGSB07, KTLS03, RZR02, SPW^{+}05, VB00, VR01, WGLT05]. fractions [SM05b]. Fracture [KCCE06]. Fragilariopsis [KMvD^{+}09]. Fragilidium [HHB05]. fragmentation [GYBA04]. framework [SCS^{+}08]. France [DAC^{+}07, MCL06, MWM^{+}04]. Francisco [CTCT03, CGL03, CG03, CGL^{+}04, CCH02, Leh00b, LRF07, LF09, MC06, NCC00, SH06, WKFP06, Kim05]. Franklin [WCB^{+}06]. Franks [SMS09]. free [BdG02, FSM^{+}07, HMJ04, LAN08, Lür03, MG08, MWH01, MSBHG01, RS01, THJ^{+}09, dGB02]. free-flowing [THJ^{+}09]. free-living [BdG02, FSM^{+}07, HMJ04, MG08, MWH01, dGB02]. free-surface [LAN08]. French [APH^{+}03]. Frequency [CC04, AWE07, BIIA03, GGIAY08, GBH^{+}02, LPB06]. Frequent [BCSW07]. fresh [HU07, dSYL^{+}08]. freshly [RFM07]. Freshwater [Cra07, AGL03, ARB^{+}04, Alv07, BSSL05, BK01a, BGE04, BSL^{+}07, BMNSK00, CT06, CWPL05, CF01, CMJ07, CLS05, DMJ09, DRSL07, FRO01, FRO04, GD00a, GJRM07, GHB^{+}05, GVT^{+}09, HZG07, HS04, HSLV05, HBS05, HBN^{+}00, KAW05, KFW09, LFS^{+}02, LBEP06, MBK^{+}09, MAC^{+}03, MGW^{+}05, MSJMN02, NPE08, NRB08, NA03, NA06, PH07, RVP^{+}09, RRH^{+}03, RH06, ROP^{+}02, RW02, RES^{+}08, RC05a, RC05b, RBJ^{+}06, SSH^{+}07, SJH06, SJW03, SS01, SS04, SdB06, WSL^{+}02, WS06, WS04, WST06, WSEL09, WBR03b, ZRL06]. freshwaters [SAE^{+}08]. Friday [CC02]. fringing [JLP^{+}08, RMK^{+}06, THA01, TMA^{+}01]. fronds [SHS01]. Front [GST^{+}01, CFO^{+}00, NDP^{+}07, SLB^{+}00]. Frontal [BTB^{+}08, GRG02, BMH^{+}08, SRZ04, SLKP05]. fronts [DMEPA04, HP03]. frustule [EH00]. Fuca [THL^{+}09]. fuel [CPL^{+}04]. Fulfilling [MW06b]. fully [LASR09]. function [Agr05, BPF^{+}08, GSFB02, Hei08, Kir03, LVSW07, MVBM04, MZV03, OAA^{+}06, PG00, TFNT07, VM05, ZLL^{+}02]. function-environment [Hei08]. Functional [AK00, GDL^{+}00, MA00, FB05, Hei08, RLM^{+}08, RHJ05]. functioning [BKM05, LLT05, Nor04]. functions [MBF^{+}03, OZM^{+}01]. Fundamental [BVN^{+}02]. fundyense [JTL01]. fungal [KVdB^{+}04, New01]. Furcilia [MAS^{+}02]. Further [HPHT00]. fusiformis [LIE06]. future [AWE07, FS09a, HFZ^{+}07, Boy01].

G [Kne02, Ris02, UGSMN01]. Gadus [KBD^{+}00]. Gaines [SG02]. galeata [BRE02, JS03a, vE02]. Galician [SGG^{+}08]. Galilee [RVP^{+}09, KKS^{+}01]. Galveston [HLCG07, THWS00, TWS01]. games [KL01]. Gammarus [DHKH05, GBL^{+}04, ZB03b]. Garda [Sal05]. Gas [BDS^{+}04, CW03, BSC^{+}09, FUG02, NDT07, TN09]. gaseous [PAF^{+}04]. gassy [TKBJ05]. gastropod [AM08, FMS04, PPL^{+}02]. gatunense [SEL^{+}02]. Gdańsk [WSBL06]. gel [DKA^{+}03, DMEPA04, OV03]. gelatinous [AWE07, KKKT07, MCB08]. gels [ET08]. gene [CHW^{+}01, FPZ09, FRO01, LPAQ^{+}05, PCDB09, ZMJ^{+}07]. general [SKM05]. generated [GGIAY08]. generation [PFL07]. generators [CTCT03]. Genetic [FRO01, GBL^{+}04, DVV^{+}07, HCH00, RA00, RNA06, SJR^{+}04, VVD^{+}09, ZRL06]. genetics [JGE^{+}09]. Geneva [ASI^{+}02, LMB05, PGM^{+}09]. Genome [SMA00]. genomic [RSS^{+}00]. genus [WGKE04]. Geochemical [BLWB08, DD09a, KCCE06, TCL^{+}09, BNM^{+}09, LvdB09, LWB^{+}01, MBC06, RMG^{+}04, ZC02]. Geochemistry [BCC^{+}04, Can00, FS00, NK05, RLC^{+}01, SFB^{+}09]. Geographic [HBG00, NFFK00]. geology [RPR^{+}08]. geometry [RS02, ZB03a]. geomorphic [SFC^{+}00]. Geophysical [Caf01, Dow02]. George [Gou01]. Georges [ODR^{+}08]. Georgia [LAN08, AW00a, AW00b, Cra07]. geostrophic [CFO^{+}00]. geothermal [SO07]. German [TKBJ05, JHW05]. Germany [CCCC07, CCC09, MLPS09, WBP^{+}06]. get [JKLR08]. Ghent [Gou01]. giant [FSB^{+}08, GRR^{+}07]. Gili [Wan02]. Gironde [DAC^{+}07]. GIS [HBG00]. Glacial [CWJ09, GMRD04, SCY00, TMUW02]. glacier [BSS^{+}01]. glacier-fed [BSS^{+}01]. glass [YWR^{+}07]. Glider [DORS08, GJT^{+}08, NEE08, NCPd08]. gliders [BF08a]. Global [HB03, Lee01, Lon00, MSVS^{+}07, MV03, BJBA03, Cal01, Cul06, DPC^{+}06, MLRMR09]. globally [XLF^{+}03]. Globigerinella [LEML09]. Globigerinoides [LEML09]. globosa [DK06, EPS00, TJV01]. Gloeotrichia [VMS09]. Glover [SGLT08]. glucose [SWSJ02]. glucosidase [AH02]. glutathione [DGL^{+}04, HMB06, TSKA05]. glycoproteins [YT03]. gobiid [OKO^{+}01]. governing [RBW09]. governs [OS01b]. gracile [TGF00]. gradient [BMT^{+}04, BDK^{+}06a, BdG02, CSK^{+}00, CS06, DMEPA04, FF04, FSS^{+}00, GCM^{+}09, GK03, JCBA08, KB07, KLWS08, KVR03, LFFC06, LZ09, MH00, PDB03, SIH^{+}06, SMP05, SNFT07, SWSJ02, SLL^{+}04, TLS01, TSMB05, dGB02]. gradients [BBCW09, CWC^{+}07, DET04, FDP^{+}08, FDH^{+}01, HOFD01, HMR^{+}07, MSH^{+}06, SMA^{+}06, SLS^{+}02, WWWY05]. grains [SrL05]. Grand [HT03]. granulata [VOG^{+}08]. granulatus [EMPI07]. grass [BZ02, DJH05]. grasses [BZH08]. grassland [SNFT07]. gravel [SA00, SFC^{+}00]. gravitoids [GDL^{+}00]. gravity [AI01, HBI01, PH03]. Grazer [TCC06, TBG01, FXL04, KDO09, SCP08, WDVA08]. Grazer-diatom [TCC06]. grazer-induced [SCP08]. grazer-mediated [WDVA08]. grazers [FA03, GNU00, TJV01, TMP02]. Grazing [AVA^{+}02, Cal01, CP03, CL04, CCS06, CPHM04, CLL^{+}09, CCC^{+}01, EKD^{+}07, EW08, GMR^{+}09, GGH07, GRB02, HBH00, HK01, JRS03, KSP07, Law03, LB00, MFF07, OLCT08, PG00, SK05, SMO07, VVD^{+}09, WL06]. Great [JBM^{+}07, MGH08, XLF^{+}03, SHGM06, ARK^{+}00, Chi00, FLWP09, GEP08, KDS03, LLvW^{+}03, PGL06, PLMPN08, SSK06]. greater [LLF^{+}09, SBR^{+}06]. Greedy [HFLD03]. Green [MLG^{+}03, GWH06, Lür03, NT08, OGR08, SBZ^{+}09, VJV^{+}03, VVD^{+}09]. Green-fluorescent [MLG^{+}03]. greenhouse [SCV04]. Greenland [KTAL04, RBJ^{+}06, SM01, VJV^{+}03]. Greifen [PWS09]. gross [AD05a, BW00, DdG09b, HHY^{+}05, SAH^{+}05, SAHS06, BW99, SMG06]. ground [Sta01]. Groundwater [CIWH09, KCV06, BSA^{+}00, CBA01, CB04, Cha07, CDBC00, DBWS06, GGM03, GMMW03, HU07, HLD^{+}03, HLK05, KM02, KH03, KC08, MSLH02, PSS^{+}06, PBGJ09, RVP^{+}09, RB09, SBC^{+}08, SSA^{+}04, SMGP06, THA01, TMA^{+}01, dSYL^{+}08]. groundwater-augmented [BSA^{+}00]. groundwater-derived [CBA01, SBC^{+}08]. groundwater-dominated [Cha07]. groundwater-fed [HLD^{+}03]. Groundwater-transported [KCV06]. groups [BRT^{+}03, CD03, KDMC05, LMSE05, YNCK04]. growing [GR01, GGP01, GM03b, WCGBG07, WWS^{+}00]. grown [EM08]. Growth [AFH07, ASAC04, BGR^{+}09, BCV^{+}08, BPF^{+}08, DBGR09, FKCF00, GD00b, HB03, RQB^{+}00, SON^{+}03, TGdB^{+}01, TvdWdB04, YNCK04, AKE04, ABJ^{+}06, ASGA^{+}07, ALRM08, Ano04, ACW02, ABFS09, ASH^{+}06, BSC^{+}09, BB05, BSR^{+}07, BMSA08, BGJ01a, BGJ01b, CL04, CC01, CF01, CLL^{+}09, CG01, CG08, CLP^{+}03, Cul06, DBS^{+}04, DTM^{+}08, DGV01b, DCMS03, DSW05, EKdF07, dSFFTD07, FMG01, GP00, HDSC09, HBH00, Har02, Har01, HHB05, HL06, HFZ^{+}07, JS04, JVL00, JTL01, KNC^{+}09, KB06, KLS^{+}06, KLS^{+}07, KLG09, KMC^{+}00, KLL04, KTMM00, KSWC^{+}03, LSPC02, LASR09, LB04a, LALE04, LLB07, LG04, LFPH07, LB00, LCPS05, LH06, Lür03, MHRW02, MGH08, Mar05, MBK^{+}09, MFG^{+}09, MML^{+}08a, MPZ08, MF01, MF02, MNM^{+}09, MC01, MB05, OIH^{+}03, OLCT08, PP04, PQP05, PG00, PSJ00, PAS^{+}05, PBGV04, RP06, RB06]. growth [RC07, SWSJ02, SP04, SSV^{+}04, SWN08, SSS08, SWH^{+}03, SEL^{+}02, SH04, SH05b, SH07, SMDG04, TMK^{+}07, TJV01, TSH^{+}06, Too01, VW01, VPA02, WGKE04, WSL^{+}02, WS06, WST06]. guidance [WW01]. guide [CK00]. Gulf [DBWS06, GGDC04, AKG^{+}05, BW01, CG09, DS07a, FPZ09, FWM^{+}05, GHM^{+}09, GL07, Has06, HVP^{+}07, KL00, KHH^{+}03, LAG^{+}03, MLC^{+}07, MKK04, MKVK05, MH04b, McM00, MBH^{+}06a, MBH^{+}06b, SBC^{+}08, SRT^{+}03, SMGP06, SMO07, SBM04, WSBL06]. Gullmar [Cro03]. gut [GM04a]. guts [BSSL05, CMQ^{+}00, LKM03]. Gypsum [PSS00]. gyre [NEE08, VD08, ZMJ^{+}07, BNK02, BK03b, CDK02, CLAK08, GAM^{+}01, HPLG09, LKAB04]. gyres [Mar05].

H [AE08, Caf01, ORH01, VTO^{+}05]. Håkon [dBSN^{+}06]. Habitat [PVD^{+}07, FL08, GRCS06, HPT^{+}09, HLNM^{+}03, KK00, MDJ^{+}04, PVD01, SWD^{+}05, SF04, VHL^{+}04, vdHPH^{+}09]. habitat-specific [VHL^{+}04]. habitats [ARO^{+}04, DJH05, HK01, JS03b, LRS^{+}03, SJW03, WGN^{+}08]. Haddock [AE08]. half [TvdWdB04]. half-saturation [TvdWdB04]. halibut [FL08]. halides [HBN^{+}00]. Halimeda [SSV^{+}04]. Halisarca [dGMH^{+}08]. Halocaridina [CRY^{+}08]. halocline [EDF^{+}03]. halogen [HM09]. hamatus [Goe08]. Hampshire [WBR^{+}03a]. Handbook [Bel01]. handling [HK06]. Haptophyceae [Leg01]. Haptophyta [Leo08]. Harbor [CHB08, CBH08, CC02, RB01b, TRF^{+}06]. hard [FLWP09, JM02, MWW06, TG09]. hard-water [FLWP09]. harmful [BGR^{+}09, BGC04a, CFG07, EJL^{+}06, PMK08, SWH^{+}05, SH03]. harpacticoid [KKT05]. Harrison [Wro01]. hatch [BCBS04]. hatchability [BSSL05]. Hatteras [GCP^{+}02]. Haunted [DLDE04]. Haven [RB01b]. Hawai’i [GGM03, Par01, PBGJ09]. Hawaiian [CRY^{+}08]. Hay [SG02]. HCO [BARS01]. head [AD05b]. headwater [HBM08, HTRM^{+}07, LVSW07, LZ09, WWMB05]. health [CK00]. Heat [JKLR08, DOE^{+}06, JLB^{+}06, SKTD04, ZJ03]. heat-induced [ZJ03]. heavy [CSK^{+}00, OFKH01, RB01b]. height [MKM06]. Helgoland [WMW^{+}08]. helgolandicus [IHH^{+}00, IHHH00]. helium [CH01]. Hemiptera [HGB00]. hemistoma [CCGH05]. Hemminga [Fou02]. hemprichii [SHBM01]. Henlopen [HU07]. herbivore [Har02, SS01, SS04]. herbivorous [dSFFTD07]. Herbivory [AP00c, WW00, BF08b, CT06, JM02]. hermatypic [LMPY00, SE06, SLDE09]. herring [AH03]. Heterocapsa [KHH^{+}03]. heterocystic [CPHM04]. heterocystous [MHKP02]. Heterogeneity [LRYS06, HSCF06, KNH00, LH05, MAM07, Ost03, PHF03, PCDB09, RM09b]. Heterosigma [BGC04a, LCS02, SNM^{+}08, TDT01]. Heterospecific [Goe08]. Heterotrophic [RlBW01, AGMP04, BRT^{+}03, BMCvE06, CKPS04, CNGV09, Coc01, CJB00, DAA^{+}01, GLP^{+}08, GAPR06, GNU00, HH04, JS04, KHP^{+}06, LN02, Nil03, OBSD04, RC07, SRZ04, SR07, SMH^{+}06, SWH^{+}03, TRT04, WNAW09]. heterotrophs [RH06]. heterotrophy [ÁÁS07, CPCK00, RBC00]. Hexactinellida [YWR^{+}07]. hexafluoride [PLK^{+}05]. hexafluoride-based [PLK^{+}05]. Hg [GKS^{+}02]. hidden [PHF03]. High [BIIA03, CMQ^{+}00, FDH^{+}01, GKJ04, GBA00, GZY^{+}08, HPE01, KJH03, LPB06, MSL^{+}03, NDT07, NAH^{+}08, SLJ00, ARK^{+}00, AY08, BM01, BRL03, BSS^{+}01, CMSVA02, CTWFP06, DZL^{+}03, ES00, EB01, FSBB05, GGIAY08, Har02, HTWG08, JGSB07, KKH06, LBW^{+}08, MPGB08, MMLH06, MSVAC04, MBPVR06, NSM09, OHUB00, RB08, RA06, Reu02, RM04, SMC03, SCM08, SHS03b, TVSS01, TMA^{+}01, TBH07, VRC07, WDVA08, ZWC^{+}04, MVA^{+}09]. high- [FSBB05]. high-affinity [JGSB07]. high-altitude [BSS^{+}01]. high-elevation [NSM09, SMC03]. High-frequency [BIIA03, LPB06, GGIAY08]. High-latitude [NAH^{+}08]. high-light [MPGB08]. high-molecular [BRL03]. high-molecular-weight [EB01, RA06, ZWC^{+}04]. high-nitrate [MMLH06]. High-resolution [FDH^{+}01, GBA00, KJH03, MSL^{+}03, NDT07, BM01, DZL^{+}03, ES00, OHUB00, SHS03b]. higher [KAM04, WAB^{+}08]. highly [CR06, FRO04, HMS^{+}04, HDVB01, TWS03, VMC^{+}07]. hillebrandii [MHDH00]. Himal [SCY00]. Himalaya [SCY00]. Himmerfjärden [BEM^{+}02]. Hiroya [Gou01]. Historic [SSK06]. historical [CCKC05, JGE^{+}09, KSS^{+}07, PLG^{+}02]. histories [BB03]. History [EBS07, BCMR00, CWPL05, CHW^{+}01, HCH00, KR08, MB00, MGVM^{+}06, RRCP03, Sak00, SD05, SPW^{+}05, TBHS01, VAS02]. Hoare [VH06]. Holarctic [TIH02]. hold [KHKS05]. Holocene [CPGC09]. hologrammetry [RWFK00]. holomixis [SKTD04]. holoplankton [BH02]. holosymbiont [DOE^{+}06]. Holthuis [CRY^{+}08]. Holzmaar [MLPS09]. Homarus [Ann05]. homing [TGF00]. homology [RSS^{+}00]. Hook [SWD^{+}05]. Horizontal [MWH01, SIK07, SI03, PF01c, RS09, SWSJ02]. host [BAPMMB01, BRE02, FBW^{+}08, GFPG09, GCG05, Hol00, TGFPT08]. host-symbiont [GCG05]. hosts [GF02]. hotspot [vODL^{+}09]. hours [TG03]. Hudson [ACD10, TWS03, dGP08]. hudsonica [BB04b]. human [BEW^{+}00, RMG^{+}04]. human-induced [BEW^{+}00]. humans [LFA^{+}09]. Humber [DNS06]. Humboldt [HHW^{+}02]. humic [AG03b, HPHT00, JMBJ01, KHP^{+}06, KJMJ03, LvdB09, OH00, SBL06, VSSTS00]. Huron [BDNF01]. Hurricane [TSML07]. huxleyi [BTdB03, DNB^{+}04, DA05, EZA^{+}05, EMWL06, EKD^{+}07, EW08, FHB^{+}09, GRV^{+}09, GD01b, LGN^{+}06, Leo08, LMT05, MAL08, RZR02, SXLM06, SWH^{+}03, TSG^{+}01, TLrR07, XTSM07, vdPVB07]. hyalina [JS03a, SD05]. Hyalodaphnia [SJR^{+}04]. hybrid [BvE07, KS04b, SBS07, WKB^{+}04]. hybridizing [KWTS07]. hybrids [HD09]. Hydraulic [FS00, LB04b, RB03]. Hydroclimatic [BRP^{+}06, LIE06]. Hydrodynamic [FAM^{+}06, HL06, JTAM09, OD08, RKK09, BNM^{+}09, HW09, JGE^{+}09]. Hydrodynamics [Bat00, CTR^{+}02, Fol05, Ano08a, MPB^{+}08, NA06, SP09]. hydrogen [CYN05, PLM^{+}09, PM08, RKGC^{+}01, TDT01]. Hydrogeomorphic [BSG^{+}08]. hydrographic [DMEPA04, SGLT08]. hydrographically [HLGI08, MHK^{+}07]. Hydrologic [Cha07, HBAK09, RB03, GMRD04, MS06, SDJT03]. hydrological [CJ08]. hydrologically [VWPP^{+}06]. hydrology [PLM^{+}09]. hydrolysis [Arn00, MGL02, ML09]. hydromedusa [CCGH05]. hydromedusae [CC02, HK06]. hydrophobic [TLM04, VOG^{+}08]. Hydrophobicity [TM05]. hydrothermal [AM08, CH01, CSP^{+}01, CMCB^{+}09, Jun01, Met04, RKGC^{+}01, RSS^{+}03, SWD^{+}05]. hydroxamate [KTMM00]. hydroxyl [PBGV04]. hyperborea [ST03]. hypereutrophic [DGV01a, DGV01b, GSH04, VOG^{+}08]. hypernutrified [DNS06]. hypersaline [LAHdBJ05]. Hyperspectral [JP03, APH^{+}03]. hypertrophic [TB05]. hyphomycete [BB04a]. Hypolimnetic [MHB03, JLB^{+}06, KNN06, ME06, RENA06]. hypolimnia [RM09a]. hypolimnion [MED^{+}08, UL05]. hypolimnion-related [MBDL01]. hyporheic [FVD01, FSSH03, HDP09, SA00]. hypothesis [AKE04, AKG^{+}05, dSFFTD07, KHKS05, KW04, KLWS08]. hypoxia [DS07a, Dow02, LBG^{+}09, MJEM06, RHCS08, SRT^{+}03].

IA [SHSH^{+}07]. Iberian [ÁSGMP01, CFG07]. Ice [Boy01, HM08, BJG^{+}00, BGV02, BGHA04, DJB^{+}07, FS09a, JBM^{+}07, KSS^{+}07, LLB04, LA09, MVA^{+}09, PTK^{+}07, PL09, PDS05, PCD^{+}08, RG04, SKC^{+}01, TS08, VGP^{+}08, WSNH06, WCM09]. ice-covered [PCD^{+}08, VGP^{+}08, WSNH06]. ichthyoplankton [STK^{+}04]. Idaho [DFBA07]. Identification [GHM^{+}09, JL08, MKK04, MHB01, Dix00]. identify [GCG^{+}02, RBW09]. Identifying [RLM^{+}08]. identity [BA01]. II [GKS^{+}02, RWK08, WKSM08, AMS^{+}02, ASH^{+}06, Coc01, JHW05, OHUB00, RSS^{+}03, SWS^{+}03]. III [BRT^{+}03, CT05, NKS^{+}02, RW02, BVWB07, MAS^{+}02]. Illinois [WGN^{+}08]. Illumination [CLL^{+}03]. Ilyanassa [FMS04]. imagery [DZL^{+}03, MSL^{+}03, SHS03b]. imaging [OHUB00, WdBD^{+}01]. Imbalanced [TMK^{+}07]. immersion [HBH^{+}06]. immigrate [KAW05]. immigration [PW01]. Impact [BZ02, PGP01, BJW^{+}02]. Impacts [GRB02, DAL02]. Implication [GM04b, HST^{+}02]. Implications [AG03b, BST^{+}05a, BGV02, Ber01, BBKC03, BRS01, BLBB07, CRJ08, DR06, Fab06, GBL^{+}04, GFM^{+}01, GSH04, HABF03, HLL^{+}06, HFZ^{+}07, IKT03, JLB^{+}06, JB00, JS03b, KHM04, MTC^{+}06, MC01, PNPA04, PIF08, RA06, RTT^{+}02, SXLM06, SABA02, SSM^{+}09, SdB06, SH00, TS01, VR01, WGD^{+}02, YPBL07, BFD^{+}04, BSSL05, CV05, CA08, Duf09, FB05, Gay00, GLVFH09, HA08, KM02, LB04a, MH00, Par01, PSJ00, PH03, RCL^{+}02, RC07, SH10, SMP^{+}01, TB05, THA01, WLRT06, WS06, WSKL06]. Importance [CMSVA02, GGRJ03, LLB07, MRM07, PW03, QSH^{+}02, ZSS^{+}08, AR06, AH01b, BFC05, Bou04, CNGV09, CCM07, GJS01, HSB00, HLP08, LFFC06, ML02, SGR08, SKR01, SMGP06, SR02, SP04, SR04, VMC^{+}07, WNP04, ZLP^{+}03, dVKT^{+}09]. important [DBWS06, Goe02a, PSS^{+}06, PN06, WA01]. impoundment [HLL^{+}06, MPA^{+}07]. impoundments [DPC^{+}06]. improve [JJP04, Sos00, VW01]. Improved [WAB^{+}08]. Improving [BMSA08, SF04, RL07, TSH^{+}06]. IMS [Ano04, KSWC^{+}03]. IMS10 [MB05, PM08]. IMS101 [MC01]. In-channel [ED05]. in-stream [HDP09]. in-water [SSMM00]. inactivated [RHW00]. inactivation [KMHN07]. including [MY05]. Incorporation [KNE08, BFD^{+}04, CD03, EH00, KJ00, TS00, VEMM07]. increase [HHS05, HW09, LHWE01, MD09, PD02, WJA^{+}07]. Increased [AG03b, BCDH09, Ham05, LNB02, NRB08, LGJ02, MSVS^{+}07, SPM04, TYK^{+}08, TKK02]. Increases [OOB09, AWE07, AE08, CDK02, Had08, SNPK^{+}09]. Increasing [VMKN03, FL06, GCM^{+}09, RG08, WK09]. incubations [MPACG01]. independence [dGP08]. index [HSB^{+}06, WIK08]. India [BMD^{+}04]. Indian [MCJ^{+}06]. indicate [HBF^{+}06, HMD^{+}07, PAS^{+}05, SMA00]. indicated [GJS01, HPAH^{+}02]. indicates [JHW^{+}07]. indicating [DST08]. indication [MBW^{+}01]. Indications [PVD^{+}07]. indicator [HP06, HPHT00]. Indicators [Kra05, DdG09a, HSR^{+}08, HMR^{+}07, IR06, KK00, RG07, TRF^{+}06]. indirect [EMPI07, HGB00, HMD^{+}07]. indirectly [Har02]. Individual [DG06, HKF01, DKKA09, GHN^{+}02, Has06, KKT^{+}07, OHUB00, PE07, RWFK00, TBF04]. individuals [HKF01]. Indonesia [AJS07, COK^{+}08, SHBM01]. induced [BEW^{+}00, BST^{+}05b, CWLH06, DMS^{+}03, ESSS01, FHDKR03, FB05, GN00, HHS05, HDP09, JVL00, KBL08, LA06, LPW05, MCJF09, MGGM06, RK02, RHN^{+}02, RH03, RMG^{+}04, SCP08, SSS08, SQTA07, UM05, VAS02, VTO^{+}05, VM06, WBH03, WAGH09, YYTH07, ZJ03, vWBT09]. induces [Gil03]. Inducible [LBT06, CF05]. Induction [SS03, GRDCS09, LV01]. Inertial [Mor06]. infaunal [MW02, TPFS09, WMWV04, ZLP^{+}03]. infected [EW08]. Infection [Hol00, BVN^{+}02, BVG^{+}07, CDD07b, WAGH09]. Infectious [EMWL06]. infer [HW07b, SGPM09]. inferences [CFO^{+}00]. Inferred [GCM^{+}09, CPB^{+}08, DPC^{+}08, RNR08, SSK06, TvG02, vdSvdSG^{+}00]. inflow [AELS04, BI07, HLK^{+}07, Kra05, LTHW04]. inflows [RM09b]. Influence [BS03, CWJ09, CF01, CPH06, CLP^{+}03, CHT02, DTL^{+}05, DFBA07, DGM^{+}05, FSM^{+}07, JDS08, LB04b, MAL08, NCO02, TSDK^{+}09, WV00, WPD^{+}01, ZJC^{+}07, dCCL09, AD01, BSJ02, BBP^{+}06, BFCLÁ08, BDNF01, CSWD07, CTR^{+}02, DGH^{+}07, EKdF07, HDP09, IHHH00, KD02, KLG09, KN07, Leh00b, LG04, MPJ08, MCB08, MTM^{+}01, MML^{+}08b, Nil03, PTLJ01, RM06, RGAS04, RHJ05, SA09, SSV^{+}04, SJW03, SC08, SŻKL08, TNM^{+}01, VPA02, WLP^{+}07, WM07, ZB03a, dlFSIN08]. Influences [CW01, EBS07, LVSW07, SSAZ01, AS00b, Ber03, BDK^{+}06b, FKS^{+}01, HFE07, HB08, KTM01, NSM09, PVP^{+}06, Rit09, WPW^{+}09, WW02]. influencing [BSM01, DKA^{+}03, HDWDD09, SNM^{+}08, ZW06]. Infochemical [VZM07, JJT07]. Infochemical-mediated [VZM07]. information [HBG00]. infrared [BS02a, DRM^{+}09, RDPM06]. Ingestion [RCD^{+}05, Fra01, NHN^{+}05, SDC00, TFNT07]. inhabiting [AVA^{+}02]. inherent [BZ03, GSO03, HDFM08]. inhibit [KNC^{+}09]. Inhibition [SWS^{+}03, SEL^{+}02, JVL00, YS06]. Inhibitory [MS05]. Inhomogeneous [MS03]. Initial [Sar07]. initiate [JCK^{+}08]. initiated [TKAW03]. initiation [PLK^{+}05, SNM^{+}08]. injection [KLF^{+}03]. Inland [TMYH02, Boo00, SPG05]. Inlet [Tun00, LB04b]. inner [GZL02, MSSW^{+}07]. inner-shelf [GZL02]. Inorganic [RlBW02, SDJT03, TRM00, TMC06, TPG^{+}08, TSML07, ASS03, BTS^{+}06, CWWS00, CLPB02, CSK^{+}07, DGH^{+}07, EPS00, GGS06, HDVB01, JR07, LTM01, LJCM08, Lee01, MOK^{+}03, MCPP06b, MPZ08, NA03, NA06, NP01, NH06b, PSS^{+}06, RBC00, SMH^{+}06, SWML08, TM02, TMA03, WC04, WKH^{+}06]. input [CBA01, Cra07, KM02, RG02, WHC^{+}06]. inputs [AMS^{+}08, DBWS06, ET08, EBS07, MSVS^{+}07, MBPVR06, NSM09, RORR^{+}09]. insect [Wur07]. insectivorous [MID^{+}07]. insects [MIDH05]. inshore [FDP^{+}08]. Insight [Cha07]. Insights [BJBA03, PWS09, SFB^{+}09, SHA08, SAP^{+}06]. inspired [BK07]. instabilities [GAM^{+}01]. instantaneous [ASAC04]. Institut [Wan02]. instrument [GTK^{+}01]. instruments [KDS03]. intact [TN09]. integrated [SAS08]. Integrating [SDJT03, SCS^{+}08, VVSM09]. integration [SS01, SS04]. integrative [Wro01]. integrator [PSS00]. integrators [Sch09, WSVS09]. Intense [CJB00, GMR^{+}09]. intensities [MBC^{+}05, TLrR07]. intensity [DKKA09, HH04, PVD01, Shi04, SCM08]. intensively [PWN07]. interact [MW06a]. Interaction [CWC^{+}07, HSS06, MPB^{+}08, CMSVA02, FMG01, LJCM08, MSAC06, Sar07, STC03, WA01, Ano08a]. Interactions [AVJ^{+}00, CGH^{+}04, FMG^{+}08, MLSS05, NHPHH08, WdMD^{+}00, BBCW09, BRE02, BRM^{+}07, BI07, GL07, HHB05, HL06, HH04, KBE^{+}04, LW02, PCvRR00, PvE00, SWH^{+}05, TJV01, TBG01, VMS^{+}04, VZM07, Mil02]. Interactive [DSW05, DGD04, Ham02, HB08, WSL^{+}02, AY08]. Interannual [HB06, LBB^{+}04, LHB04, NBW^{+}02, PBMSG04, SKJ^{+}07, SJ00, ASI^{+}02, DBTK02, LAJ^{+}04, MLPS09, PVD01, Pil04, Sal05]. Interbasin [UL05]. intercontinental [HCH00, HWA03]. Interdecadal [SRL^{+}09, MMKN08]. Interface [BMH^{+}08, BNM^{+}09, GMMW03, HFDO05, LMMW03, MCvEH08, MCJ^{+}06, SLHK02, SSAZ01, SDV^{+}06]. interfaces [CGCPA08, GTK^{+}01]. interfere [HTS^{+}03]. interfering [DGV01a]. interflow [CWLH06]. intergenerational [HR03]. intergenic [WGKE04]. interior [BMWW08, PPH03]. interlake [MM03]. interleaving [LCLP02]. intermediate [BCV^{+}08]. Intermittent [BMWW08, LA09]. intermolt [TSH^{+}06]. Internal [EMK08, LMB05, AIS00, AI01, AI03, BIIA03, BII05, GGIA06, GGIAY08, HISW00, LPB06, MCJF09, MBF^{+}03, Mor06, PDM^{+}09, SKR01, STG^{+}07, SI08, SSV^{+}04, TG03, UL05, VCCS05, VRC07, dlFSIN08]. interplay [LAG^{+}03]. interpretation [BPLH04, LW01a, SF04]. Interpreting [HR07, HL02, OHCP02, TB05, AWH07, OSDB07]. interreplacement [XTSM07]. Interspecific [RHM04, LH06]. Interstitial [All00, ZWHP04]. Intertidal [CWWS00, BBG05a, GBA00, Har02, HH03, HD03b, HR00b, KNC^{+}09, MBB^{+}00, NBW^{+}02, PDR^{+}09, PSH08, Rit09, RLJdB08, SMA^{+}06, SMP05, TMG03, VvOBM06, VEMM07, WBP^{+}06, WW02, dBWF^{+}05, vOMSM06a, vOMSM06b]. Intestinal [RCD^{+}05]. intracellular [BS03, TMC06]. intraguild [MDJ^{+}04]. intraspecific [GKJ04, LH06, RHM04]. Intrinsic [OLCT08]. introduced [AHF^{+}05]. Introduction [DIMP08]. intrusion [DCC^{+}01]. intrusions [RHN^{+}02, WRG^{+}05]. inundation [GK03]. invader [BJW^{+}02, MGH08]. invasion [CHW^{+}01, DPK^{+}00, FS09b, FHBK06, GLVFH09, HBG00, HBM06, MDJ^{+}04, MVD05]. invasions [GJRM07]. invasive [BFC07, GSA^{+}05, MTC^{+}06]. inventories [CDK02, FNLG00]. inverse [FSK^{+}04, MKV01]. Invertebrate [SA00, BDvO^{+}03, BFML05, DL00, FRO04, GJRM07, Kiø00, MD09, QLD^{+}02, TMP02]. invertebrates [BBG05a, FS09b, GM04a, HBF^{+}06, LW01a, NBW^{+}02, SLB^{+}00, TSA09, VMJ08]. investigate [PF01b]. Investigating [EKdF07, IR06]. Investigation [LMBP07, ZRL06, MCPP06b, RRH^{+}03]. investigations [SFB^{+}09]. iodate [TKAW03]. Iodine [dlM09, HM09, TKAW03]. iodocarbons [HMNL06, HMT^{+}08]. ion [TF00]. ionization [SHL^{+}05]. ions [OHT06]. Ireland [BLG^{+}07]. Iron [Ano04, BFCS^{+}01, BRS01, BDK^{+}08, BAP^{+}05, HMR^{+}07, KSWC^{+}03, LDC^{+}01, LCPS05, MBL^{+}01, MC00, MMLH06, OBSD04, RTT^{+}02, TBF04, ACW^{+}03, AWLH04, BM00, BSR^{+}07, BGC^{+}04b, BST^{+}05b, BJD05, BLBB07, BB07, CT05, Car01, CSE^{+}08, CFMH01, CW01, Coc01, CLP^{+}03, CCC^{+}03, Cul06, DGM^{+}05, ET08, EH00, ESSS01, FSBB05, GRG02, HHB^{+}09, HPL^{+}06, HB08, HHW^{+}02, JCBA08, KS04a, KTLS03, KMC^{+}00, KA06, KTMM00, KSM^{+}05, LvdB09, LTM01, LJCM08, LBC^{+}04, MAC^{+}06, MCPP06a, MH07, MW06b, MKS08, NKS^{+}02, NP01, NGS^{+}07, PP04, PQP05, PTKP00, PRKHW04, Pri05, SRNW06, SKM05, SH05a, SH04, SLS^{+}02, TGdB^{+}01, TvdWdB04, TWR04, UWA^{+}07, WJMD07, WAGH09, WTC^{+}05, WHC^{+}06, NKS^{+}02, RWK08]. iron-deficient [PP04]. iron-induced [BST^{+}05b, WAGH09]. iron-limited [Car01, Pri05]. iron-rich [PTKP00]. iron-storage [CSE^{+}08]. Iron-sulfur-phosphorus [RTT^{+}02]. IRONAGES [BVWB07]. IronEx [Coc01]. Irradiance [FQR^{+}06, BGV02, Fab06, FMG01, JTL01, Kir03, WW00, Zim03, vdPVB07]. irradiances [MW06b]. irradians [TG09]. irradiation [MSHE09, Shi04]. irrigation [GGDC04]. Isaza [OKO^{+}01]. ISBN [Bel01, Boy01, Bri01, Caf01, Dow02, Elm02, Fou02, Gou01, Jun01, Kne02, Mil02, Pom01, Ris02, SG02, Sta01, Wan02, Wro01]. Island [BBG05a, HLK05, LRS^{+}03, NK05, PBGJ09, ZVRL03, STK^{+}04, Ano08b, CSG^{+}06, WIK08]. Islands [CRY^{+}08]. isolated [BGJ01a, BGJ01b, TGFPT08]. Isolation [CT02, Goe08, HMD^{+}07, KWTS07, ZRL06]. isolation-by-distance [HMD^{+}07, ZRL06]. Isopoda [ZB03b]. isotope [BMD06, BHV04, CCKC05, CCH02, CCCC07, CCC09, CT02, CRF01, DBS^{+}04, EVW07, ES00, Fin04, GR01, GMMW03, GSNH04, GSLT08, GJS01, GKJ04, GCS06, HAA09, HLD^{+}03, HBF^{+}06, HM06a, HB06, JGSB07, KTLS03, LW01a, LBG^{+}09, LBC06, MT06, MM03, MBCD04, MCC02, NDT07, NEC03, OKO^{+}01, PDH06, PWS09, RVP^{+}09, RH01, RSS^{+}03, RZR02, Sch01, SAS^{+}03, SHSH^{+}07, STST07, SPR02, TB05, TMA^{+}01, TBH07, TTG^{+}09, TR03, VVSM09, YPBL07, vOMSM06a, vOMSM06b]. isotopes [BRB^{+}01, BCMR00, BLS^{+}06, CBA01, Cha07, CPB^{+}08, CH01, DBE08, DPC^{+}08, EPP^{+}08, HPLG09, HL02, KM02, Kra05, KH03, LBSY02, MBR08, MLPS09, NH06a, OHCP02, PLM^{+}09, RPR^{+}08, RB01a, SAH^{+}05, SAHS06, SAS08, SAP^{+}06, SVPK^{+}09, VM06]. Isotopic [DES03, LM02, MVFD04, MAM07, Sch00, Alt01, BSC^{+}09, BKM05, CYN05, EPS00, FHBK06, HVP^{+}07, IR06, KF01, LBB^{+}04, LWP^{+}07, LRYS06, MCPP06b, MVBM04, MMW00, PKKB05, PHF03, PCD^{+}08, VMS09, WKFP06, WBD02, YTVR07, ZWC^{+}04]. Israel [HPE01, HAA09, KKS^{+}01, NEC03, RVP^{+}09, SEC08, SMGP06]. ISSN [Wan02]. Issue [DIMP08]. Issyk [HPAH^{+}02, LWB^{+}01, PFH^{+}03]. Issyk-Kul [HPAH^{+}02, LWB^{+}01, PFH^{+}03]. Italiantartide [Smi00]. Italy [BMU04, HAHS^{+}09, Sal05]. Ivory [LJ09]. ix [Gou01]. Iyo [KKKT07].

Iyo-Nada [KKKT07].

J [ACD10, Bel01, Elm02, Mil02, Ris02, SMS09, Sta01, Wan02, Wro01]. Jamaica [AP00c, GGS06]. Japan [GTS^{+}09, GSB^{+}09, HHY^{+}05, KKKT07, MS05, OKO^{+}01, SAH^{+}05, SAHS06, SNM^{+}08, TMYH02, YYTH07]. Jeju [HLK05]. jellyfish [AWE07, BBJ^{+}06, DKKA09, Had08, KKT^{+}07, LHB04, RG08, WPW^{+}09, AE08]. Jersey [Agr05]. Jesogammarus [INU04]. Ji [KHK^{+}03]. Joaquin [MSJMN02]. Joint [Lon00]. Jolla [WV00]. Jones [Sta01]. jouinae [RB08]. Juan [THL^{+}09]. Juncus [KLSW00]. June [TAB^{+}00]. juvenile [BFS02, FL08, GJB^{+}04, GWR^{+}09, HGB00, HNB^{+}07, HL06, KN07, NA05, RRSC01, REL06]. juvenile-adult [NA05]. Juveniles [AMS^{+}02].

Kahana [GGM03]. Kaltenbacher [BK01b]. Karenia [KHNV05, KSP07, PMK08]. Karlodinium [JRS03]. karst [BdBS^{+}08]. karst-water [BdBS^{+}08]. karstic [CSP^{+}01]. Kattegat [CC04]. Kawanabe [Gou01]. Keddy [Bri01]. Keeling [KJJ07, MC04]. keeps [KWTS07]. kelp [FSB^{+}08, GDK03, GRR^{+}07, KNC^{+}09, SHS01]. Kenobi [Gou01]. Kenya [BDSB07]. KEOPS [ET08]. Kerguelen [ET08]. Key [SWML08, AG03a, CDD07b, KLG09]. Keys [Ris02, LWP^{+}07, SSV^{+}04, ZSS^{+}08]. kilometers [BL07]. kindti [MKG04]. kindtii [VKvD05]. kinematics [HW09]. Kinetic [RSS^{+}03, SHSH^{+}07, SC00, SKM05, WPV00]. Kinetics [HW07a, RW02, BR00, LLA05, LBC^{+}04, NKJM06, QRF06, Shi04, SY07]. Kingdom [DNS06]. kink [Cul06]. Kinneret [HPE01, HAA09, RVP^{+}09, SEC08, AI03, Ber01, BW05, DES03, GGIA06, HSB00, KKS^{+}01, LBSY02, NIE^{+}00, NEC03, Ost03]. Kirchman [Pom01]. Kirghizstan [LWB^{+}01]. Kivu [LTHW04, PDM^{+}09]. Knox [Kne02]. König [DDZ03]. Korea [HLK05, PLK^{+}05]. koreni [JGE^{+}09]. krill [ASAC04, AW00a, AW00b, AMS^{+}02, ASH^{+}06, MAS^{+}02, MFG^{+}09, ORFD^{+}09, PAS^{+}05, RFW08, SAS^{+}03, SAP^{+}06, SHS03a, TCR03, TSH^{+}06, TKM07, WCM09]. Kul [HPAH^{+}02, LWB^{+}01, PFH^{+}03]. Kuroshio [lLCChTO08]. Kyrgyzstan [HPAH^{+}02].

L [Jun01, Wan02]. L. [DDZ03, MTC^{+}06]. L1 [TF00]. L2 [TF00]. label [CT02]. labeling [BMD06, GHB^{+}05, MBB^{+}00, MMW00, VvOBM06, VEMM07, vOMSM06b]. labile [DB07, HTS^{+}03, MMS^{+}02, WWMB05]. Laboratory [BM00, Bea03, LH05, NCO02, WST06]. Lac [FAM^{+}06]. lactuca [MB01]. lacustrine [HDSC09, SNPK^{+}09, SF04, SMDG04, TTSC04]. ladderane [JHW^{+}07]. ladyfish [MH04b]. Lagoon [BMU04, MCJ^{+}06, AHF^{+}05, BM03, DACG01, HLM08, TMA03, WKH^{+}06]. Lagrangian [DIMP08, PLK^{+}05]. lagunensis [LB00]. Lake [AIK04, BDNF01, CJ08, CRBF^{+}06, DFBA07, DPK^{+}00, HBCK03, IKT03, KNH00, Kra05, LWB^{+}01, MCJF09, MS06, RHK^{+}09, RCE^{+}02, WBR^{+}03a, AVJ^{+}00, AIS00, AD01, BCC^{+}04, BD04, BPOT04, BST^{+}05a, BGV02, BDK^{+}06a, BRB^{+}01, BSS^{+}01, BSA^{+}00, CC01, CSM^{+}08, CMSVA02, CPGC09, CRC^{+}03, CCKP02, CSP^{+}01, CCCC07, CCC09, CIWH09, CW03, DTL^{+}05, DT00, DR00, EPP^{+}08, FLT02a, FLT02b, FH03, FUG02, GDB^{+}09, GLP^{+}08, GRdL^{+}08, GRB09, GHTJ00, GGIAY08, GKJ04, GSH04, GCS06, HPE01, HB05, HZG07, Hei08, HLD^{+}03, HAL^{+}09, HISW00, HLP08, HLNM^{+}03, HBN^{+}00, HTWG08, HRS04, JLB^{+}06, JSJ03, JBH^{+}04, JM08, JMBJ01, KHP^{+}06, KAJ08, KTAL04, KNN06, KHB^{+}00, KMHN07, KT02, KCP^{+}04, LVC^{+}00, LPD^{+}04, LBM^{+}04, LFFC06, LH05, LD01, LP07, LA09, LPB06, MSGK06, MFK^{+}09, MMDW09, MGH08, MED^{+}08, MHB03, MKG04, MSC^{+}04]. lake [MSVAC04, MMCW09, MLRMR09, MGMW06, MGVM^{+}06, MS05, NSM09, NHPHH08, PCC^{+}07, PBMSG04, PLG^{+}02, PSLO07, PS06, PGM^{+}09, PLM^{+}09, PL09, Pil04, PAF^{+}04, PCD^{+}08, RVP^{+}09, RA09, RM09a, RCVK09, RS01, RC05a, RC05b, RM09b, RBJ^{+}06, SI01, SCY00, SSvNH01, SEC08, SKR01, SKJ^{+}07, SKTD04, SIK07, SI08, SHGM06, SDKZ^{+}09, SDH02, SSJ07, SFL05, SI03, SJH03, SMG06, TYK^{+}08, TS06, TB00, TB05, TCD^{+}03, TMB07, UL05, VSSTS00, VHL^{+}04, WR00, YPBL07, vWSBT08, ALH01, AHHH^{+}07, ASI^{+}02, AI03, ABDS04, AC08, BHVC05, BHVC06, BVK02, BJW^{+}02, Ber01, BW05, BSS^{+}01, BLHC08, BH07, CWJ09, CH01, CCCC07, CCC09, COK^{+}08, DES03, ES00, FCJ^{+}06, GGIA06, GNU00, HPE01, HAA09, HBG00, HSS06, HSB00, Har01, HHB^{+}09, HHT01, HBS04, HPAH^{+}02, HAHS^{+}09, INU04, JGZ^{+}08]. Lake [KGYU06, KBG^{+}08, KKS^{+}01, KSS^{+}07, LPD^{+}04, LBB^{+}04, LBM^{+}04, LMB05, LTHW04, LBSY02, MRK02, MBF^{+}03, MMW06, MGH08, ME06, MSVS^{+}07, MMPA04, MMO06, MHB03, MLPS09, MS05, NIE^{+}00, NGS^{+}07, NEC03, OKO^{+}01, Ost03, PDM^{+}09, PBM^{+}07, PWS09, PFH^{+}03, PED^{+}09, PGM^{+}09, Pil04, RSC04, RHCS08, RKWG00, RKH03, RMG^{+}04, RJS^{+}03, Sal05, SAH^{+}04, Sch06a, SDG^{+}07, SEC08, Sos00, SSM^{+}04, TTSC04, UGY^{+}00, UVY^{+}01, VMH^{+}01, VH09, VBH^{+}05, VH06, VM06, WSNH06, WMRK06, WRG^{+}05, ZZPP01]. Lake-wide [RHK^{+}09]. Lakes [AOZ^{+}09, Mor06, Sch09, TDC^{+}09, WSVS09, ACD10, AG03b, AI01, AB07, ARK^{+}00, AKP^{+}09, AKH04, BCC^{+}04, BSMR08, BP03, BCSW07, BLJ09, BBJ05, BT00, BHMS06, BGE04, BIIA03, BII05, BBP^{+}06, BMWW08, BLS^{+}06, BLSQ07, BS01, CKPS04, CPV00, CSSNA02, CGCPA08, CSK^{+}00, CA08, CS06, CPCK00, CCKP02, CR10, CMC^{+}09, COK^{+}08, DGV01a, DGV01b, DET04, DHP03, DJG^{+}06, DTFL03, DP01, DPC^{+}06, DPFL09, ESSS01, EBS07, ESI01, FS09a, FHDKR03, FKS^{+}01, FLWP09, FLPL08, GM03a, GDCG08, GP06a, GA00, GA01, GGH07, Gou01, GVT06, GV09, GH00, GHJ^{+}05, HHS05, HC01, HBF^{+}06, HABF03, HBM06, HMB06, Jan03, JPJS06, JJS^{+}03, KCE^{+}00, KTG^{+}06, KJMJ03, KB05a, KJJ07, KFR^{+}01, KLWS08, Kil00, KHK^{+}03, LFZ^{+}07, LVC^{+}00, LIHS03, Law08, LCS^{+}03, LBEP06, LFA^{+}09, LB04b, LFAB06, LLK05, MNA^{+}09]. lakes [MJEM06, MCPP06a, MPAS00, MPA^{+}07, MdG08, MJ03, MIDH05, MID^{+}07, MBPVR06, MVA^{+}09, MWW06, OZM^{+}01, PC02, PGL06, PWS09, PLMPN08, PFN^{+}05, PBC02, QMS^{+}07, QLD^{+}02, RSWH04, RORR^{+}09, RAKE05, RENA06, RRD04, RB02, RS02, RS09, RYS^{+}02, RLM^{+}04, RHW00, SK05, SSvNH03, SLHK02, SFB^{+}09, SH10, SMC03, SP04, SJ00, STP^{+}07, SC08, SKC^{+}01, TPM01, TYK^{+}08, TXH05, TLS01, VJV^{+}03, VMKN03, VM06, WHHS06, WSC^{+}08, WS01, Wey04, WJA^{+}07, WK09, WDL07, WM06, WBR03b, Wur07, XLF^{+}03, vWBT09, JBM^{+}07, SSK06, SDM^{+}09, XLF^{+}03]. Lambertian [CLL^{+}03]. laminated [BVB03]. Land [WSNH06, ATJH08, AD01, BSG^{+}08, FQH^{+}01, GMMW03, HSB^{+}04, KVR03, LVB08, PWH^{+}04, PLMPN08, SDV^{+}06, UMY^{+}02]. land-derived [UMY^{+}02]. land-use [PLMPN08]. lander [FDH^{+}01]. Landscape [HSB^{+}04, LBEP06, QLD^{+}02, RPR^{+}02, AD01, MS06, NSM09, RPR^{+}08, RP06, SFB^{+}09]. Landscape-scale [LBEP06]. landscapes [KTM01, Sch09]. Langmuir [MHLN01]. Large [ASGA^{+}07, HLK05, LP07, NFFK00, PLW^{+}08, PF01a, TLS01, Aks07, AIS00, AI01, BIIA03, BDK^{+}06b, CCS06, CRC^{+}03, DD09b, DGV01a, EGP08, FSSV08, FBW^{+}08, GWH06, GM03b, GSJ^{+}09, HLP08, JMBJ01, KGYU06, KNN06, LAJ^{+}04, LPB06, MM05, MMKN08, PVP^{+}06, PLG^{+}02, RHK^{+}09, RS01, RC05a, RC05b, SB03, TGdB^{+}01, TvdWdB04, UHB^{+}08, UL05, dGP08]. large-bodied [DGV01a]. Large-scale [ASGA^{+}07, LP07, PF01a, EGP08, MMKN08]. larger [HD03b]. largest [Wey04]. Largo [SWML08]. larvae [AM08, DL00, Dix00, FSSV08, FMS04, GKJ04, HH05, MHLN01, MRM07, MAS^{+}02, PC04, PDS07, PHMR^{+}07, PDS05, PPL^{+}02, RKK09, WBvE02]. Larval [GZL02, MK00, AM08, AHF^{+}05, BPC06, BFML05, CTR^{+}02, FNM07, GCP^{+}02, HCC^{+}02, JB00, JGE^{+}09, MBDL01, Met04, MFG^{+}09, SLB^{+}00, TG09]. laser [MSL^{+}03]. Last [CWJ09]. Late [KBG^{+}08, NML^{+}05, AS06, BJG^{+}00]. Late-winter [KBG^{+}08]. Lateral [KLF^{+}03, MRK02]. latitude [LBW^{+}08, LN02, NAH^{+}08]. latitudes [LTGG03, RM04, TVSS01]. Latitudinal [GD00a, KFH^{+}09, SMA^{+}06, SRF^{+}01]. Laurentian [JBM^{+}07, MGH08]. lava [LTHW04]. Lavaca [BMU04]. law [HFDO05]. Lawrence [KCSM07, LBG^{+}09, SMJ^{+}06, GSG^{+}05]. Laws [CP03]. layer [AVJ^{+}00, BHVC05, BHVC06, BVWB07, BYF09, Bou04, CS06, GBFJ07, JWRW07, JPK^{+}05, KJH03, LHWE01, LB00, LMMW03, MRK02, PW03, PN06, RMK^{+}06, RHJ02, RHJ04, Sch06a, SSAZ01, SABA02, SR02, SR04, SMS07, SMS09, SSM^{+}09, SI03, TG03]. layer-associated [TG03]. layering [SFL05]. layers [BBCW09, BB04b, DG06, LPW05, NA09, PSJ00, RM04, SGG^{+}08, YM03]. Lazarev [MFG^{+}09]. Lead [PFN^{+}05, BRB^{+}01, HSF^{+}08, HD03b, RRD04]. leading [OHT06]. leaf [AMS^{+}08, KLSW00]. leaving [DKRZ06, RDPM06]. Ledvica [VM06]. Lee [LRS^{+}03, ZVRL03]. legacy [FCJ^{+}06]. leidyi [Ano08b, CSG^{+}06, GLVFH09]. length [ASH^{+}06, RRS06, RS09, WGKE04]. Leptocheirus [SDC00]. Leptodora [MKG04, VKvD05]. Leptolyngbya [STHG07]. Lessons [Fis00, AWH07, HR07, OSDB07]. Lethal [LWU^{+}05, LL05]. Leucine [KJ00, CD03, TS00]. Levantine [MOK^{+}03]. level [AS00a, BGC04a, GCM^{+}09, HB05, HH03, RB02, SSK06, SHL^{+}05, VWPP^{+}06]. levels [BF08b, CMH^{+}08, GCG05, HWL^{+}00, KAM04, LPD^{+}04, MGRE^{+}07, NA09, OHCP02, RP06]. liberating [LMZB08]. Life [CWPL05, HRA07, RRCP03, SD05, SMDG04, VAS02, AR06, BB03, BVB03, BCMR00, HKNG08, JKS04, KKT05, LWS01, MB00, MD04, Sak00]. Life-history [CWPL05, BCMR00]. Ligand [LCPS05]. ligands [AKH04, CMB00, SRM05, SWS^{+}06, TF00, WKSM08]. Light [Ack03, BS02a, BMFS^{+}03, BFD^{+}04, BVN^{+}00, BDK^{+}06a, ESSS01, Fin01, GD01b, Har02, JB00, LMPY00, LKAB04, MBCB08, PE07, PCA^{+}00, RZR02, SSS08, dVKJK07, Aki00, ARO^{+}04, ACHG07, AKP^{+}09, BVN^{+}02, BBAW09, BSJ02, BBG^{+}05b, BB09, BM03, DBGR09, DRM^{+}09, DKKA09, EMIP05, EDF^{+}03, FAM^{+}06, GM03b, GCG05, GMRVFP02, HFR08, HFR09, HG03a, HG03b, HB08, HMNL06, HWL^{+}00, KHKS05, LSK07, LB04a, LG04, Leo08, LKB04, MPGB08, MV00, MS05, NA09, NH06b, PED^{+}09, QMS^{+}07, RH06, RRHPT^{+}06, SMP05, SSC02, SW05b, STK^{+}04, TKM07, TGFPT08, TTG^{+}09, TLrR07, VDK04, ZJC^{+}07, vWSBT08]. Light-dependent [RZR02]. Light-induced [ESSS01, SSS08]. light-limited [Fin01, LB04a]. light-mediated [vWSBT08]. Light-scattering [PE07]. Light-stimulated [dVKJK07]. lightness [LM02]. lignin [WIK08]. Ligurian [NCPd08]. like [HOFD01, MGW^{+}05, Shi04, TLS01, TS06, WV00]. limit [LLC^{+}09, NA09]. limitation [AFL07, BBJ05, BBP^{+}06, BMNSK00, BS03, CC01, CK08, CR06, CCC^{+}03, Cul06, DTM^{+}08, DGV01a, DET04, DSW05, EM08, FAM04, FMG^{+}08, GR01, GD00b, GH00, HMR^{+}07, HB08, HHW^{+}02, JS08, KMC^{+}00, KLL04, KHM04, MACB07, MML^{+}08a, RRCP03, RQB^{+}00, SMCW02, SXLM06, SMC08, SrL05, SSM^{+}04, SSC02, SH05b, SH07, SQTA07, TB00, VD08]. Limitations [SK04, KR08, SAS^{+}03]. Limited [LTOS04, Boe00, Car01, Fin01, GW03, HBF^{+}06, LNB04, LB04a, LNB02, MP00, Pri05, SIRO05, VJ07, VAS02, Zha00]. limiting [BGE04, HM06b, vE02]. Limits [Ben00, Har02, JS03a, MC00, PDM^{+}09]. limnocorral [PW03]. Limnological [CRBF^{+}06]. limnology [AWH07, HR07, OSDB07, Ano01, DIMP08]. limpet [MTC^{+}06]. line [MSL^{+}03]. lineages [FTY^{+}06, MVC^{+}05]. Lines [McA03]. Lingulodinium [LASR09, vDBL05]. link [CHH07, GBA00, HPE01, WLTF03]. Linkages [AFB01, MdG08, MW02]. linked [ABJ^{+}06, MD09, PVD01]. Linking [FL08, NHTF01, WGLT05, WSHW05, dGB02, vWSBT08, ZMJ^{+}07]. links [LHB04]. lipid [DKt08, Has06, SHS03a, WIK08]. lipid-poor [Has06]. lipid-rich [Has06]. Lipids [CCM07, JHW^{+}07, JHGGP06, VMJ08, ZWC^{+}04]. lipophilic [LV01]. liposome [RBMN03]. Liquid [BK01b, BK01c, FH03]. Liss [Pom01]. lithistid [EKdF07]. lithogenic [Ham02]. litter [AMS^{+}08, KLSW00]. Little [WMRK06]. Littoral [BJL01, KB05a, AD00a, BBLB05b, BBLB05a, DAV04, GVT06, HB05, HHS05, Hei08, HHT01, HLP08, LAJ^{+}04, LLB07, LL05, LPB06, WHHS06]. live [LLF^{+}09]. lived [JBvHS08, KH03]. Living [BSMR08, WGC00, BdG02, Dow02, FSM^{+}07, HMJ04, MG08, MWH01, dGB02]. Livingston [Elm02]. load [DDZ03, SRT^{+}03]. loading [Caf01, JPJS06, MCPP06a, PDM^{+}09, WLS00]. loads [HU07]. Lobophora [JM02]. lobster [Ann05, BFCLÁ08, GB09]. Local [ABFS09, HH03, MMKN08, SW05a, VVD^{+}09, AH01a, AM08]. Local- [HH03]. Local-scale [ABFS09]. localized [EBS07]. located [TLS01]. Loch [GJS01]. Long [CWC^{+}06, FCJ^{+}06, GVT^{+}09, INU04, Kim05, LIE06, ME06, PCvRR00, RL07, SMH^{+}06, VW08, VWPP^{+}06, AI01, BK01c, CJWR05, JS03a, LBB^{+}04, MGRE^{+}07, MCB08, MSZ00, RBW09, RC05a, UL05, WMW^{+}08, YYTH07]. Long-term [CWC^{+}06, FCJ^{+}06, GVT^{+}09, INU04, Kim05, LIE06, ME06, PCvRR00, RL07, SMH^{+}06, VW08, VWPP^{+}06, CJWR05, JS03a, LBB^{+}04, MGRE^{+}07, MCB08, MSZ00, RBW09, WMW^{+}08, YYTH07]. longicornis [Møl07]. longimanus [HBM06]. longirostris [SCH07]. longitudinal [LN06]. longwave [OOB09]. look [BHVC06, NT08, OGR08]. loop [HSB00]. López [CR08]. Lophelia [DDL^{+}09]. loss [GRdL^{+}08, HBH^{+}06, JHW^{+}07, JTAM09, KB01, MSZ00, VJV^{+}03]. lost [PFL07]. lotic [Fin04, LW01a, PVD01]. Lough [BLG^{+}07, FL06]. Louisiana [EB01, PDTR06]. Low [BSMR08, LLvW^{+}03, AD05b, AY08, BFC07, BKB02, BGC^{+}05, CW03, FSBB05, GM03b, GBA00, HB05, HU07, HRA07, HR00a, JGSB07, LTGG03, MPGB08, MHLN01, MMLH06, NSH^{+}09, NKJM06, PSH08, RSWH04, RC07, SSK06, SLC^{+}09, TA03, TM05]. low-affinity [JGSB07]. low-chlorophyll [MMLH06]. low-cost [GBA00]. low-head [AD05b]. low-iron [FSBB05]. low-level [HB05, SSK06]. low-nutrient [MPGB08, RSWH04]. low-oxygen [MHLN01]. low-salinity [BKB02, TM05]. lower [HMT^{+}00, KCSM07, LKAB04, LBC06, NIE^{+}00, SBM04, TS01, WKB^{+}04, CF01, LBG^{+}09]. lowland [CKA03, GP06b, RP06, TSA09, THJ^{+}09]. LR [GPAP^{+}04]. Lugano [AHHH^{+}07, HAHS^{+}09, LBB^{+}04, LBM^{+}04]. lumholtzi [HCH00]. lunar [BBAW09]. Lutjanus [BW01]. Luxury [SS01, SS04]. Lymnaea [GGH07]. Lyngbya [CT06]. lysed [KGYU06]. lysis [AD00a, AD02, BW05, EKD^{+}07, LCS02, RvBB02, STHG07]. Lysogeny [WBH03]. lytic [Thi00].

M [ACD10, Caf01, Fou02, Wan02, GSL^{+}07, HBH00, KCCE06]. MAAs [Shi04, WV00]. machine [Boy01]. Macoma [PvAB^{+}03, RHM04]. Macro [CSM^{+}08, BH02]. Macro- [CSM^{+}08]. macro-holoplankton [BH02]. macroalgae [Dal03, KNC^{+}09, LLF^{+}09, LWS01, MZL04, UMY^{+}02]. macroalgal [ERH06, LW02, RTT^{+}02]. macrobenthic [FSK^{+}04]. macrobenthos [LM02]. macrobiological [DAL02]. macroconsumers [SPM04]. Macrocystis [FSB^{+}08]. macrofauna [KNE08, PDR^{+}09, SdB06, TSA09]. macroinvertebrate [CH05a, CGWK^{+}09, HL02]. macroinvertebrates [Hei08, HR03, ML02, MTM^{+}01]. macronutrients [BRS01, WDX01, WHC^{+}06]. macrophyte [BSJ02, IR06, KLSW00, LL05, NA06, NA09, RB03, RKH03, RL07]. macrophytes [BSJM06, RL07, SJW03]. macrotibialis [HGB00]. Macrourid [DPC^{+}08]. macrozoobenthos [LH05]. Magalluf [DAMM^{+}09]. magna [BB05, HW08, MCvEH08, TW09, YW02]. magnetic [ART^{+}06, CEA^{+}09, WdBD^{+}01]. Magnitude [BW00, DKA02]. main [LMSE05, VC05]. Maine [AKG^{+}05, FW07, GGDC04, Has06, MA08]. mainstream [HMT^{+}00]. maintaining [MH04b]. maintains [SMH^{+}02]. maintenance [How00, SBS07, SMS07, SSM^{+}09]. Major [KB08b, PPH03, WHHS06, ACZ00, Bat00, BGC^{+}04b, CD03, KDMC05, KMvD^{+}09, YSM^{+}03, YNCK04, ZHFD06, KB08a]. make [SSAH05]. makeup [ES00]. Malawi [HHT01, VBH^{+}05, CWJ09, Pil04]. Malawian [KK00]. males [Kiø07]. Mallorca [DAMM^{+}09]. management [CBD^{+}06, CK00, Fis00, Par01, Sch06b, Wro01]. manganese [AGL03, CS05, PP04]. manganese-54 [AGL03]. Mangrove [MPJ08, BMD^{+}04, DHWW02, KA06, SMD^{+}04]. mangroves [PVD^{+}07]. manipulation [SPM04]. manipulations [CPCK00, ED05, TMP02]. Mapping [JN08, SGLT08, ASI^{+}02, UMY^{+}02]. Mar [Wan02]. margin [ÁSGMP01, DD09b, EPD09, GTS^{+}09, GSB^{+}09, GvRE^{+}02, HKD01, SGB04, UWA^{+}07, SVPK^{+}09, WIK08]. margins [LSPC02, vODL^{+}09]. Marina [Wan02, PKKB05, DVV^{+}07, DDZ03, EW08, FG06, FHBK06, GBP03, KA06, Reu02]. Marine [BZH00, Can00, KJ01, MZL04, RM04, Aks07, All00, Alv07, AFB01, AD00b, AP00b, AR06, ACZ00, ALRM08, AMP^{+}00, AVA^{+}02, AD07, AAD08, ASP04, BMFS^{+}03, BFDC01, BVN^{+}02, BRT^{+}03, BVH02, BCV^{+}08, BK01a, BNOK^{+}04, BK03a, BRL03, BMBA07, BFBHF04, BBKC03, BMSA08, BMU04, BLSQ07, BW00, BH02, BARS01, Cal01, CL04, CT05, Car01, CSE^{+}08, Ceb02, CMH^{+}08, CFMH01, CW01, CYN05, CF05, CSK^{+}07, CTY06, CDD07b, CBvEK03, CRF01, DSD02, DL00, DRM^{+}09, DdG09b, DGL^{+}04, ES00, EH00, EPS00, FA03, Fin01, FQR^{+}06, FQC^{+}07, FMG^{+}08, GR01, GM04a, GSB00, Goe08, GD00b, GD01a, GM03b, GYBA04, GWL^{+}07, GSNH04, GJB^{+}04, GAPR06, GFJ^{+}02, HY01, HYH04, HKF01, HPAM01, HSN^{+}03, HB03, HM06b, HMNL06, HMT^{+}08, HAES02, IB05, JFH^{+}05, JHW05, JCW08, JGSB07, KCP08, KF01, KKH04]. marine [Kiø00, KIB01, KTMM00, KSM^{+}05, LJCM08, Les01, LN02, LH08, LMBP07, LWU^{+}05, MCV^{+}05, MHB01, MPACG01, MBG^{+}05, MV03, MVBM04, MPRC02, MNM^{+}09, MPdVG07, MJW03, MMW00, NP01, OH00, PWX04, PWP07, PTLJ01, PIF08, PRKHW04, PM08, QRF06, QSH^{+}02, RSS^{+}00, RRBS03, RRS06, RHJ05, RGRPD04, SC07, SRNW06, SIH^{+}06, SD00, SHSH^{+}07, SFR^{+}08, SRG00, SJH06, Smi06, SlBWC^{+}00, SKNK09, SM01, SLRM03, SAE^{+}08, SMA00, SF03a, SF03b, SF03c, SSC02, SHS03a, SH07, SMDG04, SŻKL08, TPFS09, TP03, TWR04, VB00, VHL^{+}04, WMWV04, Wan02, WBP09, WBH03, WM07, WNP04, WWS^{+}00, XW02, ZW06, ZHSJ09, dlM09, vOMSM06b, vdPAJ^{+}06, SG02]. marine-derived [BLSQ07]. marine-dominated [JCW08]. marinus [SHSH^{+}07]. Mark [SG02]. markers [HCH00, VHL^{+}04]. marsh [BCDH09, CWWS00, Cha07, FSK^{+}04, GK03, GHB^{+}05, GVT^{+}09, LN06, NA03, TMA^{+}01, TMG03, TSML07, WC04, Ham01]. marsh-dominated [WC04]. Marshall [AP00a]. marshes [Cra07, EMPI07, PDTR06, SDV^{+}06, THT^{+}09, TNO^{+}08]. Mary [BCBS04]. Mass [BBJ^{+}06, EGP08, FAM04, LNB04, LJ09, RB01b, SC00, WSHN00, BMFS^{+}03, BS04, CZRD^{+}04, CR10, DPFL09, FAC05, FAL^{+}07, FKCF00, GHTJ00, GEP08, HABF03, KKH06, KFW09, LFA^{+}09, NA06, NIE^{+}00, RB01a, RKM^{+}06, RAKE05, SHL^{+}05, SB03, STHG07, STST07, VW08]. mass-specific [BS04]. Mass-transfer [FAM04]. Mass-transfer-limited [LNB04]. Massachusetts [TT05, Too01]. masses [HLGI08, Par01, TLvAH06]. Massive [GFM^{+}01, JKLR08, TWC^{+}09]. mat [APH^{+}03, LAHdBJ05, WdBD^{+}01]. Matano [COK^{+}08]. Mate [CK08, Kiø07, KB05b]. material [Ham02, MBW^{+}01, MHB03, RJS^{+}03, TSMB05]. mating [Goe08, Kiø07]. matrices [BKK^{+}06]. matrix [Shu01, YJMT08]. mats [GFM^{+}01, SIRO05, VH06]. Mattaponi [HB06]. matter [AGL03, AVJ^{+}00, AHM^{+}08, AFB01, ACZ00, AKP^{+}09, BFC05, BFC07, BAG^{+}09, BPOT04, BLWB08, BGV02, BK03a, BT00, BGE04, BMC07, CKPS04, CSWD07, CS06, CDK02, CLG08, CSK^{+}07, DAL02, DD09a, DD09b, DGKA06, DGCMK00, DCC^{+}01, DKRZ06, FSSH03, FW07, FPPH06, GJS01, GBM^{+}04, GCS06, HAA09, HBH^{+}07, HDMB04, HMT^{+}00, HSR^{+}08, HSB^{+}06, HFE07, HLL^{+}06, KS04a, KB06, KKH06, KNN06, KUP^{+}06, LFZ^{+}07, LLB07, LRMV04, LBC06, LBD^{+}09, LNJL09, MYM^{+}05, MH00, MH04a, MVFD04, MSSB06, MBCD04, MBC06, MBW^{+}01, MSC^{+}04, MMCW09, MPD^{+}06, ME07, MTM^{+}01, MSZ00, MLPS09, MPdVG07, NSM09, NCÁSP06, NK05, ORH01, ORFD^{+}09, OZM^{+}01, OOB09, PS06, PBGV04, QSH^{+}02, RSH08, RGAS04, STC03, SMD^{+}04, SHL^{+}05, SDvEB06, SCDR07, SM01, SM05a, SM05b, SBW07, TYK^{+}08, TB00, TMUW02, TB09, TLM04, TNO^{+}08]. matter [VJ07, VW08, SVPK^{+}09, VM06, WNL^{+}08, WIK08, WW03, WHC^{+}06, WW07, WV00, WdMD^{+}00, WTM09, YJMT08, YT09, ZSS^{+}08, ZB00, ZWC^{+}04, vWSBT08, DdG09a]. maturity [ASH^{+}06]. Maud [WSNH06]. maxima [FB03, HB08, PSEL08]. Maximum [CEBMF05, GEJ^{+}02, JK08, ACG07, BVWB07, BMSA08, KHH^{+}03, LL06, CWJ09]. may [JJP04]. mayfly [CF01, MHC05]. McMurdo [DP01, GMRD04, TPM01, Vin00]. meadow [GDM02, SHBM01, ZB00]. means [DBE08, GJRM07]. measure [MPACG01, SRG00]. measured [CEA^{+}09, CW03, FDH^{+}01, HTS^{+}09, HM08, KLC06, MVW^{+}04, MHS^{+}09, RHJ05, WLRB03]. Measurement [AH01a, GGM03, GFK00, AKH04, MdG08, PP07, WdBD^{+}01]. Measurements [TRT03, VM05, ZMK^{+}08, BD04, BRW07, BZH00, DCC^{+}01, ETCP03, GTK^{+}01, GBH^{+}02, GP00, JPK^{+}05, LSPC02, PIKB08, RNR08, RDPM06, SDG^{+}07, SON^{+}03, SSMM00, ZVRL03]. measures [HMD^{+}07]. Measuring [BCMR00, TBH07]. mechanics [RG06, RL01b]. mechanism [JM08, LCLP02, PC04, PF01c, SEL^{+}02]. Mechanisms [HS04, IOG^{+}04, MH04b, MSSW^{+}07, Ben00, CHH07, DDL^{+}09, JCC02, LBSY02, MCF^{+}04, PMK08, ROP^{+}02, Thi00]. mechanistic [CMD^{+}09, Fly03]. MedFlux [RFM07]. media [LSF^{+}08]. mediate [BSG^{+}08, Sar07]. mediated [BM00, BF08b, BBL^{+}03, FNM07, JR07, MDJ^{+}04, NH06b, SH03, TDT01, VZM07, WDVA08, ZHSJ09, vWSBT08]. mediating [CPCLS04]. mediation [BdBS^{+}08, HL06]. Mediterranean [BLM^{+}06, GST^{+}01, SH05a, AD00a, AMS^{+}08, CAP09, DAV04, FNLG00, GDM02, GCM^{+}09, GWL^{+}07, HCK02, KHM04, LMSE05, LB09, LIE06, MBLP01, MOK^{+}03, MWH01, MIN^{+}05, MCB08, MBPVR06, NCPd08, RORR^{+}09, RSH08, RG02, RFM07, SL02, TRT03, TRT04, TS00, SVPK^{+}09, VCKS08, WKW09]. medium [SI03]. medium-sized [SI03]. medusa [THR00]. Meganyctiphanes [TCR03]. meio [PDR^{+}09]. meio- [PDR^{+}09]. Meiobenthos [NHN^{+}05]. Meiofauna [TSA09, BDC00, NKE08, SSAH05]. meiofaunal [RM06]. melt [BGHA04, SKC^{+}01]. meltwater [GMRD04]. Membership [NDCP08]. membrane [AS02, AGMP04]. Memoriam [Leh00a]. Memphremagog [RKH03]. meneghiana [AGL03]. mercaptopropionic [HMB06]. Mercenaria [TG09]. Mercury [SDM^{+}09, BMU04, CBG01, CGL03, CG03, CGL^{+}04, CMCB^{+}09, EBS07, FCJ^{+}06, HLCG07, LF09, PAF^{+}04, RLC^{+}01, SAH^{+}04, ZWHP04]. Meristematic [GBP03]. meromictic [AHHH^{+}07, AVJ^{+}00, HAHS^{+}09, PDM^{+}09]. meroplankton [KBL08]. Mertensia [GLVFH09]. mesnili [BRE02]. meso [HSCF06, RvAV^{+}06, TvAVH06]. meso- [HSCF06, RvAV^{+}06, TvAVH06]. mesocosm [EGPA02, EZA^{+}05, HTS^{+}03, KLS^{+}06, KLS^{+}07, LGJ02, LTT07, MJW03, TKAW03, WdMD^{+}00]. mesohaline [ZC02]. mesopelagic [CGH^{+}04, SRZ04]. Mesoscale [Ano07, MCM06, ZZPP01, AWH07, BST^{+}05b, Coc01, GAM^{+}01, GSL^{+}07, HR07, MBL^{+}01, OSDB07]. mesotrophic [JBH^{+}04, MLPS09, MS05, RS01]. Mesozooplankton [Cal01, CP03, Law03, KRZ06, TF04, VVSM09, WLTF03]. Meta [WST06]. Meta-analyses [WST06]. Metabolic [FSSH03, HDMB04, AD05a, Ano07, BFC05, CR08, DdG09b, GAM^{+}01, LU08, MCM06, PBC02, SAS08]. Metabolism [SJRB09, BCC^{+}04, BAG^{+}09, BMT^{+}04, BOC01, CKPS04, DS07b, DAV04, EF05, FVD01, Fin01, GK03, HT03, HBCK03, JN08, LRMV04, MJR08, MdG08, MSLH02, MVBM04, RG07, SP04, SSJ07, SHA08, SMG06, TKM07, TBH07, YWR^{+}07]. metabolites [PSH^{+}07]. metacommunities [LN04]. Metagenomic [HPT^{+}09]. metal [CBG01, FDH^{+}01, HWYL07, MPH08, SSM^{+}04, TSSWK^{+}06, TF00, YW02, Zha00, ZW06]. metalimnion [SI01]. metals [CSK^{+}00, CMQ^{+}00, CLS05, HA01, JS08, LRF07, MLSS05, RB01b, TPFS09, TM05, TF04, WG01, XW02, ZW06]. metamorphosis [HH05, MBDL01, PDS05, TG09]. metaphyton [KR08]. metapopulation [Ano08b, CSG^{+}06]. metazoan [LJ01, SSAH05, TCD^{+}03]. metazoans [BVB03]. metazooplankton [DGKA06]. Meteorological [FUG02]. meters [SBR^{+}06]. Methane [Ost03, SGB04, ACG07, BGJ01a, BGJ01b, CCCC07, CCC09, HPBD01, KTG^{+}06, KHP^{+}06, LM02, MS05, NEC03, SDG^{+}07, TKBJ05, THJ^{+}09, WSNH06, WBD02]. methane-derived [THJ^{+}09]. methane-supported [WBD02]. methanogenic [CCCC07, SEC08]. Methanotrophic [KHP^{+}06]. method [AG03a, ASAC04, EWE02, GMMW03, HM03a, KJJ07, MMW00, SD00, VD08, WBR^{+}03a, vE02]. methodology [Par01, TSH^{+}06]. methods [Ben00, THA01, YPBL07]. methyl [HBN^{+}00]. Methylmercury [MA08, CAP09, SKR01, WMRK06, ZWHP04]. Mexican [HKD01, SGB04]. Mexico [BW01, BFCLÁ08, CG09, DS07a, GHM^{+}09, HVP^{+}07, KL00, MKK04, MKVK05, MH04b, McM00, MBH^{+}06a, MBH^{+}06b, SBC^{+}08, SRT^{+}03, SPR02, SBM04, vHRB^{+}09]. Meyers [Caf01]. micelles [VMJ08]. Michigan [HTRM^{+}07, JBvHS08, KBG^{+}08, Mor06, TTSC04, VMH^{+}01]. micro [WLTF03]. micro- [WLTF03]. microaggregate [LLDS06]. microalgae [Dal03, HMNL06, MMW00, Nil03, PVD01]. microalgal [BDC00, HBH^{+}06, KDO09, SLRM03]. microanalysis [HSN^{+}03]. Microbial [AGDH09, BdBS^{+}08, CBRC01, CS06, FH09, GLP^{+}08, GP01, KLSW00, LFS^{+}04, NMK03, NCÁSP06, SWSJ02, APH^{+}03, BLM^{+}06, CGH^{+}04, DGH^{+}07, DGD04, FSSH03, GHM^{+}09, GFM^{+}01, GVT^{+}09, HSB00, HTS^{+}03, HPT^{+}09, Kem00, LAHdBJ05, PBM^{+}07, RW02, SEC08, SAPA^{+}02, SHA08, TRT03, TS08, TPSF01, VJ07, VW08, VMC^{+}07, VH06, WBD02, WdBD^{+}01, ZHSJ09, dBSN^{+}06, Pom01]. Microbially [vWBT09]. microbiological [MBC06]. microcosm [MAC^{+}03]. microcrustacean [LLB07]. microcystin [JJT07, Lür03, WH07, GPAP^{+}04]. microcystin-free [Lür03]. microcystin-LR [GPAP^{+}04]. microcystin-producing [Lür03]. microcystin-treated [WH07]. microcystins [RCD^{+}05]. Microcystis [JS08, WŠN^{+}07, GPAP^{+}04, HTWG08, KJ00, LKM03, Lür03, RSWH04, RB02, SWH^{+}05, SEL^{+}02]. microenvironment [Fab06, JKLR08, SdB06]. microenvironments [AVA^{+}02, Plo08]. microlayer [KLAF04, RSH08]. micronekton [BBAW09]. micronutrient [Alt01]. micronutrients [AP00b]. microorganisms [CMCB^{+}09, SŻKL08]. microparasite [BRE02]. microparasites [DLDE04]. microphytobenthic [OHUB00, SMH^{+}02, UPC^{+}05, WWS^{+}00]. microphytobenthos [CR06, MBB^{+}00, NCO02, SLL^{+}04]. microplankton [SWH^{+}03, SWS^{+}03]. microprobe [GGP01]. microsatellite [JL08]. Microsatellites [Reu02]. Microscale [MNSV03, MMW06, GSB^{+}09]. Microscopic [SIRO05]. microscopy [SF05]. microtidal [SLL^{+}04]. Microzooplankton [HZG07, JRS03, SMO07, CL04, CLL^{+}09, EKD^{+}07, OLCT08, PSHM04]. micrum [JRS03]. Mid [Dix00, MY05]. Mid-Atlantic [Dix00, MY05]. midcontinent [BHV04]. Middle [GJT^{+}08]. midge [MHLN01, OHT06]. midges [DST08]. midsummer [ZXN^{+}08]. midterm [Lon00]. midwater [THR00, TIH^{+}01]. midwestern [RH01]. might [KHKS05]. migrate [HKF01]. migration [BF08a, BKB02, BBAW09, BRM^{+}07, CTWFP06, DJO00, De 02, DKKA09, HNB^{+}07, HKF01, HFC^{+}04, HTWG08, HLH09, IJA^{+}04, KKT^{+}07, KJH03, LLS08, LBC^{+}09]. migrations [GZL02]. migratory [EDF^{+}03, UPC^{+}05]. mile [Boy01]. Millennial [RLM^{+}04]. Millennial-scale [RLM^{+}04]. millimolar [EMWL06]. Mineral [SrL05, BS04, CG01, HTS^{+}03, SWF04, SBW07]. mineral-rich [SBW07]. Mineralization [MMDW09, CWG^{+}07, GBFJ07, JMBJ01, KNN06, MMW06, OB04, VSSTS00, WBP^{+}06, dBWF^{+}05]. minerals [DFBA07]. minerogenic [PE07, PED^{+}09]. minerotrophic [KB07]. Mingulay [DDL^{+}09]. minimal [BCDH09]. minimum [ESA06, FCGC^{+}09, HLW^{+}07, JHW^{+}07, WGC00, JRS03]. mining [KT02]. Minireview [RL01b]. Minnesota [EBS07, SDJT03, SC08]. minutum [MWM^{+}04, SCP08]. Mirror [WBR^{+}03a]. mirrored [JS03b]. misinterpreted [NH06a]. Mississippi [DCC^{+}01, DS07a, SRT^{+}03, SBM04, SQTA07]. MIT9313 [SHSH^{+}07]. mitochondrial [CHW^{+}01]. mixed [Bou04, CD02, HMS^{+}04, KM02, KMMS00, KL01, LHWE01, QYEA06, SR02, SR04, TG03, UPC^{+}05]. mixed-layer [Bou04, SR02, SR04]. mixed-species [UPC^{+}05]. Mixing [RS09, SI01, AELS04, BDK^{+}06a, BLHC08, CHB08, D’A08, FLT02b, FGM^{+}03, GGIAY08, HBS05, HAHS^{+}09, HH04, JDS08, Kra05, KBD^{+}00, LAG^{+}03, LPW05, MFK^{+}09, MCJF09, MVA^{+}09, OIH^{+}03, PSLO07, PF01c, PDB03, RKWG00, Sal05, SVC^{+}07, STG^{+}07, SCM08, SSM^{+}09, UL05, WdMD^{+}00]. Mixotrophic [KHKS05, Leg01, MBC^{+}05, MBCB08, NP01, UMASG07]. mixotrophs [KHKS05]. mixotrophy [BWK^{+}07, PSHM04]. mixtures [Lür03]. ML [RCE^{+}02]. Mn [MJW03, SWCL01, SH00]. Mnemiopsis [Ano08b, CSG^{+}06, GLVFH09]. Mobile [CF01]. mobilization [Jan03, KB05a]. mobilized [TMG03]. modal [GP03]. mode [RHM04, VCCS05, VRC07]. model [CMD^{+}09, CCKP02, De 02, FB03, Fol05, FSK^{+}04, GRCS06, GHTJ00, GBFJ07, HBI01, HFC^{+}04, JPJS06, KCSM07, LNB04, MCPP06a, MMCW09, Mor03, RAKE05, RRD04, RMS^{+}08, SRNW06, SCARM00, SKM05, Shu01, SY07, WW02, ZB03a, Zim03, CR10]. Modeling [CZRD^{+}04, HISW00, KSM06, KBD^{+}00, KDS03, LIHS03, LTGG03, LA09, LSF^{+}08, MNA^{+}09, MGH08, NSBA01, PLG^{+}02, VMS^{+}04, vOMSM06a, BD04, BK07, CLF^{+}08, EMK08, HKNG08, JGE^{+}09, JPK^{+}05, KTAL04, MM05, MKV01, MC06, PED^{+}09, PHMR^{+}07, SDG^{+}07, SB03, SCS^{+}08, SF04, TVSS01, VCCS05]. Models [HBS04, ASH^{+}06, Chr05, FMG01, Fly03, McA03, MFF07, MBC01, PGM^{+}09, RFW08, SA09, STST07]. moderate [NA09]. Modern [Spe00, Cul06, HLD^{+}03]. modification [GRR^{+}07]. modified [DGKA06, HAG08]. MODIS [LAG^{+}03]. modulates [vdPAJ^{+}06]. modulating [MBCB08]. Modulation [GDK03, BRP^{+}06, STG^{+}07, UM05]. Molecular [Cur00, Dix00, FTC^{+}06, FTY^{+}06, GLVFH09, KB06, PCD^{+}08, SHL^{+}05, TFNT07, WJMD07, WBD02, BRL03, DB07, DHP03, DdG09a, EB01, HYH04, HSR^{+}08, RA06, TIH02, ZWC^{+}04]. Molecular-level [SHL^{+}05]. mollusc [LWU^{+}05]. molluscs [WSNM00]. mollusk [ASS03]. molting [TCR03]. molybdenum [TCL^{+}09, TWR04]. monasteriata [DOE^{+}06]. Monitoring [BTB^{+}08, CK00, CFO^{+}00]. Mono [BH07, CJ08, CH01, MCJF09, RCE^{+}02]. Monogononta [SS03]. monomethyl [CGL^{+}04, CG03]. monomictic [KNN06]. Monopylephorus [GHS05]. monoxide [XBD^{+}09, ZXN^{+}08]. Monsoon [RNR08]. monsoonal [Goe02b]. montane [CSM^{+}08, FVD01]. Montastraea [LMPY00, QYEA06, RRHPT^{+}06, SSL05, WCGBG07, ZJ03]. Monterey [BW99, BJBA03, MCPP06b, NDCP06]. Montipora [DOE^{+}06]. moonlight [GF02]. moraine [SCY00]. moraine-dammed [SCY00]. mordax [BCBS04]. Moreton [EM02]. morhua [KBD^{+}00]. Morone [BFS02]. Morphological [GST^{+}01, CT06, Dix00, DJT06, SJR^{+}04, TIH02, VHL^{+}04]. morphology [BD07, CMC^{+}09, FHB^{+}09, LH06, MH07, RKM^{+}06, WST06, WLP^{+}07, ZB03a]. morphometric [MJ03]. morphometry [BCC^{+}04, DACG01]. morphospecies [PCDB09]. morphotypes [Has06]. morphs [MH04b]. Mortality [EWE02, ACHG07, AAD08, BKAE01, BVWB07, DAMM^{+}09, EAO^{+}02, GRB02, GJB^{+}04, GWR^{+}09, HR03, ODR^{+}08, WBH03]. Mosby [dBSN^{+}06]. most [KKH04]. motile [SBZ^{+}09]. Motility [KB05b]. Motion [KBBJ00, AIK04, FB05, PSS00, SHS01]. motions [OH08, SIK07]. Mountain [EPP^{+}08, AMS^{+}08, AB07, CMSVA02, HBL02, HBAK09, LVC^{+}00, LLK05, MSVAC04, MBPVR06, NTM^{+}05, PFN^{+}05, VM06, BMC07, ML02]. Mountains [BSS^{+}01, TSMB05]. movement [SLB^{+}00]. mtDNA [DPK^{+}00]. mucilage [BZH00]. Mud [dBSN^{+}06, BHVC05, BHVC06, DD09a, Sch06a]. muddy [PGP01]. mudflat [PDR^{+}09]. muelleri [LG04]. Mulholland [Sta01]. Multi [AO09, MCL06]. Multi-decadal [AO09]. multi-source [MCL06]. multiannual [PGM^{+}09]. multibasin [RM09b]. multicomponent [FSK^{+}04]. Multidecadal [GDB^{+}09]. multinutrient [SY07]. multiphasic [LLC^{+}09]. Multiple [CCKC05, EMIP05, KLL04, KRM09, MBCD04, NDT07, SGPK07, SJW03]. multiple-isotope [MBCD04]. multiproxy [WIK08]. Multiscale [BBLB05b, BBLB05a, KR08]. multiseries [MHRW02]. Multispectral [DCC^{+}01, MSL^{+}03, WR03]. multivariate [FBW^{+}08, RBW09]. Multiyear [CDK02, New01]. Murphy [AD05b]. Muscle [KK00]. mussel [ALH01, BFC05, BFC07, BMF^{+}06, BFML05, BFS^{+}04, BLHC08, CMD^{+}09, Dix00, NPE08, PWX04, SWH^{+}05, SGPK07, SGPM09, WBvE02, WvE04, WSBL06]. mussels [BSL^{+}07, BDNF01, Car02, DBS^{+}04, HBG00, RSWH04, WLTF03]. mutual [HR00a]. mutualistic [BAPMMB01]. Mya [HM02]. mycosporine [MGW^{+}05, Shi04, TLS01, TS06, WV00]. mycosporine-like [MGW^{+}05, Shi04, TLS01, TS06, WV00]. myctophid [CCM07]. Myriophyllum [LL05]. mysid [TGF00]. Mysidium [TGF00]. mysids [GL07]. Mysis [BRM^{+}07, BCMR00, HLNM^{+}03]. Mystic [KNH00]. myths [TF00]. mytilid [BFML05]. Mytilidae [Dix00]. Mytilus [SGPM09, WSBL06].

N [Dow02, FMG01, HVP^{+}07, JMBJ01, SL01, WGKE04, AS00a, AD01, Ber01, BBP^{+}06, lLCChTO08, CWC^{+}07, FPZ09, FMG^{+}08, GEJ^{+}02, GRB09, GGH07, GHB^{+}05, GVT^{+}09, GHN^{+}02, HB05, HTS^{+}09, HFZ^{+}07, IR06, LBM^{+}04, LG04, MH07, MBLP01, MM05, MBCD04, MVW^{+}04, MB05, MHS^{+}09, PLW^{+}08, PCD^{+}08, RRSC01, RHM04, RG04, SRG00, SHBM01, SSL05, TMYH02, TMS^{+}00, TB00, UMY^{+}02, VR01, VEMM07]. N-irradiance [FMG01]. N-labeling [VEMM07]. N-nitrogen [GVT^{+}09]. N-tracer [HB05, TMS^{+}00]. Nada [KKKT07]. Namaqua [PN06]. Nancy [Dow02]. nano [CSSNA02]. nano-detrital [CSSNA02]. nanoflagellate [BMCvE06]. nanoflagellates [TRT04, ZHSJ09]. nanomolar [MTV^{+}02]. nanoplankton [TCCG05, WDVA08]. Narragansett [Ano08b, CSG^{+}06]. Natal [WSC^{+}06, RPR^{+}08, WSC^{+}05]. National [EPP^{+}08, HT03, How00]. native [BKB02, FS09b, GSA^{+}05, GFJ^{+}02]. Natural [ASH^{+}06, BEW^{+}00, KFR^{+}01, KMHN07, TSH^{+}06, VMS09, AP00a, AS02, AW05, BKK^{+}06, CBRC01, CW01, CLC02, CDBC00, CCC^{+}03, CS05, ET08, EGP08, FGM^{+}03, HY01, HVP^{+}07, KKH04, KLS^{+}06, KLS^{+}07, KMMS00, KH03, LLA05, MV00, MHB01, MGC^{+}07, PTK^{+}07, PBMSG04, RRSC01, RB01a, RRH^{+}03, RFM07, SD05, SLHK02, STC03, SSS02, SSS08, TGdB^{+}01, TF00, WdMD^{+}00, ZLL^{+}02]. naturally [GFM^{+}01, New01]. nature [CLS05, MHC05]. Neagh [BLG^{+}07, FL06]. neap [STG^{+}07, dSYL^{+}08]. near [BS02a, CV05, CBG01, DRM^{+}09, FLT02b, JS03b, RDPM06, SMGP06, TAB^{+}00, ZXN^{+}08]. near-bottom [CV05]. near-infrared [BS02a]. near-surface [JS03b]. Nearshore [TvG02, Has06, MC04, PWN07]. necessarily [GM03a]. necessary [TF00]. need [Fly03]. negative [GNU00]. negatively [CRC^{+}03, LFPH07, OIH^{+}03]. neglected [HPE01]. Neither [MSVAC04]. Nelson [MWE^{+}04]. Nematocysts [BH02]. Nematoda [DVV^{+}07]. nematodes [HMJ04]. Neomysis [SJ08]. Nepal [SCY00]. Ness [GJS01]. Net [ÁSGMP01, NEE08, SlBWC^{+}00, ÁÁS07, BW99, CPCK00, HHY^{+}05, HHB05, IJA^{+}04, KFR^{+}01, LAJ^{+}04, Lee01, MHK^{+}07, RNC03, RBC00]. Netherlands [SMH^{+}06]. networks [AB07]. Neuse [VWPP^{+}06]. neutral [HY01, RA06]. Nevada [SMC03]. Newfoundland [BCBS04]. Newly [AD07]. Niagara [HDFM08]. niche [BGR^{+}09, CMJ07, Ham05, KK00, MBCB08]. Nickel [EM08]. nifH [FPZ09, MPZ08]. nighttime [RKNA09]. nilotica [Har01]. Niño [Dav00, NBW^{+}02]. Nitinat [PBM^{+}07]. nitracline [AOR07]. Nitrate [GMAFP03, HTS^{+}09, LMT05, MP00, MHS^{+}09, WJA^{+}07, BSC^{+}09, BHV04, BS03, CWWS00, lLCChTO08, DBE08, FGM^{+}03, FSB^{+}08, FSBB05, GMA^{+}06, GSNH04, GSLT08, HWL^{+}00, JN08, KD02, LWP^{+}07, LG04, LKAB04, MMLH06, MVW^{+}04, SCP08, STG^{+}07, TB00, TvdWdB04, TMA^{+}01, WKFP06, YTVR07, dlM09]. Nitrate-depleted [WJA^{+}07]. nitrate-replete [dlM09]. nitrate-rich [BHV04]. Nitrification [CJ08, JR07, AJ01, ATJH08, CRJ08, KD02, Nil03, SL00]. nitrification-denitrification [AJ01, Nil03]. Nitrifiers [LL06]. nitrite [LL06]. Nitrogen [Alt01, BSC^{+}09, BFQF^{+}07, EVW07, GMA^{+}06, GTS^{+}09, GSLT08, GHB^{+}05, HAL^{+}09, HM06b, KC08, LWP^{+}07, MWM^{+}04, MML^{+}08a, MCC02, MBH^{+}06a, MBH^{+}06b, NFS^{+}07, SNFT07, StLHB^{+}07, SHBM01, VEMM07, WKFP06, AS00a, ARV^{+}07, ÁÁS07, AP00b, Ano04, ATJH08, ASAC04, AGP07, AHR^{+}01, BMD06, BRL03, BBJ05, BFCS^{+}01, BCDH09, BBP^{+}06, BCMR00, BW99, BW00, BM03, BLG^{+}07, CH05a, CPHM04, CBA01, CM03, CBK^{+}08, CCH02, DBS^{+}04, DGH^{+}07, DP04, DHWW02, Dit04, DNS06, DS07a, DBTK02, DNB^{+}04, EGPA02, EM02, EF05, FF04, FGM^{+}03, FXL04, GRdL^{+}08, GSW03, GD00b, GD01a, GFM^{+}01, GW03, GSNH04, GGS06, GKJ04, GVT^{+}09, GP01, GH00, GFJ^{+}02, HAA09, HT03, HBAK09, HPLG09, HBM08, HCK02, HHT01, HB06, HCD^{+}09, JHW^{+}07, JR07, KB08b, KTLS03, KNN06, KFH^{+}09, KFW09, KCV06, KSWC^{+}03, LNB04, LHWE01, LBEP06]. nitrogen [LTGG03, LBB^{+}04, LBM^{+}04, LPAQ^{+}05, LNB02, MPJ08, MW06a, MTC^{+}06, MAM07, MCL06, MFF07, MPZ08, MF01, MF02, MPRC02, MML^{+}08b, MBPVR06, MC01, MGL02, MB05, NCO02, NDT07, NGS^{+}07, OKO^{+}01, PKKB05, PGL06, PSS^{+}06, PWH^{+}04, RCAK05, RlBW02, SLE04, SRT^{+}03, SBL06, SDJT03, SRG00, Smi06, SlBWC^{+}00, SWML08, SAE^{+}08, SdB06, SSC02, SLL^{+}04, TMS^{+}00, TCP^{+}03, TSMB05, TB09, TN09, TR03, TMA03, UMY^{+}02, VJ07, VMS^{+}04, VMSB08, VM06, WS04, WPD^{+}01, WLS00, ZMJ^{+}07, Caf01, KB08a]. Nitrogen-fixation [BFQF^{+}07]. nitrogen-fixing [ARV^{+}07, CPHM04, CBK^{+}08]. nitrogen-limited [LNB02, VJ07]. nitrogen-phytoplankton-zooplankton [MFF07]. nitrogen-rich [DNB^{+}04]. nitrogen-sufficient [LPAQ^{+}05]. nitrogenase [ZMJ^{+}07]. nitrogenous [RlBW01]. nitrous [DNS06, FCGC^{+}09, NDT07, TMB07]. nitzschia [MHRW02, MMLH06, MH07, OLCT08, PD02, SGG^{+}08, TWC^{+}09, THL^{+}09]. NMR [HRS04]. NO [HWL^{+}00, LLC^{+}09, LV01, PvE00, Alt01]. Noctiluca [DCMS03]. Nocturnal [BBAW09, HG03a, HG03b, KJH03]. nodosa [BMT^{+}04]. Nodularia [KSEÖ^{+}02, Plo08]. nodule [KCCE06]. Non [MBLP01, ARV^{+}07, PLW^{+}08, Chr05]. non-nitrogen-fixing [ARV^{+}07]. Non-Redfield [MBLP01, Chr05]. non-Redfieldian [PLW^{+}08]. nonaxenic [SBL06]. Nondissociative [SRNW06]. nonhumic [OH00]. nonhydrostatic [dlFSIN08]. Nonionella [GSFB02]. Nonlinear [WK09, Cul06, MCJF09, dlFSIN08]. nonlocal [FAL^{+}07]. nonmarine [RSS^{+}00]. nonmotile [SBZ^{+}09]. nonnitrogenous [RlBW01]. nonpoint [Mul01, RB01b]. nonprotein [TSKA05]. Nonrandom [KS04b]. nonsinking [PDB03]. nonspherical [GD01b]. nor [MSVAC04]. North [ÁSGMP01, BK03b, DWA^{+}02, JHW05, FSM^{+}07, RYS^{+}02, SMC08, StLHB^{+}07, SKC^{+}01, WM06, ÁÁS07, AWE07, BAG^{+}09, BNK02, BST^{+}08, BTS^{+}06, BRP^{+}06, CDK02, CLAK08, DTM^{+}08, DJG^{+}06, DBTK02, DPC^{+}08, DKA02, GA00, GA01, GAM^{+}01, GP03, HPLG09, HDMB04, HSB^{+}06, HB08, KBL08, KFH^{+}09, KLAF04, LBEP06, LKAB04, LES^{+}05, LD01, LCC^{+}09, LMBP07, LHB04, MKK04, MH04b, MGRE^{+}07, MML^{+}08a, MIN^{+}05, MCC02, MML^{+}08b, NKS^{+}02, NFS^{+}07, NEE08, PDS07, PLW^{+}08, RLP^{+}06, RvAV^{+}06, RBJ^{+}06, SFB^{+}09, SAPA^{+}02, SWSJ02, SKBC01, SMC08, TvAVH06, TLvAH06, TSSWK^{+}06, TSDK^{+}09, TN09, UM05, VWPP^{+}06, VDK04, VD08, VRS09, Wey04, WMW^{+}08, ZMJ^{+}07, dCCL09, dBWF^{+}05]. north-south [StLHB^{+}07]. north-temperate [RYS^{+}02]. northeast [ASGA^{+}07, BMD^{+}04, BGC^{+}04b, BST^{+}05b, DAA^{+}01, EJL^{+}06, HCC^{+}02, HBM08, Met04, OLCT08, SBR^{+}06, SRKK08, YSO02, vODL^{+}09, ÁS07, BSSL03]. northeastern [CLP^{+}03, GvRE^{+}02, GSL^{+}07, LCPS05, PWH^{+}04]. Northern [BLG^{+}07, ARK^{+}00, BW01, BLWB08, BZH00, BDSB07, CG09, Chi00, FLWP09, GLVFH09, HP06, KJMJ03, KAW05, LSK^{+}03, LPAQ^{+}05, PGL06, PLMPN08, SPR02, TCR03, TDJ^{+}06, WJA^{+}07, MBR08]. northwest [GSG^{+}05, MKVK05, ODR^{+}08, SGT^{+}06, VCKS08, TS00]. northwestern [Aki00, CGH^{+}04, CFG07, DAV04, FNLG00, GWL^{+}07, LMSE05, LB09, LIE06, MBLP01, MIN^{+}05, MCB08, NKS^{+}02, NCPd08, TRT03, TRT04, SVPK^{+}09]. norvegica [GGP01, TCR03]. Norwegian [LFS^{+}04, NML^{+}05]. Nostoc [SJRB09]. note [MF04]. Notonectidae [HGB00]. Novel [DNB^{+}04, HRA07, JFH^{+}05, Par01, SD00, SEL^{+}02]. noxious [RSWH04]. NPZ [MFF07]. ntcA [LPAQ^{+}05]. nuclear [ART^{+}06, GGP01, WdBD^{+}01]. numerical [JPK^{+}05, MCL06]. nurseries [VNH^{+}08]. nursery [BBP^{+}06, BLSQ07, SFB^{+}09]. Nutrient [ADA00a, ADA00b, DTM^{+}08, DC05, HRGM07, HSB^{+}04, JBH^{+}04, KHM04, LB04a, MACB07, RRCP03, RH06, SBC^{+}08, SK05, SSV^{+}04, Sos00, Wur07, ZHL06, AD05b, AS06, ABFS09, AMS^{+}08, AB07, BBJ05, BGE04, BGC^{+}05, BRP^{+}06, BR00, CV05, CBRC01, CPH06, CBD^{+}06, CPCK00, CSK^{+}07, CWC^{+}07, Cra07, CJWR05, DD09a, Dal03, DTL^{+}05, DSW05, DACG01, EVW07, ED05, EGP08, FAM04, FAC05, GDM02, GGP01, GP06b, GH00, HMSI06, HB05, HMS^{+}07, HBL02, HZG07, HU07, HPV^{+}00, HK01, HH02, HTRM^{+}07, HM06b, JCC02, KHKS05, KM02, KLL04, KFW09, LSM03, LSK^{+}03, LKB04, LLF^{+}09, LZ09, MSGK06, MPGB08, MSAC06, MW06a, MSVS^{+}07, MPA^{+}07, MSSW^{+}07, MGRE^{+}07, MIDH05, MJW03, MLSS05, OAA^{+}06, PDM^{+}09, PCvRR00, PCA^{+}00, PSJ00, RSWH04, RGAS04, RTT^{+}02, SMGP06, SMC03, SHGM06, SWSJ02, SH06]. nutrient [SCS^{+}08, SL01, TMK^{+}07, VD08, WSC^{+}08, WLS00, WWS^{+}00, YS06, dSYL^{+}08]. Nutrient- [LB04a]. nutrient-color [WSC^{+}08]. nutrient-enriched [BRP^{+}06, dSYL^{+}08]. nutrient-flow [HMSI06]. nutrient-rich [JCC02, PCA^{+}00]. nutrient-saturated [SHGM06]. nutrient-stimulated [MLSS05]. Nutrients [TMUW02, BS02b, BGC^{+}04b, BGC^{+}05, BLSQ07, BDK^{+}06b, CB04, CG01, CWC^{+}06, GRB02, HTS^{+}03, JDS08, JM02, MAC^{+}03, ML05, PLW^{+}08, QMS^{+}07, SAH^{+}04, SMH^{+}06, ZGP^{+}00]. nutrified [DNS06]. nutrition [AP00b, HABF03, LM02, MKS08]. Nutritional [MCvEH08, MSJMN02, Boe00, BH02, CWPL05]. NW [AD00a, RFM07].

O [ORH01, TGdB^{+}01, BD04, DJB^{+}07, GTK^{+}01, GGRJ03, GBFJ07, GSB^{+}09, PCD^{+}08, SRL^{+}09, WG04, YPBL07]. O2 [SAS08]. O’ahu [GGM03, Par01]. obligatory [HLNM^{+}03]. obliquus [AGL03, Lür03]. Observation [VCCS05]. observational [WRG^{+}05]. Observations [BD04, BST^{+}08, CRC^{+}03, CSP^{+}01, FLT02b, HFC^{+}04, LAN08, MSSW^{+}07, PNPA04, SSM^{+}09, FT03, GJT^{+}08, HDFM08, HWL^{+}00, Mor03, PSEL08, UM05, WSNM00, WCGBG07]. observed [BF08a, BMWW08, LTB09, MJR08]. observing [CLL^{+}03]. obsoleta [FMS04]. Occurrence [SMP^{+}01, Ber01, GLVFH09]. occurring [BHL07]. OCean [BMH^{+}08, DZL^{+}03, ET08, GRV^{+}09, KTLS03, MJR08, ASP04, BMFS^{+}03, BF08a, BNC04, BL07, BST^{+}08, BB09, CLF^{+}08, Chr05, Cul06, Cur00, DH01, DR06, FH03, GTS^{+}09, GSB^{+}09, Goe02a, GJRM07, GSJ^{+}09, HBI01, HD03a, HM08, HP06, HFZ^{+}07, KB08a, KB08b, KLAF04, LB04a, LTGG03, Lon00, LVB08, MH04a, MCR^{+}07, MAM07, MWE^{+}03, MWE^{+}04, MC04, NDP^{+}07, OV03, PPH03, PBGJ09, RG08, SBC^{+}08, SXLM06, SRKK08, SVB^{+}08, SW05b, TvdWdB04, WHC^{+}06, WKW09, WCGBG07, WWWY05, ACW^{+}03, AD05a, ASGA^{+}07, ÁS07, AGDH09, AWLH04, BGC^{+}04b, BMH^{+}08, BSSL03, BJD05, BDK^{+}08, BAP^{+}05, CPL^{+}04, CBK^{+}08, CRJ08, Coc01, CCM07, CCC^{+}03, CW02, DB07, DMEPA04, DBTK02, DAV^{+}05, DKA02, ET08, FSM^{+}07, HDMB04, HPT^{+}09, HPL^{+}06, KSM06, KED^{+}07, KMvD^{+}09, KLAF04, MBL^{+}01, MKVK05, MHK^{+}07, MH04b, Met04]. Ocean [MML^{+}08b, MSPAE06, NKS^{+}02, NFS^{+}07, ORH01, OBSD04, ORFD^{+}09, PLW^{+}08, SIH^{+}06, SAS^{+}03, SF05, SRF^{+}01, SR07, StLHB^{+}07, SLS^{+}02, TPSF01, TGdB^{+}01, TvdWdB04, TM02, TPG^{+}08, TSSWK^{+}06, TSDK^{+}09, TMH^{+}06, TBF04, VGP^{+}08, WAGH09, YT09, ZJC^{+}07, dCCL09, vdSvdSG^{+}00]. oceanic [AGMP04, AWH07, BBL^{+}03, HR07, HPT^{+}09, LA06, MAC^{+}06, MMLH06, OSDB07, OCSB05, Par01, PQP05, PWN07, SH00, VM05, YT03]. oceanica [BMD06, DAMM^{+}09, Fol05, GDM02, JGSB07, LJCM08, MP00]. oceanographic [BPC06, BBG05a, HVH00, KLG09, MMKN08, OS01b, Rit09]. oceanography [AGDH09, NAH^{+}08, Wro01, Ano01, DIMP08]. oceans [GH00, HM09, Pom01]. Ochromonas [BWK^{+}07]. octocoral [JL08]. Odor [HMJ04, JWRW07, WW01, WRD01]. odor-tracking [JWRW07]. odyssey [Wan02]. off [Agr05, DCMS03, EPD09, LJ09, LNJL09, NSH^{+}09, PSEL08, TvG02, TDJ^{+}06, THH02, TWC^{+}09, WHC^{+}06, YSO02]. Offatts [HLCG07]. offshore [ÁS07, FDP^{+}08, PPL^{+}02, SHBM01]. Ohio [HBM08]. Ohrid [MSVS^{+}07]. Oikopleura [AK00, LUHS04, TFNT07]. oil [CCM07]. Oithona [Kiø07, SCB03]. old [AGKK08]. oligochaete [GHS05]. Oligopeptide [RES^{+}08]. oligotrophic [BOC01, CPV00, CBD^{+}06, Chr05, CRJ08, CCCC07, DNS06, DBTK02, GGH07, HB05, HDWDD09, LPAQ^{+}05, MYM^{+}05, MB00, MW06a, MAM07, MML^{+}08a, MCC02, MML^{+}08b, NFS^{+}07, RG02, UMASG07, UVY^{+}01]. Olympia [KLG09]. ombrotrophic [KB07]. ombrotrophic-minerotrophic [KB07]. omnivorous [SvKM01]. Omnivory [CCGH05, BCMR00]. one [PGM^{+}09, VC05]. one-dimensional [PGM^{+}09]. Onondaga [DPK^{+}00, ME06]. onset [AMS^{+}02, GPHS08, MAS^{+}02, ZXN^{+}08]. onshore [GB09, SLB^{+}00]. Ontario [BJW^{+}02, FHDKR03, FCJ^{+}06, MMPA04, Mor06, RSC04, SAH^{+}04]. onto [AHR^{+}01, FZL^{+}05]. ontogenetic [LLS08, RHM04, TS06]. ontogeny [HNB^{+}07, PWP07]. Onychopoda [GL07]. opal [PIKB08]. open [BMFS^{+}03, BB09, CAP09, GJRM07, HBI01, KNC^{+}09, KLAF04, MSLH02, MC04, RSH08, RG02, TvdWdB04, VPH^{+}07, WSC^{+}05, WSC^{+}06]. open-channel [MSLH02]. open-coast [WSC^{+}05, WSC^{+}06]. open-ocean [GJRM07, TvdWdB04]. opening [TMH^{+}06]. operated [YWR^{+}07]. Ophiuroid [LSPC02]. opossum [BCMR00]. Optical [ANSS04, AOR07, GDCG08, KMMS00, LRS^{+}03, MGC^{+}07, STHG07, SWF04, ZMK^{+}08, Agr05, BVG^{+}07, BZH00, BZ03, CFO^{+}00, CBB^{+}02, DDZ03, GSO03, HMKZ02, Kir03, LMBP07, LBD^{+}09, MLRMR09, MSZ00, NCPd08, OH00, OCSB05, SSC02, SBW07, ZB01]. optically [MS03, MZV03, SLRM03, TNO^{+}08, VMS^{+}03]. optics [SM01]. optimal [KJ01]. Optimization [SY07, De 02]. Optimization-based [SY07]. optode [FG06, GTK^{+}01, HAES02]. Oran [GST^{+}01]. Orbulina [LEML09]. ordination [RBW09]. Oregon [DD09a, D’A08, TvG02, WHC^{+}06]. Organic [DAL02, HMT^{+}00, KB01, SDKZ^{+}09, TWS01, TVL^{+}00, TB09, ART^{+}06, AVJ^{+}00, All00, AHM^{+}08, ÁSGMP01, AFB01, ACZ00, ARK^{+}00, AKP^{+}09, AHR^{+}01, BFDC01, BFC05, BFC07, BAG^{+}09, BMF^{+}06, BPOT04, BLWB08, BGV02, BNOK^{+}04, BK03a, BRL03, BLJ09, BT00, BK03b, Bou04, BW00, BMC07, CGD08, CG09, CKPS04, CSWD07, CGH^{+}04, CSSNA02, CM03, CS06, CDK02, CLG08, CCKP02, CSK^{+}07, DD09a, DD09b, DGKA06, DGH^{+}07, DB07, DGCMK00, DCC^{+}01, DH01, DKRZ06, Dit04, DdG09a, DR06, EPP^{+}08, EB01, FQH^{+}01, FSSH03, FSSP02, FW07, FPPH06, FCJ^{+}06, FXL04, GSW03, GJS01, GP01, GCS06, GFJ^{+}02, HAA09, HBH^{+}07, HD03a, HDMB04, HBCK03, HPAM01, HSR^{+}08, HSB^{+}06, HKD01, HFE07, HLL^{+}06, Jan03, JMBJ01, KS04a, KB08a, KB08b, KJMJ03, KJJ07, KAJ08, KB06, KWS06, KKH06, KNN06, KJ01, KIB01, KHK^{+}03]. organic [KCV06, KUP^{+}06, LTM01, LFZ^{+}07, LBB^{+}04, LLB07, LFFC06, LRMV04, LBC06, LBD^{+}09, LNJL09, LMZB08, MMDW09, MYM^{+}05, MPJ08, MH00, MH04a, MCF^{+}04, MVFD04, MHK^{+}07, MED^{+}08, MSSB06, MBCD04, MBC06, MBW^{+}01, MHB03, MAM07, MSC^{+}04, MMCW09, MPD^{+}06, ME07, Møl07, MTM^{+}01, MSZ00, MWE^{+}03, MC04, MLPS09, MPdVG07, NSM09, NTM^{+}05, NCÁSP06, NHN^{+}05, NK05, OH00, ORH01, OB04, OV03, ORFD^{+}09, OZM^{+}01, OOB09, OCNH00, PC02, PCC^{+}07, PWX04, PKKB05, PWH^{+}04, PS06, PBGV04, QSH^{+}02, RB08, RSH08, RA06, RCAK05, RFM07, RGAS04, RCVK09, STC03, SMD^{+}04, SHL^{+}05, SWS^{+}06, SDvEB06, SRG00, STP^{+}07, SlBWC^{+}00, SCDR07, SM01, SM05a, SM05b, SC08, SL00, TPM01, TMK^{+}07, TYK^{+}08, TIH^{+}01, TPSF01, TB00, TTSC04, TMUW02, Tun00, TLM04, TMA03, TNO^{+}08, VSSTS00, VJ07, VW08, VMH^{+}01]. organic [SVPK^{+}09, VMSB08, VM06, WNL^{+}08, WSHN00, WGLT05, WIK08, WG01, WW03, WHC^{+}06, WW07, WK09, WV00, WdMD^{+}00, WA01, WS04, WTM09, XLF^{+}03, YJMT08, YT09, ZSS^{+}08, ZB00, ZWC^{+}04, dGvD07, dGMH^{+}08, dGP08, vWSBT08, vWBT09]. organic-rich [SWS^{+}06]. organism [Dup00, Mil02]. organisms [GFK00, HSLV05, HA08, MF04, SIRO05, TS06, WM07]. organochlorine [Ber03, BSS^{+}01]. organohalogens [MMW00]. Orientation [KTM01, RG06]. Origin [HRS04, SP02, BHVC05, BHVC06, BFCLÁ08, CAP09, KB06, Sch06a, WCB^{+}06]. origins [DST08, RPR^{+}08]. Orthophosphate [TRT04, LTT07]. Örträsket [JMBJ01]. Osaka [YYTH07]. Oscillation [Dav00, DJG^{+}06, GA00, GA01, GP03, LD01, LHB04, Wey04]. oscillations [GDB^{+}09, Mor06]. oscillatory [FAC05]. Osmerus [BCBS04]. Ostreococcus [SFR^{+}08]. other [BFS02, CDH^{+}08, EPS00, GSO03, KSS^{+}07, KHNV05, LLS08, MSHE09, MBG^{+}05, MDL00, PBGV04, RL01a, dlM09]. otolith [BW01]. otoliths [RPR^{+}08, SRL^{+}09, WSC^{+}05, WSC^{+}06]. our [Boy01]. outcome [THT^{+}09, vNHM^{+}09]. outcompete [KHNV05, RL01a]. Outer [AGMP04]. Outer-membrane [AGMP04]. outgrowth [Ber01]. outnumbering [CTY06]. overeutrophication [YS06]. overgrowth [LLF^{+}09]. overlooked [MCV^{+}05]. overlying [HOFD01, RHJ04]. overturn [MVC^{+}05]. overwintering [BKAE01]. ovum [GLVFH09]. oxic [BPOT04, CGCPA08, FS00]. oxic-suboxic [FS00]. oxidation [ACG07, EPD09, HLW^{+}07, JKGT08, JNKdV09, KGP^{+}02, MHB03, MGL02, MS05, SDvEB06, TDJ^{+}06, TTSC04, TT05, TKBJ05]. Oxidative [GM04a]. oxide [BM00, DNS06, FCGC^{+}09, NDT07, RW02, TMB07]. oxides [GW03]. oxidizing [KLF^{+}03]. oxy [BPLH04]. oxy-regulatory [BPLH04]. oxyanions [HHS05]. oxycline [FCGC^{+}09]. Oxygen [FG06, PFPH04, RM09a, SLHK02, WCL^{+}07, ACZ00, BI07, BMWW08, DH01, ES00, ESA06, FS09a, FCGC^{+}09, GM03a, GSG^{+}05, GSNH04, GSLT08, GBP03, HLW^{+}07, HFDO05, JHW^{+}07, JLB^{+}06, JN08, KCSM07, KD02, KLF^{+}03, LBG^{+}09, LWP^{+}07, LBSY02, MMDW09, MHLN01, MJR08, ME06, MCPP06b, MGMW06, OH08, ORH01, PLM^{+}09, Plo01, Plo08, PIKB08, RlBW01, RHJ02, RHJ04, SAH^{+}05, SAHS06, SAS08, SDKZ^{+}09, TDJ^{+}06, TBH07, VTR03, WBP^{+}06, WGC00]. oxygen-18 [TBH07]. oxygen-deficient [DH01, TDJ^{+}06]. oxygen-isotope [ES00]. Oxyrrhis [EW08]. oyster [KLG09, KN07]. oysters [TG09]. ozone [HSF^{+}08].

P [ACD10, Caf01, SXLM06, SMS09, Sta01, Bel01, Boy01, Bri01, Dow02, Elm02, Fou02, Gou01, Jun01, Kne02, Mil02, Pom01, Ris02, SG02, Wan02, Wro01, ABJ^{+}06, ART^{+}06, AD01, BSMR08, BNK02, Ber01, Boe00, GHN^{+}02, HW08, HRS04, IR06, LG04, NKJM06, SH10, SL01, VPA02, WGKE04]. P-affected [IR06]. P-deficient [HW08]. P-limited [Boe00]. P-NMR [HRS04]. pachyptila [RSS^{+}03]. Pacific [BNK02, BK03b, CDK02, CLAK08, DBTK02, HPLG09, LKAB04, NKS^{+}02, NFS^{+}07, Aki00, Alt01, BST^{+}05b, BDK^{+}08, CBK^{+}08, Coc01, CLP^{+}03, CW02, DPC^{+}08, FCGC^{+}09, Fis00, FSS^{+}00, HPT^{+}09, HW07b, HB08, KFH^{+}09, LBW^{+}08, LM02, LCPS05, MC00, MG08, MJR08, Met04, NFFK00, NBW^{+}02, NEE08, OLCT08, RWK08, ST04, SWSJ02, SKBC01, SBR^{+}06, SRKK08, SMC08, SLS^{+}02, TRM00, TM02, TMC06, TCCG05, VDK04, VD08, VRS09, YT09, ZMJ^{+}07, vdSvdSG^{+}00]. pacifica [JHGGP06]. pacificus [PFL07]. pack [BGHA04]. Packard [Wan02]. Paerl [Caf01]. PAHs [GHTJ00]. Palatability [BH02]. Palau [WKH^{+}06]. Paleo [TRF^{+}06]. Paleo-indicators [TRF^{+}06]. paleoclimatic [JM08]. paleoecological [SFB^{+}09]. paleoecology [KW04]. Paleogenetic [DPK^{+}00]. Paleolimnological [LFA^{+}09, BPLH04, SF04]. paleolimnology [DES03, HW07b]. Paleoproductivity [MMPA04]. pallida [GCG05]. pallida/Symbiodinium [GCG05]. Pamlico [VWPP^{+}06]. Panamanian [CGWK^{+}09]. Panulirus [BFCLÁ08, GB09]. paper [Bri01]. paradigm [WSC^{+}08]. paradigms [MF01, MF02]. paradox [BHMS06, TKK^{+}03]. PARAFAC [SM05a, YJMT08]. parallel [YJMT08]. Parameterizing [LLA05]. parameters [BBG^{+}05b, GFK00, KVR03, MPACG01, RENA06, SHGM06]. Paraphysomonas [BMCvE06]. Parasite [Hol00, Duf09]. parasites [WKB^{+}04]. parasitism [KVdB^{+}04, MDJ^{+}04]. parasitized [DHTH05, TKK02]. Park [EPP^{+}08, HT03]. Parsons [Wro01]. parthenogen [HD09]. partial [Goe08, LGJ02, SKC^{+}01, YT03]. Particle [BAP^{+}05, JPK^{+}05, BMFS^{+}03, Ben00, BST^{+}08, BB09, ET08, FT03, HOFD01, HR00b, JBvHS08, NFFK00, PNPA04, PE07, Pil04, RL00, SDC00, SM02, SLD00, SVB^{+}08, TNM^{+}01, TFNT07, WSNM00, WCB^{+}06, dVKJK07]. particle-associated [SDC00, dVKJK07]. particle-reactive [SM02]. particle-rich [WCB^{+}06]. particles [ASP04, BS02a, BMFS^{+}03, BS04, BRB^{+}01, CSSNA02, CDH^{+}08, DRM^{+}09, GD01b, GWL^{+}07, GSO03, HHS05, HP03, HA01, JJP04, KB01, LMBP07, MH00, MBLP01, NTM^{+}05, NHS07, PWS09, PE07, PED^{+}09, PP07, SW05b, TS00, VB00]. Particulate [RFM07, BAG^{+}09, BCM01, BJG^{+}00, BNOK^{+}04, BK03a, Bou04, CGL03, CG03, FPPH06, GBM^{+}04, GCS06, HAA09, HD03a, LBM^{+}04, LBC06, MVFD04, MSSB06, MTM^{+}01, MWE^{+}03, OCNH00, PWX04, RCAK05, ST04, SBW07, SBM04, TG03, VMH^{+}01, WV00, WPD^{+}01, dGMH^{+}08]. Partitioning [CSK^{+}07, LBG^{+}09, MBCB08, PDH06, RS09, TLM04, WW03]. parvum [BGR^{+}09, SH03]. pass [BSSL05]. passage [GM04a, LKM03, BTB^{+}08, HMR^{+}07]. passively [SSH^{+}09]. past [BLS^{+}06, BLSQ07, DLDE04, DPK^{+}00, FS09a, HBI01, HFZ^{+}07, MMPA04, RCVK09]. Patagonian [CPB^{+}08]. patches [Fol05]. patchiness [BSS09, HM03b]. patelliferum [Leg01]. pathlength [BK01c]. Pathogen [LWU^{+}05]. paths [LAN08, RM09b]. pathway [BFBHF04, HMSI06]. Pathways [CCKP02, KB07, TTSC04, EKD^{+}07, FSB^{+}08, KGP^{+}02, LFFC06, MSGK06, MMW06, MMDW09, NDT07, VJV^{+}03, WNL^{+}08]. Patiriella [MSBHG01]. pattern [HMD^{+}07, YYB^{+}05]. patterning [HWA03]. Patterns [BFCLÁ08, CGWK^{+}09, CWC^{+}07, Fin04, Hei08, IR06, JCC02, STP^{+}07, UPC^{+}05, ASI^{+}02, ATJH08, DVV^{+}07, FMWC00, GCJ^{+}06, GRR^{+}07, GD00a, GGDC04, GP03, GBA00, HH03, HB03, JGE^{+}09, KB05b, KRM09, LKAB04, Met04, New01, PF01a, RZMB05, RYS^{+}02, SA00, VKV07, WBP^{+}06, YSO02]. Patuxent [FHBW06]. Paul [Bri01, Wro01]. Pb [JBvHS08, NSBA01, SCARM00]. PCBs [GHTJ00]. PCC [BBG^{+}05b, IB05]. pCO [BD04, DLZ^{+}09, FMG^{+}08, KDRR09, MVBM04]. PCR [SF05]. peak [RKNA09]. peaks [BCV^{+}08]. peatland [BLWB08, HPBD01, KB07]. pecos [GBL^{+}04]. Pectenotoxin [KUP^{+}06]. Pectenotoxin-2 [KUP^{+}06]. Pectinaria [JGE^{+}09]. pelagial [CS06, WŠN^{+}07]. Pelagic [HC01, SKBC01, ALH01, AGDH09, BMBA07, CNGV09, CVH^{+}05, HBF^{+}06, Jan03, JKS04, JB00, JS03b, KTG^{+}06, KBL08, KKT05, LN02, MG04, MBR08, MBCB08, MNM^{+}09, MS05, NHTF01, RHK^{+}09, RH06, TCD^{+}03, TS01]. Pelagomonas [DBGR09]. pellets [PIF08, PIKB08, RFM07]. Pellioditis [DVV^{+}07]. penetration [AKP^{+}09]. pengoi [BJW^{+}02, CHW^{+}01, GL07]. Peninsula [SPR02, CDD^{+}03, ROP^{+}02]. pennate [MMLH06, MH07]. Penobscot [MA08]. Penzé [MWM^{+}04]. Peptide [MGL02, ML09]. peptidoglycan [BK03a, NMK03, VvOBM06]. peptidoglycans [GvRE^{+}02]. perceived [DKRZ06, MF03]. perchloroethylene [MHDH00]. perennially [PCD^{+}08]. Performance [PDB03, JFH^{+}05, PSH08, RB03, VH06]. perfusion [SDJT03]. Peridinium [RL01a, SEL^{+}02]. period [HBH^{+}06, JBM^{+}07, SKTD04, TSH^{+}06]. periodic [LMMW03]. Periodicity [KKKT07, BF08a]. Periphylla [DKKA09, KKT^{+}07]. periphytic [WW00]. Periphyton [Ber03, GCJ^{+}06, FH09, HM06a, HK01, LNB04, MMO06, NRB08, QMS^{+}07, SIRO05, SPW^{+}05, SL01, TR03]. permanently [WSNH06]. permeabilities [JHW05]. permeability [AS02]. permeable [BZH08, FS00, FPPH06, FT03, HR00b, HR00c, JFH^{+}05, JHW05, MGGM06, PFPH04, RB09, VPH^{+}07, WBP^{+}06]. Perna [PWX04]. peroxide [TDT01]. Persistence [CPCK00, HBTM05, RWK08, BFC07, PW03]. persistency [GA01]. persistent [ET08]. perspective [Caf01]. perspectives [Tor00]. pertusa [CWC^{+}07, DDL^{+}09]. Peru [DGM^{+}05, HHW^{+}02, LNJL09]. Peruvian [HLW^{+}07]. pH [AKH04, HHS05, HAES02, KT02, LVB08, Plo08, WS06]. Phaeocystis [PSJ00, BRP^{+}06, DK06, EPS00, KMvD^{+}09, SG08, SWCL01, TJV01]. Phaeodactylum [CLPB02]. phage [RSS^{+}00]. phages [RSS^{+}00]. Phagotrophy [Leg01, BM00, UGY^{+}00]. phantom [DST08, MHLN01, OHT06]. phase [DPFL09, GEJ^{+}02, KB01, LH06, MAL08, OV03, WDL07]. phaseoliformis [SO07]. phenology [JBM^{+}07, SKJ^{+}07]. phenols [WIK08]. phenotypes [AD07, SCH07]. pheophorbides [GSB00]. phosphatase [DP01, DR06, NDCP06, RRH^{+}03, SXLM06, SIRO05, TXH05]. phosphatase-producing [SIRO05]. phosphate [BS03, Cul06, FZL^{+}05, FWM^{+}05, GFPG09, LG04, LMT05, MCPP06b, MTV^{+}02, NKJM06, RG02, SCP08, TRT03, TvdWdB04]. Phosphonates [BNOK^{+}04]. phosphorous [MML^{+}08b]. Phosphorus [BNK02, JNKdV09, KHB^{+}00, LKM03, MLC^{+}07, SSM^{+}04, ART^{+}06, AD00b, ASAC04, BB05, BNOK^{+}04, BK03b, BMNSK00, BLG^{+}07, CG09, CC01, CCC^{+}03, CS05, DGV01b, DAC^{+}07, DP01, DWA^{+}02, DR06, EM02, FF04, dSFFTD07, FQC^{+}07, FL06, FXL04, GM03a, GR01, GGS06, GH00, HDSC09, HBCK03, HW07a, HW08, HCK02, HFR08, HFR09, HCD^{+}09, JPJS06, JS08, JCBA08, KIB01, KMHN07, KHM04, LNB04, LH05, LSSA04, LMZB08, LTT07, MWM^{+}04, MOK^{+}03, MPAS00, MML^{+}08a, MGMW06, MBPVR06, MB05, NGS^{+}07, RP06, RB06, RKH03, RPR^{+}02, RTT^{+}02, RHW00, SSK06, SIRO05, SP04, Smi06, SMC08, SAE^{+}08, SC08, SBM04, SMH^{+}02, SQTA07, TRT03, TRT04, TXH05, TCD^{+}03, VAS02, WJMD07]. phosphorus-deficient [dSFFTD07]. phosphorus-limited [SIRO05, VAS02]. phosphorus-rich [BLG^{+}07]. Photo [SDvEB06, RMS^{+}08]. Photo-oxidation [SDvEB06]. photo-protection [RMS^{+}08]. Photoacclimation [AMP^{+}00, vdPAJ^{+}06, Fly03, MSH^{+}06, SFR^{+}08]. photoadaptation [MSH^{+}06]. photoautotrophic [KHKS05, SIRO05]. Photobehavior [CF05]. Photobiogeochemical [XBD^{+}09]. photobiology [FBW^{+}08, WLRT06]. photobleaching [DdG09a, HSR^{+}08, OZM^{+}01, OOB09, WdMD^{+}00]. Photochemical [ACZ00, BRT^{+}03, BT00, BMC07, KWS06, KHK^{+}03, STC03, MSHE09, MSZ00, NCÁSP06, ORH01, OB04, OHUB00, PAF^{+}04, VSSTS00, VW08]. Photochemically [VJ07, DGKA06]. Photodegradation [BKK^{+}06, SWS^{+}06, MPD^{+}06]. Photodissolution [MSSB06, MSHE09]. photoenzymatic [GWK01]. photoheterotrophic [SGT^{+}06]. Photoinhibition [WLRB03, HBS04, SWN08]. Photolysis [TKK^{+}03, BM00]. photolytic [UGSMN01]. photoperiod [RRS06, SH04]. photophysiological [UHB^{+}08]. Photophysiology [KMvD^{+}09]. photoproducts [MMS^{+}02]. Photoprotection [KMvD^{+}09, LSK07, LNB02, GKLF01]. photoprotective [ETCP03, MGW^{+}05]. photoreactivity [AG03b]. Photoreceptors [GF02]. photoregulation [DBGR09]. Photosynthesis [GKLF01, LAHdBJ05, RRHPT^{+}06, AJ01, BN01, BSJM06, CLPB02, CLAK08, DTM^{+}08, EMK08, GRVF^{+}00, Goe02a, GCG05, HFZ^{+}07, KT02, LEML09, MCR^{+}07, NA09, RMS^{+}08, SE06, SLDE09, SWCL01, SRF^{+}01, SWN08, SKH^{+}01, SMG06, SEL^{+}02, TLrR07, Zha00, Zim03]. Photosynthesis-controlled [LAHdBJ05]. Photosynthetic [VH06, BCV^{+}08, BBG^{+}05b, CGCPA08, CEBMF05, DDZ03, ETCP03, FTC^{+}06, GFK00, HYH04, HMR^{+}07, JP03, KVR03, Lan03, MPGB08, MCF^{+}04, NML^{+}05, SHGM06, SLS^{+}02, UPC^{+}05]. photosynthetically [BVK02, vdPAJ^{+}06]. photosystem [OHUB00]. Phototransformation [KS04a]. phototrophic [SF05]. Phraya [DBWS06]. phycocyanin [SPG05]. phycoerythrins [NTDV06]. Phylogenetic [FQC^{+}07, DHP03, MPACG01, YNCK04, dGB02]. phylogeographic [PH07]. Phylogeographical [HWA03]. phylogeography [CRY^{+}08]. Physical [FSB^{+}08, HDWDD09, LLDS06, MSGK06, OBB^{+}04, PDM^{+}09, PBM^{+}07, RHCS08, RBJ^{+}06, BSM01, BFCLÁ08, CWC^{+}07, HZ00, HLCG07, HHB^{+}09, LLK05, LCLP02, NCPd08, NIE^{+}00, PWN07, SCH07, THH02, WA01, WW02, WWWY05, ZZPP01]. physical-biogeochemical [NCPd08]. physicochemical [BB07, SWD^{+}05, SJH03]. physics [DORS08, VH09]. physiologic [dGB02]. Physiological [Has06, HA08, Leo08, MB01, BFS^{+}04, DBS^{+}04, DGM^{+}05, LTGG03, MBL^{+}01]. Physiology [MFG^{+}09, MML^{+}08b, SSV^{+}04, WJMD07]. phyto [PHF03]. phyto- [PHF03]. phytochelatin [AW05]. phytodetritus [Bea03, GHM^{+}09, NDP^{+}07]. phytodetritus-degrading [GHM^{+}09]. phytoplankter [GRV^{+}09, SWH^{+}05, SWH^{+}03]. Phytoplankton [BMD06, CL04, DP01, HZ00, HABF03, HMR^{+}07, HAG08, HHW^{+}02, KLL04, KVR03, LKB04, Mar05, MSH^{+}06, RB06, SB03, SMR^{+}01, YS06, ADA00a, ADA00b, ARV^{+}07, AD00a, AD02, AS02, ASI^{+}02, AH02, BFDC01, BFD^{+}04, BVG^{+}07, BN01, BVH02, BMBA07, BBJ05, BW05, BSR^{+}07, BEM^{+}02, BMSA08, BGC^{+}04b, CSM^{+}08, CMSVA02, CC04, CSE^{+}08, CEBMF05, CMH^{+}08, CLL^{+}09, CG01, CLC02, CLP^{+}03, CMB00, CRF01, CCC^{+}03, CS05, Cul06, CW02, DG06, DKRZ06, DSW05, DAV^{+}05, DGL^{+}04, DR06, ETCP03, EDF^{+}03, FB03, FA03, FQR^{+}06, FQC^{+}07, FMG01, Fly03, FZL^{+}05, GMR^{+}09, GSW03, Goe02b, GFM^{+}01, GSNH04, GSO03, GSJ^{+}09, HY01, HYH04, HRGM07, HHM^{+}09, HBS04, HPL^{+}06, HB08, HR00c, HH04, JK08, JDS08, JTAM09, KCE^{+}00, KGYU06, KTLS03, KNC^{+}09, KLC06, KLS^{+}06, KLS^{+}07, KMMS00, KL01, KMvD^{+}09, KHNV05]. phytoplankton [LAG^{+}03, LTM01, LFS^{+}04, LMSE05, LB04a, Leh00b, LALE04, Leo08, LES^{+}05, LH08, LCC^{+}09, LCPS05, LSSA04, LL06, LH06, LTB09, LRF07, LF09, MLC^{+}07, MBL^{+}01, MCR^{+}07, MT06, MGRE^{+}07, MFF07, MML^{+}08b, MSPAE06, MKS08, Mor03, MN04, MY05, ML09, NDCP06, NP01, NGS^{+}07, OIH^{+}03, PSLO07, PCvRR00, PCA^{+}00, PDB03, QRF06, RLM^{+}08, RBMN03, RL01a, RRH^{+}03, RvBB02, RHK^{+}09, RB02, RRBS03, RRS06, RBW09, SBL06, SVC^{+}07, SKM05, SRZ04, SRG00, Smi06, SNPK^{+}09, SON^{+}03, SKNK09, SSM^{+}09, SC08, SF03b, SF03c, SCV04, SSS08, SMO07, SMG06, SH00, SH07, SLS^{+}02, TCL^{+}09, Tor00, TRM00, TM02, TMC06, TPG^{+}08, TKAW03, UM05, UHB^{+}08, VWPP^{+}06, WG01, WAGH09, WW03, WMW^{+}08, YSO02, YTVR07, ZMK^{+}08, dlM09]. phytoplankton-based [BMBA07]. phytoplankton-nutrient [PCvRR00]. phytoplanktonic [CMJ07]. pico [CSSNA02, WDVA08]. pico- [CSSNA02, WDVA08]. picocyanobacteria [NCC00]. picocyanobacterial [FWM^{+}05, FTY^{+}06]. picocyanobacterium [BH07]. Picocystis [RCE^{+}02]. picoeukaryote [BRL03, DBGR09, FTC^{+}06, HLGI08, RV04]. picoeukaryotes [BCV^{+}08, NML^{+}05]. Picophytoplankton [LA06, BVWB07, BK01a, DTM^{+}08, WNP04]. picoplankter [SFR^{+}08]. picoplankton [ADA00a, ADA00b, BMCvE06, DMEPA04, RlBW02, SRZ04, SR07, TCCG05, WSHN00, vdSvdSG^{+}00]. Pierre [FAM^{+}06]. pigment [BST^{+}08, Leo08, SPG05]. pigmentation [APH^{+}03, Fab06, HLH09, RK02]. pigments [BCM01, BLS^{+}06, DOE^{+}06, TCB^{+}00]. pilot [Har01]. pipe [dGP08]. pistillata [GRVF^{+}00, GFPG09, GMRVFP02, GMAFP03, Shi04, WLRB03]. pit [SFL05]. Plain [BCSW07]. Plains [ARK^{+}00, FLWP09, PGL06, PLMPN08]. planar [FG06, GTK^{+}01]. Planina [VM06]. planktic [SF04]. planktivores [LL05]. planktivorous [MACB07, RH03, SLC^{+}09]. Plankton [BYF09, LN04, MGVM^{+}06, SMS09, BDK^{+}06a, BSS09, CDH^{+}08, CW01, CSK^{+}07, DMJ09, GA00, HWYL07, JS04, JS03a, JN08, LSK^{+}03, MDL00, MIN^{+}05, OAA^{+}06, PSJ00, PRKHW04, PF01c, RlBW02, RWFK00, RZMB05, SIH^{+}06, SAS08, SSS02, SRF^{+}01, SK01, SBB00, SlBWC^{+}00, SSS08, TBG01, TBF04, WDX01]. Planktonic [CPV00, KMHE04, TWS03, AD05a, AVA^{+}02, AH01b, BBCW09, CDD^{+}03, DG06, DJT06, DAV04, DdG09b, FGM^{+}03, FXL04, Goe08, GAM^{+}01, GV09, HSS06, HBS05, HB03, JGZ^{+}08, KAM04, KB05b, Kiø07, Law08, LEML09, MZL04, MG04, MBC^{+}05, OHT06, PCC^{+}07, PSHM04, RRCP03, RCVK09, RNA06, RJS^{+}03, SGT^{+}06, SS01, SS04, TJV01, TXH05, TS06, VGP^{+}08, VMS09, WPW^{+}09, vDBL05]. Planktothrix [RES^{+}08]. plant [BHMS06, KA06, NH06b, SMD^{+}04]. plant-attached [NH06b]. plant-derived [SMD^{+}04]. plant-mediated [NH06b]. plants [CCH02, TSMB05]. plaster [FAC05]. Plastic [SCH07]. Plasticity [BKB02, VK06, HH05, HLH09, LTGG03, SMDG04]. plastids [GSFB02]. plate [HAES02]. Plateau [ET08]. Platforms [DIMP08]. plays [CDD07b]. plicatilis [SS03, SS06]. plot [KJJ07]. plots [MC04, SHBM01]. Plume [JLP^{+}08, AKG^{+}05, CRC^{+}03, CSP^{+}01, DCC^{+}01, FSM^{+}07, HDFM08, KLF^{+}03, RHN^{+}02, SQTA07, TMA^{+}01, WW01]. plumes [BLBB07, FLT02a, KJ01, RKGC^{+}01, WRD01]. plumulosus [SDC00]. plutonium [FNLG00]. POC [SR02, SR04]. Pockmarks [WBP09]. point [PBGJ09]. point-source [PBGJ09]. poise [HDMB04]. Polar [GRG02, HSF^{+}08, HRA07, HM09, SRZ04, Vin00]. pollen [GVT06]. polluted [HDVB01, LFS^{+}02, WS04]. pollution [Mul01]. polonium [SF03a, SF03b, SF03c]. polonium-210 [SF03a, SF03b, SF03c]. polyacrylamide [ET08]. polychaete [HL06, JGE^{+}09, RB08, TS01]. polyedrum [LASR09, vDBL05]. Polygordius [RB08]. polykrikoides [PLK^{+}05]. polymeric [HBH^{+}06]. polymetallic [KCCE06]. polymictic [GA00, MGVM^{+}06]. polymorpha [BFC05, BFC07, BDNF01, SWH^{+}05, WBvE02, WvE04]. Polynesia [APH^{+}03]. polynya [BJG^{+}00]. polyphosphate [HRS04]. polyploidy [AMW^{+}07]. polysaccharide [Arn00, VOG^{+}08]. polysaccharides [QSH^{+}02, Tho09]. polyunsaturated [CWPL05, RBMN03, VW01, vE02]. POM [MRNT^{+}07]. pond [HGB00, VVD^{+}09, ZHFD06]. ponds [DPC^{+}06, Kil00, SCV04]. pool [MIDH05, SP09]. pools [BGHA04, CWC^{+}07, MMCW09]. poor [Has06]. poorly [KL01, LPD^{+}04]. poppei [ROP^{+}02]. Population [DCMS03, JGE^{+}09, SSH^{+}07, TS01, BGC04a, BJW^{+}02, CMH^{+}08, CK08, DG06, FL08, FMD03, GGP01, HHB05, HM03b, Hol00, JJT07, JVL00, JTL01, Kiø07, LASR09, PCDB09, Reu02, RB02, RA00, RNA06, VKvD05, WST06, WSEL09, ZRL06]. populations [ANSS04, Ano04, CBG01, DA03, FKS^{+}01, FRO01, FWM^{+}05, GRDCS09, GWH06, HY01, KCE^{+}00, KMMS00, KSWC^{+}03, LSC^{+}08, LPAQ^{+}05, MG08, MCB08, MC01, MLSS05, RES^{+}08, RL07, SD05, SGPM09, SMDG04, TPSF01, VNH^{+}08, WJMD07, ZLP^{+}03, ZLL^{+}02]. porcelain [HH05]. Pore [JFH^{+}05, JHW05, BM01, BZH08, CWG^{+}07, FS00, HKD01, LH05, MA08, MGGM06, PH03, PFPH04, RB09, RLJdB08, TPFS09]. Pore-water [JFH^{+}05, JHW05, BZH08, CWG^{+}07, HKD01, MGGM06, PH03, RLJdB08, TPFS09]. porewater [DAL02, SL02]. Porifera [YWR^{+}07]. Porites [JM02]. porous [LSF^{+}08, PHK02]. port [SNM^{+}08]. Porter [Ris02, Ris02]. ports [GJRM07]. Posidonia [BMD06, DAMM^{+}09, Fol05, GDM02]. position [BMF^{+}06, HPLG09, MS06, MM05, NSM09]. positions [BVB03]. positive [GNU00]. Possible [BTKGA07, FSS^{+}00, LSK07, SSV^{+}04, BFCLÁ08, HM09, Kiø00, MPJ08, MHDH00, NFFK00]. post [Duf09]. post-consumption [Duf09]. postglacial [FRO04]. Postindustrial [CPL^{+}04]. postlarvae [Ann05]. Postlarval [BFCLÁ08, GB09, HM02]. postsettlement [NA05, PDS05]. Potamocorbula [CTCT03]. Potamoplankton [CPH06]. Potential [TMB07, WLRT06, All00, AE08, BSSL05, BJW^{+}02, CTR^{+}02, DDZ03, EKdF07, FKCF00, Had08, HPT^{+}09, KLF^{+}03, LTOS04, MBF^{+}03, MdG08, MF04, ODR^{+}08, PCvRR00, SLS^{+}02, ZLP^{+}03, ZRL06, ZHL06]. Potentials [JS03a]. power [HFDO05]. practical [Mul01]. practice [Bia01, BK01c, FMD03]. Prairie [Ham01, KD02, MJ03, PLM^{+}09, QLD^{+}02, RLM^{+}04, WR00]. Prasinophyceae [EDF^{+}03]. precipitation [CBD^{+}06, DT00]. precipitation-dominated [DT00]. precursor [BFML05, MBW^{+}01]. precursors [GGIAY08]. Predation [CPCLS04, LUHS04, SvKM01, SA04, AH01b, DJO00, De 02, EAO^{+}02, HMSI06, HNB^{+}07, HM02, HLH09, JJP04, KLWS08, LJ01, LBC^{+}09, MDJ^{+}04, MD09, MSAC06, ODR^{+}08, PD05, VKvD05, WBR03b]. predator [BRM^{+}07, DG06, HHB05, HPV^{+}00, HDWDD09, JS03b, Law08, MD09, PSH08, RH03]. predator-prey [Law08]. predators [BJL01, DHTH05, SCH07]. predatory [BJW^{+}02, FS09b]. Predictable [KCWC06, JCK^{+}08]. Predicting [AH03, BMSA08, HD03b, SRT^{+}03, vdHPH^{+}09, BFD^{+}04, CMD^{+}09, TSH^{+}06]. Prediction [LN06]. predictions [VM05, WAB^{+}08]. predictive [SCS^{+}08, ASH^{+}06]. predictor [CMJ07]. predictors [RBJ^{+}06]. Predominance [BGHA04, RB01b, MTV^{+}02]. predominantly [AD01]. Preferential [EW08]. preferentially [TP03]. preindustrial [KMHN07]. preliminary [LWB^{+}01]. premises [Ben00]. Presence [DR06, BD07, JHW^{+}07, KS04a, KSEÖ^{+}02]. present [CWJ09, HBI01, HFZ^{+}07, RL01b]. Preservation [GvRE^{+}02]. preserved [Dix00, ME07]. Press [Bel01, Boy01, Bri01, Elm02, Fou02, Jun01, Kne02, Mil02, Ris02, Sta01]. pressure [LGJ02, SKC^{+}01, VVD^{+}09, WLS00]. pressures [FL06]. Preston [AH01a]. Preston-static [AH01a]. Pretus [Wan02]. prevent [JJS^{+}03]. preventing [GJRM07]. Prey [CC02, HK06, BF08a, BJW^{+}02, BRM^{+}07, CD02, DHTH05, HG03a, HG03b, Law08, MZL04, NH06b, TFNT07, ZW06]. Prey-specific [HK06]. prezygotic [Goe08]. Primary [BHMS06, LGJ02, MVFD04, MY05, AD05a, AKP^{+}09, Cal01, CP03, CWJ09, CFMH01, CVH^{+}05, CL02, Dal03, DdG09b, GEJ^{+}02, GRG02, GGH07, GEP08, HPE01, HSF^{+}08, HBI01, HHY^{+}05, HBS05, JCC02, KVR03, Law03, LLK05, LL06, MACB07, MBC^{+}05, PBMSG04, RCVK09, SIK07, UPC^{+}05, WPW^{+}09, YPBL07, YS06]. primers [JL08]. Princeton [Boy01, Jun01]. Principles [Bri01, BK01c]. pristine [WS04]. probable [AS06]. probe [AS02]. Probing [SdB06]. problem [LSPC02]. problems [RL01b]. procedure [AD00b]. procellariiform [CCM07]. process [AGP07, dVKJK07]. processes [BHV04, BZ03, BTS^{+}06, BB09, CGH^{+}04, Elm02, GBA00, HBS05, IKT03, KLG09, Kra05, NIE^{+}00, OB04, PWS09, PWN07, RHCS08, SDJT03, SRKK08, ST03, WSNM00]. Processing [TCP^{+}03, CV05, GHB^{+}05, WPD^{+}01, dGP08]. Prochlorococcus [VD08, BBKC03, BBG^{+}05b, CBB^{+}02, HSN^{+}03, MC00, MPRC02, SMCW02, SHSH^{+}07, ZJC^{+}07]. produce [KSEÖ^{+}02]. produced [BRT^{+}03, DNB^{+}04, GvRE^{+}02, MHB01, RFM07, VJ07, WS01]. producers [CFMH01, CVH^{+}05, MACB07, RCVK09]. producing [DMJ09, Lür03, SIRO05, SWH^{+}03]. Production [BSJ02, CMB00, FAL01, HY01, HBN^{+}00, KFN00, Møl07, PIKB08, SRM05, TN09, WG01, WW03, AP00a, AFL07, ADA00a, ADA00b, AD05a, ÁSGMP01, AKP^{+}09, BLJ09, BHMS06, BCDH09, BLSQ07, BJBA03, BJD05, Cal01, CP03, CPV00, CKA03, CF01, CH05a, CL02, lLCChTO08, CS06, CRJ08, CGWK^{+}09, CSK^{+}07, CCCC07, CCC09, CJB00, CD03, Dal03, DdG09b, DKA02, DA05, ERH06, GLP^{+}08, GEJ^{+}02, GCG^{+}02, GGH07, GHS05, GEP08, GM03b, GP00, GSJ^{+}09, HPE01, HSF^{+}08, HBH^{+}06, HD03a, HHY^{+}05, Has04, HPAM01, HBF^{+}06, HBS05, HMT^{+}08, IHH^{+}00, IHHH00, JJT07, JKS04, JCC02, KHP^{+}06, KS04b, KFR^{+}01, KBE^{+}04, KBG^{+}08, KWS06, KL00, KNH00, KVR03, KHK^{+}03, KLSW00, KSM^{+}05, LTM01, Law03, LGJ02, Lee01, LCPS05, MHRW02, MB01, MCF^{+}04, MHK^{+}07, MBG^{+}05, MBCD04, MD04, MBR08, MMCW09, MBC^{+}05, MSPAE06]. production [MY05, MMW00, NFFK00, NEE08, NSH^{+}09, PAF^{+}04, PBGV04, PM08, RNC03, RBC00, RA09, RvAV^{+}06, RSH08, RCE^{+}02, RLM^{+}04, RG04, SAH^{+}05, SAHS06, SSLE02, SCP08, SB03, SRZ04, SR07, SK01, SJW03, SSAH05, SM05b, SMH^{+}02, TIH^{+}01, TPSF01, TS01, TVSS01, TSA09, TAB^{+}00, THH02, VJ07, VHL^{+}04, WAGH09, WS06, YS06, ZHL06, dVKJK07]. Production-respiration [FAL01]. production/respiration [SK01]. Productions [Gou01]. productive [CR06, VMC^{+}07]. productivity [BSG^{+}08, CWJ09, DET04, ES00, GRG02, HBI01, HAL^{+}09, HK01, HLL^{+}06, LAJ^{+}04, MOK^{+}03, MPA^{+}07, MMO06, MML^{+}08a, MML^{+}08b, New01, PBMSG04, SIH^{+}06, TB00, TBH07, UPC^{+}05, WPW^{+}09, YPBL07]. products [HYH04, KKH04]. profile [Leo08]. profiles [BYF09, BZH08, JLB^{+}06, LN06, NDT07, NSBA01, PLG^{+}02, PGM^{+}09, SMS09]. profiling [MJR08, RNR08]. profundal [CCCC07, NEC03, SEC08]. Program [Ham01, How00]. progressive [KCSM07]. prokaryotes [CW02, GZY^{+}08]. Prokaryotic [RvAV^{+}06, BAG^{+}09, LNJL09, WSHW05]. propagating [HP03]. Propagation [LH08]. Propagule [AG03a]. proper [BEM^{+}02]. Properties [TMG03, ABDS04, BMFS^{+}03, BZ03, CFO^{+}00, CBB^{+}02, Cra07, DKA^{+}03, DDZ03, ET08, GSO03, Kir03, Lan03, LSK^{+}03, MLRMR09, MMKN08, MGC^{+}07, OCSB05, SSC02, SWF04, SBW07, TG03, UHB^{+}08, ZB01, ZMK^{+}08]. property [MSZ00]. Prorocentrum [JRS03]. protect [KVdB^{+}04]. protection [RMS^{+}08]. protective [HBG^{+}09]. protein [RS01, SR07]. proteins [CSE^{+}08, MLG^{+}03, ST04]. proteobacteria [BGHA04]. Proteolytic [OS05]. protist [JS04, LCLP02, MDL00, SWS^{+}03]. Protista [AG03a]. protists [AVA^{+}02, BVB03, RC07, SABA02, SWH^{+}03]. protobranch [ST03]. protocol [STST07]. protons [OHT06]. protozoa [BS01, LB00, TF04]. protozoan [LJ01]. Protozoans [SAP^{+}06]. provide [BBCW09, MWE^{+}03]. provided [TB00]. proxies [MSC^{+}04, TCL^{+}09]. proxy [ASS03, FHB^{+}09, PWS09]. Prymnesiophyceae [EPS00, TJV01, TSG^{+}01]. prymnesiophyte [DK06]. prymnesiophytes [vdSvdSG^{+}00]. Prymnesium [BGR^{+}09, Leg01, SH03]. Pseudo [THL^{+}09, MHRW02, MMLH06, MH07, OLCT08, PD02, SGG^{+}08, TWC^{+}09]. Pseudo-nitzschia [THL^{+}09, MHRW02, MMLH06, MH07, OLCT08, PD02, SGG^{+}08, TWC^{+}09]. Pseudocalanus [SvKM01]. Pseudomonas [CG08]. pseudonana [BS03, EH00, SWN08, SSC02]. Pseudopterogorgia [JL08]. psychrophiles [CTY06]. psychrotolerant [CTY06]. Pu [NSBA01]. public [CK00]. Puget [MMKN08]. pugettensis [DD09a]. pulex [AY08, HD09, ZB03b]. pulicaria [BAPMMB01, CWPL05, GPAP^{+}04, HD09, SW05a, WH07]. pulsations [Mor06]. pulse [KBG^{+}08]. pulsed [HPAM01]. pulsed-tracer [HPAM01]. pulses [HBTM05]. pump [BBL^{+}03, RCL^{+}02, WC04]. pupal [SrL05]. purified [GPAP^{+}04]. purifying [Alv07]. pycnocline [SFL05]. pyrifera [FSB^{+}08]. pyrophosphate [SMP^{+}01]. Pyrosoma [LJ09].

qualitative [RJS^{+}03]. quality [AFH07, BB03, Boe00, BMNSK00, BvE07, BH02, BLG^{+}07, CWPL05, HRGM07, LFPH07, MBK^{+}09, MSJMN02, NHPHH08, QMS^{+}07, RBMN03, RL07, SBS07, Sos00, SS01, SS04, TRF^{+}06, VPA02, WBvE02, WvE04]. Quantification [DSD02, FSK^{+}04, GYBA04, HPLG09, Ost03, PBGJ09, TFNT07]. quantified [AS02]. quantify [BHV04, GMMW03, KH03, SD00]. Quantifying [CH01, HFR09, MKV01, PDH06, RVP^{+}09, THA01, VHL^{+}04, WA09]. quantile [BMSA08]. Quantitative [DB07, DKRZ06, Kut04, RJS^{+}03, BLS^{+}06, SF05]. quantity [AFH07, QMS^{+}07]. quantum [HMKZ02, MMS^{+}02, Mor03, VSSTS00]. quartet [RVP^{+}09]. Quebec [RKH03, PBC02, MPAS00]. Queen [KW04]. quotas [BFCS^{+}01].

R [Caf01, Dow02, Elm02, Mil02, SMS09, Wro01, WGKE04]. Ra [BSA^{+}00, RVP^{+}09]. Rabalais [Dow02]. radiance [DKRZ06, VMS^{+}03]. radiances [MS03, MZV03]. radiata [WLP^{+}07]. radiation [ARB^{+}04, ARK^{+}00, BSMR08, BN01, BVH02, BVK02, CMSVA02, DTFL03, DSW05, FGM^{+}03, GRV^{+}09, GWK01, HR00a, HBS04, HBS05, KLWS08, KBD^{+}00, LVC^{+}00, LCS^{+}03, LW01b, LBC^{+}09, Les01, LLB04, LFPH07, LNB02, LA06, MSAC06, MGW^{+}05, NRB08, OV03, SWN08, SNPK^{+}09, TXH05, WdMD^{+}00, WWS^{+}00, ZB00, vdPAJ^{+}06]. radiation-dependent [HBS04]. radiation-nutrient [MSAC06]. radical [PBGV04]. Radically [PH07]. Radiocarbon [RA06, BW01]. radioisotope [NKJM06]. radiometric [SSMM00]. radionuclide [RFM07, Shu01]. radionuclides [JBvHS08, SM02]. Radium [BSL^{+}07, Kra05, CBA01, Cha07, KM02, KH03]. Radium-226 [BSL^{+}07]. Radon [CHB08, CBH08]. Radon-222 [CHB08, CBH08]. rainbow [BCBS04]. rainfall [DC05, GGS06, RNR08, TMG03]. Rainwater [WKSM08]. Randers [BDS^{+}04]. random [DR00, KCWC06]. range [HBH00]. ranges [FBW^{+}08]. ranging [SF05]. Rangiroa [APH^{+}03]. Raphidophyceae [BGC04a, LCS02]. raphidophyte [SNM^{+}08]. Rapid [GD01a, Har01, HPAH^{+}02, MCJ^{+}06]. raptorial [VKvD05]. Rate [BGJ01a, BGJ01b, AKE04, AFH07, AELS04, ASAC04, BCV^{+}08, CEBMF05, CG08, Com02, dSFFTD07, FKCF00, GFK00, IHHH00, Lan03, MC00, MF01, MF02, MB05, RSY04, SCARM00, SKH^{+}01, SQTA07, VPA02, WGKE04, WSL^{+}02, YNCK04]. Rates [BZH08, JKGT08, LLF^{+}09, AD02, Ano04, ASAC04, AVA^{+}02, ASH^{+}06, BF08a, BMSA08, BPF^{+}08, CRJ08, CHB08, FSBB05, Fra01, GEJ^{+}02, GRdL^{+}08, HK06, HCC^{+}02, Har01, Has04, HB03, HFZ^{+}07, JS04, KD02, KB05b, KKT05, KGP^{+}02, KSWC^{+}03, Mar05, MHB03, MHC05, MPB^{+}08, MVW^{+}04, PIKB08, PAS^{+}05, RC07, SAS08, SSAZ01, SON^{+}03, SWH^{+}03, TRT03, TSH^{+}06, TGdB^{+}01, TvdWdB04, Too01, TFNT07, TS00, VGP^{+}08, WBP^{+}06, dBWF^{+}05, Ano08a]. rather [ATJH08]. ratio [BSMR08, CCC^{+}03, CS05, FQC^{+}07, LMBP07, MBLP01, MKM06, NHPHH08, OKO^{+}01, RCVK09, Sch01, SL01, VW08, VPA02]. ratios [ALRM08, Ano04, Ber01, BBP^{+}06, CSK^{+}07, DdG09a, ETCP03, Fin04, GKJ04, HW07a, HSR^{+}08, HABF03, HM06a, HFZ^{+}07, IR06, KHQ01, KSWC^{+}03, LGN^{+}06, LTGG03, LBG^{+}09, LG04, LMT05, NH06a, RFM07, SAS08, SF04, TBHS01, ZWC^{+}04]. ray [HSN^{+}03]. rDNA [MWH01, RV04, vdSvdSG^{+}00]. Reach [BHV04, TMP02, Bat00]. Reach-scale [BHV04, TMP02]. reaction [RMS^{+}08]. reactions [HPL^{+}06]. reactive [HM09, SM02, SBM04]. reactivity [BRT^{+}03, DAL02, MH00, NCÁSP06, OH00, TM05, WR00]. reactor [dGP08]. realm [AGDH09, RvAV^{+}06]. reared [TW09]. recalcitrant [VJ07]. recently [RM06]. receptors [AGMP04]. recharge [CIWH09, KH03]. reciprocity [GWK01]. Recognizing [OHCP02]. recolonization [LTOS04, PW01, SA00]. reconciles [TIH02]. Reconsidering [AE08, Had08, MF01, MF02]. Reconstructing [HCH00, PDTR06, TBHS01]. reconstruction [JS03a]. record [ES00, FKCF00, GDB^{+}09, GSG^{+}05, HLD^{+}03, KBE^{+}04, MWE^{+}03, Pil04, RMG^{+}04, TB00, Tun00]. recorded [GWH06, SRL^{+}09]. records [SA09, TB05, TCL^{+}09, ZC02]. recovering [MED^{+}08, RRHPT^{+}06, VMKN03]. Recovery [DP04, HBH^{+}07, HBS05, KCCE06, RG07, SK04, WMRK06]. Recruitment [BBG05a, MG04, NBW^{+}02, AH03, BTKGA07, CTR^{+}02, HM02, JL08, LW02, MW02, PvAB^{+}03, SP09, SDH02, TCR03, WCM09]. recruits [JL08]. recurring [RKNA09]. recycling [LFFC06, SK05, SDV^{+}06]. Red [BTKGA07, DKRZ06, FPZ09, FWM^{+}05, KW04, MLC^{+}07, BW01, DCMS03, JVL00, KHNV05, KSP07, MHDH00, PLK^{+}05, PN06, PMK08, LPAQ^{+}05, NA05]. red-tide [JVL00]. Redfield [Chr05, KHQ01, MBLP01]. Redfieldian [PLW^{+}08]. redistribution [DR00]. redox [ESSS01, KSM06, KSM^{+}05, TIH^{+}01, UWA^{+}07]. reduce [HFLD03, HW09, JJP04]. reduced [EMWL06, HBI01, TCD^{+}03]. reduced-gravity [HBI01]. reductase [HWL^{+}00]. Reduction [QYEA06, SKNK09, GMA^{+}06, GSLT08, JPJS06, MP00, RW02, SRNW06, Sos00]. Reef [Chi00, DDL^{+}09, SGLT08, SSV^{+}04, SWML08, ALH01, AS00b, AP00c, BSM01, BPC06, BTKGA07, FAL01, FAM04, FAC05, FAL^{+}07, FMWC00, FSSV08, FB05, GMR^{+}09, HMSI06, HLM08, HG03a, HG03b, JLP^{+}08, JP03, LGJ02, LWP^{+}07, MG08, MSL^{+}03, PC04, PSS^{+}06, PVD^{+}07, RMK^{+}06, RKM^{+}06, RCAK05, SLKP05, SMH^{+}02, SSL05, TGF00, VNH^{+}08, WLRT06, WKH^{+}06, YYB^{+}05, dGMH^{+}08, GEP08, KDS03, LLvW^{+}03]. reef-building [BPC06]. reef-lagoon [WKH^{+}06]. reefs [Bat02, BF08b, BTKGA07, CCC^{+}01, FDP^{+}08, LSM03, PSH^{+}07, RKK09, Ris02, UMY^{+}02, YYG02]. reflect [BEM^{+}02]. reflectance [DKA^{+}03, Fyf03, GDCG08, MZV03, RDPM06, SLRM03, VM05, WR03, ZB03a, ZVRL03]. reflected [FHBK06]. reflects [AM08, BMNB^{+}06]. refractory [CM03]. refuge [MHLN01, MD09]. Refuges [WBP09]. refugia [Ano08b, CSG^{+}06, HBTM05, BJL01]. regeneration [ABFS09, BA01, CRJ08, DHWW02]. regime [ACHG07, BCSW07, DACG01, JSJ03, KMC^{+}00, KSS^{+}07, MGRE^{+}07, MVA^{+}09, NIE^{+}00, SSvNH01, SSvNH03, TYK^{+}08, vdPVB07]. regimes [BVH02, BRS01, CJ08, Car01, EAO^{+}02, LKB04, MJW03, OCSB05, PBM^{+}07, TKM07, WBB09, WBR03b]. region [BS02a, BZS07, JBM^{+}07, MBPVR06, SRZ04, THL^{+}09, XLF^{+}03, ZB01]. Regional [CBK^{+}08, EJL^{+}06, HWA03, Rit09, XLF^{+}03, CLF^{+}08, GBL^{+}04, HH03, LD01, MCB08, NSH^{+}09, RLP^{+}06]. regional-scale [HH03]. regions [BMH^{+}08, HSF^{+}08, PGC07, VPH^{+}07, WS01, XLF^{+}03]. Regression [McA03, BMSA08]. regrowth [PW01, TYK^{+}08]. regulate [BDK^{+}08, RH06]. regulated [BNM^{+}09, SB03, THR00, WLS00]. regulates [BLJ09]. regulating [MJW03]. Regulation [CCS06, FLWP09, GP06b, LJ01, SGPK07, SL00, DGD04, DR06, FQC^{+}07, FSS^{+}00, HW08, JKGT08, MP00, SFB^{+}09, STP^{+}07, SSJ07, TW09]. regulators [TDC^{+}09, WSVS09]. regulatory [BPLH04, MRNT^{+}07]. Rejoinder [EW03, OSDB07, TSDK^{+}09]. related [AWE07, AE08, AAD08, BMFS^{+}03, BHV04, BDK^{+}06b, DJG^{+}06, DS07b, Had08, LSK07, MFK^{+}09, MBDL01, Mor06, PvAB^{+}03, TS00, TG03]. Relating [BGC04a, CBD^{+}06, HBL02, UHB^{+}08, GRCS06]. Relation [DBS^{+}04, AS06, BBLB05b, BBLB05a, Dup00, FMWC00, FNLG00, FCJ^{+}06, GHS05, HB03, KHP^{+}06, LALE04, LMBP07, MNSV03, MGF03, SGPK07, TGdB^{+}01, TvdWdB04]. Relationship [HHY^{+}05, KTLS03, KSS^{+}07, LRMV04, ODR^{+}08, VDK04, BAPMMB01, BSS09, Cul06, ERH06, GH00, HSLV05, Hol00, HWL^{+}00, KL00, LD01, MRM07, TRT04, TCB^{+}00]. Relationships [CKPS04, HBCK03, JL08, LIE06, MS06, SH04, ZB01, BDC00, CPB^{+}08, CLC02, FAL01, Hei08, IKT03, Law08, LTB09, MPAS00, MDL00, MBC01, RLM^{+}04, RJS^{+}03, SAS^{+}03, TCC06, WBD02, Wey04]. Relative [LFFC06, ML02, MML^{+}08b, dGP08, AH01b, BOC01, BW00, CNGV09, EKD^{+}07, FF04, HLP08, JKS04, LLvW^{+}03, MBR08, PWN07, PD05, RH06, SKR01, SMC08, SHS01, TKAW03, VMC^{+}07, XTSM07]. relatively [RB08]. RElease [KTLS03, HHS05, HM09, WW07, BW99, BW00, FXL04, HMNL06, JNKdV09, KB06, MB05, MBH^{+}06a, MBH^{+}06b, PRKHW04, RTT^{+}02, SRG00, TRT03, TSKA05, YW02]. relevance [LBT06, PvE00]. Relevant [BB04b]. relict [PCD^{+}08]. relicta [BRM^{+}07, BCMR00, HLNM^{+}03]. relief [CR06]. remain [HKF01]. remains [LNJL09]. Remarkable [HSCF06]. remineralization [DD09a, DR06, GGDC04, Plo01]. Remote [BDNF01, SPG05, WR03, DZL^{+}03, GBA00, HM03a, HTWG08, KDS03, Kut04, LRS^{+}03, PFN^{+}05]. remotely [YWR^{+}07]. removal [AD05b, ATJH08, HBAK09, HTS^{+}09, HBM08, MHS^{+}09, PNPA04, RCAK05, SK04, YSM^{+}03]. Removing [KHQ01]. REMUS [SGLT08]. renewal [AELS04, HPAH^{+}02, PFH^{+}03, WRG^{+}05]. reniformis [TGFPT08, TTG^{+}09]. reoligotrophication [GP06a, JJS^{+}03]. repair [GWK01, LNB02, RMS^{+}08]. repellents [HMJ04]. repetition [CEBMF05, GFK00, Lan03, RSY04, SKH^{+}01, SQTA07]. replace [TSG^{+}01]. replete [dlM09]. replicated [CG08]. Reply [AP00a, AE08, BK01b, CP03, CR08, MWE^{+}04, SR04, SSvNH03, SRG00]. Reproduction [VC05, AY08, BB05, BB04b, DKt08, GWK01, KS04b, LLB07, MBK^{+}09, PDS05, RB06, SS03]. Reproductive [KWTS07, AD07, Goe08]. requirements [Ano04, BFC05, BFQF^{+}07, KSWC^{+}03, MWM^{+}04, MMLH06, MW06b, PQP05, SrL05]. Research [Ham01, Ack03, Bia01, IKT03, STST07]. reserves [RG07]. Reservoir [CWLH06, BI07, BSG^{+}08, CNGV09, GCG^{+}02, HKNG08, HLL^{+}06, KB08a, KB08b, PE07, SVC^{+}07, VCCS05, VRC07, VOG^{+}08, WŠN^{+}07]. reservoirs [CG01, KVR03, MNA^{+}09, RORR^{+}09, TDC^{+}09, WSVS09]. Residence [RC05b, CBH08, MRNT^{+}07, MCLM02, NTM^{+}05]. residual [BDvO^{+}03, FRO04]. Resilience [WMW^{+}08, BF08b]. resistance [AD07, RRCP03]. resolution [BM01, DZL^{+}03, ES00, FDH^{+}01, GBA00, HTWG08, KKH06, KJH03, MSL^{+}03, NDT07, OHUB00, SHS03b, WAB^{+}08]. resolves [PCDB09]. Resolving [ETCP03, SM05a, HM03a]. resonance [ART^{+}06, AI03, CEA^{+}09, WdBD^{+}01]. Resource [BB03, CMH^{+}08, CPCS01, SrL05, CPCLS04, CC02, EW07, KDO09, MSAC06, PDH06, TBG01]. resources [ABJ^{+}06, BDC00, Dow02, MSJMN02, SJRB09, TMP02, vNHM^{+}09]. respective [NIE^{+}00]. Respiration [KJJ07, MTC^{+}06, MBG^{+}05, CPV00, Com02, FAL01, GEP08, JHW05, KAJ08, LGJ02, LBG^{+}09, LEML09, PIKB08, RvAV^{+}06, RSH08, RH06, SRF^{+}01, SK01, Too01]. Respiratory [PG00, LBSY02]. respired [MdG08]. Response [ABJ^{+}06, ACD10, AWLH04, AWH07, Car01, CH05a, DOE^{+}06, LTHW04, MB00, MD09, MED^{+}08, MAC^{+}03, MSPAE06, NT08, SEC08, SMS09, TSKA05, WWWY05, AK00, BSMR08, BB04b, BI07, CMJ07, Coc01, CMB00, CJWR05, DSW05, FSSH03, FL06, GRG02, GN00, Gil03, HAG08, LDC^{+}01, LPAQ^{+}05, LLK05, LP07, LCPS05, MBL^{+}01, MGH08, MSH^{+}06, MVC^{+}05, OBSD04, PD02, RLP^{+}06, RP06, RRH^{+}03, RG08, SRT^{+}03, SS06, WK09, YCM^{+}01, YS06, vDBL05]. Responses [FF04, LG04, LSK^{+}03, ML05, Smi06, WDX01, WM06, ZLP^{+}03, BD07, BLM^{+}06, CWPL05, CGWK^{+}09, DKKA09, GCJ^{+}06, GEP08, GWK01, HB05, HMR^{+}07, HBM06, MA00, MBPVR06, OAA^{+}06, SSH^{+}07, WGD^{+}02]. responsible [PFL07]. Resting [JGZ^{+}08, AFL07, KRM09]. restoration [How00]. restoring [Bel01]. Results [CCKP02, HTS^{+}03, KCP^{+}04, MT06, vOMSM06b, AWH07, HR07, OSDB07, TVL^{+}00]. resurrection [KW04]. resuspended [KWS06]. Resuspension [Bea03, CJB00, GJT^{+}08, JJS^{+}03, KBE^{+}04, NTM^{+}05, NHPHH08, PF01b, SPZ03, SABA02]. retention [AMS^{+}08, BFS02, CSPW00, ED05, FSSP02, GM03a, GVT^{+}09, LB04b, MGMW06, PC04, SLE04, SMJ^{+}06, SHBM01, VEMM07, WLS00]. returns [HW07b]. reveal [AD07, MBR08, NDT07, Reu02]. revealed [PLM^{+}09, SF05, STK^{+}04]. reveals [FTY^{+}06, LPAQ^{+}05, MGRE^{+}07, RA00]. reversal [CCAÆ06]. Review [Bel01, Boy01, Bri01, Caf01, Dow02, Elm02, Fou02, Gou01, Jun01, Kne02, Mil02, Pom01, Ris02, SG02, Sta01, Wan02, Wro01, Ack03]. revision [TIH02]. revisited [AKG^{+}05, KHQ01]. Reworking [KF01]. Rhône [ASS03]. rhizosphere [FG06]. rhodamines [UGSMN01]. Rhode [Ano08b, CSG^{+}06]. rhodolith [FKCF00]. Rhodophyta [PGC07]. rhythms [UPC^{+}05]. Ría [ÁSGMP01, SGG^{+}08]. ribosomal [WGKE04]. Rica [SRM05]. rich [BHV04, BLG^{+}07, CGCPA08, DNB^{+}04, Has06, JCC02, PTKP00, PCA^{+}00, SWS^{+}06, SBW07, WCB^{+}06]. Richard [Boy01]. richness [BS02b, ERH06, HSLV05, RLM^{+}04, WHHS06, WSHW05]. Ridge [Dix00, SWD^{+}05]. Rift [Pil04, JM08]. Riftia [RSS^{+}03]. Rinsjön [Sos00]. riparian [FQH^{+}01]. Rise [NDP^{+}07, LRF07]. rising [BF08b]. risk [ACHG07, DJO00, De 02, HR03, JJP04]. River [CGD08, CF01, Fis00, HB06, KCSM07, MCJ^{+}06, MSJMN02, SDJT03, AJS07, BLBB07, CCS06, CL02, CPH06, CRC^{+}03, DD09b, DBWS06, FSSP02, HBTM05, HDFM08, JCW08, JBvHS08, MBCD04, PCA^{+}00, SB03, SDvEB06, TMUW02, TSSWK^{+}06, TSDK^{+}09, TSMB05, THJ^{+}09, VVSM09, WTM09, dCCL09, dGP08, AJS07, ASS03, CWWS00, CB04, DFBA07, DCC^{+}01, DS07a, FSM^{+}07, HMT^{+}00, HDFM08, MA08, NA03, RBC00, SRT^{+}03, SBM04, SQTA07, TWS03, TSMB05, TB09, WGN^{+}08, WC04, dGP08]. river- [JCW08]. river-dominated [CL02, DD09b, DBWS06]. River-estuarine [CWWS00]. Riverine [WHC^{+}06, ABJ^{+}06, BLBB07, CG09, CYN05, DGCMK00, FL06, GP00, KS04a, WGN^{+}08, WNAW09]. rivers [AS06, Ber03, CKA03, CSPW00, CH05b, Dod06, HA01, HSB^{+}04, MJEM06, PWH^{+}04, RLC^{+}01]. Rn [CDBC00]. RNA [ABJ^{+}06, VD08, VPA02]. Roads [WMW^{+}08]. Rock [WMRK06]. Rockall [vODL^{+}09]. rockfish [RRSC01]. Rocky [BSS^{+}01, ABFS09, DMS^{+}03, Har02, HD03b, HHT01, SMA^{+}06, WL06, BMC07, EPP^{+}08, ML02]. Role [CFG07, DH01, HA01, KNN06, LVB08, LF09, MHB03, PWH^{+}04, VOG^{+}08, WNAW09, ARB^{+}04, BRL03, Ber01, BZH08, CT05, CPL^{+}04, CDD07a, CDD07b, CMC^{+}09, DTFL03, EPS00, GAM^{+}01, GHJ^{+}05, HCK02, HM09, HBS05, HH02, JHGGP06, KAM04, Kiø00, KRM09, LTGG03, LN02, LMZB08, MOK^{+}03, MCvEH08, MID^{+}07, MKS08, OHCP02, PP04, RBMN03, RKH03, RHJ02, Sch01, SBS07, SVC^{+}07, SW05b, SLL^{+}04, Thi00, THJ^{+}09, TLM04, TMB07, UL05, WvE04, ZWHP04, ZGP^{+}00, vDBL05, vWSBT08]. roles [AFH07, AP00b, FBW^{+}08, GZL02, HH04, HM02, PWN07, WCL^{+}07, ZWC^{+}04]. Rolpa [SCY00]. Rolwaling [SCY00]. Rømø [dBWF^{+}05]. Roosevelt [Gou01]. roots [KA06]. Roseg [TMUW02]. Roseophage [RSS^{+}00]. Ross [BSR^{+}07, Smi00, TA03, dVKJK07, dVKT^{+}09]. rotating [dlFSIN08]. rotifer [CPCS01, CPCLS04, GRDCS09, Gil03, HBM06, KSP07, RRCP03, VZM07]. Rotifera [SS03]. Rotifers [GS04, DHP03, MA00]. roughness [FAC05, JWRW07]. round [UMASG07]. Routes [HCC^{+}02]. rRNA [MWH01, PCDB09]. rubens [MA00]. ruber [LEML09]. RubisCO [RSS^{+}03, SHSH^{+}07]. rubra [CRY^{+}08]. rubroniveus [GHS05]. runoff [DC05]. rural [GHTJ00]. Russian [SJR^{+}04].

S [AE08, Caf01, Mil02, RV04, SMS09, AHM^{+}08, MWH01, VTO^{+}05, vdSvdSG^{+}00]. Saanich [Tun00]. Sacramento [MSJMN02]. Sagami [GTS^{+}09, GSB^{+}09, HHY^{+}05, SAH^{+}05, SAHS06, SAS08]. Saginaw [BDNF01]. Saharan [RORR^{+}09, RG02]. Saint [FAM^{+}06]. Saint-Pierre [FAM^{+}06]. saline [DHP03, HL02, RVP^{+}09, RA09, RBJ^{+}06, WR00]. salinity [BKB02, BdG02, CF01, FHB^{+}09, GN00, JCBA08, KKS^{+}01, LVB08, MPD^{+}06, RB09, SP02, TM05, dGB02]. Salmon [HFE07, SAH^{+}04, BBP^{+}06, BLS^{+}06, BLSQ07, FCJ^{+}06, HW07b, ML05, RPR^{+}08, SFB^{+}09]. Salmon-derived [SAH^{+}04]. salmonids [CPB^{+}08]. Salpa [LIE06]. salps [LIE06]. Salt [THT^{+}09, BGE04, BCDH09, Cha07, FSK^{+}04, GK03, LN06, PDTR06, PL09, SH06, TSML07, ZHFD06]. saltmarsh [KGP^{+}02]. Saltmarshes [CLG08]. Salton [RA09]. sampled [PWN07]. sampler [BM01]. San [MSJMN02, CZRD^{+}04, CTCT03, CGL03, CG03, CGL^{+}04, CCH02, Kim05, Leh00b, LRF07, LF09, MC06, NDCP06, NCC00, SH06, WKFP06]. sand [RLJdB08, SFC^{+}00, WBP^{+}06, WWS^{+}00, YYG02]. sand-substratum [WWS^{+}00]. sandflat [DAL02, HR00b]. sands [GHM^{+}09, RB08]. sandy [BM01, CR06, CWG^{+}07, HU07, JHW05, dBWF^{+}05, JFH^{+}05]. Santa [BVB03, BBG05a, MSSW^{+}07]. sapidus [HFC^{+}04, JWRW07]. Sarasota [DC05]. Sarcocornia [EMPI07]. Sarcocornia-dominated [EMPI07]. sardine [MRM07]. Sargasso [Ano07, CGH^{+}04, CCF04, LSSA04, MAMC02, MCM06, TT05, TKK^{+}03, ZXN^{+}08]. Satellite [HDFM08, BST^{+}08, KSS^{+}07, Kut04, SHS03b, UM05]. Satilla [CWWS00]. saturated [SHGM06]. saturating [NA09]. saturation [CPGC09, GWR^{+}09, TvdWdB04]. Sau [VCCS05]. saxatilis [BFS02]. scalars [MF04]. Scale [SAE^{+}08, ASGA^{+}07, AIS00, AIK04, ABFS09, Bat00, BHV04, CSM^{+}08, DJH05, EGP08, GGIA06, GBA00, HH03, HISW00, HMD^{+}07, LBEP06, LES^{+}05, LH05, LP07, MRM07, MHKP02, MMKN08, NFFK00, OD08, PF01a, Plo01, Plo08, RKWG00, RHJ02, RLM^{+}04, SCB03, SI08, TMP02, Tun00, TG03, UMY^{+}02, WG04, WdBD^{+}01, YT09]. Scale-dependent [SAE^{+}08, GBA00]. scales [BB04b, Dav00, ETCP03, FSSV08, GP06a, HS04, MCLM02, PH07, RS09, SGPM09, UHB^{+}08]. Scaling [HBH00, KMHE04, MCR^{+}07, OS01a, RS09, SC07, CMH^{+}08, Fin01, HFDO05, Kiø00]. scallops [PP00, TG09]. scan [MSL^{+}03]. scattering [Agr05, BMFS^{+}03, BVN^{+}00, BVN^{+}02, BVK02, DRM^{+}09, EMIP05, GD01b, KJH03, LMBP07, PE07, PED^{+}09, SW05b, ZLL^{+}02, ZMK^{+}08]. scavenging [BL07, FZL^{+}05]. scenarios [FS09a]. Scenedesmus [AGL03, LV01, Lür03]. Scheldt [BDS^{+}04, HDVB01, SMH^{+}06]. schools [RFW08]. Science [Wro01, Bia01, Wan02]. sciences [Gai00]. Scientia [Wan02]. scintillans [DCMS03]. Scleractinia [DDL^{+}09, QYEA06, WCGBG07]. scleractinian [DOE^{+}06, GRVF^{+}00, GFPG09, GMRVFP02, GMAFP03, TGFPT08, TTG^{+}09]. Sclerolinum [SWD^{+}05]. Scotian [GPHS08]. screen [BD07]. SCUBA [GFK00]. SCUBA-based [GFK00]. scyphomedusa [RG06]. scyphomedusae [PD05]. Sea [BTKGA07, BB07, BLM^{+}06, FPZ09, MFG^{+}09, RA09, SH05a, TMYH02, YM03, dVKJK07, dVKT^{+}09, AD02, BSM01, Bat02, Bea03, BF08b, BBKC03, Bou04, BMH^{+}08, BJBA03, BZ02, BZH08, CMCB^{+}09, Dit04, DJH05, DS07b, EKdF07, EPD09, FHB^{+}09, GWH06, GMMW03, GSFB02, HHM^{+}09, HSB^{+}04, JK08, Jun01, KLAF04, LSPC02, LLB04, LLF^{+}09, MSH^{+}06, NHN^{+}05, PTK^{+}07, PCvRR00, PMPC00, RSH08, RG04, SO07, SL02, SR02, SR04, SDV^{+}06, TS08, WCM09, WCL^{+}07, YSO02, ART^{+}06, Ano07, AWE07, BZH00, BSR^{+}07, BEW^{+}00, BEM^{+}02, BBJ^{+}06, BTS^{+}06, BRP^{+}06, CGH^{+}04, lLCChTO08, CLL^{+}09, CCF04, Dup00, FDH^{+}01, FNLG00, FWM^{+}05, FTC^{+}06, FTY^{+}06, GEJ^{+}02, Goe02b, GLVFH09, GWL^{+}07, GST^{+}01, GZY^{+}08, HLK^{+}07, HVH00, HSB^{+}06, HH02, HWYL07, HWL^{+}00, JHW^{+}07, JHW05, JKGT08, KBL08, KKS^{+}01, KLF^{+}03, KUP^{+}06, LHWE01, LB09]. Sea [LIE06, LSK^{+}03, LPAQ^{+}05, LCC^{+}09, LMBP07, LSSA04, LHB04, MLC^{+}07, MAMC02, MBLP01, MT06, MGRE^{+}07, MWH01, MCM06, NDT07, NA05, OS01b, PTK^{+}07, Plo08, RVP^{+}09, RSY04, RSH08, RG02, Sch00, Sch01, SSA^{+}04, Smi00, SM01, TA03, TT05, TKK^{+}03, VJ07, SVPK^{+}09, WMW^{+}08, WKW09, WGC00, WSBL06, XBD^{+}09, ZXN^{+}08, dBWF^{+}05]. sea-levels [BF08b]. sea-surface [FHB^{+}09, RSH08]. seabed [KJH03]. seabirds [CCM07]. Seaborne [RDPM06]. seafloor [KCCE06]. Seaglider [PSEL08]. Seagrass [VNH^{+}08, Ano08a, BMD06, BMT^{+}04, CSWD07, CT02, DAMM^{+}09, DZL^{+}03, EWE02, FF04, FMD03, Fyf03, GDM02, HPV^{+}00, JNKdV09, LMZB08, MPB^{+}08, PF01a, PVD^{+}07, SCDR07, SHBM01, ZB00, Zim03, vdHPH^{+}09, Fou02]. seagrass-vegetated [LMZB08]. seagrasses [DDZ03, VMSB08]. seals [BTB^{+}08]. search [JLH01, SSS06]. searching [KDO09]. seas [BGE04, JGE^{+}09, NML^{+}05]. seashores [Kne02]. season [HFC^{+}04, KKKT07]. Seasonal [AIS00, Bat02, CC01, Car02, CMD^{+}09, lLCChTO08, CSG^{+}06, Cro03, DD09b, Dav00, DHWW02, DAMM^{+}09, DBTK02, FMWC00, GGRJ03, GJS01, GNU00, HSB00, HM03b, HST^{+}02, KM02, LLS08, LBM^{+}04, MN04, PGC07, PKKB05, PDR^{+}09, Pil04, RNC03, RZMB05, SAH^{+}05, SJ08, TS06, TVSS01, THA01, TCCG05, VGP^{+}08, VCKS08, WŠN^{+}07, WS04, WKW09, AR06, AI03, BTS^{+}06, Cha07, CMJ07, GD00a, GCS06, JPJS06, LAG^{+}03, LB09, LKAB04, LCC^{+}09, MMKN08, MLPS09, PL09, RKNA09, SBM04, TRT04, UM05, VMH^{+}01, VRC07, WAB^{+}08, WSEL09, Ano08b]. Seasonality [Com02, FPZ09, AD00a, CH05b]. Seasonally [HAA09, HLCG07, MGL02]. seaward [LWP^{+}07]. seawater [AW05, BSM01, Bat02, DLZ^{+}09, DCC^{+}01, KLC06, KLS^{+}06, KLS^{+}07, LvdB09, LGN^{+}06, MOK^{+}03, NMK03, OS05, SRL^{+}09, SWF04, SBW07, TGdB^{+}01, TG03, WKH^{+}06, WKSM08, ZB01]. seaweed [ABFS09]. SeaWiFS [LAG^{+}03]. Secondary [GHS05, REL06, SSAH05, TSA09, CL02, NA05, TS01]. secretions [LB00]. section [BW01]. sector [AO09]. SEDEX [AD00b]. Sediment [AD05b, BF08b, CGL^{+}04, DKA^{+}03, DT00, GWR^{+}09, TVL^{+}00, ZC02, ACHG07, BDvO^{+}03, Bat00, BRW07, BRB^{+}01, BNM^{+}09, CSWD07, CV05, CCCC07, CCC09, DR00, DRSL07, Dup00, FS00, FT03, GM03a, GRdL^{+}08, GRB09, GJT^{+}08, GM04b, GTS^{+}09, GSB^{+}09, GVT^{+}09, HFDO05, HR00c, KB01, KCSM07, KBE^{+}04, KA06, LSPC02, LBB^{+}04, LRMV04, LMMW03, MW02, MCJ^{+}06, ME06, MA08, MGF03, Mil02, NCO02, NHPHH08, NHTF01, PGP01, PTLJ01, PTKP00, PF01c, RKH03, RHJ02, RHJ05, RTT^{+}02, SPZ03, SPM04, SEC08, SFB^{+}09, SDJT03, SSAZ01, SA09, SDKZ^{+}09, SCDR07, SdB06, SBM04, TB00, TTSC04, TMG03, TN09, TMA03, VvOBM06, VEMM07, VTR03, WŠN^{+}07, WW02, vOMSM06a, vOMSM06b, vWSBT08]. Sediment-mediated [BF08b]. Sediment-water [CGL^{+}04]. Sedimentary [NK05, RMG^{+}04, GFJ^{+}02, MPdVG07, NSBA01, TB05, WCL^{+}07]. sedimentation [DTL^{+}05, Ham02, HHB05, WSHN00]. sedimenting [KUP^{+}06, LBB^{+}04, RJS^{+}03]. Sedimentological [PD02, ME07]. sediments [ART^{+}06, AJ01, AD00b, Arn00, AHR^{+}01, BMD06, BM01, BVB03, BEM^{+}02, BZS07, BGJ01a, BGJ01b, Bou04, BBP^{+}06, BLS^{+}06, BZ02, BZH08, CT05, CYN05, CM03, CR06, CWG^{+}07, Dal03, DSD02, DKRZ06, DNS06, EPP^{+}08, EPD09, FSSP02, FDH^{+}01, FPPH06, FSK^{+}04, GGDC04, GvRE^{+}02, HBH^{+}07, HHS05, HPAM01, HOFD01, HKD01, HR00b, HAES02, HRS04, IKT03, JFH^{+}05, JHW05, JNKdV09, JM08, KSM06, KF01, KWS06, KMHN07, KGP^{+}02, LCS02, LH05, LNJL09, LMZB08, LMMW03, MMW06, MMDW09, MSSB06, MD04, MKV01, MV03, MGGM06, ME07, MGMW06, Nil03, NEC03, OH08, OKO^{+}01, PFPH04, RB09, RW02, RHJ04, RHJ05, RTT^{+}02, RGRPD04, RJS^{+}03, RBJ^{+}06, SLE04, SSK06, SD00, SL02, SR02, SR04, SDKZ^{+}09, SDH02, SrL05, SLRM03, SL00, SMH^{+}02, TPFS09, TB00, TKBJ05, TSML07, VHL^{+}04, VPH^{+}07, VM06]. sediments [WMWV04, ZVRL03, dBWF^{+}05, dBSN^{+}06]. seed [SDH02]. seep [BFS^{+}04, TBHS01]. seepage [DT00]. seeps [LM02]. sei [BF08a]. seiche [LMB05]. seiches [RS02]. selected [XLF^{+}03]. Selection [RB08, Alv07, BD07, HMS^{+}04, LLS08, PVD^{+}07, SD05]. Selective [DHTH05, HFC^{+}04]. selectivity [HK06, IHH^{+}00, MVFD04]. selenides [BFDC01]. selenium [BFS02, BFD^{+}04, CBG01, MC06]. semi [DB07]. semi-labile [DB07]. semienclosed [GGRJ03]. semivolatile [BSS^{+}01]. sense [GF02]. sensing [BDNF01, DZL^{+}03, GBA00, HM03a, HTWG08, KDS03, Kut04, LRS^{+}03, SPG05]. sensitive [FHDKR03]. Sensitivity [dSFFTD07, BN01, ESA06, GA00, JB00, LNB02, SWN08, SNPK^{+}09, vdPVB07, vDBL05]. sensor [MMW06]. Sensors [DIMP08, JN08, MJR08]. sentinel [MVA^{+}09]. sentinels [AOZ^{+}09, MLRMR09, Sch09, WSVS09]. separating [SA09]. separation [CMJ07, FTY^{+}06, GDL^{+}00, SMA^{+}06]. sequence [RSS^{+}00]. sequences [CHW^{+}01, RV04, vdSvdSG^{+}00]. Sequential [AD00b, KBE^{+}04, SABA02]. sequestration [SBZ^{+}09, vWSBT08]. Sereda [ACD10]. series [AHHH^{+}07, GHN^{+}02, HTWG08, MCPP06b, MVC^{+}05, MN04, SKBC01, SRKK08, SON^{+}03]. serotype [TP03]. seston [BDK^{+}06a, CV05, NPE08, RKNA09, SAE^{+}08, TNM^{+}01]. sestonic [HABF03]. set [BJBA03]. Seto [TMYH02]. settlement [BFCLÁ08, CTR^{+}02, Dix00, FNM07, MBDL01, PVD^{+}07, RKK09, SLKP05]. settlement-stage [Dix00]. Settling [FSSV08, NKE08, GWL^{+}07, Gre09, GDL^{+}00, KNE08, PWS09, PF01b]. seventy [GSG^{+}05]. seventy-two-year [GSG^{+}05]. severe [FDP^{+}08]. Sex [AAD08, ASH^{+}06, DHKH05, GRDCS09]. Sex-related [AAD08]. Sexual [HD09, KS04b, SS03]. sexuality [Gil03]. shade [KNC^{+}09]. shading [WLP^{+}07]. Shallow [Gre09, Ack03, AJ01, BP03, BCSW07, BHMS06, CC04, CHH07, CFMH01, CWG^{+}07, DTFL03, EM02, Fol05, GA00, GSH04, GHJ^{+}05, HGB00, HKNG08, HTWG08, JPJS06, JJS^{+}03, JKLR08, KB01, KJH03, LTOS04, MS03, MZV03, NHPHH08, OS01a, PCvRR00, RTT^{+}02, RC05a, SLRM03, TG03, VMS^{+}03, WŠN^{+}07, WS01, WR03, WJA^{+}07, WLRB03]. shallow-water [JKLR08, WLRB03]. shape [CLC02, SCH07, TFNT07, WW00]. shaped [GCG^{+}02]. shaping [GHJ^{+}05, WNAW09]. shares [RSS^{+}00]. Shear [LPW05, AH01a, GGIAY08, JTL01, KBBJ00, MHKP02, RG06]. Shear-induced [LPW05]. shearing [BYF09, SMS09]. Shedding [BB09]. shelf [CFG07, D’A08, FLT02a, GZL02, GSO03, HCC^{+}02, HRGM07, HHM^{+}09, JGE^{+}09, LDC^{+}01, MSSW^{+}07, MBR08, MSH^{+}06, PGP01, PN06, STG^{+}07, TN09, YSO02, CPB^{+}08, DGCMK00, DCC^{+}01, GPHS08, WCB^{+}06]. shelf-slope [MBR08]. shell [HM02]. shells [ASS03, BFML05, TCL^{+}09, TBHS01]. shelves [THL^{+}09]. shield [TXH05, CPV00, HBM06]. Shift [HLK^{+}07, JJS^{+}03, Leo08, MGRE^{+}07, PPH03, RNC03]. shifting [FGM^{+}03]. Shifts [PVD01, BCSW07, GP03, HLNM^{+}03, JSJ03, MVA^{+}09, RKNA09, RH03, Sak00, SSvNH01, SSvNH03, WŠN^{+}07]. Shingobee [SDJT03]. ship [HLP08]. shoaling [AO09]. shoots [New01]. shore [HH03, Par01, SMGP06]. shoreline [PBGJ09]. shores [ABFS09, DMS^{+}03, WL06]. shoreward [Ano08b, CSG^{+}06, SLB^{+}00]. Short [JBvHS08, SAHS06, CSWD07, DOE^{+}06, GCG05, HH05, KH03, LAN08, MSAC06, MSPAE06, SKTD04, TRT04]. short-circuiting [LAN08]. Short-lived [JBvHS08, KH03]. Short-term [SAHS06, CSWD07, DOE^{+}06, GCG05, HH05, MSAC06, MSPAE06, SKTD04, TRT04]. show [TMP02]. shows [ES00]. shrimp [BCMR00, DD09a, HMD^{+}07, PH07]. shrimps [GSA^{+}05]. Si [BDK^{+}08, MH07]. Siberia [RJS^{+}03]. sibling [CPCS01, CPCLS04]. side [GP03, MBG^{+}05]. siderophore [AGMP04, MHB01]. siderophores [BRT^{+}03, IB05]. sides [FLT02b, JS03b]. Sierra [SMC03]. signal [GA01, LH08]. signals [TSML07]. signature [CWC^{+}07, MF03, MdG08]. signatures [AHM^{+}08, DKA^{+}03, HBF^{+}06, LWP^{+}07, MM03, STHG07, TR03, WSC^{+}05, WSC^{+}06]. Significance [ASP04, BA01, MCF^{+}04, ZW06, BCBS04, DK06, EKD^{+}07, JM08, TS00, WNL^{+}08, vOMSM06b]. Significant [OS05, OFKH01, SSH^{+}09, UMASG07, MPACG01, TIH^{+}01, WS01]. significantly [RFW08]. Silica [EKdF07, BBL^{+}03, BSSL03, BJBA03, BJD05, Kim05, KFN00, LVB08, SSK06, SDV^{+}06]. Silicate [HTS^{+}03, BS03, CSPW00, HA01, RCL^{+}02, TvdWdB04]. Siliceous [MCV^{+}05]. silicic [BJD05, BDK^{+}08, FGM^{+}03, FSBB05, LBC^{+}04, LLC^{+}09]. silicon [Alv07, BA01, DP04, MCV^{+}05, MVBM04]. sill [AELS04]. silt [JM08]. silver [AGL03, SDC00]. silver-110m [AGL03]. similarities [SMA00]. similarity [RDPM06]. Simple [OIH^{+}03, FB03, GRCS06, Har01, KBBJ00, Mor03]. Simulated [TKM07, DKKA09, NCO02, WW03]. Simulation [PGM^{+}09, WCGBG07]. Simulations [FS09a]. Simultaneous [TPFS09, HLH09]. Sinauer [SG02]. single [HSN^{+}03, Hol00, vE02, GGP01]. Single-cell [GGP01]. single-celled [Hol00]. sink [ACZ00, CT04, CWWS00, DBE08, KGYU06, LBG^{+}09, MCV^{+}05]. Sinking [ET08, FMS04, WGLT05, All00, BYF09, NFFK00, PG00, Plo01, PHK02, PIF08, PIKB08, PF01c, PDB03, SSH^{+}09, SMS09, ST03, SVB^{+}08]. sinks [AW05, RS01, SDG^{+}07, dGvD07]. SIO1 [RSS^{+}00]. siphonifera [LEML09]. site [GHN^{+}02, MVC^{+}05, MN04, NHTF01, RV04, VCKS08]. site-specific [NHTF01]. sites [BEM^{+}02, CTR^{+}02, JHW05, KLAF04, TBHS01]. situ [BD04, BMD06, BLM^{+}06, CDH^{+}08, Com02, DCC^{+}01, DHWW02, Dix00, ETCP03, FDH^{+}01, FBW^{+}08, GGP01, GTK^{+}01, GSB^{+}09, GFK00, GBH^{+}02, GvRE^{+}02, Har01, HLL^{+}06, JPK^{+}05, LW01b, MBB^{+}00, Mor03, NKJM06, NHN^{+}05, PSH^{+}07, PW01, RHJ05, SGPK07, SLDE09, SWML08, TPFS09, TMA^{+}01, VvOBM06, VEMM07, WLRB03, YSM^{+}03, YWR^{+}07, dBSN^{+}06, vOMSM06a, vOMSM06b]. situ-produced [GvRE^{+}02]. six [BPF^{+}08, FLWP09, LRF07]. Size [DJO00, De 02, SK01, VB00, BCV^{+}08, BR00, CSSNA02, CBRC01, CEBMF05, CPHM04, CW01, CPH06, CLC02, DPC^{+}06, Dup00, Fin01, FSBB05, GD00a, GSJ^{+}09, HBH00, HC01, HWYL07, HPL^{+}06, Hol00, HH05, KTLS03, KLG09, KRZ06, LBC^{+}04, MCR^{+}07, MBDL01, MPAS00, MBC01, MY05, PG00, RLP^{+}06, SMA00, TCC06, TG09, TNM^{+}01, TFNT07, VMS09, WPD^{+}01]. Size-dependent [DJO00, De 02, FSBB05]. size-fractionated [CEBMF05, HWYL07]. size-spectra [Dup00]. sized [Ham02, SI03]. sizes [KFR^{+}01]. Skagerrak [PF01b]. skeletal [ACW02]. skeleton [TMK^{+}07]. Skeletonema [AMP^{+}00, BB04b, SNM^{+}08]. skeletons [EMIP05, LRYS06, WCGBG07]. skvortzowii [JGZ^{+}08]. sleep [GHWG04]. sleep-swimming [GHWG04]. slipper [MTC^{+}06]. slope [DdG09a, GP03, HSR^{+}08, MBR08, TN09, YSO02]. slope-water [GP03]. slopes [DdG09a, HSR^{+}08, LBD^{+}09, LPW05]. sloping [BII05, MS03]. sloppy [Møl07]. Slough [BNM^{+}09]. Slovenian [VM06]. slow [YPBL07]. Small [MHKP02, Plo01, Plo08, RKWG00, WG04, CC01, CSM^{+}08, CCKP02, CCGH05, DET04, FSSV08, FUG02, GWH06, GHTJ00, HPV^{+}00, HMD^{+}07, LH05, LH08, MSGK06, MKG04, RHJ02, RM09b, SCB03, SHBM01, TGdB^{+}01, UMASG07, WSL^{+}02]. Small-scale [MHKP02, Plo01, Plo08, RKWG00, WG04, LH05, RHJ02, SCB03]. smelt [BCBS04]. smelters [CBG01]. smooth [New01]. snail [GGH07]. snapper [BW01]. snapshots [BTB^{+}08]. snow [FS09a, GYBA04, Kiø00, KJ01, LWU^{+}05, PIF08]. snowmelt [PVD01]. sociale [KTMM00]. sockeye [BBP^{+}06, SFB^{+}09]. SOFeX [BJD05, BAP^{+}05, OBSD04, TBF04]. soft [HM02, RHJ05]. softwater [GCS06]. soil [Cra07, HLL^{+}06, PS06, TSML07]. soil-derived [PS06]. soils [TSMB05]. SOIREE [KTLS03]. Solar [MSAC06, CMSVA02, DTFL03, OZM^{+}01, ZSS^{+}08, ZB00]. sole [CD02]. solid [SC00]. solid-water [SC00]. solitary [DK06]. SOLOPC [CDH^{+}08]. solubility [NKS^{+}02]. soluble [CT05]. solute [CWG^{+}07, JFH^{+}05, JHW05, KJ01, MV03]. solutions [FB03]. Somali [IOG^{+}04]. somatic [MBK^{+}09, RB06, VPA02]. Some [SGR08, SCARM00, SAS^{+}03, HKF01, JSJ03, MF01, MF02, KKH04]. Sorptive [AHR^{+}01]. Sound [NK05, TT05, MMKN08]. source [BMD06, BGJ01a, BGJ01b, CWWS00, CLG08, CWC^{+}07, DdG09a, FCJ^{+}06, GLP^{+}08, GJS01, HSR^{+}08, JR07, LBC06, MHDH00, MCL06, Mul01, NDT07, PSS^{+}06, PBGJ09, SAP^{+}06, SDKZ^{+}09, TIH^{+}01, TNO^{+}08, WNL^{+}08, WKSM08, YSM^{+}03]. source-age [LBC06]. sourcebook [Ris02]. Sources [AW05, CLPB02, CCF04, DNS06, EB01, LW01a, MH00, RS01, SDG^{+}07, TM02, WIK08, AM08, BFCLÁ08, DBTK02, GSB00, HWYL07, KTM01, KKS^{+}01, KCP^{+}04, MBCD04, MdG08, ML02, MPRC02, MC04, OCNH00, RB01b, SSS02, SKR01, TMYH02, TSML07, WKFP06]. South [HLD^{+}03, Mil02, PF01a, StLHB^{+}07, AW00a, AW00b, BZS07, lLCChTO08, CLL^{+}09, DAL02, FCGC^{+}09, GHS05, HWYL07, MJR08, SHBM01, WJMD07, WBB09]. Southeast [HCK02, HMS^{+}07, HCC^{+}02, JBvHS08, ML05, NBW^{+}02, SL02]. southeastern [CWWS00, NSH^{+}09, OCNH00, PBC02, XBD^{+}09, ZZPP01]. Southern [Dav00, GRG02, AO09, CSM^{+}08, CG01, CPB^{+}08, DORS08, EB01, HSB^{+}06, HMR^{+}07, KBG^{+}08, LLvW^{+}03, LMBP07, MPAS00, PLK^{+}05, Par01, PF01b, Pil04, PN06, WSBL06, AD05a, AWLH04, BMH^{+}08, BJD05, BAP^{+}05, CLG08, CCM07, CCC^{+}03, DMEPA04, DAV^{+}05, ET08, HPL^{+}06, KTLS03, KMvD^{+}09, MBL^{+}01, MSPAE06, OBSD04, ORFD^{+}09, SAS^{+}03, SR07, TGdB^{+}01, TvdWdB04, TPG^{+}08, TBF04, WAGH09]. southwest [EMPI07, HRGM07, KTAL04, MBPVR06]. sp [BVG^{+}07, BMCvE06, BWK^{+}07, EDF^{+}03, IB05, NCC00, PSJ00, Plo08, RG06, SWD^{+}05, SEL^{+}02, WŠN^{+}07]. space [RLM^{+}08]. spacer [WGKE04]. Spain [DAMM^{+}09]. Spartina [SBL06, GK03]. Spartina-derived [SBL06]. Spatial [BFML05, BLS^{+}06, Fyf03, GRR^{+}07, GGIA06, JBM^{+}07, KTM01, KRZ06, LIHS03, MRK02, MOK^{+}03, Met04, NCC00, ORH01, PGL06, PLMPN08, PLM^{+}09, SAS08, SSA^{+}04, WBP^{+}06, ZSS^{+}08, vNHM^{+}09, BSJ02, BBLB05b, BBLB05a, CSM^{+}08, FLWP09, FSSV08, FHBK06, FTY^{+}06, HTRM^{+}07, HTWG08, JN08, LSK07, MVC^{+}05, OD08, Ost03, PKKB05, RWFK00, RZMB05, RM09b, RYS^{+}02, SPW^{+}05, SGPM09, SSL05, VWPP^{+}06, WAB^{+}08, WG04]. Spatial-temporal [MRK02]. Spatially [Ano08a, MPB^{+}08, DJG^{+}06, LD01, MCPP06a]. Spatio [RLC^{+}01]. Spatio-temporal [RLC^{+}01]. Spatiotemporal [GGDC04, LSC^{+}08, Tho09, WBB09, YYTH07]. Spawning [GWH06, AH03, EGP08, GEP08, HFC^{+}04, ML05, SAH^{+}04, TCR03, YM03]. Special [DIMP08]. specialist [KHKS05]. Speciation [TLM04, BMU04, BLBB07, BB07, Cro03, HLCG07, HHB^{+}09, HD09, MJW03, PFN^{+}05, SDvEB06, TF00, UWA^{+}07, ZWHP04]. Species [KRM09, MBDL01, NHN^{+}05, UPC^{+}05, ARV^{+}07, BA01, CT04, CGD08, CPCS01, CPCLS04, CHT02, DGV01a, DVV^{+}07, ERH06, FAL^{+}07, Fyf03, GBL^{+}04, HZ00, Ham05, HC01, HSLV05, HMNL06, HH04, JS03a, KGYU06, KS04b, KWTS07, KLG09, KMMS00, KAW05, MDJ^{+}04, MMLH06, MPH08, MG04, MCL06, NPE08, PH07, RB02, RBW09, RLM^{+}04, SBS07, SP09, SS06, WHHS06, WM06, ZRL06, ZMK^{+}08]. Species-specific [NHN^{+}05, UPC^{+}05]. specific [BS04, DWA^{+}02, EAO^{+}02, EPP^{+}08, HMS^{+}04, HK06, KRM09, LALE04, MLC^{+}07, MPZ08, NHN^{+}05, NHTF01, RRH^{+}03, SCS^{+}08, TS00, UPC^{+}05, VHL^{+}04, ZWC^{+}04]. Specificity [Gil03, LBW^{+}08, SS06, Arn00]. specimens [CCKC05, OKO^{+}01]. spectra [CBRC01, Dup00, KRZ06, MMS^{+}02, MBC01, SLRM03]. Spectral [DRM^{+}09, MV00, NTDV06, BS02a, CMSVA02, CLC02, DKA^{+}03, DdG09a, ETCP03, Fyf03, HSR^{+}08, KDS03, LBD^{+}09, LRS^{+}03, MF03, OZM^{+}01, SWN08, VMS^{+}03]. spectrally [Fyf03]. Spectrofluorometric [MBW^{+}01]. spectrometry [FKCF00, KKH06, SHL^{+}05]. spectroscopy [BK01b, BK01c, HFE07]. Spectrum [VSSTS00, RDPM06]. speed [CW03, RNR08]. spherical [SJRB09]. spicatum [LL05]. spinicauda [XW02]. spinifera [JHGGP06]. spiny [BFCLÁ08, GB09]. spionid [HL06]. spiraling [EVW07]. split [GDL^{+}00]. SPLITT [GDL^{+}00]. Sponge [PSH^{+}07, EKdF07, SWML08, YSM^{+}03, dGMH^{+}08]. Sponges [JR07, RCAK05, HMSI06, MCV^{+}05, YWR^{+}07]. spores [BB04a, Duf09]. spp [BFBHF04, EAO^{+}02, JL08, MSBHG01, MBH^{+}06a, MBH^{+}06b, OLCT08, SvKM01, SGPM09, SMC08, SGG^{+}08, TSSWK^{+}06, WSKL06]. Spring [LFS^{+}04, RNA06, STG^{+}07, BJG^{+}00, BSSL03, DJB^{+}07, DPFL09, DKA02, GA00, GPHS08, GBM^{+}04, HNB^{+}07, KFN00, LRF07, PSLO07, PF01b, ST03, SLJ00, SKC^{+}01, SKH^{+}01, SMG06, WMW^{+}08, XBD^{+}09, ZXN^{+}08]. Spring-neap [STG^{+}07]. springs [KR08]. springtime [CTY06, SP09]. spumigena [KSEÖ^{+}02, Plo08]. SSU [PCDB09]. St [LBG^{+}09]. St. [BCBS04, GSG^{+}05, KCSM07, SMJ^{+}06]. Stability [Law08, GBH^{+}02, JLB^{+}06, LBW^{+}08, MRM07]. stabilizing [WKSM08]. Stable [CCH02, HBF^{+}06, KF01, LW01a, LBG^{+}09, NH06a, NEC03, SPR02, YTVR07, BP03, BRB^{+}01, BCMR00, BLS^{+}06, CCKC05, CPB^{+}08, CCCC07, CCC09, DBE08, DBS^{+}04, DPC^{+}08, EVW07, EPP^{+}08, Fin04, GR01, GJS01, GKJ04, GCS06, HPLG09, HL02, HM06a, HB06, HVP^{+}07, IR06, MM03, MBR08, MLPS09, OHCP02, PKKB05, PDH06, PLM^{+}09, PCD^{+}08, RH01, SAS^{+}03, SAP^{+}06, STST07, TR03, SVPK^{+}09, VM06, vOMSM06b]. Stacey [Caf01]. Stage [EAO^{+}02, ASH^{+}06, DLDE04, Dix00]. Stage-specific [EAO^{+}02]. stages [JGZ^{+}08, LWS01, MSBHG01, SvKM01]. stagnant [AVJ^{+}00]. stamukhi [GLP^{+}08]. standard [STST07]. Standing [KED^{+}07, PCA^{+}00]. star [VTR03]. starvation [DP04, RRCP03]. state [AFB01, CNGV09, Dod06, FB03, GWR^{+}09, HAG08, JJS^{+}03, Leo08, TWC^{+}09]. states [BP03, BCSW07, LCLP02, MWW06, BHV04, HCC^{+}02, PDS07]. static [AH01a]. Station [BK03b, CDK02, Gre09]. statistical [LW01a]. statistics [SGPM09]. statoblast [FRO01]. status [BK01a, BS01, CSSNA02, DGM^{+}05, FWM^{+}05, GGP01, GLVFH09, MLC^{+}07, SSH^{+}07, SSV^{+}04]. Staying [Duf09]. Steady [FB03, Leo08]. Steady-state [FB03]. stealth [HW09]. Stechlin [CCCC07]. stella [GSFB02]. Stentor [MBC^{+}05]. Stephanodiscus [Hol00]. sterol [TGFPT08]. sterols [CYN05, MCvEH08, TTG^{+}09]. Steven [SG02]. stimulants [CT06]. stimulate [GRVF^{+}00]. stimulated [CJB00, MLSS05, dVKJK07]. stimulates [VJ07]. Stimulation [CM03, DGKA06]. stimuli [SCH07]. Stinson [dSYL^{+}08]. stock [AH03]. Stocking [LRS^{+}03, ZVRL03]. stocks [Aks07, KED^{+}07]. Stoichiometric [AP00b, HW08, dSFFTD07]. stoichiometries [TBF04]. Stoichiometry [MC01, AKE04, ABJ^{+}06, ASAC04, BDK^{+}06a, BBKC03, CG08, DTL^{+}05, FQR^{+}06, FH09, FXL04, HAL^{+}09, HK01, KLL04, KFW09, NPE08, Pri05, QMS^{+}07, SBS07, SAE^{+}08, SSS08, SBZ^{+}09, WGKE04, WSKL06]. stomach [CCM07, DPC^{+}08, SAP^{+}06]. storage [CSE^{+}08, ED05, GMRD04, HBL02, HBAK09, SBM04, WLS00]. storm [CWLH06, GJT^{+}08]. storm-induced [CWLH06]. stormflow [WTM09]. storms [KBE^{+}04]. stormwater [DC05]. story [HSS06, Sos00]. strain [BBG^{+}05b, CLPB02, RCE^{+}02]. strains [BPF^{+}08, HSN^{+}03]. Strait [SWD^{+}05]. Stramenopiles [PCDB09]. strategies [BFQF^{+}07, BL07, BHL07, CT04, JS03b, MD04]. strategy [MPGB08, RL01a, SD05, VK06]. Stratification [WS01, BD04, BLHC08, CR10, LTHW04, MSGK06, RAKE05, RS09, SKTD04, VMH^{+}01]. stratified [AIS00, AIK04, BIIA03, BI07, CGCPA08, CR10, DET04, EDF^{+}03, ESI01, GGIAY08, HISW00, KHP^{+}06, RAKE05, RENA06, SI01, SVC^{+}07, SIK07, SI08, SFL05, WPV00, ZHFD06, dlFSIN08]. stratospheric [HSF^{+}08]. Stream [MVW^{+}04, AS06, AMS^{+}08, AB07, AGKK08, BS02b, BHV04, CH05a, CJWR05, EW07, EB01, FVD01, GMRD04, HBAK09, HTS^{+}09, HFC^{+}04, HDP09, HFR09, HFE07, HPBD01, KDO09, KTM01, KD02, KKH06, LNB04, LVSW07, MSLH02, ML02, MHS^{+}09, NSM09, NTM^{+}05, RH01, RP06, RKNA09, RM06, RM09b, SAH^{+}04, SA00, SPM04, SHL^{+}05, SPW^{+}05, SSAH05, SL01, TMS^{+}00, TSA09, WW00, WWMB05, WPD^{+}01]. streambed [Bat00, SrL05]. streamflow [GN00]. Streamlining [AD00b]. streams [AH01a, ATJH08, BMC07, CGWK^{+}09, Dod06, EVW07, FVD01, FQH^{+}01, GP06b, HBL02, HT03, HBM08, HTRM^{+}07, LFZ^{+}07, LRMV04, LZ09, MJEM06, MW06a, ML05, MTM^{+}01, MBC01, PWH^{+}04, RB03, RM06, RGAS04, RPR^{+}02, SNFT07, SFC^{+}00, SLJ00, TMP02, TNM^{+}01, Sta01]. strength [Car02, EDF^{+}03]. strengths [HSS06]. stress [AH01a, CTR^{+}02, CMB00, DOE^{+}06, DWA^{+}02, HH03, HMR^{+}07, HMNL06, HCD^{+}09, LPAQ^{+}05, LSSA04, TSKA05, WCGBG07]. stressors [HR00a]. Strike [HW09]. striped [BFS02]. stromatolite [BdBS^{+}08]. Strong [AD00a, Alv07, LWS01, AKH04, MVFD04, SWS^{+}06]. stronger [SLC^{+}09]. strongly [JBH^{+}04, SIK07, SI08, SFL05, VVD^{+}09]. Strongylocentrotus [GWH06]. strontium [LGN^{+}06]. structural [FSSH03, GCJ^{+}06]. Structure [IB05, Ano08a, BKM05, CSWD07, CGH^{+}04, CWJ09, CCKC05, CPH06, CGWK^{+}09, DVV^{+}07, FH09, FWM^{+}05, FTC^{+}06, FTY^{+}06, Goe02b, GGIA06, LJ01, Law08, LFAB06, LVSW07, MDJ^{+}04, MKG04, MPB^{+}08, MGVM^{+}06, MY05, NSM09, NHTF01, OAA^{+}06, PCvRR00, RMK^{+}06, Rit09, SABA02, SIK07, SK01, SJ00, SL01, UHB^{+}08, UVY^{+}01, VDK04, WBB09, WWWY05, WLS00, ZLP^{+}03]. structures [BTB^{+}08, BJL01, Cra07, DAL02, HDP09]. structuring [DTFL03, HTS^{+}03, PH07, RM06, WA01]. studied [DMEPA04, GRdL^{+}08, GRB09, HAHS^{+}09, YWR^{+}07]. Studies [Dow02, BM00, Ben00, DGL^{+}04, HAES02, LA09, RSY04, SLD00, SF03b, SF03c, VR01]. Study [ET08, GHN^{+}02, Lon00, MN04, ART^{+}06, BGR^{+}09, Bea03, BRB^{+}01, BSJ02, BDK^{+}06b, CCAÆ06, Cha07, CG08, COK^{+}08, Cur00, EWE02, EKD^{+}07, FPPH06, FG06, GSB^{+}09, GHB^{+}05, HTWG08, HRS04, JCW08, LBB^{+}04, LH05, LBSY02, MMW06, MSAC06, MBB^{+}00, NCO02, NSH^{+}09, OAA^{+}06, PLK^{+}05, PHMR^{+}07, RVP^{+}09, RH01, RRSC01, RB02, SKBC01, SHS03a, TB00, TVSS01, TMA^{+}01, TWC^{+}09, VvOBM06, VEMM07, WdMD^{+}00, ZRL06, vOMSM06b, MVC^{+}05]. studying [RL00]. Stylophora [GRVF^{+}00, GFPG09, GMRVFP02, GMAFP03, Shi04, WLRB03]. sub [MML^{+}08b, RES^{+}08]. sub-populations [RES^{+}08]. sub-tropical [MML^{+}08b]. subalpine [BSS^{+}01, GP06a, KLWS08, PBMSG04]. subarctic [Aki00, AHM^{+}08, BST^{+}05b, CLP^{+}03, FSS^{+}00, HSB^{+}04, KB05a, KAJ08, LCPS05, RLP^{+}06, RK02, RCVK09, RWK08, SJR^{+}04, SLS^{+}02, TMC06, VDK04]. subcellular [ZW06]. subestuary [CB04]. Subglacial [CRBF^{+}06]. subglobosum [HHB05]. Subitaneous [BSSL05]. sublethal [LL05]. sublittoral [GHM^{+}09]. Submarine [CB04, PSS^{+}06, dSYL^{+}08, Cha07, GGM03, HLK05, KC08, SSA^{+}04, SMGP06]. submerged [RKH03]. Submersed [BP03]. submersible [SON^{+}03, YWR^{+}07]. Submesoscale [NCPd08]. submicron [VB00]. Submillimeter [BVB03]. suboxia [WGC00]. suboxic [FS00]. subpixel [HM03a]. subsaline [DHP03]. subsidies [Wur07]. subsidize [MIDH05, PCC^{+}07]. subsidy [MID^{+}07]. substance [HBH^{+}06]. substances [HPHT00, LvdB09]. Substantial [KB01]. substrata [KD02]. Substrate [Arn00, BZ03, PIF08, WR03]. substrates [JP03, PBGV04, RlBW01]. substratum [RGAS04, WWS^{+}00]. Subsurface [FB03, HB08, LWP^{+}07, PSEL08]. subterranean [SBC^{+}08, SCM08]. subtidal [CWG^{+}07, NH06a, SLRM03, SHS01]. Subtropical [BNK02, BK03b, CDK02, CLAK08, GAM^{+}01, HPLG09, LKAB04, TM02, ACW^{+}03, ASGA^{+}07, ÁS07, BAG^{+}09, CWLH06, CPL^{+}04, DBTK02, DAA^{+}01, EM02, EF05, GDB^{+}09, HZG07, HDWDD09, HOFD01, KFH^{+}09, MYM^{+}05, Mar05, MML^{+}08a, MPZ08, NEE08, ORH01, TSSWK^{+}06, TSDK^{+}09, TCCG05, VD08, VEMM07, ZMJ^{+}07, ZB00, dCCL09]. subzero [YCM^{+}01]. success [BWK^{+}07, LED06]. successfully [KAW05]. succession [AR06, ASI^{+}02, LFS^{+}04, RB02, RNA06]. successions [dGB02]. suction [HW09]. suction-feeding [HW09]. Sudbury [CBG01]. sufficient [LPAQ^{+}05]. sugars [BK03a, RA06]. suggest [AWE07, FRO01]. suitability [vdHPH^{+}09]. Sulawesi [SHBM01]. sulfate [LFS^{+}02]. sulfate-polluted [LFS^{+}02]. sulfide [CGCPA08, CCF04, EMWL06, FHBK06, JNKdV09, PMPC00, RKGC^{+}01, RA09, SSS06]. sulfide-based [PMPC00]. sulfide-rich [CGCPA08]. sulfite [HTAS06]. sulfur [CT05, DBS^{+}04, FHBK06, GK03, KA06, LPD^{+}04, PLK^{+}05, RTT^{+}02, VCKS08, ZHFD06]. sulpho [UGSMN01]. Sumatra [AJS07]. Summer [CHH07, AGL03, BGHA04, CC04, HM08, HDP09, KKKT07, KTAL04, LHWE01, LSK^{+}03, LP07, MSGK06, MAC^{+}03, MLSS05, NML^{+}05, PTK^{+}07, PBC02, STG^{+}07, SKTD04, SR07, SLS^{+}02, TKK^{+}03, WDVA08]. Sundarban [BMD^{+}04]. sunlight [PBGV04]. superba [ASAC04, AW00a, AW00b, AMS^{+}02, ASH^{+}06, MAS^{+}02, MFG^{+}09, RQB^{+}00, SAP^{+}06, SHS03a, TSH^{+}06, TKM07]. Superior [AC08, HHB^{+}09, MHB03, PED^{+}09, SSM^{+}04, ZZPP01]. superoxide [KSM^{+}05, PP04]. Supersaturated [PCD^{+}08]. suppl [Wan02]. Supplementation [Has04, VW01, RBMN03]. Supply [AM08, BDK^{+}06a, BCDH09, BLBB07, CV05, DDL^{+}09, FSSH03, FSB^{+}08, FNM07, GM03a, HK01, KB05a, LG04, MSGK06, MMDW09, MW06a, Met04, MBCB08, MTV^{+}02, RGAS04, SCARM00, SMGP06, SKBC01, SBR^{+}06, SRKK08, SH05b, TCD^{+}03]. supplying [HCK02]. support [GMA^{+}06, PED^{+}09, ZHL06]. supported [Ano04, KSWC^{+}03, WBD02]. supporting [DBTK02, MBCD04]. suppression [BF08b]. surf [DMS^{+}03, Gay00]. surf-zone [Gay00]. Surface [ARB^{+}04, RS02, vHRB^{+}09, ASGA^{+}07, BRW07, BZS07, BLS^{+}06, CV05, CFO^{+}00, DLZ^{+}09, DJH05, FAC05, FHB^{+}09, GM03a, GSL^{+}07, GDL^{+}00, HPT^{+}09, JS03b, KFR^{+}01, KTAL04, KLAF04, Lee01, LAN08, LD01, LP07, LNJL09, MRK02, MMW06, MW02, MF04, MLSS05, OD08, OS05, ORH01, PPH03, PN06, Plo08, PH03, RSH08, RHJ05, RWK08, RJS^{+}03, ST04, SLE04, SVC^{+}07, SAPA^{+}02, SI03, THWS00, TWS01, TWS03, TP03, UGSMN01, WBH03, WS01, YSO02, dBSN^{+}06]. surface-water [LP07]. surfaces [FZL^{+}05]. surges [Mor06]. surprise [HW09]. surrogate [BCM01]. surrounding [LB00]. surroundings [MHC05]. surveillance [DORS08]. surveys [NEE08]. Survival [BSSL05, PDS05, AY08, GWK01, SMR^{+}01, TG09]. Suspended [PE07, AGL03, BS04, BAG^{+}09, DKRZ06, HHS05, JJP04, KUP^{+}06, NHS07, SWF04, SBW07, WPD^{+}01, YYG02]. suspension [BB04a, CWLH06, Com02, NHTF01, RHM04, SLD00, TS01, WSNM00, WLTF03]. suspension-feeding [NHTF01, RHM04, SLD00, WSNM00]. suspensions [BVN^{+}00]. sustains [VTO^{+}05]. SW [GST^{+}01]. Sweden [BRB^{+}01, CSPW00, Cro03, JMBJ01, KJMJ03, RMG^{+}04, Wey04]. Swedish [BBJ05, HSB^{+}04, WGLT05]. sweep [OH08]. swim [CT04]. swimming [BGC04a, BYF09, GHWG04, GYBA04, RG06, SMS09]. swinhoei [YSM^{+}03]. Swiss [WW00]. Switzerland [AHHH^{+}07, HAHS^{+}09, PWS09, TB00, TB05, TMUW02]. Switzerland/Italy [HAHS^{+}09]. Sydney [DCMS03]. Sylt [WBP^{+}06, dBWF^{+}05]. Symbiodinium [GCG05, MG08]. Symbiont [SMP05, FBW^{+}08, GCG05, LLvW^{+}03, LEML09, YSM^{+}03]. symbiont-bearing [LEML09, YSM^{+}03]. symbioses [LBW^{+}08]. symbiosis [GCG05, PSH^{+}07, PMPC00, VTO^{+}05]. symbiotic [EMIP05, FMWC00, FSM^{+}07, GKLF01, GF02, WLRT06]. sympatric [CPCS01, GRDCS09, MPH08, MH04b]. synchronised [CTWFP06]. Synchronization [TCR03]. Synchronized [BCBS04]. Synchronous [PC02]. Synchrony [CH05b, Wey04]. synechobactins [IB05]. Synechococcus [BVG^{+}07, BBKC03, CBvEK03, HSN^{+}03, IB05, LPAQ^{+}05, MKVK05, MPRC02, MTV^{+}02, NCC00, SRM05, TP03]. Synergistic [CLF^{+}08]. synergy [WTC^{+}05]. synoptic [LP07]. synoptic-scale [LP07]. synthesis [Lon00, LSF^{+}08, SDG^{+}07]. synthetic [Bia01]. system [ÁSGMP01, ÁÁS07, ABFS09, BMBA07, BGC^{+}05, CSWD07, CLF^{+}08, CCH02, CIWH09, DTL^{+}05, DGM^{+}05, GCJ^{+}06, HBG00, HBTM05, HLGI08, HRA07, HLM08, JFH^{+}05, JLP^{+}08, JBvHS08, MNA^{+}09, MSJMN02, NCÁSP06, PWN07, PN06, RB02, TMUW02, TB09, UMASG07, VVD^{+}09, WKH^{+}06, YTVR07, vHRB^{+}09, AO09, DORS08]. system-level [RB02]. systematics [SPR02, TIH02]. systems [BK01a, BW00, CL04, CDD07b, DAC^{+}07, HLL^{+}06, LN04, MSC^{+}04, MC04, NSH^{+}09, Nor04, PNPA04, Thi00, Elm02].

T [Caf01, Wan02]. tactile [SA04]. Taiwan [CWLH06]. Tana [BDSB07]. Tanganyika [MMO06, VH09]. tank [FPPH06]. tannins [MPJ08]. Tapajós [RLC^{+}01]. Tattersall [INU04]. taxa [KMvD^{+}09, MPZ08, PDB03]. taxa-specific [MPZ08]. taxifolia [CM03]. taxiformi [MHDH00]. taxiformi [?]FalkenbergiaMarshall:2000:RAA. taxiformis [PGC07]. Taxon [LALE04, MLC^{+}07, RRH^{+}03]. Taxon-specific [LALE04, MLC^{+}07, RRH^{+}03]. Taxonomic [BS02b, LSSA04, CPH06, DJT06, HSLV05, RYS^{+}02, VWPP^{+}06]. TDN [PWH^{+}04]. technique [EPS00, GBA00, Har01, MCL06, MSBHG01, NKJM06]. techniques [RL00, RBW09]. Temora [Møl07]. temperate [AR06, BVH02, BGE04, BdG02, Com02, CH05b, FLPL08, HMS^{+}07, HFE07, HPBD01, JR07, KCE^{+}00, KKH06, LN02, NFS^{+}07, OZM^{+}01, RB01a, RYS^{+}02, SM05a, SKC^{+}01, TYK^{+}08, TMB07, TMA03, WM06, XLF^{+}03, YWR^{+}07, dGP08, vdPAJ^{+}06, vdHPH^{+}09]. Temperature [Ann05, DAMM^{+}09, LEML09, MW06a, WGD^{+}02, ADA00a, ADA00b, ACHG07, AY08, ASH^{+}06, AC08, BMSA08, BRM^{+}07, BPF^{+}08, CF01, CG08, CHT02, DSW05, Fab06, FS09a, GCM^{+}09, HDSC09, HB03, JLB^{+}06, KSS^{+}07, LD01, MBK^{+}09, MAC^{+}03, MF01, MF02, MB05, PLG^{+}02, PSH08, RHK^{+}09, RC07, SSC02, VMKN03, WSL^{+}02, YSO02, ZJC^{+}07]. Temperature-dependent [WGD^{+}02]. temperatures [GCG05, HRA07, HDP09, KTAL04, LD01, LP07, MW06b, WDVA08, WK09, YCM^{+}01]. templates [PVD01]. Temporal [Agr05, ASI^{+}02, ARO^{+}04, CGD08, CLAK08, EGPA02, EF05, FNLG00, FHBK06, MVC^{+}05, RKNA09, RYS^{+}02, SSJ07, SSL05, BFML05, CSM^{+}08, FLWP09, FRO01, Fyf03, HTRM^{+}07, KRZ06, LIHS03, MRK02, MOK^{+}03, MM05, Met04, NCC00, Ost03, PLM^{+}09, RV04, RLC^{+}01, SPW^{+}05, VWPP^{+}06, VC05, WG04, WBP^{+}06, ZSS^{+}08]. TEP [ASP04]. term [CSWD07, CWC^{+}06, CJWR05, DOE^{+}06, FCJ^{+}06, GCG05, GVT^{+}09, HH05, INU04, JS03a, Kim05, LBB^{+}04, LIE06, MSAC06, ME06, MGRE^{+}07, MCB08, MSZ00, MSPAE06, PCvRR00, RBW09, RL07, SAHS06, SKTD04, SMH^{+}06, TRT04, VW08, VWPP^{+}06, WMW^{+}08, YYTH07]. termination [BST^{+}05b]. terrains [HA01]. Terrestrial [AKP^{+}09, AHM^{+}08, CCH02, HSB^{+}06, MIDH05, MPD^{+}06, PCC^{+}07, WIK08, Wur07]. terrestrially [SM01]. terrigenous [Dit04, JM08, SBW07]. test [AKE04, AD05a, BEM^{+}02, CMD^{+}09, DAV^{+}05, GWK01, LED06, MSBHG01, RBMN03, WSL^{+}02, WH07]. testate [vHRB^{+}09]. Testing [CD02, DET04, EZA^{+}05, dSFFTD07, GAPR06, HFC^{+}04]. testudinum [DDZ03, NKJM06, PF01a]. Teton [HT03]. Tetraselmis [EDF^{+}03]. Texas [BMU04, GMA^{+}06]. Th [MWE^{+}03, QSH^{+}02]. Th/ [MWE^{+}03]. Thailand [DBWS06]. Thalassia [DDZ03, NKJM06, PF01a, SHBM01]. Thalassinidae [DD09a]. Thalassiosira [BS03, DP04, EH00, JGSB07, LJCM08, LLA05, LB04a, MP00, SWN08, SSC02, TSKA05, vdPVB07]. Thalia [LIE06]. Thames [BDS^{+}04]. thecamoebians [vHRB^{+}09]. thecate [ZGP^{+}00]. Their [Gou01, MPJ08, AS06, All00, ASI^{+}02, BMF^{+}06, BF08a, BMNB^{+}06, BDC00, CSM^{+}08, CK00, DHTH05, FNLG00, GM03a, GRR^{+}07, KAM04, KFH^{+}09, LG04, LD01, LMBP07, MZL04, PKKB05, RKK09, RFM07, WS06, WCL^{+}07, ZWHP04, ZB03b]. Theonella [YSM^{+}03]. theoretical [BSJ02, GBFJ07]. Theory [FMD03, ZB03a, AP00b, CR08, EMK08, dSFFTD07, LU08, Thi00]. There [Kiø07, GH00]. Thermal [MCJ^{+}06, Ano08b, ABDS04, CSG^{+}06, HH03, JLB^{+}06, JKLR08, PGM^{+}09, RSC04, SJ00, TYK^{+}08, WBB09, WCGBG07]. thermally [CR10, ESI01, RAKE05, RENA06]. thermocline [HHM^{+}09, SMR^{+}01]. these [HMKZ02]. thickness [QYEA06]. Thin [SGG^{+}08, BBCW09, BB04b, DG06, GDL^{+}00, Has06, SMS07, SSM^{+}09]. thin-cell [GDL^{+}00]. Thiols [ZWHP04, DNB^{+}04, DA05, HMB06, THWS00, TSKA05]. thiosulfate [HTAS06]. Thiotaurine [PMPC00]. third [VCCS05]. Thomas [Leh00a]. those [LLvW^{+}03]. thoughts [SGR08]. threats [HLH09]. Three [RWFK00, SGR08, SF04, BDS^{+}04, Car01, Com02, DGV01a, DGV01b, HTRM^{+}07, HM06b, HMNL06, JHW05, KRM09, MMCW09, PVD^{+}07, SP09, WGN^{+}08]. Three-dimensional [RWFK00, SF04]. Threshold [AD05a]. Thresholds [DdG09b, SCS^{+}08]. throughout [CPGC09]. thymidine [CD03]. Thysanoessa [JHGGP06]. Tidal [GBM^{+}04, SDV^{+}06, TNO^{+}08, Bel01, BKM05, CBH08, Cra07, EB01, FNM07, GHS05, GK03, GHB^{+}05, GVT^{+}09, HBH^{+}06, HFC^{+}04, JCC02, LTOS04, MSJMN02, NA03, PF01c, SGPK07, SABA02, STK^{+}04, Tho09, VVSM09]. tidal-stream [HFC^{+}04]. Tidally [BNM^{+}09, LLDS06]. Tide [RLJdB08, DCMS03, GRB02, HU07, JVL00, KHNV05, KSP07, PLK^{+}05, PSH08, PN06, PMK08, SP09, STG^{+}07]. Tide-driven [RLJdB08]. tide-pool [SP09]. tidepool [Dav00]. tides [ARO^{+}04, DKRZ06, SSV^{+}04, dSYL^{+}08]. tight [VMH^{+}01]. Time [GHN^{+}02, MVC^{+}05, MN04, AHHH^{+}07, BEM^{+}02, Boy01, CBH08, HTWG08, LB04b, LTB09, MRNT^{+}07, MCPP06b, MCLM02, OHCP02, PDS05, RC05b, SAS08, SKBC01, SRKK08, SDKZ^{+}09, SON^{+}03, TG03]. time-integrated [SAS08]. Time-series [GHN^{+}02, MVC^{+}05, MN04, HTWG08, SKBC01, SON^{+}03]. timescales [FAL01, MMKN08]. timing [DJO00, De 02, TVT04]. Timothy [Wro01]. tintinnid [JJP04]. tintinnids [DJT06]. Tissue [DBS^{+}04, ACW02, FMWC00, QYEA06, SSL05, TMK^{+}07, TGFPT08]. titres [EMWL06]. TN [NHPHH08]. together [BDK^{+}08]. tolerance [NRB08]. tonsa [Møl07]. too [Kiø07]. tool [RB02, WPV00]. tools [Sos00]. Top [CCS06, Goe02b, MMO06, RH03, SCDR07, WPW^{+}09, DHTH05, HDWDD09, KB05a, LWS01, PSH08, SMO07]. Top-down [Goe02b, MMO06, SCDR07, WPW^{+}09, DHTH05, LWS01, SMO07]. Top-predator [RH03]. Topographic [WLP^{+}07]. Topographical [DMS^{+}03]. topography [BII05, OD08, RHJ02, RHJ05]. Total [GH00, HTS^{+}09, MPAS00, CGL^{+}04, HU07, KHP^{+}06, LHWE01, Lee01, MVW^{+}04, SL01, TKAW03, WKH^{+}06, CGL03]. total-iodine [TKAW03]. Toxic [AKG^{+}05, CK00, TDT01, AD07, AAD08, CD02, DA03, KSEÖ^{+}02, LCS02, MGVM^{+}06, RCD^{+}05, RES^{+}08, SW05a]. Toxicity [DMJ09, RL01a, DGD04, EM08, LFS^{+}02, Leg01, MAMC02, WTC^{+}05, WST06]. toxigenic [OLCT08, TWC^{+}09]. toxin [BKK^{+}06, JTL01, SCP08]. toxins [SH03]. TP [NHPHH08]. Trace [YW02, AVJ^{+}00, BFS02, BFML05, FCJ^{+}06, HHS05, HWYL07, LRF07, MLSS05, SDvEB06, SSM^{+}04, TSSWK^{+}06, TM05, TF04, WG01, WSC^{+}05, Zha00]. trace-elemental [WSC^{+}05]. trace=elemental [WSC^{+}06]. traced [SVPK^{+}09]. tracer [AHHH^{+}07, BHV04, DCMS03, EVW07, HB05, HPAM01, MVW^{+}04, NHN^{+}05, RvBB02, TMS^{+}00, TMA^{+}01, VMS^{+}04, vOMSM06a]. tracers [CDBC00, HPAH^{+}02, HAHS^{+}09, JBvHS08, UGSMN01, YPBL07]. Tracing [AHM^{+}08, SRKK08, SM05b, dGMH^{+}08, KKS^{+}01, NDT07, WBD02]. track [DL00, EPP^{+}08]. Tracking [AGKK08, RPR^{+}08, TMA^{+}01, BFML05, JWRW07, MCL06, WRD01]. tracks [HPT^{+}09]. tract [LWP^{+}07, SSL05]. traits [MB00, RES^{+}08]. transect [CW02, FTC^{+}06, FTY^{+}06, StLHB^{+}07, TN09, WW02]. transfer [BDS^{+}04, CLS05, CW03, DH01, DPFL09, FAM04, FAC05, FAL^{+}07, HM09, LNB04, LFA^{+}09, MBF^{+}03, MRNT^{+}07, NA06, OH08, PHF03, RKM^{+}06, SIH^{+}06, SC00, SLHK02, TF04, ZW06]. transformation [BT00]. transformations [BM03, DC05, KHK^{+}03, SDJT03]. transient [Dal03, ED05, GMRD04, HBL02, HPAH^{+}02]. Transition [Shu01, GJT^{+}08, MVFD04, SMJ^{+}06, TIH^{+}01, ZLP^{+}03]. Transition-matrix [Shu01]. transitions [TPFS09]. Translocation [HLL^{+}06]. Transmission [LLB04, RHJ04, BGV02]. transparency [GDB^{+}09, KLWS08]. transparency-gradient [KLWS08]. transparent [ASP04, BSMR08, MBLP01, PP07]. transplant [GCG^{+}02]. transplanted [DBS^{+}04]. Transport [HR00c, JM08, NA03, dBWF^{+}05, AM08, AG03a, BTB^{+}08, CBA01, CSPW00, DCMS03, FLPL08, FNLG00, GB09, HCC^{+}02, HFC^{+}04, HSB^{+}06, Hes00, HA01, HR00b, HM02, JGSB07, LSM03, LMMW03, LTB09, LTOS04, MP00, MAC^{+}06, MT06, MV03, MGGM06, MTM^{+}01, MF04, MCLM02, PDS07, PPL^{+}02, PF01c, SDM^{+}09, SLB^{+}00, SLD00, SI03, TWS03, WIK08, WW02, Wur07, dGP08]. transported [KCV06]. transporters [Alv07]. transports [HDMB04]. trap [KB01, LBB^{+}04, TB00, TVL^{+}00]. trapped [BMD06]. trapping [TFNT07]. traps [KBE^{+}04, LSPC02]. trawling [PGP01]. treated [WH07]. treatments [RHW00]. tree [HGBB08]. trend [BDK^{+}06b, CCAÆ06, VBH^{+}05]. Trends [AS06, BEM^{+}02, CWC^{+}06, DS07b, JBM^{+}07, VWPP^{+}06]. Trepagnier [EB01]. tributaries [HMT^{+}00]. trichloroethylene [AP00a, MHDH00]. Trichodesmium [ACW^{+}03, Ano04, BFCS^{+}01, BFBHF04, lLCChTO08, DWA^{+}02, HVP^{+}07, HFZ^{+}07, HCD^{+}09, KDRR09, KSWC^{+}03, LDC^{+}01, MC01, MB05, MBH^{+}06a, MBH^{+}06b, NT08, OGR08, PM08, SMC08, TSSWK^{+}06, WJMD07, WSKL06]. tricornutum [CLPB02]. trigger [PP00, Sak00]. triggers [SSS06].

Trimethylamine [Sak00, PvE00]. triple [SAH^{+}05, SAHS06, SAS08]. tripos [HHB05]. triquetra [KHH^{+}03]. Trophic [BMF^{+}06, BMCvE06, CPB^{+}08, CLS05, DK06, SAS^{+}03, TF04, AS00a, AS06, BCSW07, BK01a, BS01, CSSNA02, CNGV09, CCKC05, CDD07a, DGH^{+}07, Dod06, GL07, HPLG09, HAG08, KAM04, LN02, MM05, MBC01, MWW06, OHCP02, OCSB05, PHF03, Sal05, SK05, SSH^{+}07, TJV01, VR01, VZM07, WBD02, ZW06, ZHSJ09, vOMSM06b]. trophodynamics [FCJ^{+}06]. trophogenic [CS06]. tropical [AMW^{+}07, AJS07, Boo00, CKA03, CPL^{+}04, COK^{+}08, DTM^{+}08, FDP^{+}08, FCGC^{+}09, FSM^{+}07, HBI01, HM09, KKH06, KFH^{+}09, MML^{+}08b, NTDV06, RNC03, RP06, RPR^{+}02, SHGM06, SCM08, TSSWK^{+}06, TSDK^{+}09, WNL^{+}08, WTM09, dCCL09]. trossulus [WSBL06]. Tsho [SCY00]. tube [AH01a]. tubeworm [RSS^{+}03]. tuna [SSV^{+}04]. Tuning [vDBL05]. turbid [BKM05, DAC^{+}07, RDPM06, SPG05, TM05]. turbidity [ACG07, ARO^{+}04, CLL^{+}03, NH06b, PGP01]. Turbinaria [TGFPT08, TTG^{+}09]. Turbulence [Fra01, ESI01, FAL^{+}07, FMS04, FNM07, HHB05, JWRW07, LMMW03, MK00, RKWG00, RMK^{+}06, SCB03]. turbulence-enhanced [Fra01]. turbulence-mediated [FNM07]. Turbulent [MCJF09, PSLO07, SFL05, WPV00, CTR^{+}02, HM08, MNSV03, RSC04, RKK09, SI01, SSM^{+}09, VOG^{+}08, WW01, WRD01]. turf [QYEA06]. Turner [Caf01, Dow02]. turnover [BD04, DBS^{+}04, KL00, PG00, PIF08, RP06, SR07]. Twenty [DJG^{+}06]. twilight [BB09, SVB^{+}08]. Two [DHKH05, BHMS06, BMU04, BdG02, Boy01, BLBB07, CSSNA02, CG01, CWG^{+}07, CH05b, Dav00, FVD01, FMG01, FG06, GSG^{+}05, Ham05, HOFD01, JS03a, KGYU06, KMvD^{+}09, Law08, LH05, MBK^{+}09, MH04b, MBCB08, MTM^{+}01, OH00, OCNH00, PED^{+}09, Reu02, RLM^{+}04, SDM^{+}09, SPR02, SBB00, SC08, TLM04, VM06, WBB09, WH07, ZB03a, ZRL06]. two-component [PED^{+}09]. two-dimensional [FG06, LH05, ZB03a]. two-mile [Boy01]. type [DK06, RGAS04]. types [RLM^{+}08, SHS03b]. Typhoons [JCK^{+}08]. typicus [Goe08, Møl07].

U [MWE^{+}03]. U.K. [GHTJ00]. U.S [Cra07]. U.S. [AS06, CWWS00, FS09a, PWH^{+}04, RH01, YSO02]. U.S.A. [CL02, CCKC05, HBM08, LAN08, OCNH00, VWPP^{+}06, WHC^{+}06]. ubiquitous [TCB^{+}00]. Uca [KA06]. ultra [KKH06, UVY^{+}01]. ultra-high [KKH06]. ultra-oligotrophic [UVY^{+}01]. ultraoligotrophic [CMC^{+}09, MSGK06]. Ultrasonic [GBH^{+}02]. Ultraviolet [BN01, BJG^{+}00, KLWS08, LFPH07, OV03, ROP^{+}02, ARB^{+}04, ARK^{+}00, BSMR08, BM03, DSW05, FGM^{+}03, HBS04, HBS05, KCE^{+}00, KBD^{+}00, LVC^{+}00, LCS^{+}03, LW01b, LLB04, LNB02, MPD^{+}06, MGW^{+}05, NRB08, Shi04, SWN08, SNPK^{+}09, TXH05, WdMD^{+}00, WGD^{+}02, WDL07, ZSS^{+}08, vdPAJ^{+}06]. ultraviolet-B [FGM^{+}03, WdMD^{+}00]. ultraviolet-visible [MPD^{+}06]. Ulva [CWC^{+}07, MB01, MCL06]. unaffected [DKt08]. unarmored [MA00]. uncertainties [AD02]. Uncertainty [BNC04, MWE^{+}04]. unchelated [MKS08]. unconnected [MGRE^{+}07]. Uncoupled [BD07]. underestimation [YPBL07]. undersaturation [CPL^{+}04]. understand [GL07]. understanding [BSS09, RL01b, Sch06b]. underwater [CLF^{+}08, FAM^{+}06, HM08, JN08, LSK07, NAH^{+}08, RWFK00, SGLT08]. undulatus [EKdF07]. Unexpected [YPBL07]. unfished [PGP01]. unicellular [ARV^{+}07, FMG^{+}08, KCP08, LB09, NFS^{+}07]. unidirectional [MKM06]. uniform [CLL^{+}03]. Union [Caf01, Dow02]. Unique [GS04]. United [BHV04, DNS06, HCC^{+}02, PDS07]. universa [LEML09]. Universal [HFDO05, PSS00]. University [Boy01, Bri01, Fou02, Jun01, Mil02]. unmanipulated [RYS^{+}02]. unmask [PHF03]. unproductive [BLJ09, BBJ05, KJJ07]. unsaturated [WBvE02]. unsolved [RL01b]. unstable [BCSW07]. unsteadiness [HDFM08]. unsteady [LASR09]. unsustainable [THT^{+}09]. Unsynchronised [CTWFP06]. Untersee [WSNH06]. Unusual [SO07, UVY^{+}01]. upgrading [BMCvE06]. upland [CGWK^{+}09]. Upogebia [DD09a]. upon [MSHE09]. Upper [KNH00, ASP04, BNC04, CS06, CPL^{+}04, FCGC^{+}09, LHWE01, LBC06, MWE^{+}03, MWE^{+}04, SRKK08, ZXN^{+}08, AIK04, XLF^{+}03]. upper-ocean [SRKK08]. upside [CBD^{+}06]. upside-down [CBD^{+}06]. Upstream [KDO09, lLCChTO08, LFZ^{+}07]. Uptake [BFDC01, BFC05, CBvEK03, GMRVFP02, SBL06, Ano08a, ATJH08, AB07, BFD^{+}04, BW99, BW00, BSSL03, BDK^{+}08, BARS01, CEBMF05, lLCChTO08, CT02, DHWW02, FAM04, FGM^{+}03, FSBB05, GGRJ03, GFPG09, GKS^{+}02, GD01a, GMAFP03, GP06b, HBL02, HT03, HTS^{+}09, HTRM^{+}07, HWL^{+}00, KD02, Kim05, KTMM00, LJCM08, LNB04, LKM03, LBC^{+}04, LLC^{+}09, MBL^{+}01, MPB^{+}08, MGL02, MVW^{+}04, ML09, NKJM06, PWX04, PSJ00, QRF06, RCD^{+}05, SMCW02, SC00, SWSJ02, SH03, SY07, SC08, SH05b, SH07, TRT04, TvAVH06, TMC06, TPG^{+}08, VMSB08]. Upwelling [RRSC01, AOR07, ÁSGMP01, ÁS07, ABFS09, BZS07, BRS01, CFG07, DKA^{+}03, DGM^{+}05, KMC^{+}00, LAG^{+}03, MS03, MZV03, NSH^{+}09, NCÁSP06, PN06, PPL^{+}02, SRM05, SLB^{+}00, SGG^{+}08, TvG02, TAB^{+}00, VMC^{+}07, VMS^{+}03, WBB09, HHW^{+}02]. upwelling-downwelling [SGG^{+}08]. urban [CB04, LBEP06, MED^{+}08, TYK^{+}08]. urbanization [QLD^{+}02, YYTH07]. urbanization-induced [YYTH07]. urbanized [DC05, TWS03, YS06]. urchin [GWH06, LLB04, YM03]. urchins [NA05]. urea [EM08, LV01]. urea-grown [EM08]. Urease [DA03]. Urotricha [WSL^{+}02]. Urrutia [CR08]. US$100 [Elm02]. US$139 [Pom01]. US$200 [Ris02]. US$24.95 [Boy01]. US$39.50 [Jun01]. US$52.95 [Bri01]. US$56 [Caf01]. US$60 [Mil02]. US$62.95 [SG02]. US$68 [Dow02]. US$76 [Gou01]. US$80 [Fou02]. US$90 [Kne02]. USA [CSM^{+}08, MA08]. Use [BW01, BK01c, HBI01, MBC01, MC04, PF01b, RBW09, SRG00, ATJH08, AD01, AGP07, BMF^{+}06, BSG^{+}08, CH01, CC02, FQH^{+}01, GGP01, HHB^{+}09, HCH00, HW07b, HMKZ02, IR06, KVR03, KHNV05, MCLM02, NP01, PWH^{+}04, PLMPN08, RvBB02, SrL05, SHS03a, SKH^{+}01, UGSMN01, XTSM07]. use/cover [BSG^{+}08]. used [DNS06, EWE02]. usefulness [WRD01]. uses [SAS^{+}03]. Using [KH03, PHF03, SGPM09, AH01a, BNK02, BRB^{+}01, BMSA08, CEA^{+}09, CD02, CT02, DCMS03, DPK^{+}00, ETCP03, EPS00, FDH^{+}01, FKCF00, FSK^{+}04, GWH06, GFK00, HPLG09, HSN^{+}03, HSB^{+}06, HAHS^{+}09, HTWG08, JL08, JN08, KM02, LBD^{+}09, MVW^{+}04, NAH^{+}08, NKJM06, NCPd08, OHUB00, PLM^{+}09, RVP^{+}09, RPR^{+}08, RRSC01, RBMN03, RB01a, RWFK00, SAS08, SHL^{+}05, SDJT03, SA09, STST07, SM05a, TMS^{+}00, VD08, VHL^{+}04, WIK08, WKH^{+}06, WAB^{+}08, WSL^{+}02, WR03, vE02]. Utah [DFBA07, KR08]. Utility [CBA01, BCM01, WSNM00]. Utilization [GFJ^{+}02, MPRC02, Aki00, Alt01, CS06, CCKP02, EPS00, FSSP02, FL08, FSB^{+}08, KCP08, LMPY00, MC01, PPH03, RlBW01, RRHPT^{+}06, RGAS04, TB00]. utilizing [SO07]. UV [BVH02, BVK02, GRV^{+}09, GWK01, HR00a, HBG^{+}09, HLH09, LBC^{+}09, Les01, LA06, MV00, MBC^{+}05, MN04, RK02, WV00]. UV-B [MBC^{+}05]. UV-enhanced [LBC^{+}09]. UV-induced [RK02]. UVB [WWS^{+}00]. UVR [CMC^{+}09, NRB08]. UVR-acclimatized [NRB08].

Val [TMUW02]. Valen [KW04]. Valid [MSC^{+}04]. validate [BW01]. Valigura [Caf01]. Valley [DP01, Pil04, TPM01]. Valleys [GMRD04, Vin00]. Vallisneria [NA06, NA09]. values [LW01a, TB00, TB05]. Vancouver [NK05, WIK08]. Variability [Ceb02, NSH^{+}09, SHGM06, THL^{+}09, TG03, Agr05, ASGA^{+}07, Ano07, ABDS04, AC08, Bat02, BNC04, BFS^{+}04, BMH^{+}08, BLS^{+}06, BBG05a, CLAK08, DD09b, DBTK02, EF05, FLWP09, FCJ^{+}06, GBP03, GKJ04, GST^{+}01, HKF01, HPT^{+}09, HST^{+}02, HH05, JN08, KFR^{+}01, KRZ06, LVC^{+}00, LIE06, LSSA04, LHB04, MRK02, MOK^{+}03, MM03, MM05, MIN^{+}05, MCB08, MWE^{+}04, MCM06, NBW^{+}02, NCC00, OS01b, PBMSG04, PLMPN08, PLM^{+}09, RZMB05, RV04, RYS^{+}02, Sal05, SSH^{+}07, SM05a, SMA00, TLS01, THA01, VC05, VCKS08, WG04, ZSS^{+}08, ZZPP01]. Variable [BLSQ07, CT04, WH07, BVH02, SHS03b, VWPP^{+}06]. variables [BBLB05b, BBLB05a, BDK^{+}06b, RM06]. variance [LH06]. Variation [FW07, GGP01, NKS^{+}02, SIH^{+}06, VR01, ABJ^{+}06, ARO^{+}04, BFML05, CG09, CC01, Car02, Dav00, DHKH05, FHBK06, Fyf03, GSB^{+}09, HTRM^{+}07, HB06, KN07, LLS08, LW01a, Leg01, LBB^{+}04, LBM^{+}04, MM05, OD08, PC02, PKKB05, PDR^{+}09, PGL06, PVD01, PCDB09, RLP^{+}06, Rit09, RPR^{+}02, RHM04, RNA06, SAHS06, SRL^{+}09, SRF^{+}01, SP09, SPW^{+}05, SJ00, SSL05, TWR04, UPC^{+}05, VRS09]. Variations [BS04, MG04, SBW07, SMO07, BTS^{+}06, BBG^{+}05b, CGD08, CBB^{+}02, DLZ^{+}09, DHWW02, DRM^{+}09, FSSH03, FNLG00, GCS06, HSB^{+}04, LIHS03, MFK^{+}09, NFFK00, PCA^{+}00, SAH^{+}05, SAS08, SJ08, SRT^{+}03, Sch01, SP02, TRT04, TCCG05, YYTH07]. variegata [JM02]. varies [CNGV09, CG08]. variopedatus [TS01]. various [HSN^{+}03, ZMK^{+}08]. varved [GRdL^{+}08, GRB09, JM08]. varying [ATJH08, HAA09, LG04, RS09, TLrR07, VM05]. vegetated [LMZB08]. vegetation [BP03, LN06, OS01a]. vehicle [HM08, JN08, NAH^{+}08, SGLT08]. vehicles [CLF^{+}08]. veligers [BMF^{+}06]. velocimetry [CEA^{+}09]. velocities [BDS^{+}04, CW03, DMS^{+}03, GWL^{+}07]. velocity [CT02, GDL^{+}00, LN06, PG00, PIF08, PIKB08, TR03]. Venice [BMU04]. vent [Dix00, RSS^{+}03, SWD^{+}05]. ventilated [LPD^{+}04]. ventilation [MGGM06]. vents [AM08, CMCB^{+}09, Met04, Jun01]. versus [AFH07, BNC04, BS01, CWPL05, CD02, HH03, HWYL07, KHKS05, KMC^{+}00, KCP^{+}04, MG04, MBCB08, MWE^{+}04, SC00, SK05, TSG^{+}01, TSA09, TCB^{+}00, VMJ08, VM05]. Vertical [BKAE01, BLHC08, GSL^{+}07, HTAS06, KLWS08, TRT04, Agr05, Ann05, BF08a, BGC04a, BKB02, BJW^{+}02, BRM^{+}07, CMJ07, CLAK08, CMC^{+}09, CTWFP06, Cra07, DJO00, De 02, DKKA09, FLT02a, FGM^{+}03, FNLG00, GZL02, GNU00, HKF01, HFC^{+}04, HST^{+}02, HTWG08, HLH09, IJA^{+}04, JB00, KKT^{+}07, KSM06, Kir03, KL01, KJH03, LLS08, MWH01, RKWG00, Sal05, SJ08, SVC^{+}07, STG^{+}07, SBB00, TPFS09, TKK02, VCCS05, VRC07, WPV00]. vertical-mode [VRC07]. vesicomyid [TBHS01]. via [BFBHF04, CDBC00, GSA^{+}05, HTS^{+}03, HDP09, TN09]. Viability [BDvO^{+}03, AS02]. Vibrio [GKS^{+}02]. vicinity [ZVRL03]. Victoria [MRK02, Har01]. views [HM06b]. Vigo [ÁSGMP01]. vii [Dow02]. viii [Caf01]. Villefranche [TRT03, TRT04]. Vineyard [TT05]. violacein [DMJ09]. violacein-producing [DMJ09]. Viral [CDD07b, MNM^{+}09, PRKHW04, BVN^{+}00, BVG^{+}07, EKD^{+}07, STHG07, SMA00, WAGH09, WSHW05, YCM^{+}01]. Virginia [HB06]. virginica [KN07, TG09]. viridis [PWX04]. Virus [GM04b, HMSI06, HOFD01, MD04, EMWL06, EW08, WBH03]. virus-induced [WBH03]. Virus-like [HOFD01]. Viruses [BVWB07, LCS02, CW02, MGF03, Thi00]. visible [DRM^{+}09, MPD^{+}06]. visual [Aks07, DJO00, De 02, SA04]. visualization [TPFS09]. Vitamin [BSR^{+}07, TS08]. vitro [SLD00]. vivo [OHUB00, WSNM00]. vocalization [BF08a]. vocans [KA06]. volatile [HMNL06, HMT^{+}08, MMW00]. Volcanic [SKTD04, HLK05]. Volcano [dBSN^{+}06]. Volume [Ano01, Agr05, MDL00, MF01, MF02, TSML07, WS06, ZLL^{+}02]. Vostok [CRBF^{+}06]. vs [SVB^{+}08].

W [Caf01, Gou01, Ris02, SMS09]. W. [Leh00a]. Wadden [dBWF^{+}05]. wake [STK^{+}04]. wall [GvRE^{+}02]. Waquoit [YTVR07]. Ward [SRG00]. warfare [PSH^{+}07]. warm [HGBB08, HZG07]. warm-water [HGBB08]. warmer [JLB^{+}06, JKLR08]. warming [BST^{+}05a, DJG^{+}06, HAG08, JBM^{+}07, JSJ03, LTHW04, MGH08, MSVS^{+}07, MSPAE06, SSvNH01, SJH03, SCV04, VH09, VBH^{+}05, WSEL09, WCGBG07]. warms [TYK^{+}08]. Washington [Dow02, WIK08, ABDS04, CC02, EPD09, HSS06, KF01, PSEL08, RHN^{+}02, RHW00, THH02, TWC^{+}09]. Water [HLGI08, LCLP02, MRNT^{+}07, SH06, SHS01, Agr05, Aki00, AJS07, AELS04, Arn00, BS04, BPOT04, BM01, BEM^{+}02, BdBS^{+}08, BMD^{+}04, BKK^{+}06, BNM^{+}09, BDNF01, BM03, BLG^{+}07, BZH08, Caf01, CBD^{+}06, CGL^{+}04, CK00, CWG^{+}07, CT02, DDL^{+}09, DMS^{+}03, DKRZ06, DPFL09, DKKA09, EDF^{+}03, FS00, FS09a, FLWP09, Fol05, FB05, GSG^{+}05, GJRM07, GP03, Gre09, HGBB08, HLD^{+}03, HDP09, HST^{+}02, HPAH^{+}02, HKD01, HFDO05, JDS08, JFH^{+}05, JHW05, JKLR08, KJMJ03, KAJ08, Kir03, KAW05, KL01, KRM09, Lee01, LH05, LP07, LMMW03, MCJ^{+}06, MGGM06, MWW06, MS05, NIE^{+}00, ORH01, PCC^{+}07, PGP01, Par01, PFH^{+}03, PSS00, PH03, PFPH04, RCAK05, RG02, RLJdB08, RDPM06, RL07, SC00, SLHK02, SHL^{+}05, SKR01, SDvEB06, SSAZ01, SPG05, Sos00, ST03, SSMM00, SHS03b, TPFS09, TLvAH06, TR03, TRF^{+}06, TG03]. water [TMA03, VBH^{+}05, Wey04, WDL07, WLRB03, WBR03b, WRG^{+}05, vODL^{+}09, CHB08]. water-column [Aki00, JDS08]. water-leaving [DKRZ06, RDPM06]. waterbirds [GSA^{+}05]. waterfowl [RL07]. waters [Ack03, ACZ00, ARK^{+}00, BMFS^{+}03, BAG^{+}09, BOC01, Boo00, BRP^{+}06, CG09, CDD^{+}03, CLG08, CAP09, DCMS03, DHWW02, DJH05, DRM^{+}09, EDF^{+}03, FLT02a, FSBB05, GLP^{+}08, GZL02, GSO03, GZY^{+}08, GDL^{+}00, HOFD01, HPT^{+}09, KFR^{+}01, LFS^{+}04, MMLH06, MV00, MA08, MGC^{+}07, MLSS05, NFFK00, NTDV06, OAA^{+}06, OOB09, OCSB05, PLK^{+}05, PPH03, RB09, RC07, RLC^{+}01, RWK08, RDPM06, ST04, SSH^{+}09, STC03, SWS^{+}06, SAPA^{+}02, THWS00, TWS01, TWS03, TvAVH06, TDJ^{+}06, TBH07, TP03, TF00, TMH^{+}06, UGSMN01, VMS^{+}03, VM05, WBH03, WCB^{+}06, WR03, WPV00, YT03, YSO02, Zha00, ZWHP04, ZWC^{+}04, Sta01]. watershed [AD01, EBS07, HBM08, KVR03, KHB^{+}00, RMG^{+}04, TCP^{+}03, XLF^{+}03]. watershed-derived [TCP^{+}03]. watersheds [KCV06, Sch09]. Wave [FPPH06, FB05, VTO^{+}05, AIS00, AI03, ABFS09, CMD^{+}09, DMS^{+}03, DRSL07, EMK08, FAM04, GDK03, HMS^{+}07, HH03, HD03b, HLM08, JLB^{+}06, OD08, PFPH04, RKK09, SPZ03, VCCS05, WLP^{+}07]. wave-dominated [DRSL07, FAM04, HMS^{+}07]. wave-driven [HLM08, PFPH04, RKK09]. wave-exposed [ABFS09]. Wave-induced [FB05, VTO^{+}05, DMS^{+}03]. waveguide [BK01b]. waveguides [BK01c]. wavelength [LBD^{+}09]. waves [AI01, BIIA03, BII05, GGIA06, GGIAY08, HD03b, HISW00, HLP08, LPB06, LSF^{+}08, MCJF09, MF04, PH03, RKM^{+}06, SI08, TG03, UL05, VRC07, dlFSIN08]. Weak [LLT05, BLHC08]. weakly [dlFSIN08]. weather [GV09, HGFS^{+}05]. web [AGKK08, BMBA07, CSWD07, CSK^{+}00, CPCK00, CGWK^{+}09, DPC^{+}08, FH09, HB05, Ham05, HSS06, HSB00, HTS^{+}03, KAM04, Law08, MGVM^{+}06, OAA^{+}06, PCC^{+}07, PSHM04, SAPA^{+}02, STST07, SHS03a, TRT03, TJV01, THJ^{+}09, VJ07, VR01, WWMB05, WLTF03, WLS00, ZHSJ09, vODL^{+}09]. webs [BFD^{+}04, BL07, BSG^{+}08, CLS05, DGH^{+}07, HMSI06, JSJ03, KTG^{+}06, LFFC06, NH06a, OHCP02, QMS^{+}07, SSvNH01, SPM04, VMC^{+}07]. Weddell [PTK^{+}07]. week [AGKK08]. weight [BRL03, DdG09a, EB01, HYH04, HSR^{+}08, HB03, MSBHG01, RA06, ZWC^{+}04]. weighting [OZM^{+}01]. weissflogii [DP04, LLA05, TSKA05, vdPVB07]. well [KM02, TCP^{+}03]. well-flushed [TCP^{+}03]. well-mixed [KM02]. West [LJ09, CDD^{+}03, FSS^{+}00, PDS07, SLS^{+}02, DGCMK00, DCC^{+}01, LDC^{+}01, RBJ^{+}06]. western [ACW^{+}03, ALH01, AKG^{+}05, BP03, BLHC08, BST^{+}08, CLL^{+}09, DB07, FSM^{+}07, HHB^{+}09, HVP^{+}07, KED^{+}07, KFH^{+}09, MHK^{+}07, MH04b, RvAV^{+}06, RSH08, RG02, RWK08, SGB04, TMYH02, TCCG05, WJMD07, DWA^{+}02, SSA^{+}04]. wetland [BP03, Ham01, IR06, LAN08, MYM^{+}05, RW02, SMD^{+}04, SJW03, THA01, VW08, ZWHP04, ZHL06, Bri01]. wetland-derived [VW08]. wetland/estuarine [MYM^{+}05]. wetlands [ARK^{+}00, HL02, LFS^{+}02, NHS07, OS01a, PWH^{+}04, THA01, WS04, Bel01]. WH7803 [CBvEK03]. whale [BF08a, Sch00]. Where [WGN^{+}08]. while [TYK^{+}08]. whitefish [GP06a]. Whole [FVD01, JMBJ01, BRB^{+}01, BLM^{+}06, CCKP02, GHB^{+}05, KCP^{+}04, RlBW02]. whole-basin [BRB^{+}01]. whole-community [BLM^{+}06]. whole-ecosystem [GHB^{+}05]. Whole-lake [JMBJ01, CCKP02, KCP^{+}04]. Whole-stream [FVD01]. Wide [SF05, RHK^{+}09]. Wide-ranging [SF05]. widely [STP^{+}07]. widespread [LCS02]. Wiley [Pom01]. Wiley-Liss [Pom01]. Willapa [RHN^{+}02]. Wind [GGIAY08, RHN^{+}02, UM05, AI03, BSS09, BI07, CW03, DR00, GZL02, HLP08, NSH^{+}09, REL06, RNR08, WPV00]. wind-driven [GZL02, REL06]. wind-forced [NSH^{+}09, WPV00]. Wind-induced [RHN^{+}02, UM05]. wind-inflow [BI07]. Wind-shear-generated [GGIAY08]. wind/internal [AI03]. Winter [KAJ08, KBE^{+}04, AMS^{+}02, CJB00, FLT02a, FLT02b, Har02, KBG^{+}08, MAC^{+}03, MJ03, MAS^{+}02, MFG^{+}09, RL01a, SJH03, WHC^{+}06]. wintry [HGFS^{+}05]. within [All00, AHR^{+}01, BMD06, FS00, FAL01, FCGC^{+}09, FRO01, GRR^{+}07, GGP01, HBH00, HHM^{+}09, Hol00, LSPC02, LVC^{+}00, LAN08, LZ09, New01, PH07, PCDB09, PHK02, PP07, PIF08, RKM^{+}06, REL06, SMP05, TCB^{+}00, WNAW09]. within-lake [LVC^{+}00]. without [LSF^{+}08, MSVAC04]. wolfi [HMD^{+}07]. wollei [CT06]. Wood [EW07]. Woodin [Mil02]. world [GD00a, GSJ^{+}09, JLB^{+}06, TYK^{+}08, SF05]. world-ocean [GSJ^{+}09]. worm [vNHM^{+}09]. worms [KKH04]. WT [UGSMN01]. Wyoming [IKT03, HT03].

X [Bel01, HSN^{+}03]. X-ray [HSN^{+}03]. xi [Fou02]. xii [Wro01]. xiv [Bri01]. xix [Sta01]. xx [Jun01]. xxi [Mil02].

Yaquina [DD09a]. year [AH03, BJG^{+}00, ES00, GSG^{+}05, SKBC01, SRKK08, UMASG07]. year-round [UMASG07]. years [BLSQ07, DJG^{+}06, GRdL^{+}08, Ham05, LD01, MMPA04, PLG^{+}02, RCVK09, Tun00]. Yellowstone [IKT03]. yield [HMKZ02, MMS^{+}02, Mor03, VSSTS00]. Yoldia [ST03]. York [DPK^{+}00, GRB02, Kra05, NA03, RBC00]. Young [SMS09, AH03, LED06, RQB^{+}00]. young-of-the-year [AH03]. Yucatan [SPR02].

Zealand [NDP^{+}07, WLP^{+}07]. Zebra [BDNF01, ALH01, BFC05, BFC07, BMF^{+}06, BLHC08, NPE08, RSWH04, SWH^{+}05, WBvE02, WvE04, WLTF03]. Zedler [Bel01]. zetterstedtii [SJRB09]. Zinc [BTdB03, JGSB07, PWS09, SG08, SXLM06, XTSM07, CLP^{+}03, CS05, DA05, EM08, EH00, FSS^{+}00, LCPS05, RRD04, SDC00, SH04, SH05b, TSG^{+}01, TLM04]. Zinc-bicarbonate [BTdB03]. Zinc-cobalt [SG08]. Zn [MJW03, SXLM06, SH00, WDX01]. Zn-P [SXLM06]. zonation [HH03, JB00, dBSN^{+}06]. Zone [GRG02, KCCE06, AO09, BKB02, BK03b, BW99, BB09, BJL01, CGH^{+}04, CS06, CWC^{+}06, CIWH09, DMS^{+}03, DH01, FS00, FCGC^{+}09, FVD01, Gay00, HB05, HLW^{+}07, HD03b, HLP08, JHW^{+}07, JB00, LKAB04, LPB06, MVFD04, NEE08, RHK^{+}09, SA00, SMJ^{+}06, SVB^{+}08, TIH^{+}01, TCB^{+}00, TM05, WGC00]. zones [HHT01, TAB^{+}00, ZLP^{+}03]. zooplanktivores [SA04]. Zooplanktivory [HG03a, HG03b]. Zooplankton [ET08, GWK01, HBH00, HLH09, JKS04, SK04, SMJ^{+}06, ANSS04, ARB^{+}04, BB04b, BMNSK00, CPHM04, CD02, CTWFP06, DTL^{+}05, DET04, DJO00, DHTH05, dSFFTD07, GJS01, GHJ^{+}05, HGB00, HPLG09, HNB^{+}07, HS04, HABF03, HBG^{+}09, JJT07, KKKT07, KJMJ03, KHKS05, KLWS08, Kiø00, LLS08, LBC^{+}09, MNSV03, MM03, MM05, MBG^{+}05, MKG04, MFF07, MCC02, MSJMN02, OBB^{+}04, PHF03, PW03, PIF08, RHK^{+}09, RFM07, RH03, RYS^{+}02, SK05, STST07, SVB^{+}08, SLC^{+}09, SCV04, STK^{+}04, TA03, TLS01, TCD^{+}03, VC05, VK06, VKV07, WGN^{+}08, WHHS06, WGD^{+}02, WDL07, WST06, WBR03b, WLTF03, YYB^{+}05, ZHSJ09]. zooplankton-mediated [ZHSJ09]. zooxanthellae [GFPG09, QYEA06, SSL05, TGFPT08]. zooxanthellate [GRVF^{+}00]. Zostera [DDZ03, FG06, FHBK06, GBP03, PKKB05, Reu02]. Zygorhizidium [Hol00].

References

Avery:2008:SRD

[AAD08] David E. Avery, KristaJoy K. Altland, and Hans G. Dam. Sexrelated differential mortality of a marine copepod exposed to a toxic dinoflagellate. Limnology and Oceanography, 53(6):2627– 2635, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Alvarez:2007:BBE

[ÁÁS07] Marta Álvarez and XosÉ Antín Álvarez-Salgado. Biogeochemical budgets in the eastern boundary current system of the North Atlantic: Evidence of net heterotrophy and nitrogen fixation. Limnology and Oceanography, 52(4):1328–1335, July 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Arp:2007:DSN

[AB07] Christopher D. Arp and Michelle A. Baker. Discontinuities in stream nutrient uptake below lakes in mountain drainage networks. Limnology and Oceanography, 52(5):1978–1990, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Arhonditsis:2004:ECV

[ABDS04] George B. Arhonditsis, Michael T. Brett, Curtis L. DeGasperi, and Daniel E. Schindler. Effects of climatic variability on the thermal properties of Lake Washington. Limnology and Oceanography, 49(1):256–270, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Aquilino:2009:LSN

[ABFS09] Kristin M. Aquilino, Matthew E. S. Bracken, Melinda N. Faubel, and John J. Stachowicz. Local-scale nutrient regeneration facilitates seaweed growth on wave-exposed rocky shores in an upwelling system. Limnology and Oceanography, 54(1):309–317, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Acharya:2006:CGL

[ABJ^{+}06] Kumud Acharya, Paul A. Bukaveckas, Jeffrey D. Jack, Marcia Kyle, and James J. Elser. Consumer growth linked to diet and RNA–P stoichiometry: Response of Bosmina to variation in riverine food resources. Limnology and Oceanography, 53(2): 1859–1869, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Austin:2008:CTV

[AC08] Jay Austin and Steve Colman. A century of temperature variability in Lake Superior. Limnology and Oceanography, 53(6): 2724–2730, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Amyot:2010:RCE

[ACD10] Marc Amyot, Saad Chidam, and Éric Demers. Response to “Comparative estimate of P fluxes in lakes” by J. M. Sereda and J. J. Hudson. Limnology and Oceanography, 55(1):466–468, January 2010. CODEN LIOCAH. ISSN 0024-3590. See [SH10].

Abril:2007:EMO

[ACG07] Gwenaël Abril, Marc-Vincent Commarieu, and Frédéric Guérin. Enhanced methane oxidation in an estuarine turbidity maximum. Limnology and Oceanography, 52(1):470–475, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Anthony:2007:BEC

[ACHG07] Kenneth R. N. Anthony, Sean R. Connolly, and Ove Hoegh-Guldberg. Bleaching, energetics, and coral mortality risk: Effects of temperature, light, and sediment regime. Limnology and Oceanography, 52(2):716–726, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Ackleson:2003:LSW

[Ack03] Steven G. Ackleson. Light in shallow waters: a brief research review. Limnology and Oceanography, 48(1 (part 2)):323–328, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Anthony:2002:CAE

[ACW02] Kenneth R. N. Anthony, Sean R. Connolly, and Bette L. Willis. Comparative analysis of energy allocation to tissue and skeletal growth in corals. Limnology and Oceanography, 47(5):1417– 1429, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Achilles:2003:BIT

[ACW^{+}03] Katherine M. Achilles, Thomas M. Church, Steven W. Wilhelm, George W. Luther III, and David A. Hutchins. Bioavailability of iron to Trichodesmium colonies in the western subtropical Atlantic Ocean. Limnology and Oceanography, 48(6):2250–2255, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Andrews:2000:POC

[ACZ00] Steven S. Andrews, Sigalit Caron, and Oliver C. Zafiriou. Photochemical oxygen consumption in marine waters: a major sink for colored dissolved organic matter? Limnology and Oceanography, 45(2):267–277, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Agusti:2000:SSP

[AD00a] Susana Agustí and Carlos M. Duarte. Strong seasonality in phytoplankton cell lysis in the NW Mediterranean littoral. Limnology and Oceanography, 45(4):940–947, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2000:SEA

[AD00b] L. D. Anderson and M. L. Delaney. Sequential extraction and analysis of phosphorus in marine sediments: Streamlining of the SEDEX procedure. Limnology and Oceanography, 60(3):509– 515, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Arbuckle:2001:IWL

[AD01] Kelly E. Arbuckle and John A. Downing. The influence of watershed land use on lake N : P in a predominantly agricultural landscape. Limnology and Oceanography, 46(4):970–975, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Agusti:2002:AUA

[AD02] Susana Agusti and Carlos M. Duarte. Addressing uncertainties in the assessment of phytoplankton lysis rates in the sea. Limnology and Oceanography, 47(3):921–924, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Agusti:2005:TGP

[AD05a] Susana Agustí and Carlos M. Duarte. Threshold of gross primary production for planktonic metabolic balance in the Southern Ocean: an experimental test. Limnology and Oceanography, 50(4):1334–1339, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Ahearn:2005:SND

[AD05b] Dylan S. Ahearn and Randy A. Dahlgren. Sediment and nutrient dynamics following a low-head dam removal at Murphy creek, California. Limnology and Oceanography, 50(6):1752– 1762, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Avery:2007:NDR

[AD07] David E. Avery and Hans G. Dam. Newly discovered reproductive phenotypes of a marine copepod reveal the costs and advantages of resistance to a toxic dinoflagellate. Limnology and Oceanography, 52(5):2099–2108, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Agawin:2000:NTCa

[ADA00a] Nona S. R. Agawin, Carlos M. Duarte, and Susana Agustí. Nutrient and temperature control of the contribution of picoplankton to phytoplankton biomass and production. Limnology and Oceanography, 45(3):591–600, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Agawin:2000:NTCb

[ADA00b] Nona S. R. Agawin, Carlos M. Duarte, and Susana Agustí. Nutrient and temperature control of the contribution of picoplankton to phytoplankton biomass and production. Limnology and Oceanography, 45(8):1891, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Attrill:2008:RHH

[AE08] Martin J. Attrill and Martin Edwards. Reply to Haddock, S. H. D., Reconsidering evidence for potential climate-related increases in jellyfish. Limnology and Oceanography, 53(6):2763– 2766, November 2008. CODEN LIOCAH. ISSN 0024-3590. See [Had08].

Arneborg:2004:RIM

[AELS04] Lars Arneborg, Carina P. Erlandsson, Bengt Liljebladh, and Anders Stigebrandt. The rate of inflow and mixing during deepwater renewal in a sill fjord. Limnology and Oceanography, 49 (3):768–777, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Amon:2001:LAB

[AFB01] Rainer M. W. Amon, Hans-Peter Fitznar, and Ronald Benner. Linkages among the bioreactivity, chemical composition, and diagenetic state of marine dissolved organic matter. Limnology and Oceanography, 46(2):287–297, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Andersen:2007:GRV

[AFH07] Tom Andersen, Per J. Færøvig, and Dag O. Hessen. Growth rate versus biomass accumulation: Different roles of food quality and quantity for consumers. Limnology and Oceanography, 52 (5):2128–2134, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Abrusan:2007:BLR

[AFL07] György Abrusán, Patrick Fink, and Winfried Lampert. Biochemical limitation of resting egg production in Daphnia. Limnology and Oceanography, 52(4):1724–1728, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Alve:2003:PTK

[AG03a] Elisabeth Alve and Susan T. Goldstein. Propagule transport as a key method of dispersal in benthic Foraminifera (Protista). Limnology and Oceanography, 48(6):2163–2170, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Anesio:2003:IPD

[AG03b] Alexandre Magno Anesio and Wilhelm Granéli. Increased photoreactivity of DOC by acidification: Implications for the carbon cycle in humic lakes. Limnology and Oceanography, 48(2): 735–744, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Aristegui:2009:MOD

[AGDH09] Javier Arístegui, Josep M. Gasol, Carlos M. Duarte, and Gerhard J. Herndl. Microbial oceanography of the Dark Ocean’s pelagic realm. Limnology and Oceanography, 54(5):1501–1529, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Augspurger:2008:TCF

[AGKK08] Clemens Augspurger, Gerd Gleixner, Christiane Kramer, and Kirsten Küsel. Tracking carbon flow in a 2-week-old and 6-week-old stream biofilm food web. Limnology and Oceanography, 53 (2):642–650, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Adam:2003:BSC

[AGL03] Christelle Adam and Jacqueline Garnier-Laplace. Bioaccumulation of silver-110m, cobalt-60, cesium-137, and manganese-54 by the freshwater algae Scenedesmus obliquus and Cyclotella meneghiana and by suspended matter collected during a summer bloom event. Limnology and Oceanography, 48(6):2303–2313, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Armstrong:2004:OMS

[AGMP04] Evelyn Armstrong, Julie Granger, Elizabeth L. Mann, and Neil M. Price. Outer-membrane siderophore receptors of heterotrophic oceanic bacteria. Limnology and Oceanography, 49 (2):579–587, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Arnon:2007:BPC

[AGP07] Shai Arnon, Kimberly A. Gray, and Aaron I. Packman. Biophysicochemical process coupling controls nitrogen use by benthic biofilms. Limnology and Oceanography, 52(4):1665–1671, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Agrawal:2005:OVS

[Agr05] Yogesh C. Agrawal. The optical volume scattering function: Temporal and vertical variability in the water column off the New Jersey coast. Limnology and Oceanography, 50(6):1787– 1794, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Ackerman:2001:MLB

[AH01a] Josef Daniel Ackerman and Trent M. Hoover. Measurement of local bed shear stress in streams using a Preston-static tube. Limnology and Oceanography, 47(4):2080–2087, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Attayde:2001:RIF

[AH01b] José Luiz Attayde and Lars-Anders Hansson. The relative importance of fish predation and excretion effects on planktonic communities. Limnology and Oceanography, 46(5):1001–1012, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Arrieta:2002:CBG

[AH02] Jesùs M. Arrieta and Gerhard J. Herndl. Changes in bacterial β-glucosidase diversity during a coastal phytoplankton bloom. Limnology and Oceanography, 47(2):594–599, March 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Axenrot:2003:PHR

[AH03] Thomas Axenrot and Sture Hansson. Predicting herring recruitment from young-of-the-year densities, spawning stock biomass, and climate. Limnology and Oceanography, 48(4):1716–1720, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Arnold:2005:DIL

[AHF^{+}05] William S. Arnold, Gary L. Hitchcock, Marc E. Frischer, Rik Wanninkhof, and Y. Peter Sheng. Dispersal of an introduced larval cohort in a coastal lagoon. Limnology and Oceanography, 50(4):587–597, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Aeschbach-Hertig:2007:TSE

[AHHH^{+}07] W. Aeschbach-Hertig, C. P. Holzner, M. Hofer, M. Simona, A. Barbieri, and R. Kipfer. A time series of environmental tracer data from deep meromictic Lake Lugano, Switzerland. Limnology and Oceanography, 52(1):257–273, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Alling:2008:TTO

[AHM^{+}08] Vanja Alling, Christoph Humborg, Carl-Magnus Mörth, Lars Rahm, and Falk Pollehne. Tracing terrestrial organic matter by δ ^{34}S and δ ^{13}C signatures in a subarctic estuary. Limnology and Oceanography, 53(6):2594–2602, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Aufdenkampe:2001:SFD

[AHR^{+}01] Anthony K. Aufdenkampe, John I. Hedges, Jeffery E. Richey, Alex V. Krusche, and Carlos A. Llerena. Sorptive fractionation of dissolved organic nitrogen and amino acids onto fine sediments within the Amazon Basin. Limnology and Oceanography, 46(8): 1921–1935, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Antenucci:2001:ELI

[AI01] Jason P. Antenucci and Jörg Imberger. Energetics of long internal gravity waves in large lakes. Limnology and Oceanography, 46(7):1760–1773, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Antenucci:2003:SEW

[AI03] Jason P. Antenucci and Jörg Imberger. The seasonal evolution of wind/internal wave resonance in Lake Kinneret. Limnology and Oceanography, 49(4):2055–2061, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Appt:2004:BSM

[AIK04] Jochen Appt, Jörg Imberger, and Helmut Kobus. Basin-scale motion in stratified Upper Lake Constance. Limnology and Oceanography, 49(4):919–933, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Antenucci:2000:SEB

[AIS00] Jason P. Antenucci, Jörg Imberger, and Angelo Saggio. Seasonal evolution of the basin-scale internal wave field in a large stratified lake. Limnology and Oceanography, 45(7):1621–1638, November 2000. CODEN LIOCAH. ISSN 0024-3590.

An:2001:ECN

[AJ01] Soonmo An and Samantha B. Joye. Enhancement of coupled nitrification-denitrification by benthic photosynthesis in shallow estuarine sediments. Limnology and Oceanography, 46(1):62–74, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Alkhatib:2007:BDR

[AJS07] M. Alkhatib, T. C. Jennerjahn, and Joko Samiaji. Biogeochemistry of the Dumai River estuary, Sumatra, Indonesia, a tropical black-water river. Limnology and Oceanography, 52(6):2410– 2417, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Acuna:2000:FRA

[AK00] José Luis Acuña and Markus Kiefer. Functional response of the appendicularian Oikopleura dioica. Limnology and Oceanography, 45(3):608–618, May 2000. CODEN LIOCAH. ISSN 0024-

Acharya:2004:BSD

[AKE04] Kumud Acharya, Marcia Kyle, and James J. Elser. Biological stoichiometry of Daphnia growth: an ecophysiological test of the growth rate hypothesis. Limnology and Oceanography, 49 (3):656–665, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2005:TAB

[AKG^{+}05] D. M. Anderson, B. A. Keafer, W. R. Geyer, R. P. Signell, and T. C. Loder. Toxic alexandrium blooms in the western Gulf of Maine: The plume advection hypothesis revisited. Limnology and Oceanography, 55(3):328–345, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Averyt:2004:EPM

[AKH04] Kristen B. Averyt, Jonathan P. Kim, and Keith A. Hunter. Effect of pH on measurement of strong copper binding ligands in lakes. Limnology and Oceanography, 49(1):20–27, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Akihiro:2000:EWC

[Aki00] Shiomoto Akihiro. Efficiency of water-column light utilization in the subarctic northwestern Pacific. Limnology and Oceanography, 45(4):982–987, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Ask:2009:TOM

[AKP^{+}09] Jenny Ask, Jan Karlsson, Lennart Persson, Per Ask, Pär Byström, and Mats Jansson. Terrestrial organic matter and light penetration: Effects on bacterial and primary production in lakes. Limnology and Oceanography, 54(6):2034–2040, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Aksnes:2007:EVC

[Aks07] Dag L. Aksnes. Evidence for visual constraints in large marine fish stocks. Limnology and Oceanography, 52(1):198–203, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Ackerman:2001:BPC

[ALH01] Josef Daniel Ackerman, Mark R. Loewen, and Paul F. Hamblin. Benthic–pelagic coupling over a zebra mussel reef in western Lake Erie. Limnology and Oceanography, 46(4):892–904, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:2000:IDO

[All00] Alice L. Alldredge. Interstitial dissolved organic carbon (DOC) concentrations within sinking marine aggregates and their potential contribution to carbon flux. Limnology and Oceanography, 45(6):1245–1253, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Annett:2008:ECF

[ALRM08] Amber L. Annett, Suzanne Lapi, Thomas J. Ruth, and Maria T. Maldonado. The effects of Cu and Fe availability on the growth and Cu : C ratios of marine diatoms. Limnology and Oceanography, 53(6):2451–2461, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Altabet:2001:NIE

[Alt01] Mark A. Altabet. Nitrogen isotopic evidence for micronutrient control of fractional NO_{3}^{−} utilization in the equatorial Pacific. Limnology and Oceanography, 46(2):368–380, March 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Alverson:2007:SPS

[Alv07] Andrew J. Alverson. Strong purifying selection in the silicon transporters of marine and freshwater diatoms. Limnology and Oceanography, 52(4):1420–1429, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Adams:2008:SGL

[AM08] Diane K. Adams and Lauren S. Mullineaux. Supply of gastropod larvae to hydrothermal vents reflects transport from local larval sources. Limnology and Oceanography, 53(5):1945–1955, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Anning:2000:PMD

[AMP^{+}00] Tracy Anning, Hugh L. MacIntyre, Sandra M. Pratt, Pippa J. Sammes, Stuart Gibb, and Richard J. Geider. Photoacclimation in the marine diatom Skeletonema costatum. Limnology and Oceanography, 45(8):1807–1817, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:2002:FEB

[AMS^{+}02] A. Atkinson, B. Meyer, D. Stübing, W. Hagen, K. Schmidt, and U. V. Bathmann. Feeding and energy budgets of Antarctic krill Euphausia superba at the onset of winter — II. Juveniles and adults. Limnology and Oceanography, 47(4):953–966, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Argerich:2008:CEL

[AMS^{+}08] Alba Argerich, Eugènia Martí, Francesc Sabater, Miquel Ribot, Daniel von Schiller, and Joan L. Riera. Combined effects of leaf litter inputs and a flood on nutrient retention in a Mediterranean mountain stream during fall. Limnology and Oceanography, 53 (2):631–641, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Aguilera:2007:APD

[AMW^{+}07] Ximena Aguilera, Joachim Mergeay, An Wollebrants, Steven Declerck, and Luc De Meester. Asexuality and polyploidy in Daphnia from the tropical Andes. Limnology and Oceanography, 52 (5):2079–2088, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Annis:2005:TEV

[Ann05] Eric R. Annis. Temperature effects on the vertical distribution of lobster postlarvae (Homarus americanus). Limnology and Oceanography, 50(6):1972–1982, November 2005. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2001:ALO

[Ano01] Anonymous. Authors in Limnology and Oceanography volume 46, 2001. Limnology and Oceanography, 47(4):2095, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2004:IRD

[Ano04] Anonymous. Iron requirements for dinitrogen- and ammoniumsupported growth in cultures of Trichodesmium (IMS 101): Comparison with nitrogen fixation rates and iron : carbon ratios of field populations. Limnology and Oceanography, 49(4):1224, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2007:EMV

[Ano07] Anonymous. Erratum: Mesoscale variability in the metabolic balance of the Sargasso Sea. Limnology and Oceanography, 52 (6):2722, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2008:EIB

[Ano08a] Anonymous. Erratum: Interaction between hydrodynamics and seagrass canopy structure: Spatially explicit effects on ammonium uptake rates. Limnology and Oceanography, 53(5):2056, September 2008. CODEN LIOCAH. ISSN 0024-3590. See [MPB^{+}08].

Anonymous:2008:ESR

[Ano08b] Anonymous. Erratum: Seasonal refugia, shoreward thermal amplification, and metapopulation dynamics of the ctenophore Mnemiopsis leidyi in Narragansett Bay, Rhode Island. Limnology and Oceanography, 57(3):873, March 2008. CODEN LI-OCAH. ISSN 0024-3590. See [CSG^{+}06].

Aksnes:2004:OCF

[ANSS04] Dag L. Aksnes, Jens Nejstgaard, Eivind Sædberg, and Tom Sørnes. Optical control of fish and zooplankton populations. Limnology and Oceanography, 49(1):233–238, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Aksnesa:2009:MDS

[AO09] Dag L. Aksnesa and Mark D. Ohman. Multi-decadal shoaling of the euphotic zone in the southern sector of the California Current System. Limnology and Oceanography, 54(4):1272–1281, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Aksnes:2007:OEN

[AOR07] Dag L. Aksnes, Mark D. Ohman, and Pascal Rivière. Optical effect on the nitracline in a coastal upwelling area. Limnology and Oceanography, 52(3):1179–1187, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Adrian:2009:LSC

[AOZ^{+}09] Rita Adrian, Catherine M. O’Reilly, Horacio Zagarese, Stephen B. Baines, Dag O. Hessen, Wendel Keller, David M. Livingstone, Ruben Sommaruga, Dietmar Straile, Ellen Van Donk, Gesa A. Weyhenmeyer, and Monika Winder. Lakes as sentinels of climate change. Limnology and Oceanography, 54(6 (part 2)):2283–2297, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Abrahamsson:2000:ENP

[AP00a] Katarina Abrahamsson and Marianne Pedersén. Evidence of the natural production of trichloroethylene (reply to the comment by Marshall et al.). Limnology and Oceanography, 60(3):520– 522, March 2000. CODEN LIOCAH. ISSN 0024-3590. See [MHDH00].

Anderson:2000:STE

[AP00b] Thomas R. Anderson and David W. Pond. Stoichiometric theory extended to micronutrients: Comparison of the roles of essential fatty acids, carbon, and nitrogen in the nutrition of marine copepods. Limnology and Oceanography, 45(5):1162–1167, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Aronson:2000:HAD

[AP00c] Richard B. Aronson and William F. Precht. Herbivory and algal dynamics on the coral reef at Discovery Bay, Jamaica. Limnology and Oceanography, 45(1):251–255, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Andrefouet:2003:AHD

[APH^{+}03] Serge Andréfouët, Claude Payri, Eric J. Hochberg, Lydie Mao Che, and Marlin J. Atkinson. Airborne hyperspectral detection of microbial mat pigmentation in Rangiroa Atoll (French Polynesia). Limnology and Oceanography, 48(1 (part 2)):426–430, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2006:CAS

[AR06] D. M. Anderson and K. Rengefors. Community assembly and seasonal succession of marine dinoflagellates in a temperate estuary: The importance of life cycle events. Limnology and Oceanography, 51(2):860–873, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Alonso:2004:SAF

[ARB^{+}04] Cristina Alonso, Vanina Rocco, Juan Pablo Barriga, Miguel Ángel Battini, and Horacio Zagarese. Surface avoidance by freshwater zooplankton: Field evidence on the role of ultraviolet radiation. Limnology and Oceanography, 49(1):225–232, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Arts:2000:AUR

[ARK^{+}00] Michael T. Arts, Richard D. Robarts, Fumie Kasai, Marley J. Waiser, Vijay P. Tumber, Amanda J. Plante, Hakumat Rai, and Hendrika J. de Lange. The attenuation of ultraviolet radiation in high dissolved organic carbon waters of wetlands and lakes on the northern Great Plains. Limnology and Oceanography, 45 (2):292–299, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Arnosti:2000:SSP

[Arn00] C. Arnosti. Substrate specificity in polysaccharide hydrolysis: Contrasts between bottom water and sediments. Limnology and Oceanography, 45(5):1112–1119, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Anthony:2004:TVL

[ARO^{+}04] Kenneth R. N. Anthony, Peter V. Ridd, Alan R. Orpin, Piers Larcombe, and Janice Lough. Temporal variation of light availability in coastal benthic habitats: Effects of clouds, turbidity, and tides. Limnology and Oceanography, 49(6):2201–2211, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Ahlgren:2006:DOP

[ART^{+}06] Joakim Ahlgren, Kasper Reitzel, Lars Tranvik, Adolf Gogoll, and Emil Rydin. Degradation of organic phosphorus compounds in anoxic Baltic Sea sediments: a ^{31}P nuclear magnetic resonance study. Limnology and Oceanography, 51(5):2341–2348, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Agawin:2007:CFB

[ARV^{+}07] Nona S. R. Agawin, Sophie Rabouille, Marcel J. W. Veldhuis, Lidewij Servatius, Suzanne Hol, Harriët M. J. van Overzee, and Jef Huisman. Competition and facilitation between unicellular nitrogen-fixing cyanobacteria and non-nitrogen-fixing phytoplankton species. Limnology and Oceanography, 52(5):2233– 2248, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Adams:2000:EDN

[AS00a] Thomas S. Adams and Robert W. Sterner. The effect of dietary nitrogen content on trophic level^{15}N enrichment. Limnology and Oceanography, 45(3):601–607, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Adey:2000:CEI

[AS00b] Walter H. Adey and Allegra M. Small. Community and environmental influences on reef coral calcification. Limnology and Oceanography, 45(7):1667–1671, November 2000. CODEN LI-OCAH. ISSN 0024-3590.

Agusti:2002:CVN

[AS02] Susana Agustí and M. Carmen Sánchez. Cell viability in natural phytoplankton communities quantified by a membrane permeability probe. Limnology and Oceanography, 47(3):818–828, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Alexander:2006:TNE

[AS06] Richard B. Alexander and Richard A. Smith. Trends in the nutrient enrichment of U.S. rivers during the late 20th century and their relation to changes in probable stream trophic conditions. Limnology and Oceanography, 51(1 (part 2)):639–654, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Alvarez-Salgado:2007:CUF

[ÁS07] X. A. Álvarez-Salgado. Contribution of upwelling filaments to offshore carbon export in the subtropical Northeast Atlantic Ocean. Limnology and Oceanography, 52(3):1287–1292, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Arnold:2004:GRA

[ASAC04] Katharine H. Arnold, Rachael S. Shreeve, Angus Atkinson, and Andrew Clarke. Growth rates of Antarctic krill, Euphausia superba: Comparison of the instantaneous growth rate method with nitrogen and phosphorus stoichiometry. Limnology and Oceanography, 49(6):2152–2161, November 2004. CODEN LI-OCAH. ISSN 0024-3590.

Alonso-Saez:2007:LSV

[ASGA^{+}07] Laura Alonso-Sáez, Josep M. Gasol, Javier Arístegui, Juan C. Vilas, Dolors Vaqué, Carlos M. Duarte, and Susana Agustí. Large-scale variability in surface bacterial carbon demand and growth efficiency in the subtropical northeast Atlantic Ocean. Limnology and Oceanography, 52(2):533–546, March 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Alvarez-Salgado:2001:NEP

[ÁSGMP01] X. A. Álvarez-Salgado, J. Gago, B. M. Míguez, and F. F. Pérez. Net ecosystem production of dissolved organic carbon in a coastal upwelling system: the Ría de Vigo, Iberian margin of the North Atlantic. Limnology and Oceanography, 46(1): 135–146, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:2006:NGR

[ASH^{+}06] Angus Atkinson, Rachael S. Shreeve, Andrew G. Hirst, Peter Rothery, Geraint A. Tarling, David W. Pond, Rebecca E. Korb, Eugene J. Murphy, and Jonathon L. Watkins. Natural growth rates in Antarctic krill (Euphausia superba): II. Predictive models based on food, temperature, body length, sex, and maturity stage. Limnology and Oceanography, 51(2):973–987, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Anneville:2002:TMP

[ASI^{+}02] Orlane Anneville, Sami Souissi, Frédéric Ibanez, Vincent Ginot, Jean Claude Druart, and Nadine Angeli. Temporal mapping of phytoplankton assemblages in Lake Geneva: Annual and interannual changes in their patterns of succession. Limnology and Oceanography, 47(5):1355–1366, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Azetsu-Scott:2004:AMP

[ASP04] Kumiko Azetsu-Scott and Uta Passow. Ascending marine particles: Significance of transparent exopolymer particles (TEP) in the upper ocean. Limnology and Oceanography, 49(3):741–748, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Aucour:2003:FMS

[ASS03] Anne-Marie Aucour, Simon M. F. Sheppard, and Régine Savoye. δ ^{13}C of fluvial mollusk shells (Rhône River): a proxy for dissolved inorganic carbon? Limnology and Oceanography, 48(6): 2186–2193, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Arango:2008:AUR

[ATJH08] C. P. Arango, J. L. Tank, L. T. Johnson, and S. K. Hamilton. Assimilatory uptake rather than nitrification and denitrification determines nitrogen removal patterns in streams of varying land use. Limnology and Oceanography, 53(6):2558–2572, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Artolozaga:2002:GRB

[AVA^{+}02] Itxaso Artolozaga, María Valcárcel, Begoña Ayo, Ainhoa Latatu, and Juan Iriberri. Grazing rates of bacterivorous protists inhabiting diverse marine planktonic microenvironments. Limnology and Oceanography, 47(1):142–150, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Alberic:2000:IBT

[AVJ^{+}00] Patrick Albéric, Eric Voillier, Didier Jézéquel, Cécile Grosbois, and Gil Michard. Interactions between trace elements and dissolved organic matter in the stagnant anoxic deep layer of a meromictic lake. Limnology and Oceanography, 45(5):1088– 1096, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:2000:AEAa

[AW00a] Angus Atkinson and Michael J. Whitehouse. Ammonium excretion by Antarctic krill Euphausia superba at South Georgia. Limnology and Oceanography, 45(1):55–63, January 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Atkinson:2000:AEAb

[AW00b] Angus Atkinson and Michael J. Whitehouse. Ammonium excretion by Antarctic krill Euphausia superba at South Georgia. Limnology and Oceanography, 45(4):1012, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Ahner:2005:SSD

[AW05] Beth A. Ahner and Liping Wei. Sources and sinks of dissolved phytochelatin in natural seawater. Limnology and Oceanography, 50(1):13–22, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Attrill:2007:CRI

[AWE07] Martin J. Attrill, Jade Wright, and Martin Edwards. Climaterelated increases in jellyfish frequency suggest a more gelatinous future for the North Sea. Limnology and Oceanography, 52(1): 480–485, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Arrieta:2007:RIR

[AWH07] Jesús M. Arrieta, Markus G. Weinbauer, and G. J. Herndl. Response to: “Interpreting the results of oceanic mesoscale enrichment experiments: Caveats and lessons from limnology and coastal ecology”. Limnology and Oceanography, 52(2):916–918, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Arrieta:2004:RBI

[AWLH04] Jesús M. Arrieta, Markus G. Weinbauer, Carolien Lute, and Gerhard J. Herndl. Response of bacterioplankton to iron fertilization in the Southern Ocean. Limnology and Oceanography, 49(3):799–808, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Ashforth:2008:IEC

[AY08] Dawn Ashforth and Norman D. Yan. The interactive effects of calcium concentration and temperature on the survival and reproduction of Daphnia pulex at high and low food concentrations. Limnology and Oceanography, 53(2):420–432, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Bidle:2001:BCS

[BA01] Kay D. Bidle and Farooq Azam. Bacterial control of silicon regeneration from diatom detritus: Significance of bacterial ectohydrolases and species identity. Limnology and Oceanography, 46 (7):1606–1623, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Baltar:2009:EPM

[BAG^{+}09] Federico Baltar, Javier Arístegui, Josep M. Gasol, Eva Sintes, and Gerhard J. Herndl. Evidence of prokaryotic metabolism on suspended particulate organic matter in the dark waters of the subtropical North Atlantic. Limnology and Oceanography, 54(1): 182–193, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Buesseler:2005:PED

[BAP^{+}05] K. O. Buesseler, J. E. Andrews, S. M. Pike, M. A. Charette, L. E. Goldson, M. A. Brzezinski, and V. P. Lance. Particle export during the Southern Ocean iron experiment (SOFeX). Limnology and Oceanography, 55(3):311–327, January 2005. CODEN LI-OCAH. ISSN 0024-3590.

Barea-Arco:2001:EMR

[BAPMMB01] J. Barea-Arco, C. Pérez-Martínez, and R. Morales-Baquero. Evidence of a mutualistic relationship between an algal epibiont and its host, Daphnia pulicaria. Limnology and Oceanography, 46(4):871–881, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Burkhardt:2001:CHU

[BARS01] Steffen Burkhardt, Gabi Amoroso, Ulf Riebesell, and Dieter Sültemeyer. CO_{2} and HCO_{3}^{−} uptake in marine diatoms acclimated to different CO_{2} concentrations. Limnology and Oceanography, 46(6):1378–1391, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Battin:2000:HMD

[Bat00] Tom J. Battin. Hydrodynamics is a major determinant of streambed biofilm activity: From the sediment to the reach scale. Limnology and Oceanography, 45(6):1308–1319, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Bates:2002:SVE

[Bat02] Nicholas R. Bates. Seasonal variability of the effect of coral reefs on seawater CO_{2} and air–sea CO_{2} exchange. Limnology and Oceanography, 47(1):43–52, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Becker:2003:RQE

[BB03] Claes Becker and Maarten Boersma. Resource quality effects on life histories of Daphnia. Limnology and Oceanography, 48(2): 700–706, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Barlocher:2004:CAH

[BB04a] Felix Bärlocher and Heinz Brendelberger. Clearance of aquatic hyphomycete spores by a benthic suspension feeder. Limnology and Oceanography, 49(6):2292–2296, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Bochdansky:2004:RSZ

[BB04b] Alexander B. Bochdansky and Stephen M. Bollens. Relevant scales in zooplankton ecology: Distribution, feeding, and reproduction of the copepod Acartia hudsonica in response to thin layers of the diatom Skeletonema costatum. Limnology and Oceanography, 49(3):625–636, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Becker:2005:DEP

[BB05] Claes Becker and Maarten Boersma. Differential effects of phosphorus and fatty acids on Daphnia magna growth and reproduction. Limnology and Oceanography, 55(3):388–397, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Buck:2007:PSD

[BB07] Kristen N. Buck and Kenneth W. Bruland. The physicochemical speciation of dissolved iron in the Bering Sea, Alaska. Limnology and Oceanography, 52(5):1800–1808, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Buesseler:2009:SLP

[BB09] Ken O. Buesseler and Philip W. Boyd. Shedding light on processes that control particle export and flux attenuation in the twilight zone of the open ocean. Limnology and Oceanography, 54(4):1210–1232, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Benoit-Bird:2009:NLL

[BBAW09] Kelly J. Benoit-Bird, Whitlow W. L. Au, and Daniel W. Wisdoma. Nocturnal light and lunar cycle effects on diel migration of micronekton. Limnology and Oceanography, 54(5):1789–1800, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Benoit-Bird:2009:EGP

[BBCW09] Kelly J. Benoit-Bird, Timothy J. Cowles, and Christopher E. Wingard. Edge gradients provide evidence of ecological interactions in planktonic thin layers. Limnology and Oceanography, 54 (4):1382–1392, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Broitman:2005:RII

[BBG05a] Bernardo R. Broitman, Carol A. Blanchette, and Steven D. Gaines. Recruitment of intertidal invertebrates and oceanographic variability at Santa Cruz Island, California. Limnology and Oceanography, 50(5):1473–1479, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Bruyant:2005:DVP

[BBG^{+}05b] F. Bruyant, M. Babin, B. Genty, O. Prasil, M. J. Behrenfeld, H. Claustre, A. Bricaud, L. Garczarek, J. Holtzendorff, M. Koblizek, H. Dousova, and F. Partensky. Diel variations in the photosynthetic parameters of Prochlorococcus strain PCC 9511: Combined effects of light and cell cycle. Limnology and Oceanography, 50(3):850–863, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Bergstrom:2005:EAN

[BBJ05] Ann-Kristin Bergström, Peter Blomqvist, and Mats Jansson. Effects of atmospheric nitrogen deposition on nutrient limitation and phytoplankton biomass in unproductive Swedish lakes. Limnology and Oceanography, 50(3):987–994, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Billett:2006:MDJ

[BBJ^{+}06] D. S. M. Billett, B. J. Bett, C. L. Jacobs, I. P. Rouse, and B. D. Wigham. Mass deposition of jellyfish in the deep Arabian Sea. Limnology and Oceanography, 51(5):2077–2083, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Bertilsson:2003:ECM

[BBKC03] S. Bertilsson, O. Berglund, D. M. Karl, and S. W. Chisholm. Elemental composition of marine Prochlorococcus and Synechococcus: Implications for the ecological stoichiometry of the sea. Limnology and Oceanography, 48(5):1721–1731, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Bidle:2003:DEO

[BBL^{+}03] Kay D. Bidle, Mark A. Brzezinski, Richard A. Long, Janice L. Jones, and Farooq Azam. Diminished efficiency in the oceanic silica pump caused by bacteria-mediated silica dissolution. Limnology and Oceanography, 48(5):1855–1868, September 2003. CODEN LIOCAH. ISSN 0024-3590.

BrindAmour:2005:MSDb

[BBLB05a] A. Brind’Amour, D. Boisclair, P. Legendre, and D. Borcard. Multiscale spatial distribution of a littoral fish community in relation to environmental variables. Limnology and Oceanography, 50(4):1344, July 2005. CODEN LIOCAH. ISSN 0024-3590.

BrindAmour:2005:MSDa

[BBLB05b] Anik Brind’Amour, Daniel Boisclair, Pierre Legendre, and Daniel Borcard. Multiscale spatial distribution of a littoral fish community in relation to environmental variables. Limnology and Oceanography, 50(2):465–479, March 2005. CODEN LI-OCAH. ISSN 0024-3590.

Brahney:2006:INL

[BBP^{+}06] Janice Brahney, Darren G. Bos, Marlow G. Pellatt, Thomas W. D. Edwards, and Richard Routledge. The influence of nitrogen limitation on δ ^{15}N and carbon : nitrogen ratios in sediments from sockeye salmon nursery lakes in British Columbia, Canada. Limnology and Oceanography, 51(5):2333– 2340, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Bradbury:2004:SHE

[BCBS04] Ian R. Bradbury, Steven E. Campana, Paul Bentzen, and Paul V. R. Snelgrove. Synchronized hatch and its ecological significance in rainbow smelt Osmerus mordax in St. Mary’s Bay, Newfoundland. Limnology and Oceanography, 49(6):2310–2315, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Bade:2004:CDL

[BCC^{+}04] Darren L. Bade, Stephen R. Carpenter, Jonathan J. Cole, Paul C. Hanson, and Raymond H. Hesslein. Controls of δ ^{13}C-DIC in lakes: Geochemistry, lake metabolism, and morphometry. Limnology and Oceanography, 49(4):1160–1172, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Bowen:2009:ISA

[BCDH09] J. L. Bowen, B. C. Crump, L. A. Deegan, and J. E. Hobbiec. Increased supply of ambient nitrogen has minimal effect on salt marsh bacterial production. Limnology and Oceanography, 54 (3):713–722, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Belviso:2001:EUC

[BCM01] S. Belviso, H. Claustre, and J-C. Marty. Evaluation of the utility of chemotaxonomic pigments as a surrogate for particulate DMSP. Limnology and Oceanography, 46(4):989–995, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Branstrator:2000:MLH

[BCMR00] Donn K. Branstrator, Gilbert Cabana, Asit Mazumder, and Joseph B. Rasmussen. Measuring life-history omnivory in the opossum shrimp, Mysis relicta, with stable nitrogen isotopes. Limnology and Oceanography, 45(2):463–467, March 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Bayley:2007:FRS

[BCSW07] S. E. Bayley, I. F. Creed, G. Z. Sass, and A. S. Wong. Frequent regime shifts in trophic states in shallow lakes on the boreal plain: Alternative “unstable” states? Limnology and Oceanography, 52(5):2002–2012, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Bec:2008:GRP

[BCV^{+}08] BÉatrice Bec, Yves Collos, André Vaquer, David Mouillot, and Philippe Souchu. Growth rate peaks at intermediate cell size in marine photosynthetic picoeukaryotes. Limnology and Oceanography, 57(3):863–867, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Baehr:2004:SPM

[BD04] Matthew M. Baehr and Michael D. DeGrandpre. In situ pCO_{2} and O_{2} measurements in a lake during turnover and stratification: Observations and modeling. Limnology and Oceanography, 49(2):330–340, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Bednarska:2007:CFS

[BD07] Anna Bednarska and Piotr Dawidowicz. Change in filter-screen morphology and depth selection: Uncoupled responses of Daphnia to the presence of filamentous cyanobacteria. Limnology and Oceanography, 52(6):2358–2363, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Bissett:2008:MMS

[BdBS^{+}08] Andrew Bissett, Dirk de Beer, Raphaela Schoon, Fumito Shiraishi, Andreas Reimer, and Gernot Arp. Microbial mediation of stromatolite formation in karst-water creeks. Limnology and Oceanography, 54(3):1159–1168, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Buffan-Dubau:2000:DFB

[BDC00] Evelyne Buffan-Dubau and Kevin R. Carman. Diel feeding behavior of meiofauna and their relationships with microalgal resources. Limnology and Oceanography, 45(2):381–395, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Bouvier:2002:CCF

[BdG02] Thierry C. Bouvier and Paul A. del Giorgio. Compositional changes in free-living bacterial communities along a salinity gradient in two temperate estuaries. Limnology and Oceanography, 47(2):453–470, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Berger:2006:LSP

[BDK^{+}06a] Stella A. Berger, Sebastian Diehl, Thomas J. Kunz, Dieter Albrecht, Amine M. Oucible, and Sylvie Ritzer. Light supply, plankton biomass, and seston stoichiometry in a gradient of lake mixing depths. Limnology and Oceanography, 53(2):1898–1905, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:2006:CTA

[BDK^{+}06b] JoAnn M. Burkholder, David A. Dickey, Carol A. Kinder, Robert E. Reed, Michael A. Mallin, Matthew R. McIver, Lawrence B. Cahoon, Greg Melia, Cavell Brownie, Joy Smith, Nora Deamer, Jeffrey Springer, Howard B. Glasgow, and David Toms. Comprehensive trend analysis of nutrients and related variables in a large eutrophic estuary: a decadal study of anthropogenic and climatic influences. Limnology and Oceanography, 51(1 (part 2)):463–487, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:2008:ISA

[BDK^{+}08] Mark A. Brzezinski, Cynthia Dumousseaud, Jeffrey W. Krause, Christopher I. Measures, and David M. Nelson. Iron and silicic acid concentrations together regulate Si uptake in the equatorial Pacific Ocean. Limnology and Oceanography, 53(3):875–889, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Budd:2001:RSB

[BDNF01] Judith W. Budd, Thomas D. Drummer, Thomas F. Nalepa, and Gary L. Fahnenstiel. Remote sensing of biotic effects: Zebra mussels (Dreissena polymorpha) influence on water clarity in Saginaw Bay, Lake Huron. Limnology and Oceanography, 46(2): 213–223, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Borges:2004:GTV

[BDS^{+}04] Alberto Vieira Borges, Bruno Delille, Laure-Sophie Schiettecatte, Frédéric Gazeau, Gwenaöl Abril, and Michel Frankignoulle. Gas transfer velocities of CO_{2} in three European estuaries (Randers Fjord, Scheldt, and Thames). Limnology and Oceanography, 49(5):1630–1641, September 2004. CODEN LI-OCAH. ISSN 0024-3590.

Bouillon:2007:BTE

[BDSB07] Steven Bouillon, Frank Dehairs, Laure-Sophie Schiettecatte, and Alberto Vieira Borges. Biogeochemistry of the Tana estuary and delta (northern Kenya). Limnology and Oceanography, 52(1):46– 59, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Bailey:2003:VID

[BDvO^{+}03] Sarah A. Bailey, Ian C. Duggan, Colin D. A. van Overdijk, Philip T. Jenkins, and Hugh J. MacIsaac. Viability of invertebrate diapausing eggs collected from residual ballast sediment. Limnology and Oceanography, 48(4):1701–1710, July 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Beaulieu:2003:RPS

[Bea03] Stace E. Beaulieu. Resuspension of phytodetritus from the sea floor: a laboratory flume study. Limnology and Oceanography, 48 (4):1235–1244, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Bell:2001:BRZ

[Bel01] Susan S. Bell. Book review: Zedler, J. B. 2000. Handbook for restoring tidal wetlands. CRC Press. 439 p. $89.95. ISBN 0-8493-9063-X. Limnology and Oceanography, 46(3):1583, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Bianchi:2002:DSC

[BEM^{+}02] Thomas S. Bianchi, Erika Engelhaupt, Brent A. McKee, Stephen Miles, Ragnar Elmgren, Susanna Hajdu, Candida Savage, and Mark Baskaran. Do sediments from coastal sites accurately reflect time trends in water column phytoplankton? A test from Himmerfjärden Bay (Baltic Sea proper). Limnology and Oceanography, 47(5):1537–1544, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Beninger:2000:LCC

[Ben00] Peter G. Beninger. Limits and constraints: a comment on premises and methods in recent studies of particle capture mechanisms in bivalves. Limnology and Oceanography, 45(5):1196– 1199, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Berman:2001:RDE

[Ber01] Tom Berman. The role of DON and the effect of N : P ratios on occurrence of cyanobacterial blooms: Implications from the outgrowth of aphanizomenon in Lake Kinneret. Limnology and Oceanography, 49(1):443–447, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Berglund:2003:PDI

[Ber03] Olof Berglund. Periphyton density influences organochlorine accumulation in rivers. Limnology and Oceanography, 48(6):2106– 2116, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Bianchi:2000:CBB

[BEW^{+}00] Thomas S. Bianchi, Erika Engelhaupt, Per Westman, Thomas Andrén, Carl Rolff, and Ragnar Elmgren. Cyanobacterial blooms in the Baltic Sea: Natural or human-induced? Limnology and Oceanography, 45(3):716–726, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Baumgartner:2008:DPB

[BF08a] Mark F. Baumgartner and David M. Fratantoni. Diel periodicity in both sei whale vocalization rates and the vertical migration of their copepod prey observed from ocean gliders. Limnology and Oceanography, 53(5 (part 2)):2197–2209, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Bellwood:2008:SMS

[BF08b] David R. Bellwood and Christopher J. Fulton. Sedimentmediated suppression of herbivory on coral reefs: Decreasing resilience to rising sea-levels and climate change? Limnology and Oceanography, 53(6):2695–2701, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Berman-Frank:2004:DMC

[BFBHF04] Ilana Berman-Frank, Kay D. Bidle, Liti Haramaty, and Paul G. Falkowski. The demise of the marine cyanobacterium, Trichodesmium spp., via an autocatalyzed cell death pathway. Limnology and Oceanography, 49(4):997–1005, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Baines:2005:UDO

[BFC05] Stephen B. Baines, Nicholas S. Fisher, and Jonathan J. Cole. Uptake of dissolved organic matter (DOM) and its importance to metabolic requirements of the zebra mussel, Dreissena polymorpha. Limnology and Oceanography, 50(1):36–47, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Baines:2007:DOM

[BFC07] Stephen B. Baines, Nicholas S. Fisher, and Jonathan J. Cole. Dissolved organic matter and persistence of the invasive zebra mussel (Dreissena polymorpha) under low food conditions. Limnology and Oceanography, 52(1):70–78, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Briones-FourzAn:2008:PSS

[BFCLÁ08] Patricia Briones-FourzÁn, Julio Candela, and Enrique Lozano-Álvarez. Postlarval settlement of the spiny lobster Panulirus argus along the Caribbean coast of Mexico: Patterns, influence of physical factors, and possible sources of origin. Limnology and Oceanography, 54(3):970–985, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Berman-Frank:2001:IAC

[BFCS^{+}01] Ilana Berman-Frank, Jay T. Cullen, Yeala Shaked, Robert M. Sherrell, and Paul G. Falkowski. Iron availability, cellular iron quotas, and nitrogen fixation in Trichodesmium. Limnology and Oceanography, 46(6):1249–1260, September 2001. CODEN LI-OCAH. ISSN 0024-3590.

Baines:2004:LDS

[BFD^{+}04] Stephen B. Baines, Nicholas S. Fisher, Martina A. Doblin, Gregory A. Cutter, Lynda S. Cutter, and Brian Cole. Light dependence of selenium uptake by phytoplankton and implications for predicting selenium incorporation into food webs. Limnology and Oceanography, 49(2):566–578, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Baines:2001:UDO

[BFDC01] Stephen B. Baines, Nicholas S. Fisher, Martina A. Doblin, and Gregory A. Cutter. Uptake of dissolved organic selenides by marine phytoplankton. Limnology and Oceanography, 46(8):1936– 1944, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Becker:2005:STV

[BFML05] Bonnie J. Becker, F. Joel Fodrie, Pat A. McMillan, and Lisa A. Levin. Spatial and temporal variation in trace elemental fingerprints of mytilid mussel shells: a precursor to invertebrate larval tracking. Limnology and Oceanography, 50(1):48–61, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Berman-Frank:2007:NFS

[BFQF^{+}07] Ilana Berman-Frank, Antonietta Quigg, Zoe V. Finkel, Andrew J. Irwin, and Liti Haramaty. Nitrogen-fixation strategies and Fe requirements in cyanobacteria. Limnology and Oceanography, 52(5):2260–2269, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Baines:2002:ARS

[BFS02] Stephen B. Baines, Nicholas S. Fisher, and Robin Stewart. Assimilation and retention of selenium and other trace elements from crustacean food by juvenile striped bass (Morone saxatilis). Limnology and Oceanography, 47(3):646–655, May 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Bergquist:2004:EDP

[BFS^{+}04] Derk C. Bergquist, Clint Fleckenstein, Emily B. Szalai, Julie Knisel, and Charles R. Fisher. Environment drives physiological variability in the cold seep mussel Bathymodiolus childressi. Limnology and Oceanography, 49(3):706–715, May 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Bearon:2004:RCL

[BGC04a] R. N. Bearon, D. Grünbaum, and R. A. Cattolico. Relating celllevel swimming behaviors to vertical population distributions in Heterosigma akashiwo (Raphidophyceae), a harmful alga. Limnology and Oceanography, 49(2):607–613, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Blain:2004:AIM

[BGC^{+}04b] Stèphane Blain, Cècile Guieu, Hervè Claustre, Karine Leblanc, Thierry Moutin, Bernard Quèguiner, Josèphine Ras, and Gèraldine Sarthou. Availability of iron and major nutrients for phytoplankton in the northeast Atlantic Ocean. Limnology and Oceanography, 49(6):2095–2104, November 2004. CODEN LI-OCAH. ISSN 0024-3590.

Bonnet:2005:EAN

[BGC^{+}05] Sophie Bonnet, Cécile Guieu, Jacques Chiaverini, Joséphine Ras, and Agnès Stock. Effect of atmospheric nutrients on the autotrophic communities in a low nutrient, low chlorophyll system. Limnology and Oceanography, 50(6):1810–1819, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Blomqvist:2004:WLN

[BGE04] Sven Blomqvist, Anneli Gunnars, and Ragnar Elmgren. Why the limiting nutrient differs between temperate coastal seas and freshwater lakes: a matter of salt. Limnology and Oceanography, 49(6):2236–2241, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Brinkmeyer:2004:PPS

[BGHA04] Robin Brinkmeyer, Frank-Oliver Glöckner, Elisabeth Helmke, and Rudolf Amann. Predominance of β-proteobacteria in summer melt pools on Arctic pack ice. Limnology and Oceanography, 49(4):1013–1021, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Boudreau:2001:RGIa

[BGJ01a] Bernard P. Boudreau, Bruce S. Gardiner, and Bruce D. Johnson. Rate of growth of isolated bubbles in sediments with a diagenetic source of methane. Limnology and Oceanography, 46(3):616–622, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Boudreau:2001:RGIb

[BGJ01b] Bernard P. Boudreau, Bruce S. Gardiner, and Bruce D. Johnson. Rate of growth of isolated bubbles in sediments with a diagenetic source of methane. Limnology and Oceanography, 46(6):1578, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Baker:2009:GEN

[BGR^{+}09] Jason W. Baker, James P. Grover, Ratheesh Ramachandrannair, Cody Black, Theodore W. Valenti, Jr., Bryan W. Brooks, and Daniel L. Roelkec. Growth at the edge of the niche: an experimental study of the harmful alga Prymnesium parvum. Limnology and Oceanography, 54(5):1679–1687, September 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Belzile:2002:CDO

[BGV02] Claude Belzile, John A. E. Gibson, and Warwick F. Vincent. Colored dissolved organic matter and dissolved organic carbon exclusion from lake ice: Implications for irradiance transmission and carbon cycling. Limnology and Oceanography, 47(5):1283– 1293, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Bullard:2002:PMM

[BH02] Stephan G. Bullard and Mark E. Hay. Palatability of marine macro-holoplankton: Nematocysts, nutritional quality, and chemistry as defenses against consumers. Limnology and Oceanography, 47(5):1456–1467, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Budinoff:2007:EML

[BH07] Charles R. Budinoff and James T. Hollibaugh. Ecophysiology of a Mono Lake picocyanobacterium. Limnology and Oceanography, 52(6):2484–2496, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Burdick:2007:ESC

[BHL07] Daniel S. Burdick, Daniel K. Hartline, and Petra H. Lenz. Escape strategies in co-occurring calanoid copepods. Limnology and Oceanography, 52(6):2373–2385, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Blindow:2006:PPT

[BHMS06] I. Blindow, A. Hargeby, J. Meyercordt, and H. Schubert. Primary production in two shallow lakes with contrasting plant form dominance: a paradox of enrichment? Limnology and Oceanography, 52(6):2711–2721, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Bohlke:2004:RSI

[BHV04] John Karl Böhlke, Judson W. Harvey, and Mary A. Voytek. Reach-scale isotope tracer experiment to quantify denitrification and related processes in a nitrate-rich stream, midcontinent United States. Limnology and Oceanography, 49(3):821– 838, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Bachmann:2005:OFM

[BHVC05] Roger W. Bachmann, Mark V. Hoyer, Susana B. Vinzon, and Daniel, Jr. E. Canfield. The origin of the fluid mud layer in Lake Apopka, Florida. Limnology and Oceanography, 50(4):629–635, March 2005. CODEN LIOCAH. ISSN 0024-3590. See comment [Sch06a].

Bachmann:2006:ALO

[BHVC06] Roger W. Bachmann, Mark V. Hoyer, Susana B. Vinzon, and Daniel E. Canfield, Jr. Another look at the origin of the fluid mud layer in Lake Apopka, Florida. Limnology and Oceanography, 51(5):2480–2483, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Botelho:2007:DOR

[BI07] Daniel A. Botelho and Jörg Imberger. Dissolved oxygen response to wind-inflow interactions in a stratified reservoir. Limnology and Oceanography, 52(5):2027–2052, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Bianchi:2001:ESS

[Bia01] Thomas S. Bianchi. Estuarine science: a synthetic approach to research and practice. Limnology and Oceanography, 46(6):746, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Boegman:2005:DIW

[BII05] L. Boegman, G. N. Ivey, and J. Imberger. The degeneration of internal waves in lakes with sloping topography. Limnology and Oceanography, 50(5):1620–1637, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Boegman:2003:HFI

[BIIA03] L. Boegman, J. Imberger, G. N. Ivey, and J. P. Antenucci. Highfrequency internal waves in large stratified lakes. Limnology and Oceanography, 48(2):895–919, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:2003:BBS

[BJBA03] Mark A. Brzezinski, Janice L. Jones, Kay D. Bidle, and Farooq Azam. The balance between silica production and silica dissolution in the sea: Insights from Monterey Bay, California, applied to the global data set. Limnology and Oceanography, 48(5):1846– 1854, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:2005:CSP

[BJD05] Mark A. Brzezinski, Janice L. Jones, and Mark S. Demarest. Control of silica production by iron and silicic acid during the Southern Ocean iron experiment (SOFeX). Limnology and Oceanography, 50(3):810–824, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Belzile:2000:UAD

[BJG^{+}00] Claude Belzile, Sophia C. Johannessen, Michel Gosselin, Serge Demers, and William L. Miller. Ultraviolet attenuation by dissolved and particulate constituents of first-year ice during late spring in an Arctic polynya. Limnology and Oceanography, 45 (6):1265–1273, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Burks:2001:LZS

[BJL01] Romi L. Burks, Erik Jeppesen, and David M. Lodge. Littoral zone structures as Daphnia refugia against fish predators. Limnology and Oceanography, 46(2):230–237, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Benoit:2002:AIR

[BJW^{+}02] Hugues P. Benoît, Ora E. Johannsson, David M. Warner, W. Gary Sprules, and Lars G. Rudstam. Assessing the impact of a recent predatory invader: The population dynamics, vertical distribution, and potential prey of Cercopagis pengoi in Lake Ontario. Limnology and Oceanography, 47(3):626–635, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Bell:2001:CPM

[BK01a] Thomas Bell and Jacob Kalff. The contribution of picophytoplankton in marine and freshwater systems of different trophic status and depth. Limnology and Oceanography, 46(5):1243– 1248, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Byrne:2001:LCW

[BK01b] Robert H. Byrne and Eric Kaltenbacher. Liquid capillary waveguide application in absorbance spectroscopy (reply to the comment by Byrne and Kaltenbacher). Limnology and Oceanography, 46(6):742–745, May 2001. CODEN LIOCAH. ISSN 0024-3590. See [BK01c].

Byrne:2001:ULC

[BK01c] Robert H. Byrne and Eric Kaltenbacher. Use of liquid core waveguides for long pathlength absorbance spectroscopy: Principles and practice. Limnology and Oceanography, 46(6):740– 742, May 2001. CODEN LIOCAH. ISSN 0024-3590. See reply [BK01b].

Benner:2003:AAS

[BK03a] Ronald Benner and Karl Kaiser. Abundance of amino sugars and peptidoglycan in marine particulate and dissolved organic matter. Limnology and Oceanography, 48(1):118–128, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Bjorkman:2003:BDO

[BK03b] Karin M. Björkman and David M. Karl. Bioavailability of dissolved organic phosphorus in the euphotic zone at Station ALOHA, North Pacific Subtropical Gyre. Limnology and Oceanography, 48(3):1049–1057, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Bruggeman:2007:BIA

[BK07] Jorn Bruggeman and Sebastiaan A. L. M. Kooijman. A biodiversity-inspired approach to aquatic ecosystem modeling. Limnology and Oceanography, 52(4):1533–1544, July 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Bagoien:2001:VDM

[BKAE01] Espen Bagøien, Stein Kaartvedt, Dag L. Aksnes, and Ketil Eiane. Vertical distribution and mortality of overwintering Calanus. Limnology and Oceanography, 46(6):1494–1510, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Bennett:2002:PVM

[BKB02] William A. Bennett, Wim J. Kimmerer, and Jon R. Burau. Plasticity in vertical migration by native and exotic estuarine fishes in a dynamic low-salinity zone. Limnology and Oceanography, 47(5):1496–1507, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Bouillon:2006:PAT

[BKK^{+}06] René-Christian Bouillon, Tika L. Knierim, Robert J. Kieber, Stephen A. Skrabal, and Jeffrey L. C. Wright. Photodegradation of the algal toxin domoic acid in natural water matrices. Limnology and Oceanography, 51(1):321–330, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Boschker:2005:BCI

[BKM05] H. T. S. Boschker, J. C. Kromkamp, and J. J. Middelburg. Biomarker and carbon isotopic constraints on bacterial and algal community structure and functioning in a turbid, tidal estuary. Limnology and Oceanography, 50(1):70–80, January 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Blankenship:2007:EFW

[BL07] Lesley E. Blankenship and Lisa A. Levin. Extreme food webs: Foraging strategies and diets of scavenging amphipods from the ocean’s deepest 5 kilometers. Limnology and Oceanography, 52 (4):1685–1697, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Buck:2007:DIS

[BLBB07] Kristen N. Buck, Maeve C. Lohan, Carolyn J. M. Berger, and Kenneth W. Bruland. Dissolved iron speciation in two distinct river plumes and an estuary: Implications for riverine iron supply. Limnology and Oceanography, 52(2):843–855, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Bunting:2007:DWQ

[BLG^{+}07] Lynda Bunting, Peter R. Leavitt, Christopher E. Gibson, Edward J. McGee, and Valerie A. Hall. Degradation of water quality in Lough Neagh, Northern Ireland, by diffuse nitrogen flux from a phosphorus-rich catchment. Limnology and Oceanography, 52(1):354–369, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Boegman:2008:VMW

[BLHC08] L. Boegman, M. R. Loewen, P. F. Hamblin, and D. A. Culver. Vertical mixing and weak stratification over zebra mussel colonies in western Lake Erie. Limnology and Oceanography, 54 (3):1093–1110, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Berggren:2009:AAO

[BLJ09] Martin Berggren, Hjalmar Laudon, and Mats Jansson. Aging of allochthonous organic carbon regulates bacterial production in unproductive boreal lakes. Limnology and Oceanography, 54(4): 1333–1342, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Buhring:2006:BMW

[BLM^{+}06] S. I. Bühring, N. Lampadariou, L. Moodley, A. Tselepides, and U. Witte. Benthic microbial and whole-community responses to different amounts of ^{13}C-enriched algae: In situ experiments in the deep Cretan Sea (Eastern Mediterranean). Limnology and Oceanography, 51(1):157–165, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Brock:2006:SVS

[BLS^{+}06] Curtis S. Brock, Peter R. Leavitt, Daniel E. Schindler, Susan P. Johnson, and Jonathan W. Moore. Spatial variability of stable isotopes and fossil pigments in surface sediments of Alaskan coastal lakes: Constraints on quantitative estimates of past salmon abundance. Limnology and Oceanography, 52(2): 1637–1647, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Brock:2007:VEM

[BLSQ07] Curtis S. Brock, Peter R. Leavitt, Daniel E. Schindler, and Paul D. Quay. Variable effects of marine-derived nutrients on algal production in salmon nursery lakes of Alaska during the past 300 years. Limnology and Oceanography, 52(4):1588–1598, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Beer:2008:GCA

[BLWB08] Julia Beer, Kern Lee, Michael Whiticar, and Christian Blodau. Geochemical controls on anaerobic organic matter decomposition in a northern peatland. Limnology and Oceanography, 53 (4):1393–1407, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Barbeau:2000:LFS

[BM00] K. Barbeau and J. W. Moffett. Laboratory and field studies of colloidal iron oxide dissolution as mediated by phagotrophy and photolysis. Limnology and Oceanography, 45(4):827–835, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Berg:2001:HRP

[BM01] Peter Berg and Karen J. McGlathery. A high-resolution pore water sampler for sandy sediments. Limnology and Oceanography, 46(1):203–210, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Buffam:2003:EUL

[BM03] Ishi Buffam and Karen J. McGlathery. Effect of ultraviolet light on dissolved nitrogen transformations in coastal lagoon water. Limnology and Oceanography, 48(2):723–734, March 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Berglund:2007:EPB

[BMBA07] Johnny Berglund, Umut Müren, Ulf Båmstedt, and Agneta Andersson. Efficiency of a phytoplankton-based and a bacterialbased food web in a pelagic marine system. Limnology and Oceanography, 52(1):121–131, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Brooks:2007:PCC

[BMC07] Marjorie L. Brooks, Diane M. McKnight, and William H. Clements. Photochemical control of copper complexation by dissolved organic matter in Rocky Mountain streams, Colorado. Limnology and Oceanography, 52(2):766–779, March 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Bec:2006:TUA

[BMCvE06] Alexandre Bec, Dominik Martin-Creuzburg, and Eric von Elert. Trophic upgrading of autotrophic picoplankton by the heterotrophic nanoflagellate Paraphysomonas sp. Limnology and Oceanography, 52(2):1699–1707, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Biswas:2004:BCA

[BMD^{+}04] H. Biswas, S. K. Mukhopadhyay, T. K. De, S. Sen, and T. K. Jana. Biogenic controls on the air–water carbon dioxide exchange in the Sundarban mangrove environment, northeast coast of Bay of Bengal, India. Limnology and Oceanography, 49 (1):95–101, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Barrin:2006:PTW

[BMD06] Cristina Barrín, Jack J. Middelburg, and Carlos M. Duarte. Phytoplankton trapped within seagrass (Posidonia oceanica) sediments are a nitrogen source: an in situ isotope labeling experiment. Limnology and Oceanography, 52(2):1648–1653, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Barnard:2006:TPZ

[BMF^{+}06] Christine Barnard, Christine Martineau, Jean-Jacques Frenette, Julian J. Dodson, and Warwick F. Vincent. Trophic position of zebra mussel veligers and their use of dissolved organic carbon. Limnology and Oceanography, 51(3):1473–1484, May 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Babin:2003:LSP

[BMFS^{+}03] Marcel Babin, André Morel, Vincent Fournier-Sicre, Frank Fell, and Dariusz Stramski. Light scattering properties of marine particles in coastal and open ocean waters as related to the particle mass concentration. Limnology and Oceanography, 48(2): 843–859, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Boutin:2008:ASC

[BMH^{+}08] J. Boutin, L. Merlivat, C. Hénocq, N. Martin, and J. B. Sallée. Air–sea CO_{2} flux variability in frontal regions of the Southern Ocean from CARbon Interface OCean Atmosphere drifters. Limnology and Oceanography, 53(5 (part 2)):2062–2079, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Brett:2006:DFA

[BMNB^{+}06] Michael T. Brett, Dörthe C. Müller-Navarra, Ashley P. Ballantyne, Joseph L. Ravet, and Charles R. Goldman. Daphnia fatty acid composition reflects that of their diet. Limnology and Oceanography, 51(5):2428–2437, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Brett:2000:EAE

[BMNSK00] Michael T. Brett, Dörthe C. Müller-Navarra, and Park Sang-Kyu. Empirical analysis of the effect of phosphorus limitation on algal food quality for freshwater zooplankton. Limnology and Oceanography, 45(7):1564–1575, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Bissinger:2008:PMP

[BMSA08] Jan E. Bissinger, David J. S. Montagnes, Jonathan Sharples, and David Atkinson. Predicting marine phytoplankton maximum growth rates from temperature: Improving on the Eppley curve using quantile regression. Limnology and Oceanography, 53(2): 487–493, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Barron:2004:CMC

[BMT^{+}04] Cristina Barrón, Nùria Marbé, Jorge Terrados, Hilary Kennedy, and Carlos M. Duarte. Community metabolism and carbon budget along a gradient of seagrass (Cymodocea nodosa) colonization. Limnology and Oceanography, 49(5):1642–1651, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Bloom:2004:CSF

[BMU04] Nicolas S. Bloom, Ligia M. Moretto, and Paolo Ugo. A comparison of the speciation and fate of mercury in two contaminated coastal marine ecosystems: The Venice Lagoon (Italy) and Lavaca Bay (Texas). Limnology and Oceanography, 49(2): 367–375, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Brand:2008:IOF

[BMWW08] Andreas Brand, Daniel F. McGinnis, Bernhard Wehrli, and Alfred Wüest. Intermittent oxygen flux from the interior into the bottom boundary of lakes as observed by eddy correlation. Limnology and Oceanography, 53(5):1997–2006, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Banaszak:2001:URS

[BN01] Anastazia T. Banaszak and Patrick J. Neale. Ultraviolet radiation sensitivity of photosynthesis in phytoplankton from an estuarine environment. Limnology and Oceanography, 46(3):592– 603, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Benitez-Nelson:2004:UVV

[BNC04] Claudia R. Benitez-Nelson and Matthew A. Charette. Uncertainty versus variability in upper ocean carbon flux estimates. Limnology and Oceanography, 49(4):1218–1220, July 2004. CO-DEN LIOCAH. ISSN 0024-3590. See reply [MWE^{+}04].

Benitez-Nelson:2002:PCN

[BNK02] Claudia R. Benitez-Nelson and David M. Karl. Phosphorus cycling in the North Pacific Subtropical Gyre using cosmogenic 32P and ^{33}P. Limnology and Oceanography, 47(3):762–770, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Breier:2009:TRC

[BNM^{+}09] J. A. Breier, N. Nidzieko, S. Monismith, W. Moore, and A. Paytan. Tidally regulated chemical fluxes across the sediment–water interface in Elkhorn Slough, California: Evidence from a coupled geochemical and hydrodynamic approach. Limnology and Oceanography, 54(6):1964–1980, November 2009. CODEN LI-OCAH. ISSN 0024-3590.

Benitez-Nelson:2004:PPO

[BNOK^{+}04] C. R. Benitez-Nelson, Lauren O’Neill, Lauren C. Kolowith, Perry Pellechia, and Robert Thunell. Phosphonates and particulate organic phosphorus cycling in an anoxic marine basin. Limnology and Oceanography, 49(5):1593–1604, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Biddanda:2001:DBM

[BOC01] Bopaiah Biddanda, Megan Ogdahl, and James Cotner. Dominance of bacterial metabolism in oligotrophic relative to eutrophic waters. Limnology and Oceanography, 46(6):730–739, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Boersma:2000:NQP

[Boe00] Maarten Boersma. The nutritional quality of P-limited algae for Daphnia. Limnology and Oceanography, 45(5):1157–1161, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Bootsma:2000:EDT

[Boo00] Harvey A. Bootsma. Ecological dynamics of tropical inland waters. Limnology and Oceanography, 60(3):523, March 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Boudreau:2004:WCM

[Bou04] Bernard P. Boudreau. What controls the mixed-layer depth in deep-sea sediments? The importance of particulate organic carbon flux. Limnology and Oceanography, 49(2):620–622, March 2004. CODEN LIOCAH. ISSN 0024-3590. See reply [SR04].

Boyle:2001:BRA

[Boy01] Edward A. Boyle. Book review: Alley, Richard B. 2000. The twomile time machine: Ice cores, abrupt climate change, and our future. Princeton University Press. 229 p. US$24.95. ISBN 0-691-00493-5. Limnology and Oceanography, 46(3):1580, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Bayley:2003:DWL

[BP03] Suzanne E. Bayley and Colleen M. Prather. Do wetland lakes exhibit alternative stable states? Submersed aquatic vegetation and chlorophyll in western boreal shallow lakes. Limnology and Oceanography, 48(6):2335–2345, November 2003. CODEN LI-OCAH. ISSN 0024-3590.

Baums:2006:BOF

[BPC06] Iliana B. Baums, Claire B. Paris, and Laurent M. Chérubin. A bio-oceanographic filter to larval dispersal in a reef-building coral. Limnology and Oceanography, 51(5):1969–1981, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Buitenhuis:2008:GRS

[BPF^{+}08] Erik T. Buitenhuis, Tanja Pangerc, Daniel J. Franklin, Corinne Le Quéré, and Gill Malin. Growth rates of six coccolithophorid strains as a function of temperature. Limnology and Oceanography, 54(3):1181–1185, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Brodersen:2004:CDO

[BPLH04] Klaus Peter Brodersen, Ole Pedersen, Claus Lindegaard, and Kirsten Hamburger. Chironomids (Diptera) and oxy-regulatory capacity: an experimental approach to paleolimnological interpretation. Limnology and Oceanography, 49(5):1549–1559, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Bastviken:2004:DDO

[BPOT04] David Bastviken, Linn Persson, Göran Odham, and Lars Tranvik. Degradation of dissolved organic matter in oxic and anoxic lake water. Limnology and Oceanography, 49(1):109–116, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Button:2000:ENK

[BR00] D. K. Button and Betsy Robertson. Effect of nutrient kinetics and cytoarchitecture on bacterioplankter size. Limnology and Oceanography, 60(3):499–505, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Bindler:2001:WBS

[BRB^{+}01] Richard Bindler, Ingemar Renberg, Maja-Lena Brännvall, Ove Emteryd, and Farid El-Daoushy. A whole-basin study of sediment accumulation using stable lead isotopes and flyash particles in an acidified lake, Sweden. Limnology and Oceanography, 46 (1):178–188, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Bittner:2002:EIM

[BRE02] Kerstin Bittner, Karl-Otto Rothhaupt, and Dieter Ebert. Ecological interactions of the microparasite Caullerya mesnili and its host Daphnia galeata. Limnology and Oceanography, 47(1): 300–305, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Brinson:2001:BRK

[Bri01] Mark M. Brinson. Book review: Keddy, Paul A. 2000. Wetland ecology: Principles and conservation. Cambridge University Press, xiv + 614 p. US$52.95, paper. ISBN 0-521-78367-4. Limnology and Oceanography, 46(3):1581, September 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Berg:2003:RPA

[BRL03] Gry Mine Berg, Daniel J. Repeta, and Julie LaRoche. The role of the picoeukaryote Aureococcus anophagefferens in cycling of marine high-molecular weight dissolved organic nitrogen. Limnology and Oceanography, 48(5):1825–1830, September 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Boscarino:2007:ETP

[BRM^{+}07] Brent T. Boscarino, Lars G. Rudstam, Shylene Mata, Gideon Gal, Ora E. Johannsson, and Edward L. Mills. The effects of temperature and predator–prey interactions on the migration behavior and vertical distribution of Mysis relicta. Limnology and Oceanography, 52(4):1599–1613, July 2007. CODEN LI-OCAH. ISSN 0024-3590.

Breton:2006:HMD

[BRP^{+}06] Elsa Breton, Véronique Rousseau, Jean-Yves Parent, José Ozer, and Christiane Lancelot. Hydroclimatic modulation of diatom/ Phaeocystis blooms in nutrient-enriched Belgian coastal waters (North Sea). Limnology and Oceanography, 51(3):1401–1409, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Bruland:2001:IMC

[BRS01] Kenneth W. Bruland, Eden L. Rue, and Geoffrey J. Smith. Iron and macronutrients in California coastal upwelling regimes: Implications for diatom blooms. Limnology and Oceanography, 46 (7):1661–1674, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Barbeau:2003:PRS

[BRT^{+}03] Katherine Barbeau, Eden L. Rue, Charles G. Trick, Kenneth W. Bruland, and Alison Butler. Photochemical reactivity of siderophores produced by marine heterotrophic bacteria and cyanobacteria based on characteristic Fe(III) binding groups. Limnology and Oceanography, 48(3):1069–1078, May 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Berg:2007:ECF

[BRW07] Peter Berg, Hans Røy, and Patricia L. Wiberg. Eddy correlation flux measurements: The sediment surface area that contributes to the flux. Limnology and Oceanography, 52(4):1672–1684, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Burns:2001:CCV

[BS01] Carolyn W. Burns and Marc Schallenberg. Calanoid copepods versus cladocerans: Consumer effects on protozoa in lakes of different trophic status. Limnology and Oceanography, 46(6): 1558–1565, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Babin:2002:LAA

[BS02a] Marcel Babin and Dariusz Stramski. Light absorption by aquatic particles in the near-infrared spectral region. Limnology and Oceanography, 47(3):911–915, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Biggs:2002:TRS

[BS02b] Barry J. F. Biggs and Robert A. Smith. Taxonomic richness of stream benthic algae: Effects of flood disturbance and nutrients. Limnology and Oceanography, 47(4):1175–1186, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Bucciarelli:2003:ICN

[BS03] Eva Bucciarelli and William G. Sunda. Influence of CO_{2}, nitrate, phosphate, and silicate limitation on intracellular dimethylsulfoniopropionate in batch cultures of the coastal diatom Thalassiosira pseudonana. Limnology and Oceanography, 48(6):2256– 2265, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Babin:2004:VMS

[BS04] Marcel Babin and Dariusz Stramski. Variations in the massspecific absorption coefficient of mineral particles suspended in water. Limnology and Oceanography, 49(3):756–767, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Brenner:2000:BAG

[BSA^{+}00] Mark Brenner, Joseph M. Smoak, Micheal S. Allen, Claire L. Schelske, and Douglas A. Leeper. Biological accumulation of 226Ra in a groundwater-augmented Florida lake. Limnology and Oceanography, 45(3):710–715, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Bauersachs:2009:NIF

[BSC^{+}09] Thorsten Bauersachs, Stefan Schouten, Justine Compaoré, Ute Wollenzien, Lucas J. Stal, and Jaap S. Sinninghe Damsteé. Nitrogen isotopic fractionation associated with growth on dinitrogen gas and nitrate by cyanobacteria. Limnology and Oceanography, 54(4):1403–1411, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Bremigan:2008:HFM

[BSG^{+}08] Mary T. Bremigan, Patricia A. Soranno, Maria J. Gonzalez, David B. Bunnell, Kristi K. Arend, William H. Renwick, Roy A. Stein, and Michael J. Vanni. Hydrogeomorphic features mediate the effects of land use/cover on reservoir productivity and food webs. Limnology and Oceanography, 53(4):1420–1433, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Binzer:2002:PAM

[BSJ02] Thomas Binzer and Kaj Sand-Jensen. Production in aquatic macrophyte communities: a theoretical and empirical study of the influence of spatial light distribution. Limnology and Oceanography, 47(6):1742–1750, November 2002. CODEN LI-OCAH. ISSN 0024-3590.

Binzer:2006:CPA

[BSJM06] Thomas Binzer, Kaj Sand-Jensen, and Anne-Lise Middelboe. Community photosynthesis of aquatic macrophytes. Limnology and Oceanography, 52(6):2722–2733, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Brenner:2007:RAF

[BSL^{+}07] Mark Brenner, Joseph M. Smoak, Douglas A. Leeper, Matt Streubert, and Shirley M. Baker. Radium-226 accumulation in Florida freshwater mussels. Limnology and Oceanography, 52(4): 1614–1623, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Bates:2001:BPF

[BSM01] Nicholas R. Bates, Leone Samuels, and Liliane Merlivat. Biogeochemical and physical factors influencing seawater fCO_{2} and air-sea CO_{2} exchange on the Bermuda coral reef. Limnology and Oceanography, 46(4):833–846, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Balseiro:2008:LTL

[BSMR08] Esteban Balseiro, María Sol Souza, Beatriz Modenutti, and Mariana Reissig. Living in transparent lakes: Low food P : C ratio decreases antioxidant response to ultraviolet radiation in Daphnia. Limnology and Oceanography, 53(6):2383–2390, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Bertrand:2007:VBI

[BSR^{+}07] Erin M. Bertrand, Mak A. Saito, Julie M. Rose, Christina R. Riesselman, Maeve C. Lohan, Abigail E. Noble, Peter A. Lee, and Giacomo R. DiTullio. Vitamin B_{12} and iron colimitation of phytoplankton growth in the Ross Sea. Limnology and Oceanography, 52(3):1079–1093, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Blais:2001:FSO

[BSS^{+}01] Jules M. Blais, David W. Schindler, Martin Sharp, Eric Braekevelt, Melissa Lafreniěre, Karen McDonald, Derek C. G. Muir, and William M. J. Strachan. Fluxes of semivolatile organochlorine compounds in Bow Lake, a high-altitude, glacierfed, subalpine lake in the Canadian Rocky Mountains. Limnology and Oceanography, 47(4):2019–2031, November 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Blukacz:2009:TUR

[BSS09] E. A. Blukacz, B. J. Shuter, and W. G. Sprules. Towards understanding the relationship between wind conditions and plankton patchiness. Limnology and Oceanography, 54(5):1530–1540, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Brown:2003:USD

[BSSL03] Louise Brown, Richard Sanders, Graham Savidge, and Cathy H. Lucas. The uptake of silica during the spring bloom in the Northeast Atlantic Ocean. Limnology and Oceanography, 48(5):1831– 1845, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Bartholmee:2005:SEF

[BSSL05] Silvia Bartholmeé, Laryssa Samchyshyna, Barbara Santer, and Winfried Lampert. Subitaneous eggs of freshwater copepods pass through fish guts: Survival, hatchability, and potential ecological implications. Limnology and Oceanography, 50(3):923– 929, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Baulch:2005:EWB

[BST^{+}05a] H. M. Baulch, D. W. Schindler, M. A. Turner, D. L. Findlay, M. J. Paterson, and R. D. Vinebrooke. Effects of warming on benthic communities in a boreal lake: Implications of climate change. Limnology and Oceanography, 50(5):1377–1392, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Boyd:2005:ETI

[BST^{+}05b] Philip W. Boyd, Robert Strzepek, Shigenobu Takeda, George Jackson, C. S. Wong, R. Mike McKay, Cliff Law, Hiroshi Kiyosawa, Hiroaki Saito, Nelson Sherry, Keith Johnson, Jim Gower, and Neelam Ramaiah. The evolution and termination of an iron-induced mesoscale bloom in the northeast subarctic Pacific. Limnology and Oceanography, 50(6):1872–1886, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Boss:2008:OPP

[BST^{+}08] E. Boss, D. Swift, L. Taylor, P. Brickley, R. Zaneveld, S. Riser, M. J. Perry, and P. G. Strutton. Observations of pigment and particle distributions in the western North Atlantic from an autonomous float and ocean color satellite. Limnology and Oceanography, 53(5 (part 2)):2112–2122, September 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Bertilsson:2000:PTD

[BT00] Stefan Bertilsson and Lars J. Tranvik. Photochemical transformation of dissolved organic matter in lakes. Limnology and Oceanography, 45(4):753–762, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Boehme:2008:MDP

[BTB^{+}08] L. Boehme, S. E. Thorpe, M. Biuw, M. Fedak, and M. P. Meredith. Monitoring Drake Passage with elephant seals: Frontal structures and snapshots of transport. Limnology and Oceanography, 53(5 (part 2)):2350–2360, September 2008. CODEN LI-OCAH. ISSN 0024-3590.

Buitenhuis:2003:ZBC

[BTdB03] Erik T. Buitenhuis, Klaas R. Timmermans, and Hein J. W. de Baar. Zinc-bicarbonate colimitation of Emiliania huxleyi. Limnology and Oceanography, 48(4):1575–1582, July 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Ben-Tzvi:2007:PED

[BTKGA07] Ofer Ben-Tzvi, Moshe Kiflawi, Hezi Gildor, and Avigdor Abelson. Possible effects of downwelling on the recruitment of coral reef fishes to the Eilat (Red Sea) coral reefs. Limnology and Oceanography, 52(6):2618–2628, November 2007. CODEN LI-OCAH. ISSN 0024-3590.

Bozec:2006:APC

[BTS^{+}06] Yann Bozec, Helmuth Thomas, Laure-Sophie Schiettecatte, Alberto V. Borges, Khalid Elkalay, and Hein J. W. de Baar. Assessment of the processes controlling the seasonal variations of dissolved inorganic carbon in the North Sea. Limnology and Oceanography, 52(6):2746–2762, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Bernhard:2003:SLP

[BVB03] Joan M. Bernhard, Pieter T. Visscher, and Samuel S. Bowser. Submillimeter life positions of bacteria, protists, and metazoans in laminated sediments of the Santa Barbara Basin. Limnology and Oceanography, 48(2):813–828, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:2007:BFQ

[BvE07] Tomasz Brzeziński and Eric von Elert. Biochemical food quality effects on a Daphnia hybrid complex. Limnology and Oceanography, 52(6):2350–2357, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Balch:2007:BOC

[BVG^{+}07] W. M. Balch, J. M. Vaughn, J. I. Goes, J. F. Novotny, D. T. Drapeau, E. S. Booth, and C. L. Vining. Bio-optical consequences of viral infection of phytoplankton: I. Experiments with the cyanobacterium, Synechococcus sp. Limnology and Oceanography, 52(2):727–738, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Barbieri:2002:EAU

[BVH02] Elena S. Barbieri, Virginia E. Villafanñ, and E. Walter Helbling. Experimental assessment of UV effects on temperate marine phytoplankton when exposed to variable radiation regimes. Limnology and Oceanography, 47(6):1648–1655, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Belzile:2002:CAS

[BVK02] Claude Belzile, Warwick F. Vincent, and Michio Kumagai. Contribution of absorption and scattering to the attenuation of UV and photosynthetically available radiation in Lake Biwa. Limnology and Oceanography, 47(1):95–107, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Balch:2000:LSV

[BVN^{+}00] William M. Balch, James Vaughn, James Novotny, David T. Drapeau, Robert Vaillancourt, Janeen Lapierre, and Amanda Ashe. Light scattering by viral suspensions. Limnology and Oceanography, 45(2):492–498, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Balch:2002:FCL

[BVN^{+}02] William M. Balch, James M. Vaughn, Jr., James F. Novotny, David T. Drapeau, Joaquim I. Goes, Emily Booth, Janeen M. Lapierre, Cathy L. Vining, and Amanda Ashe. Fundamental changes in light scattering associated with infection of marine bacteria by bacteriophage. Limnology and Oceanography, 47(5): 1554–1561, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Baudoux:2007:VMA

[BVWB07] Anne-Claire Baudoux, Marcel J. W. Veldhuis, Harry J. Witte, and Corina P. D. Brussaard. Viruses as mortality agents of picophytoplankton in the deep chlorophyll maximum layer during IRONAGES III. Limnology and Oceanography, 52(6):2519–2529, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Bronk:1999:GNN

[BW99] D. A. Bronk and B. B. Ward. Gross and net nitrogen uptake and DON release in the euphotic zone of Monterey Bay, California. Limnology and Oceanography, 44(3):573–585, May 1999. CODEN LIOCAH. ISSN 0024-3590. See reply [SRG00].

Bronk:2000:MDO

[BW00] Deborah A. Bronk and Bess B. Ward. Magnitude of dissolved organic nitrogen release relative to gross nitrogen uptake in marine systems. Limnology and Oceanography, 45(8):1879–1883, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Baker:2001:UBR

[BW01] Scott, Jr. M. Baker and Charles A. Wilson. Use of bomb radiocarbon to validate otolith section ages of red snapper Lutjanus campechanus from the northern Gulf of Mexico. Limnology and Oceanography, 46(7):1819–1824, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Berman:2005:APL

[BW05] Tom Berman and David Wynne. Assessing phytoplankton lysis in Lake Kinneret. Limnology and Oceanography, 50(4):526–537, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Boechat:2007:BES

[BWK^{+}07] Iola G. Boëchat, Guntram Weithoff, Angela Krüger, Björn Gücker, and Rita Adrian. A biochemical explanation for the success of mixotrophy in the flagellate Ochromonas sp. Limnology and Oceanography, 52(4):1624–1632, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Birch:2009:PLP

[BYF09] Daniel A. Birch, William R. Young, and Peter J. S. Franks. Plankton layer profiles as determined by shearing, sinking, and swimming. Limnology and Oceanography, 54(1):397–399, January 2009. CODEN LIOCAH. ISSN 0024-3590. See response [SMS09].

Burdige:2002:ISG

[BZ02] David J. Burdige and Richard C. Zimmerman. Impact of sea grass density on carbonate dissolution in Bahamian sediments. Limnology and Oceanography, 47(6):1751–1763, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Boss:2003:EBS

[BZ03] Emmanuel Boss and J. Ronald V. Zaneveld. The effect of bottom substrate on inherent optical properties: Evidence of biogeochemical processes. Limnology and Oceanography, 48(1 (part 2)):346–354, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Berthon:2000:MOM

[BZH00] Jean-Francois Berthon, Giuseppe Zibordi, and Stanford B. Hooker. Marine optical measurements of a mucilage event in the northern Adriatic Sea. Limnology and Oceanography, 45(2): 322–327, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Burdige:2008:RCD

[BZH08] David J. Burdige, Richard C. Zimmerman, and Xinping Hu. Rates of carbonate dissolution in permeable sediments estimated from pore-water profiles: The role of sea grasses. Limnology and Oceanography, 53(2):549–565, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Bockelmann:2007:AEC

[BZS07] Frank-D. Bockelmann, Karin A. F. Zonneveld, and Martin Schmidt. Assessing environmental control on dinoflagellate cyst distribution in surface sediments of the Benguela upwelling region (eastern South Atlantic). Limnology and Oceanography, 52 (6):2582–2594, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Chidami:2008:FDB

[CA08] Saad Chidami and Marc Amyot. Fish decomposition in boreal lakes and biogeochemical implications. Limnology and Oceanography, 53(5):1988–1996, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Caffrey:2001:BRV

[Caf01] Jane M. Caffrey. Book review: Valigura, R. A., R. B. Alexander, M. S. Castro, T. P. Meyers, H. W. Paerl, P. E. Stacey, and R. E. Turner [Eds.]. 2001. Nitrogen loading in coastal water bodies: an atmospheric perspective. American Geophysical Union. viii + 254 p. US$56. ISBN 0-87590-271-5. Limnology and Oceanography, 46(3):1584, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Calbet:2001:MGE

[Cal01] Albert Calbet. Mesozooplankton grazing effect on primary production: a global comparative analysis in marine ecosystems. Limnology and Oceanography, 46(7):1824–1830, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Canfield:2000:MG

[Can00] Donald E. Canfield. Marine geochemistry. Limnology and Oceanography, 45(7):1680, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Cossa:2009:OMO

[CAP09] Daniel Cossa, Bernard Averty, and Nicola Pirrone. The origin of methylmercury in open Mediterranean waters. Limnology and Oceanography, 54(3):837–844, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Cary:2001:RMB

[Car01] S. Craig Cary. Response of marine bacterial community composition to iron additions in three iron-limited regimes. Limnology and Oceanography, 46(6):1535–1545, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Carrington:2002:SVA

[Car02] Emily Carrington. Seasonal variation in the attachment strength of blue mussels: Causes and consequences. Limnology and Oceanography, 47(6):1723–1733, November 2002. CODEN LI-OCAH. ISSN 0024-3590.

Charette:2004:SGD

[CB04] Matthew A. Charette and Ken O. Buesseler. Submarine groundwater discharge of nutrients and copper to an urban subestuary of Chesapeake Bay (Elizabeth River). Limnology and Oceanography, 49(2):376–385, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Charette:2001:URI

[CBA01] Matthew A. Charette, Ken O. Buesseler, and John E. Andrews. Utility of radium isotopes for evaluating the input and transport of groundwater-derived nitrogen to a Cape Cod estuary. Limnology and Oceanography, 49(1):465–470, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Claustre:2002:DVP

[CBB^{+}02] Hervé Claustre, Annick Bricaud, Marcel Babin, Flavienne Bruyant, Laure Guillou, Florence Le Gall, Dominique Marie, and Frédéric Partensky. Diel variations in Prochlorococcus optical properties. Limnology and Oceanography, 47(6):1637–1647, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Childers:2006:RPW

[CBD^{+}06] Daniel L. Childers, Joseph N. Boyer, Stephen E. Davis, Christopher J. Madden, David T. Rudnick, and Fred H. Sklar. Relating precipitation and water management to nutrient concentrations in the oligotrophic “upside-down” estuaries of the Florida Everglades. Limnology and Oceanography, 51(1 (part 2)):602–616, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Chen:2001:AES

[CBG01] Yu-Wei Chen, Nelson Belzile, and John M. Gunn. Antagonistic effect of selenium on mercury assimilation by fish populations near Sudbury metal smelters? Limnology and Oceanography, 46 (7):1814–1818, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Colbert:2008:RBCb

[CBH08] S. L. Colbert, W. M. Berelson, and D. E. Hammond. Radon-222 budget in Catalina harbor, California: 2. Flow dynamics and residence time in a tidal beach. Limnology and Oceanography, 53 (2):659–665, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Church:2008:RDN

[CBK^{+}08] Matthew J. Church, Karin M. Björkman, David M. Karl, Mak A. Saito, and Jonathan P. Zehr. Regional distributions of nitrogen-fixing bacteria in the Pacific Ocean. Limnology and Oceanography, 53(1):63–77, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Cavender-Bares:2001:MSS

[CBRC01] Kent K. Cavender-Bares, Andrea Rinaldo, and Sallie W. Chisholm. Microbial size spectra from natural and nutrient enriched ecosystems. Limnology and Oceanography, 46(4):778–789, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Croot:2003:UEM

[CBvEK03] Peter L. Croot, Karlson Bengt, J. T. van Elteren, and J. J. Kroon. Uptake and efflux of^{64}Cu by the marine cyanobacterium Synechococcus (WH7803). Limnology and Oceanography, 48(1): 179–188, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Carlsson:2001:SVP

[CC01] Per Carlsson and David A. Caron. Seasonal variation of phosphorus limitation of bacterial growth in a small lake. Limnology and Oceanography, 46(1):108–120, January 2001. CODEN LI-OCAH. ISSN 0024-3590.

Costello:2002:PRU

[CC02] J. H. Costello and S. P. Colin. Prey resource use by coexistent hydromedusae from Friday Harbor, Washington. Limnology and Oceanography, 47(4):934–942, July 2002. CODEN LI-OCAH. ISSN 0024-3590.

Carstensen:2004:FCC

[CC04] Jacob Carstensen and Daniel J. Conley. Frequency, composition, and causes of summer phytoplankton blooms in a shallow coastal ecosystem, the Kattegat. Limnology and Oceanography, 49(1): 191–201, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Carstensen:2006:CET

[CCAÆ06] Jacob Carstensen, Daniel J. Conley, Jesper H. Andersen, and Gunni Ærtebjerg. Coastal eutrophication and trend reversal: a Danish case study. Limnology and Oceanography, 51(1 (part 2)): 398–408, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Crossman:2001:DFG

[CCC^{+}01] David J. Crossman, Howard J. Choat, Kendall D. Clements, Tom Hardy, and Jason McConochie. Detritus as food for grazing fishes on coral reefs. Limnology and Oceanography, 46(7):1596– 1605, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Cullen:2003:EIL

[CCC^{+}03] Jay T. Cullen, Zanna Chase, Kenneth H. Coale, Steve E. Fitzwater, and Robert M. Sherrell. Effect of iron limitation on the cadmium to phosphorus ratio of natural phytoplankton assemblages from the Southern Ocean. Limnology and Oceanography, 48(3): 1079–1087, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Conrad:2009:CSI

[CCC09] Ralf Conrad, Peter Claus, and Peter Casper. Characterization of stable isotope fractionation during methane production in the sediment of a eutrophic lake, Lake Dagow, Germany. Limnology and Oceanography, 54(2):457–471, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Conrad:2007:CMA

[CCCC07] Ralf Conrad, On-Chim Chan, Peter Claus, and Peter Casper. Characterization of methanogenic Archaea and stable isotope fractionation during methane production in the profundal sediment of an oligotrophic lake (Lake Stechlin, Germany). Limnology and Oceanography, 52(4):1393–1406, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Cutter:2004:SCC

[CCF04] Gregory A. Cutter, Lynda S. Cutter, and Katherine C. Filippino. Sources and cycling of carbonyl sulfide in the Sargasso Sea. Limnology and Oceanography, 49(2):555–565, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Colin:2005:OSC

[CCGH05] Sean P. Colin, John H. Costello, William M. Graham, and John Higgins III. Omnivory by the small cosmopolitan hydromedusa Aglaura hemistoma. Limnology and Oceanography, 50(4):1264– 1268, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Cloern:2002:SCN

[CCH02] James E. Cloern, Elizabeth A. Canuel, and David Harris. Stable carbon and nitrogen isotope composition of aquatic and terrestrial plants of the San Francisco Bay estuarine system. Limnology and Oceanography, 47(3):713–729, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Chasar:2005:EEE

[CCKC05] Lia C. Chasar, Jeffrey P. Chanton, Christopher C. Koenig, and Felicia C. Coleman. Evaluating the effect of environmental disturbance on the trophic structure of Florida Bay, U.S.A.: Multiple stable isotope analyses of contemporary and historical specimens. Limnology and Oceanography, 50(4):1059–1072, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Cole:2002:POC

[CCKP02] Jonathan J. Cole, Stephen R. Carpenter, James F. Kitchell, and Michael L. Pace. Pathways of organic carbon utilization in small lakes: Results from a whole-lake ^{13}C addition and coupled model. Limnology and Oceanography, 47(6):1664–1675, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Connan:2007:LSO

[CCM07] Maëlle Connan, Yves Cherel, and Patrick Mayzaud. Lipids from stomach oil of procellariiform seabirds document the importance of myctophid fish in the Southern Ocean. Limnology and Oceanography, 52(6):2445–2455, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Caraco:2006:TCB

[CCS06] N. F. Caraco, J. J. Cole, and D. L. Strayer. Top down control from the bottom: Regulation of eutrophication in a large river by benthic grazing. Limnology and Oceanography, 51(1 (part 2)): 664–670, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Colin:2002:TTE

[CD02] Sean P. Colin and Hans G. Dam. Testing for toxic effects of prey on zooplankton using sole versus mixed diets. Limnology and Oceanography, 47(5):1430–1437, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Cottrell:2003:CMB

[CD03] Matthew T. Cottrell and Kirchman L. David. Contribution of major bacterial groups to bacterial biomass production (thymidine and leucine incorporation) in the Delaware estuary. Limnology and Oceanography, 48(1):168–178, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Corbett:2000:EGC

[CDBC00] D. Reide Corbett, Kevin Dillon, William Burnett, and Jeffrey Chanton. Estimating the groundwater contribution into Florida Bay via natural tracers, ^{222}Rn and CH_{4}. Limnology and Oceanography, 45(7):1546–1557, November 2000. CODEN LI-OCAH. ISSN 0024-3590.

Church:2003:ADP

[CDD^{+}03] Matthew J. Church, Edward F. DeLong, Hugh W. Ducklow, Markus B. Karner, Christina M. Preston, and David M. Karl.

Abundance and distribution of planktonic Archaea and Bacteria in the waters west of the Antarctic Peninsula. Limnology and Oceanography, 49(4):1893–1902, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Corinaldesi:2007:EDT

[CDD07a] Cinzia Corinaldesi, Antonio Dell’Anno, and Roberto Danovaro. Early diagenesis and trophic role of extracellular DNA in different benthic ecosystems. Limnology and Oceanography, 52(4): 1710–1717, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Corinaldesi:2007:VIP

[CDD07b] Cinzia Corinaldesi, Antonio Dell’Anno, and Roberto Danovaro. Viral infection plays a key role in extracellular DNA dynamics in marine anoxic systems. Limnology and Oceanography, 52(2): 508–516, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Checkley:2008:APO

[CDH^{+}08] D. M. Checkley,, Jr., R. E. Davis, A. W. Herman, G. A. Jackson, B. Beanlands, and L. A. Regier. Assessing plankton and other particles in situ with the SOLOPC. Limnology and Oceanography, 53(5 (part 2)):2123–2136, September 2008. CODEN LI-OCAH. ISSN 0024-3590.

Church:2002:MID

[CDK02] Matthew J. Church, Hugh W. Ducklow, and David M. Karl. Multiyear increases in dissolved organic matter inventories at Station ALOHA in the North Pacific Subtropical Gyre. Limnology and Oceanography, 47(1):1–10, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Chang:2009:FIC

[CEA^{+}09] Sandy Chang, Chris Elkins, Marcus Alley, John Eaton, and Stephen Monismith. Flow inside a coral colony measured using magnetic resonance velocimetry. Limnology and Oceanography, 54(5):1819–1827, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Cebrian:2002:VCC

[Ceb02] Just Cebrian. Variability and control of carbon consumption, export, and accumulation in marine communities. Limnology and Oceanography, 47(1):11–22, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Cermeno:2005:MPE

[CEBMF05] Pedro Cermeño, Patricia Estévez-Blanco, Emilio Marañón, and Emilio Fernóndez. Maximum photosynthetic efficiency of sizefractionated phytoplankton assessed by ^{14}C uptake and fast repetition rate fluorometry. Limnology and Oceanography, 50(5): 1438–1446, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Chadwick:2001:IST

[CF01] Michael A. Chadwick and Jack W. Feminella. Influence of salinity and temperature on the growth and production of a freshwater mayfly in the Lower Mobile River, Alabama. Limnology and Oceanography, 46(3):532–542, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Cohen:2005:PID

[CF05] Jonathan H. Cohen and Richard B,. Jr. Forward. Photobehavior as an inducible defense in the marine copepod Calanopia americana. Limnology and Oceanography, 50(4):1269–1277, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Crespo:2007:RAS

[CFG07] B. G. Crespo, F. G. Figueiras, and S. Groom. Role of across-shelf currents in the dynamics of harmful dinoflagellate blooms in the northwestern Iberian upwelling. Limnology and Oceanography, 52(6):2668–2678, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Chambers:2001:BEI

[CFMH01] Randolph M. Chambers, James W. Fourqurean, Stephen A. Macko, and Regis Hoppenot. Biogeochemical effects of iron availability on primary producers in a shallow marine carbonate environment. Limnology and Oceanography, 46(6):1278–1286, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Claustre:2000:CMS

[CFO^{+}00] Hervé Claustre, Frank Fell, Kadija Oubelkheir, Louis Prieur, Antoine Sciandra, Bernard Gentili, and Marcel Babin. Continuous monitoring of surface optical properties across a geostrophic front: Biogeochemical inferences. Limnology and Oceanography, 45(2):309–321, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Chrzanowski:2001:EMN

[CG01] Thomas H. Chrzanowski and James P. Grover. Effects of mineral nutrients on the growth of bacterio- and phytoplankton in two southern reservoirs. Limnology and Oceanography, 46(6):1319– 1330, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Choe:2003:DPCb

[CG03] Key-Young Choe and Gary A. Gill. Distribution of particulate, colloidal, and dissolved mercury in San Francisco Bay estuary.

2. Monomethyl mercury. Limnology and Oceanography, 48(4): 1547–1556, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Chrzanowski:2008:ECP

[CG08] Thomas H. Chrzanowski and James P. Grover. Element content of Pseudomonas fluorescens varies with growth rate and temperature: a replicated chemostat study addressing ecological stoichiometry. Limnology and Oceanography, 53(4):1242–1251, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Cai:2009:AVC

[CG09] Yihua Cai and Laodong Guo. Abundance and variation of colloidal organic phosphorus in riverine, estuarine, and coastal waters in the northern Gulf of Mexico. Limnology and Oceanography, 54(4):1393–1402, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Casamayor:2008:CDF

[CGCPA08] Emilio O. Casamayor, Josefina García-Cantizano, and Carlos Pedrós-Alió. Carbon dioxide fixation in the dark by photosynthetic bacteria in sulfide-rich stratified lakes with oxic–anoxic interfaces. Limnology and Oceanography, 53(4):1193–1203, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Cai:2008:TVO

[CGD08] Yihua Cai, Laodong Guo, and Thomas A. Douglas. Temporal variations in organic carbon species and fluxes from the Chena River, Alaska. Limnology and Oceanography, 53(4):1408–1419, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Carlson:2004:IAD

[CGH^{+}04] Craig A. Carlson, Stephen J. Giovannoni, Dennis A. Hansell, Stuart J. Goldberg, Rachel Parsons, and Kevin Vergin. Interactions among dissolved organic carbon, microbial processes, and community structure in the mesopelagic zone of the northwestern Sargasso Sea. Limnology and Oceanography, 49(4):1073– 1083, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Choe:2003:DPCa

[CGL03] Key-Young Choe, Gary A. Gill, and Ronald Lehman. Distribution of particulate, colloidal, and dissolved mercury in San Francisco Bay estuary. 1. Total mercury. Limnology and Oceanography, 48(4):1535–1546, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Choe:2004:SWE

[CGL^{+}04] Key-Young Choe, Gary A. Gill, Ronald D. Lehman, Seunghee Han, Wesley A. Heim, and Kenneth H. Coale. Sediment-water exchange of total mercury and monomethyl mercury in the San Francisco Bay-Delta. Limnology and Oceanography, 49(5):1512– 1527, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Colon-Gaud:2009:AER

[CGWK^{+}09] Checo Colón-Gaud, Matt R. Whiles, Susan S. Kilham, Karen R. Lips, Cathy M. Pringle, Scott Connelly, and Scot D. Peterson. Assessing ecological responses to catastrophic amphibian declines: Patterns of macroinvertebrate production and food web structure in upland Panamanian streams. Limnology and Oceanography, 54(1):331–343, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Clark:2001:QFH

[CH01] Jordan F. Clark and G. Bryant Hudson. Quantifying the flux of hydrothermal fluids into Mono Lake by use of helium isotopes. Limnology and Oceanography, 46(1):189–196, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Chadwick:2005:RSM

[CH05a] Michael A. Chadwick and Alexander D. Huryn. Response of stream macroinvertebrate production to atmospheric nitrogen deposition and channel drying. Limnology and Oceanography, 55(3):228–236, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Crump:2005:SSB

[CH05b] Byron C. Crump and John E. Hobbie. Synchrony and seasonality in bacterioplankton communities of two temperate rivers.

Limnology and Oceanography, 50(6):1718–1729, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Charette:2007:HFS

[Cha07] Matthew A. Charette. Hydrologic forcing of submarine groundwater discharge: Insight from a seasonal study of radium isotopes in a groundwater-dominated salt marsh estuary. Limnology and Oceanography, 52(1):230–239, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Colbert:2008:RBCa

[CHB08] S. L. Colbert, D. E. Hammond, and W. M. Berelson. Radon-222 budget in Catalina harbor, California: 1. Water mixing rates. Limnology and Oceanography, 53(2):651–658, March 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Carstensen:2007:SAB

[CHH07] Jacob Carstensen, Peter Henriksen, and Anna-Stiina Heiskanen. Summer algal blooms in shallow estuaries: Definition, mechanisms, and link to eutrophication. Limnology and Oceanography, 52(1):370–384, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Chisholm:2000:CCC

[Chi00] John R. M. Chisholm. Calcification by crustose coralline algae on the northern Great Barrier Reef, Australia. Limnology and Oceanography, 45(7):1476–1484, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Christian:2005:BCO

[Chr05] James Robert Christian. Biogeochemical cycling in the oligotrophic ocean: Redfield and non-Redfield models. Limnology and Oceanography, 50(4):646–657, March 2005. CODEN LI-OCAH. ISSN 0024-3590.

Croteau:2002:ITC

[CHT02] Marie-Noële Croteau, Landis Hare, and André Tessier. Influence of temperature on Cd accumulation by species of the biomonitor Chaoborus. Limnology and Oceanography, 47(2):505–514, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Cristescu:2001:IHC

[CHW^{+}01] Melania E. A. Cristescu, Paul D. N. Hebert, Jonathan D. S. Witt, Hugh J. MacIsaac, and Igor A. Grigorovich. An invasion history for Cercopagis pengoi based on mitochondrial gene sequences. Limnology and Oceanography, 46(2):224–229, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Costelloe:2009:GRD

[CIWH09] Justin F. Costelloe, Elizabeth C. Irvine, Andrew W. Western, and Andrew L. Herczeg. Groundwater recharge and discharge dynamics in an arid-zone ephemeral lake system, Australia. Limnology and Oceanography, 54(1):86–100, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Carini:2008:NML

[CJ08] Stephen A. Carini and Samantha B. Joye. Nitrification in Mono Lake, California: Activity and community composition during contrasting hydrological regimes. Limnology and Oceanography, 53(6):2546–2557, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Cotner:2000:IWH

[CJB00] James B. Cotner, Thomas H. Johengen, and Bopaiah A. Biddanda. Intense winter heterotrophic production stimulated by benthic resuspension. Limnology and Oceanography, 45(7):1672– 1676, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Cross:2005:CRS

[CJWR05] W. F. Cross, B. R. Johnson, J. B. Wallace, and A. D. Rosemond. Contrasting response of stream detritivores to long-term nutrient enrichment. Limnology and Oceanography, 50(6):1730–1739, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Christoffersen:2000:TCW

[CK00] Kirsten Christoffersen and Hanne Kaas. Toxic cyanobacteria in water. A guide to their public health consequences, monitoring, and management. Limnology and Oceanography, 45(5): 1212, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Choi:2008:MLE

[CK08] Keun-Hyung Choi and Wim J. Kimmerer. Mate limitation in an estuarine population of copepods. Limnology and Oceanography, 53(5):1656–1664, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Castillo:2003:BCB

[CKA03] María M. Castillo, George W. Kling, and J. David Allan. Bottom-up controls on bacterial production in tropical lowland rivers. Limnology and Oceanography, 48(4):1466–1475, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Cammack:2004:FDO

[CKPS04] W. K. Levi Cammack, Jacob Kalff, Yves T. Prairie, and Erik M. Smith. Fluorescent dissolved organic matter in lakes: Relationships with heterotrophic metabolism. Limnology and Oceanography, 49(6):2034–2045, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Chanton:2002:ECB

[CL02] Jeffrey Chanton and F. Graham Lewis. Examination of coupling between primary and secondary production in a riverdominated estuary: Apalachicola Bay, Florida, U.S.A. Limnology and Oceanography, 47(3):683–697, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Calbet:2004:PGM

[CL04] Albert Calbet and Michael R. Landry. Phytoplankton growth, microzooplankton grazing, and carbon cycling in marine systems. Limnology and Oceanography, 49(1):51–57, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Corno:2008:TVV

[CLAK08] Guido Corno, Ricardo M. Letelier, Mark R. Abbott, and David M. Karl. Temporal and vertical variability in photosynthesis in the North Pacific Subtropical Gyre. Limnology and Oceanography, 53(4):1252–1265, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Ciotti:2002:ARB

[CLC02] Áurea M. Ciotti, Marlon R. Lewis, and John J. Cullen. Assessment of the relationships between dominant cell size in natural phytoplankton communities and the spectral shape of the absorption coefficient. Limnology and Oceanography, 47(2):404– 417, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Chao:2008:SAA

[CLF^{+}08] Yi Chao, Zhijin Li, John D. Farrara, Mark A. Moline, Oscar M. E. Schofield, and Sharanya J. Majumdar. Synergistic applications of autonomous underwater vehicles and the regional ocean modeling system in coastal ocean forecasting. Limnology and Oceanography, 53(5 (part 2)):2251–2263, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Clark:2008:SSC

[CLG08] Catherine D. Clark, Liannea P. Litz, and Stanley B. Grant. Saltmarshes as a source of chromophoric dissolved organic matter (CDOM) to Southern California coastal waters. Limnology and Oceanography, 53(5):1923–1933, September 2008. CODEN LI-OCAH. ISSN 0024-3590.

Carder:2003:ITE

[CLL^{+}03] Kendall L. Carder, Cheng-Chien Liu, Zhongping Lee, David C. English, James Patten, F. Robert Chen, James E. Ivey, and Curtiss O. Davis. Illumination and turbidity effects on observing faceted bottom elements with uniform Lambertian albedos. Limnology and Oceanography, 48(1 (part 2)):355–363, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Chen:2009:CCB

[CLL^{+}09] Bingzhang Chen, Hongbin Liu, Michael R. Landry, Minhan Dai, Bangqin Huang, and Jun Sun. Close coupling between phytoplankton growth and microzooplankton grazing in the western South China Sea. Limnology and Oceanography, 54(4):1084– 1097, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Crawford:2003:IZI

[CLP^{+}03] D. W. Crawford, M. S. Lipsen, D. A. Purdie, M. C. Lohan, P. J. Statham, F. A. Whitney, J. N. Putland, W. K. Johnson, N. Sutherland, T. D. Peterson, P. J. Harrison, and C. S. Wong. Influence of zinc and iron enrichments on phytoplankton growth in the northeastern subarctic Pacific. Limnology and Oceanography, 48(4):1583–1600, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Cassar:2002:SIC

[CLPB02] Nicolas Cassar, Edward A. Laws, Brian N. Popp, and Robert R. Bidigare. Sources of inorganic carbon for photosynthesis in a strain of Phaeodactylum tricornutum. Limnology and Oceanography, 47(4):1192–1197, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Croteau:2005:TTM

[CLS05] Marie-Noële Croteau, Samuel N. Luoma, and A. Robin Stewart. Trophic transfer of metals along freshwater food webs: Evidence of cadmium biomagnification in nature. Limnology and Oceanography, 50(5):1511–1519, September 2005. CODEN LI-OCAH. ISSN 0024-3590.

Chisholm:2003:SNF

[CM03] John R. M. Chisholm and Pascale Moulin. Stimulation of nitrogen fixation in refractory organic sediments by Caulerpa taxifolia (Chlorophyta). Limnology and Oceanography, 48(2):787–794, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Croot:2000:PEC

[CMB00] Peter L. Croot, James W. Moffett, and Larry E. Brand. Production of extracellular Cu complexing ligands by eucaryotic phytoplankton in response to Cu stress. Limnology and Oceanography, 45(3):619–627, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Corno:2009:BDM

[CMC^{+}09] Gianluca Corno, Beatriz E. Modenutti, Cristiana Callieri, Esteban G. Balseiro, Roberto Bertoni, and Emanuele Caravatia. Bacterial diversity and morphology in deep ultraoligotrophic Andean lakes: The role of UVR on vertical distribution. Limnology and Oceanography, 54(4):1098–1112, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Crepo-Medina:2009:ACM

[CMCB^{+}09] Melitza Crepo-Medina, Aspassia D. Chatziefthimiou, Nicolas S. Bloom, George W. Luther III, Derek D. Wright, John R. Reinfelder, Costantino Vetriani, and Tamar Barkay. Adaptation of chemosynthetic microorganisms to elevated mercury concentrations in deep-sea hydrothermal vents. Limnology and Oceanography, 54(1):41–49, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Carrington:2009:SDM

[CMD^{+}09] Emily Carrington, Gretchen M. Moeser, James Dimond, Joseph J. Mello, and Michael L. Boller. Seasonal disturbance to mussel beds: Field test of a mechanistic model predicting wave dislodgment. Limnology and Oceanography, 54(3):978–986, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Cermeno:2008:RLA

[CMH^{+}08] Pedro Cermeño, Emilio Marañín, Derek Harbour, Francisco G. Figueiras, Bibiana G. Crespo, María Huete-Ortega, Manuel Varela, and Roger P. Harris. Resource levels, allometric scaling of population abundance, and marine phytoplankton diversity. Limnology and Oceanography, 53(1):312–318, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Clegg:2007:BRP

[CMJ07] Mark R. Clegg, Stephen C. Maberly, and Roger I. Jones. Behavioral response as a predictor of seasonal depth distribution and vertical niche separation in freshwater phytoplanktonic flagellates. Limnology and Oceanography, 52(1):441–455, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Chen:2000:HCC

[CMQ^{+}00] Zhen Chen, Lawrence M. Mayer, Christophe Quétel, Olivier F. X. Donard, Robert F. L. Self, Peter A. Jumars, and Donald P. Weston. High concentrations of complexed metals in the guts of deposit feeders. Limnology and Oceanography, 45(6):1358–1367, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Carrillo:2002:IPB

[CMSVA02] Presentación Carrillo, Juan Manuel Medina-Sánchez, and Manuel Villar-Argaiz. The interaction of phytoplankton and bacteria in a high mountain lake: Importance of the spectral composition of solar radiation. Limnology and Oceanography, 47(5):1294–1306, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Caston:2009:RIH

[CNGV09] C. Brad Caston, Weston H. Nowlin, Alicia Gaulke, and Michael J. Vanni. The relative importance of heterotrophic bacteria to pelagic ecosystem dynamics varies with reservoir trophic state. Limnology and Oceanography, 54(6):2143–2156, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Cochlan:2001:HBR

[Coc01] William P. Cochlan. The heterotrophic bacterial response during a mesoscale iron enrichment experiment (IronEx II) in the eastern equatorial Pacific Ocean. Limnology and Oceanography, 49(1):428–435, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Crowe:2008:BTL

[COK^{+}08] Sean A. Crowe, Andrew H. O’Neill, Sergei Katsev, Peter Hehanussa, G. Douglas Haffner, Bjørn Sundby, Alfonso Mucci, and David A. Fowle. The biogeochemistry of tropical lakes: a case study from Lake Matano, Indonesia. Limnology and Oceanography, 53(1):319–331, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Coma:2002:SSR

[Com02] Rafel Coma. Seasonality of in situ respiration rate in three temperate benthic suspension feeders. Limnology and Oceanography, 47(1):324–331, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Calbet:2003:MGP

[CP03] Albert Calbet and Yves T. Prairie. Mesozooplankton grazing and primary production: Reply to the comment by Laws. Limnology and Oceanography, 48(4):1359–1362, May 2003. CODEN LIOCAH. ISSN 0024-3590. See [Law03].

Ciancio:2008:TRE

[CPB^{+}08] Javier E. Ciancio, Miguel A. Pascual, Florencia Botto, Esteban Frere, and Oscar Iribarne. Trophic relationships of exotic anadromous salmonids in the southern Patagonian Shelf as inferred from stable isotopes. Limnology and Oceanography, 53(2): 788–798, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Cole:2000:PNH

[CPCK00] Jonathan J. Cole, Michael L. Pace, Stephen R. Carpenter, and James F. Kitchell. Persistence of net heterotrophy in lakes during nutrient addition and food web manipulations. Limnology and Oceanography, 45(8):1718–1730, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Ciros-Perez:2004:PFM

[CPCLS04] Jorge Ciros-Pérez, María José Carmona, Sara Lapesa, and Manuel Serra. Predation as a factor mediating resource competition among rotifer sibling species. Limnology and Oceanography, 49(1):40–50, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Ciros-Perez:2001:RCB

[CPCS01] Jorge Ciros-Pérez, María José Carmona, and Manuel Serra. Resource competition between sympatric sibling rotifer species. Limnology and Oceanography, 46(6):1511–1523, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Catalan:2009:CCS

[CPGC09] Jordi Catalan, Sergi Pla, Joan García, and Lluís Camarero. Climate and CO_{2} saturation in an alpine lake throughout the Holocene. Limnology and Oceanography, 54(6 (part 2)):2542– 2552, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Chetelat:2006:PSS

[CPH06] John Chételat, Frances R. Pick, and Paul B. Hamilton. Potamoplankton size structure and taxonomic composition: Influence of river size and nutrient concentrations. Limnology and Oceanography, 51(1 (part 2)):681–689, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Chan:2004:BFH

[CPHM04] Francis Chan, Michael L. Pace, Robert W. Howarth, and Roxanne M. Marino. Bloom formation in heterocystic nitrogenfixing cyanobacteria: The dependence on colony size and zooplankton grazing. Limnology and Oceanography, 49(6):2171– 2178, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Chung:2004:PEA

[CPL^{+}04] Sook-Nye Chung, Geun-Ha Park, Kitack Lee, Robert M. Key, Frank J. Millero, Richard A. Feely, Christopher L. Sabine, and Paul G. Falkowski. Postindustrial enhancement of aragonite undersaturation in the upper tropical and subtropical Atlantic Ocean: The role of fossil fuel CO_{2}. Limnology and Oceanography, 49(2):315–321, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Carignan:2000:PPR

[CPV00] Richard Carignan, Dolors Planas, and Chantal Vis. Planktonic production and respiration in oligotrophic Shield lakes. Limnology and Oceanography, 45(1):189–199, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Cook:2006:ARC

[CR06] Perran L. M. Cook and Hans Røy. Advective relief of CO_{2} limitation in microphytobenthos in highly productive sandy sediments. Limnology and Oceanography, 51(4):1594–1601, July 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Caron:2008:MTE

[CR08] David A. Caron and Julie M. Rose. The metabolic theory of ecology and algal bloom formation (reply to comment by López–Urrutia). Limnology and Oceanography, 53(5):2048–2049, September 2008. CODEN LIOCAH. ISSN 0024-3590. See [LU08].

Cook:2010:ECS

[CR10] Freeman Cook and Alon Rimmer. Erratum: Chemical stratification in thermally stratified lakes: a chloride mass balance model. Limnology and Oceanography, 55(3):1463–1465, May 2010. CO-DEN LIOCAH. ISSN 0024-3590. See [RAKE05].

Craft:2007:FIS

[Cra07] Christopher Craft. Freshwater input structures soil properties, vertical accretion, and nutrient accumulation of Georgia and U.S tidal marshes. Limnology and Oceanography, 52(3):1220–1230, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Christner:2006:LCS

[CRBF^{+}06] Brent C. Christner, George Royston-Bishop, Christine M. Foreman, Brianna R. Arnold, Martyn Tranter, Kathleen A. Welch, W. Berry Lyons, Alexandre I. Tsapin, Michael Studinger, and John C. Priscu. Limnological conditions in Subglacial Lake Vostok, Antarctica. Limnology and Oceanography, 51(6):2485–2501, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Churchill:2003:ONB

[CRC^{+}03] James H. Churchill, Elise A. Ralph, Angela M. Cates, Judith W. Budd, and Noel R. Urban. Observations of a negatively buoyant river plume in a large lake. Limnology and Oceanography, 48(2): 884–894, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Cullen:2001:EAC

[CRF01] Jay T. Cullen, Yair Rosenthal, and Paul G. Falkowski. The effect of anthropogenic CO_{2} on the carbon isotope composition of marine phytoplankton. Limnology and Oceanography, 44(4): 996–998, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Clark:2008:ARN

[CRJ08] Darren R. Clark, Andrew P. Rees, and Ian Joint. Ammonium regeneration and nitrification rates in the oligotrophic Atlantic Ocean: Implications for new production estimates. Limnology and Oceanography, 53(1):52–62, January 2008. CODEN LI-OCAH. ISSN 0024-3590.

Croot:2003:SCC

[Cro03] Peter L. Croot. Seasonal cycle of copper speciation in gullmar fjord, Sweden. Limnology and Oceanography, 48(2):764– 776, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Craft:2008:IUI

[CRY^{+}08] Jonathan D. Craft, Atlantis D. Russ, Mike N. Yamamoto, Thomas Y. Iwai, Jr., Skippy Hau, John Kahiapo, Charlie T. Chong, Sharon Ziegler-Chong, Cam Muir, Yoshihisa Fujita, Dan A. Polhemus, Robert A. Kinzie III, and Scott R. Santos. Islands under islands: The phylogeography and evolution of Halocaridina rubra Holthuis, 1963 (Crustacean: Decapoda: Atyidae) in the Hawaiian archipelago. Limnology and Oceanography, 53 (2):675–689, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Cullen:2005:EDC

[CS05] Jay T. Cullen and Robert M. Sherrell. Effects of dissolved carbon dioxide, zinc, and manganese on the cadmium to phosphorus ratio in natural phytoplankton assemblages. Limnology and Oceanography, 50(4):1193–1204, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Chrost:2006:MPU

[CS06] Ryszard J. Chróst and Waldemar Siuda. Microbial production, utilization, and enzymatic degradation of organic matter in the upper trophogenic layer in the pelagial zone of lakes along a eutrophication gradient. Limnology and Oceanography, 51(1 (part 2)):749–762, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Castruita:2008:AII

[CSE^{+}08] Madeli Castruita, Yeala Shaked, Lauren A. Elmegreen, Edward I. Stiefel, and François M. M. Morel. Availability of iron from iron-storage proteins to marine phytoplankton. Limnology and Oceanography, 53(3):890–899, May 2008. CODEN LI-OCAH. ISSN 0024-3590.

Costello:2006:SRS

[CSG^{+}06] J. H. Costello, B. K. Sullivan, D. J. Gifford, D. Van Keuren, and L. J. Sullivan. Seasonal refugia, shoreward thermal amplification, and metapopulation dynamics of the ctenophore Mnemiopsis leidyi in Narragansett Bay, Rhode Island. Limnology and Oceanography, 51(5):1819–1831, July 2006. CODEN LI-OCAH. ISSN 0024-3590. See erratum [Ano08b].

Chen:2000:AHM

[CSK^{+}00] Celia Y. Chen, Richard S. Stemberger, Bjorn Klaue, Joel D. Blum, Paul C. Pickhardt, and Carol L. Folt. Accumulation of heavy metals in food web components across a gradient of lakes. Limnology and Oceanography, 45(7):1525–1536, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Conan:2007:POP

[CSK^{+}07] Pascal Conan, Morten Søndergaard, Theis Kragh, Frede Thingstad, Mireille Pujo-Pay, Peter J. le B. Williams, Stiig Markager, Gustave Cauwet, Niels Henrik Borch, Dylan Evans, and Bo Riemann. Partitioning of organic production in marine plankton communities: The effects of inorganic nutrient ratios and community composition on new dissolved organic matter. Limnology and Oceanography, 52(2):753–765, March 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Caron:2008:MFS

[CSM^{+}08] David A. Caron, Beth Stauffer, Stefanie Moorthi, Amarjeet Singh, Maxim Batalin, Eric A. Graham, Mark Hansen, William J. Kaiser, Jnaneshwar Das, Arvind Pereira, Amit Dhariwal, Bin Zhang, Carl Oberg, and Gaurav S. Sukhatme. Macroto fine-scale spatial and temporal distributions and dynamics of phytoplankton and their environmental driving forces in a small montane lake in southern California, USA. Limnology and Oceanography, 53(5 (part 2)):2333–2349, September 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Colomer:2001:OHP

[CSP^{+}01] Jordi Colomer, Teresa Serra, Jaume Piera, Elena Roget, and Xavier Casamitjana. Observations of a hydrothermal plume in a karstic lake. Limnology and Oceanography, 46(1):197–203, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Conley:2000:TRD

[CSPW00] Daniel J. Conley, Per Stålnacke, Heikki Pitkänen, and Anders Wilander. The transport and retention of dissolved silicate by rivers in Sweden and Finland. Limnology and Oceanography, 45 (8):1850–1853, December 2000. CODEN LIOCAH. ISSN 0024-

Carrias:2002:DSB

[CSSNA02] Jean-François Carrias, Jean-Philippe Serre, Télesphore Sime-Ngando, and Christian Amblard. Distribution, size, and bacterial colonization of pico- and nano-detrital organic particles (DOP) in two lakes of different trophic status. Limnology and Oceanography, 47(4):1202–1209, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Canuel:2007:BFW

[CSWD07] Elizabeth A. Canuel, Amanda C. Spivak, Elizabeth J. Waterson, and J. Emmett Duffy. Biodiversity and food web structure influence short-term accumulation of sediment organic matter in an experimental seagrass system. Limnology and Oceanography, 52 (2):590–602, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Cornelisen:2002:AUS

[CT02] Christopher D. Cornelisen and Florence I. M. Thomas. Ammonium uptake by seagrass epiphytes: Isolation of the effects of water velocity using an isotope label. Limnology and Oceanography, 47(4):1223–1229, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Caceres:2004:SSV

[CT04] Carla E. Cáceres and Alan J. Tessier. To sink or swim: Variable diapause strategies among Daphnia species. Limnology and Oceanography, 49(4 (part 2)):1333–1340, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Carey:2005:RSF

[CT05] Elizabeth Carey and Martial Taillefert. The role of soluble Fe(III) in the cycling of iron and sulfur in coastal marine sediments. Limnology and Oceanography, 50(4):1129–1141, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Camacho:2006:AHF

[CT06] Frank A. Camacho and Robert W. Thacker. Amphipod herbivory on the freshwater cyanobacterium Lyngbya wollei: Chemical stimulants and morphological defenses. Limnology and Oceanography, 53(2):1870–1875, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Chauvaud:2003:CCG

[CTCT03] Laurent Chauvaud, Janet K. Thompson, James E. Cloern, and Gérard Thouzeau. Clams as CO_{2} generators: The Potamocorbula amurensis example in San Francisco Bay. Limnology and Oceanography, 48(6):2086–2092, November 2003. CODEN LI-OCAH. ISSN 0024-3590.

Crimaldi:2002:HLS

[CTR^{+}02] John P. Crimaldi, Janet K. Thompson, Johanna H. Rosman, Ryan J. Lowe, and Jeffrey R. Koseff. Hydrodynamics of larval settlement: The influence of turbulent stress events at potential recruitment sites. Limnology and Oceanography, 47(4): 1137–1151, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Cottier:2006:USV

[CTWFP06] Finlo R. Cottier, Geraint A. Tarling, Anette Wold, and Stig Falk-Petersen. Unsynchronised and synchronised vertical migration of zooplankton in a high Arctic fjord. Limnology and Oceanography, 51(6):2586–2599, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Connelly:2006:EPO

[CTY06] T. L. Connelly, C. M. Tilburg, and P. L. Yager. Evidence for psychrophiles outnumbering psychrotolerant marine bacteria in the springtime coastal Arctic. Limnology and Oceanography, 51 (2):1205–1210, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Cullen:2006:NRB

[Cul06] Jay T. Cullen. On the nonlinear relationship between dissolved cadmium and phosphate in the modern global ocean: Could chronic iron limitation of phytoplankton growth cause the kink? Limnology and Oceanography, 51(3):1369–1380, May 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Curtis:2000:MAS

[Cur00] Suttle Curtis. Molecular approaches to the study of the ocean. Limnology and Oceanography, 45(7):1677, November 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Carmichael:2005:CNB

[CV05] R. H. Carmichael and I. Valiela. Coupling of near-bottom seston and surface sediment composition: Changes with nutrient enrichment and implications for estuarine food supply and biogeochemical processing. Limnology and Oceanography, 50(1): 97–105, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Chandra:2005:ECE

[CVH^{+}05] Sudeep Chandra, M. Jake Vander Zanden, Alan C. Heyvaert, Bob C. Richards, Brant C. Allen, and Charles R. Goldman. The effects of cultural eutrophication on the coupling between pelagic primary producers and benthic consumers. Limnology and Oceanography, 50(5):1368–1376, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Chen:2001:BNC

[CW01] Min Chen and Wen-Xiong Wang. Bioavailability of natural colloid-bound iron to marine plankton: Influences of colloidal size and aging. Limnology and Oceanography, 46(8):1956–1967, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Culley:2002:ADC

[CW02] Alexander I. Culley and Nicholas A. Welschmeyer. The abundance, distribution, and correlation of viruses, phytoplankton, and prokaryotes along a Pacific Ocean transect. Limnology and Oceanography, 47(5):1508–1513, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Crusius:2003:GTV

[CW03] John Crusius and Rik Wanninkhof. Gas transfer velocities measured at low wind speed over a lake. Limnology and Oceanography, 48(3):1010–1017, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Clarke:2006:LTT

[CWC^{+}06] A. L. Clarke, K. Weckström, D. J. Conley, N. J. Anderson, F. Adser, E. Andrén, V. N. de Jonge, M. Ellegaard, S. Juggins, P. Kauppila, A. Korhola, N. Reuss, R. J. Telford, and S. Vaalgamaa. Long-term trends in eutrophication and nutrients in the coastal zone. Limnology and Oceanography, 51(1 (part 2)): 385–397, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Cornelisen:2007:PSU

[CWC^{+}07] Christopher D. Cornelisen, Stephen R. Wing, Kim L. Clark, M. Hamish Bowman, Russell D. Frew, and Catriona L. Hurd. Patterns in the δ ^{13}C and δ ^{15}N signature of Ulva pertusa: Interaction between physical gradients and nutrient source pools. Limnology and Oceanography, 52(2):820–832, March 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Cook:2007:BSE

[CWG^{+}07] Perran L. M. Cook, Frank Wenzhöfer, Ronnie N. Glud, Felix Janssen, and Markus Huettel. Benthic solute exchange and carbon mineralization in two shallow subtidal sandy sediments: Effect of advective pore-water exchange. Limnology and Oceanography, 52(5):1943–1963, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Castaneda:2009:ICC

[CWJ09] Isla S. Castañeda, Josef P. Werne, and Thomas C. Johnson. Influence of climate change on algal community structure and primary productivity of Lake Malawi (East Africa) from the Last Glacial Maximum to present. Limnology and Oceanography, 54(6 (part 2)):2431–2447, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Chen:2006:BSI

[CWLH06] Yi-Jing Cherry Chen, Shian-Chee Wu, Bing-Shiou Lee, and Chang-Che Hung. Behavior of storm-induced suspension interflow in subtropical Feitsui Reservoir, Taiwan. Limnology and Oceanography, 51(2):1125–1133, March 2006. CODEN LI-OCAH. ISSN 0024-3590.

Carotenuto:2005:LHR

[CWPL05] Ylenia Carotenuto, Thomas Wichard, Georg Pohnert, and Winfried Lampert. Life-history responses of Daphnia pulicaria to diets containing freshwater diatoms: Effects of nutritional quality versus polyunsaturated aldehydes. Limnology and Oceanography, 50(2):449–454, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Cai:2000:IMS

[CWWS00] Wei-Jun Cai, William J. Wiebe, Yongchen Wang, and Joan E. Sheldon. Intertidal marsh as a source of dissolved inorganic carbon and a sink of nitrate in the Satilla River-estuarine complex in the southeastern U.S. Limnology and Oceanography, 45(8): 1743–1752, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Chikaraishi:2005:CHI

[CYN05] Yoshito Chikaraishi, Yusuke Yamada, and Hiroshi Naraoka. Carbon and hydrogen isotopic compositions of sterols from riverine and marine sediments. Limnology and Oceanography, 50(6): 1763–1770, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Chadwick:2004:MMB

[CZRD^{+}04] D. B. Chadwick, A. Zirino, I. Rivera-Duarte, C. N. Katz, and A. C. Blake. Modeling the mass balance and fate of copper in San Diego Bay. Limnology and Oceanography, 49(2):355–366, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Dyhrman:2003:UAC

[DA03] Sonya T. Dyhrman and Donald M. Anderson. Urease activity in cultures and field populations of the toxic dinoflagellate Alexandrium. Limnology and Oceanography, 48(2):647–655, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Dupont:2005:ECC

[DA05] Christopher L. Dupont and Beth A. Ahner. Effects of copper, cadmium, and zinc on the production and exudation of thiols by Emiliania huxleyi. Limnology and Oceanography, 50(4):508–515, March 2005. CODEN LIOCAH. ISSN 0024-3590.

DAsaro:2008:DMB

[D’A08] Eric A. D’Asaro. A diapycnal mixing budget on the Oregon shelf. Limnology and Oceanography, 53(5 (part 2)):2137–2150, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Duarte:2001:EHS

[DAA^{+}01] Carlos M. Duarte, Susana Agustí, Javier Arístegui, Natalia González, and Ricardo Anadón. Evidence for a heterotrophic subtropical northeast Atlantic. Limnology and Oceanography, 49 (1):425–428, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Deborde:2007:DPT

[DAC^{+}07] Jonathan Deborde, Pierre Anschutz, Gwénaëlle Chaillou, Henri Etcheber, Marc-Vincent Commarieu, Pascal Lecroart, and Gwenaël Abril. The dynamics of phosphorus in turbid estuarine systems: Example of the Gironde estuary (France). Limnology and Oceanography, 52(2):862–872, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Dufour:2001:AMC

[DACG01] Philippe Dufour, Serge Andréfouët, Loic Charpy, and Nicole Garcia. Atoll morphometry controls lagoon nutrient regime. Limnology and Oceanography, 49(1):456–461, March 2001. CO-DEN LIOCAH. ISSN 0024-3590.

DAndrea:2002:OMF

[DAL02] Anthony F. D’Andrea, Robert C. Aller, and Glenn R. Lopez. Organic matter flux and reactivity on a South Carolina sandflat: The impacts of porewater advection and macrobiological structures. Limnology and Oceanography, 47(4):1056–1070, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Dalsgaard:2003:BPP

[Dal03] Tage Dalsgaard. Benthic primary production and nutrient cycling in sediments with benthic microalgae and transient accumulation of macroalgae. Limnology and Oceanography, 48(6): 2138–2150, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Diaz-Almela:2009:SDP

[DAMM^{+}09] Elena Díaz-Almela, Núria Marbà, Regino Martínez, Rocío Santiago, and Carlos M. Duarte. Seasonal dynamics of Posidonia oceanica in Magalluf Bay (Mallorca, Spain): Temperature effects on seagrass mortality. Limnology and Oceanography, 54 (6):2170–2182, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Davis:2000:CTF

[Dav00] Jana L. D. Davis. Changes in a tidepool fish assemblage on two scales of environmental variation: Seasonal and El Niño southern oscillation. Limnology and Oceanography, 45(6):1368– 1379, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Duarte:2004:CPM

[DAV04] Carlos M. Duarte, Susana Agustí, and Dolors Vaqué. Controls on planktonic metabolism in the Bay of Blanes, northwestern Mediterranean littoral. Limnology and Oceanography, 49(6): 2162–2170, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Duarte:2005:ETB

[DAV^{+}05] Carlos M. Duarte, Susana Agustí, Dolors Vaqué, Nona S. R. Agawin, Jordi Felipe, Emilio O. Casamayor, and Josep M. Gasol. Experimental test of bacteria-phytoplankton coupling in the Southern Ocean. Limnology and Oceanography, 50(6):1844– 1854, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Davis:2007:QEL

[DB07] Jenny Davis and Ronald Benner. Quantitative estimates of labile and semi-labile dissolved organic carbon in the western Arctic Ocean: a molecular approach. Limnology and Oceanography, 52 (6):2434–2444, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Dahnke:2008:NSE

[DBE08] Kirstin Dähnke, Enno Bahlmann, and Kay Emeis. A nitrate sink in estuaries? An assessment by means of stable nitrate isotopes in the Elbe estuary. Limnology and Oceanography, 53 (4):1504–1511, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Dimier:2009:GPD

[DBGR09] Céline Dimier, Christophe Brunet, Richard Geider, and John Raven. Growth and photoregulation dynamics of the picoeukaryote Pelagomonas calceolata in fluctuating light. Limnology and Oceanography, 54(3):823–836, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Dattagupta:2004:TCN

[DBS^{+}04] S. Dattagupta, D. C. Bergquist, E. B. Szalai, S. A. Macko, and C. R. Fisher. Tissue carbon, nitrogen, and sulfur stable isotope turnover in transplanted Bathymodiolus childressi mussels: Relation to growth and physiological condition. Limnology and Oceanography, 49(4):1144–1151, July 2004. CODEN LIOCAH. ISSN 0024-3590.

deBeer:2006:SFZ

[dBSN^{+}06] Dirk de Beer, Eberhard Sauter, Helge Niemann, Norbert Kaul, Jean-Paul Foucher, Ursula Witte, Michael Schlüter, and Antje Boetius. In situ fluxes and zonation of microbial activity in surface sediments of the Håkon Mosby Mud Volcano. Limnology and Oceanography, 51(3):1315–1331, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Dore:2002:SIV

[DBTK02] John E. Dore, Jennifer R. Brum, Luis M. Tupas, and David M. Karl. Seasonal and interannual variability in sources of nitrogen supporting export in the oligotrophic subtropical North Pacific Ocean. Limnology and Oceanography, 47(6):1595–1607, November 2002. CODEN LIOCAH. ISSN 0024-3590.

deBeer:2005:TMR

[dBWF^{+}05] Dirk de Beer, Frank Wenzhöfer, Timothy G. Ferdelman, Susan E. Boehme, Markus Huettel, Justus E. E. van Beusekom, Michael E. Böttcher, Niculina Musat, and Nicole Dubilier. Transport and mineralization rates in North Sea sandy intertidal sediments, Sylt–Rømø Basin, Wadden Sea. Limnology and Oceanography, 50(1):113–127, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Dulaiova:2006:GIR

[DBWS06] H. Dulaiova, W. C. Burnett, G. Wattayakorn, and P. Sojisuporn. Are groundwater inputs into river-dominated areas important? The Chao Phraya River — Gulf of Thailand. Limnology and Oceanography, 51(5):2232–2247, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Dillon:2005:NTB

[DC05] Kevin S. Dillon and Jeffrey P. Chanton. Nutrient transformations between rainfall and stormwater runoff in an urbanized coastal environment: Sarasota Bay, Florida. Limnology and Oceanography, 50(1):62–69, January 2005. CODEN LIOCAH. ISSN 0024-3590.

DelCastillo:2001:MSM

[DCC^{+}01] Carlos E. Del Castillo, Paula G. Coble, Robyn N. Conmy, Frank E. Müller-Karger, Lisa Vanderbloemen, and Gabriel A. Vargo. Multispectral in situ measurements of organic matter and chlorophyll fluorescence in seawater: Documenting the intrusion of the Mississippi River plume in the West Florida Shelf. Limnology and Oceanography, 46(7):1836–1843, November 2001. CODEN LIOCAH. ISSN 0024-3590.

daCunha:2009:IRD

[dCCL09] Leticia Cotrim da Cunha, Peter Croot, and Julie LaRoche. Influence of river discharge in the tropical and subtropical North Atlantic Ocean. Limnology and Oceanography, 54(2):644–648, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Dela-Cruz:2003:PGT

[DCMS03] Jocelyn Dela-Cruz, Jason H. Middleton, and Iain M. Suthers. Population growth and transport of the red tide dinoflagellate, Noctiluca scintillans, in the coastal waters off Sydney Australia, using cell diameter as a tracer. Limnology and Oceanography, 48 (2):656–674, March 2003. CODEN LIOCAH. ISSN 0024-3590.

DAndrea:2009:GEE

[DD09a] Anthony F. D’Andrea and Theodore H. DeWitt. Geochemical ecosystem engineering by the mud shrimp Upogebia pugettensis (Crustacea: Thalassinidae) in Yaquina Bay, Oregon: Densitydependent effects on organic matter remineralization and nutrient cycling. Limnology and Oceanography, 54(6):1911–1932, November 2009. CODEN LIOCAH. ISSN 0024-3590.

DSa:2009:SVC

[DD09b] Eurico J. D’Sa and Steven F. DiMarco. Seasonal variability and controls on chromophoric dissolved organic matter in a large river-dominated coastal margin. Limnology and Oceanography, 54(6):2233–2242, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Duarte:2009:EAS

[DdG09a] Carlos M. Duarte and Aurore Regaudie de Gioux. Erratum: Absorption spectral slopes and slope ratios as indicators of molecular weight, source, and photobleaching of chromophoric dissolved organic matter. Limnology and Oceanography, 55(6):1023, May 2009. CODEN LIOCAH. ISSN 0024-3590. See [HSR^{+}08].

Duarte:2009:TGP

[DdG09b] Carlos M. Duarte and Aurore Regaudie de Gioux. Thresholds of gross primary production for the metabolic balance of marine planktonic communities. Limnology and Oceanography, 54(3): 1015–1022, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Davies:2009:DDW

[DDL^{+}09] Andrew J. Davies, Gerard C. A. Duineveld, Marc S. S. Lavaleye, Magda J. N. Bergman, Hans van Haren, and J. Murray Roberts. Downwelling and deep-water bottom currents as food supply mechanisms to the cold-water coral Lophelia pertusa (Scleractinia) at the Mingulay Reef Complex. Limnology and Oceanography, 54(2):620–629, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Drake:2003:EEL

[DDZ03] Lisa A. Drake, Fred C. Dobbs, and Richard C. Zimmerman. Effects of epiphyte load on optical properties and photosynthetic potential of the seagrasses Thalassia testudinum Banks ex König and Zostera marina L. Limnology and Oceanography, 48(1 (part 2)):456–463, January 2003. CODEN LIOCAH. ISSN 0024-3590.

DeRobertis:2002:SDV

[De 02] Alex De Robertis. Size-dependent visual predation risk and the timing of vertical migration: an optimization model. Limnology and Oceanography, 47(4):925–933, July 2002. CODEN LI-OCAH. ISSN 0024-3590.

Dubowski:2003:IPL

[DES03] Y. Dubowski, J. Erez, and M. Stiller. Isotopic paleolimnology of Lake Kinneret. Limnology and Oceanography, 48(1):68–78, January 2003. CODEN LIOCAH. ISSN 0024-3590.

DeMott:2004:TZF

[DET04] William R. DeMott, Jennifer R. Edington, and Alan J. Tessier. Testing zooplankton food limitation across gradients of depth and productivity in small stratified lakes. Limnology and Oceanography, 49(4 (part 2)):1408–1416, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Dean:2007:IDB

[DFBA07] Walter E. Dean, Richard M. Forester, Jordon Bright, and Roger Y. Anderson. Influence of the diversion of Bear River into Bear Lake (Utah and Idaho) on the environment of deposition of carbonate minerals. Limnology and Oceanography, 52 (3):1094–1111, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Deuer:2006:IFB

[DG06] Susanne Menden Deuer and Daniel Grünbaum. Individual foraging behaviors and population distributions of a planktonic predator aggregating to phytoplankton thin layers. Limnology and Oceanography, 51(1):109–116, January 2006. CODEN LI-OCAH. ISSN 0024-3590.

delGiorgio:2002:LPP

[dGB02] Paul A. del Giorgio and Thierry C. Bouvier. Linking the physiologic and phylogenetic successions in free-living bacterial communities along an estuarine salinity gradient. Limnology and Oceanography, 47(2):471–486, March 2002. CODEN LIOCAH. ISSN 0024-3590.

DelCastillo:2000:DRC

[DGCMK00] Carlos E. Del Castillo, Fernando Gilbes, Paula G. Coble, and Frank E. Müller-Karger. On the dispersal of riverine colored dissolved organic matter over the West Florida Shelf. Limnology and Oceanography, 45(6):1425–1432, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Dryden:2004:IRD

[DGD04] Christina L. Dryden, Andrew S. Gordon, and John R. Donat. Interactive regulation of dissolved copper toxicity by an estuarine microbial community. Limnology and Oceanography, 49(4): 1115–1122, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Davidson:2007:IBI

[DGH^{+}07] Keith Davidson, Linda C. Gilpin, Mark C. Hart, Eric Fouilland, Elaine Mitchell, Ignacio Álvarez Calleja, Céline Laurent, Axel E. J. Miller, and Raymond J. G. Leakey. The influence of the balance of inorganic and organic nitrogen on the trophic dynamics of microbial food webs. Limnology and Oceanography, 52(5): 2147–2163, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Daniel:2006:SMP

[DGKA06] Cesar Daniel, Wilhelm Granéli, Emma S. Kritzberg, and Alexandre M. Anesio. Stimulation of metazooplankton by photochemically modified dissolved organic matter. Limnology and Oceanography, 51(1):101–108, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Dupont:2004:DCG

[DGL^{+}04] Christopher L. Dupont, Tyler J. Goepfert, Peter Lo, Liping Wei, and Beth A. Ahner. Diurnal cycling of glutathione in marine phytoplankton: Field and culture studies. Limnology and Oceanography, 49(4):991–996, July 2004. CODEN LIOCAH. ISSN 0024-3590.

DiTullio:2005:IIA

[DGM^{+}05] Giacomo R. DiTullio, Mark E. Geesey, Jennifer M. Maucher, Melissa B. Alm, Sarah F. Riseman, and Kenneth W. Bruland. Influence of iron on algal community composition and physiological status in the Peru upwelling system. Limnology and Oceanography, 50(6):1887–1907, November 2005. CODEN LI-OCAH. ISSN 0024-3590.

deGoeij:2008:TED

[dGMH^{+}08] Jasper M. de Goeij, Leon Moodley, Marco Houtekamer, Néstor M. Carballeira, and Fleur C. van Duyl. Tracing 13C-enriched dissolved and particulate organic carbon in the bacteria-containing coral reef sponge Halisarca caerulea: Evidence for DOM-feeding. Limnology and Oceanography, 53(4): 1376–1386, July 2008. CODEN LIOCAH. ISSN 0024-3590.

delGiorgio:2008:RIO

[dGP08] Paul A. del Giorgio and Michael L. Pace. Relative independence of organic carbon transport and processing in a large temperate river: The Hudson River as both pipe and reactor. Limnology and Oceanography, 53(1):185–197, January 2008. CODEN LIOCAH. ISSN 0024-3590.

DeMott:2001:DFL

[DGV01a] William R. DeMott, Ramesh D. Gulati, and Ellen Van Donk. Daphnia food limitation in three hypereutrophic Dutch lakes: Evidence for exclusion of large-bodied species by interfering filaments of cyanobacteria. Limnology and Oceanography, 47(4): 2054–2060, November 2001. CODEN LIOCAH. ISSN 0024-3590.

DeMott:2001:EDP

[DGV01b] William R. DeMott, Ramesh D. Gulati, and Ellen Van Donk. Effects of dietary phosphorus deficiency on the abundance, phosphorus balance, and growth of Daphnia cucullata in three hypereutrophic Dutch lakes. Limnology and Oceanography, 46(8): 1871–1880, November 2001. CODEN LIOCAH. ISSN 0024-3590.

deGoeij:2007:CCS

[dGvD07] Jasper M. de Goeij and Fleur C. van Duyl. Coral cavities are sinks of dissolved organic carbon (DOC). Limnology and Oceanography, 52(6):2608–2617, November 2007. CODEN LI-OCAH. ISSN 0024-3590.

Devol:2001:ROD

[DH01] Allan H. Devol and Hilairy E. Hartnett. Role of the oxygendeficient zone in transfer of organic carbon to the deep ocean. Limnology and Oceanography, 46(7):1684–1690, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Dunn:2005:TCS

[DHKH05] Alison M. Dunn, John C. Hogg, Andrew Kelly, and Melanie J. Hatcher. Two cues for sex determination in Gammarus duebeni: Adaptive variation in environmental sex determination? Limnology and Oceanography, 55(3):346–353, January 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Derry:2003:ERS

[DHP03] Alison M. Derry, Paul D. N. Hebert, and Ellie E. Prepas. Evolution of rotifers in saline and subsaline lakes: a molecular phylogenetic approach. Limnology and Oceanography, 48(2):675–685, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Duffy:2005:SPT

[DHTH05] Meghan A. Duffy, Spencer R. Hall, Alan J. Tessier, and Marianne Huebner. Selective predators and their parasitized prey: Are epidemics in zooplankton under top-down control? Limnology and Oceanography, 50(2):412–420, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Dham:2002:SVU

[DHWW02] Vivek V. Dham, Anjali Menezes Heredia, Sayeeda Wafar, and Mohideen Wafar. Seasonal variations in uptake and in situ regeneration of nitrogen in mangrove waters. Limnology and Oceanography, 47(1):241–254, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Dickey:2008:ILO

[DIMP08] T. D. Dickey, E. C. Itsweire, M. A. Moline, and M. J. Perry. Introduction to the Limnology and Oceanography special issue on Autonomous and Lagrangian Platforms and Sensors (ALPS). Limnology and Oceanography, 53(5 (part 2)):2057– 2061, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Dittmar:2004:ETD

[Dit04] Thorsten Dittmar. Evidence for terrigenous dissolved organic nitrogen in the Arctic deep sea. Limnology and Oceanography, 49(1):148–156, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Dixon:2000:MMI

[Dix00] David R. Dixon. Molecular and morphological identification of settlement-stage vent mussel larvae, Bathymodiolus azoricus (Bivalvia: Mytilidae), preserved in situ at active vent fields on the Mid-Atlantic Ridge. Limnology and Oceanography, 45(7): 1655–1661, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Delille:2007:BCD

[DJB^{+}07] Bruno Delille, Bruno Jourdain, Alberto V. Borges, Jean-Louis Tison, and Daniel Delille. Biogas (CO_{2}, O_{2}, dimethylsulfide) dynamics in spring Antarctic fast ice. Limnology and Oceanography, 52(4):1367–1379, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Dokulil:2006:TYS

[DJG^{+}06] Martin T. Dokulil, Albert Jagsch, Glen D. George, Orlane Anneville, Thomas Jankowski, Bernd Wahl, Brigitte Lenhart, Thorsten Blenckner, and Katrin Teubner. Twenty years of spatially coherent deepwater warming in lakes across Europe related to the North Atlantic Oscillation. Limnology and Oceanography, 52(6):2787–2793, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Dorval:2005:CSW

[DJH05] Emmanis Dorval, Cynthia M. Jones, and Robyn Hannigan. Chemistry of surface waters: Distinguishing fine-scale differences in sea grass habitats of Chesapeake Bay. Limnology and Oceanography, 50(4):1073–1083, July 2005. CODEN LIOCAH. ISSN 0024-3590.

DeRobertis:2000:SDV

[DJO00] Alex De Robertis, Jules S. Jaffe, and Mark D. Ohman. Sizedependent visual predation risk and the timing of vertical migration in zooplankton. Limnology and Oceanography, 45(8): 1838–1844, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Dolan:2006:CTM

[DJT06] J. R. Dolan, S. Jacquet, and J.-P. Torréton. Comparing taxonomic and morphological biodiversity of tintinnids (planktonic ciliates) of New Caledonia. Limnology and Oceanography, 51(2): 950–958, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Dutz:2006:TSS

[DK06] Jörg Dutz and Marja Koski. Trophic significance of solitary cells of the prymnesiophyte Phaeocystis globosa depends on cell type. Limnology and Oceanography, 51(3):1230–1238, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:2002:MSB

[DKA02] H. W.. Ducklow, D. L. Kirchman, and T. R. Anderson. The magnitude of spring bacterial production in the North Atlantic Ocean. Limnology and Oceanography, 47(6):1684–1693, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Decho:2003:SPI

[DKA^{+}03] Alan W. Decho, Tomohiro Kawaguchi, Mead A. Allison, Eric M. Louchard, R. Pamela Reid, F. Carol Stephens, Kenneth J. Voss, Robert A. Wheatcroft, and Brenton B. Taylor. Sediment properties influencing upwelling spectral reflectance signatures: The “biofilm gel effect”. Limnology and Oceanography, 48(1 (part 2)): 431–443, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Dupont:2009:DVM

[DKKA09] Nicolas Dupont, T. A. Klevjer, S. Kaartvedt, and D. L. Aksnes. Diel vertical migration of the deep-water jellyfish Periphylla periphylla simulated as individual responses to absolute light intensity. Limnology and Oceanography, 54(5):1765–1775, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Dierssen:2006:RBT

[DKRZ06] Heidi M. Dierssen, Raphael M. Kudela, John P. Ryan, and Richard C. Zimmerman. Red and black tides: Quantitative analysis of water-leaving radiance and perceived color for phytoplankton, colored dissolved organic matter, and suspended sediments. Limnology and Oceanography, 51(6):2646–2659, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Dutz:2008:CRU

[DKt08] Jörg Dutz, Marja Koski, and Sigrún H. Jínasdí ttir. Copepod reproduction is unaffected by diatom aldehydes or lipid composition. Limnology and Oceanography, 53(1):225–235, January 2008. CODEN LIOCAH. ISSN 0024-3590.

DiBacco:2000:DAE

[DL00] Claudio DiBacco and Lisa A. Levin. Development and application of elemental fingerprinting to track the dispersal of marine invertebrate larvae. Limnology and Oceanography, 45(4):871– 880, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Decaestecker:2004:HPE

[DLDE04] Ellen Decaestecker, Christophe Lefever, Luc De Meester, and Dieter Ebert. Haunted by the past: Evidence for dormant stage banks of microparasites and epibionts of Daphnia. Limnology and Oceanography, 49(4 (part 2)):1355–1364, July 2004. CO-DEN LIOCAH. ISSN 0024-3590.

delaFuente:2008:EIW

[dlFSIN08] Alberto de la Fuente, Kenji Shimizu, Jörg Imberger, and Yarko Niño. The evolution of internal waves in a rotating, stratified, circular basin and the influence of weakly nonlinear and nonhydrostatic accelerations. Limnology and Oceanography, 53(6): 2738–2748, November 2008. CODEN LIOCAH. ISSN 0024-3590.

delaCuesta:2009:IAM

[dlM09] John L. de laCuesta and Steven L. Manley. Iodine assimilation by marine diatoms and other phytoplankton in nitratereplete conditions. Limnology and Oceanography, 54(5):1653– 1664, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Dai:2009:DVS

[DLZ^{+}09] Minhan Dai, Zhongming Lu, Weidong Zhai, Baoshan Chen, Zhimian Cao, Kuanbo Zhou, Wei-Jun Cai, and Chen-Tung Arthur Chenc. Diurnal variations of surface seawater pCO_{2} in contrasting coastal environments. Limnology and Oceanography, 54(3):735–745, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Diiez:2004:DEP

[DMEPA04] Beatriz Diíez, Ramon Massana, Marta Estrada, and Carlos Pedrós-Alió. Distribution of eukaryotic picoplankton assemblages across hydrographic fronts in the Southern Ocean, studied by denaturing gradient gel electrophoresis. Limnology and Oceanography, 49(4):1022–1034, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Deines:2009:TVP

[DMJ09] Peter Deines, Carsten Matz, and Klaus Jürgens. Toxicity of violacein-producing bacteria fed to bacterivorous freshwater plankton. Limnology and Oceanography, 54(4):1343–1352, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Denny:2003:EWV

[DMS^{+}03] M. W. Denny, L. P. Miller, M. D. Stokes, L. J. H. Hunt, and B. S. T. Helmuth. Extreme water velocities: Topographical amplification of wave-induced flow in the surf zone of rocky shores. Limnology and Oceanography, 48(1):1–8, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Dupont:2004:NCB

[DNB^{+}04] Christopher L. Dupont, Robert K. Nelson, Saj Bashir, James W. Moffett, and Beth A. Ahner. Novel copper-binding and nitrogenrich thiols produced and exuded by Emiliania huxleyi. Limnology and Oceanography, 49(5):1754–1762, September 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Dong:2006:SNU

[DNS06] Liang F. Dong, David B. Nedwell, and Andrew Stott. Sources of nitrogen used for denitrification and nitrous oxide formation in sediments of the hypernutrified Colne, the nutrified Humber, and the oligotrophic Conwy estuaries, United Kingdom. Limnology and Oceanography, 51(1 (part 2)):545–557, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Dodds:2006:ETS

[Dod06] Walter K. Dodds. Eutrophication and trophic state in rivers and streams. Limnology and Oceanography, 51(1 (part 2)):671–680, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Dove:2006:RHP

[DOE^{+}06] Sophie Dove, Juan Carlos Ortiz, Susana Enríquez, Maoz Fine, Paul Fisher, Roberto Iglesias-Prieto, Dan Thornhill, and Ove Hoegh-Guldberg. Response of holosymbiont pigments from the scleractinian coral Montipora monasteriata to short-term heat stress. Limnology and Oceanography, 51(2):1149–1158, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Davis:2008:GSP

[DORS08] Russ E. Davis, Mark D. Ohman, Daniel L. Rudnick, and Jeff T. Sherman. Glider surveillance of physics and biology in the southern California Current System. Limnology and Oceanography, 53(5 (part 2)):2151–2168, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Downing:2002:BRR

[Dow02] John A. Downing. Book review: Rabalais, Nancy N., and R. Eugene Turner [eds.]. 2001. Coastal hypoxia: Consequences for living resources and ecosystems. Coastal and Estuarine Studies,

58. American Geophysical Union, Washington, D.C. vii + 463 p. US$68. ISBN 0-87590-272-3. Limnology and Oceanography, 47(4):1269, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Dore:2001:PPD

[DP01] John E. Dore and John C. Priscu. Phytoplankton phosphorus deficiency and alkaline phosphatase activity in the McMurdo Dry Valley lakes, Antarctica. Limnology and Oceanography, 46 (6):1331–1346, September 2001. CODEN LIOCAH. ISSN 0024-3590.

DeLaRocha:2004:RTW

[DP04] Christina L. De La Rocha and Uta Passow. Recovery of Thalassiosira weissflogii from nitrogen and silicon starvation. Limnology and Oceanography, 49(1):245–255, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Downing:2006:GAS

[DPC^{+}06] J. A. Downing, Y. T. Prairie, J. J. Cole, C. M. Duarte, L. J. Tranvik, R. G. Striegl, W. H. McDowell, P. Kortelainen, N. F. Caraco, J. M. Melack, and J. J. Middelburg. The global abundance and size distribution of lakes, ponds, and impoundments. Limnology and Oceanography, 51(5):2388–2397, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Drazen:2008:BAB

[DPC^{+}08] Jeffrey C. Drazen, Brian N. Popp, C. Anela Choy, Tara Clemente, Lisa De Forest, and Kenneth L. Smith, Jr. Bypassing the abyssal benthic food web: Macrourid diet in the eastern North Pacific inferred from stomach content and stable isotopes analyses. Limnology and Oceanography, 53(6):2644–2654, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Droscher:2009:CCS

[DPFL09] Iris Dröscher, Alain Patoine, Kerri Finlay, and Peter R. Leavitt. Climate control of the spring clear-water phase through the transfer of energy and mass to lakes. Limnology and Oceanography, 54(6 (part 2)):2469–2480, November 2009. CODEN LI-OCAH. ISSN 0024-3590.

Duffy:2000:PEP

[DPK^{+}00] Meghan A. Duffy, Linda J. Perry, Colleen M. Kearns, Lawrence J. Weider, and Nelson G. Hairston, Jr. Paleogenetic evidence for a past invasion of Onondaga Lake, New York, by exotic Daphnia curvirostris using mtDNA from dormant eggs. Limnology and Oceanography, 45(6):1409–1414, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Douglas:2000:RRS

[DR00] Richard W. Douglas and Brian Rippey. The random redistribution of sediment by wind in a lake. Limnology and Oceanography, 45(3):686–694, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Dyhrman:2006:PRA

[DR06] Sonya T. Dyhrman and Kathleen C. Ruttenberg. Presence and regulation of alkaline phosphatase activity in eukaryotic phytoplankton from the coastal ocean: Implications for dissolved organic phosphorus remineralization. Limnology and Oceanography, 51(3):1381–1390, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Doxaran:2009:SVL

[DRM^{+}09] David Doxaran, Kevin Ruddick, David McKee, Bernard Gentili, Dominique Tailliez, Malik Chami, and Marcel Babin. Spectral variations of light scattering by marine particles in coastal waters, from the visible to the near infrared. Limnology and Oceanography, 54(4):1257–1271, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Droppo:2007:BCS

[DRSL07] Ian G. Droppo, Nathalie Ross, Michael Skafel, and Steven N. Liss. Biostabilization of cohesive sediment beds in a freshwater wave-dominated environment. Limnology and Oceanography, 52 (2):577–589, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Donner:2007:HCC

[DS07a] Simon D. Donner and Donald Scavia. How climate controls the flux of nitrogen by the Mississippi River and the development of hypoxia in the Gulf of Mexico. Limnology and Oceanography, 52 (2):856–861, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Drazen:2007:DRT

[DS07b] Jeffrey C. Drazen and Brad A. Seibel. Depth-related trends in metabolism of benthic and benthopelagic deep-sea fishes. Limnology and Oceanography, 52(5):2306–2316, September 2007. CODEN LIOCAH. ISSN 0024-3590.

DellAnno:2002:QBC

[DSD02] Antonio Dell’Anno, Bompadre Stefano, and Roberto Danovaro. Quantification, base composition, and fate of extracellular DNA in marine sediments. Limnology and Oceanography, 47(3):899– 905, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Ferrao-Filho:2007:SHZ

[dSFFTD07] Aloysio da S. Ferrão-Filho, Alan J. Tessier, and William R. De-Mott. Sensitivity of herbivorous zooplankton to phosphorusdeficient diets: Testing stoichiometric theory and the growth rate hypothesis. Limnology and Oceanography, 52(1):407–415, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Dupuis:2008:COE

[DST08] Derrick D. Dupuis, Jan-Erik Svensson, and Derek J. Taylor. The cryptic origins of environment-indicating phantom midges (Chaoborus). Limnology and Oceanography, 53(1):236–243, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Doyle:2005:IET

[DSW05] Shaina A. Doyle, Jasmine E. Saros, and Craig E. Williamson. Interactive effects of temperature and nutrient limitation on the response of alpine phytoplankton growth to ultraviolet radiation. Limnology and Oceanography, 50(5):1362–1367, September 2005. CODEN LIOCAH. ISSN 0024-3590.

deSieyes:2008:SDN

[dSYL^{+}08] Nicholas R. de Sieyes, Kevan M. Yamahara, Blythe A. Layton, Elizabeth H. Joyce, and Alexandria B. Boehm. Submarine discharge of nutrient-enriched fresh groundwater at Stinson Beach, California is enhanced during neap tides. Limnology and Oceanography, 53(4):1434–1445, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Doskey:2000:SCA

[DT00] Paul V. Doskey and Robert W. Talbot. Sediment chronologies of atmospheric deposition in a precipitation-dominated seepage lake. Limnology and Oceanography, 45(4):895–904, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Donahue:2003:RSR

[DTFL03] William F. Donahue, Michael A. Turner, David L. Findlay, and Peter R. Leavitt. The role of solar radiation in structuring the shallow benthic communities of boreal forest lakes. Limnology and Oceanography, 48(1):31–47, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Darchambeau:2005:IZS

[DTL^{+}05] François Darchambeau, Isabelle Thys, Bruno Leporcq, Lucien Hoffmann, and Jean-Pierre Descy. Influence of zooplankton stoichiometry on nutrient sedimentation in a lake system. Limnology and Oceanography, 50(3):905–913, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Davey:2008:NLP

[DTM^{+}08] Margaret Davey, Glen A. Tarran, Matthew M. Mills, Celine Ridame, Richard J. Geider, and Julie LaRoche. Nutrient limitation of picophytoplankton photosynthesis and growth in the tropical North Atlantic. Limnology and Oceanography, 53(5):1722–1733, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Duffy:2009:SAP

[Duf09] Meghan A. Duffy. Staying alive: The post-consumption fate of parasite spores and its implications for disease dynamics. Limnology and Oceanography, 54(3):770–773, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Duplisea:2000:BOB

[Dup00] Daniel E. Duplisea. Benthic organism biomass size-spectra in the Baltic Sea in relation to the sediment environment. Limnology and Oceanography, 45(3):558–568, May 2000. CODEN LIOCAH. ISSN 0024-3590.

delValle:2007:LSP

[dVKJK07] Daniela A. del Valle, David J. Kieber, Bisgrove John, and Ronald P. Kiene. Light-stimulated production of dissolved DMSO by a particle-associated process in the Ross Sea, Antarctica. Limnology and Oceanography, 52(6):2456–2466, November 2007. CODEN LIOCAH. ISSN 0024-3590.

delValle:2009:BCD

[dVKT^{+}09] Daniela A. del Valle, David J. Kieber, Dierdre A. Toole, Jordan Brinkley, and Ronald P. Kienea. Biological consumption of dimethylsulfide (DMS) and its importance in DMS dynamics in the Ross Sea, Antarctica. Limnology and Oceanography, 54(3): 785–798, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Derycke:2007:CPN

[DVV^{+}07] S. Derycke, R. Van Vynckt, J. Vanoverbeke, M. Vincx, and T. Moens. Colonization patterns of Nematoda on decomposing algae in the estuarine environment: Community assembly and genetic structure of the dominant species Pellioditis marina. Limnology and Oceanography, 52(3):992–1001, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Dyhrman:2002:CSD

[DWA^{+}02] Sonya T. Dyhrman, Eric A. Webb, Donald M. Anderson, James W. Moffett, and John B. Waterbury. Cell-specific detection of phosphorus stress in Trichodesmium from the Western North Atlantic. Limnology and Oceanography, 47(6):1832–1836, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Dierssen:2003:OCR

[DZL^{+}03] Heidi M. Dierssen, Richard C. Zimmerman, Robert A. Leathers, T. Valerie Downes, and Curtiss O. Davis. Ocean color remote sensing of seagrass and bathymetry in the Bahamas Banks by high-resolution airborne imagery. Limnology and Oceanography, 48(1 (part 2)):444–455, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Eiane:2002:SSM

[EAO^{+}02] Ketil Eiane, Dag L. Aksnes, Mark D. Ohman, Simon Wood, and Monica B. Martinussen. Stage-specific mortality of Calanus spp. under different predation regimes. Limnology and Oceanography, 47(3):636–645, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Engelhaupt:2001:SCH

[EB01] Erika Engelhaupt and Thomas S. Bianchi. Sources and composition of high-molecular-weight dissolved organic carbon in a southern Louisiana tidal stream (Bayou Trepagnier). Limnology and Oceanography, 46(4):917–926, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Engstrom:2007:HMI

[EBS07] Daniel R. Engstrom, Steven J. Balogh, and Edward B. Swain. History of mercury inputs to Minnesota lakes: Influences of watershed disturbance and localized atmospheric deposition. Limnology and Oceanography, 52(6):2467–2483, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Ensign:2005:CTS

[ED05] Scott H. Ensign and Martin W. Doyle. In-channel transient storage and associated nutrient retention: Evidence from experimental manipulations. Limnology and Oceanography, 50(6): 1740–1751, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Erga:2003:NAM

[EDF^{+}03] Svein Rune Erga, Maren Dybwad, Øyvind Frette, Jon Kåre Lotsberg, and Kjetil Aursland. New aspects of migratory behavior of phytoplankton in stratified waters: Effects of halocline strength and light on Tetraselmis sp. (Prasinophyceae) in an artificial water column. Limnology and Oceanography, 48(4): 1202–1213, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Eyre:2005:BMN

[EF05] Bradley D. Eyre and Angus J. P. Ferguson. Benthic metabolism and nitrogen cycling in a subtropical east Australian estuary (Brunswick): Temporal variability and controlling factors. Limnology and Oceanography, 50(1):81–96, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Eyre:2008:MCS

[EGP08] Bradley D. Eyre, Ronnie N. Glud, and Nicole Patten. Mass coral spawning: a natural large-scale nutrient addition experiment. Limnology and Oceanography, 54(3):997–1013, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Engel:2002:TDC

[EGPA02] Anja Engel, Sarah Goldthwait, Uta Passow, and Alice Alldredge. Temporal decoupling of carbon and nitrogen dynamics in a mesocosm diatom bloom. Limnology and Oceanography, 47(3): 753–761, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Ellwood:2000:IZI

[EH00] Michael J. Ellwood and Keith A. Hunter. The incorporation of zinc and iron into the frustule of the marine diatom Thalassiosira pseudonana. Limnology and Oceanography, 45(7):1517– 1524, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Edwards:2006:RCC

[EJL^{+}06] M. Edwards, D. G. Johns, S. C. Leterme, E. Svendsen, and A. J. Richardson. Regional climate change and harmful algal blooms in the northeast Atlantic. Limnology and Oceanography, 51(2): 820–829, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Evans:2007:RSV

[EKD^{+}07] Claire Evans, Susanne V. Kadner, Louise J. Darroch, William H. Wilson, Peter S. Liss, and Gillian Malin. The relative significance of viral lysis and microzooplankton grazing as pathways of dimethylsulfoniopropionate (DMSP) cleavage: an Emiliania huxleyi culture study. Limnology and Oceanography, 52(3):1036– 1045, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Ellwood:2007:SBD

[EKdF07] Michael J. Ellwood, Michelle Kelly, and Bertrand Richer de Forges. Silica banding in the deep-sea lithistid sponge Corallistes undulatus: Investigating the potential influence of diet and environment on growth. Limnology and Oceanography, 52(5): 1865–1873, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Elmgren:2002:BRL

[Elm02] Ragnar Elmgren. Book review: Livingston, R. J. 2001. Eutrophication processes in coastal systems. CRC Press. 318 p. US$100. ISBN 0-8493-9062-1. Limnology and Oceanography, 47(2):612, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Eyre:2002:CNP

[EM02] Bradley D. Eyre and Lester J. McKee. Carbon, nitrogen, and phosphorus budgets for a shallow subtropical coastal embayment (Moreton Bay, Australia). Limnology and Oceanography, 47(4): 1043–1055, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Egleston:2008:NLZ

[EM08] Eric S. Egleston and François M. M. Morel. Nickel limitation and zinc toxicity in a urea-grown diatom. Limnology and Oceanography, 53(6):2462–2471, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Enriquez:2005:MSC

[EMIP05] Susana Enríquez, Eugenio R. Méndez, and Roberto Iglesias-Prieto. Multiple scattering on coral skeletons enhances light absorption by symbiotic algae. Limnology and Oceanography, 50 (4):1025–1032, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Evans:2008:IWE

[EMK08] Mary Anne Evans, Sally MacIntyre, and George W. Kling. Internal wave effects on photosynthesis: Experiments, theory, and modeling. Limnology and Oceanography, 53(1):339–353, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Escapa:2007:DIE

[EMPI07] Mauricio Escapa, Darío R. Minkoff, Gerardo M. E. Perillo, and Oscar Iribarne. Direct and indirect effects of burrowing crab Chasmagnathus granulatus activities on erosion of southwest Atlantic Sarcocornia-dominated marshes. Limnology and Oceanography, 52(6):2340–2349, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Evans:2006:ITE

[EMWL06] Claire Evans, Gillian Malin, William H. Wilson, and Peter S. Liss. Infectious titres of Emiliania huxleyi virus 86 are reduced by exposure to millimolar dimethyl sulfide and acrylic acid. Limnology and Oceanography, 51(5):2468–2471, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Engstrom:2009:AAO

[EPD09] Pia Engström, C. Ryan Penton, and Allan H. Devol. Anaerobic ammonium oxidation in deep-sea sediments off the Washington margin. Limnology and Oceanography, 54(5):1643–1652, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Enders:2008:CSS

[EPP^{+}08] Sara K. Enders, Mark Pagani, Silvio Pantoja, Jill S. Baron, Alexander P. Wolfe, Nikolai Pedentchouk, and Lilian Nuñez. Compound-specific stable isotopes of organic compounds from lake sediments track recent environmental changes in an alpine ecosystem, Rocky Mountain National Park, Colorado. Limnology and Oceanography, 53(4):1468–1478, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Elzenga:2000:REC

[EPS00] J. Theo M. Elzenga, Hidde B. A. Prins, and Jacqueline Stefels. The role of extracellular carbonic anhydrase activity in inorganic carbon utilization of Phaeocystis globosa (Prymnesiophyceae): a comparison with other marine algae using the isotopic disequilibrium technique. Limnology and Oceanography, 45(2):372–380, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Eriksson:2006:CDC

[ERH06] Britas Klemens Eriksson, Anja Rubach, and Helmut Hillebrand. Community dominance by a canopy species controls the relationship between macroalgal production and species richness. Limnology and Oceanography, 51(5):1813–1818, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Edlund:2000:YHR

[ES00] Mark B. Edlund and Eugene F. Stoermer. A 200,000-year, high-resolution record of diatom productivity and community makeup from Lake Baikal shows high correspondence to the marine oxygen-isotope record of climate change. Limnology and Oceanography, 45(4):948–962, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Erlandsson:2006:SMO

[ESA06] Carina P. Erlandsson, Anders Stigebrandt, and Lars Arneborg. The sensitivity of minimum oxygen concentrations in a fjord to changes in biotic and abiotic external forcing. Limnology and Oceanography, 51(1 (part 2)):631–638, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Etemad-Shahidi:2001:ATT

[ESI01] Amir Etemad-Shahidi and Jörg Imberger. Anatomy of turbulence in thermally stratified lakes. Limnology and Oceanography, 46(5):1158–1170, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Emmenegger:2001:LIR

[ESSS01] Lukas Emmenegger, René Schönenberger, Laura Sigg, and Barbara Sulzberger. Light-induced redox cycling of iron in circumneutral lakes. Limnology and Oceanography, 46(1):49–61, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Ebersbach:2008:SPP

[ET08] F. Ebersbach and T. W. Trull. Sinking particle properties from polyacrylamide gels during the Kerguelen Ocean and Plateau compared Study (KEOPS): Zooplankton control of carbon export in an area of persistent natural iron inputs in the Southern Ocean. Limnology and Oceanography, 53(1):212–224, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Eisner:2003:RPP

[ETCP03] Lisa B. Eisner, Michael S. Twardowski, Timothy J. Cowles, and Mary Jane. Perry. Resolving phytoplankton photoprotective : photosynthetic carotenoid ratios on fine scales using in situ spectral absorption measurements. Limnology and Oceanography, 48 (2):632–646, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Earl:2007:NSS

[EVW07] Stevan R. Earl, H. Maurice Valett, and Jackson R. Webster. Nitrogen spiraling in streams: Comparisons between stable isotope tracer and nutrient addition experiments. Limnology and Oceanography, 52(4):1718–1723, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Ebert:2003:RFA

[EW03] Thomas A. Ebert and Susan L. Williams. Rejoinder to Fourqurean et al. (2003). Limnology and Oceanography, 49(4): 2074–2075, September 2003. CODEN LIOCAH. ISSN 0024-3590. See [FMD03].

Eggert:2007:WBF

[EW07] S. L. Eggert and J. B. Wallace. Wood biofilm as a food resource for stream detritivores. Limnology and Oceanography, 52(3): 1239–1245, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Evans:2008:PGO

[EW08] Claire Evans and William H. Wilson. Preferential grazing of Oxyrrhis marina on virus infected Emiliania huxleyi. Limnology and Oceanography, 53(5):2035–2040, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Ebert:2002:MEA

[EWE02] Thomas A. Ebert, Susan L. Williams, and Patrick J. Ewanchuk. Mortality estimates from age distributions: Critique of a method used to study seagrass dynamics. Limnology and Oceanography, 47(2):600–603, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Engel:2005:TDE

[EZA^{+}05] Anja Engel, Ingrid Zondervan, Katrien Aerts, Luc Beaufort, Albert Benthien, Lei Chou, Bruno Delille, Jean-Pierre Gattuso, Jerome Harlay, Christel Heemann, Linn Hoffmann, Stéphan Jacquet, Jens Nejstgaard, Marie-Dominique Pizay, Emma Rochelle-Newall, Uta Schneider, Anja Terbrueggen, and Ulf Riebesell. Testing the direct effect of CO_{2} concentration on a bloom of the coccolithophorid Emiliania huxleyi in mesocosm experiments. Limnology and Oceanography, 50(4):493–507, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Fernandez:2003:EMP

[FA03] Diego Fernández and José Luis Acuña. Enhancement of marine phytoplankton blooms by appendicularian grazers. Limnology and Oceanography, 48(2):587–593, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Fabricius:2006:EIF

[Fab06] Katharina E. Fabricius. Effects of irradiance, flow, and colony pigmentation on the temperature microenvironment around corals: Implications for coral bleaching? Limnology and Oceanography, 51(1):30–37, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Falter:2005:ESR

[FAC05] James L. Falter, Marlin J. Atkinson, and Carlos F. M. Coimbra. Effects of surface roughness and oscillatory flow on the dissolution of plaster forms: Evidence for nutrient mass transfer to coral reef communities. Limnology and Oceanography, 55(3): 246–254, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Falter:2001:PRR

[FAL01] James L. Falter, M. J. Atkinson, and Chris Langdon. Production-respiration relationships at different timescales within the Biosphere 2 coral reef biome. Limnology and Oceanography, 46(7):1653–1660, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Falter:2007:ENT

[FAL^{+}07] James L. Falter, Marlin J. Atkinson, Ryan J. Lowe, Stephen G. Monismith, and Jeffrey R. Koseff. Effects of nonlocal turbulence on the mass transfer of dissolved species to reef corals. Limnology and Oceanography, 52(1):274–285, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Falter:2004:MTL

[FAM04] James L. Falter, Marlin J. Atkinson, and Mark A. Merrifield. Mass-transfer limitation of nutrient uptake by a wave-dominated reef flat community. Limnology and Oceanography, 49(5):1820– 1831, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Frenette:2006:HCU

[FAM^{+}06] Jean-Jacques Frenette, Michael T. Arts, Jean Morin, Denis Gratton, and Carl Martin. Hydrodynamic control of the underwater light climate in fluvial Lac Saint-Pierre. Limnology and Oceanography, 51(6):2632–2645, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Fennel:2003:SMP

[FB03] Katja Fennel and Emmanuel Boss. Subsurface maxima of phytoplankton and chlorophyll: Steady-state solutions from a simple model. Limnology and Oceanography, 48(4):1521–1534, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Fulton:2005:WIW

[FB05] Christopher J. Fulton and David R. Bellwood. Wave-induced water motion and the functional implications for coral reef fish assemblages. Limnology and Oceanography, 55(3):255–264, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Frade:2008:SPC

[FBW^{+}08] P. R. Frade, P. Bongaerts, A. J. S. Winkelhagen, L. Tonk, and R. P. M. Bak. In situ photobiology of corals over large depth ranges: a multivariate analysis on the roles of environment, host, and algal symbiont. Limnology and Oceanography, 53(6):2711– 2723, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Farias:2009:DNO

[FCGC^{+}09] Laura Farías, Maribeb Castro-González, Marcela Cornejo, José Charpentier, Juan Faúndez, Narin Boontanon, and Naohiro Yoshida. Denitrification and nitrous oxide cycling within the upper oxycline of the eastern tropical South Pacific oxygen minimum zone. Limnology and Oceanography, 54(1):132–144, January 2009. CODEN LIOCAH. ISSN 0024-3590.

French:2006:LTC

[FCJ^{+}06] Todd D. French, Linda M. Campbell, Donald A. Jackson, John M. Casselman, Wolfgang A. Scheider, and Al Hayton. Long-term changes in legacy trace organic contaminants and mercury in Lake Ontario salmon in relation to source controls, trophodynamics, and climatic variability. Limnology and Oceanography, 52(6):2794–2807, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Fones:2001:HRM

[FDH^{+}01] Gary R. Fones, William Davison, Ola Holby, Bo Barker Jorgensen, and Bo Thamdrup. High-resolution metal gradients measured by in situ DGT/DET deployment in Black Sea sediments using an autonomous benthic lander. Limnology and Oceanography, 46(4):982–988, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Fabricius:2008:DGI

[FDP^{+}08] Katharina E. Fabricius, Glenn De’ath, Marji Lee Puotinen, Terry Done, Timothy F. Cooper, and Scott C. Burgess. Disturbance gradients on inshore and offshore coral reefs caused by a severe tropical cyclone. Limnology and Oceanography, 53(2): 690–704, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Ferdie:2004:RSC

[FF04] Meredith Ferdie and James W. Fourqurean. Responses of seagrass communities to fertilization along a gradient of relative availability of nitrogen and phosphorus in a carbonate environment. Limnology and Oceanography, 49(6):2082–2094, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Frederiksen:2006:ODR

[FG06] Morten S. Frederiksen and Ronnie N. Glud. Oxygen dynamics in the rhizosphere of Zostera marina: a two-dimensional planar optode study. Limnology and Oceanography, 51(2):1072–1083, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Fouilland:2003:EUB

[FGM^{+}03] Eric Fouilland, Michel Gosselin, Behzad Mostajir, Maurice Levasseur, Jean-Pierre Chanut, Serge Demers, and Stephen de Mora. Effects of ultraviolet-B radiation and vertical mixing on nitrogen uptake by a natural planktonic community shifting from nitrate to silicic acid deficiency. Limnology and Oceanography, 48(1):18–30, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Fer:2003:DLC

[FH03] Ilker Fer and Peter M. Haugan. Dissolution from a liquid CO_{2} lake disposed in the deep ocean. Limnology and Oceanography, 48(2):872–883, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Fitter:2009:MFW

[FH09] Anja Fitter and Helmut Hillebrand. Microbial food web structure affects bottom-up effects and elemental stoichiometry in periphyton assemblages. Limnology and Oceanography, 54(6): 2183–2200, November 2009. CODEN LIOCAH. ISSN 0024-

Fielding:2009:AAE

[FHB^{+}09] Samuel R. Fielding, Jens O. Herrle, Jórg Bollmann, Richard H. Worden, and David J. S. Montagnes. Assessing the applicability of Emiliania huxleyi coccolith morphology as a sea-surface salinity proxy. Limnology and Oceanography, 54(5):1475–1480, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Frederiksen:2006:TSV

[FHBK06] Morten S. Frederiksen, Marianne Holmer, Jens Borum, and Hilary Kennedy. Temporal and spatial variation of sulfide invasion in eelgrass (Zostera marina) as reflected by its sulfur isotopic composition. Limnology and Oceanography, 51(5):2308–2318, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Fisher:2006:CEC

[FHBW06] T. R. Fisher, J. D. Hagy III, W. R. Boynton, and M. R. Williams. Cultural eutrophication in the Choptank and Patuxent estuaries of Chesapeake Bay. Limnology and Oceanography, 51(1 (part 2)):435–447, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Faulkenham:2003:EDI

[FHDKR03] Shari E. Faulkenham, Roland I. Hall, Peter J. Dillon, and Tammy Karst-Riddoch. Effects of drought-induced acidification on diatom communities in acid-sensitive Ontario lakes. Limnology and Oceanography, 48(4):1662–1673, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Finkel:2001:LAS

[Fin01] Zoe Vanessa Finkel. Light absorption and size scaling of lightlimited metabolism in marine diatoms. Limnology and Oceanography, 46(1):86–94, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Finlay:2004:PCL

[Fin04] Jacques C. Finlay. Patterns and controls of lotic algal stable carbon isotope ratios. Limnology and Oceanography, 49(3):850– 861, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Fisher:2000:REM

[Fis00] Stuart G. Fisher. River ecology and management: Lessons from the Pacific coastal ecoregion. Limnology and Oceanography, 45 (6):1445, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Frantz:2000:GRP

[FKCF00] Brian R. Frantz, Michaele Kashgarian, Kenneth H. Coale, and Michael S. Foster. Growth rate and potential climate record from a rhodolith using ^{14}C accelerator mass spectrometry. Limnology and Oceanography, 45(8):1773–1777, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Findlay:2001:CIA

[FKS^{+}01] D. L. Findlay, S. E. M. Kasian, M. P. Stainton, K. Beaty, and M. Lyng. Climatic influences on algal populations of boreal forest lakes in the Experimental Lakes Area. Limnology and Oceanography, 46(7):1784–1793, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Foy:2006:EDR

[FL06] R. H. Foy and S. D. Lennox. Evidence for a delayed response of riverine phosphorus exports from increasing agricultural catchment pressures in the Lough Neagh catchment. Limnology and Oceanography, 51(1 (part 2)):655–663, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Fodrie:2008:LJH

[FL08] F. Joel Fodrie and Lisa A. Levin. Linking juvenile habitat utilization to population dynamics of California halibut. Limnology and Oceanography, 53(2):799–812, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Forrest:2008:CDT

[FLPL08] Alexander L. Forrest, Bernard E. Laval, Roger Pieters, and Darlene S. S. Lim. Convectively driven transport in temperate lakes. Limnology and Oceanography, 53(5 (part 2)):2321–2332, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Fer:2002:CEV

[FLT02a] Ilker Fer, Ulrich Lemmin, and S. A. Thorpe. Contribution of entrainment and vertical plumes to the winter cascading of cold shelf waters in a deep lake. Limnology and Oceanography, 47(2): 576–580, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Fer:2002:OMN

[FLT02b] Ilker Fer, Ulrich Lemmin, and S. A. Thorpe. Observations of mixing near the sides of a deep lake in winter. Limnology and Oceanography, 47(2):535–544, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Finlay:2009:RST

[FLWP09] K. Finlay, P. R. Leavitt, B. Wissel, and Y. T. Prairie. Regulation of spatial and temporal variability of carbon flux in six hard-water lakes of the northern Great Plains. Limnology and Oceanography, 54(6 (part 2)):2553–2564, November 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Flynn:2003:DWN

[Fly03] Kevin J. Flynn. Do we need complex mechanistic photoacclimation models for phytoplankton? Limnology and Oceanography, 48(6):2243–2249, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Fourqurean:2003:ESP

[FMD03] James W. Fourqurean, Núria Marbà, and Carlos M. Duarte. Elucidating seagrass population dynamics: Theory, constraints, and practice. Limnology and Oceanography, 49(4):2070–2074, September 2003. CODEN LIOCAH. ISSN 0024-3590. See rejoinder [EW03].

Flynn:2001:CTI

[FMG01] Kevin J. Flynn, Helen Marshall, and Richard J. Geider. A comparison of two n-irradiance interaction models of phytoplankton growth. Limnology and Oceanography, 46(7):1794–1802, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Fu:2008:IBC

[FMG^{+}08] Fei-Xue Fu, Margaret R. Mulholland, Nathan S. Garcia, Aaron Beck, Peter W. Bernhardt, Mark E. Warner, Sergio A. Sañudo-Wilhelmy, and David A. Hutchins. Interactions between changing pCO_{2}, N_{2} fixation, and Fe limitation in the marine unicellular cyanobacterium crocosphaera. Limnology and Oceanography, 53(6):2472–2484, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Fuchs:2004:SBG

[FMS04] Heidi L. Fuchs, Lauren S. Mullineaux, and Andrew R. Solow. Sinking behavior of gastropod larvae (Ilyanassa obsoleta) in turbulence. Limnology and Oceanography, 49(6):1937–1948, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Fitt:2000:SPT

[FMWC00] W. K. Fitt, F. K. McFarland, M. E. Warner, and G. C. Chilcoat. Seasonal patterns of tissue biomass and densities of symbiotic dinoflagellates in reef corals and relation to coral bleaching. Limnology and Oceanography, 45(3):677–685, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Fowler:2000:TVP

[FNLG00] Scott W. Fowler, Victor E. Noshkin, Jacques La Rosa, and Janine Gastaud. Temporal variations in plutonium and americium inventories and their relation to vertical transport in the northwestern Mediterranean Sea. Limnology and Oceanography, 45 (2):446–458, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Fuchs:2007:ETM

[FNM07] Heidi L. Fuchs, Michael G. Neubert, and Lauren S. Mullineaux. Effects of turbulence-mediated larval behavior on larval supply and settlement in tidal currents. Limnology and Oceanography, 52(3):1156–1165, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Folkard:2005:HMP

[Fol05] Andrew M. Folkard. Hydrodynamics of model Posidonia oceanica patches in shallow water. Limnology and Oceanography, 50 (5):1592–1600, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Fourqurean:2002:BRH

[Fou02] James W. Fourqurean. Book review: Hemminga, M. A., and C. M. Duarte. 2000. Seagrass ecology. Cambridge University Press. xi + 298 p. US$80. ISBN 0-521-66184-6. Limnology and Oceanography, 47(2):611, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Franke:2006:WTS

[FPPH06] Ulrich Franke, Lubos Polerecky, Elimar Precht, and Markus Huettel. Wave tank study of particulate organic matter degradation in permeable sediments. Limnology and Oceanography, 51(2):1084–1096, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Foster:2009:SFN

[FPZ09] R. A. Foster, A. Paytan, and J. P. Zehr. Seasonality of N_{2} fixation and nifH gene diversity in the Gulf of Aqaba (Red Sea). Limnology and Oceanography, 54(1):219–233, January 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Finkel:2007:PDC

[FQC^{+}07] Z. V. Finkel, A. S. Quigg, R. K. Chiampi, O. E. Schofield, and P. G. Falkowski. Phylogenetic diversity in cadmium : phosphorus ratio regulation by marine phytoplankton. Limnology and Oceanography, 52(3):1131–1138, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Findlay:2001:ELU

[FQH^{+}01] Stuart Findlay, John M. Quinn, Christopher W. Hickey, Greg Burrell, and Malcolm Downes. Effects of land use and riparian flowpath on delivery of dissolved organic carbon to streams. Limnology and Oceanography, 46(2):345–355, March 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Finkel:2006:IES

[FQR^{+}06] Z. V. Finkel, A. Quigg, J. A. Raven, J. R. Reinfelder, O. E. Schofield, and P. G. Falkowski. Irradiance and the elemental stoichiometry of marine phytoplankton. Limnology and Oceanography, 52(6):2690–2701, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Franks:2001:TAA

[Fra01] Peter J. S. Franks. Turbulence avoidance: an alternate explanation of turbulence-enhanced ingestion rates in the field. Limnology and Oceanography, 46(4):959–963, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Freeland:2001:GCW

[FRO01] Joanna R. Freeland, Vivien K. Rimmer, and Beth Okamura. Genetic changes within freshwater bryozoan populations suggest temporal gene flow from statoblast banks. Limnology and Oceanography, 46(5):1121–1129, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Freeland:2004:ERP

[FRO04] Joanna R. Freeland, Vivien K. Rimmer, and Beth Okamura. Evidence for a residual postglacial founder effect in a highly dispersive freshwater invertebrate. Limnology and Oceanography, 49(3):879–883, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Falter:2000:HCP

[FS00] James L. Falter and Francis J. Sansone. Hydraulic control of pore water geochemistry within the oxic-suboxic zone of a permeable sediment. Limnology and Oceanography, 45(3):550–557, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Fang:2009:SCE

[FS09a] Xing Fang and Heinz G. Stefan. Simulations of climate effects on water temperature, dissolved oxygen, and ice and snow covers in lakes of the contiguous U.S. under past and future climate scenarios. Limnology and Oceanography, 54(6 (part 2)):2359– 2370, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Foster:2009:EBI

[FS09b] Sophie E. Foster and Gary Sprule. Effects of the Bythotrephes invasion on native predatory invertebrates. Limnology and Oceanography, 54(3):757–769, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Fram:2008:PPU

[FSB^{+}08] Jonathan P. Fram, Hannah L. Stewart, Mark A. Brzezinski, Brian Gaylord, Daniel C. Reed, Susan L. Williams, and Sally MacIntyre. Physical pathways and utilization of nitrate supply to the giant kelp, Macrocystis pyrifera. Limnology and Oceanography, 53(5):1589–1603, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Franck:2005:CSD

[FSBB05] Valerie M. Franck, Geoffrey J. Smith, Kenneth W. Bruland, and Mark A. Brzezinski. Comparison of size-dependent carbon, nitrate, and silicic acid uptake rates in high- and low-iron waters. Limnology and Oceanography, 50(3):825–838, May 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Furukawa:2004:QME

[FSK^{+}04] Yoko Furukawa, April C. Smith, Joel E. Kostka, Janet Watkins, and Clark R. Alexander. Quantification of macrobenthic effects on diagenesis using a multicomponent inverse model in salt marsh sediments. Limnology and Oceanography, 49(6):2058– 2072, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Foster:2007:IAR

[FSM^{+}07] R. A. Foster, A. Subramaniam, C. Mahaffey, E. J. Carpenter, D. G. Capone, and J. P. Zehr. Influence of the Amazon River plume on distributions of free-living and symbiotic cyanobacteria in the western tropical north Atlantic Ocean. Limnology and Oceanography, 52(2):517–532, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Fukuda:2000:EWG

[FSS^{+}00] Rumi Fukuda, Yoshiki Sohrin, Nobue Saotome, Hideki Fukuda, Toshi Nagata, and Isao Koike. East–west gradient in ectoenzyme activities in the subarctic Pacific: Possible regulation by zinc. Limnology and Oceanography, 45(4):930–939, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Findlay:2003:MSR

[FSSH03] Stuart E. G. Findlay, Robert L. Sinsabaugh, William V. Sobczak, and M. Hoostal. Metabolic and structural response of hyporheic microbial communities to variations in supply of dissolved organic matter. Limnology and Oceanography, 48(4): 1608–1617, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Fischer:2002:DRU

[FSSP02] Helmut Fischer, Anke Sachse, Christian E. W. Steinberg, and Martin Pusch. Differential retention and utilization of dissolved organic carbon by bacteria in river sediments. Limnology and Oceanography, 47(6):1702–1711, November 2002. CODEN LI-OCAH. ISSN 0024-3590.

Forrester:2008:SLS

[FSSV08] Graham E. Forrester, Mark A. Steele, Jameal F. Samhouri, and Richard R. Vance. Settling larvae of a small coral-reef fish discriminate reef features at large, but not small, spatial scales. Limnology and Oceanography, 53(5):1956–1962, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Fries:2003:FOE

[FT03] J. Stephen Fries and John H. Trowbridge. Flume observations of enhanced fine-particle deposition to permeable sediment beds. Limnology and Oceanography, 48(2):802–812, March 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Fuller:2006:MAPa

[FTC^{+}06] Nicholas J. Fuller, Glen A. Tarran, Denise G. Cummings, E. Malcolm S. Woodward, Karen M. Orcutt, Marian Yallop, Florence Le Gall, and David J. Scanlan. Molecular analysis of photosynthetic picoeukaryote community structure along an Arabian Sea transect. Limnology and Oceanography, 51(6):2502– 2514, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Fuller:2006:MAPb

[FTY^{+}06] Nicholas J. Fuller, Glen A. Tarran, Marian Yallop, Karen M. Orcutt, and David J. Scanlan. Molecular analysis of picocyanobacterial community structure along an Arabian Sea transect reveals distinct spatial separation of lineages. Limnology and Oceanography, 51(6):2515–2526, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Frost:2002:MCG

[FUG02] Thomas Frost and Robert C. Upstill-Goddard. Meteorological controls of gas exchange at a small English lake. Limnology and Oceanography, 47(4):1165–1174, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Fellows:2001:WSM

[FVD01] Christine S. Fellows, Maurice H. Valett, and Clifford N. Dahm. Whole-stream metabolism in two montane streams: Contribution of the hyporheic zone. Limnology and Oceanography, 46(3): 523–531, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Floge:2007:VCC

[FW07] Sheri A. Floge and Mark L. Wells. Variation in colloidal chromophoric dissolved organic matter in the damariscotta estuary, Maine. Limnology and Oceanography, 52(1):32–45, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Fuller:2005:DCS

[FWM^{+}05] Nicholas J. Fuller, Nyree J. West, Dominique Marie, Marian Yallop, Tanya Rivlin, Anton F. Post, and David J. Scanlan. Dynamics of community structure and phosphate status of picocyanobacterial populations in the Gulf of Aqaba, Red Sea. Limnology and Oceanography, 55(3):363–375, January 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Frost:2004:SDO

[FXL04] Paul C. Frost, Marguerite A. Xenopoulos, and James H. Larson. The stoichiometry of dissolved organic carbon, nitrogen, and phosphorus release by a planktonic grazer, Daphnia. Limnology and Oceanography, 49(5):1802–1808, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Fyfe:2003:STV

[Fyf03] S. K. Fyfe. Spatial and temporal variation in spectral reflectance: Are seagrass species spectrally distinct? Limnology and Oceanography, 48(1 (part 2)):464–479, January 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Fu:2005:BBC

[FZL^{+}05] Fei-Xue Fu, Yaohong Zhang, Karine Leblanc, Sergio A. Sañudo-Wilhelmy, and David A. Hutchins. The biological and biogeochemical consequences of phosphate scavenging onto phytoplankton cell surfaces. Limnology and Oceanography, 50(5): 1459–1472, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Gerten:2000:CDC

[GA00] Dieter Gerten and Rita Adrian. Climate-driven changes in spring plankton dynamics and the sensitivity of shallow polymictic lakes to the North Atlantic Oscillation. Limnology and Oceanography, 45(5):1058–1066, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Gerten:2001:DPN

[GA01] Dieter Gerten and Rita Adrian. Differences in the persistency of the North Atlantic Oscillation signal among lakes. Limnology and Oceanography, 49(1):448–455, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Gaiser:2000:DAE

[Gai00] Evelyn E. Gaiser. The diatoms: Applications for the environmental and earth sciences. Limnology and Oceanography, 45(3): 751, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Gonzale:2001:MBP

[GAM^{+}01] Natalia Gonzále, Ricardo Anadón, Beatriz Mouriño, Emilio Fernández, Bablu Sinha, José Escáne, and Demetrio de Armas. The metabolic balance of the planktonic community in the North Atlantic Subtropical Gyre: The role of mesoscale instabilities. Limnology and Oceanography, 46(4):946–952, June 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Grossart:2006:TEC

[GAPR06] Hans-Peter Grossart, Martin Allgaier, Uta Passow, and Ulf Riebesell. Testing the effect of CO_{2} concentration on the dynamics of marine heterotrophic bacterioplankton. Limnology and Oceanography, 51(1):1–11, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Gaylord:2000:BIS

[Gay00] Brian Gaylord. Biological implications of surf-zone flow complexity. Limnology and Oceanography, 45(1):174–188, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Goldsteina:2009:BEO

[GB09] Jason S. Goldsteina and Mark J. Butler I.V. Behavioral enhancement of onshore transport by postlarval Caribbean spiny lobster (Panulirus argus). Limnology and Oceanography, 54(5): 1669–1678, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Guichard:2000:HRR

[GBA00] Frédéric Guichard, Edwin Bourget, and Jean-Paul Agnard. High-resolution remote sensing of intertidal ecosystems: a lowcost technique to link scale-dependent patterns and processes. Limnology and Oceanography, 45(2):328–338, March 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Glud:2007:EDB

[GBFJ07] Ronnie N. Glud, Peter Berg, Henrik Fossing, and Bo B. Jørgensen. Effect of the diffusive boundary layer on benthic mineralization and O_{2} distribution: a theoretical model analysis. Limnology and Oceanography, 52(2):547–557, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Grafe:2002:USM

[GBH^{+}02] Holger Gräfe, Bertram Boehrer, Niels Hoppe, Stefan C. Müller, and Peter Hauptmann. Ultrasonic in situ measurements of density, adiabatic compressibility, and stability frequency. Limnology and Oceanography, 47(4):1255–1260, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Gervasio:2004:GDG

[GBL^{+}04] Vivianaluxa Gervasio, David J. Berg, Brian K. Lang, Nathan L. Allan, and Sheldon I. Guttman. Genetic diversity in the Gammarus pecos species complex: Implications for conservation and regional biogeography in the chihuahuan desert. Limnology and Oceanography, 49(2):520–531, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Grossart:2004:TDD

[GBM^{+}04] Hans-Peter Grossart, Thorsten Brinkhoff, Torben Martens, Claus Duerselen, Gerd Liebezeit, and Meinhard Simon. Tidal dynamics of dissolved and particulate matter and bacteria in a tidal flat ecosystem in spring and fall. Limnology and Oceanography, 49(6):2212–2222, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Greve:2003:MOV

[GBP03] Tina Maria Greve, Jens Borum, and Ole Pedersen. Meristematic oxygen variability in eelgrass (Zostera marina). Limnology and Oceanography, 48(1):210–216, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Gasol:2002:TEI

[GCG^{+}02] Josep M. Gasol, Marta Comerma, Juan Carlos García, Joan Armengol, Emilio O. Casamayor, Petra Kojecká, and Karel Šimek. A transplant experiment to identify the factors controlling bacterial abundance, activity, production, and community composition in a eutrophic canyon-shaped reservoir. Limnology and Oceanography, 47(1):62–77, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Goulet:2005:EST

[GCG05] Tamar L. Goulet, Clayton B. Cook, and Denis Goulet. Effect of short-term exposure to elevated temperatures and light levels on photosynthesis of different host-symbiont combinations in the Aiptasia pallida/Symbiodinium symbiosis. Limnology and Oceanography, 50(5):1490–1498, September 2005. CODEN LI-OCAH. ISSN 0024-3590.

Gaiser:2006:PRE

[GCJ^{+}06] Evelyn E. Gaiser, Daniel L. Childers, Ronald D. Jones, Jennifer H. Richards, Leonard J. Scinto, and Joel C. Trexler. Periphyton responses to eutrophication in the Florida Everglades:

Cross-system patterns of structural and compositional change. Limnology and Oceanography, 51(1 (part 2)):617–630, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Goffredo:2009:ILC

[GCM^{+}09] Stefano Goffredo, Erik Caroselli, Guido Mattioli, Elettra Pignotti, Zvy Dubinsky, and Francesco Zaccanti. Inferred level of calcification decreases along an increasing temperature gradient in a Mediterranean endemic coral. Limnology and Oceanography, 54(3):930–937, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Grothues:2002:FLF

[GCP^{+}02] Thomas M. Grothues, Robert K. Cowen, Leonard J. Pietrafesa, Francesco Bignami, Georges L. Weatherly, and Charles N. Flagg. Flux of larval fish around Cape Hatteras. Limnology and Oceanography, 47(1):165–175, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Gu:2006:FCS

[GCS06] Binhe Gu, Andrew D. Chapman, and Claire L. Schelske. Factors controlling seasonal variations in stable isotope composition of particulate organic matter in a softwater eutrophic lake. Limnology and Oceanography, 52(6):2837–2848, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Gillooly:2000:LPS

[GD00a] James F. Gillooly and Stanley I. Dodson. Latitudinal patterns in the size distribution and seasonal dynamics of new world, freshwater cladocerans. Limnology and Oceanography, 45(1):22– 30, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Goldman:2000:GMB

[GD00b] Joel C. Goldman and Mark R. Dennett. Growth of marine bacteria in batch and continuous culture under carbon and nitrogen limitation. Limnology and Oceanography, 45(4):789–800, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Goldman:2001:RNU

[GD01a] Joel C. Goldman and Mark R. Dennett. Rapid nitrogen uptake by marine bacteria. Limnology and Oceanography, 46(5):1195– 1198, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Gordon:2001:LSN

[GD01b] Howard R. Gordon and Tao Du. Light scattering by nonspherical particles: Application to coccoliths detached from Emiliania huxleyi. Limnology and Oceanography, 46(6):1438–1454, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Gaiser:2009:MCO

[GDB^{+}09] Evelyn E. Gaiser, Nancy D. Deyrup, Roger W. Bachmann, Larry E. Battoe, and Hilary M. Swain. Multidecadal climate oscillations detected in a transparency record from a subtropical Florida lake. Limnology and Oceanography, 54(6):2228–2232, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Gallegos:2008:OCL

[GDCG08] C. L. Gallegos, R. J. Davies-Colley, and M. Gall. Optical closure in lakes with contrasting extremes of reflectance. Limnology and Oceanography, 53(5):2021–2034, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Gaylord:2003:MWF

[GDK03] Brian Gaylord, Mark W. Denny, and Mimi A. R. Koehl. Modulation of wave forces on kelp canopies by alongshore currents. Limnology and Oceanography, 48(2):860–871, March 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Gustafsson:2000:FSC

[GDL^{+}00] Örjan Gustafsson, Anders Düker, Jenny Larsson, Per Andersson, and Johan Ingri. Functional separation of colloids and gravitoids in surface waters based on differential settling velocity: Coupled cross-flow filtration — split flow thin-cell fractionation (CFF-SPLITT). Limnology and Oceanography, 45(8):1731– 1742, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Gacia:2002:CND

[GDM02] Esperança Gacia, Carlos M. Duarte, and Jack J. Middelburg. Carbon and nutrient deposition in a Mediterranean seagrass (Posidonia oceanica) meadow. Limnology and Oceanography, 47 (1):23–32, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Gallon:2002:MRF

[GEJ^{+}02] J. R. Gallon, A. M. Evans, D. A. Jones, P. Albertano, R. Congestri, B. Bergman, K. Gundersen, K. M. Orcutt, K. von Bröckel, P. Fritsche, M. Meyerhöfer, K. Nachtigall, U. Ohlendieck, S. te Lintel Hekkert, K. Sivonen, S. Repka, L. J. Stal, and M. Staal. Maximum rates of N_{2} fixation and primary production are out of phase in a developing cyanobacterial bloom in the Baltic Sea. Limnology and Oceanography, 47 (5):1514–1521, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Glud:2008:BRM

[GEP08] Ronnie N. Glud, Bradley D. Eyre, and Nicole Patten. Biogeochemical responses to mass coral spawning at the Great Barrier Reef: Effects on respiration and primary production. Limnology and Oceanography, 54(3):1014–1024, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Gorbunov:2002:PCH

[GF02] Maxim Y. Gorbunov and Paul G. Falkowski. Photoreceptors in the cnidarian hosts allow symbiotic corals to sense blue moonlight. Limnology and Oceanography, 47(1):309–315, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Guldberg:2002:UMS

[GFJ^{+}02] Lise Bonne Guldberg, Kai Finster, Niels O. G. Jørgensen, Mathias Middelboe, and Bente Aa. Lomstein. Utilization of marine sedimentary dissolved organic nitrogen by native anaerobic bacteria. Limnology and Oceanography, 47(6):1712–1722, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Gorbunov:2000:MPP

[GFK00] Maxim Y. Gorbunov, Paul G. Falkowski, and Zbigniew S. Kolber. Measurement of photosynthetic parameters in benthic organisms in situ using a SCUBA-based fast repetition rate fluorometer. Limnology and Oceanography, 45(1):242–245, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Graco:2001:MDM

[GFM^{+}01] Michelle Graco, Laura Farías, Verónica Molina, Dimitri Gutiérrez, and Lars Peter Nielsen. Massive developments of microbial mats following phytoplankton blooms in a naturally eutrophic bay: Implications for nitrogen cycling. Limnology and Oceanography, 46(4):821–832, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Godinot:2009:CPU

[GFPG09] C. Godinot, C. Ferrier-Pagés, and R. Grover. Control of phosphate uptake by zooxanthellae and host cells in the scleractinian coral Stylophora pistillata. Limnology and Oceanography, 54(5): 1627–1633, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Green:2004:SPC

[GGDC04] Mark A. Green, Jeanne D. Gulnick, Nathaniel Dowse, and Patience Chapman. Spatiotemporal patterns of carbon remineralization and bio-irrigation in sediments of Casco Bay Estuary, Gulf of Maine. Limnology and Oceanography, 49(2):396–407, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Gettel:2007:EGS

[GGH07] Gretchen M. Gettel, Anne E. Giblin, and Robert W. Howarth. The effects of grazing by the snail, Lymnaea elodes, on benthic N_{2} fixation and primary production in oligotrophic, Arctic lakes. Limnology and Oceanography, 52(6):2398–2409, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Gomez-Giraldo:2006:SSD

[GGIA06] Andrés Gómez-Giraldo, Jörg Imberger, and Jason P. Antenucci. Spatial structure of the dominant basin-scale internal waves in Lake Kinneret. Limnology and Oceanography, 51(1):229–246, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Gomez-Giraldo:2008:WSG

[GGIAY08] Andrés Gómez-Giraldo, Jörg Imberger, Jason P. Antenucci, and Peter S. Yeates. Wind-shear-generated high-frequency internal waves as precursors to mixing in a stratified lake. Limnology and Oceanography, 53(1):354–367, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Garrison:2003:MSG

[GGM03] G. H. Garrison, C. R. Glenn, and G. M. McMurtry. Measurement of submarine groundwater discharge in Kahana Bay, O’ahu, Hawai’i. Limnology and Oceanography, 48(2):920–928, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Gisselson:2001:VCN

[GGP01] Lars-Åke Gisselson, Edna Granéli, and Jan Pallon. Variation in cellular nutrient status within a population of Dinophysis norvegica (Dinophyceae) growing in situ: Single-cell elemental analysis by use of a nuclear microprobe. Limnology and Oceanography, 46(5):1237–1242, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Glud:2003:SDB

[GGRJ03] Ronnie N. Glud, Jens K. Gundersen, Hans Røy, and Bo Barker Jørgensen. Seasonal dynamics of benthic O_{2} uptake in a semienclosed bay: Importance of diffusion and faunal activity. Limnology and Oceanography, 48(4):1265–1276, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Greenaway:2006:ERD

[GGS06] Anthony M. Greenaway and Debbie-Ann Gordon-Smith. The effects of rainfall on the distribution of inorganic nitrogen and phosphorus in Discovery Bay, Jamaica. Limnology and Oceanography, 51(5):2206–2220, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Guildford:2000:TNT

[GH00] Stephanie J. Guildford and Robert E. Hecky. Total nitrogen, total phosphorus, and nutrient limitation in lakes and oceans: Is there a common relationship? Limnology and Oceanography, 45(6):1213–1223, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Gribsholt:2005:NPT

[GHB^{+}05] Britta Gribsholt, T. Henricus, S. Boschker, Eric Struyf, Maria Andersson, Anton Tramper, Loreto De Brabandere, Stefan van Damme, Natacha Brion, Patrick Meire, Frank Dehairs, Jack J. Middelburg, and Carlo H. R. Heip. Nitrogen processing in a tidal freshwater marsh: a whole-ecosystem ^{15}N labeling study. Limnology and Oceanography, 50(6):1945–1959, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Gyllstrom:2005:RCS

[GHJ^{+}05] M. Gyllström, L.-A. Hansson, E. Jeppesen, F. García Criado, E. Gross, K. Irvine, T. Kairesalo, R. Kornijow, M. R. Miracle, M. Nykänen, T. Nõges, S. Romo, D. Stephen, E. Van Donk, and B. Moss. The role of climate in shaping zooplankton communities of shallow lakes. Limnology and Oceanography, 50(6): 2008–2021, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Gihring:2009:IPD

[GHM^{+}09] Thomas M. Gihring, Mike Humphrys, Heath J. Mills, Markus Huette, and Joel E. Kostka. Identification of phytodetritusdegrading microbial communities in sublittoral Gulf of Mexico sands. Limnology and Oceanography, 54(4):1073–1083, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Gundersen:2002:ECP

[GHN^{+}02] Kjell Gundersen, Mikal Heldal, Svein Norland, Duncan A. Purdie, and Anthony H. Knap. Elemental C, N, and P cell content of individual bacteria collected at the Bermuda Atlantic Timeseries Study (BATS) site. Limnology and Oceanography, 47(5): 1525–1530, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Gillett:2005:SPD

[GHS05] David J. Gillett, A. Frederick Holland, and Denise M. Sanger. Secondary production of a dominant oligochaete (Monopylephorus rubroniveus) in the tidal creeks of South Carolina and its relation to ecosystem characteristics. Limnology and Oceanography, 50(4):566–577, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Gevao:2000:TCM

[GHTJ00] Bondi Gevao, John Hamilton-Taylor, and Kevin C. Jones. Towards a complete mass balance and model for PCBs and PAHs in a small rural lake, Cumbria U.K. Limnology and Oceanography, 45(4):881–894, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Goldshmid:2004:ACS

[GHWG04] Rotem Goldshmid, Roi Holzman, Daniel Weihs, and Amatzia Genin. Aeration of corals by sleep-swimming fish. Limnology and Oceanography, 49(5):1832–1839, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:2003:SCR

[Gil03] John J. Gilbert. Specificity of crowding response that induces sexuality in the rotifer Brachionus. Limnology and Oceanography, 48(4):1297–1303, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Green:2004:DMJ

[GJB^{+}04] Mark A. Green, Michael E. Jones, Carrie L. Boudreau, Richard L. Moore, and Benjamin A. Westman. Dissolution mortality of juvenile bivalves in coastal marine deposits. Limnology and Oceanography, 49(3):727–734, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Gray:2007:EOO

[GJRM07] Derek K. Gray, Thomas H. Johengen, David F. Reid, and Hugh J. MacIsaac. Efficacy of open-ocean ballast water exchange as a means of preventing invertebrate invasions between freshwater ports. Limnology and Oceanography, 52(6):2386–2397, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Grey:2001:SCI

[GJS01] Jonathan Grey, Roger I. Jones, and Darren Sleep. Seasonal changes in the importance of the source of organic matter to the diet of zooplankton in Loch Ness, as indicated by stable isotope analysis. Limnology and Oceanography, 46(3):505–513, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Glenn:2008:GOS

[GJT^{+}08] Scott Glenn, Clayton Jones, Michael Twardowski, Louis Bowers, John Kerfoot, Josh Kohut, Doug Webb, and Oscar Schofield. Glider observations of sediment resuspension in a Middle Atlantic Bight fall transition storm. Limnology and Oceanography, 53(5 (part 2)):2180–2196, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Gribsholt:2003:BMS

[GK03] Britta Gribsholt and Erik Kristensen. Benthic metabolism and sulfur cycling along an inundation gradient in a tidal Spartina anglica salt marsh. Limnology and Oceanography, 48(6):2151– 2162, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Grey:2004:HIV

[GKJ04] Jonathan Grey, Andrew Kelly, and Roger I. Jones. High intraspecific variability in carbon and nitrogen stable isotope ratios of lake chironomid larvae. Limnology and Oceanography, 49 (1):239–244, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Gorbunov:2001:PPS

[GKLF01] Maxim Y. Gorbunov, Zbigniew S. Kolber, Michael P. Lesser, and Paul G. Falkowski. Photosynthesis and photoprotection in symbiotic corals. Limnology and Oceanography, 46(1):75–85, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Golding:2002:EFU

[GKS^{+}02] George R. Golding, Carol A. Kelly, Richard Sparling, Peter C. Loewen, John W. M. Rudd, and Tamar Barkay. Evidence for facilitated uptake of Hg(II) by Vibrio anguillarum and Escherichia coli under anaerobic and aerobic conditions. Limnology and Oceanography, 47(4):967–975, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Gorokhova:2007:CAU

[GL07] Elena Gorokhova and Maiju Lehtiniemi. A combined approach to understand trophic interactions between Cercopagis pengoi (Cladocera: Onychopoda) and mysids in the Gulf of Finland. Limnology and Oceanography, 52(2):685–695, March 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Galand:2008:MCD

[GLP^{+}08] Pierre E. Galand, Connie Lovejoy, Jérémie Pouliot, Marie ève Garneau, and Warwick F. Vincent. Microbial community diversity and heterotrophic production in a coastal Arctic ecosystem: a stamukhi lake and its source waters. Limnology and Oceanography, 53(2):813–823, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Gorokhova:2009:MEO

[GLVFH09] Elena Gorokhova, Maiju Lehtiniemi, Satu Viitasalo-Frösen, and Steven H. D. Haddock. Molecular evidence for the occurrence of ctenophore Mertensia ovum in the northern Baltic Sea and implications for the status of the Mnemiopsis leidyi invasion. Limnology and Oceanography, 54(6):2025–2033, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Gachter:2003:WPR

[GM03a] René Gächter and Beat Müller. Why the phosphorus retention of lakes does not necessarily depend on the oxygen supply to their sediment surface. Limnology and Oceanography, 48(2):929–933, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Goldman:2003:ELM

[GM03b] Joel C. Goldman and Dennis, Jr. J. McGillicuddy. Effect of large marine diatoms growing at low light on episodic new production. Limnology and Oceanography, 48(3):1176–1182, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Giessing:2004:OCD

[GM04a] Anders M. B. Giessing and Lawrence M. Mayer. Oxidative coupling during gut passage in marine deposit-feeding invertebrates. Limnology and Oceanography, 49(3):716–726, May 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Glud:2004:VBD

[GM04b] Ronnie N. Glud and Middelboe Mathias. Virus and bacteria dynamics of a coastal sediment: Implication for benthic carbon cycling. Limnology and Oceanography, 49(6):2073–2081, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Gardner:2006:NFD

[GMA^{+}06] Wayne S. Gardner, Mark J. McCarthy, Soonmo An, Dmitri Sobolev, Karen S. Sell, and David Brock. Nitrogen fixation and dissimilatory nitrate reduction to ammonium (DNRA) support nitrogen dynamics in Texas estuaries. Limnology and Oceanography, 51(1 (part 2)):558–568, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Grover:2003:NUS

[GMAFP03] Renaud Grover, Jean-François Maguer, Denis Allemand, and Christine Ferrier-Pagés. Nitrate uptake in the scleractinian coral Stylophora pistillata. Limnology and Oceanography, 48(6):2266– 2274, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Gramling:2003:CIM

[GMMW03] C. M. Gramling, D. C. McCorkle, A. E. Mulligan, and T. L. Woods. A carbon isotope method to quantify groundwater discharge at the land–sea interface. Limnology and Oceanography, 48(3):957–970, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Genin:2009:IBG

[GMR^{+}09] Amatzia Genin, Stephen G. Monismith, Matthew A. Reidenbach, Gitai Yahel, and Jeffrey R. Koseff. Intense benthic grazing of phytoplankton in a coral reef. Limnology and Oceanography, 54(3):938–951, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Gooseff:2004:DHT

[GMRD04] Michael N. Gooseff, Diane M. McKnight, Robert L. Runkel, and John H. Duff. Denitrification and hydrologic transient storage in a glacial meltwater stream, McMurdo Dry Valleys, Antarctica. Limnology and Oceanography, 49(5):1884–1895, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Grover:2002:UAS

[GMRVFP02] Renaud Grover, Jean-François Maguer, Stephanie Reynaud-Vaganay, and Christine Ferrier-Pagès. Uptake of ammonium by the scleractinian coral Stylophora pistillata: Effect of feeding, light, and ammonium concentrations. Limnology and Oceanography, 47(3):782–790, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Gibson:2000:RCB

[GN00] Jody R. Gibson and Raymond G. Najjar. The response of Chesapeake Bay salinity to climate-induced changes in streamflow. Limnology and Oceanography, 45(8):1764–1772, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Gurung:2000:SVD

[GNU00] Tek Bahadur Gurung, Masami Nakanishi, and Jotaro Urabe. Seasonal and vertical difference in negative and positive effects of grazers on heterotrophic bacteria in Lake Biwa. Limnology and Oceanography, 45(8):1689–1696, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Goericke:2002:BOA

[Goe02a] Ralf Goericke. Bacteriochlorophyll-a in the ocean: Is anoxygenic bacterial photosynthesis important? Limnology and Oceanography, 47(1):290–295, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Goericke:2002:TCP

[Goe02b] Ralf Goericke. Top-down control of phytoplankton biomass and community structure in the monsoonal Arabian Sea. Limnology and Oceanography, 47(5):1307–1323, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Goetze:2008:HMP

[Goe08] Erica Goetze. Heterospecific mating and partial prezygotic reproductive isolation in the planktonic marine copepods Centropages typicus and Centropages hamatus. Limnology and Oceanography, 53(2):433–445, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Goulden:2001:BRK

[Gou01] Clyde Goulden. Book review: Kawanabe, Hiroya, George W. Coulter, and Anna C. Roosevelt [Eds.]. 1999. Ancient lakes: Their cultural and biological diversity. Kenobi Productions, Ghent, Belgium. ix + 340 p. US$76. ISBN 90-804341-2-4. Limnology and Oceanography, 46(3):1579–1580, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Grossart:2000:BPG

[GP00] Hans-Peter Grossart and Helle Ploug. Bacterial production and growth efficiencies: Direct measurements on riverine aggregates. Limnology and Oceanography, 45(2):436–445, March 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Grossart:2001:MDO

[GP01] Hans-Peter Grossart and Helle Ploug. Microbial degradation of organic carbon and nitrogen on diatom aggregates. Limnology and Oceanography, 46(2):267–277, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Greene:2003:FSN

[GP03] Charles H. Greene and Andrew J. Pershing. The flip-side of the North Atlantic Oscillation and modal shifts in slope-water circulation patterns. Limnology and Oceanography, 48(1):319– 322, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Gerdeaux:2006:CWS

[GP06a] Daniel Gerdeaux and Marie-Elodie Perga. Changes in whitefish scales δ ^{13}C during eutrophication and reoligotrophication of subalpine lakes. Limnology and Oceanography, 51(1 (part 2)): 772–780, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Gucker:2006:RNU

[GP06b] Björn Gücker and Martin T. Pusch. Regulation of nutrient uptake in eutrophic lowland streams. Limnology and Oceanography, 51(3):1443–1453, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Ghadouani:2004:EMA

[GPAP^{+}04] Anas Ghadouani, Bernadette Pinel-Alloul, Klaus Plath, Geoffrey A. Codd, and Winfried Lampert. Effects of Microcystis aeruginosa and purified microcystin-LR on the feeding behavior of Daphnia pulicaria. Limnology and Oceanography, 49(3): 666–679, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Greenan:2008:OES

[GPHS08] B. J. W. Greenan, B. D. Petrie, W. G. Harrison, and P. M. Strain. The onset and evolution of a spring bloom on the Scotian Shelf. Limnology and Oceanography, 53(5):1759–1775, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Gervais:2001:EPL

[GR01] Frank Gervais and Ulf Riebesell. Effect of phosphorus limitation on elemental composition and stable carbon isotope fractionation in a marine diatom growing under different CO_{2} concentrations. Limnology and Oceanography, 46(3):497–504, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Gobler:2002:ING

[GRB02] Christopher J. Gobler, Marc J. Renaghan, and Nathaniel J. Buck. Impacts of nutrients and grazing mortality on the abundance of Aureococcus anophagefferens during a New York brown tide bloom. Limnology and Oceanography, 47(1):129–141, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Galman:2009:DDE

[GRB09] Veronika Gälman, Johan Rydberg, and Christian Bigler. Decadal diagenetic effects on δ ^{13}C and δ ^{15}N studied in varved lake sediment. Limnology and Oceanography, 54(3):917–924, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Garcia-Roger:2006:SMR

[GRCS06] Eduardo M. García-Roger, María José Carmona, and Manuel Serra. A simple model relating habitat features to a diapause egg bank. Limnology and Oceanography, 51(3):1542–1547, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Garcia-Roger:2009:CIS

[GRDCS09] Eduardo M. García-Roger, Natália Dias, María José Carmona, and Manuel Serra. Crossed induction of sex in sympatric congeneric rotifer populations. Limnology and Oceanography, 54(6): 1845–1854, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Galman:2008:CNL

[GRdL^{+}08] Veronika Gälman, Johan Rydberg, Sara Sjöstedt de Luna, Richard Bindler, and Ingemar Renberg. Carbon and nitrogen loss rates during aging of lake sediment: Changes over 27 years studied in varved lake sediment. Limnology and Oceanography, 54(3):1076–1082, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Greenwood:2009:SWD

[Gre09] J. Greenwood. Shallow water dissolution of settling calcite at Station ALOHA. Limnology and Oceanography, 54(5):1420– 1424, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Gervais:2002:CPP

[GRG02] Frank Gervais, Ulf Riebesell, and Maxim Y. Gorbunov. Changes in primary productivity and chlorophyll-a in response to iron fertilization in the southern polar frontal zone. Limnology and Oceanography, 47(5):1324–1335, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Gaylord:2007:SPF

[GRR^{+}07] Brian Gaylord, Johanna H. Rosman, Daniel C. Reed, Jeffrey R. Koseff, Jonathan Fram, Sally MacIntyre, Katie Arkema, Cameron McDonald, Mark A. Brzezinski, John L. Largier, Stephen G. Monismith, Peter T. Raimondi, and Brent Mardian. Spatial patterns of flow and their modification within and around a giant kelp forest. Limnology and Oceanography, 52(5): 1838–1852, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Gao:2009:OAE

[GRV^{+}09] Kunshan Gao, Zuoxi Ruan, Virginia E. Villafañe, Jean-Pierre Gattuso, and E. Walter Helbling. Ocean acidification exacerbates the effect of UV radiation on the calcifying phytoplankter Emiliania huxleyi. Limnology and Oceanography, 54(6):1855– 1862, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Gattuso:2000:CDS

[GRVF^{+}00] J.-P. Gattuso, S. Reynaud-Vaganay, P. Furla, S. Romaine-Lioud, J. Jaubert, I. Bourge, and M. Frankignoulle. Calcification does not stimulate photosynthesis in the zooxanthellate scleractinian coral Stylophora pistillata. Limnology and Oceanography, 45(1): 246–250, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:2004:RDF

[GS04] John J. Gilbert and Thomas Schröder. Rotifers from diapausing, fertilized eggs: Unique features and emergence. Limnology and Oceanography, 49(4 (part 2)):1341–1354, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Green:2005:DIN

[GSA^{+}05] Andy J. Green, Marta I. Sánchez, Francisco Amat, Jordi Figuerola, Francisco Hontoria, Olga Ruiz, and Francisco Hortas. Dispersal of invasive and native brine shrimps Artemia (Anostraca) via waterbirds. Limnology and Oceanography, 50 (4):737–742, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Goericke:2000:DSC

[GSB00] Ralf Goericke, Suzanne L. Strom, and Marnie A. Bell. Distribution and sources of cyclic pheophorbides in the marine environment. Limnology and Oceanography, 45(1):200–211, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Glud:2009:SMV

[GSB^{+}09] Ronnie N. Glud, Henrik Stahl, Peter Berg, Frank Wenzhöfer, Kazumasa Oguri, and Hiroshi Kitazato. In situ microscale variation in distribution and consumption of O_{2}: a case study from a deep ocean margin sediment (Sagami Bay, Japan). Limnology and Oceanography, 54(1):1–12, January 2009. CODEN LI-OCAH. ISSN 0024-3590.

Grzymski:2002:FPD

[GSFB02] Joe Grzymski, Oscar M. Schofield, Paul G. Falkowski, and Joan M. Bernhard. The function of plastids in the deep-sea benthic foraminifer, Nonionella stella. Limnology and Oceanography, 47(6):1569–1580, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:2005:STY

[GSG^{+}05] Denis Gilbert, Bjorn Sundby, Charles Gobeil, Alfonso Mucci, and Gilles-H. Tremblay. A seventy-two-year record of diminishing deep-water oxygen in the St. Lawrence estuary: The northwest Atlantic connection. Limnology and Oceanography, 50(5): 1654–1666, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Gu:2004:EES

[GSH04] Binhe Gu, Claire L. Schelske, and David A. Hodell. Extreme^{13}C enrichments in a shallow hypereutrophic lake: Implications for carbon cycling. Limnology and Oceanography, 49(4):1152–1159, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Guidi:2009:EPC

[GSJ^{+}09] Lionel Guidi, Lars Stemmann, George A. Jackson, Frédéric Ibanez, Hervé Claustre, Louis Legendre, Marc Picheral, and Gabriel Gorskya. Effects of phytoplankton community on production, size, and export of large aggregates: a world-ocean analysis. Limnology and Oceanography, 54(6):1951–1963, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Guidi:2007:VDA

[GSL^{+}07] Lionel Guidi, Lars Stemmann, Louis Legendre, Marc Picheral, Louis Prieur, and Gabriel Gorsky. Vertical distribution of aggregates (> 110 μ m) and mesoscale activity in the northeastern Atlantic: Effects on the deep vertical export of surface carbon. Limnology and Oceanography, 52(1):7–18, January 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Granger:2008:NOI

[GSLT08] Julie Granger, Daniel M. Sigman, Moritz F. Lehmann, and Philippe D. Tortell. Nitrogen and oxygen isotope fractionation during dissimilatory nitrate reduction by denitrifying bacteria. Limnology and Oceanography, 53(6):2533–2545, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Granger:2004:CNO

[GSNH04] Julie Granger, Daniel M. Sigman, Joseph A. Needoba, and Paul J. Harrison. Coupled nitrogen and oxygen isotope fractionation of nitrate during assimilation by cultures of marine phytoplankton. Limnology and Oceanography, 49(5):1763–1773, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Green:2003:CPO

[GSO03] Rebecca E. Green, Heidi M. Sosik, and Robert J. Olson. Contributions of phytoplankton and other particles to inherent optical properties in New England continental shelf waters. Limnology and Oceanography, 48(6):2377–2391, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Grout:2001:MCV

[GST^{+}01] H. Grout, R. Sempere, A. Thill, A. Calafat, L. Prieur, and M. Canals. Morphological and chemical variability of colloids in the almeria-oran front in the eastern Alboran Sea (SW Mediterranean Sea). Limnology and Oceanography, 46(6):1347–1357, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Gobler:2003:CCO

[GSW03] Christopher J. Gobler and Sergio A. Sañudo-Wilhelmy. Cycling of colloidal organic carbon and nitrogen during an estuarine phytoplankton bloom. Limnology and Oceanography, 48(6):2314– 2320, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Glud:2001:SIP

[GTK^{+}01] Ronnie N. Glud, Anders Tengberg, Michael Kühl, Per O. J. Hall, and Ingo Klimant. An in situ instrument for planar O_{2} optode measurements at benthic interfaces. Limnology and Oceanography, 47(4):2073–2080, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Glud:2009:NCD

[GTS^{+}09] Ronnie N. Glud, Bo Thamdrup, Henrik Stahl, Frank Wenzhoefer, Anni Glud, Hidetaka Nomaki, Kazumasa Oguri, Niels Peter Revsbech, and Hiroshi Kitazato. Nitrogen cycling in a deep ocean margin sediment (Sagami Bay, Japan). Limnology and Oceanography, 54(3):723–734, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Graham:2009:EWE

[GV09] Mark D. Graham and Rolf D. Vinebrooke. Extreme weather events alter planktonic communities in boreal lakes. Limnology and Oceanography, 54(6 (part 2)):2481–2492, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Grutters:2002:PAA

[GvRE^{+}02] Mark Grutters, Wim van Raaphorst, Eric Epping, Willem Helder, Jan W. de Leeuw, Daniel P. Glavin, and Jeffrey Bada. Preservation of amino acids from in situ-produced bacterial cell wall peptidoglycans in northeastern Atlantic continental margin sediments. Limnology and Oceanography, 47(5):1521–1524, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Graham:2006:CBF

[GVT06] Mark D. Graham, Rolf D. Vinebrooke, and Michael Turner. Coupling of boreal forests and lakes: Effects of conifer pollen on littoral communities. Limnology and Oceanography, 51(3): 1524–1529, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Gribsholt:2009:LTN

[GVT^{+}09] Britta Gribsholt, Bart Veuger, Anton Tramper, Jack J. Middelburg, and Henricus T. S. Boschker. Long-term ^{15}N-nitrogen retention in tidal freshwater marsh sediment: Elucidating the microbial contribution. Limnology and Oceanography, 54(1):13– 22, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Granger:2003:ANO

[GW03] Julie Granger and Bess B. Ward. Accumulation of nitrogen oxides in copper-limited cultures of denitrifying bacteria. Limnology and Oceanography, 48(1):313–318, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Gaudette:2006:SES

[GWH06] Julien Gaudette, Richard A. Wahle, and John H. Himmelman. Spawning events in small and large populations of the green sea urchin Strongylocentrotus droebachiensis as recorded using fertilization assays. Limnology and Oceanography, 51(3):1485– 1496, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Grad:2001:ZSR

[GWK01] Gabriella Grad, Craig E. Williamson, and Danielle M. Karapelou. Zooplankton survival and reproduction responses to damaging UV radiation: a test of reciprocity and photoenzymatic repair. Limnology and Oceanography, 46(3):584–591, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Goutx:2007:CDM

[GWL^{+}07] Madeleine Goutx, Stuart G. Wakeham, Cindy Lee, Marie Duflos, Catherine Guigue, Zhanfei Liu, Brivaëla Moriceau, Richard Sempére, Marc Tedetti, and Jianhong Xue. Composition and degradation of marine particles with different settling velocities in the northwestern Mediterranean Sea. Limnology and Oceanography, 52(4):1645–1664, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Green:2009:DDS

[GWR^{+}09] Mark A. Green, George G. Waldbusser, Shannon L. Reilly, Karla Emerson, and Scott O’Donnell. Death by dissolution: Sediment saturation state as a mortality factor for juvenile bivalves. Limnology and Oceanography, 54(4):1037–1047, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Goldthwait:2004:QMS

[GYBA04] Sarah Goldthwait, Jeannette Yen, Jason Brown, and Alice Alldredge. Quantification of marine snow fragmentation by swimming euphausiids. Limnology and Oceanography, 49(4):940–952, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Garland:2002:LDI

[GZL02] Elizabeth D. Garland, Cheryl Ann Zimmer, and Steven J. Lentz. Larval distributions in inner-shelf waters: The roles of winddriven cross-shelf currents and diel vertical migrations. Limnology and Oceanography, 47(3):803–817, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Gunter:2008:HAD

[GZY^{+}08] Jost Günter, Mikhail V. Zubkov, Evgeniy Yakushev, Matthias Labrenz, and Klaus Jürgens. High abundance and dark CO_{2} fixation of chemolithoautotrophic prokaryotes in anoxic waters of the Baltic Sea. Limnology and Oceanography, 53(1):14–22, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Hill:2001:RCF

[HA01] David M. Hill and Andrew C. Aplin. Role of colloids and fine particles in the transport of metals in rivers draining carbonate and silicate terrains. Limnology and Oceanography, 46(2):331– 344, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Hessen:2008:ECA

[HA08] Dag O. Hessen and Thomas R. Anderson. Excess carbon in aquatic organisms and ecosystems: Physiological, ecological, and evolutionary implications. Limnology and Oceanography, 53 (5):1685–1696, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Hadas:2009:SVN

[HAA09] Ora Hadas, Mark A. Altabet, and Rajesh Agnihotri. Seasonally varying nitrogen isotope biogeochemistry of particulate organic matter in Lake Kinneret, Israel. Limnology and Oceanography, 54(1):75–85, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Hessen:2003:PCS

[HABF03] Dag O. Hessen, Tom Andersen, Pål Brettum, and Bjørn A. Faafeng. Phytoplankton contribution to sestonic mass and elemental ratios in lakes: Implications for zooplankton nutrition.

Limnology and Oceanography, 48(4):1289–1296, May 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Haddock:2008:REP

[Had08] Steven H. D. Haddock. Reconsidering evidence for potential climate-related increases in jellyfish. Limnology and Oceanography, 53(6):2759–2762, November 2008. CODEN LIOCAH. ISSN 0024-3590. See reply [AE08].

Hulth:2002:PPF

[HAES02] Stefan Hulth, Robert C. Aller, Pia Engström, and Erik Selander. A pH plate fluorosensor (optode) for early diagenetic studies of marine sediments. Limnology and Oceanography, 47(1):212–220, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Huber:2008:PRC

[HAG08] Veronika Huber, Rita Adrian, and Dieter Gerten. Phytoplankton response to climate warming modified by trophic state. Limnology and Oceanography, 53(1):1–13, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Holzner:2009:EME

[HAHS^{+}09] C. P. Holzner, W. Aeschbach-Hertig, M. Simona, M. Veronesi, D. M. Imboden, and R. Kipfer. Exceptional mixing events in meromictic Lake Lugano (Switzerland/Italy), studied using environmental tracers. Limnology and Oceanography, 54(4):1113– 1124, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Hessen:2009:NDC

[HAL^{+}09] Dag O. Hessen, Tom Andersen, Søren Larsen, Brit Lisa Skjelkvåle, and Heleen A. de Wit. Nitrogen deposition, catchment productivity, and climate as determinants of lake stoichiometry. Limnology and Oceanography, 54(6 (part 2)):2520– 2528, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:2001:PWE

[Ham01] Stephen K. Hamilton. Prairie wetland ecology. The contribution of the Marsh Ecology Research Program. Limnology and Oceanography, 46(6):747–748, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Hamm:2002:IAS

[Ham02] Christian E. Hamm. Interactive aggregation and sedimentation of diatoms and clay-sized lithogenic material. Limnology and Oceanography, 47(6):1790–1795, November 2002. CODEN LI-OCAH. ISSN 0024-3590.

Hampton:2005:IND

[Ham05] Stephanie E. Hampton. Increased niche differentiation between two Conochilus species over 33 years of climate change and food web alteration. Limnology and Oceanography, 50(2):421–426, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Hart:2001:RES

[Har01] Rob C. Hart. Rapid estimation of in situ growth rates of Caridina nilotica (Crustacea: Decapoda) in Lake Victoria: Description and pilot application of a simple, field-compatible technique. Limnology and Oceanography, 46(6):692–698, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Harley:2002:LAI

[Har02] Christopher D. G. Harley. Light availability indirectly limits herbivore growth and abundance in a high rocky intertidal community during the winter. Limnology and Oceanography, 47(4): 1217–1222, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Hassett:2004:SDD

[Has04] R. Patrick Hassett. Supplementation of a diatom diet with cholesterol can enhance copepod egg-production rates. Limnology and Oceanography, 49(2):488–494, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Hassett:2006:PCL

[Has06] R. Patrick Hassett. Physiological characteristics of lipid-rich “fat” and lipid-poor “thin” morphotypes of individual Calanus finmarchicus C5 copepodites in nearshore Gulf of Maine. Limnology and Oceanography, 51(2):997–1003, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Hirst:2003:GMP

[HB03] A. G. Hirst and A. J. Bunker. Growth of marine planktonic copepods: Global rates and patterns in relation to chlorophylla, temperature, and body weight. Limnology and Oceanography, 49(4):1988–2010, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Hadwen:2005:FWR

[HB05] Wade L. Hadwen and Stuart E. Bunn. Food web responses to low-level nutrient and ^{15}N-tracer additions in the littoral zone of an oligotrophic dune lake. Limnology and Oceanography, 50 (4):1096–1105, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Hoffman:2006:IVS

[HB06] J. C. Hoffman and D. A. Bronk. Interannual variation in stable carbon and nitrogen isotope biogeochemistry of the Mattaponi River, Virginia. Limnology and Oceanography, 51(5):2319–2332, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Hopkinson:2008:III

[HB08] Brian M. Hopkinson and Katherine A. Barbeau. Interactive influences of iron and light limitation on phytoplankton at subsurface chlorophyll maxima in the eastern North Pacific. Limnology and Oceanography, 53(4):1303–1318, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Hall:2009:HCN

[HBAK09] Robert, Jr. O. Hall, Michelle A. Baker, Christopher D. Arp, and Benjamin J. Koch. Hydrologic control of nitrogen removal, storage, and export in a mountain stream. Limnology and Oceanography, 54(6):2128–2142, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Hanson:2003:LMR

[HBCK03] Paul C. Hanson, Darren L. Bade, Stephen R. Carpenter, and Timothy K. Kratz. Lake metabolism: Relationships with dissolved organic carbon and phosphorus. Limnology and Oceanography, 48(3):1112–1119, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Hershey:2006:SIS

[HBF^{+}06] A. E. Hershey, S. Beaty, K. Fortino, S. Kelly, M. Keyse, C. Luecke, W. J. O’Brien, and S. C. Whalen. Stable isotope signatures of benthic invertebrates in Arctic lakes indicate limited coupling to pelagic production. Limnology and Oceanography, 51(1):177–188, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Haltuch:2000:GIS

[HBG00] Melissa A. Haltuch, Paul Arthur Berkman, and David W. Garton. Geographic information system (GIS) analysis of ecosystem invasion: Exotic mussels in Lake Erie. Limnology and Oceanography, 45(8):1778–1787, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Hylander:2009:CUP

[HBG^{+}09] Samuel Hylander, Wiebke J. Boeing, Wilhelm Granéli, Jan Karlsson, Jessica von Einem, Kelly Gutseit, and Lars-Anders Hanssoa. Complementary UV protective compounds in zooplankton. Limnology and Oceanography, 54(6):1883–1893, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Hansen:2000:ZGG

[HBH00] Per Juel Hansen, Peter Koefoed BjØrnsen, and Benni Winding Hansen. Zooplankton grazing and growth: Scaling within the 2–2,000-μ m body size range. Limnology and Oceanography, 45 (8):1891, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Hanlon:2006:DEP

[HBH^{+}06] A. R. M. Hanlon, B. Bellinger, K. Haynes, G. Xiao, T. A. Hofmann, M. R. Gretz, A. S. Ball, A. M. Osborn, and G. J. C. Underwood. Dynamics of extracellular polymeric substance (EPS) production and loss in an estuarine, diatom-dominated, microalgal biofilm over a tidal emersion–immersion period. Limnology and Oceanography, 51(1):79–93, January 2006. CODEN LI-OCAH. ISSN 0024-3590.

Hall:2007:DOM

[HBH^{+}07] Per O. J. Hall, Jenny Brunnegård, Gustaf Hulthe, William R. Martin, Henrik Stahl, and Anders Tengberg. Dissolved organic matter in abyssal sediments: Core recovery artifacts. Limnology and Oceanography, 52(1):19–31, January 2007. CODEN LI-OCAH. ISSN 0024-3590.

Handoh:2001:URG

[HBI01] Itsuki C. Handoh, Grant R. Bigg, and Masamichi Inoue. Use of a reduced-gravity model to evaluate present and past primary productivity in the tropical open ocean. Limnology and Oceanography, 46(7):1632–1641, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Hall:2002:RNU

[HBL02] Robert, Jr. O. Hall, Emily S. Bernhardt, and Gene E. Likens. Relating nutrient uptake with transient storage in forested mountain streams. Limnology and Oceanography, 47(1):255–265, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Hovius:2006:ERC

[HBM06] Jonathan T. Hovius, Beatrix E. Beisner, and Kevin S. Mc-Cann. Epilimnetic rotifer community responses to Bythotrephes longimanus invasion in Canadian Shield lakes. Limnology and Oceanography, 51(2):1004–1012, March 2006. CODEN LI-OCAH. ISSN 0024-3590.

Herrman:2008:ANR

[HBM08] Kyle S. Herrman, Virginie Bouchard, and Richard H. Moore. An assessment of nitrogen removal from headwater streams in an agricultural watershed, northeast Ohio, U.S.A. Limnology and Oceanography, 53(6):2573–2582, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Huang:2000:PCM

[HBN^{+}00] Wenlin Huang, Xin Bu, Lisa Nguyen, Richard H. Gammon, and John L. Bullister. Production and consumption of methyl halides in a freshwater lake. Limnology and Oceanography, 45(7):1537– 1545, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Hiriart-Baer:2004:MUR

[HBS04] Véronique P. Hiriart-Baer and Ralph E. H. Smith. Models for ultraviolet radiation-dependent photoinhibition of Lake Erie phytoplankton. Limnology and Oceanography, 49(1):202–214, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Hiriart-Baer:2005:EUR

[HBS05] Véronique P. Hiriart-Baer and Ralph E. H. Smith. The effect of ultraviolet radiation on freshwater planktonic primary production: The role of recovery and mixing processes. Limnology and Oceanography, 50(5):1352–1361, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:2005:PAR

[HBTM05] Stephen K. Hamilton, Stuart E. Bunn, Martin C. Thoms, and Jonathan C. Marshall. Persistence of aquatic refugia between flow pulses in a dryland river system (Cooper Creek, Australia). Limnology and Oceanography, 50(3):743–754, May 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Havlicek:2001:PSS

[HC01] Tanya D. Havlicek and Stephen R. Carpenter. Pelagic species size distributions in lakes: Are they discontinuous? Limnology and Oceanography, 46(5):1021–1033, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Hare:2002:RRL

[HCC^{+}02] Jonathan A. Hare, James H. Churchill, Robert K. Cowen, Thomas J. Berger, Peter C. Cornillon, Paul Dragos, Scott M. Glenn, John J. Govoni, and Thomas N. Lee. Routes and rates of larval fish transport from the southeast to the northeast United States continental shelf. Limnology and Oceanography, 47(6): 1774–1789, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Hynes:2009:CBC

[HCD^{+}09] Annette M. Hynes, P. Dreux Chappell, Sonya T. Dyhrman, Scott C. Doney, and Eric A. Webbd. Cross-basin comparison of phosphorus stress and nitrogen fixation in Trichodesmium. Limnology and Oceanography, 54(5):1438–1448, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Havel:2000:RHI

[HCH00] John E. Havel, John K. Colbourne, and Paul D. N. Hebert. Reconstructing the history of intercontinental dispersal in Daphnia lumholtzi by use of genetic markers. Limnology and Oceanography, 45(6):1414–1419, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Herut:2002:RDS

[HCK02] Barak Herut, Robert Collier, and Michael D. Krom. The role of dust in supplying nitrogen and phosphorus to the southeast Mediterranean. Limnology and Oceanography, 47(3):870–878, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Hansell:2003:BDP

[HD03a] D. A. Hansell and H. W. Ducklow. Bacterioplankton distribution and production in the bathypelagic ocean: Directly coupled to particulate organic carbon export? Limnology and Oceanography, 48(1):150–156, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Helmuth:2003:PWE

[HD03b] Brian Helmuth and Mark W. Denny. Predicting wave exposure in the rocky intertidal zone: Do bigger waves always lead to larger forces? Limnology and Oceanography, 48(4):1338–1345, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Heier:2009:ESC

[HD09] Christopher R. Heier and Jeffry L. Dudycha. Ecological speciation in a cyclic parthenogen: Sexual capability of experimental hybrids between Daphnia pulex and Daphnia pulicaria. Limnology and Oceanography, 54(2):492–502, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Horner-Devine:2008:EIU

[HDFM08] Alexander R. Horner-Devine, Derek A. Fong, and Stephen G. Monismith. Evidence for the inherent unsteadiness of a river plume: Satellite observations of the Niagara River discharge. Limnology and Oceanography, 53(6):2731–2737, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Hansell:2004:MPN

[HDMB04] Dennis A. Hansell, Hugh W. Ducklow, Alison M. Macdonald, and Molly O-Neil Baringer. Metabolic poise in the North Atlantic Ocean diagnosed from organic matter transports. Limnology and Oceanography, 49(4):1084–1094, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Hester:2009:ISS

[HDP09] Erich T. Hester, Martin W. Doyle, and Geoffrey C. Poole. The influence of in-stream structures on summer water temperatures via induced hyporheic exchange. Limnology and Oceanography, 54(1):355–367, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Hall:2009:ETC

[HDSC09] Edward K. Hall, Andrew R. Dzialowski, Samuel M. Stoxen, and James B. Cotner. The effect of temperature on the coupling between phosphorus and growth in lacustrine bacterioplankton communities. Limnology and Oceanography, 54(3):880–889, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Hellings:2001:DIC

[HDVB01] L. Hellings, F. Dehairs, S. Van Damme, and W. Baeyens. Dissolved inorganic carbon in a highly polluted estuary (the Scheldt). Limnology and Oceanography, 46(6):1406–1414, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Heithaus:2009:PFI

[HDWDD09] Michael R. Heithaus, Bryan K. Delius, Aaron J. Wirsing, and Meagan M. Dunphy-Daly. Physical factors influencing the distribution of a top predator in a subtropical oligotrophic estuary. Limnology and Oceanography, 54(2):472–482, March 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Heino:2008:PFB

[Hei08] Jani Heino. Patterns of functional biodiversity and functionenvironment relationships in lake littoral macroinvertebrates. Limnology and Oceanography, 53(4):1446–1455, July 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Hesslein:2000:CFT

[Hes00] Raymond H. Hesslein. Chemical fate and transport in the environment. Limnology and Oceanography, 45(7):1678, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Hench:2004:TST

[HFC^{+}04] James L. Hench, Richard B. Forward, Jr., Sarah D. Carr, Daniel Rittschof, and Richard A. Luettich, Jr. Testing a selective tidalstream transport model: Observations of female blue crab (Callinectes sapidus) vertical migration during the spawning season. Limnology and Oceanography, 49(5):1857–1870, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Hondzo:2005:USD

[HFDO05] Miki Hondzo, Tom Feyaerts, Richard Donovan, and Ben L. O’Connor. Universal scaling of dissolved oxygen distribution at the sediment–water interface: a power law. Limnology and Oceanography, 50(5):1667–1676, September 2005. CODEN LI-OCAH. ISSN 0024-3590.

Hood:2007:SID

[HFE07] Eran Hood, Jason Fellman, and Rick T. Edwards. Salmon influences on dissolved organic matter in a coastal temperate brownwater stream: an application of fluorescence spectroscopy. Limnology and Oceanography, 52(4):1580–1587, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Hellweger:2003:GAR

[HFLD03] Ferdi L. Hellweger, Kevin J. Farley, Upmanu Lall, and Dominic M. Di Toro. Greedy algae reduce arsenate. Limnology and Oceanography, 48(6):2275–2288, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Hill:2008:DBA

[HFR08] W. R. Hill, S. E. Fanta, and B. J. Roberts. ^{13}C dynamics in benthic algae: Effects of light, phosphorus, and biomass development. Limnology and Oceanography, 53(4):1217–1226, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Hill:2009:QPL

[HFR09] Walter R. Hill, Shari E. Fanta, and Brian J. Roberts. Quantifying phosphorus and light effects in stream algae. Limnology and Oceanography, 54(1):368–380, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Hutchins:2007:CCT

[HFZ^{+}07] D. A. Hutchins, F.-X. Fu, Y. Zhang, M. E. Warner, Y. Feng, K. Portune, P. W. Bernhardt, and M. R. Mulholland. CO_{2} control of Trichodesmium N_{2} fixation, photosynthesis, growth rates, and elemental ratios: Implications for past, present, and future ocean biogeochemistry. Limnology and Oceanography, 52 (4):1293–1304, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Holzman:2003:ZNCa

[HG03a] Roi Holzman and Amatzia Genin. Zooplanktivory by a nocturnal coral-reef fish: Effects of light, flow, and prey density. Limnology and Oceanography, 48(4):1367–1375, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Holzman:2003:ZNCb

[HG03b] Roi Holzman and Amatzia Genin. Zooplanktivory by a nocturnal coral-reef fish: Effects of light, flow, and prey density. Limnology and Oceanography, 48(4):1376–1387, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Hampton:2000:DIE

[HGB00] Stephanie E. Hampton, John J. Gilbert, and Carolyn W. Burns. Direct and indirect effects of juvenile Buenoa macrotibialis (Hemiptera: Notonectidae) on the zooplankton of a shallow pond. Limnology and Oceanography, 45(4):1006–1012, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Habura:2008:BTE

[HGBB08] Andrea Habura, Susan T. Goldstein, Sarah Broderick, and Samuel S. Bowser. A bush, not a tree: The extraordinary diversity of cold-water basal foraminiferans extends to warm-water environments. Limnology and Oceanography, 53(4):1339–1351, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Hoegh-Guldberg:2005:CBF

[HGFS^{+}05] Ove Hoegh-Guldberg, Maoz Fine, William Skirving, Ron Johnstone, Sophie Dove, and Alan Strong. Coral bleaching following wintry weather. Limnology and Oceanography, 55(3):265–271, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Hjerne:2002:RFF

[HH02] Olle Hjerne and Sture Hansson. The role of fish and fisheries in Baltic Sea nutrient dynamics. Limnology and Oceanography, 47 (4):1023–1032, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Harley:2003:LRS

[HH03] Christopher D. G. Harley and Brian S. T. Helmuth. Local- and regional-scale effects of wave exposure, thermal stress, and absolute versus effective shore level on patterns of intertidal zonation. Limnology and Oceanography, 48(4):1498–1508, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Hulot:2004:AIB

[HH04] Florence D. Hulot and Jef Huisman. Allelopathic interactions between phytoplankton species: The roles of heterotrophic bacteria and mixing intensity. Limnology and Oceanography, 49(4 (part 2)):1424–1434, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Howard:2005:EST

[HH05] Shelby C. Howard and Brian T. Hentschel. Effects of short-term food variability on the plasticity of age and size at metamorphosis of porcelain crab larvae. Limnology and Oceanography, 50 (6):1960–1971, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Havskum:2005:ETS

[HHB05] Harry Havskum, Per Juel Hansen, and Elisa Berdalet. Effect of turbulence on sedimentation and net population growth of the dinoflagellate Ceratium tripos and interactions with its predator, Fragilidium subglobosum. Limnology and Oceanography, 50(5): 1543–1551, September 2005. CODEN LIOCAH. ISSN 0024-

Hassler:2009:EIB

[HHB^{+}09] Christel S. Hassler, Sonya M. Havens, George S. Bullerjahn, R. Michael L. McKay, and Michael R. Twissa. An evaluation of iron bioavailability and speciation in western Lake Superior with the use of combined physical, chemical, and biological assessment. Limnology and Oceanography, 54(3):987–1001, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Hickman:2009:DCA

[HHM^{+}09] Anna E. Hickman, Patrick M. Holligan, C. Mark Moore, Jonathan Sharples, Vladimir Krivtsov, and Matthew R. Palmer. Distribution and chromatic adaptation of phytoplankton within a shelf sea thermocline. Limnology and Oceanography, 54(2): 525–536, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Harita:2005:RTO

[HHS05] Yunosuke Harita, Toshitaka Hori, and Masahito Sugiyama. Release of trace oxyanions from littoral sediments and suspended particles induced by pH increase in the epilimnion of lakes. Limnology and Oceanography, 50(4):636–645, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Higgins:2001:ENF

[HHT01] Scott N. Higgins, Robert E. Hecky, and William D. Taylor. Epilithic nitrogen fixation in the rocky littoral zones of Lake Malawi, Africa. Limnology and Oceanography, 46(4):976–982, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Hutchins:2002:PIL

[HHW^{+}02] D. A. Hutchins, C. E. Hare, R. S. Weaver, Y. Zhang, G. F. Firme, G. R. DiTullio, M. B. Alm, S. F. Riseman, J. M. Maucher, M. E. Geesey, C. G. Trick, G. J. Smith, E. L. Rue, J. Conn, and K. W. Bruland. Phytoplankton iron limitation in the Humboldt Current and Peru Upwelling. Limnology and Oceanography, 47 (4):997–1011, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Hashimoto:2005:RBN

[HHY^{+}05] S. Hashimoto, N. Horimoto, Y. Yamaguchi, T. Ishimaru, and T. Saino. Relationship between net and gross primary production in the Sagami Bay, Japan. Limnology and Oceanography, 50 (6):1830–1835, November 2005. CODEN LIOCAH. ISSN 0024-

Hodges:2000:MBS

[HISW00] Ben R. Hodges, Jörg Imberger, Angelo Saggio, and K. B. Winters. Modeling basin-scale internal waves in a stratified lake. Limnology and Oceanography, 45(7):1603–1620, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Hillebrand:2001:EGN

[HK01] Helmut Hillebrand and Maria Kahlert. Effect of grazing and nutrient supply on periphyton biomass and nutrient stoichiometry in habitats of different productivity. Limnology and Oceanography, 46(8):1881–1898, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Hansson:2006:PSE

[HK06] Lars Johan Hansson and Thomas Kiørboe. Prey-specific encounter rates and handling efficiencies as causes of prey selectivity in ambush-feeding hydromedusae. Limnology and Oceanography, 53(2):1849–1858, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Holcombe:2001:DPW

[HKD01] Brook L. Holcombe, Richard G. Keil, and Allan H. Devol. Determination of pore-water dissolved organic carbon fluxes from Mexican margin sediments. Limnology and Oceanography, 46 (2):298–308, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Hays:2001:IVD

[HKF01] G. C. Hays, H. Kennedy, and B. W. Frost. Individual variability in diel vertical migration of a marine copepod: Why some individuals remain at depth when others migrate. Limnology and Oceanography, 47(4):2050–2054, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Hellweger:2008:ABM

[HKNG08] Ferdi L. Hellweger, Elena S. Kravchuk, Vladimir Novotny, and Michail I. Gladyshev. Agent-based modeling of the complex life cycle of a cyanobacterium (Anabaena) in a shallow reservoir. Limnology and Oceanography, 53(4):1227–1241, July 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Hart:2002:ISI

[HL02] E. Andrew Hart and James R. Lovvorn. Interpreting stable isotopes from macroinvertebrate foodwebs in saline wetlands. Limnology and Oceanography, 47(2):580–584, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Hentschel:2006:HMD

[HL06] Brian T. Hentschel and Amy A. Larson. Hydrodynamic mediation of density-dependent growth and adult-juvenile interactions of a spionid polychaete. Limnology and Oceanography, 51(2): 1031–1037, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Han:2007:CPS

[HLCG07] Seunghee Han, Ronald D. Lehman, Key-Young Choe, and Gary A. Gill. Chemical and physical speciation of mercury in Offatts Bayou: a seasonally anoxic bayou in Galveston Bay. Limnology and Oceanography, 52(4):1380–1392, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Herczeg:2003:MIR

[HLD^{+}03] A. L. Herczeg, F. W. Leaney, J. C. Dighton, S. Lamontagne, S. L. Schiff, A. L. Telfer, and M. C. English. A modern isotope record of changes in water and carbon budgets in a groundwaterfed lake: Blue Lake, South Australia. Limnology and Oceanography, 48(6):2093–2105, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:2008:WMB

[HLGI08] Andrew K. Hamilton, Connie Lovejoy, Pierre E. Galand, and R. Grant Ingram. Water masses and biogeography of picoeukaryote assemblages in a cold hydrographically complex system. Limnology and Oceanography, 53(3):922–935, May 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Hylander:2009:ZVM

[HLH09] Samuel Hylander, Niklas Larsson, and Lars-Anders Hansson. Zooplankton vertical migration and plasticity of pigmentation arising from simultaneous UV and predation threats. Limnology and Oceanography, 54(2):483–491, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Hwang:2005:LSG

[HLK05] Dong-Woon Hwang, Yong-Woo Lee, and Guebuem Kim. Large submarine groundwater discharge and benthic eutrophication in Bangdu Bay on volcanic Jeju Island, Korea. Limnology and Oceanography, 50(5):1393–1403, September 2005. CODEN LI-OCAH. ISSN 0024-3590.

Hannig:2007:SDA

[HLK^{+}07] M. Hannig, G. Lavik, M. M. M. Kuypers, D. Woebken, W. Martens-Habbena, and K. Jürgens. Shift from denitrification to anammox after inflow events in the central Baltic Sea. Limnology and Oceanography, 52(4):1336–1345, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Houel:2006:TSO

[HLL^{+}06] Stephane Houel, Patrick Louchouarn, Marc Lucotte, Rene Canuel, and Bassam Ghaleb. Translocation of soil organic matter following reservoir impoundment in boreal systems: Implications for in situ productivity. Limnology and Oceanography, 51 (3):1497–1513, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Hench:2008:ECE

[HLM08] James L. Hench, James J. Leichter, and Stephen G. Monismith. Episodic circulation and exchange in a wave-driven coral reef and lagoon system. Limnology and Oceanography, 53(6):2681– 2694, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Horppila:2003:MRE

[HLNM^{+}03] Jukka Horppila, Anne Liljendahl-Nurminen, Tommi Malinen, Mikko Salonen, Antti Tuomaala, Laura Uusitalo, and Mika Vinni. Mysis relicta in a eutrophic lake: Consequences of obligatory habitat shifts. Limnology and Oceanography, 48(4):1214– 1222, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Hofmann:2008:RIW

[HLP08] Hilmar Hofmann, Andreas Lorke, and Frank Peeters. The relative importance of wind and ship waves in the littoral zone of a large lake. Limnology and Oceanography, 53(1):368–380, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Hamersley:2007:AAO

[HLW^{+}07] M. Robert Hamersley, Gaute Lavik, Dagmar Woebken, Jayne E. Rattray, Phyllis Lam, Ellen C. Hopmans, Jaap S. Sinninghe Damsté, Siegfried Krüger, Michelle Graco, Dimitri Gutiérrez, and Marcel M. M. Kuypers. Anaerobic ammonium oxidation in the Peruvian oxygen minimum zone. Limnology and Oceanography, 52(3):923–933, May 2007. CODEN LIOCAH. ISSN 0024-

Hunt:2002:RPP

[HM02] Heather L. Hunt and Lauren S. Mullineaux. The roles of predation and postlarval transport in recruitment of the soft shell clam (Mya arenaria). Limnology and Oceanography, 47(1):151– 164, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Hedley:2003:RSM

[HM03a] John D. Hedley and Peter J. Mumby. A remote sensing method for resolving depth and subpixel composition of aquatic benthos. Limnology and Oceanography, 48(1 (part 2)):480–488, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Hembre:2003:SDP

[HM03b] Leif K. Hembre and Robert O. Megard. Seasonal and diel patchiness of a Daphnia population: an acoustic analysis. Limnology and Oceanography, 48(6):2221–2233, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Hill:2006:CCS

[HM06a] Walter R. Hill and R. Gerry Middleton. Changes in carbon stable isotope ratios during periphyton development. Limnology and Oceanography, 51(5):2360–2369, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Howarth:2006:NLN

[HM06b] Robert W. Howarth and Roxanne Marino. Nitrogen as the limiting nutrient for eutrophication in coastal marine ecosystems: Evolving views over three decades. Limnology and Oceanography, 51(1 (part 2)):364–376, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Hayes:2008:IOT

[HM08] Daniel R. Hayes and James Morison. Ice–ocean turbulent exchange in the Arctic summer measured by an autonomous underwater vehicle. Limnology and Oceanography, 53(5 (part 2)): 2287–2308, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Hill:2009:RRB

[HM09] Valerie L. Hill and Steven L. Manley. Release of reactive bromine and iodine from diatoms and its possible role in halogen transfer in polar and tropical oceans. Limnology and Oceanography, 54 (3):812–822, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Hu:2006:DTG

[HMB06] Haiying Hu, Steven E. Mylon, and Gaboury Benoit. Distribution of the thiols glutathione and 3-mercaptopropionic acid in Connecticut lakes. Limnology and Oceanography, 52(6):2763–2774, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Hulsmans:2007:DIM

[HMD^{+}07] Ann Hulsmans, Kelle Moreau, Luc De Meester, Bruce J. Riddoch, and Luc Brendonck. Direct and indirect measures of dispersal in the fairy shrimp Branchipodopsis wolfi indicate a small scale isolation-by-distance pattern. Limnology and Oceanography, 52(2):676–684, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Hockelmann:2004:OCC

[HMJ04] Claudia Höckelmann, Tom Moens, and Friedrich Jüttner. Odor compounds from cyanobacterial biofilms acting as attractants and repellents for free-living nematodes. Limnology and Oceanography, 49(5):1809–1819, September 2004. CODEN LI-OCAH. ISSN 0024-3590.

Hu:2002:AAC

[HMKZ02] Chuanmin Hu, Frank E. Muller-Karger, and Richard G. Zepp. Absorbance, absorption coefficient, and apparent quantum yield: a comment on common ambiguity in the use of these optical concepts. Limnology and Oceanography, 47(4):1261–1267, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Hughes:2006:ELS

[HMNL06] Claire Hughes, Gill Malin, Philip D. Nightingale, and Peter S. Liss. The effect of light stress on the release of volatile iodocarbons by three species of marine microalgae. Limnology and Oceanography, 52(6):2849–2854, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Hopkinson:2007:ILA

[HMR^{+}07] Brian M. Hopkinson, B. Greg Mitchell, Rick A. Reynolds, Haili Wang, Karen E. Selph, Christopher I. Measures, Christopher D. Hewes, Osmund Holm-Hansen, and Katherine A. Barbeau. Iron limitation across chlorophyll gradients in the southern Drake Passage: Phytoplankton responses to iron addition and photosynthetic indicators of iron stress. Limnology and Oceanography, 52(6):2540–2554, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Hamels:2004:ECH

[HMS^{+}04] Ilse Hamels, Heidi Mussche, Koen Sabbe, Koenraad Muylaert, and Wim Vyverman. Evidence for constant and highly specific active food selection by benthic ciliates in mixed diatoms assemblages. Limnology and Oceanography, 49(1):58–68, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Haese:2007:DCN

[HMS^{+}07] Ralf R. Haese, Emma J. Murray, Craig S. Smith, Jodie Smith, Lesley Clementson, and David T. Heggie. Diatoms control nutrient cycles in a temperate, wave-dominated estuary (southeast Australia). Limnology and Oceanography, 52(6):2686–2700, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Hadas:2006:VPS

[HMSI06] Eran Hadas, Dominique Marie, Muki Shpigel, and Micha Ilan. Virus predation by sponges is a new nutrient-flow pathway in coral reef food webs. Limnology and Oceanography, 51(3):1548– 1550, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Hedges:2000:OMB

[HMT^{+}00] John I. Hedges, Emilio Mayorga, Elizabeth Tsamakis, Michael E. McClain, Anthony Aufdenkampe, Paul Quay, Jeffrey E. Richey, Ron Benner, Steve Opsahl, Brenda Black, Tania Pimentel, Jorge Quintanilla, and Laurence Maurice. Organic matter in Bolivian tributaries of the Amazon River: a comparison to the lower mainstream. Limnology and Oceanography, 45(7):1449–1466, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Hughes:2008:PVI

[HMT^{+}08] C. Hughes, G. Malin, C. M. Turley, B. J. Keely, and P. D. Nightingale. The production of volatile iodocarbons by biogenic marine aggregates. Limnology and Oceanography, 57(3): 867–872, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Hansson:2007:CFP

[HNB^{+}07] Lars-Anders Hansson, Alice Nicolle, Jakob Brodersen, Pia Romare, P. Anders Nilsson, Christer Brönmark, and Christian Skov. Consequences of fish predation, migration, and juvenile ontogeny on zooplankton spring dynamics. Limnology and Oceanography, 52(2):696–706, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Hewson:2001:VLP

[HOFD01] Ian Hewson, Judith M. O’Neil, Jed A. Fuhrman, and William C. Dennison. Virus-like particle distribution and abundance in sediments and overlying waters along eutrophication gradients in two subtropical estuaries. Limnology and Oceanography, 46(7): 1734–1746, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Holfeld:2000:ISC

[Hol00] Harald Holfeld. Infection of the single-celled diatom Stephanodiscus alpinus by the chytrid Zygorhizidium: Parasite distribution within host population, changes in host cell size, and host–parasite size relationship. Limnology and Oceanography, 45(6):1440–1444, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Howarth:2000:ERM

[How00] Robert W. Howarth. Estuary restoration and maintenance: The National Estuary Program. Limnology and Oceanography, 45(8): 1889, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Helfrich:2003:APP

[HP03] Karl R. Helfrich and Jesús Pineda. Accumulation of particles in propagating fronts. Limnology and Oceanography, 48(4):1509– 1520, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Hooff:2006:CBI

[HP06] Rian C. Hooff and William T. Peterson. Copepod biodiversity as an indicator of changes in ocean and climate conditions of the northern California current ecosystem. Limnology and Oceanography, 51(6):2607–2620, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Hofer:2002:RDW

[HPAH^{+}02] Markus Hofer, Frank Peeters, Werner Aeschbach-Hertig, Matthias Brennwald, Johannes Holocher, David M. Livingstone, Vladimir Romanovski, and Rolf Kipfer. Rapid deep-water renewal in Lake Issyk-Kul (Kyrgyzstan) indicated by transient tracers. Limnology and Oceanography, 47(4):1210–1216, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Hee:2001:DOC

[HPAM01] Carol A. Hee, Tamara K. Pease, Marc J. Alperin, and Christopher S. Martens. Dissolved organic carbon production and consumption in anoxic marine sediments: a pulsed-tracer experiment. Limnology and Oceanography, 46(8):1908–1920, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Hope:2001:CDM

[HPBD01] Diane Hope, Sheila M. Palmer, Michael F. Billett, and Julian J. C. Dawson. Carbon dioxide and methane evasion from a temperate peatland stream. Limnology and Oceanography, 46 (4):847–857, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Hadas:2001:HCP

[HPE01] Ora Hadas, Riki Pinkas, and Jonathan Erez. High chemoautotrophic primary production in Lake Kinneret, Israel: a neglected link in the carbon cycle of the lake. Limnology and Oceanography, 46(8):1968–1976, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Hunt:2000:FEE

[HPHT00] A. P. Hunt, J. D. Parry, and J. Hamilton-Taylor. Further evidence of elemental composition as an indicator of the bioavailability of humic substances to bacteria. Limnology and Oceanography, 45(1):237–241, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Hoffmann:2006:DRS

[HPL^{+}06] Linn J. Hoffmann, Ilka Peeken, Karin Lochte, Philipp Assmy, and Marcel Veldhuis. Different reactions of Southern Ocean phytoplankton size classes to iron fertilization. Limnology and Oceanography, 51(3):1217–1229, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Hannides:2009:QZT

[HPLG09] Cecelia C. S. Hannides, Brian N. Popp, Michael R. Landry, and Brittany S. Graham. Quantification of zooplankton trophic position in the North Pacific Subtropical Gyre using stable nitrogen isotopes. Limnology and Oceanography, 54(1):50–61, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Hewson:2009:MPM

[HPT^{+}09] Ian Hewson, Ryan W. Paerl, H. James Tripp, Jonathan P. Zehr, and David M. Karl. Metagenomic potential of microbial assemblages in the surface waters of the central Pacific Ocean tracks variability in oceanic habitat. Limnology and Oceanography, 54 (6):1981–1994, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Heck:2000:ENE

[HPV^{+}00] K. L. Heck, Jr., J. R. Pennock, J. F. Valentine, L. D. Coen, and S. A. Sklenar. Effects of nutrient enrichment and small predator density on seagrass ecosystems: an experimental assessment. Limnology and Oceanography, 45(5):1041–1057, July 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Hessen:2000:URL

[HR00a] Dag O. Hessen and Nina Alstad Rukke. UV radiation and low calcium as mutual stressors for Daphnia. Limnology and Oceanography, 45(8):1834–1838, December 2000. CODEN LI-OCAH. ISSN 0024-3590.

Huettel:2000:APT

[HR00b] Markus Huettel and Antje Rusch. Advective particle transport into permeable sediments — evidence from experiments in an intertidal sandflat. Limnology and Oceanography, 45(3):525–533, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Huettel:2000:TDP

[HR00c] Markus Huettel and Antje Rusch. Transport and degradation of phytoplankton in permeable sediment. Limnology and Oceanography, 45(3):534–549, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Humphries:2003:EID

[HR03] Stuart Humphries and Graeme D. Ruxton. Estimation of intergenerational drift dispersal distances and mortality risk for aquatic macroinvertebrates. Limnology and Oceanography, 48 (6):2117–2124, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Hale:2007:IRO

[HR07] Michelle S. Hale and Richard B. Rivkin. Interpreting the results of oceanic mesoscale enrichment experiments: Caveats and lessons from limnology and coastal ecology. Limnology and Oceanography, 52(2):912–916, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Hebert:2007:LLT

[HRA07] Paul D. N. Hebert, Chad L. Rowe, and Sarah J. Adamowicz. Life at low temperatures: a novel breeding-system adjustment in a polar cladoceran. Limnology and Oceanography, 52(6):2507– 2518, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Heil:2007:NQD

[HRGM07] Cynthia A. Heil, Marta Revilla, Patricia M. Glibert, and Sue Murasko. Nutrient quality drives differential phytoplankton community composition on the southwest Florida shelf. Limnology and Oceanography, 52(3):1067–1078, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Hupfer:2004:ODP

[HRS04] Michael Hupfer, Bernhard Ruübe, and Peter Schmieder. Origin and diagenesis of polyphosphate in lake sediments: a ^{31}P-NMR study. Limnology and Oceanography, 49(1):1–10, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Havel:2004:MES

[HS04] John E. Havel and Jonathan B. Shurin. Mechanisms, effects, and scales of dispersal in freshwater zooplankton. Limnology and Oceanography, 49(4 (part 2)):1229–1238, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Hart:2000:SDL

[HSB00] Deborah R. Hart, Lewi Stone, and Tom Berman. Seasonal dynamics of the Lake Kinneret food web: The importance of the microbial loop. Limnology and Oceanography, 45(2):350–361, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Humborg:2004:NVB

[HSB^{+}04] Christoph Humborg, Erik Smedberg, Sven Blomqvist, Carl-Magnus Mörth, Jenni Brink, Lars Rahm, Åsa Danielsson, and Jörgen Sahlberg. Nutrient variations in boreal and subarctic Swedish rivers: Landscape control of land–sea fluxes. Limnology and Oceanography, 49(5):1871–1883, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Herfort:2006:CTD

[HSB^{+}06] Lydie Herfort, Stefan Schouten, Jan P. Boon, Martijn Woltering, Marianne Baas, Johan W. H. Weijers, and Jaap S. Sinninghe Damsté. Characterization of transport and deposition of terrestrial organic matter in the southern North Sea using the BIT index. Limnology and Oceanography, 51(5):2196–2205, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Hewson:2006:RHM

[HSCF06] Ian Hewson, Joshua A. Steele, Douglas G. Capone, and Jed A. Fuhrman. Remarkable heterogeneity in meso- and bathypelagic bacterioplankton assemblage composition. Limnology and Oceanography, 51(3):1274–1283, May 2006. CODEN LI-OCAH. ISSN 0024-3590.

Hamre:2008:CSO

[HSF^{+}08] Børge Hamre, Jakob J. Stamnes, Bøyvind Frette, Svein Rune Erga, and Knut Stamnes. Could stratospheric ozone depletion lead to enhanced aquatic primary production in the polar regions? Limnology and Oceanography, 53(1):332–338, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Heino:2005:RBS

[HSLV05] Jani Heino, Janne Soininen, Jyrki Lappalainen, and Risto Virtanen. The relationship between species richness and taxonomic distinctness in freshwater organisms. Limnology and Oceanography, 50(3):978–986, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Heldal:2003:ECS

[HSN^{+}03] M. Heldal, D. J. Scanlan, S. Norland, F. Thingstad, and N. H. Mann. Elemental composition of single cells of various strains of marine Prochlorococcus and Synechococcus using X-ray microanalysis. Limnology and Oceanography, 48(5):1732–1743, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Helms:2008:ASS

[HSR^{+}08] John R. Helms, Aron Stubbins, Jason D. Ritchie, Elizabeth C. Minor, David J. Kieber, and Kenneth Mopper. Absorption spectral slopes and slope ratios as indicators of molecular weight, source, and photobleaching of chromophoric dissolved organic matter. Limnology and Oceanography, 53(3):955–969, May 2008. CODEN LIOCAH. ISSN 0024-3590. See erratum [DdG09a].

Hampton:2006:CLW

[HSS06] Stephanie E. Hampton, Mark D. Scheuerell, and Daniel E. Schindler. Coalescence in the Lake Washington story: Interaction strengths in a planktonic food web. Limnology and Oceanography, 51(5):2042–2051, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Ho:2002:ACW

[HST^{+}02] Tung-Yuan Ho, Mary I. Scranton, Gordon T. Taylor, Ramon Varela, Robert C. Thunell, and Frank Muller-Karger. Acetate cycling in the water column of the Cariaco Basin: Seasonal and vertical variability and implication for carbon cycling. Limnology and Oceanography, 47(4):1119–1128, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Hall:2003:EMC

[HT03] Robert, Jr. O. Hall and Jennifer L. Tank. Ecosystem metabolism controls nitrogen uptake in streams in Grand Teton National Park, Wyoming. Limnology and Oceanography, 48(3):1120– 1128, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Hayes:2006:VDT

[HTAS06] Meredith K. Hayes, Gordon T. Taylor, Yrene Astor, and Mary I. Scranton. Vertical distributions of thiosulfate and sulfite in the Cariaco Basin. Limnology and Oceanography, 51(1):280–287, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Hoellein:2007:CST

[HTRM^{+}07] Timothy J. Hoellein, Jennifer L. Tank, Emma J. Rosi-Marshall, Sally A. Entrekin, and Gary A. Lamberti. Controls on spatial and temporal variation of nutrient uptake in three Michigan headwater streams. Limnology and Oceanography, 52(5):1964– 1977, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Havskum:2003:SLD

[HTS^{+}03] Harry Havskum, T. Frede Thingstad, Renate Scharek, Francesc Peters, Elisa Berdalet, M. Montserrat Sala, Miquel Alcaraz, Jan C. Bangsholt, Ulla Li Zweifel, Åke Hagström, Maite Perez, and John R. Dolan. Silicate and labile DOC interfere in structuring the microbial food web via algal–bacterial competition for mineral nutrients: Results of a mesocosm experiment. Limnology and Oceanography, 48(1):129–140, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Hall:2009:NRS

[HTS^{+}09] Robert O. Hall, Jr., Jennifer L. Tank, Daniel J. Sobota, Patrick J. Mulholland, Jonathan M. O’Brien, Walter K. Dodds, Jackson R. Webster, H. Maurice Valett, Geoffrey C. Poole, Bruce J. Peterson, Judy L. Meyer, William H. McDowell, Sherri L. Johnson, Stephen K. Hamilton, Nancy B. Grimm, Stanley V. Gregory, Clifford N. Dahm, Lee W. Cooper, Linda R. Ashkenas, Suzanne M. Thomas, Richard W. Sheibley, Jody D. Potter, B. R. Niederlehner, Laura T. Johnson, Ashley M. Helton, Chelsea M. Crenshaw, Amy J. Burgin, Melody J. Bernot, Jake J. Beaulieu, and Clay P. Arangob. Nitrate removal in stream ecosystems measured by ^{15}N addition experiments: Total uptake. Limnology and Oceanography, 54(3):653–665, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Hunter:2008:SDV

[HTWG08] P. D. Hunter, A. N. Tyler, N. J. Willby, and D. J. Gilvear. The spatial dynamics of vertical migration by Microcystis aeruginosa in a eutrophic shallow lake: a case study using high spatial resolution time-series airborne remote sensing. Limnology and Oceanography, 53(6):2391–2406, November 2008. CODEN LI-OCAH. ISSN 0024-3590.

Hays:2007:DDT

[HU07] Rebecca L. Hays and William J. Ullman. Direct determination of total and fresh groundwater discharge and nutrient loads from a sandy beachface at low tide (Cape Henlopen, Delaware). Limnology and Oceanography, 52(1):240–247, January 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Hanninen:2000:CFA

[HVH00] Jari Hänninen, Ilppo Vuorinen, and Pekka Hjelt. Climatic factors in the Atlantic control the oceanographic and ecological changes in the Baltic Sea. Limnology and Oceanography, 45(3): 703–710, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Holl:2007:TWG

[HVP^{+}07] Carolyn M. Holl, Tracy A. Villareal, Christopher D. Payne, Tonya D. Clayton, Cassandra Hart, and Joseph P. Montoya. Trichodesmium in the western Gulf of Mexico: ^{15}N_{2}-fixation and natural abundance stable isotopic evidence. Limnology and Oceanography, 52(5):2249–2259, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

He:2007:KPD

[HW07a] Xuejia He and Wen-Xiong Wang. Kinetics of phosphorus in Daphnia at different food concentrations and carbon : phosphorus ratios. Limnology and Oceanography, 52(1):395–406, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Hobbs:2007:CUP

[HW07b] William O. Hobbs and Alexander P. Wolfe. Caveats on the use of paleolimnology to infer Pacific salmon returns. Limnology and Oceanography, 52(5):2053–2061, September 2007. CODEN LIOCAH. ISSN 0024-3590.

He:2008:SRC

[HW08] Xuejia He and Wen-Xiong Wang. Stoichiometric regulation of carbon and phosphorus in P-deficient Daphnia magna. Limnology and Oceanography, 53(1):244–254, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Holzman:2009:HSC

[HW09] Roi Holzman and Peter C. Wainwright. How to surprise a copepod: Strike kinematics reduce hydrodynamic disturbance and increase stealth of suction-feeding fish. Limnology and Oceanography, 54(6):2201–2212, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Hebert:2003:PPD

[HWA03] Paul D. N. Hebert, Jonathan D. S. Witt, and Sarah J. Adamowicz. Phylogeographical patterning in Daphnia ambigua: Regional divergence and intercontinental cohesion. Limnology and Oceanography, 48(1):261–268, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Hung:2000:ELN

[HWL^{+}00] Chin-Chang Hung, George T. F. Wong, Kon-Kee Liu, Shiah Fuh-Kwo, and Gong Gwo-Ching. The effects of light and nitrate levels on the relationship between nitrate reductase activity and 15NO_{3}^{−} uptake: Field observations in the East China Sea. Limnology and Oceanography, 45(4):836–848, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Ho:2007:TMC

[HWYL07] Tung-Yuan Ho, Liang-Saw Wen, Chen-Feng You, and Der-Chuen Lee. The trace metal composition of size-fractionated plankton in the South China Sea: Biotic versus abiotic sources. Limnology and Oceanography, 52(5):1776–1788, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Hama:2001:PNA

[HY01] Takeo Hama and Katsumi Yanagi. Production and neutral aldose composition of dissolved carbohydrates excreted by natural marine phytoplankton populations. Limnology and Oceanography, 46(8):1945–1955, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Hama:2004:DMW

[HYH04] Takeo Hama, Katsumi Yanagi, and Junko Hama. Decrease in molecular weight of photosynthetic products of marine phytoplankton during early diagenesis. Limnology and Oceanography, 49(2):471–481, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Hambright:2000:PSD

[HZ00] K. David Hambright and Tamar Zohary. Phytoplankton species diversity control through competitive exclusion and physical disturbances. Limnology and Oceanography, 45(1):110–122, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Hambright:2007:MDC

[HZG07] K. David Hambright, Tamar Zohary, and Hans Güde. Microzooplankton dominate carbon flow and nutrient cycling in a warm subtropical freshwater lake. Limnology and Oceanography, 52 (3):1018–1025, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Ito:2005:SSN

[IB05] Yusai Ito and Alison Butler. Structure of synechobactins, new siderophores of the marine cyanobacterium Synechococcus sp. PCC 7002. Limnology and Oceanography, 50(6):1918–1923, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Irigoien:2000:FSE

[IHH^{+}00] X. Irigoien, R. N. Head, R. P. Harris, D. Cummings, D. Harbour, and B. Meyer-Harms. Feeding selectivity and egg production of Calanus helgolandicus in the English Channel. Limnology and Oceanography, 45(1):44–54, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Irigoien:2000:IDA

[IHHH00] X. Irigoien, R. P. Harris, R. N. Head, and D. Harbour. The influence of diatom abundance on the egg production rate of Calanus helgolandicus in the English Channel. Limnology and Oceanography, 45(6):1433–1439, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Ianson:2004:ENA

[IJA^{+}04] Debby Ianson, George A. Jackson, Martin V. Angel, Richard S. Lampitt, and Adrian B. Burd. Effect of net avoidance on estimates of diel vertical migration. Limnology and Oceanography, 49(6):2297–2303, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Interlandi:2003:DCR

[IKT03] Sebastian J. Interlandi, Susan S. Kilham, and Edward C. Theriot. Diatom–chemistry relationships in Yellowstone Lake (Wyoming) sediments: Implications for climatic and aquatic processes research. Limnology and Oceanography, 48(1):79–92, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Ishikawa:2004:LTC

[INU04] Toshiyuki Ishikawa, Tetsuya Narita, and Jotaro Urabe. Longterm changes in the abundance of Jesogammarus annandalei (Tattersall) in Lake Biwa. Limnology and Oceanography, 49(5): 1840–1847, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Idrisi:2004:MED

[IOG^{+}04] Nasseer Idrisi, M. Josefina Olascoaga, Zulema Garraffo, Donald B. Olson, and Sharon L. Smith. Mechanisms for emergence from diapause of Calanoides carinatus in the Somali Current. Limnology and Oceanography, 49(4 (part 2)):1262–1268, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Inglett:2006:IUM

[IR06] Patrick W. Inglett and K. Ramesh Reddy. Investigating the use of macrophyte stable C and N isotopic ratios as indicators of wetland eutrophication: Patterns in the P-affected Everglades. Limnology and Oceanography, 51(5):2380–2387, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Jansson:2003:AOC

[Jan03] Mats Jansson. Allochthonous organic carbon decreases pelagic energy mobilization in lakes. Limnology and Oceanography, 48 (4):1711–1716, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Job:2000:LSL

[JB00] Suresh D. Job and David R. Bellwood. Light sensitivity in larval fishes: Implications for vertical zonation in the pelagic zone.

Limnology and Oceanography, 45(2):362–371, March 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Jiri:2004:NCS

[JBH^{+}04] Kopáček Jiří, Martina Brzákova, Josef Hejzlar, Jiří Nedoma, Petr Porcal, and Jaroslav Vrba. Nutrient cycling in a strongly acidified mesotrophic lake. Limnology and Oceanography, 49(4): 1202–1213, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Jensen:2007:SAI

[JBM^{+}07] Olaf P. Jensen, Barbara J. Benson, John J. Magnuson, Virginia M. Card, Martyn N. Futter, Patricia A. Soranno, and Kenton M. Stewart. Spatial analysis of ice phenology trends across the Laurentian Great Lakes region during a recent warming period. Limnology and Oceanography, 52(5):2013–2026, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Jweda:2008:SLR

[JBvHS08] Jason Jweda, Mark Baskaran, Ed van Hees, and Linda Schweitzer. Short-lived radionuclides (^{7}Be and ^{210}Pb) as tracers of particle dynamics in a river system in southeast Michigan. Limnology and Oceanography, 53(5):1934–1944, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Jordan:2008:CPB

[JCBA08] Thomas E. Jordan, Jeffrey C. Cornwell, Walter R. Boynton, and Jon T. Anderson. Changes in phosphorus biogeochemistry along an estuarine salinity gradient: The iron conveyer belt. Limnology and Oceanography, 53(1):172–184, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Jassby:2002:APP

[JCC02] Alan D. Jassby, James E. Cloern, and Brian E. Cole. Annual primary production: Patterns and mechanisms of change in a nutrient-rich tidal ecosystem. Limnology and Oceanography, 47 (3):698–712, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Jones:2008:TIP

[JCK^{+}08] Stuart E. Jones, Chih-Yu Chiu, Timothy K. Kratz, Jiunn-Tzong Wu, Ashley Shade, and Katherine D. McMahon. Typhoons initiate predictable change in aquatic bacterial communities. Limnology and Oceanography, 53(4):1319–1326, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2008:CSC

[JCW08] Li-Qing Jiang, Wei-Jun Cai, and Yongchen Wang. A comparative study of carbon dioxide degassing in river- and marinedominated estuaries. Limnology and Oceanography, 53(6):2603– 2615, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Jager:2008:IWC

[JDS08] Christoph G. Jäger, Sebastian Diehl, and Gertraud M. Schmidt. Influence of water-column depth and mixing on phytoplankton biomass, community composition, and nutrients. Limnology and Oceanography, 53(6):2361–2373, November 2008. CODEN LI-OCAH. ISSN 0024-3590.

Janssen:2005:PWAa

[JFH^{+}05] F. Janssen, P. Faerber, M. Huettel, V. Meyer, and U. Witte. Pore-water advection and solute fluxes in permeable marine sediments (i): Calibration and performance of the novel benthic chamber system Sandy. Limnology and Oceanography, 50(3): 768–778, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Jolly:2009:PGH

[JGE^{+}09] M. T. Jolly, P. Guyard, C. Ellien, F. Gentil, F. Viard, E. Thiébaut, and D. Jollivet. Population genetics and hydrodynamic modeling of larval dispersal dissociate contemporary patterns of connectivity from historical expansion into European shelf seas in the polychaete Pectinaria koreni. Limnology and Oceanography, 54(6):2089–2106, November 2009. CODEN LIOCAH. ISSN 0024-3590.

John:2007:ZIF

[JGSB07] Seth G. John, Robert W. Geis, Mak A. Saito, and Edward A. Boyle. Zinc isotope fractionation during high-affinity and lowaffinity zinc transport by the marine diatom Thalassiosira oceanica. Limnology and Oceanography, 52(6):2710–2714, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Jewson:2008:RSE

[JGZ^{+}08] David H. Jewson, Nick G. Granin, Andre A. Zhdanov, Ludmila A. Gorbunova, Nina A. Bondarenko, and Ruslan Yu. Gnatovsky. Resting stages and ecology of the planktonic diatom Aulacoseira skvortzowii in Lake Baikal. Limnology and Oceanography, 54(3):1125–1136, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Ju:2006:RLD

[JHGGP06] Se-Jong Ju, H. Rodger Harvey, Jaime Gómez-Gutiérrez, and William T. Peterson. The role of lipids during embryonic development of the euphausiids Euphausia pacifica and Thysanoessa spinifera. Limnology and Oceanography, 51(5):2398– 2408, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Janssen:2005:PWAb

[JHW05] F. Janssen, M. Huettel, and U. Witte. Pore-water advection and solute fluxes in permeable marine sediments (II): Benthic respiration at three sandy sites with different permeabilities (German Bight, North Sea). Limnology and Oceanography, 50(3):779–792, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Jaeschke:2007:PLL

[JHW^{+}07] Andrea Jaeschke, Ellen C. Hopmans, Stuart G. Wakeham, Stefan Schouten, and Jaap S. Sinninghe Damsté. The presence of ladderane lipids in the oxygen minimum zone of the Arabian Sea indicates nitrogen loss through anammox. Limnology and Oceanography, 52(2):780–786, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Jonsson:2004:ASP

[JJP04] Per R. Jonsson, Mona Johansson, and Richard W. Pierce. Attachment to suspended particles may improve foraging and reduce predation risk for tintinnid ciliates. Limnology and Oceanography, 49(6):1907–1914, November 2004. CODEN LI-OCAH. ISSN 0024-3590.

Jeppesen:2003:DRP

[JJS^{+}03] Erik Jeppesen, Jens Peder Jensen, Martin Søndergaard, Kjeld Sandby Hansen, Poul Hald Møller, Helle Utoft Rasmussen, Vibeke Norby, and Søren E. Larsen. Does resuspension prevent a shift to a clear state in shallow lakes during reoligotrophication? Limnology and Oceanography, 49(4):1913–1919, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Jang:2007:CMP

[JJT07] Min-Ho Jang, Jong-Mun Jung, and Noriko Takamura. Changes in microcystin production in cyanobacteria exposed to zooplankton at different population densities and infochemical concentrations. Limnology and Oceanography, 52(4):1454–1466, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Jackson:2008:MPC

[JK08] George A. Jackson and Thomas Kiørboe. Maximum phytoplankton concentrations in the sea. Limnology and Oceanography, 53 (1):395–399, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Jensen:2008:RRA

[JKGT08] Marlene M. Jensen, Marcel M. M. Kuypers, Lavik Gaute, and Bo Thamdrup. Rates and regulation of anaerobic ammonium oxidation and denitrification in the Black Sea. Limnology and Oceanography, 53(1):23–36, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Jimenez:2008:HBT

[JKLR08] Isabel M. Jimenez, Michael Kühl, A. W. D. Larkum, and P. J. Ralph. Heat budget and thermal microenvironment of shallowwater corals: Do massive corals get warmer than branching corals? Limnology and Oceanography, 53(5):1548–1561, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Jarnagin:2004:ZLC

[JKS04] S. Taylor Jarnagin, W. Charles Kerfoot, and Brandon K. Swan. Zooplankton life cycles: Direct documentation of pelagic births and deaths relative to diapausing egg production. Limnology and Oceanography, 49(4 (part 2)):1317–1332, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Jamison:2008:IOR

[JL08] Jennifer L. Jamison and Howard R. Lasker. Identification of octocoral recruits using microsatellite primers: Relationships between recruitment and adult distribution of Pseudopterogorgia spp. Limnology and Oceanography, 53(5):1963–1972, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Jankowski:2006:CEH

[JLB^{+}06] Thomas Jankowski, David M. Livingstone, Heinrich Bührer, Richard Forster, and Pius Niederhauser. Consequences of the 2003 European heat wave for lake temperature profiles, thermal stability, and hypolimnetic oxygen depletion: Implications for a warmer world. Limnology and Oceanography, 51(2):815–819, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Jensen:2001:DFS

[JLH01] Knut Helge Jensen, Petter Larsson, and Göran Högstedt. Detecting food search in Daphnia in the field. Limnology and Oceanography, 46(5):1013–1020, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Jones:2008:PDF

[JLP^{+}08] Nicole L. Jones, Ryan J. Lowe, Geno Pawlak, Derek A. Fong, and Stephen G. Monismith. Plume dispersion on a fringing coral reef system. Limnology and Oceanography, 53(5 (part 2)):2273– 2286, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Jompa:2002:ENH

[JM02] Jamaluddin Jompa and Laurence J. McCook. The effects of nutrients and herbivory on competition between a hard coral (Porites cylindrica) and a brown alga (Lobophora variegata). Limnology and Oceanography, 47(2):527–534, March 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Johnson:2008:TMP

[JM08] Thomas C. Johnson and I. N. McCave. Transport mechanism and paleoclimatic significance of terrigenous silt deposited in varved sediments of an African rift lake. Limnology and Oceanography, 53(5):1622–1632, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Jonsson:2001:WLM

[JMBJ01] Anders Jonsson, Markus Meili, Ann-Kristin Bergström, and Mats Jansson. Whole-lake mineralization of allochthonous and autochthonous organic carbon in a large humic lake (Örträsket, N. Sweden). Limnology and Oceanography, 46(7):1691–1700, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2008:MSV

[JN08] Kenneth S. Johnson and Joseph A. Needoba. Mapping the spatial variability of plankton metabolism using nitrate and oxygen sensors on an autonomous underwater vehicle. Limnology and Oceanography, 53(5 (part 2)):2237–2250, September 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Jensen:2009:PRC

[JNKdV09] Henning S. Jensen, Ole I. Nielsen, Marguerite S. Koch, and Inmaculada de Vicentea. Phosphorus release with carbonate dissolution coupled to sulfide oxidation in Florida Bay seagrass sediments. Limnology and Oceanography, 54(5):1753–1764, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Joyce:2003:HAC

[JP03] Karen E. Joyce and Stuart R. Phinn. Hyperspectral analysis of chlorophyll content and photosynthetic capacity of coral reef substrates. Limnology and Oceanography, 48(1 (part 2)):489– 496, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Jensen:2006:EMD

[JPJS06] Jens Peder Jensen, Asger Roer Pedersen, Erik Jeppesen, and Martin Søndergaard. An empirical model describing the seasonal dynamics of phosphorus in 16 shallow eutrophic lakes after external loading reduction. Limnology and Oceanography, 51 (1 (part 2)):791–800, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Jonsson:2005:PDA

[JPK^{+}05] Per R. Jonsson, Jens K. Petersen, Örjan Karlsson, Lars-Ove Loo, and Stefan Nilsson. Particle depletion above experimental bivalve beds: In situ measurements and numerical modeling of bivalve filtration in the boundary layer. Limnology and Oceanography, 50(6):1989–1998, November 2005. CODEN LI-OCAH. ISSN 0024-3590.

Jimenez:2007:SSD

[JR07] Eroteida Jiménez and Marta Ribes. Sponges as a source of dissolved inorganic nitrogen: Nitrification mediated by temperate sponges. Limnology and Oceanography, 52(3):948–958, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2003:MGP

[JRS03] Matthew D. Johnson, Michelle Rome, and Diane K. Stoecker. Microzooplankton grazing on Prorocentrum minimum and Karlodinium micrum in Chesapeake Bay. Limnology and Oceanography, 48(1):238–248, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Jankowski:2003:CEB

[JS03a] Thomas Jankowski and Dietmar Straile. A comparison of eggbank and long-term plankton dynamics of two Daphnia species, D. hyalina and D. galeata: Potentials and limits of reconstruction. Limnology and Oceanography, 49(4):1948–1955, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Johnsen:2003:CCM

[JS03b] Sönke Johnsen and Heidi M. Sosik. Cryptic coloration and mirrored sides as camouflage strategies in near-surface pelagic habitats: Implications for foraging and predator avoidance. Limnology and Oceanography, 48(4):1277–1288, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Jakobsen:2004:CCG

[JS04] Hans H. Jakobsen and Suzanne L. Strom. Circadian cycles in growth and feeding rates of heterotrophic protist plankton. Limnology and Oceanography, 49(6):1915–1922, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Ji:2008:DEP

[JS08] Yongcheng Ji and Robert M. Sherrell. Differential effects of phosphorus limitation on cellular metals in chlorella and microcystis. Limnology and Oceanography, 53(5):1790–1804, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Jeppesen:2003:CWR

[JSJ03] Erik Jeppesen, Martin Søndergaard, and Jens Peder Jensen. Climatic warming and regime shifts in lake food webs — some comments. Limnology and Oceanography, 48(4):1346–1349, May 2003. CODEN LIOCAH. ISSN 0024-3590. See reply [SSvNH03].

Jones:2009:HCP

[JTAM09] Nicole L. Jones, Janet K. Thompson, Kevin R. Arrigo, and Stephen G. Monismith. Hydrodynamic control of phytoplankton loss to the benthos in an estuarine environment. Limnology and Oceanography, 54(3):952–969, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Juhl:2001:EFS

[JTL01] Andrew R. Juhl, Vera L. Trainer, and Michael I. Latz. Effect of fluid shear and irradiance on population growth and cellular toxin content of the dinoflagellate Alexandrium fundyense. Limnology and Oceanography, 46(4):758–764, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Juniper:2001:BRV

[Jun01] S. K. Juniper. Book review: Van Dover, C. L. 2000. The ecology of deep-sea hydrothermal vents. Princeton University Press. xx + 412 p. US$39.50. ISBN 0-691-04929-7. Limnology and Oceanography, 47(4):2094, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Juhl:2000:EGC

[JVL00] Andrew R. Juhl, Vivianna Velazquez, and Michael I. Latz. Effect of growth conditions on flow-induced inhibition of population growth of a red-tide dinoflagellate. Limnology and Oceanography, 45(4):905–915, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Jackson:2007:BRE

[JWRW07] J. L. Jackson, D. R. Webster, S. Rahman, and M. J. Weissburg. Bed roughness effects on boundary-layer turbulence and consequences for odor-tracking behavior of blue crabs (Callinectes sapidus). Limnology and Oceanography, 52(5):1883– 1897, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Kristensen:2006:CFC

[KA06] Erik Kristensen and Daniel M. Alongi. Control by fiddler crabs (Uca vocans) and plant roots (Avicennia marina) on carbon, iron, and sulfur biogeochemistry in mangrove sediment. Limnology and Oceanography, 51(4):1557–1571, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Karlsson:2008:WRA

[KAJ08] Jan Karlsson, Jenny Ask, and Mats Jansson. Winter respiration of allochthonous and autochthonous organic carbon in a subarctic clear-water lake. Limnology and Oceanography, 53(3): 948–954, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Kainz:2004:EFA

[KAM04] Martin Kainz, Michael T. Arts, and Asit Mazumder. Essential fatty acids in the planktonic food web and their ecological role for higher trophic levels. Limnology and Oceanography, 49(5): 1784–1793, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Kisand:2005:BFS

[KAW05] Veljo. Kisand, Nina. Andersson, and Johan. Wikner. Bacterial freshwater species successfully immigrate to the brackish water environment in the northern Baltic. Limnology and Oceanography, 50(3):945–956, May 2005. CODEN LIOCAH. ISSN 0024-

Kahler:2001:OPS

[KB01] Paul Kähler and Eduard Bauerfeind. Organic particles in a shallow sediment trap: Substantial loss to the dissolved phase. Limnology and Oceanography, 46(6):719–723, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Karlsson:2005:LEM

[KB05a] Jan Karlsson and Pär Byström. Littoral energy mobilization dominates energy supply for top consumers in subarctic lakes. Limnology and Oceanography, 50(4):538–543, March 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Kiorboe:2005:MPM

[KB05b] Thomas Kiørboe and Espen Bagøien. Motility patterns and mate encounter rates in planktonic copepods. Limnology and Oceanography, 50(6):1999–2007, November 2005. CODEN LI-OCAH. ISSN 0024-3590.

Kawasaki:2006:BRD

[KB06] Nobu Kawasaki and Ronald Benner. Bacterial release of dissolved organic matter during cell growth and decline: Molecular origin and composition. Limnology and Oceanography, 51(5): 2170–2180, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Keller:2007:PAC

[KB07] Jason K. Keller and Scott D. Bridgham. Pathways of anaerobic carbon cycling across an ombrotrophic-minerotrophic peatland gradient. Limnology and Oceanography, 52(1):96–107, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Kaiser:2008:EMB

[KB08a] Karl Kaiser and Ronald Benner. Erratum: Major bacterial contribution to the ocean reservoir of detrital organic carbon and nitrogen. Limnology and Oceanography, 54(3):1192, May 2008. CODEN LIOCAH. ISSN 0024-3590. See [KB08b].

Kaiser:2008:MBC

[KB08b] Karl Kaiser and Ronald Benner. Major bacterial contribution to the ocean reservoir of detrital organic carbon and nitrogen. Limnology and Oceanography, 53(1):99–112, January 2008. CODEN LIOCAH. ISSN 0024-3590. See erratum [KB08a].

Karp-Boss:2000:MDS

[KBBJ00] Lee Karp-Boss, Emmanuel Boss, and Peter A. Jumars. Motion of dinoflagellates in a simple shear flow. Limnology and Oceanography, 45(7):1594–1602, November 2000. CODEN LI-OCAH. ISSN 0024-3590.

Kuhn:2000:MEU

[KBD^{+}00] Penny S. Kuhn, Howard I. Browman, Richard F. Davis, John J. Cullen, and Bruce L. McArthur. Modeling the effects of ultraviolet radiation on embryos of Calanus finmarchicus and Atlantic cod (Gadus morhua) in a mixing environment. Limnology and Oceanography, 45(8):1797–1806, December 2000. CODEN LI-OCAH. ISSN 0024-3590.

Kerfoot:2004:WSS

[KBE^{+}04] W. Charles Kerfoot, Judy Wells Budd, Brian J. Eadie, Henry A. Vanderploeg, and Megan Agy. Winter storms: Sequential sediment traps record Daphnia ephippial production, resuspension, and sediment interactions. Limnology and Oceanography, 49(4 (part 2)):1365–1381, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:2008:DDL

[KBG^{+}08] W. Charles Kerfoot, Judith W. Budd, Sarah A. Green, James B. Cotner, Bopaiah A. Biddanda, David J. Schwab, and Henry A. Vanderploeg. Doughnut in the desert: Late-winter production pulse in southern Lake Michigan. Limnology and Oceanography, 53(2):589–604, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Kirby:2008:CIE

[KBL08] Richard R. Kirby, Gregory Beaugrand, and John A. Lindley. Climate-induced effects on the meroplankton and the benthicpelagic ecology of the North Sea. Limnology and Oceanography, 53(5):1805–1815, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Kroeger:2008:NBS

[KC08] K. D. Kroeger and M. A. Charette. Nitrogen biogeochemistry of submarine groundwater discharge. Limnology and Oceanography, 54(3):1025–1039, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Khripounoff:2006:GBR

[KCCE06] Alexis Khripounoff, Jean-Claude Caprais, Philippe Crassous, and Joël Etoubleau. Geochemical and biological recovery of the disturbed seafloor in polymetallic nodule fields of the clippertonclarion fracture zone (CCFZ) at 5,000-m depth. Limnology and Oceanography, 51(5):2033–2041, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Kaczmarska:2000:EUB

[KCE^{+}00] Irena Kaczmarska, Thomas A. Clair, James M. Ehrman, Sara L. MacDonald, David Lean, and Kristin E. Day. The effect of ultraviolet B on phytoplankton populations in clear and brown temperate Canadian lakes. Limnology and Oceanography, 45(3): 651–663, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Kritzberg:2004:AVA

[KCP^{+}04] Emma S. Kritzberg, Jonathan J. Cole, Michael L. Pace, Wilhelm Granéli, and Darren L. Bade. Autochthonous versus allochthonous carbon sources of bacteria: Results from whole-lake 13C addition experiments. Limnology and Oceanography, 49(2): 588–596, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Kamennaya:2008:CUC

[KCP08] Nina A. Kamennaya, Mark Chernihovsky, and Anton F. Post. The cyanate utilization capacity of marine unicellular cyanobacteria. Limnology and Oceanography, 53(6):2485–2494, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Katsev:2007:EPO

[KCSM07] Sergei Katsev, Gwénaëlle Chaillou, Bjørn Sundby, and Alfonso Mucci. Effects of progressive oxygen depletion on sediment diagenesis and fluxes: a model for the lower St. Lawrence River Estuary. Limnology and Oceanography, 52(6):2555–2568, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Kroeger:2006:GTD

[KCV06] K. D. Kroeger, M. L. Cole, and I. Valiela. Groundwatertransported dissolved organic nitrogen exports from coastal watersheds. Limnology and Oceanography, 51(5):2248–2261, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Kan:2006:BCC

[KCWC06] Jinjun Kan, Byron C. Crump, Kui Wang, and Feng Chen. Bacterioplankton community in Chesapeake Bay: Predictable or random assemblages. Limnology and Oceanography, 51(5):2157– 2169, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Kemp:2002:IAN

[KD02] Melody J. Kemp and Walter K. Dodds. The influence of ammonium, nitrate, and dissolved oxygen concentrations on uptake, nitrification, and denitrification rates associated with prairie stream substrata. Limnology and Oceanography, 47(5):1380– 1393, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:2005:BMB

[KDMC05] David L. Kirchman, Ana I. Dittel, Rex R. Malmstrom, and Matthew T. Cottrell. Biogeography of major bacterial groups in the Delaware Estuary. Limnology and Oceanography, 50(5): 1697–1706, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Katano:2009:URA

[KDO09] Izumi Katano, Hideyuki Doi, and Tadashi Oishi. Upstream resource abundance determines the food searching behavior of a stream grazer: Effect of microalgal cues. Limnology and Oceanography, 54(4):1162–1166, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Kranz:2009:CAT

[KDRR09] Sven A. Kranz, Sültemeyer Dieter, Klaus-Uwe Richter, and Björn Rost. Carbon acquisition by Trichodesmium: the effect of pCO_{2} and diurnal changes. Limnology and Oceanography, 54 (2):548–559, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Kutser:2003:MSD

[KDS03] Tiit Kutser, Arnold G. Dekker, and William Skirving. Modeling spectral discrimination of Great Barrier Reef benthic communities by remote sensing instruments. Limnology and Oceanography, 48(1 (part 2)):497–510, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:2007:SSA

[KED^{+}07] David L. Kirchman, Hila Elifantz, Ana I. Dittel, Rex R. Malmstrom, and Matthew T. Cottrell. Standing stocks and activity of archaea and bacteria in the western Arctic Ocean. Limnology and Oceanography, 52(2):495–507, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Kemp:2000:AME

[Kem00] Paul F. Kemp. Aquatic microbial ecology. Limnology and Oceanography, 45(5):1211, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Keil:2001:RAA

[KF01] Richard G. Keil and Marilyn L. Fogel. Reworking of amino acid in marine sediments: Stable carbon isotopic composition of amino acids in sediments along the Washington coast. Limnology and Oceanography, 46(1):14–23, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Kitajima:2009:LDD

[KFH^{+}09] Satoshi Kitajima, Ken Furuya, Fuminori Hashihama, Shigenobu Takeda, and Jota Kanda. Latitudinal distribution of diazotrophs and their nitrogen fixation in the tropical and subtropical western North Pacific. Limnology and Oceanography, 54(2):537–547, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Kristiansen:2000:PBS

[KFN00] Svein Kristiansen, Tove Farbrot, and Lars-Johan Naustvoll. Production of biogenic silica by spring diatoms. Limnology and Oceanography, 45(2):472–478, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Kelly:2001:NVC

[KFR^{+}01] Carol A. Kelly, Everett Fee, Patricia S. Ramlal, John W. M. Rudd, Raymond H. Hesslein, Cory Anema, and Eva U. Schindler. Natural variability of carbon dioxide and net epilimnetic production in the surface waters of boreal lakes of different sizes. Limnology and Oceanography, 46(5):1054–1064, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Klump:2009:BBC

[KFW09] J. Val Klump, Sharon A. Fitzgerald, and James T. Waplesa. Benthic biogeochemical cycling, nutrient stoichiometry, and carbon and nitrogen mass balances in a eutrophic freshwater bay. Limnology and Oceanography, 54(3):692–712, May 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Kostka:2002:RPC

[KGP^{+}02] Joel E. Kostka, Britta Gribsholt, Ellen Petrie, Dava Dalton, Hayley Skelton, and Erik Kristensen. The rates and pathways of carbon oxidation in bioturbated saltmarsh sediments. Limnology and Oceanography, 47(1):230–240, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Kagami:2006:SLC

[KGYU06] Maiko Kagami, Tek Bahadur Gurung, Takehito Yoshida, and Jotaro Urabe. To sink or to be lysed? Contrasting fate of two large phytoplankton species in Lake Biwa. Limnology and Oceanography, 52(6):2775–2786, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Krest:2003:UND

[KH03] James M. Krest and Judson W. Harvey. Using natural distributions of short-lived radium isotopes to quantify groundwater discharge and recharge. Limnology and Oceanography, 48(1): 290–298, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:2000:PAA

[KHB^{+}00] Jiří Kopáček, Josef Hejzlar, Jakub Borovec, Petr Porcal, and Ivana Kotorova. Phosphorus availability in an acidified watershed–lake ecosystem. Limnology and Oceanography, 45(1): 212–225, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Kononen:2003:DDC

[KHH^{+}03] Kaisa Kononen, Maija Huttunen, Seija Hällfors, Patrick Gentien, Michel Lunven, Timo Huttula, Jaan Laanemets, Madis Lilover, Juss Pavelson, and Adolf Stips. Development of a deep chlorophyll maximum of Heterocapsa triquetra ehrenb. at the entrance to the Gulf of Finland. Limnology and Oceanography, 48 (2):594–607, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:2003:PCB

[KHK^{+}03] Jiŕí Kopáček, Josef Hejzlar, Jiŕí Kaña, Petr Porcal, and Šárkárka Klementová. Photochemical, chemical, and biological transformations of dissolved organic carbon and its effect on alkalinity production in acidified lakes ji. Limnology and Oceanography, 48 (1):106–117, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Katechakis:2005:MVP

[KHKS05] Alexis Katechakis, Tanja Haseneder, Robert Kling, and Herwig Stibor. Mixotrophic versus photoautotrophic specialist algae as food for zooplankton: The light : nutrient hypothesis might not hold for mixotrophs. Limnology and Oceanography, 50(4):1290– 1299, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Krom:2004:NBE

[KHM04] M. D. Krom, B. Herut, and R. F. C. Mantoura. Nutrient budget for the Eastern Mediterranean: Implications for phosphorus limitation. Limnology and Oceanography, 49(5):1582–1592, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Kubanek:2005:DRT

[KHNV05] Julia Kubanek, Melissa K. Hicks, Jerome Naar, and Tracy A. Villareal. Does the red tide dinoflagellate Karenia brevis use allelopathy to outcompete other phytoplankton? Limnology and Oceanography, 50(3):883–895, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Kankaala:2006:MAR

[KHP^{+}06] Paula Kankaala, Jussi Huotari, Elina Peltomaa, Tuomo Saloranta, and Anne Ojala. Methanotrophic activity in relation to methane efflux and total heterotrophic bacterial production in a stratified, humic, boreal lake. Limnology and Oceanography, 51 (2):1195–1204, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Kortzinger:2001:RRR

[KHQ01] Arne Körtzinger, John I. Hedges, and Paul D. Quay. Redfield ratios revisited: Removing the biasing effect of anthropogenic CO_{2}. Limnology and Oceanography, 46(4):964–970, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Kolowith:2001:CCM

[KIB01] Lauren Clark Kolowith, Ellery D. Ingall, and Ronald Benner. Composition and cycling of marine organic phosphorus. Limnology and Oceanography, 46(2):309–320, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Kilham:2000:BLP

[Kil00] Susan S. Kilham. The biology of lakes and ponds. Limnology and Oceanography, 45(3):752, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:2005:LTC

[Kim05] Wim Kimmerer. Long-term changes in apparent uptake of silica in the San Francisco estuary. Limnology and Oceanography, 50 (3):793–798, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Kiorboe:2000:CMS

[Kiø00] Thomas Kiørboe. Colonization of marine snow aggregates by invertebrate zooplankton: Abundance, scaling, and possible role. Limnology and Oceanography, 45(2):479–484, March 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Kiorboe:2007:MFM

[Kiø07] Thomas Kiørboe. Mate finding, mating, and population dynamics in a planktonic copepod Oithona davisae: There are too few males. Limnology and Oceanography, 52(4):1511–1522, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Kirk:2003:VAI

[Kir03] John T. O. Kirk. The vertical attenuation of irradiance as a function of the optical properties of the water. Limnology and Oceanography, 48(1):9–17, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Kamjunke:2000:LIM

[KJ00] Norbert Kamjunke and Sabine Jähnichen. Leucine incorporation by Microcystis aeruginosa. Limnology and Oceanography, 45(3): 741–743, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Kiorboe:2001:MSO

[KJ01] Thomas Kiørboe and George A. Jackson. Marine snow, organic solute plumes, and optimal chemosensory behavior of bacteria. Limnology and Oceanography, 46(6):1309–1318, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Kringel:2003:SSL

[KJH03] Keli Kringel, Peter A. Jumars, and D. V. Holliday. A shallow scattering layer: High-resolution acoustic analysis of nocturnal vertical migration from the seabed. Limnology and Oceanography, 48(4):1223–1234, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Karlsson:2007:RAO

[KJJ07] Jan Karlsson, Mats Jansson, and Anders Jonsson. Respiration of allochthonous organic carbon in unproductive forest lakes determined by the Keeling plot method. Limnology and Oceanography, 52(2):603–608, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Karlsson:2003:CZD

[KJMJ03] Jan Karlsson, Anders Jonsson, Markus Meili, and Mats Jansson. Control of zooplankton dependence on allochthonous organic carbon in humic and clear-water lakes in northern Sweden. Limnology and Oceanography, 48(1):269–276, January 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Kuusipalo:2000:MFA

[KK00] Liisa Kuusipalo and Reijo Käkelä. Muscle fatty acids as indicators of niche and habitat in Malawian cichlids. Limnology and Oceanography, 45(4):996–1000, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Kicklighter:2004:DBN

[KKH04] C. E. Kicklighter, J. Kubanek, and M. E. Hay. Do brominated natural products defend marine worms from consumers? Some do, most don’t. Limnology and Oceanography, 49(2):430–441, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Kim:2006:BDO

[KKH06] Sunghwan Kim, Louis A. Kaplan, and Patrick G. Hatcher. Biodegradable dissolved organic matter in a temperate and a tropical stream determined from ultra-high resolution mass spectrometry. Limnology and Oceanography, 51(2):1054–1063, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Kaneda:2007:PAG

[KKKT07] Atsushi Kaneda, Takeshi Kohama, Yuji Kawamura, and Hidetaka Takeoka. Periodicity in the accumulation of gelatinous zooplankton during the summer season in the coastal area of iyo-nada, Japan. Limnology and Oceanography, 52(2):707–715, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Kolodny:2001:CTS

[KKS^{+}01] Y. Kolodny, A. Katz, A. Starinsky, T. Moise, and E. Simon. Chemical tracing of salinity sources in Lake Kinneret (Sea of Galilee), Israel. Limnology and Oceanography, 46(4):995, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Koski:2005:BLP

[KKT05] Marja Koski, Thomas Kiørboe, and Kazutaka Takahashi. Benthic life in the pelagic: Aggregate encounter and degradation rates by pelagic harpacticoid copepods. Limnology and Oceanography, 50(4):1254–1263, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Kaartvedt:2007:DVM

[KKT^{+}07] Stein Kaartvedt, Thor A. Klevjer, Thomas Torgersen, Tom A. Sørnes, and Anders Røstad. Diel vertical migration of individual jellyfish (Periphylla periphylla). Limnology and Oceanography, 52(3):975–983, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Kiene:2000:DTD

[KL00] Ronald P. Kiene and Laura J. Linn. Distribution and turnover of dissolved DMSP and its relationship with bacterial production and dimethylsulfide in the Gulf of Mexico. Limnology and Oceanography, 45(4):849–861, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Klausmeier:2001:AGV

[KL01] Christopher A. Klausmeier and Elena Litchman. Algal games: The vertical distribution of phytoplankton in poorly mixed water columns. Limnology and Oceanography, 46(8):1998–2007, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Kuznetsova:2004:EAA

[KLAF04] Marina Kuznetsova, Cindy Lee, Josephine Aller, and Nelson Frew. Enrichment of amino acids in the sea surface microlayer at coastal and open ocean sites in the North Atlantic Ocean. Limnology and Oceanography, 49(5):1605–1619, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Kim:2006:CPB

[KLC06] Hyun-Cheol Kim, Kitack Lee, and Wonyong Choi. Contribution of phytoplankton and bacterial cells to the measured alkalinity of seawater. Limnology and Oceanography, 51(1):331–338, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Konovalov:2003:LIO

[KLF^{+}03] Sergey K. Konovalov, George W. Luther III, Gernot E. Friederich, Donald B. Nuzzio, Bradley M. Tebo, James W. Murray, Temel Oguz, Brian Glazer, Robert E. Trouwborst, Brian Clement, Karen J. Murray, and Alexander S. Romanov. Lateral injection of oxygen with the Bosporus plume — fingers of oxidizing potential in the Black Sea. Limnology and Oceanography, 48(6):2369–2376, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Kimbro:2009:COP

[KLG09] David L. Kimbro, John Largier, and Edwin D. Grosholz. Coastal oceanographic processes influence the growth and size of a key estuarine species, the Olympia oyster. Limnology and Oceanography, 54(5):1425–1437, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Klausmeier:2004:PGS

[KLL04] Christopher A. Klausmeier, Elena Litchman, and Simon A. Levin. Phytoplankton growth and stoichiometry under multiple nutrient limitation. Limnology and Oceanography, 49(4 (part 2)):1463–1470, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Kim:2006:ESC

[KLS^{+}06] Ja-Myung Kim, Kitack Lee, Kyoungsoon Shin, Jung-Hoon Kang, Hyun-Woo Lee, Miok Kim, Pung-Guk Jang, and Min-Chul Jang. The effect of seawater CO_{2} concentration on growth of a natural phytoplankton assemblage in a controlled mesocosm experiment. Limnology and Oceanography, 51(4):1629– 1636, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Kim:2007:EES

[KLS^{+}07] Ja-Myung Kim, Kitack Lee, Kyoungsoon Shin, Jung-Hoon Kang, Hyun-Woo Lee, Miok Kim, Pung-Guk Jang, and Min-Chul Jang. Erratum: The effect of seawater CO_{2} concentration on growth of a natural phytoplankton assemblage in a controlled mesocosm experiment. Limnology and Oceanography, 52(2):922– 921, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Kuehn:2000:MBP

[KLSW00] K. A. Kuehn, M. J. Lemke, K. Suberkropp, and R. G. Wetzel. Microbial biomass and production associated with decaying leaf litter of the emergent macrophyte Juncus effusus. Limnology and Oceanography, 45(4):862–870, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Kessler:2008:VDZ

[KLWS08] Kirsten Kessler, Ryan S. Lockwood, Craig E. Williamson, and Jasmine E. Saros. Vertical distribution of zooplankton in subalpine and alpine lakes: Ultraviolet radiation, fish predation, and the transparency-gradient hypothesis. Limnology and Oceanography, 53(6):2374–2382, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Kelly:2002:SCG

[KM02] R. P. Kelly and S. B. Moran. Seasonal changes in groundwater input to a well-mixed estuary estimated using radium isotopes and implications for coastal nutrient budgets. Limnology and Oceanography, 47(6):1796–1807, November 2002. CODEN LI-OCAH. ISSN 0024-3590.

Kirchman:2000:CVI

[KMC^{+}00] David L. Kirchman, Benedikt Meon, Matthew T. Cottrell, David A. Hutchins, Debra Weeks, and Kenneth W. Bruland. Carbon versus iron limitation of bacterial growth in the California upwelling regime. Limnology and Oceanography, 45(8): 1681–1688, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:2004:PBS

[KMHE04] W. Charles Kerfoot, Gary G. Mittelbach, Nelson G. Hairston, Jr., and James J. Elser. Planktonic biodiversity: Scaling up and down. Limnology and Oceanography, 49(4 (part 2)):1225–1228, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:2007:NIP

[KMHN07] Jiři Kopáček, Marie Marešová, Josef Hejzlar, and Stephen A. Norton. Natural inactivation of phosphorus by aluminum in preindustrial lake sediments. Limnology and Oceanography, 52 (3):1147–1155, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Kirkpatrick:2000:ODP

[KMMS00] Gary J. Kirkpatrick, David F. Millie, Mark A. Moline, and Oscar Schofield. Optical discrimination of a phytoplankton species in natural mixed populations. Limnology and Oceanography, 45(2): 467–471, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Kropuenske:2009:PTM

[KMvD^{+}09] Lindsey R. Kropuenske, Matthew M. Mills, Gert L. van Dijken, Shaun Bailey, Dale H. Robinson, Nicholas A. Welschmeyer, and Kevin R. Arrigo. Photophysiology in two major Southern Ocean phytoplankton taxa: Photoprotection in Phaeocystis antarctica and Fragilariopsis cylindrus. Limnology and Oceanography, 54 (4):1176–1196, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Kimmel:2007:ICV

[KN07] David G. Kimmel and Roger I. E. Newell. The influence of climate variation on eastern oyster (Crassostrea virginica) juvenile abundance in Chesapeake Bay. Limnology and Oceanography, 52(3):959–965, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Kavanaugh:2009:EAE

[KNC^{+}09] Maria T. Kavanaugh, Karina J. Nielsen, Francis T. Chan, Bruce A. Menge, Ricardo M. Letelier, and Lea M. Goodrich. Experimental assessment of the effects of shade on an intertidal kelp: Do phytoplankton blooms inhibit growth of open coast macroalgae? Limnology and Oceanography, 54(1):276–288, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Kneib:2002:BRK

[Kne02] Ronald T. Kneib. Book review: Knox, G. A. 2001. The ecology of seashores. CRC Press. 557 p. US$90. ISBN 0-8493-0008-8. Limnology and Oceanography, 47(4):1268, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Karlson:2008:IBC

[KNE08] Agnes M. L. Karlson, Francisco J. A. Nascimento, and Ragnar Elmgren. Incorporation and burial of carbon from settling cyanobacterial blooms by deposit-feeding macrofauna. Limnology and Oceanography, 53(6):2754–2758, November 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Knauer:2000:PCH

[KNH00] K. Knauer, H. M. Nepf, and H. F. Hemond. The production of chemical heterogeneity in upper mystic lake. Limnology and Oceanography, 45(7):1647–1654, November 2000. CODEN LI-OCAH. ISSN 0024-3590.

Kim:2006:RDO

[KNN06] Chulgoo Kim, Yoko Nishimura, and Toshi Nagata. Role of dissolved organic matter in hypolimnetic mineralization of carbon and nitrogen in a large, monomictic lake. Limnology and Oceanography, 51(1):70–78, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Keleher:2008:DLH

[KR08] Mary Jane Keleher and Russell B. Rader. Dispersal limitations and history explain community composition of metaphyton in desert springs of the Bonneville Basin, Utah: a multiscale analysis. Limnology and Oceanography, 53(5):1604–1613, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Kraemer:2005:RIC

[Kra05] Thomas F. Kraemer. Radium isotopes in Cayuga Lake, New York: Indicators of inflow and mixing processes. Limnology and Oceanography, 55(3):158–168, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Kremp:2009:SSE

[KRM09] Anke Kremp, Karin Rengefors, and Marina Montresor. Species specific encystment patterns in three Baltic cold-water dinoflagellates: The role of multiple cues in resting cyst formation. Limnology and Oceanography, 54(4):1125–1138, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Kimmel:2006:STV

[KRZ06] David G. Kimmel, Michael R. Roman, and Xinsheng Zhang. Spatial and temporal variability in factors affecting mesozooplankton dynamics in Chesapeake Bay: Evidence from biomass size spectra. Limnology and Oceanography, 51(1):131–141, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Kaiser:2004:PRD

[KS04a] Edith Kaiser and Barbara Sulzberger. Phototransformation of riverine dissolved organic matter (DOM) in the presence of abundant iron: Effect on DOM bioavailability. Limnology and Oceanography, 49(2):540–554, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Keller:2004:NSR

[KS04b] Barbara Keller and Piet Spaak. Nonrandom sexual reproduction and diapausing egg production in a Daphnia hybrid species complex. Limnology and Oceanography, 49(4 (part 2)):1393– 1400, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Koski:2002:CCF

[KSEÖ^{+}02] Marja Koski, Katrin Schmidt, Jonna Engström-Öst, Markku Viitasalo, Sigrún Jónasdóttir, Sari Repka, and Kaarina Sivonen.

Calanoid copepods feed and produce eggs in the presence of toxic cyanobacteria Nodularia spumigena. Limnology and Oceanography, 47(3):878–885, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Kustka:2005:EPS

[KSM^{+}05] Adam B. Kustka, Yeala Shaked, Allen J. Milligan, D. Whitney King, and François M. M. Morel. Extracellular production of superoxide by marine diatoms: Contrasting effects on iron redox chemistry and bioavailability. Limnology and Oceanography, 50 (4):1172–1180, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Katsev:2006:MVE

[KSM06] Sergei Katsev, Bjørn Sundby, and Alfonso Mucci. Modeling vertical excursions of the redox boundary in sediments: Application to deep basins of the Arctic Ocean. Limnology and Oceanography, 51(4):1581–1593, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Kubanek:2007:CDR

[KSP07] Julia Kubanek, Terry W. Snell, and Christen Pirkle. Chemical defense of the red tide dinoflagellate Karenia brevis against rotifer grazing. Limnology and Oceanography, 52(3):1026–1035, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Kouraev:2007:IRL

[KSS^{+}07] Alexei V. Kouraev, Sergei V. Semovski, Michail N. Shimaraev, Nelly M. Mognard, Benoît Legrésy, and Frédérique Rémy. The ice regime of Lake Baikal from historical and satellite data: Relationship to air temperature, dynamical, and other factors. Limnology and Oceanography, 52(3):1268–1286, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Kustka:2003:IRD

[KSWC^{+}03] Adam B. Kustka, Sergio A. Sañudo-Wilhelmy, Edward J. Carpenter, Douglas Capone, James Burns, and William G. Sunda. Iron requirements for dinitrogen- and ammonium-supported growth in cultures of Trichodesmium (IMS 101): Comparison with nitrogen fixation rates and iron: carbon ratios of field populations. Limnology and Oceanography, 48(5):1869–1884, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Koschorreck:2002:BPA

[KT02] Matthias Koschorreck and Jörg Tittel. Benthic photosynthesis in an acidic mining lake (pH 2.6). Limnology and Oceanography, 47(4):1197–1201, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Kettle:2004:EMS

[KTAL04] Helen Kettle, Roy Thompson, N. John Anderson, and David M. Livingstone. Empirical modeling of summer lake surface temperatures in southwest Greenland. Limnology and Oceanography, 49 (1):271–282, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Kankaala:2006:EEC

[KTG^{+}06] Paula Kankaala, Sami Taipale, Jonathan Grey, Eloni Sonninen, Lauri Arvola, and Roger I. Jones. Experimental δ ^{13}C evidence for a contribution of methane to pelagic food webs in lakes. Limnology and Oceanography, 52(6):2821–2827, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Karsh:2003:RNI

[KTLS03] K. L. Karsh, T. W. Trull, M. J. Lourey, and D. M. Sigman. Relationship of nitrogen isotope fractionation to phytoplankton size and iron availability during the Southern Ocean Iron RElease Experiment (SOIREE). Limnology and Oceanography, 48 (3):1058–1068, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Keller:2001:OCC

[KTM01] Troy A. Keller, Abbie M. Tomba, and Paul A. Moore. Orientation in complex chemical landscapes: Spatial arrangement of chemical sources influences crayfish food-finding efficiency in artificial stream. Limnology and Oceanography, 46(2):238–247, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Kuma:2000:EHF

[KTMM00] Kenshi Kuma, Junya Tanaka, Kensuke Matsunaga, and Katsuhiko Matsunaga. Effect of hydroxamate ferrisiderophore complex (ferrichrome) on iron uptake and growth of a coastal marine diatom, Chaetoceros sociale. Limnology and Oceanography, 45 (6):1235–1244, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Kuuppo:2006:PDS

[KUP^{+}06] Pirjo Kuuppo, Pauliina Uronen, Anika Petermann, Timo Tamminen, and Edna Granéli. Pectenotoxin-2 and dinophysistoxin-1 in suspended and sedimenting organic matter in the Baltic Sea. Limnology and Oceanography, 51(5):2300–2307, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Kutser:2004:QDC

[Kut04] Tiit Kutser. Quantitative detection of chlorophyll in cyanobacterial blooms by satellite remote sensing. Limnology and Oceanography, 49(6):2179–2189, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Kagami:2004:DCP

[KVdB^{+}04] Maiko Kagami, Ellen Van Donk, Arnout de Bruin, Machteld Rijkeboer, and Bas W. Ibelings. Daphnia can protect diatoms from fungal parasitism. Limnology and Oceanography, 49(3): 680–685, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Knoll:2003:PPP

[KVR03] Lesley B. Knoll, Michael J. Vanni, and William H. Renwick. Phytoplankton primary production and photosynthetic parameters in reservoirs along a gradient of watershed land use. Limnology and Oceanography, 48(2):608–617, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:2004:EPR

[KW04] W. Charles Kerfoot and Lawrence J. Weider. Experimental paleoecology (resurrection ecology): Chasing Van Valen’s red Queen hypothesis. Limnology and Oceanography, 49(4 (part 2)):1300– 1316, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Kieber:2006:PPD

[KWS06] Robert J. Kieber, Robert F. Whitehead, and Stephen A. Skrabal. Photochemical production of dissolved organic carbon from resuspended sediments. Limnology and Oceanography, 51(5): 2187–2195, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Keller:2007:RIK

[KWTS07] Barbara Keller, Justyna Wolinska, Christoph Tellenbach, and Piet Spaak. Reproductive isolation keeps hybridizing Daphnia species distinct. Limnology and Oceanography, 52(3):984–991, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Llabres:2006:PCD

[LA06] Moira Llabrés and Susana Agustí. Picophytoplankton cell death induced by UV radiation: Evidence for oceanic Atlantic communities. Limnology and Oceanography, 51(1):21–29, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:2009:MDI

[LA09] David M. Livingstone and Rita Adrian. Modeling the duration of intermittent ice cover on a lake for climate-change studies. Limnology and Oceanography, 54(5):1709–1722, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Labiosa:2003:IBU

[LAG^{+}03] Rochelle G. Labiosa, Kevin R. Arrigo, Amatzia Genin, Stephen G. Monismith, and Gert van Dijken. The interplay between upwelling and deep convective mixing in determining the seasonal phytoplankton dynamics in the Gulf of Aqaba: Evidence from SeaWiFS and MODIS. Limnology and Oceanography, 48(6):2355–2368, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Ludwig:2005:PCC

[LAHdBJ05] Rebecca Ludwig, Fuad A. Al-Horani, Dirk de Beer, and Henk M. Jonkers. Photosynthesis-controlled calcification in a hypersaline microbial mat. Limnology and Oceanography, 50(6):1836–1843, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Larmola:2004:FCC

[LAJ^{+}04] Tuula Larmola, Jukka Alm, Sari Juutinen, Sanna Saarnio, Pertti J. Martikainen, and Jouko Silvola. Floods can cause large interannual differences in littoral net ecosystem productivity. Limnology and Oceanography, 49(5):1896–1906, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Lehman:2004:FBT

[LALE04] John T. Lehman, Sally E. B. Abella, Arni H. Litt, and W. Thomas Edmondson. Fingerprints of biocomplexity: Taxonspecific growth of phytoplankton in relation to environmental factors. Limnology and Oceanography, 49(4 (part 2)):1446–1456, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Laney:2003:AEP

[Lan03] S. R. Laney. Assessing the error in photosynthetic properties determined by fast repetition rate fluorometry. Limnology and Oceanography, 48(6):2234–2242, November 2003. CODEN LI-OCAH. ISSN 0024-3590.

Lightbody:2008:OSC

[LAN08] Anne F. Lightbody, Margaret E. Avener, and Heidi M. Nepf. Observations of short-circuiting flow paths within a free-surface wetland in Augusta, Georgia, U.S.A. Limnology and Oceanography, 54(3):1040–1053, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Latz:2009:EFC

[LASR09] Michael I. Latz, Jeffrey Allen, Sutanu Sarkar, and Jim Rohra. Effect of fully characterized unsteady flow on population growth of the dinoflagellate Lingulodinium polyedrum. Limnology and Oceanography, 54(4):1243–1256, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Laws:2003:MGP

[Law03] Edward A. Laws. Mesozooplankton grazing and primary production: an alternative assessment. Limnology and Oceanography, 48(4):1357–1359, May 2003. CODEN LIOCAH. ISSN 0024-3590. See reply [CP03].

Laws:2008:FWS

[Law08] Edward A. Laws. Food web structure and planktonic predatorprey relationships in two eutrophic European lakes: Stability constraints on carbon fluxes. Limnology and Oceanography, 53 (2):760–772, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Liu:2000:ESE

[LB00] Hongbin Liu and Edward J. Buskey. The exopolymer secretions (EPS) layer surrounding Aureoumbra lagunensis cells affects growth, grazing, and behavior of protozoa. Limnology and Oceanography, 45(5):1187–1191, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Laws:2004:NLL

[LB04a] E. A. Laws and T. T. Bannister. Nutrient- and light-limited growth of Thalassiosira fluviatilis in continuous culture with implications for phytoplankton growth in the ocean. Limnology and Oceanography, 49(6):2316, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Lindstrom:2004:IIB

[LB04b] Eva S. Lindström and Ann-Kristin Bergström. Influence of inlet bacteria on bacterioplankton assemblage composition in lakes of different hydraulic retention time. Limnology and Oceanography, 49(1):125–136, January 2004. CODEN LIOCAH. ISSN 0024-3590.

LeMoal:2009:DUC

[LB09] Morgane Le Moal and Isabelle C. Biegala. Diazotrophic unicellular cyanobacteria in the northwestern Mediterranean Sea: a seasonal cycle. Limnology and Oceanography, 54(3):845–855, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Lehmann:2004:IVI

[LBB^{+}04] Moritz F. Lehmann, Stefano M. Bernasconi, Alberto Barbieri, Marco Simona, and Judith A. McKenzie. Interannual variation of the isotopic composition of sedimenting organic carbon and nitrogen in Lake Lugano: a long-term sediment trap study. Limnology and Oceanography, 49(3):839–849, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Leynaert:2004:EID

[LBC^{+}04] A. Leynaert, E. Bucciarelli, P. Claquin, R. C. Dugdale, V. Martin-Jézéquel, P. Pondaven, and O. Ragueneau. Effect of iron deficiency on diatom cell size and silicic acid uptake kinetics. Limnology and Oceanography, 49(4):1134–1143, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Loh:2006:DPO

[LBC06] Ai Ning Loh, James E. Bauer, and Elizabeth A. Canuel. Dissolved and particulate organic matter source-age characterization in the upper and lower Chesapeake Bay: a combined isotope and biochemical approach. Limnology and Oceanography, 51(3): 1421–1431, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Leech:2009:UEF

[LBC^{+}09] Dina M. Leech, Wiebke J. Boeing, Sandra L. Cooke, Craig E. Williamson, and Lisette Torres. UV-enhanced fish predation and the differential migration of zooplankton to UV radiation and fish. Limnology and Oceanography, 54(4):1152–1161, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Loiselle:2009:OCC

[LBD^{+}09] Steven A. Loiselle, Luca Bracchini, Arduino M. Dattilo, Maso Ricci, Antonio Tognazzi, Andres Cózar, and Claudio Rossi. The optical characterization of chromophoric dissolved organic matter using wavelength distribution of absorption spectral slopes. Limnology and Oceanography, 54(2):590–597, March 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Leavitt:2006:LSE

[LBEP06] Peter R. Leavitt, Curtis S. Brock, Cortney Ebel, and Alain Patoine. Landscape-scale effects of urban nitrogen on a chain of freshwater lakes in central North America. Limnology and Oceanography, 51(5):2262–2277, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Lehmann:2009:ARH

[LBG^{+}09] Moritz F. Lehmann, Bruce Barnett, Y. Gélinas, Denis Gilbert, Roxane J. Maranger, Alfonso Mucci, Bjorn Sundby, and Benoit Thibodeau. Aerobic respiration and hypoxia in the Lower St. Lawrence Estuary: Stable isotope ratios of dissolved oxygen constrain oxygen sink partitioning. Limnology and Oceanography, 54(6):2157–2169, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Lehmann:2004:SVP

[LBM^{+}04] Moritz F. Lehmann, Stefano M. Bernasconi, Judith A. McKenzie, Alberto Barbieri, Marco Simona, and Mauro Veronesi. Seasonal variation of the δ ^{13}C and δ ^{15}N of particulate and dissolved carbon and nitrogen in Lake Lugano: Constraints on biogeochemical cycling in a eutrophic lake. Limnology and Oceanography, 49(2):415–429, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Luz:2002:ECR

[LBSY02] Boaz Luz, Eugeni Barkan, Yftach Sagi, and Yosef Z. Yacobi. Evaluation of community respiratory mechanisms with oxygen isotopes: a case study in Lake Kinneret. Limnology and Oceanography, 47(1):33–42, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Laforsch:2006:IDR

[LBT06] Christian Laforsch, Laura Beccara, and Ralph Tollrian. Inducible defenses: The relevance of chemical alarm cues in Daphnia. Limnology and Oceanography, 51(3):1466–1472, May 2006. CODEN LIOCAH. ISSN 0024-3590.

LaJeunesse:2008:SSH

[LBW^{+}08] T. C. LaJeunesse, H. Reyes Bonilla, M. E. Warner, M. Wills, G. W. Schmidt, and W. K. Fitt. Specificity and stability in high latitude eastern Pacific coral-algal symbioses. Limnology and Oceanography, 53(2):719–727, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Llope:2009:EEC

[LCC^{+}09] M. Llope, K.-S. Chan, L. Ciannelli, P. C. Reid, L. C. Stige, and N. C. Stenseth. Effects of environmental conditions on the seasonal distribution of phytoplankton biomass in the North Sea. Limnology and Oceanography, 54(2):512–524, March 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Lovejoy:2002:WCI

[LCLP02] Connie Lovejoy, Eddy C. Carmack, Louis Legendre, and Neil M. Price. Water column interleaving: a new physical mechanism determining protist communities and bacterial states. Limnology and Oceanography, 47(6):1819–1831, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Lohan:2005:IZE

[LCPS05] Maeve C. Lohan, David W. Crawford, Duncan A. Purdie, and Peter J. Statham. Iron and zinc enrichments in the northeastern subarctic Pacific: Ligand production and zinc availability in response to phytoplankton growth. Limnology and Oceanography, 50(5):1427–1437, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Lawrence:2002:VCL

[LCS02] Janice E. Lawrence, Amy M. Chan, and Curtis A. Suttle. Viruses causing lysis of the toxic bloom-forming alga Heterosigma akashiwo (Raphidophyceae) are widespread in coastal sediments of British Columbia, Canada. Limnology and Oceanography, 47(2):545–550, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Leavitt:2003:CCU

[LCS^{+}03] Peter R. Leavitt, B. F. Cumming, J. P. Smol, M. Reasoner, R. Pienitz, and D. A. Hodgson. Climatic control of ultraviolet radiation effects on lakes. Limnology and Oceanography, 49(4):

2062–2069, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:2001:EYS

[LD01] David M. Livingstone and Martin T. Dokulil. Eighty years of spatially coherent Austrian lake surface temperatures and their relationship to regional air temperature and the North Atlantic Oscillation. Limnology and Oceanography, 46(5):1220– 1227, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Lenes:2001:IFT

[LDC^{+}01] Jason M. Lenes, Brian P. Darrow, Christopher Cattrall, Cynthia A. Heil, Michael Callahan, Gabriel A. Vargo, Robert H. Byrne, Joseph M. Prospero, David E. Bates, Kent A. Fanning, and John J. Walsh. Iron fertilization and the Trichodesmium response on the West Florida shelf. Limnology and Oceanography, 46(6):1261–1277, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Louette:2006:ESY

[LED06] Gerald Louette, Mieke Vander Elst, and Luc De Meester. Establishment success in young cladoceran communities: an experimental test. Limnology and Oceanography, 51(2):1021–1030, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Lee:2001:GNC

[Lee01] Kitack Lee. Global net community production estimated from the annual cycle of surface water total dissolved inorganic carbon. Limnology and Oceanography, 46(6):1287–1297, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Legrand:2001:PTV

[Leg01] C. Legrand. Phagotrophy and toxicity variation in the mixotrophic Prymnesium patelliferum (Haptophyceae). Limnology and Oceanography, 46(5):1208–1214, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Lehman:2000:MWT

[Leh00a] John T. Lehman. In memoriam: W. Thomas Edmondson (1916– 2000). Limnology and Oceanography, 45(6):1448, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Lehman:2000:ICP

[Leh00b] P. W. Lehman. The influence of climate on phytoplankton community biomass in San Francisco Bay Estuary. Limnology and Oceanography, 45(3):580–590, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Lombard:2009:TER

[LEML09] F. Lombard, J. Erez, E. Michel, and L. Labeyrie. Temperature effect on respiration and photosynthesis of the symbiont-bearing planktonic Foraminifera Globigerinoides ruber, Orbulina universa, and Globigerinella siphonifera. Limnology and Oceanography, 54(1):210–218, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Leonardos:2008:PSS

[Leo08] Nikos Leonardos. Physiological steady state of phytoplankton in the field? An example based on pigment profile of Emiliania huxleyi (Haptophyta) during a light shift. Limnology and Oceanography, 53(1):306–311, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Lesser:2001:EUR

[Les01] Michael P. Lesser. The effects of UV radiation in the marine environment. Limnology and Oceanography, 46(1):211, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Leterme:2005:DBS

[LES^{+}05] S. C. Leterme, M. Edwards, L. Seuront, M. J. Attrill, P. C. Reid, and A. W. G. John. Decadal basin-scale changes in diatoms, dinoflagellates, and phytoplankton color across the North Atlantic. Limnology and Oceanography, 50(4):1244–1253, July 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Luengen:2009:RPM

[LF09] Allison C. Luengen and A. Russell Flegal. Role of phytoplankton in mercury cycling in the San Francisco Bay estuary. Limnology and Oceanography, 54(1):23–40, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Leavitt:2009:PEE

[LFA^{+}09] Peter R. Leavitt, S. C. Fritz, N. J. Anderson, P. A. Baker, T. Blenckner, L. Bunting, J. Catalan, D. J. Conley, W. O. Hobbs, E. Jeppesen, A. Korhola, S. McGowan, K. Rühland, J. A. Rusak, G. L. Simpson, N. Solovieva, and J. Werne. Paleolimnological evidence of the effects on lakes of energy and mass transfer from climate and humans. Limnology and Oceanography, 54(6 (part 2)):2330–2348, November 2009. CODEN LI-OCAH. ISSN 0024-3590.

Lindstrom:2006:ECB

[LFAB06] Eva S. Lindström, Markus Forslund, Grete Algesten, and Ann-Kristin Bergström. External control of bacterial community structure in lakes. Limnology and Oceanography, 51(1):339–342, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Lennon:2006:RIC

[LFFC06] Jay T. Lennon, Anthony M. Faiia, Xiahong Feng, and Kathryn L. Cottingham. Relative importance of CO_{2} recycling and CH_{4} pathways in lake food webs along a dissolved organic carbon gradient. Limnology and Oceanography, 51(4): 1602–1613, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Leu:2007:URN

[LFPH07] Eva Leu, Stig Falk-Petersen, and Dag O. Hessen. Ultraviolet radiation negatively affects growth but not food quality of Arctic diatoms. Limnology and Oceanography, 52(2):787–797, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Lamers:2002:FCE

[LFS^{+}02] Leon P. M. Lamers, Sarah-J. Falla, Edyta M. Samborska, Ivo A. R. van Dulken, Gijs van Hengstum, and Jan G. M. Roelofs. Factors controlling the extent of eutrophication and toxicity in sulfate-polluted freshwater wetlands. Limnology and Oceanography, 47(2):585–593, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Larsen:2004:SPB

[LFS^{+}04] Aud Larsen, Gro A. Fonnes Flaten, Ruth-Anne Sandaa, Tonje Castberg, Runar Thyrhaug, Svein Rune Erga, Stéphan Jacquet, and Gunnar Bratbak. Spring phytoplankton bloom dynamics in Norwegian coastal waters: Microbial community succession and diversity. Limnology and Oceanography, 49(1):180–190, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Larson:2007:EUL

[LFZ^{+}07] James H. Larson, Paul C. Frost, Zhiyu Zheng, Carol A. Johnston, Scott D. Bridgham, David M. Lodge, and Gary A. Lamberti. Effects of upstream lakes on dissolved organic matter in streams. Limnology and Oceanography, 52(1):60–69, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Leonardos:2004:REB

[LG04] Nikos Leonardos and Richard J. Geider. Responses of elemental and biochemical composition of Chaetoceros muelleri to growth under varying light and nitrate : phosphate supply ratios and their influence on critical N : P. Limnology and Oceanography, 49 (6):2105–2114, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Leclercq:2002:PPR

[LGJ02] Nicolas Leclercq, Jean-Pierre Gattuso, and Jean Jaubert. Primary production, respiration, and calcification of a coral reef mesocosm under increased CO_{2} partial pressure. Limnology and Oceanography, 47(2):558–564, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Langer:2006:CSC

[LGN^{+}06] Gerald Langer, Nikolaus Gussone, Gernot Nehrke, Ulf Riebesell, Anton Eisenhauer, Henning Kuhnert, Björn Rost, Scarlett Trimborn, and Silke Thoms. Coccolith strontium to calcium ratios in Emiliania huxleyi: The dependence on seawater strontium and calcium concentrations. Limnology and Oceanography, 51(1):310–320, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Lewandowski:2005:EMT

[LH05] Jörg Lewandowski and Michael Hupfer. Effect of macrozoobenthos on two-dimensional small-scale heterogeneity of pore water phosphorus concentrations in lake sediments: a laboratory study. Limnology and Oceanography, 50(4):1106–1118, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Long:2006:WIE

[LH06] Jeremy D. Long and Mark E. Hay. When intraspecific exceeds interspecific variance: Effects of phytoplankton morphology and growth phase on copepod feeding and fitness. Limnology and Oceanography, 51(2):988–996, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Li:2008:PAC

[LH08] William K. W. Li and W. Glen Harrison. Propagation of an atmospheric climate signal to phytoplankton in a small marine basin. Limnology and Oceanography, 53(5):1734–1745, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Lynam:2004:IVA

[LHB04] Christopher P. Lynam, Stephen J. Hay, and Andrew S. Brierley. Interannual variability in abundance of North Sea jellyfish and links to the North Atlantic Oscillation. Limnology and Oceanography, 49(3):637–643, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Larsson:2001:BSN

[LHWE01] Ulf Larsson, Susanna Hajdu, Jakob Walve, and Ragnar Elmgren. Baltic Sea nitrogen fixation estimated from the summer increase in upper mixed layer total nitrogen. Limnology and Oceanography, 46(4):811–820, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Licandro:2006:LTF

[LIE06] Priscilla Licandro, Frédéric Ibañez, and Michèle Etienne. Longterm fluctuations (1974-99) of the salps Thalia democratica and Salpa fusiformis in the northwestern Mediterranean Sea: Relationships with hydroclimatic variability. Limnology and Oceanography, 53(2):1832–1848, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Laval:2003:MCL

[LIHS03] Bernard Laval, Jörg Imberger, Ben R. Hodges, and Roman Stocker. Modeling circulation in lakes: Spatial and temporal variations. Limnology and Oceanography, 48(3):983–994, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Langenheder:2001:RBB

[LJ01] Silke Langenheder and Klaus Jürgens. Regulation of bacterial biomass and community structure by metazoan and protozoan predation. Limnology and Oceanography, 46(1):121–134, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Lebrato:2009:MDE

[LJ09] M. Lebrato and D. O. B. Jones. Mass deposition event of Pyrosoma atlanticum carcasses off Ivory Coast (West Africa). Limnology and Oceanography, 54(4):1197–1209, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Lane:2008:IBI

[LJCM08] Erin S. Lane, Kira Jang, Jay T. Cullen, and Maria T. Maldonado. The interaction between inorganic iron and cadmium uptake in the marine diatom Thalassiosira oceanica. Limnology and Oceanography, 53(5):1784–1789, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Letelier:2004:LDS

[LKAB04] Ricardo M. Letelier, David M. Karl, Mark R. Abbott, and Robert R. Bidigare. Light driven seasonal patterns of chlorophyll and nitrate in the lower euphotic zone of the North Pacific Subtropical Gyre. Limnology and Oceanography, 49(2):508–519, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Litchman:2004:PNC

[LKB04] Elena Litchman, Christopher A. Klausmeier, and Peter Bossard. Phytoplankton nutrient competition under dynamic light regimes. Limnology and Oceanography, 49(4 (part 2)):1457– 1462, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Lewin:2003:PUM

[LKM03] Wolf-Christian Lewin, Norbert Kamjunke, and Thomas Mehner. Phosphorus uptake by Microcystis during passage through fish guts. Limnology and Oceanography, 48(6):2392–2396, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Linden:2005:LSE

[LL05] Eveliina Lindén and Maiju Lehtiniemi. The lethal and sublethal effects of the aquatic macrophyte Myriophyllum spicatum on Baltic littoral planktivores. Limnology and Oceanography, 50 (2):405–411, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Lomas:2006:FPN

[LL06] Michael W. Lomas and Fredric Lipschultz. Forming the primary nitrite maximum: Nitrifiers or phytoplankton? Limnology and Oceanography, 51(5):2453–2467, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Laney:2005:PNF

[LLA05] S. R. Laney, R. M. Letelier, and M. R. Abbott. Parameterizing the natural fluorescence kinetics of Thalassiosira weissflogii. Limnology and Oceanography, 50(5):1499–1510, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Lesser:2004:TUR

[LLB04] Michael P. Lesser, Miles D. Lamare, and Michael F. Barker. Transmission of ultraviolet radiation through the Antarctic annual sea ice and its biological effects on sea urchin embryos. Limnology and Oceanography, 49(6):1957–1963, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Lemke:2007:IDA

[LLB07] A. Maria Lemke, Michael J. Lemke, and Arthur C. Benke. Importance of detrital algae, bacteria, and organic matter to littoral microcrustacean growth and reproduction. Limnology and Oceanography, 52(5):2164–2176, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Leynaert:2009:NLM

[LLC^{+}09] Aude Leynaert, Sorcha Ní Longphuirt, Pascal Claquin, Laurent Chauvaud, and Olivier Ragueneau. No limit? The multiphasic uptake of silicic acid by benthic diatoms. Limnology and Oceanography, 54(2):571–576, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Chen:2008:SDN

[lLCChTO08] Yuh ling Lee Chen, Houng-Yung Chen, Sing how Tuo, and Kaori Ohki. Seasonal dynamics of new production from Trichodesmium N_{2} fixation and nitrate uptake in the upstream Kuroshio and South China Sea basin. Limnology and Oceanography, 53(5):1705–1721, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Lunau:2006:PBC

[LLDS06] Mirko Lunau, Andreas Lemke, Olaf Dellwig, and Meinhard Simon. Physical and biogeochemical controls of microaggregate dynamics in a tidally affected coastal ecosystem. Limnology and Oceanography, 51(2):847–859, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Liu:2009:ROM

[LLF^{+}09] Pi-Jen Liu, Shiue-Ming Lin, Tung-Yung Fan, Pei-Jie Meng, Kwang-Tsao Shao, and Hsing-Juh Lina. Rates of overgrowth by macroalgae and attack by sea anemones are greater for live coral than dead coral under conditions of nutrient enrichment. Limnology and Oceanography, 54(4):1167–1175, July 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Livingstone:2005:ADD

[LLK05] David M. Livingstone, André F. Lotter, and Helen Kettle. Altitude-dependent differences in the primary physical response of mountain lakes to climatic forcing. Limnology and Oceanography, 50(4):1313–1325, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Lagergren:2008:SOV

[LLS08] Ragnar Lagergren, Karolina Leberfinger, and Jan A. E. Stenson. Seasonal and ontogenetic variation in diel vertical migration of Chaoborus flavicans and its effect on depth-selection behavior of other zooplankton. Limnology and Oceanography, 54(3):1083– 1092, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Langenheder:2005:WCB

[LLT05] Silke Langenheder, Eva S. Lindström, and Lars J. Tranvik. Weak coupling between community composition and functioning of aquatic bacteria. Limnology and Oceanography, 50(3):957–967, May 2005. CODEN LIOCAH. ISSN 0024-3590.

LaJeunesse:2003:LSD

[LLvW^{+}03] Todd C. LaJeunesse, William K. W. Loh, Robert van Woesik, Ove Hoegh-Guldberg, Gregory W. Schmidt, and William K. Fitt. Low symbiont diversity in southern Great Barrier Reef corals, relative to those of the Caribbean. Limnology and Oceanography, 49(4):2046–2054, September 2003. CODEN LI-OCAH. ISSN 0024-3590.

Levin:2002:IEC

[LM02] Lisa A. Levin and Robert H. Michener. Isotopic evidence for chemosynthesis-based nutrition of macrobenthos: The lightness of being at Pacific methane seeps. Limnology and Oceanography, 47(5):1336–1345, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Lemmin:2005:ISD

[LMB05] U. Lemmin, C. H. Mortimer, and E. Bäuerle. Internal seiche dynamics in Lake Geneva. Limnology and Oceanography, 55(3): 207–216, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Loisel:2007:IOB

[LMBP07] Hubert Loisel, Xavier Mériaux, Jean-François Berthon, and Antoine Poteau. Investigation of the optical backscattering to scattering ratio of marine particles in relation to their biogeochemical composition in the eastern English Channel and southern North Sea. Limnology and Oceanography, 52(2):739–752, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Lorke:2003:BSC

[LMMW03] Andreas Lorke, Beat Müller, Martin Maerki, and Alfred Wüest. Breathing sediments: The control of diffusive transport across the sediment–water interface by periodic boundary-layer turbulence. Limnology and Oceanography, 48(6):2077–2085, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Lesser:2000:LAU

[LMPY00] Michael P. Lesser Lesser, Charles Mazel, David Phinney, and Charles S. Yentsch. Light absorption and utilization by colonies of the congeneric hermatypic corals Montastraea faveolata and Montastraea cavernosa. Limnology and Oceanography, 45(1): 76–86, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Latasa:2005:ECF

[LMSE05] Mikel Latasa, Xosé Anxelu G. Morán, Renate Scharek, and Marta Estrada. Estimating the carbon flux through main phytoplankton groups in the northwestern Mediterranean. Limnology and Oceanography, 50(5):1447–1458, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Lessard:2005:NPR

[LMT05] Evelyn J. Lessard, Agostino Merico, and Toby Tyrrell. Nitrate : phosphate ratios and Emiliania huxleyi blooms. Limnology and Oceanography, 50(3):1020–1024, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Long:2008:ROA

[LMZB08] Matthew H. Long, Karen J. McGlathery, Joseph C. Zieman, and Peter Berg. The role of organic acid exudates in liberating phosphorus from seagrass-vegetated carbonate sediments. Limnology and Oceanography, 53(6):2616–2626, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Levinsen:2002:TRM

[LN02] Henrik Levinsen and Torkel Gissel Nielsen. The trophic role of marine pelagic ciliates and heterotrophic dinoflagellates in Arctic and temperate coastal ecosystems: a cross-latitude comparison. Limnology and Oceanography, 47(2):427–439, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Leibold:2004:BMP

[LN04] Mathew A. Leibold and Jon Norberg. Biodiversity in metacommunities: Plankton as complex adaptive systems? Limnology and Oceanography, 49(4 (part 2)):1278–1289, July 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Lightbody:2006:PVP

[LN06] Anne F. Lightbody and Heidi M. Nepf. Prediction of velocity profiles and longitudinal dispersion in salt marsh vegetation. Limnology and Oceanography, 51(1):218–228, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Litchman:2002:ISU

[LNB02] Elena Litchman, Patrick J. Neale, and Anastazia T. Banaszak. Increased sensitivity to ultraviolet radiation in nitrogen-limited dinoflagellates: Photoprotection and repair. Limnology and Oceanography, 47(1):86–94, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Larned:2004:MTL

[LNB04] Scott T. Larned, Vladimir I. Nikora, and Barry J. F. Biggs. Mass-transfer-limited nitrogen and phosphorus uptake by stream periphyton: a conceptual model and experimental evidence. Limnology and Oceanography, 49(6):1992–2000, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Lomstein:2009:APR

[LNJL09] Bente Aa. Lomstein, Jutta Niggemann, Bo B. Jørgensen, and Alice T. Langerhuusa. Accumulation of prokaryotic remains during organic matter diagenesis in surface sediments off Peru. Limnology and Oceanography, 54(4):1139–1151, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:2000:COC

[Lon00] Alan Longhurst. The changing ocean carbon cycle — a midterm synthesis of the joint global ocean flux study. Limnology and Oceanography, 45(6):1446–1447, September 2000. CODEN LI-OCAH. ISSN 0024-3590.

Livingstone:2007:LSC

[LP07] David M. Livingstone and Judit Padisák. Large-scale coherence in the response of lake surface-water temperatures to synopticscale climate forcing during summer. Limnology and Oceanography, 52(2):896–902, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Lindell:2005:ENS

[LPAQ^{+}05] Debbie Lindell, Sigrid Penno, Mutaz Al-Qutob, Efrat David, Tanya Rivlin, Boaz Lazar, and Anton F. Post. Expression of the nitrogen stress response gene ntca reveals nitrogen-sufficient Synechococcus populations in the oligotrophic northern Red Sea. Limnology and Oceanography, 50(6):1932–1944, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Lorke:2006:HFI

[LPB06] Andreas Lorke, Frank Peeters, and Erich Bäuerle. Highfrequency internal waves in the littoral zone of a large lake. Limnology and Oceanography, 53(2):1935–1939, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Lee:2004:ELD

[LPD^{+}04] Peter A. Lee, John C. Priscu, Giacomo R. DiTullio, Sarah F. Riseman, Nicole Tursich, and Stephen J. de Mora. Elevated levels of dimethylated-sulfur compounds in Lake Bonney, a poorly ventilated Antarctic lake. Limnology and Oceanography, 49(4): 1044–1055, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Lorke:2005:SIC

[LPW05] Andreas Lorke, Frank Peeters, and Alfred Wüest. Shear-induced convective mixing in bottom boundary layers on slopes. Limnology and Oceanography, 50(5):1612–1619, September 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Luengen:2007:CBS

[LRF07] Allison C. Luengen, Peter T. Raimondi, and A. Russell Flegal. Contrasting biogeochemistry of six trace metals during the rise and decay of a spring phytoplankton bloom in San Francisco Bay. Limnology and Oceanography, 52(3):1112–1130, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Logue:2004:RBS

[LRMV04] Jürg Brendan Logue, Christopher T. Robinson, Christoph Meier, and Jan Roelof Van der Meer. Relationship between sediment organic matter, bacteria composition, and the ecosystem metabolism of alpine streams. Limnology and Oceanography, 49 (6):2001–2010, November 2004. CODEN LIOCAH. ISSN 0024-

Louchard:2003:ORS

[LRS^{+}03] Eric M. Louchard, R. Pamela Reid, F. Carol Stephens, Curtiss O. Davis, Robert A. Leathers, and Downes T. Valerie. Optical remote sensing of benthic habitats and bathymetry in coastal environments at Lee Stocking Island, Bahamas: a comparative spectral classification approach. Limnology and Oceanography, 48(1 (part 2)):511–521, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Levy:2006:HCS

[LRYS06] Oren Levy, Michael Rosenfeld, Ruth Yam, and Aldo Shemesh. Heterogeneity of coral skeletons isotopic compositions during the 1998 bleaching event. Limnology and Oceanography, 51(2):1142– 1148, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Lin:2008:SDB

[LSC^{+}08] Xueju Lin, Mary I. Scranton, Andrei Y. Chistoserdov, Ramon Varela, and Gordon T. Taylor. Spatiotemporal dynamics of bacterial populations in the anoxic Cariaco Basin. Limnology and Oceanography, 53(1):37–51, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Lowe:2008:MFC

[LSF^{+}08] Ryan J. Lowe, Uri Shavit, James L. Falter, Jeffrey R. Koseff, and Stephen G. Monismith. Modeling flow in coral communities with and without waves: a synthesis of porous media and canopy flow approaches. Limnology and Oceanography, 53(6): 2668–2680, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Lignell:2003:BBP

[LSK^{+}03] Risto Lignell, Jukka Seppälä, Pirjo Kuuppo, Timo Tamminen, Tom Andersen, and Ingrid Gismervik. Beyond bulk properties: Responses of coastal summer plankton communities to nutrient enrichment in the northern Baltic Sea. Limnology and Oceanography, 48(1):189–209, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Lavaud:2007:PCD

[LSK07] Johann Lavaud, Robert F. Strzepek, and Peter G. Kroth. Photoprotection capacity differs among diatoms: Possible consequences on the spatial distribution of diatoms related to fluctuations in the underwater light climate. Limnology and Oceanography, 52(3):1188–1194, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Leichter:2003:ENT

[LSM03] James J. Leichter, Hannah L. Stewart, and Steven L. Miller. Episodic nutrient transport to Florida coral reefs. Limnology and Oceanography, 48(4):1394–1407, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Lampitt:2002:OGW

[LSPC02] Richard S. Lampitt, Paulo Y. G. Sumida, and Fernando Pérez-Castillo. Ophiuroid growth within deep-sea sediment traps: a problem for carbon flux measurements at continental margins. Limnology and Oceanography, 47(2):571–575, March 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Lomas:2004:TVP

[LSSA04] Michael W. Lomas, Ashley Swain, Ryan Shelton, and James W. Ammerman. Taxonomic variability of phosphorus stress in Sargasso Sea phytoplankton. Limnology and Oceanography, 49(6): 2303–2309, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Lucas:2009:WDR

[LTB09] Lisa V. Lucas, Janet K. Thompson, and Larry R. Brown. Why are diverse relationships observed between phytoplankton biomass and transport time? Limnology and Oceanography, 54 (1):381–390, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Lefevre:2003:MCN

[LTGG03] Nathalie Lefèvre, Arnold H. Taylor, Francis J. Gilbert, and Richard J. Geider. Modeling carbon to nitrogen and carbon to chlorophyll-a ratios in the ocean at low latitudes: Evaluation of the role of physiological plasticity. Limnology and Oceanography, 48(5):1796–1807, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Lorke:2004:RLK

[LTHW04] Andreas Lorke, Klaus Tietze, Michel Halbwachs, and Alfred Wüest. Response of Lake Kivu stratification to lava inflow and climate warming. Limnology and Oceanography, 49(3):778–783, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Lam:2001:DEI

[LTM01] Phoebe J. Lam, Philippe D. Tortell, and François M. M. Morel. Differential effects of iron additions on organic and inorganic carbon production by phytoplankton. Limnology and Oceanography, 46(5):1199–1202, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Lundquist:2004:LTR

[LTOS04] Carolyn J. Lundquist, Simon F. Thrush, John W. Oldman, and Alastair K. Senior. Limited transport and recolonization potential in shallow tidal estuaries. Limnology and Oceanography, 49 (2):386–395, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Lovdal:2007:ABC

[LTT07] Trond Løvdal, Tsuneo Tanaka, and T. Frede Thingstad. Algal– bacterial competition for phosphorus from dissolved DNA, ATP, and orthophosphate in a mesocosm experiment. Limnology and Oceanography, 52(4):1407–1419, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Urrutia:2008:MTE

[LU08] Ángel López-Urrutia. The metabolic theory of ecology and algal bloom formation. Limnology and Oceanography, 53(5):2046– 2047, September 2008. CODEN LIOCAH. ISSN 0024-3590. See reply [CR08].

Lopez-Urrutia:2004:PCC

[LUHS04] Ángel López-Urrutia, Roger P. Harris, and Tania Smith. Predation by calanoid copepods on the appendicularian Oikopleura dioica. Limnology and Oceanography, 49(1):303–307, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Lurling:2003:DGM

[Lür03] Miquel Lürling. Daphnia growth on microcystin-producing and microcystin-free Microcystis aeruginosa in different mixtures with the green alga Scenedesmus obliquus. Limnology and Oceanography, 48(6):2214–2220, November 2003. CODEN LI-OCAH. ISSN 0024-3590.

Lurling:2001:CFS

[LV01] Miquel Lürling and Eric Von Elert. Colony formation in Scenedesmus: No contribution of urea in induction by a lipophilic Daphnia exudate. Limnology and Oceanography, 46 (7):1809–1813, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Loucaide:2008:DBS

[LVB08] Socratis Loucaide, Philippe Van Cappelle, and Thilo Behrends. Dissolution of biogenic silica from land to ocean: Role of salinity and pH. Limnology and Oceanography, 53(5):1614–1621, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Laurion:2000:AUR

[LVC^{+}00] Isabelle Laurion, Marc Ventura, Jordi Catalan, Roland Psenner, and Ruben Sommaruga. Attenuation of ultraviolet radiation in mountain lakes: Factors controlling the among- and withinlake variability. Limnology and Oceanography, 45(6):1274–1288, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Laglera:2009:EGC

[LvdB09] Luis M. Laglera and Constant M. G. van den Berg. Evidence for geochemical control of iron by humic substances in seawater. Limnology and Oceanography, 54(2):610–619, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Lottig:2007:FAC

[LVSW07] Noah R. Lottig, H. Maurice Valett, Madeline E. Schreiber, and Jackson R. Webster. Flooding and arsenic contamination: Influences on ecosystem structure and function in an Appalachian headwater stream. Limnology and Oceanography, 52(5):1991– 2001, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Lancaster:2001:SIV

[LW01a] Jill Lancaster and Susan Waldron. Stable isotope values of lotic invertebrates: Sources of variation, experimental design, and statistical interpretation. Limnology and Oceanography, 46(6): 723–730, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Leech:2001:SEU

[LW01b] Dina M. Leech and Craig E. Williamson. In situ exposure to ultraviolet radiation alters the depth distribution of Daphnia. Limnology and Oceanography, 49(1):416–420, March 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Lotze:2002:CIC

[LW02] Heike K. Lotze and Boris Worm. Complex interactions of climatic and ecological controls on macroalgal recruitment. Limnology and Oceanography, 47(6):1734–1741, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Lyons:2001:PAG

[LWB^{+}01] W. B. Lyons, K. A. Welch, J.-C. Bonzongo, E. Y. Graham, G. Shabunin, H. E. Gaudette, and R. J. Poreda. A preliminary assessment of the geochemical dynamics of Issyk-Kul Lake, Kirghizstan. Limnology and Oceanography, 46(6):713–718, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Leichter:2007:NOI

[LWP^{+}07] James J. Leichter, Scott Wankel, Adina Paytan, Katharine Hanson, Steven Miller, and Mark A. Altabet. Nitrogen and oxygen isotopic signatures of subsurface nitrate seaward of the Florida Keys reef tract. Limnology and Oceanography, 52(3):1258–1267, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Lotze:2001:SBT

[LWS01] Heike K. Lotze, Boris Worm, and Ulrich Sommer. Strong bottom-up and top-down control of early life stages of macroalgae. Limnology and Oceanography, 46(4):749–757, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Lyons:2005:LMS

[LWU^{+}05] M. Maille Lyons, J. Evan Ward, Kevin R. Uhlinger, Rebecca J. Gast, and Roxanna Smolowitz. Lethal marine snow: Pathogen of bivalve mollusc concealed in marine aggregates. Limnology and Oceanography, 50(6):1983–1988, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Lyon:2009:CCW

[LZ09] David R. Lyon and Susan E. Ziegler. Carbon cycling within epilithic biofilm communities across a nutrient gradient of headwater streams. Limnology and Oceanography, 54(2):439–449, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Mohr:2000:FRR

[MA00] Silvia Mohr and Rita Adrian. Functional responses of the rotifers Brachionus calyciflorus and Brachionus rubens feeding on armored and unarmored ciliates. Limnology and Oceanography, 45(5):1175–1179, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Merritt:2008:MCE

[MA08] Karen A. Merritt and Aria Amirbahman. Methylmercury cycling in estuarine sediment pore waters (Penobscot River estuary, Maine, USA). Limnology and Oceanography, 54(3):1064– 1075, May 2008. CODEN LIOCAH. ISSN 0024-3590.

McKee:2003:RFM

[MAC^{+}03] D. McKee, D. Atkinson, S. E. Collings, J. W. Eaton, A. B. Gill, I. Harvey, K. Hatton, T. Heyes, D. Wilson, and B. Moss. Response of freshwater microcosm communities to nutrients, fish, and elevated temperature during winter and summer. Limnology and Oceanography, 48(2):707–722, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2006:CDI

[MAC^{+}06] Maria T. Maldonado, Andrew E. Allen, Joanne S. Chong, Kevin Lin, Dan Leus, Nataliya Karpenko, and Shannon L. Harris. Copper-dependent iron transport in coastal and oceanic diatoms. Limnology and Oceanography, 52(2):1729–1743, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Malzahn:2007:NLP

[MACB07] Arne Michael Malzahn, Nicole Aberle, Catriona Clemmesen, and Maarten Boersma. Nutrient limitation of primary producers affects planktivorous fish condition. Limnology and Oceanography, 52(5):2062–2071, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Muller:2008:ICC

[MAL08] Marius N. Müller, Avan N. Antia, and Julie LaRoche. Influence of cell cycle phase on calcification in the coccolithophore Emiliania huxleyi. Limnology and Oceanography, 53(2):506–512, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Meador:2007:IHC

[MAM07] T. B. Meador, L. I. Aluwihare, and C. Mahaffey. Isotopic heterogeneity and cycling of organic nitrogen in the oligotrophic ocean. Limnology and Oceanography, 52(3):934–947, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Mann:2002:CTC

[MAMC02] Elizabeth L. Mann, Nathan Ahlgren, James W. Moffett, and Sallie W. Chisholm. Copper toxicity and cyanobacteria ecology in the Sargasso Sea. Limnology and Oceanography, 47(4):976– 988, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Maranon:2005:PGR

[Mar05] Emilio Marañón. Phytoplankton growth rates in the Atlantic subtropical gyres. Limnology and Oceanography, 55(3):299–310, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Meyer:2002:FEB

[MAS^{+}02] B. Meyer, A. Atkinson, D. Stöbing, B. Oettl, W. Hagen, and U. V. Bathmann. Feeding and energy budgets of Antarctic krill Euphausia superba at the onset of winter — I. Furcilia III larvae. Limnology and Oceanography, 47(4):943–952, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Makino:2000:RLH

[MB00] Wataru Makino and Syuhei Ban. Response of life history traits to food conditions in a cyclopoid copepod from an oligotrophic environment. Limnology and Oceanography, 45(2):396– 407, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Manley:2001:PCB

[MB01] Steven L. Manley and Pebble E. Barbero. Physiological constraints on bromoform (CHBr_{3}) production by Ulva lactuca (Chlorophyta). Limnology and Oceanography, 46(6):1392–1399, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2005:EGR

[MB05] Margaret R. Mulholland and Peter W. Bernhardt. The effect of growth rate, phosphorus concentration, and temperature on N_{2} fixation, carbon fixation, and nitrogen release in continuous cultures of Trichodesmium IMS10. Limnology and Oceanography, 50(3):839–849, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Middelburg:2000:FIM

[MBB^{+}00] Jack J. Middelburg, Christiane Barranguet, Henricus T. S. Boschker, Peter M. J. Herman, Tom Moens, and Carlo H. R. Heip. The fate of intertidal microphytobenthos carbon: an in situ ^{13}C-labeling study. Limnology and Oceanography, 45(6): 1224–1234, September 2000. CODEN LIOCAH. ISSN 0024-

Morin:2001:USS

[MBC01] Antoine Morin, Nathalie Bourassa, and Antonella Cattaneo. Use of size spectra and empirical models to evaluate trophic relationships in streams. Limnology and Oceanography, 46(4):935–940, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Modenutti:2005:EUB

[MBC^{+}05] Beatriz E. Modenutti, Esteban G. Balseiro, Cristiana Callieri, Roberto Bertoni, and Claudia P. Queimaliños. Effect of UV-B and different PAR intensities on the primary production of the mixotrophic planktonic ciliate Stentor araucanus. Limnology and Oceanography, 50(3):864–871, May 2005. CODEN LI-OCAH. ISSN 0024-3590.

McCallister:2006:BED

[MBC06] S. Leigh McCallister, James E. Bauer, and Elizabeth A. Canuel. Bioreactivity of estuarine dissolved organic matter: a combined geochemical and microbiological approach. Limnology and Oceanography, 51(1):94–100, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Modenutti:2008:LVF

[MBCB08] Beatriz E. Modenutti, Esteban G. Balseiro, Cristiana Callieri, and Roberto Bertoni. Light versus food supply as factors modulating niche partitioning in two pelagic mixotrophic ciliates. Limnology and Oceanography, 53(2):446–455, March 2008. CO-DEN LIOCAH. ISSN 0024-3590.

McCallister:2004:ASA

[MBCD04] S. Leigh McCallister, James E. Bauer, Jennifer E. Cherrier, and Hugh W. Ducklow. Assessing sources and ages of organic matter supporting river and estuarine bacterial production: a multipleisotope (Δ ^{14}C, Δ ^{13}C, and Δ ^{15}N) approach. Limnology and Oceanography, 49(5):1687–1702, September 2004. CODEN LI-OCAH. ISSN 0024-3590.

Martel:2001:SEH

[MBDL01] André L. Martel, Brad S. Baldwin, Ronald M. Dermott, and Richard A. Lutz. Species and epilimnion/hypolimnion-related differences in size at larval settlement and metamorphosis in Dreissena (Bivalvia). Limnology and Oceanography, 46(6):707– 713, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Mackay:2003:APD

[MBF^{+}03] Anson W. Mackay, Richard W. Battarbee, Roger J. Flower, Nicholas G. Granin, David H. Jewson, David B. Ryves, and Michael Sturm. Assessing the potential for developing internal diatom-based transfer functions for Lake Baikal. Limnology and Oceanography, 48(3):1183–1192, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Mayzaud:2005:RMZ

[MBG^{+}05] Patrick Mayzaud, Marc Boutoute, Stéphane Gasparini, Laure Mousseau, and Dominique Lefevre. Respiration in marine zooplankton — the other side of the coin: CO_{2} production. Limnology and Oceanography, 55(3):291–298, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2006:ENF

[MBH^{+}06a] Margaret R. Mulholland, Peter W. Bernhardt, Cynthia A. Heil, Deborah A. Bronk, and Judith M. O’Neil. Erratum: Nitrogen fixation and release of fixed nitrogen by Trichodesmium spp. in the Gulf of Mexico. Limnology and Oceanography, 51(5):2484, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2006:NFR

[MBH^{+}06b] Margaret R. Mulholland, Peter W. Bernhardt, Cynthia A. Heil, Deborah A. Bronk, and Judith M. O’Neil. Nitrogen fixation and release of fixed nitrogen by Trichodesmium spp. in the Gulf of Mexico. Limnology and Oceanography, 52(2):1762–1776, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Masclaux:2009:CEF

[MBK^{+}09] Hélène Masclaux, Alexandre Bec, Martin J. Kainz, Christian Desvilettes, Lionel Jouve, and Gilles Bourdier. Combined effects of food quality and temperature on somatic growth and reproduction of two freshwater cladocerans. Limnology and Oceanography, 54(4):1323–1332, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2001:IUP

[MBL^{+}01] Maria T. Maldonado, Philip W. Boyd, Julie LaRoche, Robert Strzepek, Anya Waite, Andrew R. Bowie, Peter L. Croot, Russell D. Frew, and Neil M. Price. Iron uptake and physiological response of phytoplankton during a mesoscale Southern Ocean iron enrichment. Limnology and Oceanography, 46(7):

1802–1808, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Mari:2001:NRC

[MBLP01] Xavier Mari, Sophie Beauvais, Rodolphe Lemée, and Maria Luiza Pedrotti. Non-Redfield C : N ratio of transparent exopolymeric particles in the northwestern Mediterranean Sea. Limnology and Oceanography, 46(7):1831–1836, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Morales-Baquero:2006:AIP

[MBPVR06] Rafael Morales-Baquero, Elvira Pulido-Villena, and Isabel Reche. Atmospheric inputs of phosphorus and nitrogen to the southwest Mediterranean region: Biogeochemical responses of high mountain lakes. Limnology and Oceanography, 51(2):830– 837, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Miller:2008:CSI

[MBR08] Todd W. Miller, Richard D. Brodeur, and Greg H. Rau. Carbon stable isotopes reveal relative contribution of shelf-slope production to the Northern California Current pelagic community. Limnology and Oceanography, 53(4):1493–1503, July 2008. CO-DEN LIOCAH. ISSN 0024-3590.

McKnight:2001:SCD

[MBW^{+}01] Diane M. McKnight, Elizabeth W. Boyer, Paul K. Westerhoff, Peter T. Doran, Thomas Kulbe, and Dale T. Andersen. Spectrofluorometric characterization of dissolved organic matter for indication of precursor organic material and aromaticity. Limnology and Oceanography, 46(1):38–48, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Mann:2000:ILC

[MC00] Elizabeth L. Mann and Sallie W. Chisholm. Iron limits the cell division rate of Prochlorococcus in the eastern equatorial Pacific. Limnology and Oceanography, 45(5):1067–1076, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2001:SNC

[MC01] Margaret R. Mulholland and Douglas G. Capone. Stoichiometry of nitrogen and carbon utilization in cultured populations of Trichodesmium IMS101: Implications for growth. Limnology and Oceanography, 49(1):436–443, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Mortazavi:2004:UKP

[MC04] Behzad Mortazavi and Jeffrey P. Chanton. Use of Keeling plots to determine sources of dissolved organic carbon in nearshore and open ocean systems. Limnology and Oceanography, 49(1): 102–108, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Meseck:2006:EBC

[MC06] Shannon L. Meseck and Gregory A. Cutter. Evaluating the biogeochemical cycle of selenium in San Francisco Bay through modeling. Limnology and Oceanography, 51(5):2018–2032, September 2006. CODEN LIOCAH. ISSN 0024-3590.

McArdle:2003:LME

[McA03] Brian H. McArdle. Lines, models, and errors: Regression in the field. Limnology and Oceanography, 48(4):1363–1366, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Molinero:2008:IAR

[MCB08] Juan Carlos Molinero, Michele Casini, and Emmanuelle Buecher. The influence of the Atlantic and regional climate variability on the long-term changes in gelatinous carnivore populations in the northwestern Mediterranean. Limnology and Oceanography, 53(4):1456–1467, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Montoya:2002:NFN

[MCC02] Joseph P. Montoya, Edward J. Carpenter, and Douglas G. Capone. Nitrogen fixation and nitrogen isotope abundances in zooplankton of the oligotrophic North Atlantic. Limnology and Oceanography, 47(6):1617–1628, November 2002. CODEN LI-OCAH. ISSN 0024-3590.

Maranon:2004:SMP

[MCF^{+}04] Emilio Marañón, Pedro Cermeño, Emilio Fernández, Jaime Rodríguez, and Laura Zabala. Significance and mechanisms of photosynthetic production of dissolved organic carbon in a coastal eutrophic ecosystem. Limnology and Oceanography, 49 (5):1652–1666, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Martin:2006:TCE

[MCJ^{+}06] Jonathan B. Martin, Jaye E. Cable, John Jaeger, Kevin Hartl, and Christopher G. Smith. Thermal and chemical evidence for rapid water exchange across the sediment–water interface by bioirrigation in the Indian River Lagoon, Florida. Limnology and Oceanography, 51(3):1332–1341, May 2006. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:2009:TMI

[MCJF09] Sally MacIntyre, Jordan F. Clark, Robert Jellison, and Jonathan P. Fram. Turbulent mixing induced by nonlinear internal waves in Mono Lake, California. Limnology and Oceanography, 54(6):2255–2272, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Menesguen:2006:NNT

[MCL06] Alain Ménesguen, Philippe Cugier, and Isabelle Leblond. A new numerical technique for tracking chemical species in a multisource, coastal ecosystem, applied to nitrogen causing Ulva blooms in the Bay of Brest (France). Limnology and Oceanography, 51(1 (part 2)):591–601, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Monsen:2002:CUF

[MCLM02] Nancy E. Monsen, James E. Cloern, Lisa V. Lucas, and Stephen G. Monismith. A comment on the use of flushing time, residence time, and age as transport time scales. Limnology and Oceanography, 47(5):1545–1553, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

McManus:2000:BGM

[McM00] George B. McManus. Biogeochemistry of Gulf of Mexico estuaries. Limnology and Oceanography, 60(3):524, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Mourino-Carballido:2006:MVM

[MCM06] Beatriz Mouriño-Carballido and Dennis J. McGillicuddy, Jr. Mesoscale variability in the metabolic balance of the Sargasso Sea. Limnology and Oceanography, 51(6):2675–2689, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Maranger:2006:SEM

[MCPP06a] Roxane Maranger, Charles D. Canham, Michael L. Pace, and Michael J. Papaik. A spatially explicit model of iron loading to lakes. Limnology and Oceanography, 51(1):247–256, January 2006. CODEN LIOCAH. ISSN 0024-3590.

McLaughlin:2006:TSI

[MCPP06b] Karen McLaughlin, Francisco Chavez, J. Timothy Pennington, and Adina Paytan. A time series investigation of the oxygen isotopic composition of dissolved inorganic phosphate in Monterey Bay, California. Limnology and Oceanography, 51(5):2370–2379, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Maraoon:2007:SPP

[MCR^{+}07] Emilio Maraóón, Pedro Cermeóo, Jaime Rodríguez, Mikhail V. Zubkov, and Roger P. Harris. Scaling of phytoplankton photosynthesis and cell size in the ocean. Limnology and Oceanography, 52(5):2190–2198, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2005:SSS

[MCV^{+}05] Manuel Maldonado, Ma Carmen Carmona, Zoila Velásquez, Angels Puig, Antonio Cruzado, Angel López, and Craig M. Young. Siliceous sponges as a silicon sink: an overlooked aspect of benthopelagic coupling in the marine silicon cycle. Limnology and Oceanography, 50(3):799–809, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Martin-Creuzburg:2008:NCC

[MCvEH08] Dominik Martin-Creuzburg, Eric von Elert, and Klaus H. Hoffmann. Nutritional constraints at the cyanobacteria–Daphnia magna interface: The role of sterols. Limnology and Oceanography, 53(2):456–468, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Mei:2004:VPL

[MD04] M. L. Mei and R. Danovaro. Virus production and life strategies in aquatic sediments. Limnology and Oceanography, 49(2):459– 470, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Manca:2009:RIP

[MD09] Marina Manca and William R. DeMott. Response of the invertebrate predator bythotrephes to a climate-linked increase in the duration of a refuge from fish predation. Limnology and Oceanography, 54(6 (part 2)):2506–2512, November 2009. CO-DEN LIOCAH. ISSN 0024-3590.

McCallister:2008:DMS

[MdG08] S. Leigh McCallister and Paul A. del Giorgio. Direct measurement of the δ ^{13}C signature of carbon respired by bacteria in lakes: Linkages to potential carbon sources, ecosystem baseline metabolism, and CO_{2} fluxes. Limnology and Oceanography, 53 (4):1204–1216, July 2008. CODEN LIOCAH. ISSN 0024-3590.

MacNeil:2004:SIM

[MDJ^{+}04] Calum MacNeil, Jaimie T. A. Dick, Mark P. Johnson, Melanie J. Hatcher, and Alison M. Dunn. A species invasion mediated through habitat structure, intraguild predation, and parasitism. Limnology and Oceanography, 49(5):1848–1856, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Menden-Deuer:2000:CVR

[MDL00] Susanne Menden-Deuer and Evelyn J. Lessard. Carbon to volume relationships for dinoflagellates, diatoms, and other protist plankton. Limnology and Oceanography, 45(3):569–579, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Matthews:2006:LTC

[ME06] David A. Matthews and Steven W. Effler. Long-term changes in the areal hypolimnetic oxygen deficit (AHOD) of Onondaga Lake: Evidence of sediment feedback. Limnology and Oceanography, 51(1 (part 2)):702–714, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Mollenhauer:2007:DSC

[ME07] Gesine Mollenhauer and Timothy I. Eglinton. Diagenetic and sedimentological controls on the composition of organic matter preserved in California Borderland Basin sediments. Limnology and Oceanography, 52(2):558–576, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Matthews:2008:EBH

[MED^{+}08] David A. Matthews, Steven W. Effler, Charles T. Driscoll, Susan M. O’Donnell, and Carol M. Matthews. Electron budgets for the hypolimnion of a recovering urban lake, 1989–2004: Response to changes in organic carbon deposition and availability of electron acceptors. Limnology and Oceanography, 53(2):743– 759, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Metaxas:2004:STP

[Met04] Anna Metaxas. Spatial and temporal patterns in larval supply at hydrothermal vents in the northeast Pacific Ocean. Limnology and Oceanography, 49(6):1949–1956, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Montagnes:2001:ETD

[MF01] David J. S. Montagnes and Montagnes Franklin. Effect of temperature on diatom volume, growth rate, and carbon and nitrogen content: Reconsidering some paradigms. Limnology and Oceanography, 46(8):2008–2018, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Montagnes:2002:ETD

[MF02] David J. S. Montagnes and Daniel J. Franklin. Effect of temperature on diatom volume, growth rate, and carbon and nitrogen content: Reconsidering some paradigms. Limnology and Oceanography, 47(4):1272, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Mazel:2003:CFS

[MF03] Charles H. Mazel and Eran Fuchs. Contribution of fluorescence to the spectral signature and perceived color of corals. Limnology and Oceanography, 48(1 (part 2)):390–401, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Monismith:2004:NPT

[MF04] Stephen G. Monismith and Derek A. Fong. A note on the potential transport of scalars and organisms by surface waves. Limnology and Oceanography, 49(4):1214–1217, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Mitra:2007:AGD

[MFF07] Aditee Mitra, Kevin J. Flynn, and Michael J. R. Fasham. Accounting for grazing dynamics in nitrogen-phytoplankton-zooplankton (NPZ) models. Limnology and Oceanography, 52 (2):649–661, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Meyer:2009:PGD

[MFG^{+}09] Bettina Meyer, Veronica Fuentes, Citlali Guerra, Katrin Schmidt, Angus Atkinson, Susanne Spahic, Boris Cisewski, Ulrich Freier, Alejandro Olariaga, and Ulrich Bathmann. Physiology, growth, and development of larval krill Euphausia superba in autumn and winter in the Lazarev Sea, Antarctica. Limnology and Oceanography, 54(5):1595–1614, September 2009. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:2009:CRV

[MFK^{+}09] Sally MacIntyre, Jonathan P. Fram, Paul J. Kushner, Neil D. Bettez, W. J. O’Brien, J. E. Hobbie, and George W. Kling.

Climate-related variations in mixing dynamics in an Alaskan Arctic lake. Limnology and Oceanography, 54(6 (part 2)):2401– 2417, November 2009. CODEN LIOCAH. ISSN 0024-3590.

McQuoid:2004:RCP

[MG04] Melissa R. McQuoid and Anna Godhe. Recruitment of coastal planktonic diatoms from benthic versus pelagic cells: Variations in bloom development and species composition. Limnology and Oceanography, 49(4):1123–1133, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Manning:2008:DPF

[MG08] Mackenzie M. Manning and Ruth D. Gates. Diversity in populations of free-living Symbiodinium from a Caribbean and Pacific reef. Limnology and Oceanography, 53(5):1853–1861, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Morel:2007:OPC

[MGC^{+}07] André Morel, Bernard Gentili, Hervé Claustre, Marcel Babin, Annick Bricaud, Joséphine Ras, and Fanny Tièche. Optical properties of the “clearest” natural waters. Limnology and Oceanography, 52(1):217–229, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Middelboe:2003:DVB

[MGF03] Mathias Middelboe, Ronnie N. Glud, and Kai Finster. Distribution of viruses and bacteria in relation to diagenetic activity in an estuarine sediment. Limnology and Oceanography, 48(4): 1447–1456, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Meysman:2006:BPS

[MGGM06] Filip J. R. Meysman, Oleksiy S. Galaktionov, Britta Gribsholt, and Jack J. Middelburg. Bioirrigation in permeable sediments: Advective pore-water transport induced by burrow ventilation. Limnology and Oceanography, 51(1):142–156, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Malkin:2008:MGR

[MGH08] Sairah Y. Malkin, Stephanie J. Guildford, and Robert E. Hecky. Modeling the growth response of Cladophora in a Laurentian Great Lake to the exotic invader Dreissena and to lake warming. Limnology and Oceanography, 54(3):1111–1124, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2002:PHA

[MGL02] Margaret R. Mulholland, Christopher J. Gobler, and Cindy Lee. Peptide hydrolysis, amino acid oxidation, and nitrogen uptake in communities seasonally dominated by Aureococcus anophagefferens. Limnology and Oceanography, 47(4):1094–1108, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Moosmann:2006:PRA

[MGMW06] Lorenz Moosmann, René Gächter, Beat Müller, and Alfred Wüest. Is phosphorus retention in autochthonous lake sediments controlled by oxygen or phosphorus? Limnology and Oceanography, 51(1 (part 2)):763–771, January 2006. CODEN LIOCAH. ISSN 0024-3590.

McQuatters-Gollop:2007:LTC

[MGRE^{+}07] Abigail McQuatters-Gollop, Dionysios E. Raitsos, Martin Edwards, Yaswant Pradhan, Laurence D. Mee, Samantha J. Lavender, and Martin J. Attrill. A long-term chlorophyll dataset reveals regime shift in North Sea phytoplankton biomass unconnected to nutrient levels. Limnology and Oceanography, 52(2): 635–648, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Moustaka-Gouni:2006:PFW

[MGVM^{+}06] Maria Moustaka-Gouni, Elisabeth Vardaka, Evangelia Michaloudi, Konstantinos Ar. Kormas, Eleni Tryfon, Helen Mihalatou, Spyros Gkelis, and Tom Lanaras. Plankton food web structure in a eutrophic polymictic lake with a history of toxic cyanobacterial blooms. Limnology and Oceanography, 51(1 (part 2)):715–727, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Moeller:2005:DAP

[MGW^{+}05] Robert E. Moeller, Shawna Gilroy, Craig E. Williamson, Gabriella Grad, and Ruben Sommaruga. Dietary acquisition of photoprotective compounds (mycosporine-like amino acids, carotenoids) and acclimation to ultraviolet radiation in a freshwater copepod. Limnology and Oceanography, 50(2):427–439, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Mannino:2000:BCP

[MH00] Antonio Mannino and H. Rodger Harvey. Biochemical composition of particles and dissolved organic matter along an estuarine gradient: Sources and implications for DOM reactivity. Limnology and Oceanography, 45(4):775–788, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Mannino:2004:BCE

[MH04a] Antonio Mannino and H. Rodger Harvey. Black carbon in estuarine and coastal ocean dissolved organic matter. Limnology and Oceanography, 49(3):735–740, May 2004. CODEN LIOCAH. ISSN 0024-3590.

McBride:2004:MMS

[MH04b] Richard S. McBride and Andrij Z. Horodysky. Mechanisms maintaining sympatric distributions of two ladyfish (Elopidae: Elops) morphs in the Gulf of Mexico and western North Atlantic Ocean. Limnology and Oceanography, 49(4):1173–1181, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Marchetti:2007:CCC

[MH07] Adrian Marchetti and Paul J. Harrison. Coupled changes in the cell morphology and elemental (C, N, and Si) composition of the pennate diatom Pseudo-nitzschia due to iron deficiency. Limnology and Oceanography, 52(5):2270–2284, September 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Martinez:2001:IND

[MHB01] J. S. Martinez, M. G. Haygood, and Alison Butler. Identification of a natural desferrioxamine siderophore produced by a marine bacterium. Limnology and Oceanography, 49(1):420–424, March 2001. CODEN LIOCAH. ISSN 0024-3590.

McManus:2003:HOR

[MHB03] James McManus, Erik A. Heinen, and Matthew M. Baehr. Hypolimnetic oxidation rates in Lake Superior: Role of dissolved organic material on the lake’s carbon budget. Limnology and Oceanography, 48(4):1624–1632, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Michaud:2005:ERC

[MHC05] Annick L. Michaud, Landis Hare, and Peter G. C. Campbell. Exchange rates of cadmium between a burrowing mayfly and its surroundings in nature. Limnology and Oceanography, 50(6): 1707–1717, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Marshall:2000:RAA

[MHDH00] Rhoda A. Marshall, John T. G. Hamilton, Matthew J. Dring, and David B. Harper. The red alga Asparagopsis taxiformi/ Falkenbergia hillebrandii — a possible source of trichloroethylene and perchloroethylene? Limnology and Oceanography, 60(3): 516–519, March 2000. CODEN LIOCAH. ISSN 0024-3590. See reply [AP00a].

Mathis:2007:DND

[MHK^{+}07] Jeremy T. Mathis, Dennis A. Hansell, David Kadko, Nicholas R. Bates, and Lee W. Cooper. Determining net dissolved organic carbon production in the hydrographically complex western Arctic Ocean. Limnology and Oceanography, 52(5):1789– 1799, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Moisander:2002:SSS

[MHKP02] Pia H. Moisander, James L. Hench, Kaisa Kononen, and Hans W. Paerl. Small-scale shear effects on heterocystous cyanobacteria. Limnology and Oceanography, 47(1):108–119, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Malinen:2001:LCD

[MHLN01] Tommi Malinen, Jukka Horppila, and Anne Liljendahl-Nurminen. Langmuir circulations disturb the low-oxygen refuge of phantom midge larvae. Limnology and Oceanography, 46(6): 689–692, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2002:EFC

[MHRW02] Maria T. Maldonado, Margaret P. Hughes, Eden L. Rue, and Mark L. Wells. The effect of Fe and Cu on growth and domoic acid production by Pseudo-nitzschia multiseries and Pseudonitzschia australis. Limnology and Oceanography, 47(2):515– 526, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2009:NRS

[MHS^{+}09] Patrick J. Mulholland, Robert O. Hall, Jr., Daniel J. Sobota, Walter K. Dodds, Stuart E. G. Findlay, Nancy B. Grimm, Stephen K. Hamilton, William H. McDowell, Jonathan M. O’Brien, Jennifer L. Tank, Linda R. Ashkenas, Lee W. Cooper, Clifford N. Dahm, Stanley V. Gregory, Sherri L. Johnson, Judy L. Meyer, Bruce J. Peterson, Geoffrey C. Poole, H. Maurice Valett, Jackson R. Webster, Clay P. Arango, Jake J. Beaulieu, Melody J. Bernot, Amy J. Burgin, Chelsea L. Crenshaw, Ashley M. Helton, Laura T. Johnson, B. R. Niederlehner, Jody D. Potter, Richard W. Sheibley, and Suzanne M. Thomasn. Nitrate removal in stream ecosystems measured by ^{15}N addition experiments: Denitrification. Limnology and Oceanography, 54 (3):666–680, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Mehner:2007:RIF

[MID^{+}07] Thomas Mehner, Jan Ihlau, Hendrik Dörner, Michael Hupfer, and Franz Hölker. The role of insectivorous fish in fostering the allochthonous subsidy of lakes. Limnology and Oceanography, 52 (6):2718–2721, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Mehner:2005:CFF

[MIDH05] Thomas Mehner, Jan Ihlau, Hendrik Dörner, and Franz Hölker. Can feeding of fish on terrestrial insects subsidize the nutrient pool of lakes? Limnology and Oceanography, 50(6):2022–2031, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Miller:2002:BRA

[Mil02] Douglas C. Miller. Book review: Aller, J. Y., S. A. Woodin, and R. C. Aller [Eds.]. 2001. Organism–sediment interactions. University of South Carolina Press. xxi + 403 p. US$60. ISBN 1-57003-431-1. Limnology and Oceanography, 47(6):1856, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Molinero:2005:NAC

[MIN^{+}05] Juan Carlos Molinero, Frédéric Ibanez, Paul Nival, Emmanuelle Buecher, and Sami Souissi. North Atlantic climate and northwestern Mediterranean plankton variability. Limnology and Oceanography, 50(4):1213–1220, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Meding:2003:BCM

[MJ03] Marianne E. Meding and Leland J. Jackson. Biotic, chemical, and morphometric factors contributing to winter anoxia in prairie lakes. Limnology and Oceanography, 48(4):1633–1642, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Mallin:2006:FCH

[MJEM06] Michael A. Mallin, Virginia L. Johnson, Scott H. Ensign, and Tara A. MacPherson. Factors contributing to hypoxia in rivers, lakes, and streams. Limnology and Oceanography, 51(1 (part 2)): 690–701, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Martz:2008:OMO

[MJR08] Todd R. Martz, Kenneth S. Johnson, and Stephen C. Riser. Ocean metabolism observed with oxygen sensors on profiling floats in the South Pacific. Limnology and Oceanography, 53(5 (part 2)):2094–2111, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Muller:2003:BFR

[MJW03] François L. L. Muller, Stéphan Jacquet, and William H. Wilson. Biological factors regulating the chemical speciation of Cu, Zn, and Mn under different nutrient regimes in a marine mesocosm experiment. Limnology and Oceanography, 48(6):2289– 2302, November 2003. CODEN LIOCAH. ISSN 0024-3590.

MacKenzie:2000:LFF

[MK00] Brian R. MacKenzie and Thomas KiØrboe. Larval fish feeding and turbulence: a case for the downside. Limnology and Oceanography, 45(1):1–10, January 2000. CODEN LIOCAH. ISSN 0024-3590.

McNaught:2004:CZC

[MKG04] A. Scott McNaught, Richard L. Kiesling, and Anas Ghadouani. Changes to zooplankton community structure following colonization of a small lake by Leptodora kindti. Limnology and Oceanography, 49(4 (part 2)):1239–1249, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Malmstrom:2004:IEB

[MKK04] Rex R. Malmstrom, Ronald P. Kiene, and David L. Kirchman. Identification and enumeration of bacteria assimilating dimethylsulfoniopropionate (DMSP) in the North Atlantic and Gulf of Mexico. Limnology and Oceanography, 49(2):597–606, March 2004. CODEN LIOCAH. ISSN 0024-3590.

McDonald:2006:EDC

[MKM06] C. B. McDonald, J. R. Koseff, and S. G. Monismith. Effects of the depth to coral height ratio on drag coefficients for unidirectional flow over coral. Limnology and Oceanography, 51(3): 1294–1301, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Morel:2008:RUF

[MKS08] F. M. M. Morel, A. B. Kustka, and Y. Shaked. The role of unchelated Fe in the iron nutrition of phytoplankton. Limnology and Oceanography, 53(1):400–404, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Meile:2001:QBA

[MKV01] Christof Meile, Carla M. Koretsky, and Philippe Van Cappellen. Quantifying bioirrigation in aquatic sediments: an inverse modeling approach. Limnology and Oceanography, 46(1):164–177, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Malmstrom:2005:DDA

[MKVK05] Rex R. Malmstrom, Ronald P. Kiene, Maria Vila, and David L. Kirchman. Dimethylsulfoniopropionate (DMSP) assimilation by Synechococcus in the Gulf of Mexico and northwest Atlantic Ocean. Limnology and Oceanography, 50(6):1924–1931, November 2005. CODEN LIOCAH. ISSN 0024-3590.

McCutchan:2002:RIC

[ML02] James H. McCutchan, Jr. and William M. Lewis, Jr. Relative importance of carbon sources for macroinvertebrates in a Rocky Mountain stream. Limnology and Oceanography, 47(3):742–752, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Mitchell:2005:RDN

[ML05] Nicole L. Mitchell and Gary A. Lamberti. Responses in dissolved nutrients and epilithon abundance to spawning salmon in southeast Alaska streams. Limnology and Oceanography, 55(3): 217–227, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2009:PHU

[ML09] Margaret R. Mulholland and Cindy Lee. Peptide hydrolysis and the uptake of dipeptides by phytoplankton. Limnology and Oceanography, 54(3):856–868, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Mackey:2007:PAP

[MLC^{+}07] Katherine R. M. Mackey, Rochelle G. Labiosa, Mike Calhoun, Joseph H. Street, Anton F. Post, and Adina Paytan. Phosphorus availability, phytoplankton community dynamics, and taxonspecific phosphorus status in the Gulf of Aqaba, Red Sea. Limnology and Oceanography, 52(2):873–885, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Mazel:2003:GFP

[MLG^{+}03] Charles H. Mazel, Michael P. Lesser, Maxim Y. Gorbunov, Thomas M. Barry, Julianne H. Farrell, Kevin D. Wyman, and Paul G. Falkowski. Green-fluorescent proteins in Caribbean corals. Limnology and Oceanography, 48(1 (part 2)):402–411, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Moschen:2009:CSI

[MLPS09] Robert Moschen, Andreas Lücke, Jörn Parplies, and Gerhard H. Schleser. Controls on the seasonal and interannual dynamics of organic matter stable carbon isotopes in mesotrophic Lake Holzmaar, Germany. Limnology and Oceanography, 54(1):194– 209, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Mladenov:2009:ALO

[MLRMR09] N. Mladenov, J. Lípez-Ramos, D. M. McKnight, and I. Rechea. Alpine lake optical properties as sentinels of dust deposition and global change. Limnology and Oceanography, 54(6 (part 2)): 2386–2400, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Muller:2005:IBA

[MLSS05] François L. L. Muller, Aud Larsen, Colin A. Stedmon, and Morten Søndergaard. Interactions between algal-bacterial populations and trace metals in fjord surface waters during a nutrientstimulated summer bloom. Limnology and Oceanography, 50(6): 1855–1871, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Matthews:2003:CIV

[MM03] Blake Matthews and Asit Mazumder. Compositional and interlake variability of zooplankton affect baseline stable isotope signatures. Limnology and Oceanography, 49(4):1977–1987, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Matthews:2005:CLT

[MM05] Blake Matthews and Asit Mazumder. Consequences of large temporal variability of zooplankton δ ^{15}N for modeling fish trophic position and variation. Limnology and Oceanography, 50(5):1404–1414, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Miller:2009:MDP

[MMCW09] Matthew P. Miller, Diane M. McKnight, Steven C. Chapra, and Mark W. Williams. A model of degradation and production of three pools of dissolved organic matter in an alpine lake. Limnology and Oceanography, 54(6):2213–2227, November 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Maerki:2009:MPL

[MMDW09] Martin Maerki, Beat Müller, Christian Dinkel, and Bernhard Wehrli. Mineralization pathways in lake sediments with different oxygen and organic carbon supply. Limnology and Oceanography, 54(2):428–438, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Moore:2008:LLS

[MMKN08] Stephanie K. Moore, Nathan J. Mantua, Jonathan P. Kellogg, and Jan A. Newton. Local and large-scale climate forcing of Puget Sound oceanographic properties on seasonal to interdecadal timescales. Limnology and Oceanography, 53(5):1746– 1758, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Mills:2008:NPC

[MML^{+}08a] M. M. Mills, C. M. Moore, R. Langlois, A. Milne, E. Achterberg, K. Nachtigall, K. Lochte, R. J. Geider, and Roche J. La. Nitrogen and phosphorus co-limitation of bacterial productivity and growth in the oligotrophic subtropical North Atlantic. Limnology and Oceanography, 53(2):824–834, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Moore:2008:RIN

[MML^{+}08b] C. Mark Moore, Matthew M. Mills, Rebecca Langlois, Angela Milne, Eric P. Achterberg, Julie La Roche, and Richard J. Geider. Relative influence of nitrogen and phosphorous availability on phytoplankton physiology and productivity in the oligotrophic sub-tropical North Atlantic Ocean. Limnology and Oceanography, 53(1):291–305, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Marchetti:2006:IRP

[MMLH06] Adrian Marchetti, Maria T. Maldonado, Erin S. Lane, and Paul J. Harrison. Iron requirements of the pennate diatom Pseudo-nitzschia: Comparison of oceanic (high-nitrate, lowchlorophyll waters) and coastal species. Limnology and Oceanography, 51(5):2092–2101, September 2006. CODEN LIOCAH. ISSN 0024-3590.

McIntyre:2006:TBC

[MMO06] Peter B. McIntyre, Ellinor Michel, and Michelle Olsgard. Topdown and bottom-up controls on periphyton biomass and productivity in Lake Tanganyika. Limnology and Oceanography, 51 (3):1514–1523, May 2006. CODEN LIOCAH. ISSN 0024-3590.

McFadden:2004:PEL

[MMPA04] Melany A. McFadden, Henry T. Mullins, William P. Patterson, and William T. Anderson. Paleoproductivity of eastern Lake Ontario over the past 10,000 years. Limnology and Oceanography, 49(5):1570–1581, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Miller:2002:DAQ

[MMS^{+}02] William L. Miller, Miller Moran, Wade M. Sheldon, Richard G. Zepp, and Stephen Opsahl. Determination of apparent quantum yield spectra for the formation of biologically labile photoproducts. Limnology and Oceanography, 47(2):343–352, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Murphy:2000:ILM

[MMW00] Cormac D. Murphy, Robert M. Moore, and Robert L. White. An isotopic labeling method for determining production of volatile organohalogens by marine microalgae. Limnology and Oceanography, 45(8):1868–1871, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Maerki:2006:MMP

[MMW06] Martin Maerki, Beat Müller, and Bernhard Wehrli. Microscale mineralization pathways in surface sediments: a chemical sensor study in Lake Baikal. Limnology and Oceanography, 51(3):1342– 1354, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Morrison:2004:SCP

[MN04] J. Ruairidh Morrison and Norman B. Nelson. Seasonal cycle of phytoplankton UV absorption at the Bermuda Atlantic timeseries study (BATS) site. Limnology and Oceanography, 49(1): 215–224, January 2004. CODEN LIOCAH. ISSN 0024-3590.

MacKay:2009:MLR

[MNA^{+}09] M. D. MacKay, P. J. Neale, C. D. Arp, L. N. De Senerpont Domis, X. Fang, G. Gal, K. D. Jöhnk, G. Kirillin, J. D. Lenters, E. Litchman, S. MacIntyre, P. Marsh, J. Melack, W. M. Mooij, F. Peeters, A. Quesada, S. G. Schladow, M. Schmid, C. Spence, and S. L. Stokes. Modeling lakes and reservoirs in the climate system. Limnology and Oceanography, 54(6 (part 2)):2315–2329, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Motegi:2009:VCB

[MNM^{+}09] Chiaki Motegi, Toshi Nagata, Takeshi Miki, Markus G. Weinbauer, Louis Legendre, and Fereidoun Rassoulzadegan. Viral control of bacterial growth efficiency in marine pelagic environments. Limnology and Oceanography, 54(6):1901–1910, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Maar:2003:MDZ

[MNSV03] Marie Maar, Torkel Gissel Nielsen, Adolf Stips, and AndréW. Visser. Microscale distribution of zooplankton in relation to turbulent diffusion. Limnology and Oceanography, 48(4):1312– 1325, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Markaki:2003:ADI

[MOK^{+}03] Z. Markaki, K. Oikonomou, M. Kocak, G. Kouvarakis, A. Chaniotaki, N. Kubilay, and N. Mihalopoulos. Atmospheric deposition of inorganic phosphorus in the Levantine Basin, eastern Mediterranean: Spatial and temporal variability and its role in seawater productivity. Limnology and Oceanography, 48(4): 1557–1568, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Moller:2007:PDO

[Møl07] Eva Friis Møller. Production of dissolved organic carbon by sloppy feeding in the copepods Acartia tonsa, Centropages typicus, and Temora longicornis. Limnology and Oceanography, 52 (1):79–84, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Morrison:2003:SDQ

[Mor03] J. Ruairidh Morrison. In situ determination of the quantum yield of phytoplankton chlorophyll-a fluorescence: a simple algorithm, observations, and a model. Limnology and Oceanography, 48(2): 618–631, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Mortimer:2006:IOR

[Mor06] C. H. Mortimer. Inertial oscillations and related internal beat pulsations and surges in lakes Michigan and Ontario. Limnology and Oceanography, 51(5):1941–1955, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2000:NRF

[MP00] Maria T. Maldonado and Neil M. Price. Nitrate regulation of Fe reduction and transport by Fe-limited Thalassiosira oceanica. Limnology and Oceanography, 45(4):814–826, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Matzinger:2007:EIN

[MPA^{+}07] Andreas Matzinger, Roger Pieters, Ken I. Ashley, Gregory A. Lawrence, and Alfred Wüest. Effects of impoundment on nutrient availability and productivity in lakes. Limnology and Oceanography, 52(6):2629–2640, November 2007. CODEN LI-OCAH. ISSN 0024-3590.

Massana:2001:CMB

[MPACG01] Ramon Massana, Carlos Pedrós-Alió, Emilio O. Casamayor, and Josep M. Gasol. Changes in marine bacterioplankton phylogenetic composition during incubations designed to measure biogeochemically significant parameters. Limnology and Oceanography, 46(5):1181–1188, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Masson:2000:TPC

[MPAS00] Stéphane Masson, Bernadette Pinel-Alloul, and Val H. Smith. Total phosphorus–chlorophyll-a size fraction relationships in southern Québec lakes. Limnology and Oceanography, 45(3): 732–740, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Morris:2008:IBH

[MPB^{+}08] E. P. Morris, G. Peralta, F. G. Brun, L. van Duren, T. J. Bouma, and J. L. Perez-Llorens. Interaction between hydrodynamics and seagrass canopy structure: Spatially explicit effects on ammonium uptake rates. Limnology and Oceanography, 53(4):1531– 1539, July 2008. CODEN LIOCAH. ISSN 0024-3590. See erratum [Ano08a].

Minor:2006:ESC

[MPD^{+}06] E. C. Minor, J. Pothen, B. J. Dalzell, H. Abdulla, and K. Mopper. Effects of salinity changes on the photodegradation and ultraviolet-visible absorbance of terrestrial dissolved organic matter. Limnology and Oceanography, 51(5):2181–2186, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Mouradian:2007:DDE

[MPdVG07] Maggy Mouradian, Robert J. Panetta, Anne de Vernal, and Yves Gélinas. Dinosterols or dinocysts to estimate dinoflagellate contributions to marine sedimentary organic matter? Limnology and Oceanography, 52(6):2569–2581, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Mackey:2008:PSC

[MPGB08] Katherine R. M. Mackey, Adina Paytan, Arthur R. Grossman, and Shaun Bailey. A photosynthetic strategy for coping in a high-light, low-nutrient environment. Limnology and Oceanography, 53(3):900–913, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Martin:2008:EMC

[MPH08] Sylvain Martin, Isabelle Proulx, and Landis Hare. Explaining metal concentrations in sympatric Chironomus species. Limnology and Oceanography, 53(2):411–419, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Maie:2008:MTA

[MPJ08] Nagamitsu Maie, Oliva Pisani, and Rudolf Jaffé. Mangrove tannins in aquatic ecosystems: Their fate and possible influence on dissolved organic carbon and nitrogen cycling. Limnology and Oceanography, 53(1):160–171, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Moore:2002:UDN

[MPRC02] Lisa R. Moore, Anton F. Post, Gabrielle Rocap, and Sallie W. Chisholm. Utilization of different nitrogen sources by the marine cyanobacteria Prochlorococcus and Synechococcus. Limnology and Oceanography, 47(4):989–996, July 2002. CODEN LI-OCAH. ISSN 0024-3590.

Moisander:2008:EIN

[MPZ08] Pia H. Moisander, Hans W. Paerl, and Jonathan P. Zehr. Effects of inorganic nitrogen on taxa-specific cyanobacterial growth and nifH expression in a subtropical estuary. Limnology and Oceanography, 53(6):2519–2532, November 2008. CODEN LI-OCAH. ISSN 0024-3590.

MacIntyre:2002:STV

[MRK02] Sally MacIntyre, José R. Romero, and George W. Kling. Spatialtemporal variability in surface layer deepening and lateral advection in an embayment of Lake Victoria, East Africa. Limnology and Oceanography, 47(3):656–671, May 2002. CODEN LIOCAH. ISSN 0024-3590.

McClatchie:2007:ISR

[MRM07] Sam McClatchie, Paul J. Rogers, and Lachie McLeay. Importance of scale to the relationship between abundance of sardine larvae, stability, and food. Limnology and Oceanography, 52(4): 1570–1579, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Mari:2007:WRT

[MRNT^{+}07] Xavier Mari, Emma Rochelle-Newall, Jean-Pascal Torréton, Olivier Pringault, Aymeric Jouon, and Christophe Migon. Water residence time: a regulatory factor of the DOM to POM transfer efficiency. Limnology and Oceanography, 52(2):808–819, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Mobley:2003:EOSa

[MS03] Curtis D. Mobley and Lydia K. Sundman. Effects of optically shallow bottoms on upwelling radiances: Inhomogeneous and sloping bottoms. Limnology and Oceanography, 48(1 (part 2)): 329–336, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Murase:2005:IEL

[MS05] Jun Murase and Atsuko Sugimoto. Inhibitory effect of light on methane oxidation in the pelagic water column of a mesotrophic lake (Lake Biwa, Japan). Limnology and Oceanography, 50(4): 1339–1343, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Martin:2006:LLP

[MS06] Sherry L. Martin and Patricia A. Soranno. Lake landscape position: Relationships to hydrologic connectivity and landscape features. Limnology and Oceanography, 51(2):801–814, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Manuel:2006:SRN

[MSAC06] Juan Manuel, Medina Sánchez, Manuel Villar Argaiz, and Presentaciín Carrillo. Solar radiation-nutrient interaction enhances the resource and predation algal control on bacterioplankton: a short-term experimental study. Limnology and Oceanography, 51 (2):913–924, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Moreno:2001:TAF

[MSBHG01] Guillermo Moreno, Paulina Selvakumaraswamy, Maria Byrne, and Ove Hoegh-Guldberg. A test of the ash-free dry weight technique on the developmental stages of Patiriella spp. (Echinodermata: Asteroidea). Limnology and Oceanography, 46(5): 1214–1220, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Meckler:2004:NOM

[MSC^{+}04] A. N. Meckler, C. J. Schubert, G. L. Cowie, S. Peiffer, and M. Dittrich. New organic matter degradation proxies: Valid in lake systems? Limnology and Oceanography, 49(6):2023–2033, November 2004. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:2006:PPN

[MSGK06] Sally MacIntyre, James O. Sickman, Sarah A. Goldthwait, and George W. Kling. Physical pathways of nutrient supply in a small, ultraoligotrophic Arctic lake during summer stratification. Limnology and Oceanography, 51(2):1107–1124, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Moore:2006:PPP

[MSH^{+}06] C. Mark Moore, David J. Suggett, Anna E. Hickman, Young-Nam Kim, Jacqueline F. Tweddle, Jonathan Sharples, Richard J. Geider, and Patrick M. Holligan. Phytoplankton photoacclimation and photoadaptation in response to environmental gradients in a shelf sea. Limnology and Oceanography, 51 (2):936–949, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Mayer:2009:POP

[MSHE09] Lawrence M. Mayer, Linda L. Schick, Kathleen R. Hardy, and Margaret L. Estapa. Photodissolution and other photochemical changes upon irradiation of algal detritus. Limnology and Oceanography, 54(5):1688–1698, September 2009. CODEN LI-OCAH. ISSN 0024-3590.

Muller-Solger:2002:NQF

[MSJMN02] Anke B. Müller-Solger, Alan D. Jassby, and Dörthe C. Müller-Navarra. Nutritional quality of food resources for zooplankton (Daphnia) in a tidal freshwater system (Sacramento–San Joaquin River Delta). Limnology and Oceanography, 47(5):1468– 1476, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Mazel:2003:HRD

[MSL^{+}03] Charles H. Mazel, Michael P. Strand, Michael P. Lesser, Michael P. Crosby, Bryan Coles, and Andrew J. Nevis. Highresolution determination of coral reef bottom cover from multispectral fluorescence laser line scan imagery. Limnology and Oceanography, 48(1 (part 2)):522–534, January 2003. CODEN LIOCAH. ISSN 0024-3590.

McCutchan:2002:EGF

[MSLH02] James H. McCutchan, Jr., James F. Saunders III, William. M. Lewis, Jr., and Matthew G. Hayden. Effects of groundwater flux on open-channel estimates of stream metabolism. Limnology and Oceanography, 47(1):321–324, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Moran:2006:RSO

[MSPAE06] Xosé Anxelu G. Morán, Marta Sebastián, Carlos Pedrís-Alií, and Marta Estrada. Response of Southern Ocean phytoplankton and bacterioplankton production to short-term experimental warming. Limnology and Oceanography, 51(5):1791–1800, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Mayer:2006:PPO

[MSSB06] Lawrence M. Mayer, Linda L. Schick, Krysia Skorko, and Emmanuel Boss. Photodissolution of particulate organic matter from sediments. Limnology and Oceanography, 51(2):1064–1071, March 2006. CODEN LIOCAH. ISSN 0024-3590.

McPhee-Shaw:2007:MND

[MSSW^{+}07] Erika E. McPhee-Shaw, David A. Siegel, Libe Washburn, Mark A. Brzezinski, Janice L. Jones, Al Leydecker, and John Melack. Mechanisms for nutrient delivery to the inner shelf: Observations from the Santa Barbara Channel. Limnology and Oceanography, 52(5):1748–1766, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Medina-Sanchez:2004:NNY

[MSVAC04] Juan Manuel Medina-Sánchez, Manuel Villar-Argaiz, and Presentación Carrillo. Neither with nor without you: a complex algal control on bacterioplankton in a high mountain lake. Limnology and Oceanography, 49(5):1722–1733, September 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Matzinger:2007:EAL

[MSVS^{+}07] Andreas Matzinger, Martin Schmid, Elizabeta Veljanoska-Sarafiloska, Suzana Patceva, Dafina Guseska, Bernd Wagner, Beat Müller, Michael Sturm, and Alfred Wüest. Eutrophication of ancient Lake Ohrid: Global warming amplifies detrimental effects of increased nutrient inputs. Limnology and Oceanography, 52(1):338–353, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Moran:2000:CLO

[MSZ00] Mary Ann Moran, Wade M. Sheldon, Jr., and Richard G. Zepp. Carbon loss and optical property changes during long-term photochemical and biological degradation of estuarine dissolved organic matter. Limnology and Oceanography, 45(6):1254–1264, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Martin:2006:BTE

[MT06] Cheryl L. Martin and Philippe D. Tortell. Bicarbonate transport and extracellular carbonic anhydrase activity in Bering Sea phytoplankton assemblages: Results from isotope disequilibrium experiments. Limnology and Oceanography, 51(5):2111–2121, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Martin:2006:RCE

[MTC^{+}06] Sophie Martin, Gérard Thouzeau, Laurent Chauvaud, Frédéric Jean, Laurent Guérin, and Jacques Clavier. Respiration, calcification, and excretion of the invasive slipper limpet, Crepidula fornicata L.: Implications for carbon, carbonate, and nitrogen fluxes in affected areas. Limnology and Oceanography, 51(5): 1996–2007, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Monaghan:2001:IFF

[MTM^{+}01] Michael T. Monaghan, Steven A. Thomas, G. Wayne Minshall, J. Denis Newbold, and Colbert E. Cushing. The influence of filter-feeding benthic macroinvertebrates on the transport and deposition of particulate organic matter and diatoms in two streams. Limnology and Oceanography, 46(5):1091–1099, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Moutin:2002:DCN

[MTV^{+}02] Thierry Moutin, Tron F. Thingstad, France Van Wambeke, Dominique Marie, Gerd Slawyk, Patrick Raimbault, and Herveé Claustre. Does competition for nanomolar phosphate supply explain the predominance of the cyanobacterium Synechococcus? Limnology and Oceanography, 47(5):1562–1567, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2001:ACN

[Mul01] Patrick J. Mulholland. Assessment and control of nonpoint source pollution of aquatic ecosystems. A practical approach. Limnology and Oceanography, 46(1):212, January 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Markager:2000:SLA

[MV00] Stiig Markager and Warwick F. Vincent. Spectral light attenuation and the absorption of UV and blue light in natural waters. Limnology and Oceanography, 45(3):642–650, May 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Meile:2003:GEE

[MV03] Christof Meile and Philippe Van Cappellen. Global estimates of enhanced solute transport in marine sediments. Limnology and Oceanography, 48(2):777–786, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Mueller:2009:HAL

[MVA^{+}09] Derek R. Mueller, Patrick. Van Hove, Dermot Antoniades, Martin O. Jeffries, and Warwick F. Vincent. High Arctic lakes as sentinel ecosystems: Cascading regime shifts in climate, ice cover, and mixing. Limnology and Oceanography, 54(6 (part 2)):2371– 2385, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Milligan:2004:DSM

[MVBM04] Allen J. Milligan, Diana E. Varela, Mark A. Brzezinski, and François M. M. Morel. Dynamics of silicon metabolism and silicon isotopic discrimination in a marine diatomas a function of pCO_{2}. Limnology and Oceanography, 49(2):322–329, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Morris:2005:TSR

[MVC^{+}05] Robert M. Morris, Kevin L. Vergin, Jang-Cheon Cho, Michael S. Rappé, Craig A. Carlson, and Stephen J. Giovannoni. Temporal and spatial response of bacterioplankton lineages to annual convective overturn at the Bermuda Atlantic Time-series Study site. Limnology and Oceanography, 50(5):1687–1696, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Mergeay:2005:CID

[MVD05] Joachim Mergeay, Dirk Verschuren, and Luc De Meester. Cryptic invasion and dispersal of an American Daphnia in East Africa. Limnology and Oceanography, 50(4):1278–1283, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Martineau:2004:PCP

[MVFD04] Christine Martineau, Warwick F. Vincent, Jean-Jacques Frenette, and Julian J. Dodson. Primary consumers and particulate organic matter: Isotopic evidence of strong selectivity in the estuarine transition zone. Limnology and Oceanography, 49(5):1679–1686, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2004:SDT

[MVW^{+}04] Patrick J. Mulholland, H. Maurice Valett, Jackson R. Webster, Steven A. Thomas, Lee W. Cooper, Stephen K. Hamilton, and Bruce J. Peterson. Stream denitrification and total nitrate uptake rates measured using a field ^{15}N tracer addition approach. Limnology and Oceanography, 49(3):809–820, May 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Marinelli:2002:EEL

[MW02] Roberta L. Marinelli and Sarah Ann Woodin. Experimental evidence for linkages between infaunal recruitment, disturbance, and sediment surface chemistry. Limnology and Oceanography, 47(1):221–229, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Marcarelli:2006:TNS

[MW06a] Amy M. Marcarelli and Wayne A. Wurtsbaugh. Temperature and nutrient supply interact to control nitrogen fixation in oligotrophic streams: an experimental examination. Limnology and Oceanography, 51(5):2278–2289, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Miao:2006:FIR

[MW06b] Ai-Jun Miao and Wen-Xiong Wang. Fulfilling iron requirements by a coastal diatom under different temperatures and irradiances. Limnology and Oceanography, 51(2):925–935, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Moran:2003:DDP

[MWE^{+}03] S. B. Moran, S. E. Weinstein, H. N. Edmonds, J. N. Smith, R. P. Kelly, M. E. Q. Pilson, and W. G. Harrison. Does ^{234}Th/ 238U disequilibrium provide an accurate record of the export flux of particulate organic carbon from the upper ocean? Limnology and Oceanography, 48(3):1018–1029, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Moran:2004:UVV

[MWE^{+}04] S. B. Moran, S. E. Weinstein, H. N. Edmonds, J. N. Smith, R. P. Kelly, M. E. Q. Pilson, and W. G. Harrison. Uncertainty versus variability in upper ocean carbon fluxes (reply to comment by Benitez-Nelson and Charette). Limnology and Oceanography, 49 (4):1221–1223, July 2004. CODEN LIOCAH. ISSN 0024-3590. See [BNC04].

Moeseneder:2001:HVC

[MWH01] Markus M. Moeseneder, Christian. Winter, and Gerhard J. Herndl. Horizontal and vertical complexity of attached and freeliving bacteria of the eastern Mediterranean Sea, determined by 16S rDNA and ^{16}S rRNA fingerprints. Limnology and Oceanography, 46(1):95–107, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Maguer:2004:NPR

[MWM^{+}04] Jean-François Maguer, Mohideen Wafar, Christian Madec, Pascal Morin, and Evelyne Erard-Le Denn. Nitrogen and phosphorus requirements of an Alexandrium minutum bloom in the Penzé Estuary, France. Limnology and Oceanography, 49(4): 1108–1114, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Muller:2006:CCH

[MWW06] Beat Müller, Ying Wang, and Bernhard Wehrli. Cycling of calcite in hard water lakes of different trophic states. Limnology and Oceanography, 52(2):1678–1688, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Mouw:2005:PPC

[MY05] Colleen B. Mouw and James A. Yoder. Primary production calculations in the Mid-Atlantic Bight, including effects of phytoplankton community size structure. Limnology and Oceanography, 50(4):1232–1243, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Maie:2005:CCD

[MYM^{+}05] Nagamitsu Maie, Chengyong Yang, Toshikazu Miyoshi, Kathleen Parish, and Rudolf Jaffé. Chemical characteristics of dissolved organic matter in an oligotrophic subtropical wetland/ estuarine. Limnology and Oceanography, 50(1):23–35, January 2005. CODEN LIOCAH. ISSN 0024-3590.

McManus:2004:MPC

[MZL04] George B. McManus, Huan Zhang, and Senjie Lin. Marine planktonic ciliates that prey on macroalgae and enslave their chloroplasts. Limnology and Oceanography, 49(1):308–313, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Mobley:2003:EOSb

[MZV03] Curtis D. Mobley, Hao Zhang, and Kenneth J. Voss. Effects of optically shallow bottoms on upwelling radiances: Bidirectional reflectance distribution function effects. Limnology and Oceanography, 48(1 (part 2)):337–345, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Neubauer:2003:TDI

[NA03] Scott C. Neubauer and Iris C. Anderson. Transport of dissolved inorganic carbon from a tidal freshwater marsh to the York River estuary. Limnology and Oceanography, 48(1):299–307, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Nishizaki:2005:SCC

[NA05] Michael T. Nishizaki and Josef Daniel Ackerman. A secondary chemical cue facilitates juvenile-adult postsettlement associations in Red Sea urchins. Limnology and Oceanography, 55(3): 354–362, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Nishihara:2006:EHM

[NA06] Gregory N. Nishihara and Josef D. Ackerman. The effect of hydrodynamics on the mass transfer of dissolved inorganic carbon to the freshwater macrophyte Vallisneria americana. Limnology and Oceanography, 52(6):2734–2745, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Nishihara:2009:DBL

[NA09] Gregory N. Nishihara and Josef D. Ackerman. Diffusive boundary layers do not limit the photosynthesis of the aquatic macrophyte, Vallisneria americana, at moderate flows and saturating light levels. Limnology and Oceanography, 54(6):1874–1882, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Nicholls:2008:HLO

[NAH^{+}08] Keith W. Nicholls, E. Povl Abrahamsen, Karen J. Heywood, Kate Stansfield, and Svein østerhus. High-latitude oceanography using the Autosub autonomous underwater vehicle. Limnology and Oceanography, 53(5 (part 2)):2309–2320, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Navarrete:2002:RII

[NBW^{+}02] Sergio A. Navarrete, Bernardo Broitman, Evie A. Wieters, Gerhard R. Finke, Roberto M. Venegas, and Alvaro Sotomayor. Recruitment of intertidal invertebrates in the southeast Pacific: Interannual variability and the 1997–1998 El Niño. Limnology and Oceanography, 47(3):791–802, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Nieto-Cid:2006:MPR

[NCÁSP06] M. Nieto-Cid, X. A. Álvarez-Salgado, and F. F. Pérez. Microbial and photochemical reactivity of fluorescent dissolved organic matter in a coastal upwelling system. Limnology and Oceanography, 51(3):1391–1400, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Ning:2000:STV

[NCC00] Xiuren Ning, James E. Cloern, and Brian E. Cole. Spatial and temporal variability of picocyanobacteria Synechococcus sp. in San Francisco Bay. Limnology and Oceanography, 45(3):695– 702, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Newell:2002:ISB

[NCO02] Roger I. E. Newell, Jeffrey C. Cornwell, and Michael S. Owens. Influence of simulated bivalve biodeposition and microphytobenthos on sediment nitrogen dynamics: a laboratory study. Limnology and Oceanography, 47(5):1367–1379, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Niewiadomska:2008:SPB

[NCPd08] Katarzyna Niewiadomska, Hervé Claustre, Louis Prieur, and Fabrizio d’Ortenzio. Submesoscale physical-biogeochemical coupling across the Ligurian Current (northwestern Mediterranean) using a bio-optical glider. Limnology and Oceanography, 53(5 (part 2)):2210–2225, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Nicholson:2006:APA

[NDCP06] David Nicholson, Sonya Dyhrman, Francisco Chavez, and Adina Paytan. Alkaline phosphatase activity in the phytoplankton communities of Monterey Bay and San Francisco Bay. Limnology and Oceanography, 51(2):874–883, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Nicholson:2008:AMA

[NDCP08] David Nicholson, Sonya Dyhrman, Francisco Chavez, and Adina Paytan. ASLO membership application 2008. Limnology and Oceanography, 51(4):874, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Nodder:2007:FPD

[NDP^{+}07] Scott D. Nodder, Gerard C. A. Duineveld, Conrad A. Pilditch, Philip J. Sutton, P. Keith Probert, Marc S. S. Lavaleye, Rob Witbaard, F. Hoe Chang, Julie A. Hall, and Ken M. Richardson. Focusing of phytodetritus deposition beneath a deep-ocean front, chatham rise, New Zealand. Limnology and Oceanography, 52(1):299–314, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Nicholls:2007:HRP

[NDT07] Joanna Claire Nicholls, Christian Andrew Davies, and Mark Trimmer. High-resolution profiles and nitrogen isotope tracing reveal a dominant source of nitrous oxide and multiple pathways of nitrogen gas formation in the central Arabian Sea. Limnology and Oceanography, 52(1):156–168, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Nusslein:2003:SIB

[NEC03] B. Nüsslein, W. Eckert, and R. Conrad. Stable isotope biogeochemistry of methane formation in profundal sediments of Lake Kinneret (Israel). Limnology and Oceanography, 48(4): 1439–1446, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Nicholson:2008:NCP

[NEE08] David Nicholson, Steven Emerson, and Charles C. Eriksen. Net community production in the deep euphotic zone of the subtropical North Pacific gyre from glider surveys. Limnology and Oceanography, 53(5 (part 2)):2226–2236, September 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Newell:2001:MPF

[New01] Steven Y. Newell. Multiyear patterns of fungal biomass dynamics and productivity within naturally decaying smooth cordgrass shoots. Limnology and Oceanography, 46(3):573–583, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Nagata:2000:BDP

[NFFK00] Toshi Nagata, Hideki Fukuda, Rumi Fukuda, and Isao Koike. Bacterioplankton distribution and production in deep Pacific waters: Large–scale geographic variations and possible coupling with sinking particle fluxes. Limnology and Oceanography, 45 (2):426–435, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Needoba:2007:NFU

[NFS^{+}07] Joseph A. Needoba, Rachel A. Foster, Carole Sakamoto, Jonathan P. Zehr, and Kenneth S. Johnson. Nitrogen fixation by unicellular diazotrophic cyanobacteria in the temperate oligotrophic North Pacific Ocean. Limnology and Oceanography, 52 (4):1317–1327, July 2007. CODEN LIOCAH. ISSN 0024-3590.

North:2007:EPN

[NGS^{+}07] R. L. North, S. J. Guildford, R. E. H. Smith, S. M. Havens, and M. R. Twiss. Evidence for phosphorus, nitrogen, and iron colimitation of phytoplankton communities in Lake Erie. Limnology and Oceanography, 52(1):315–328, January 2007. CODEN LI-OCAH. ISSN 0024-3590.

Nadon:2006:SIS

[NH06a] Marc-Olivier Nadon and John H. Himmelman. Stable isotopes in subtidal food webs: Have enriched carbon ratios in benthic consumers been misinterpreted? Limnology and Oceanography, 52 (6):2828–2836, November 2006. CODEN LIOCAH. ISSN 0024-

Nurminen:2006:EFF

[NH06b] Leena Nurminen and Jukka Horppila. Efficiency of fish feeding on plant-attached prey: Effects of inorganic turbidity and plant-mediated changes in the light environment. Limnology and Oceanography, 51(3):1550–1555, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Nomaki:2005:SSI

[NHN^{+}05] Hidetaka Nomaki, Petra Heinz, Takeshi Nakatsuka, Motohiro Shimanaga, and Hiroshi Kitazato. Species-specific ingestion of organic carbon by deep-sea benthic Foraminifera and Meiobenthos: In situ tracer experiments. Limnology and Oceanography, 50(1):134–146, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Niemisto:2008:IBS

[NHPHH08] Juha Niemistö, Heidi Holmroos, Zeynep Pekcan-Hekim, and Jukka Horppila. Interactions between sediment resuspension and sediment quality decrease the TN : TP ratio in a shallow lake. Limnology and Oceanography, 53(6):2407–2415, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Noe:2007:CSP

[NHS07] Gregory B. Noe, Judson W. Harvey, and James E. Saiers. Characterization of suspended particles in Everglades wetlands. Limnology and Oceanography, 52(3):1166–1178, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Norkko:2001:BPC

[NHTF01] Alf Norkko, Judi E. Hewitt, Simon F. Thrush, and Thrush Funnell. Benthic-pelagic coupling and suspension-feeding bivalves: Linking site-specific sediment flux and biodeposition to benthic community structure. Limnology and Oceanography, 47(4):2067– 2072, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Nishri:2000:PRR

[NIE^{+}00] A. Nishri, J. Imberger, W. Eckert, I. Ostrovsky, and Y. Geifman. The physical regime and the respective biogeochemical processes in the lower water mass of Lake Kinneret. Limnology and Oceanography, 45(4):972–981, June 2000. CODEN LI-OCAH. ISSN 0024-3590.

Nils:2003:CND

[Nil03] Risgaard-Petersen Nils. Coupled nitrification-denitrification in autotrophic and heterotrophic estuarine sediments: On the influence of benthic microalgae. Limnology and Oceanography, 48 (1):93–105, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Nuwer:2005:SOM

[NK05] Jonathan M. Nuwer and Richard G. Keil. Sedimentary organic matter geochemistry of Clayoquot Sound, Vancouver Island, British Columbia. Limnology and Oceanography, 50(4):1119– 1128, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Nascimento:2008:SBF

[NKE08] Francisco J. A. Nascimento, Agnes M. L. Karlson, and Ragnar Elmgren. Settling blooms of filamentous cyanobacteria as food for meiofauna assemblages. Limnology and Oceanography, 53(6): 2636–2643, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:2006:TTP

[NKJM06] Ole I. Nielsen, Marguerite S. Koch, Henning S. Jensen, and Christopher J. Madden. Thalassia testudinum phosphate uptake kinetics at low in situ concentrations using a ^{33}P radioisotope technique. Limnology and Oceanography, 51(1):208–217, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Nakabayashi:2002:VII

[NKS^{+}02] Shigeto Nakabayashi, Kenshi Kuma, Kohsei Sasaoka, Seiichi Saitoh, Mamiko Mochizuki, Naonobu Shiga, and Masashi Kusakabe. Variation in iron(III) solubility and iron concentration in the northwestern North Pacific Ocean. Limnology and Oceanography, 47(3):885–892, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Nagata:2003:MDP

[NMK03] T. Nagata, B. Meon, and D. L. Kirchman. Microbial degradation of peptidoglycan in seawater. Limnology and Oceanography, 48 (2):745–754, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Not:2005:LSC

[NML^{+}05] Fabrice Not, Ramon Massana, Mikel Latasa, Dominique Marie, Céline Colson, Wenche Eikrem, Carlos Pedrós-Alió, Daniel Vaulot, and Nathalie Simon. Late summer community composition and abundance of photosynthetic picoeukaryotes in Norwegian and Barents Seas. Limnology and Oceanography, 50(5): 1677–1686, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Norberg:2004:BEF

[Nor04] Jon Norberg. Biodiversity and ecosystem functioning: a complex adaptive systems approach. Limnology and Oceanography, 49(4 (part 2)):1269–1277, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Nodwell:2001:DUI

[NP01] Lisa M. Nodwell and Neil M. Price. Direct use of inorganic colloidal iron by marine mixotrophic phytoplankton. Limnology and Oceanography, 46(4):765–777, June 2001. CODEN LI-OCAH. ISSN 0024-3590.

Naddafi:2008:EZM

[NPE08] Rahmat Naddafi, Kurt Pettersson, and Peter Eklöv. Effects of the zebra mussel, an exotic freshwater species, on seston stoichiometry. Limnology and Oceanography, 53(5):1973–1987, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Navarro:2008:ITU

[NRB08] E. Navarro, C. T. Robinson, and R. Behra. Increased tolerance to ultraviolet radiation (UVR) and cotolerance to cadmium in UVR-acclimatized freshwater periphyton. Limnology and Oceanography, 54(3):1149–1158, May 2008. CODEN LI-OCAH. ISSN 0024-3590.

Nie:2001:MDS

[NSBA01] Yonghong Nie, Ismail B. Suayah, Larry K. Benninger, and Marc J. Alperin. Modeling detailed sedimentary ^{210}Pb and fallout ^{239,240}Pu profiles to allow episodic events: an application in Chesapeake Bay. Limnology and Oceanography, 46(6):1425– 1437, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Nieblas:2009:VBP

[NSH^{+}09] Anne-Elise Nieblas, Bernadette M. Sloyan, Alistair J. Hobday, Richard Coleman, and Anthony J. Richardson. Variability of biological production in low wind-forced regional upwelling systems: a case study off southeastern Australia. Limnology and Oceanography, 54(5):1548–1558, September 2009. CODEN LI-OCAH. ISSN 0024-3590.

Nelson:2009:CIS

[NSM09] Craig E. Nelson, Steven Sadro, and John M. Melack. Contrasting the influences of stream inputs and landscape position on bacterioplankton community structure and dissolved organic matter composition in high-elevation lake chains. Limnology and Oceanography, 54(4):1292–1305, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Neveux:2008:RAL

[NT08] Jacques Neveux and Marcio M. B. Tenório. Response to “Another look at green Trichodesmium colonies”. Limnology and Oceanography, 53(5):2052–2055, September 2008. CODEN LI-OCAH. ISSN 0024-3590. See [OGR08].

Neveux:2006:SDP

[NTDV06] Jacques Neveux, Marcio M. B. Tenírio, Cécile Dupouy, and Tracy A. Villareal. Spectral diversity of phycoerythrins and diazotroph abundance in tropical waters. Limnology and Oceanography, 52(2):1689–1698, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Newbold:2005:DBR

[NTM^{+}05] J. Denis Newbold, Steven A. Thomas, G. Wayne Minshall, Colbert E. Cushing, and Ted Georgian. Deposition, benthic residence, and resuspension of fine organic particles in a mountain stream. Limnology and Oceanography, 50(5):1571–1580, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Olsen:2006:CSR

[OAA^{+}06] Y. Olsen, S. Agustí, T. Andersen, C. M. Duarte, J. M. Gasol, I. Gismervik, A.-S. Heiskanen, E. Hoell, P. Kuuppo, R. Lignell, H. Reinertsen, U. Sommer, H. Stibor, T. Tamminen, O. Vadstein, D. Vaqué, and M. Vidal. A comparative study of responses in plankton food web structure and function in contrasting European coastal waters exposed to experimental nutrient addition. Limnology and Oceanography, 51(1 (part 2)):488–503, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Obernosterer:2004:CBB

[OB04] Ingrid Obernosterer and Ronald Benner. Competition between biological and photochemical processes in the mineralization of dissolved organic carbon. Limnology and Oceanography, 49(1): 117–124, January 2004. CODEN LIOCAH. ISSN 0024-3590.

OBrien:2004:PCB

[OBB^{+}04] W. John O’Brien, Michael Barfield, Neil D. Bettez, Gretchen M. Gettel, Anne E. Hershey, Michael E. McDonald, Michael C. Miller, Howard Mooers, John Pastor, Carl Richards, and Jeff Schuldt. Physical, chemical, and biotic effects on Arctic zooplankton communities and diversity. Limnology and Oceanography, 49(4 (part 2)):1250–1261, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Oliver:2004:HBR

[OBSD04] Jacques L. Oliver, Richard T. Barber, Walker O. Smith, Jr., and Hugh W. Ducklow. The heterotrophic bacterial response during the Southern Ocean iron experiment (SOFeX). Limnology and Oceanography, 49(6):2129–2140, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Otero:2000:APO

[OCNH00] Ernesto Otero, Randolph Culp, John E. Noakes, and Robert E. Hodson. Allocation of particulate organic carbon from different sources in two contrasting estuaries of southeastern U.S.A. Limnology and Oceanography, 45(8):1753–1763, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Oubelkheir:2005:BOB

[OCSB05] Kadija Oubelkheir, Hervé Claustre, Antoine Sciandra, and Marcel Babin. Bio-optical and biogeochemical properties of different trophic regimes in oceanic waters. Limnology and Oceanography, 50(6):1795–1809, November 2005. CODEN LIOCAH. ISSN 0024-3590.

ODonnell:2008:HFS

[OD08] Michael J. O’Donnell and Mark W. Denny. Hydrodynamic forces and surface topography: Centimeter-scale spatial variation in wave forces. Limnology and Oceanography, 53(2):579– 588, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Ohman:2008:RPP

[ODR^{+}08] Mark D. Ohman, Edward G. Durbin, Jeffrey A. Runge, Barbara K. Sullivan, and David B. Field. Relationship of predation potential to mortality of Calanus finmarchicus on Georges Bank, northwest Atlantic. Limnology and Oceanography, 53(5): 1643–1655, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Omori:2001:SDF

[OFKH01] Makoto Omori, Hironobu Fukami, Hisao Kobinata, and Masayuki Hatta. Significant drop of fertilization of Acropora corals in 1999: an after-effect of heavy coral bleaching? Limnology and Oceanography, 46(6):704–706, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Orcutt:2008:ALG

[OGR08] Karen M. Orcutt, Kjell Gundersen, and Ulla Rasmussen. Another look at green Trichodesmium colonies. Limnology and Oceanography, 53(5):2049–2051, September 2008. CODEN LI-OCAH. ISSN 0024-3590. See response [NT08].

Obernosterer:2000:DOB

[OH00] Ingrid Obernosterer and Gerhard J. Herndl. Differences in the optical and biological reactivity of the humic and nonhumic dissolved organic carbon component in two contrasting coastal marine environments. Limnology and Oceanography, 45(5):1120– 1129, July 2000. CODEN LIOCAH. ISSN 0024-3590.

OConnor:2008:DOT

[OH08] Ben L. O’Connor and Miki Hondzo. Dissolved oxygen transfer to sediments by sweep and eject motions in aquatic environments. Limnology and Oceanography, 53(2):566–578, March 2008. CO-DEN LIOCAH. ISSN 0024-3590.

OReilly:2002:ISI

[OHCP02] C. M. O’Reilly, R. E. Hecky, A. S. Cohen, and P.-D. Plisnier. Interpreting stable isotopes in food webs: Recognizing the role of time averaging at different trophic levels. Limnology and Oceanography, 47(1):306–309, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Orvoine:2006:CBP

[OHT06] Jord Orvoine, Landis Hare, and André Tessier. Competition between protons and cadmium ions in the planktonic food chain leading to the phantom midge chaoborus. Limnology and Oceanography, 51(2):1013–1020, March 2006. CODEN LI-OCAH. ISSN 0024-3590.

Oxborough:2000:VEP

[OHUB00] K. Oxborough, A. R. M. Hanlon, G. J. C. Underwood, and N. R. Baker. In vivo estimation of the photosystem II photochemical efficiency of individual microphytobenthic cells using high-resolution imaging of chlorophyll-a fluorescence. Limnology and Oceanography, 45(6):1420–1425, September 2000. CODEN LIOCAH. ISSN 0024-3590.

OBrien:2003:SMC

[OIH^{+}03] Katherine R. O’Brien, Gregory N. Ivey, David P. Hamilton, Anya M. Waite, and Petra M. Visser. Simple mixing criteria for the growth of negatively buoyant phytoplankton. Limnology and Oceanography, 48(4):1326–1337, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Ogawa:2001:FNI

[OKO^{+}01] Nanako O. Ogawa, Tadatoshi Koitabashi, Hirotaka Oda, Toshio Nakamura, Naohiko Ohkouchi, and Eitaro Wada. Fluctuations of nitrogen isotope ratio of gobiid fish (Isaza) specimens and sediments in Lake Biwa, Japan, during the 20th century. Limnology and Oceanography, 46(5):1228–1236, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Olson:2008:IGM

[OLCT08] M. Brady Olson, Evelyn J. Lessard, Lessard Cochlan, and Vera L. Trainer. Intrinsic growth and microzooplankton grazing on toxigenic Pseudo-nitzschia spp. diatoms from the coastal northeast Pacific. Limnology and Oceanography, 53(4):1352– 1368, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Osburn:2009:ILP

[OOB09] Christopher L. Osburn, Daniel W. O’Sullivan, and Thomas J. Boyd. Increases in the longwave photobleaching of chromophoric dissolved organic matter in coastal waters. Limnology and Oceanography, 54(1):145–159, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Ortega-Retuerta:2009:BCD

[ORFD^{+}09] Eva Ortega-Retuerta, Thomas K. Frazer, Carlos M. Duarte, Sergio Ruiz-Halpern, Antonio Tovar-Sánchez, Jesús M. Arrieta, and Isabel Reche. Biogeneration of chromophoric dissolved organic matter by bacteria and krill in the Southern Ocean. Limnology and Oceanography, 54(6):1941–1950, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Obernosterer:2001:SDD

[ORH01] Ingrid Obernosterer, Piet Ruardij, and Gerhard J. Herndl. Spatial and diurnal dynamics of dissolved organic matter (DOM) fluorescence and H_{2}O_{2} and the photochemical oxygen demand of surface water DOM across the subtropical Atlantic Ocean. Limnology and Oceanography, 46(6):632–643, May 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Oldham:2001:EEV

[OS01a] C. E. Oldham and J. J. Sturman. The effect of emergent vegetation on convective flushing in shallow wetlands: Scaling and experiments. Limnology and Oceanography, 46(6):1486–1493, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Ottersen:2001:ACG

[OS01b] Geir Ottersen and Nils Chr. Stenseth. Atlantic climate governs oceanographic and ecological variability in the Barents Sea. Limnology and Oceanography, 46(7):1774–1780, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Obayashi:2005:PEC

[OS05] Yumiko Obayashi and Satoru Suzuki. Proteolytic enzymes in coastal surface seawater: Significant activity of endopeptidases and exopeptidases. Limnology and Oceanography, 50(4):722– 726, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Oliver:2007:RIR

[OSDB07] Jacques L. Oliver, Walker O. Smith, Hugh W. Ducklow, and Richard T. Barber. Rejoinder to: “Interpreting the results of oceanic mesoscale enrichment experiments: Caveats and lessons from limnology and coastal ecology”. Limnology and Oceanography, 52(2):919–920, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Ostrovsky:2003:MBL

[Ost03] Ilia Ostrovsky. Methane bubbles in Lake Kinneret: Quantification and temporal and spatial heterogeneity. Limnology and Oceanography, 48(3):1030–1036, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Orellana:2003:URB

[OV03] Mónica V. Orellana and Pedro Verdugo. Ultraviolet radiation blocks the organic carbon exchange between the dissolved phase and the gel phase in the ocean. Limnology and Oceanography, 48 (4):1618–1623, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Osburn:2001:CSW

[OZM^{+}01] Christopher L. Osburn, Horacio E. Zagarese, Donald P. Morris, Bruce R. Hargreaves, and Walter E. Cravero. Calculation of spectral weighting functions for the solar photobleaching of chromophoric dissolved organic matter in temperate lakes. Limnology and Oceanography, 46(6):1455–1467, September 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Poulain:2004:BPP

[PAF^{+}04] A. J. Poulain, M. Amyot, D. Findlay, S. Telor, T. Barkay, and H. Hintelmann. Biological and photochemical production of dissolved gaseous mercury in a boreal lake. Limnology and Oceanography, 49(6):2265–2275, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Parnell:2001:DEO

[Par01] P. Ed Parnell. The distribution of estuarine and oceanic water masses on the southern shore of O’ahu, Hawai’i: Ecological and coastal management implications, and novel methodology. Limnology and Oceanography, 46(6):1468–1485, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Pond:2005:DFA

[PAS^{+}05] David W. Pond, Angus Atkinson, Rachael S. Shreeve, Geraint Tarling, and Peter Ward. Diatom fatty acid biomarkers indicate recent growth rates in Antarctic krill. Limnology and Oceanography, 50(4):732–736, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Prairie:2002:SMB

[PBC02] Yves T. Prairie, David F. Bird, and Jonathan J. Cole. The summer metabolic balance in the epilimnion of southeastern Quebec lakes. Limnology and Oceanography, 47(1):316–321, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Peterson:2009:QPS

[PBGJ09] Richard N. Peterson, William C. Burnett, Craig R. Glenn, and Adam G. Johnson. Quantification of point-source groundwater discharges to the ocean from the shoreline of the Big Island, Hawaii. Limnology and Oceanography, 54(3):890–904, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Pullin:2004:ESH

[PBGV04] Michael J. Pullin, Stefan Bertilsson, Jared V. Goldstone, and Bettina M. Voelker. Effects of sunlight and hydroxyl radical on dissolved organic matter: Bacterial growth efficiency and production of carboxylic acids and other substrates. Limnology and Oceanography, 49(6):2011–2022, November 2004. CODEN LI-OCAH. ISSN 0024-3590.

Pawlowicz:2007:PCM

[PBM^{+}07] Rich Pawlowicz, Susan A. Baldwin, Annette Muttray, Jana Schmidtova, Bernard Laval, and Grant Lamont. Physical, chemical, and microbial regimes in an anoxic fjord (Nitinat Lake). Limnology and Oceanography, 52(3):1002–1017, May 2007. CO-DEN LIOCAH. ISSN 0024-3590.

Park:2004:CFP

[PBMSG04] Sangkyu Park, Michael T. Brett, Anke Müller-Solger, and Charles R. Goldman. Climatic forcing and primary productivity in a subalpine lake: Interannual variability as a natural experiment. Limnology and Oceanography, 49(2):614–619, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Pace:2002:SVD

[PC02] Michael L. Pace and Jonathan J. Cole. Synchronous variation of dissolved organic carbon and color in lakes. Limnology and Oceanography, 47(2):333–342, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Paris:2004:DEB

[PC04] Claire B. Paris and Robert K. Cowen. Direct evidence of a biophysical retention mechanism for coral reef fish larvae. Limnology and Oceanography, 49(6):1964–1979, November 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Phlips:2000:LAV

[PCA^{+}00] E. J. Phlips, M. Cichra, F. J. Aldridge, J. Jembeck, J. Hendrickson, and R. Brody. Light availability and variations in phytoplankton standing crops in a nutrient-rich blackwater river. Limnology and Oceanography, 45(4):916–929, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Pace:2007:DTO

[PCC^{+}07] Michael L. Pace, Stephen R. Carpenter, Jonathan J. Cole, James J. Coloso, James F. Kitchell, James R. Hodgson, Jack J. Middelburg, Nicholas D. Preston, Christopher T. Solomon, and Brian C. Weidel. Does terrestrial organic carbon subsidize the planktonic food web in a clear-water lake? Limnology and Oceanography, 52(5):2177–2189, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Priscu:2008:SPI

[PCD^{+}08] John C. Priscu, Brent C. Christner, John E. Dore, Brian N. Popp, Karen L. Casciotti, and W. Berry Lyons. Supersaturated N_{2}O in a perennially ice-covered Antarctic lake: Molecular and stable isotopic evidence for a biogeochemical relict. Limnology and Oceanography, 53(6):2439–2450, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Pfandl:2009:SRG

[PCDB09] Karin Pfandl, Antonis Chatzinotas, Patricia Dyal, and Jens Boenigk. SSU rRNA gene variation resolves population heterogeneity and ecophysiological differentiation within a morphospecies (Stramenopiles, Chrysophyceae). Limnology and Oceanography, 54(1):171–181, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Philippart:2000:LTP

[PCvRR00] Catharina J. M. Philippart, Gerhard C. Cadée, Wim van Raaphorst, and Roel Riegman. Long-term phytoplanktonnutrient interactions in a shallow coastal sea: Algal community structure, nutrient budgets, and denitrification potential. Limnology and Oceanography, 45(1):131–144, January 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Parsons:2002:SEI

[PD02] Michael L. Parsons and Quay Dortch. Sedimentological evidence of an increase in Pseudo-nitzschia (Bacillariophyceae) abundance in response to coastal eutrophication. Limnology and Oceanography, 47(2):551–558, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Purcell:2005:ECR

[PD05] J. E. Purcell and M. B. Decker. Effects of climate on relative predation by scyphomedusae and ctenophores on copepods in Chesapeake Bay during 1987–2000. Limnology and Oceanography, 55(3):376–387, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Ptacnik:2003:PSN

[PDB03] Robert Ptacnik, Sebastian Diehl, and Stella Berger. Performance of sinking and nonsinking phytoplankton taxa in a gradient of mixing depths. Limnology and Oceanography, 49(4): 1903–1912, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Paterson:2006:QRP

[PDH06] Gordon Paterson, Kenneth G. Drouillard, and G. Douglas Haffner. Quantifying resource partitioning in centrarchids with stable isotope analysis. Limnology and Oceanography, 51(2): 1038–1044, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Pasche:2009:PBL

[PDM^{+}09] Natacha Pasche, Christian Dinkel, Beat Müller, Martin Schmid, Alfred Wüest, and Bernhard Wehrli. Physical and biogeochemical limits to internal nutrient loading of meromictic Lake Kivu. Limnology and Oceanography, 54(6):1863–1873, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Pascal:2009:SVC

[PDR^{+}09] Pierre-Yves Pascal, Christine Dupuy, Pierre Richard, Clarisse Mallet, Eric Armynot du Châ telet, and Nathalie Niquilb. Seasonal variation in consumption of benthic bacteria by meio- and macrofauna in an intertidal mudflat. Limnology and Oceanography, 54(4):1048–1059, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Pineda:2005:BLI

[PDS05] Jesús Pineda, Claudio DiBacco, and Victoria Starczak. Barnacle larvae in ice: Survival, reproduction, and time to postsettlement metamorphosis. Limnology and Oceanography, 50(5):1520–1528, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Park:2007:NAE

[PDS07] Wongyu Park, David C. Douglas, and Thomas C. Shirley. North to Alaska: Evidence for conveyor belt transport of Dungeness crab larvae along the west coast of the United States and Canada. Limnology and Oceanography, 52(1):248–256, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Parsons:2006:RDE

[PDTR06] Michael L. Parsons, Quay Dortch, R. Eugene Turner, and Nancy R. Rabalais. Reconstructing the development of eutrophication in Louisiana salt marshes. Limnology and Oceanography, 51(1 (part 2)):534–544, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Peng:2007:SMP

[PE07] Feng Peng and Steven W. Effler. Suspended minerogenic particles in a reservoir: Light-scattering features from individual particle analysis. Limnology and Oceanography, 52(1):204–216, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Peng:2009:CMP

[PED^{+}09] Feng Peng, Steve W. Effler, David O’ Donnell, Alan D. Weidemann, and Martin T. Auer. Characterizations of minerogenic particles in support of modeling light scattering in Lake Superior through a two-component approach. Limnology and Oceanography, 54(4):1369–1381, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Peterson:2001:LSP

[PF01a] Bradley J. Peterson and James W. Fourqurean. Large-scale patterns in seagrass (Thalassia testudinum) demographics in south Florida. Limnology and Oceanography, 46(5):1077–1090, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Petersson:2001:UAA

[PF01b] Mona Petersson and Sören Floderus. Use of amino acid composition to investigate settling and resuspension of a spring bloom in the southern Skagerrak. Limnology and Oceanography, 46(5): 1111–1120, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Pringle:2001:AMT

[PF01c] James M. Pringle and Peter J. S. Franks. Asymmetric mixing transport: a horizontal transport mechanism for sinking plankton and sediment in tidal flows. Limnology and Oceanography, 46(2):381–391, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Peeters:2003:DWR

[PFH^{+}03] Frank Peeters, David Finger, Markus Hofer, Matthias Brennwald, David M. Livingstone, and Rolf Kipfer. Deep-water renewal in Lake Issyk-Kul driven by differential cooling. Limnology and Oceanography, 48(4):1419–1431, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Pierson:2007:LGC

[PFL07] James J. Pierson, Bruce W. Frost, and Andrew W. Leising. The lost generation of Calanus pacificus: Is the diatom effect responsible? Limnology and Oceanography, 52(5):2089–2098, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Ploger:2005:LCS

[PFN^{+}05] Anja Plöger, Eric Fischer, Hans-Peter Nirmaier, Luis M. Laglera, Damiano Monticelli, and Constant M. G. van den Berg. Lead and copper speciation in remote mountain lakes. Limnology and Oceanography, 50(3):995–1010, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Precht:2004:ODP

[PFPH04] Elimar Precht, Ulrich Franke, Lubos Polerecky, and Markus Huettel. Oxygen dynamics in permeable sediments with wavedriven pore water exchange. Limnology and Oceanography, 49 (3):693–705, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2000:BGG

[PG00] Helle Ploug and Hans-Peter Grossart. Bacterial growth and grazing on diatom aggregates: Respiratory carbon turnover as a function of aggregate size and sinking velocity. Limnology and Oceanography, 45(7):1467–1475, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Padilla-Gamino:2007:SAA

[PGC07] Jacqueline L. Padilla-Gamiño and Robert C. Carpenter. Seasonal acclimatization of Asparagopsis taxiformis (Rhodophyta) from different biogeographic regions. Limnology and Oceanography, 52(2):833–842, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Patoine:2006:SVN

[PGL06] Alain Patoine, Mark D. Graham, and Peter R. Leavitt. Spatial variation of nitrogen fixation in lakes of the northern Great Plains. Limnology and Oceanography, 52(2):1665–1677, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Perroud:2009:SMT

[PGM^{+}09] Marjorie Perroud, Stéphane Goyette, Andrey Martynov, Martin Beniston, and Orlane Annevillec. Simulation of multiannual thermal profiles in deep Lake Geneva: a comparison of onedimensional lake models. Limnology and Oceanography, 54(5): 1574–1594, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Palanques:2001:IBT

[PGP01] A. Palanques, J. Guillén, and P. Puig. Impact of bottom trawling on water turbidity and muddy sediment of an unfished continental shelf. Limnology and Oceanography, 46(5):1100–1110, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Precht:2003:APW

[PH03] Elimar Precht and Markus Huettel. Advective pore-water exchange driven by surface gravity waves and its ecological implications. Limnology and Oceanography, 48(4):1674–1684, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Page:2007:RDS

[PH07] Timothy J. Page and Jane M. Hughes. Radically different scales of phylogeographic structuring within cryptic species of freshwater shrimp (atyidae: Caridina). Limnology and Oceanography, 52(3):1055–1066, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Pel:2003:UHI

[PHF03] Roel Pel, Hans Hoogveld, and Virgilio Floris. Using the hidden isotopic heterogeneity in phyto- and zooplankton to unmask disparity in trophic carbon transfer. Limnology and Oceanography, 48(6):2200–2207, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2002:DAW

[PHK02] Helle Ploug, Susanna Hietanen, and Jorma Kuparinen. Diffusion and advection within and around sinking, porous diatom aggregates. Limnology and Oceanography, 47(4):1129–1136, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Pfeiffer-Herbert:2007:DBL

[PHMR^{+}07] Anna S. Pfeiffer-Herbert, Margaret A. McManus, Peter T. Raimondi, Yi Chao, and Fei Chai. Dispersal of barnacle larvae along the central California coast: a modeling study. Limnology and Oceanography, 52(4):1559–1569, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2008:BSV

[PIF08] Helle Ploug, Morten Hvitfeldt Iversen, and Gerhard Fischer. Ballast, sinking velocity, and apparent diffusivity within marine snow and zooplankton fecal pellets: Implications for substrate turnover by attached bacteria. Limnology and Oceanography, 53 (5):1878–1886, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2008:POR

[PIKB08] Helle Ploug, Morten Hvitfeldt Iversen, Marja Koski, and Erik Theodoor Buitenhuis. Production, oxygen respiration rates, and sinking velocity of copepod fecal pellets: Direct measurements of ballasting by opal and calcite. Limnology and Oceanography, 53(2):469–476, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Pilskaln:2004:SIP

[Pil04] Cynthia H. Pilskaln. Seasonal and interannual particle export in an African Rift Valley lake: a 5-yr record from Lake Malawi, southern East Africa. Limnology and Oceanography, 49(4):964– 977, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Papadimitriou:2005:SSV

[PKKB05] S. Papadimitriou, H. Kennedy, D. P. Kennedy, and J. Borum. Seasonal and spatial variation in the organic carbon and nitrogen concentration and their stable isotopic composition in Zostera marina (Denmark). Limnology and Oceanography, 50(4):1084– 1095, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Pieters:2009:ESE

[PL09] Roger Pieters and Gregory A. Lawrence. Effect of salt exclusion from lake ice on seasonal circulation. Limnology and Oceanography, 54(2):401–412, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Peeters:2002:MYH

[PLG^{+}02] Frank Peeters, David M. Livingstone, Gerrit-Hein Goudsmit, Rolf Kipfer, and Richard Forster. Modeling 50 years of historical temperature profiles in a large central European lake. Limnology and Oceanography, 47(1):186–197, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Park:2005:SHB

[PLK^{+}05] Geun-Ha Park, Kitack Lee, Chul-Min Koo, Hyun-Woo Lee, Chang-Kyu Lee, Jun-Seok Koo, Tongsup Lee, Shin-Hong Ahn, Hak-Gyoon Kim, and Byong-Kwon Park. A sulfur hexafluoridebased Lagrangian study on initiation and accumulation of the red tide Cochlodinium polykrikoides in southern coastal waters of Korea. Limnology and Oceanography, 50(4):578–586, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Pham:2009:STV

[PLM^{+}09] Samantha V. Pham, Peter R. Leavitt, Suzanne McGowan, Björn Wissel, and Leonard I. Wassenaar. Spatial and temporal variability of prairie lake hydrology as revealed using stable isotopes of hydrogen and oxygen. Limnology and Oceanography, 54(1): 101–118, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Pham:2008:SVC

[PLMPN08] Samantha V. Pham, Peter R. Leavitt, Suzanne McGowan, and Pedro Peres-Neto. Spatial variability of climate and land-use effects on lakes of the northern Great Plains. Limnology and Oceanography, 53(2):728–742, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2001:SSO

[Plo01] Helle Ploug. Small-scale oxygen fluxes and remineralization in sinking aggregates. Limnology and Oceanography, 46(7):1624– 1631, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2008:CSB

[Plo08] Helle Ploug. Cyanobacterial surface blooms formed by Aphanizomenon sp. and Nodularia spumigena in the Baltic Sea: Small-scale fluxes, pH, and oxygen microenvironments. Limnology and Oceanography, 53(3):914–921, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Park:2008:LNR

[PLW^{+}08] Geun-Ha Park, Kitack Lee, Rik Wanninkhof, Jia-Zhong Zhang, Dennis A. Hansell, and Richard A. Feely. Large, non-Redfieldian drawdown of nutrients and carbon in the extratropical North Atlantic Ocean (46^{o}N): Evidence for dinitrogen fixation? Limnology and Oceanography, 53(5):1697–1704, September 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Punshon:2008:AHP

[PM08] Stephen Punshon and Robert M. Moore. Aerobic hydrogen production and dinitrogen fixation in the marine cyanobacterium Trichodesmium erythraeum IMS10. Limnology and Oceanography, 53(6):2749–2753, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Prince:2008:EHA

[PMK08] Emily K. Prince, Tracey L. Myers, and Julia Kubanek. Effects of harmful algal blooms on competitors: Allelopathic mechanisms of the red tide dinoflagellate Karenia brevis. Limnology and Oceanography, 53(2):531–541, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Pruski:2000:TBS

[PMPC00] A. M. Pruski, A. Fiala Médioni, R. Prodon, and J. C. Colomines. Thiotaurine is a biomarker of sulfide-based symbiosis in deep-sea bivalves. Limnology and Oceanography, 45(8):1860–1867, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Pitcher:2006:CSB

[PN06] Grant C. Pitcher and Greville Nelson. Characteristics of the surface boundary layer important to the development of red tide on the southern Namaqua shelf of the Benguela upwelling system. Limnology and Oceanography, 51(6):2660–2674, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Palmer:2004:OPC

[PNPA04] Molly R. Palmer, Heidi M. Nepf, Thomas J. R. Pettersson, and Josef D. Ackerman. Observations of particle capture on a cylindrical collector: Implications for particle accumulation and removal in aquatic systems. Limnology and Oceanography, 49(1): 76–85, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:2001:BRK

[Pom01] Lawrence R. Pomeroy. Book review: Kirchman, D. [Ed.]. 2000. Microbial ecology of the oceans. Wiley-Liss. 542 p. US$139. ISBN 0-471-29993-6. Limnology and Oceanography, 49(1):471–472, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Powers:2000:CDD

[PP00] Sean P. Powers and Charles H. Peterson. Conditional density dependence: The flow trigger to expression of density-dependent emigration in bay scallops. Limnology and Oceanography, 45(3): 727–732, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Peers:2004:RMS

[PP04] Graham Peers and Neil M. Price. A role for manganese in superoxide dismutases and growth of iron-deficient diatoms. Limnology and Oceanography, 49(5):1774–1783, September 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Ploug:2007:DMD

[PP07] Helle Ploug and Uta Passow. Direct measurement of diffusivity within diatom aggregates containing transparent exopolymer particles. Limnology and Oceanography, 52(1):1–6, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Perez:2003:MSB

[PPH03] Maria T. Pèrez, Clemens Pausz, and Gerhard J. Herndl. Major shift in bacterioplankton utilization of enantiomeric amino acids between surface waters and the ocean’s interior. Limnology and Oceanography, 48(2):755–763, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Poulin:2002:AOT

[PPL^{+}02] Elie Poulin, Alvaro T. Palma, Germán Leiva, Diego Narvaez, Rodrigo Pacheco, Sergio A. Navarrete, and Juan C. Castilla. Avoiding offshore transport of competent larvae during upwelling events: The case of the gastropod Concholepas concholepas in Central Chile. Limnology and Oceanography, 47(4): 1248–1255, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Peers:2005:CRI

[PQP05] Graham Peers, Sarah-Ann Quesnel, and Neil M. Price. Copper requirements for iron acquisition and growth of coastal and oceanic diatoms. Limnology and Oceanography, 50(4):1149– 1158, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Price:2005:ESC

[Pri05] Neil M. Price. The elemental stoichiometry and composition of an iron-limited diatom. Limnology and Oceanography, 50(4): 1159–1171, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Poorvin:2004:VRI

[PRKHW04] Leo Poorvin, Johanna M. Rinta-Kanto, David A. Hutchins, and Steven W. Wilhelm. Viral release of iron and its bioavailability to marine plankton. Limnology and Oceanography, 49(5):1734– 1741, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Perez:2006:DEA

[PS06] María Teresa Pérez and Ruben Sommaruga. Differential effect of algal- and soil-derived dissolved organic matter on alpine lake bacterial community composition and activity. Limnology and Oceanography, 51(6):2527–2537, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Perry:2008:SOB

[PSEL08] M. J. Perry, B. S. Sackmann, C. C. Eriksen, and C. M. Lee. Seaglider observations of blooms and subsurface chlorophyll maxima off the Washington coast. Limnology and Oceanography, 53(5 (part 2)):2169–2179, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Pawlik:2007:CWC

[PSH^{+}07] Joseph R. Pawlik, Laura Steindler, Timothy P. Henkel, Sven Beer, and Micha Ilan. Chemical warfare on coral reefs: Sponge metabolites differentially affect coral symbiosis in situ. Limnology and Oceanography, 52(2):907–911, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Pincebourde:2008:BTD

[PSH08] Sylvain Pincebourde, Eric Sanford, and Brian Helmuth. Body temperature during low tide alters the feeding performance of a top intertidal predator. Limnology and Oceanography, 53(5): 1562–1573, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Ptacnik:2004:EMM

[PSHM04] Robert Ptacnik, Ulrich Sommer, Thomas Hansen, and Volker Martens. Effects of microzooplankton and mixotrophy in an experimental planktonic food web. Limnology and Oceanography, 49(4 (part 2)):1435–1445, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Ploug:2000:DBL

[PSJ00] Helle Ploug, Willem Stolte, and Bo Barker Jørgensen. Diffusive boundary layers of the colony-forming plankton alga, phaeocystis sp. — implications for nutrient uptake and cellular growth. Limnology and Oceanography, 45(4):1012, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Peeters:2007:TMP

[PSLO07] Frank Peeters, Dietmar Straile, Andreas Lorke, and Dieter Ollinger. Turbulent mixing and phytoplankton spring bloom development in a deep lake. Limnology and Oceanography, 52 (1):286–298, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Porter:2000:GDU

[PSS00] Elka T. Porter, Lawrence P. Sanford, and Steven E. Suttles. Gypsum dissolution is not a universal integrator of ‘water motion’. Limnology and Oceanography, 45(1):145–158, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Paytan:2006:SGD

[PSS^{+}06] Adina Paytan, Gregory G. Shellenbarger, Joseph H. Street, Meagan E. Gonneea, Kristen Davis, Megan B. Young, and Willard S. Moore. Submarine groundwater discharge: an important source of new inorganic nitrogen to coral reef ecosystems. Limnology and Oceanography, 51(1):343–348, January 2006. CODEN LI-OCAH. ISSN 0024-3590.

Papadimitriou:2007:BCN

[PTK^{+}07] S. Papadimitriou, D. N. Thomas, H. Kennedy, C. Haas, H. Kuosa, A. Krell, and G. S. Dieckmann. Biogeochemical composition of natural sea ice brines from the Weddell Sea during early austral summer. Limnology and Oceanography, 52(5):1809–1823, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Peine:2000:EFI

[PTKP00] Ariane Peine, Andrea Tritschler, Kirsten Küsel, and Stefan Peiffer. Electron flow in an iron-rich acidic sediment — evidence for an acidity-driven iron cycle. Limnology and Oceanography, 45 (5):1077–1087, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Pedersen:2001:BIA

[PTLJ01] Anna-Grethe U. Pedersen, Trine R. Thomsen, Bente Aa. Lomstein, and Niels O. G. Jørgensen. Bacterial influence on amino acid enantiomerization in a coastal marine sediment. Limnology and Oceanography, 46(6):1358–1369, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Philippart:2003:CRC

[PvAB^{+}03] Catharina J. M. Philippart, Hendrik M. van Aken, Jan J. Beukema, Oscar G. Bos, Gerhard C. Cadée, and Rob Dekker.

Climate-related changes in recruitment of the bivalve Macoma balthica. Limnology and Oceanography, 48(6):2171–2185, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Peterson:2001:SHT

[PVD01] Christopher G. Peterson, H. Maurice Valett, and Clifford N. Dahm. Shifts in habitat templates for lotic microalgae linked to interannual variation in snowmelt intensity. Limnology and Oceanography, 46(4):858–870, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Pollux:2007:HSD

[PVD^{+}07] B. J. A. Pollux, W. C. E. P. Verberk, M. Dorenbosch, E. Cocheret de la Morinière, I. Nagelkerken, and G. van der Velde. Habitat selection during settlement of three Caribbean coral reef fishes: Indications for directed settlement to seagrass beds and mangroves. Limnology and Oceanography, 52(2):903– 907, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Pohnert:2000:NER

[PvE00] Georg Pohnert and Eric von Elert. No ecological relevance of trimethylamine in fish — Daphnia interactions. Limnology and Oceanography, 45(5):1153–1156, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Paerl:2006:ACI

[PVP^{+}06] Hans W. Paerl, Lexia M. Valdes, Benjamin L. Peierls, Jason E. Adolf, and Lawrence, Jr. W. Harding. Anthropogenic and climatic influences on the eutrophication of large estuarine ecosystems. Limnology and Oceanography, 51(1 (part 2)):448–462, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Plante:2001:BRD

[PW01] Craig J. Plante and Susan B. Wilde. Bacterial recolonization of deposit-feeder egesta: In situ regrowth or immigration? Limnology and Oceanography, 46(5):1171–1181, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Pilati:2003:IZP

[PW03] Alberto Pilati and Wayne A. Wurtsbaugh. Importance of zooplankton for the persistence of a deep chlorophyll layer: a limnocorral experiment. Limnology and Oceanography, 48(1):249–260, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Pellerin:2004:RWD

[PWH^{+}04] Brian A. Pellerin, Wilfred M. Wollheim, Charles S. Hopkinson, William H. McDowell, Michael R. Williams, Charles J. Vörösmarty, and Michelle L. Daley. Role of wetlands and developed land use on dissolved organic nitrogen concentrations and DON/TDN in northeastern U.S. rivers and streams. Limnology and Oceanography, 49(4):910–918, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Pfister:2007:RRC

[PWN07] Catherine A. Pfister, J. Timothy Wootton, and Christopher J. Neufeld. The relative roles of coastal and oceanic processes in determining physical and chemical characteristics of an intensively sampled nearshore system. Limnology and Oceanography, 52(5):1767–1775, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Parrish:2007:DAE

[PWP07] Christopher C. Parrish, Michael Whiticar, and Velmurugu Puvanendran. Is ω6 docosapentaenoic acid an essential fatty acid during early ontogeny in marine fauna? Limnology and Oceanography, 52(1):476–479, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Peel:2009:ZIC

[PWS09] Kate Peel, Dominik Weiss, and Laura Sigg. Zinc isotope composition of settling particles as a proxy for biogeochemical processes in lakes: Insights from the eutrophic Lake Greifen, Switzerland. Limnology and Oceanography, 54(5):1699–1708, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Pan:2004:DUD

[PWX04] Jin-Fen Pan and Wang Wen-Xiong. Differential uptake of dissolved and particulate organic carbon by the marine mussel Perna viridis. Limnology and Oceanography, 49(6):1980–1991, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Quinlan:2002:LEC

[QLD^{+}02] Roberto Quinlan, Peter R. Leavitt, Aruna S. Dixit, Roland I. Hall, and John P. Smol. Landscape effects of climate, agriculture, and urbanization on benthic invertebrate communities of Canadian prairie lakes. Limnology and Oceanography, 47(2): 378–391, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Qin:2007:ESB

[QMS^{+}07] Peibing Qin, Christine M. Mayer, Kimberly L. Schulz, Xinli Ji, and Mark E. Ritchie. Ecological stoichiometry in benthic food webs: Effect of light and nutrients on periphyton food quantity and quality in lakes. Limnology and Oceanography, 52(4):1728– 1734, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Quigg:2006:CUK

[QRF06] Antonietta Quigg, John R. Reinfelder, and Nicholas S. Fisher. Copper uptake kinetics in diverse marine phytoplankton. Limnology and Oceanography, 51(2):893–899, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Quigley:2002:IAP

[QSH^{+}02] Matthew S. Quigley, Peter H. Santschi, Chin-Chang Hung, Laodong Guo, and Bruce D. Honeyman. Importance of acid polysaccharides for ^{234}Th complexation to marine organic matter. Limnology and Oceanography, 47(2):367–377, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Quan-Young:2006:RZD

[QYEA06] L. I. Quan-Young and J. Espinoza-Avalos. Reduction of zooxanthellae density, chlorophyll-a concentration, and tissue thickness of the coral Montastraea faveolata (Scleractinia) when competing with mixed turf algae. Limnology and Oceanography, 51(2): 1159–1166, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Rynearson:2000:DFR

[RA00] Tatiana A. Rynearson and E. Virginia Armbrust. DNA fingerprinting reveals extensive genetic diversity in a field population of the centric diatom Ditylum brightwellii. Limnology and Oceanography, 45(6):1329–1340, September 2000. CODEN LI-OCAH. ISSN 0024-3590.

Repeta:2006:RAN

[RA06] Daniel J. Repeta and Lihini I. Aluwihare. Radiocarbon analysis of neutral sugars in high-molecular-weight dissolved organic carbon: Implications for organic carbon cycling. Limnology and Oceanography, 51(2):1045–1053, March 2006. CODEN LI-OCAH. ISSN 0024-3590.

Reese:2009:DSP

[RA09] Brandi Kiel Reese and Michael A. Anderson. Dimethyl sulfide production in a saline eutrophic lake, Salton Sea, California. Limnology and Oceanography, 54(1):250–261, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Rimmer:2005:CST

[RAKE05] Alon Rimmer, Yasuaki Aota, Michio Kumagai, and Werner Eckert. Chemical stratification in thermally stratified lakes: a chloride mass balance model. Limnology and Oceanography, 50(1): 147–157, January 2005. CODEN LIOCAH. ISSN 0024-3590. See erratum [CR10].

Raymond:2001:DCT

[RB01a] Peter A. Raymond and James E. Bauer. DOC cycling in a temperate estuary: a mass balance approach using natural ^{14}C and ^{13}C isotopes. Limnology and Oceanography, 46(6):655–667, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Rozan:2001:MBH

[RB01b] Timothy F. Rozan and Gaboury Benoit. Mass balance of heavy metals in New Haven Harbor, Connecticut: Predominance of nonpoint sources. Limnology and Oceanography, 47(4):2032– 2049, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Roelke:2002:DPS

[RB02] Daniel Roelke and Yesim Buyukates. Dynamics of phytoplankton succession coupled to species diversity as a system-level tool for study of Microcystis population dynamics in eutrophic lakes. Limnology and Oceanography, 47(4):1109–1118, July 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Riis:2003:HHC

[RB03] Tenna Riis and Barry J. F. Biggs. Hydrologic and hydraulic control of macrophyte establishment and performance in streams. Limnology and Oceanography, 48(4):1488–1497, July 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Ravet:2006:PEF

[RB06] Joseph L. Ravet and Michael T. Brett. Phytoplankton essential fatty acid and phosphorus content constraints on Daphnia somatic growth and reproduction. Limnology and Oceanography, 51(5):2438–2452, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Ramey:2008:SDF

[RB08] Patricia A. Ramey and Erica Bodnar. Selection by a depositfeeding polychaete, Polygordius jouinae, for sands with relatively high organic content. Limnology and Oceanography, 53 (4):1512–1520, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Rapaglia:2009:EGA

[RB09] John P. Rapaglia and Henry J. Bokuniewicz. The effect of groundwater advection on salinity in pore waters of permeable sediments. Limnology and Oceanography, 54(2):630–643, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Raymond:2000:ACE

[RBC00] Peter A. Raymond, James E. Bauer, and Jonathan J. Cole. Atmospheric CO_{2} evasion, dissolved inorganic carbon production, and net heterotrophy in the York River estuary. Limnology and Oceanography, 45(8):1707–1717, December 2000. CODEN LI-OCAH. ISSN 0024-3590.

Ryves:2006:PCP

[RBJ^{+}06] David B. Ryves, Richard W. Battarbee, Stephen Juggins, Sherilyn C. Fritz, and N. John Anderson. Physical and chemical predictors of diatom dissolution in freshwater and saline lake sediments in North America and West Greenland. Limnology and Oceanography, 51(3):1355–1368, May 2006. CODEN LI-OCAH. ISSN 0024-3590.

Ravet:2003:TRP

[RBMN03] Joseph L. Ravet, Michael T. Brett, and Dörthe C. Müller-Navarra. A test of the role of polyunsaturated fatty acids in phytoplankton food quality for Daphnia using liposome supplementation. Limnology and Oceanography, 49(4):1938–1947, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Rothenberger:2009:ULT

[RBW09] Meghan B. Rothenberger, JoAnn M. Burkholder, and Thomas R. Wentworth. Use of long-term data and multivariate ordination techniques to identify environmental factors governing estuarine phytoplankton species dynamics. Limnology and Oceanography, 54(6):2107–2127, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Rueda:2005:EBF

[RC05a] Francisco J. Rueda and Edwin A. Cowen. Exchange between a freshwater embayment and a large lake through a long, shallow channel. Limnology and Oceanography, 55(3):169–183, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Rueda:2005:RTF

[RC05b] Francisco J. Rueda and Edwin A. Cowen. Residence time of a freshwater embayment connected to a large lake. Limnology and Oceanography, 50(5):1638–1653, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Rose:2007:DLT

[RC07] Julie M. Rose and David A. Caron. Does low temperature constrain the growth rates of heterotrophic protists? Evidence and implications for algal blooms in cold waters. Limnology and Oceanography, 52(2):886–895, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Ribes:2005:SAC

[RCAK05] Marta Ribes, Rafel Coma, Marlin J. Atkinson, and Robert A. Kinzie III. Sponges and ascidians control removal of particulate organic nitrogen from coral reef water. Limnology and Oceanography, 50(5):1480–1489, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Rohrlack:2005:IMD

[RCD^{+}05] Thomas Rohrlack, Kirsten Christoffersen, Elke Dittmann, Isabel Nogueira, Vitor Vasconcelos, and Thomas Börner. Ingestion of microcystins by Daphnia: Intestinal uptake and toxic effects. Limnology and Oceanography, 50(2):440–448, March 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Roesler:2002:DPE

[RCE^{+}02] Collin S. Roesler, Charles W. Culbertson, Stacey M. Etheridge, Ralf Goericke, Ronald P. Kiene, Laurence G. Miller, and Ronald S. Oremland. Distribution, production, and ecophysiology of Picocystis strain ML in Mono Lake, California. Limnology and Oceanography, 47(2):440–452, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Ragueneau:2002:DEB

[RCL^{+}02] Olivier Ragueneau, Laurent Chauvaud, Aude Leynaert, Gérard Thouzeau, Yves-Marie Paulet, Sophie Bonnet, Anne Lorrain, Jacques Grall, Rudolph Corvaisier, Maryvonne Le Hir, Frédéric Jean, and Jacques Clavier. Direct evidence of a biologically active coastal silicate pump: Ecological implications. Limnology and Oceanography, 47(6):1849–1854, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Rosen:2009:ECO

[RCVK09] Peter Rosén, Laura Cunningham, Jorien Vonk, and Jan Karlssona. Effects of climate on organic carbon and the ratio of planktonic to benthic primary producers in a subarctic lake during the past 45 years. Limnology and Oceanography, 54 (5):1723–1732, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Ruddick:2006:SMN

[RDPM06] Kevin G. Ruddick, Vera De Cauwer, Young-Je Park, and Gerald Moore. Seaborne measurements of near infrared water-leaving reflectance: The similarity spectrum for turbid waters. Limnology and Oceanography, 51(2):1167–1179, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Reyns:2006:SDE

[REL06] Nathalie B. Reyns, David B. Eggleston, and Richard, Jr. A. Luettich. Secondary dispersal of early juvenile blue crabs within a wind-driven estuary. Limnology and Oceanography, 51(5): 1982–1995, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Rimmer:2006:EHD

[RENA06] Alon Rimmer, Werner Eckert, Aminadav Nishri, and Yehuda Agnon. Evaluating hypolimnetic diffusion parameters in thermally stratified lakes. Limnology and Oceanography, 53(2):1906– 1914, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Rohrlack:2008:OCT

[RES^{+}08] Thomas Rohrlack, Bente Edvardsen, Randi Skulberg, Camilla B. Halstvedt, Hans C. Utkilen, Robert Ptacnik, and Olav M. Skulberg. Oligopeptide chemotypes of the toxic freshwater cyanobacterium Planktothrix can form sub-populations with dissimilar ecological traits. Limnology and Oceanography, 53(4):1279–1293, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Reusch:2002:MRH

[Reu02] Thorsten B. H. Reusch. Microsatellites reveal high population connectivity in eelgrass (Zostera marina) in two contrasting coastal areas. Limnology and Oceanography, 47(1):78–85, January 2002. CODEN LIOCAH. ISSN 0024-3590.

RodriguezyBaena:2007:POC

[RFM07] Alessia M. Rodriguez y Baena, Scott W. Fowler, and Juan Carlos Miquel. Particulate organic carbon : natural radionuclide ratios in zooplankton and their freshly produced fecal pellets from the NW Mediterranean (MedFlux 2005). Limnology and Oceanography, 52(3):966–974, May 2007. CODEN LIOCAH. ISSN 0024-3590.

RodriguezyBaena:2008:CKS

[RFW08] Alessia M. Rodriguez y Baena, Scott W. Fowler, and Michel Warnau. Could krill schools significantly bias 234Th-based carbon flux models? Limnology and Oceanography, 54(3):1186– 1191, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Ridame:2002:SIP

[RG02] Céline Ridame and Cécile Guieu. Saharan input of phosphate to the oligotrophic water of the open western Mediterranean Sea. Limnology and Oceanography, 47(3):856–869, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Rysgaard:2004:APA

[RG04] Søren Rysgaard and Ronnie Nøhr Glud. Anaerobic N_{2} production in Arctic sea ice. Limnology and Oceanography, 49(1):86–94, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Rakow:2006:OSM

[RG06] Kelly C. Rakow and William M. Graham. Orientation and swimming mechanics by the scyphomedusa aurelia sp. in shear flow. Limnology and Oceanography, 51(2):1097–1106, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Rodrigues:2007:ERM

[RG07] Lisa J. Rodrigues and Andréa G. Grottoli. Energy reserves and metabolism as indicators of coral recovery from bleaching. Limnology and Oceanography, 52(5):1874–1882, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Richardson:2008:JIR

[RG08] Anthony J. Richardson and Mark J. Gibbons. Are jellyfish increasing in response to ocean acidification? Limnology and Oceanography, 53(5):2040–2045, September 2008. CODEN LI-OCAH. ISSN 0024-3590.

Romani:2004:IST

[RGAS04] Anna M. Romaní, Adonis Giorgi, Vicenç Acuña, and Sergi Sabater. The influence of substratum type and nutrient supply on biofilm organic matter utilization in streams. Limnology and Oceanography, 49(5):1713–1721, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Rysgaard:2004:DAA

[RGRPD04] Søren Rysgaard, Ronnie Nøhr Glud, Nils Risgaard-Petersen, and Tage Dalsgaard. Denitrification and anammox activity in Arctic marine sediments. Limnology and Oceanography, 49(5):1493– 1502, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Raikow:2001:BDM

[RH01] David F. Raikow and Stephen K. Hamilton. Bivalve diets in a midwestern U.S. stream: a stable isotope enrichment study. Limnology and Oceanography, 46(3):514–522, May 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Romare:2003:BCT

[RH03] Pia Romare and Lars-Anders Hansson. A behavioral cascade: Top-predator induced behavioral shifts in planktivorous fish and zooplankton. Limnology and Oceanography, 49(4):1956–1964, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Roberts:2006:NLA

[RH06] Brian J. Roberts and Robert W. Howarth. Nutrient and light availability regulate the relative contribution of autotrophs and heterotrophs to respiration in freshwater pelagic ecosystems. Limnology and Oceanography, 51(1):288–298, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Rao:2008:PPH

[RHCS08] Yerubandi R. Rao, Nathan Hawley, Murray N. Charlton, and William M. Schertzer. Physical processes and hypoxia in the central basin of Lake Erie. Limnology and Oceanography, 53(5):

2007–2020, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Roy:2002:RSS

[RHJ02] Hans Røy, Markus Hüttel, and Bo Barker Jørgensen. The role of small-scale sediment topography for oxygen flux across the diffusive boundary layer. Limnology and Oceanography, 47(3): 837–847, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Roy:2004:TOC

[RHJ04] Hans Røy, Markus Huettel, and Bo Barker Jørgensen. Transmission of oxygen concentration fluctuations through the diffusive boundary layer overlying aquatic sediments. Limnology and Oceanography, 49(3):686–692, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Roy:2005:ITF

[RHJ05] Hans Røy, Markus Huettel, and Bo Barker Jørgensen. The influence of topography on the functional exchange surface of marine soft sediments, assessed from sediment topography measured in situ. Limnology and Oceanography, 50(1):106–112, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Rinke:2009:LWD

[RHK^{+}09] Karsten Rinke, Andrea M. R. Huber, Sebastian Kempke, Magdalena Eder, Thomas Wolf, Wolfgang N. Probst, and Karl-Otto Rothhaupt. Lake-wide distributions of temperature, phytoplankton, zooplankton, and fish in the pelagic zone of a large lake. Limnology and Oceanography, 54(4):1306–1322, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Rossi:2004:IIV

[RHM04] F. Rossi, P. M. J. Herman, and J. J. Middelburg. Interspecific and intraspecific variation of δ ^{13}C and δ ^{15}N in depositand suspension-feeding bivalves (Macoma balthica and Cerastoderma edule): Evidence of ontogenetic changes in feeding mode of Macoma balthica. Limnology and Oceanography, 49(2):408– 414, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Roegner:2002:WIP

[RHN^{+}02] G. Curtis Roegner, Barbara M. Hickey, Jan A. Newton, Alan L. Shanks, and David A. Armstrong. Wind-induced plume and bloom intrusions into Willapa Bay, Washington. Limnology and Oceanography, 47(4):1033–1042, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Rydin:2000:API

[RHW00] Emil Rydin, Brian Huser, and Eugene B. Welch. Amount of phosphorus inactivated by alum treatments in Washington lakes. Limnology and Oceanography, 45(1):226–230, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Risk:2002:BRP

[Ris02] Michael Risk. Book review: Porter, J. W., and K. G. Porter. 2002. The Everglades, Florida Bay, and coral reefs of the Florida Keys: an ecosystem sourcebook. CRC Press. 1,000 p. US$200. ISBN 0-8493-2026-7. Limnology and Oceanography, 47(6):1855, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Ritter:2009:RVO

[Rit09] Amy F. Ritter. Regional variation in oceanographic conditions influences intertidal fish assemblage structure. Limnology and Oceanography, 54(5):1559–1573, September 2009. CODEN LI-OCAH. ISSN 0024-3590.

Ryves:2003:QQR

[RJS^{+}03] David B. Ryves, David H. Jewson, Michael Sturm, Richard W. Battarbee, Roger J. Flower, Anson W. Mackay, and Nikolai G. Granin. Quantitative and qualitative relationships between planktonic diatom communities and diatom assemblages in sedimenting material and surface sediments in Lake Baikal, Siberia. Limnology and Oceanography, 48(4):1643–1661, July 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Rautio:2002:UIP

[RK02] Milla Rautio and Atte Korhola. UV-induced pigmentation in subarctic Daphnia. Limnology and Oceanography, 47(1):295– 299, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Radford-Knoery:2001:DBD

[RKGC^{+}01] Joël Radford-Knoery, C. R. German, J.-L. Charlou, J.-P. Donval, and Y. Fouquet. Distribution and behavior of dissolved hydrogen sulfide in hydrothermal plumes. Limnology and Oceanography, 49(1):461–464, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Rooney:2003:RSM

[RKH03] Neil Rooney, Jacob Kalff, and Catherine Habel. The role of submerged macrophyte beds in phosphorus and sediment accumulation in Lake Memphremagog, Quebec, Canada. Limnology and Oceanography, 49(4):1927–1937, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Reidenbach:2009:HFL

[RKK09] Matthew A. Reidenbach, Jeffrey R. Koseff, and M. A. R. Koehl. Hydrodynamic forces on larvae affect their settlement on coral reefs in turbulent, wave-driven flow. Limnology and Oceanography, 54(1):318–330, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Reidenbach:2006:EWM

[RKM^{+}06] Matthew A. Reidenbach, Jeffrey R. Koseff, Stephen G. Monismith, Jonah V. Steinbuck, and Amatzia Genin. The effects of waves and morphology on mass transfer within branched reef corals. Limnology and Oceanography, 51(2):1134–1141, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Richardson:2009:TDS

[RKNA09] David C. Richardson, Louis A. Kaplan, J. Denis Newbold, and Anthony K. Aufdenkampe. Temporal dynamics of seston: a recurring nighttime peak and seasonal shifts in composition in a stream ecosystem. Limnology and Oceanography, 54(1):344–354, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Ravens:2000:SST

[RKWG00] Thomas M. Ravens, Otti Kocsis, Alfred Wüest, and N. Granin. Small-scale turbulence and vertical mixing in Lake Baikal. Limnology and Oceanography, 45(1):159–173, January 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Riisgaard:2000:CET

[RL00] Hans Ulrik Riisgård and Poul. S. Larsen. A comment on experimental techniques for studying particle capture in filter-feeding bivalves. Limnology and Oceanography, 45(5):1192–1195, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Rengefors:2001:TPA

[RL01a] Karin Rengefors and Catherine Legrand. Toxicity in Peridinium aciculiferum — an adaptive strategy to outcompete other winter phytoplankton? Limnology and Oceanography, 46(8):1990–1997, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Riisgaard:2001:MCF

[RL01b] Hans Ulrik Riisgård and Poul S. Larsen. Minireview: Ciliary filter feeding and bio-fluid mechanics — present understanding and unsolved problems. Limnology and Oceanography, 46(4): 882–891, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Rybicki:2007:LTC

[RL07] Nancy B. Rybicki and Jurate M. Landwehr. Long-term changes in abundance and diversity of macrophyte and waterfowl populations in an estuary with exotic macrophytes and improving water quality. Limnology and Oceanography, 52(3):1195–1207, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Rodrigues:2001:HBU

[RlBW01] Rubina M. N. V. Rodrigues and Peter J. le B. Williams. Heterotrophic bacterial utilization of nitrogenous and nonnitrogenous substrates, determined from ammonia and oxygen fluxes. Limnology and Oceanography, 46(7):1675–1683, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Rodrigues:2002:INA

[RlBW02] Rubina M. N. V. Rodrigues and Peter J. le B. Williams. Inorganic nitrogen assimilation by picoplankton and whole plankton in a coastal ecosystem. Limnology and Oceanography, 47(6): 1608–1616, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Roulet:2001:STG

[RLC^{+}01] M. Roulet, M. Lucotte, R. Canuel, N. Farella, Y. G. De Freitos Goch, J. R. Pacheco Peleja, J.-R. D. Guimarães, D. Mergler, and M. Amorim. Spatio-temporal geochemistry of mercury in waters of the Tapajós and Amazon rivers, Brazil. Limnology and Oceanography, 46(5):1141–1157, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Roy:2008:TDD

[RLJdB08] Hans Røy, Jae Seong Lee, Stefan Jansen, and Dirk de Beer. Tide-driven deep pore-water flow in intertidal sand flats. Limnology and Oceanography, 53(4):1521–1530, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Rusak:2004:MSR

[RLM^{+}04] James A. Rusak, Peter R. Leavitt, Suzanne McGowan, Gemai Chen, Olaf Olson, Sybille Wunsam, and Brian F. Cumming. Millennial-scale relationships of diatom species richness and production in two prairie lakes. Limnology and Oceanography, 49 (4 (part 2)):1290–1299, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Raitsos:2008:IFP

[RLM^{+}08] Dionysios E. Raitsos, Samantha J. Lavender, Christos D. Maravelias, John Haralabous, Anthony J. Richardson, and Philip C. Reid. Identifying four phytoplankton functional types from space: an ecological approach. Limnology and Oceanography, 53 (2):605–613, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Raitsos:2006:CBS

[RLP^{+}06] Dionysios E. Raitsos, Samantha J. Lavender, Yaswant Pradhan, Toby Tyrrell, Philip C. Reid, and Martin Edwards. Coccolithophore bloom size variation in response to the regional environment of the subarctic North Atlantic. Limnology and Oceanography, 51(5):2122–2130, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Roberts:2004:MDB

[RM04] Jason Roberts and Andrew McMinn. Marine diffusive boundary layers at high latitudes. Limnology and Oceanography, 49(4): 934–939, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Robertson:2006:ISA

[RM06] A. L. Robertson and A. M. Milner. The influence of stream age and environmental variables in structuring meiofaunal assemblages in recently deglaciated streams. Limnology and Oceanography, 51(3):1454–1465, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Rippey:2009:ODL

[RM09a] Brian Rippey and Catherine McSorley. Oxygen depletion in lake hypolimnia. Limnology and Oceanography, 54(3):905–916, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Rueda:2009:FPS

[RM09b] Francisco J. Rueda and Sally MacIntyre. Flow paths and spatial heterogeneity of stream inflows in a small multibasin lake.

Limnology and Oceanography, 54(6):2041–2057, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Routh:2004:SGR

[RMG^{+}04] Joyanto Routh, Philip A. Meyers, Örjan Gustafsson, Mark Baskaran, Rolf Hallberg, and Anna Schöldström. Sedimentary geochemical record of human–induced environmental changes in the Lake Brunnsviken watershed, Sweden. Limnology and Oceanography, 49(5):1560–1569, September 2004. CODEN LI-OCAH. ISSN 0024-3590.

Reidenbach:2006:BLT

[RMK^{+}06] Matthew A. Reidenbach, Stephen G. Monismith, Jeffrey R. Koseff, Gitai Yahel, and Amatzia Genin. Boundary layer turbulence and flow structure over a fringing coral reef. Limnology and Oceanography, 51(5):1956–1968, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Ross:2008:MPP

[RMS^{+}08] Oliver N. Ross, C. Mark Moore, Dave J. Suggett, Hugh L. MacIntyre, and Richard J. Geider. A model of photosynthesis and photo-protection based on reaction center damage and repair. Limnology and Oceanography, 53(5):1835–1852, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Rynearson:2006:SBD

[RNA06] T. A. Rynearson, J. A. Newton, and E. V. Armbrust. Spring bloom development, genetic variation, and population succession in the planktonic diatom Ditylum brightwellii. Limnology and Oceanography, 51(3):1249–1261, May 2006. CODEN LI-OCAH. ISSN 0024-3590.

Ram:2003:SSN

[RNC03] A. S. Pradeep Ram, Shanta Nair, and D. Chandramohan. Seasonal shift in net ecosystem production in a tropical estuary. Limnology and Oceanography, 48(4):1601–1607, July 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Riser:2008:MEB

[RNR08] S. C. Riser, J. Nystuen, and A. Rogers. Monsoon effects in the Bay of Bengal inferred from profiling float-based measurements of wind speed and rainfall. Limnology and Oceanography, 53(5 (part 2)):2080–2093, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Rocco:2002:UDC

[ROP^{+}02] Vanina E. Rocco, Oscar Oppezzo, Ramón Pizarro, Ruben Sommaruga, Marcela Ferraro, and Horacio E. Zagarese. Ultraviolet damage and counteracting mechanisms in the freshwater copepod Boeckella poppei from the Antarctic Peninsula. Limnology and Oceanography, 47(3):829–836, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Reche:2009:ESD

[RORR^{+}09] Isabel Reche, Eva Ortega-Retuerta, Otilia Romera, Elvira Pulido Villena, Rafael Morales Baquero, and Emilio O. Casamayor. Effect of Saharan dust inputs on bacterial activity and community composition in Mediterranean lakes and reservoirs. Limnology and Oceanography, 54(3):869–879, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Ramirez:2006:FGT

[RP06] Alonso Ramírez and Catherine M. Pringle. Fast growth and turnover of chironomid assemblages in response to stream phosphorus levels in a tropical lowland landscape. Limnology and Oceanography, 51(1):189–196, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Rosemond:2002:LVP

[RPR^{+}02] Amy D. Rosemond, Catherine M. Pringle, Alonso Ramírez, Michael J. Paul, and Judy L. Meyer. Landscape variation in phosphorus concentration and effects on detritus-based tropical streams. Limnology and Oceanography, 47(1):278–289, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Rachel:2008:TNO

[RPR^{+}08] Barnett-Johnson Rachel, Thomas E. Pearson, Frank C. Ramos, Churchill B. Grimes, and R. Bruce MacFarlane. Tracking natal origins of salmon using isotopes, otoliths, and landscape geology. Limnology and Oceanography, 53(5):1633–1642, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Ross:2000:GLY

[RQB^{+}00] Robin M. Ross, Langdon B. Quetin, Karen S. Baker, Maria Vernet, and Raymond C. Smith. Growth limitation in young Euphausia superba under field conditions. Limnology and Oceanography, 45(1):31–43, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Rost:2003:CAB

[RRBS03] Björn Rost, Ulf Riebesell, Steffen Burkhardt, and Dieter Sültemeyer. Carbon acquisition of bloom-forming marine phytoplankton. Limnology and Oceanography, 48(1):55–67, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Ramos-Rodriguez:2003:NLP

[RRCP03] EloÍsa Ramos-Rodríguez and José M. Conde-Porcuna. Nutrient limitation on a planktonic rotifer: Life history consequences and starvation resistance. Limnology and Oceanography, 48(2):933– 938, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Rippey:2004:MLZ

[RRD04] Brian Rippey, Catherine L. Rose, and Richard W. Douglas. A model for lead, zinc, and copper in lakes. Limnology and Oceanography, 49(6):2256–2264, November 2004. CODEN LI-OCAH. ISSN 0024-3590.

Rengefors:2003:EIT

[RRH^{+}03] K. Rengefors, K. C. Ruttenberg, C. L. Haupert, C. Taylor, B. L. Howes, and D. M. Anderson. Experimental investigation of taxon-specific response of alkaline phosphatase activity in natural freshwater phytoplankton. Limnology and Oceanography, 48 (3):1167–1175, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Rodriguez-Roman:2006:PLU

[RRHPT^{+}06] Aimé Rodríguez-Román, Xavier Hernández-Pech, Patricia E. Thome, Susana Enríquez, and Roberto Iglesias-Prieto. Photosynthesis and light utilization in the Caribbean coral Montastraea faveolata recovering from a bleaching event. Limnology and Oceanography, 52(6):2702–2710, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Rost:2006:CAM

[RRS06] Björn Rost, Ulf Riebesell, and Dieter Sültemeyer. Carbon acquisition of marine phytoplankton: Effect of photoperiod length. Limnology and Oceanography, 51(1):12–20, January 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Rau:2001:UCD

[RRSC01] Greg H. Rau, Stephen Ralston, John R. Southon, and Francisco P. Chavez. Upwelling and the condition and diet of juvenile rockfish: a study using ^{14}C, ^{13}C, and ^{15}N natural abundances. Limnology and Oceanography, 46(6):1565–1570, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Rosenstock:2001:SSD

[RS01] Bernd Rosenstock and Meinhard Simon. Sources and sinks of dissolved free amino acids and protein in a large and deep mesotrophic lake. Limnology and Oceanography, 46(6):644–654, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Rueda:2002:SSL

[RS02] Francisco J. Rueda and S. Geoffrey Schladow. Surface seiches in lakes of complex geometry. Limnology and Oceanography, 47 (3):906–910, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Rueda:2009:MSL

[RS09] Francisco Rueda and Geoffrey Schladow. Mixing and stratification in lakes of varying horizontal length scales: Scaling arguments and energy partitioning. Limnology and Oceanography, 54 (6):2003–2017, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Rao:2004:CTE

[RSC04] Yerubandi R. Rao, Michael G. Skafel, and Murray N. Charlton. Circulation and turbulent exchange characteristics during the thermal bar in Lake Ontario. Limnology and Oceanography, 49 (6):2190–2200, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Reinthaler:2008:DOM

[RSH08] Thomas Reinthaler, Eva Sintes, and Gerhard J. Herndl. Dissolved organic matter and bacterial production and respiration in the sea-surface microlayer of the open Atlantic and the western Mediterranean Sea. Limnology and Oceanography, 53(1): 122–136, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Rohwer:2000:CGS

[RSS^{+}00] Forest Rohwer, Anca Segall, Griegx Steward, Victor Seguritan, Mya Breitbart, Felise Wolven, and Farooqx Farooq Azam. The complete genomic sequence of the marine phage Roseophage SIO1 shares homology with nonmarine phages. Limnology and Oceanography, 45(2):408–418, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Robinson:2003:KIE

[RSS^{+}03] Jonathan J. Robinson, Kathleen M. Scott, S. Todd Swanson, Marion H. O’Leary, Kempton Horken, F. Robert Tabita, and Colleen M. Cavanaugh. Kinetic isotope effect and characterization of form II RubisCO from the chemoautotrophic endosymbionts of the hydrothermal vent tubeworm Riftia pachyptila. Limnology and Oceanography, 48(1):48–54, January 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Raikow:2004:DNC

[RSWH04] David F. Raikow, Orlando Sarnelle, Alan E. Wilson, and Stephen K. Hamilton. Dominance of the noxious cyanobacterium Microcystis aeruginosa in low-nutrient lakes is associated with exotic zebra mussels. Limnology and Oceanography, 49(2): 482–487, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Raateoja:2004:FRR

[RSY04] Mika Raateoja, Jukka Seppälä, and Pasi Ylöstalo. Fast repetition rate fluorometry is not applicable to studies of filamentous cyanobacteria from the Baltic Sea. Limnology and Oceanography, 49(4):1006–1012, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Rozan:2002:ISP

[RTT^{+}02] Tim F. Rozan, Martial Taillefert, Robert E. Trouwborst, Brian T. Glazer, Shufen Ma, Julian Herszage, Lexia M. Valdes, Kent S. Price, and George W. Luther III. Iron-sulfur-phosphorus cycling in the sediments of a shallow coastal bay: Implications for sediment nutrient release and benthic macroalgal blooms. Limnology and Oceanography, 47(5):1346–1354, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Romari:2004:CTV

[RV04] Khadidja Romari and Daniel Vaulot. Composition and temporal variability of picoeukaryote communities at a coastal site of the English Channel from ^{18}S rDNA sequences. Limnology and Oceanography, 49(3):784–798, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Reinthaler:2006:PRP

[RvAV^{+}06] Thomas Reinthaler, Hendrik van Aken, Cornelis Veth, Javier Arístegui, Carol Robinson, Peter J. le B. Williams, Philippe Lebaron, and Gerhard J. Herndl. Prokaryotic respiration and production in the meso- and bathypelagic realm of the eastern and western North Atlantic basin. Limnology and Oceanography, 51(3):1262–1273, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Riegman:2002:UDE

[RvBB02] Roel Riegman, Judith Diana Louise van Bleijswijk, and Cornelia Paula Dorothea Brussaard. The use of dissolved esterase activity as a tracer of phytoplankton lysis. Limnology and Oceanography, 47(3):916–920, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Raanan:2009:QSG

[RVP^{+}09] Hadas Raanan, Avner Vengosh, Adina Paytan, Ami Nishri, and Zbigniew Kabala. Quantifying saline groundwater flow into a freshwater lake using the Ra isotope quartet: a case study from the Sea of Galilee (Lake Kinneret), Israel. Limnology and Oceanography, 54(1):119–131, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Roden:2002:KMF

[RW02] Eric E. Roden and Robert G. Wetzel. Kinetics of microbial Fe(III) oxide reduction in freshwater wetland sediments. Limnology and Oceanography, 47(1):198–211, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Rogerson:2000:TDS

[RWFK00] A. Rogerson, J. Watson, Xiao Fang, and E. P. Krantz. Threedimensional spatial coordinates of individual plankton determined using underwater hologrammetry. Limnology and Oceanography, 45(5):1167–1174, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Roy:2008:PII

[RWK08] Eric G. Roy, Mark L. Wells, and D. Whitney King. Persistence of iron(II) in surface waters of the western subarctic Pacific. Limnology and Oceanography, 53(1):89–98, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Rusak:2002:TST

[RYS^{+}02] J. A. Rusak, N. D. Yan, K. M. Somers, K. L. Cottingham, F. Micheli, S. R. Carpenter, T. M. Frost, M. J. Paterson, and D. J. McQueen. Temporal, spatial, and taxonomic patterns of crustacean zooplankton variability in unmanipulated northtemperate lakes. Limnology and Oceanography, 47(3):613–625, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Roman:2005:SAV

[RZMB05] M. Roman, X. Zhang, C. McGilliard, and W. Boicourt. Seasonal and annual variability in the spatial patterns of plankton biomass in Chesapeake Bay. Limnology and Oceanography, 50 (4):480–492, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Rost:2002:LDC

[RZR02] Björn Rost, Ingrid Zondervan, and Ulf Riebesell. Lightdependent carbon isotope fractionation in the coccolithophorid Emiliania huxleyi. Limnology and Oceanography, 47(1):120–128, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Schmid-Araya:2000:IRP

[SA00] J. M. Schmid-Araya. Invertebrate recolonization patterns in the hyporheic zone of a gravel stream. Limnology and Oceanography, 45(4):1000–1005, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Sornes:2004:PEV

[SA04] Tom A. Sørnes and Dag L. Aksnes. Predation efficiency in visual and tactile zooplanktivores. Limnology and Oceanography, 49 (1):69–75, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Simpsona:2009:DEC

[SA09] Gavin L. Simpsona and N. J. Anderson. Deciphering the effect of climate change and separating the influence of confounding factors in sediment core records using additive models. Limnology and Oceanography, 54(6 (part 2)):2529–2541, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Shimeta:2002:SRP

[SABA02] Jeff Shimeta, Carl L. Amos, Stace E. Beaulieu, and Oladipo M. Ashiru. Sequential resuspension of protists by accelerating tidal flow: Implications for community structure in the benthic boundary layer. Limnology and Oceanography, 47(4):1152–1164, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Sterner:2008:SDC

[SAE^{+}08] Robert W. Sterner, Tom Andersen, James J. Elser, Dag O. Hessen, James M. Hood, Edward McCauley, and Jotaro Urabe.

Scale-dependent carbon : nitrogen : phosphorus seston stoichiometry in marine and freshwaters. Limnology and Oceanography, 54(3):1169–1180, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Sarica:2004:SDM

[SAH^{+}04] José Sarica, Marc Amyot, Landis Hare, Marie-Renée Doyon, and Les William Stanfield. Salmon-derived mercury and nutrients in a Lake Ontario spawning stream. Limnology and Oceanography, 49(4):891–899, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Sarma:2005:SVT

[SAH^{+}05] V. V. S. S. Sarma, O. Abe, S. Hashimoto, A. Hinuma, and T. Saino. Seasonal variations in triple oxygen isotopes and gross oxygen production in the Sagami Bay, central Japan. Limnology and Oceanography, 50(4):544–552, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Sarma:2006:STV

[SAHS06] V. V. S. S. Sarma, O. Abe, A. Hinuma, and T. Saino. Short-term variation of triple oxygen isotopes and gross oxygen production in the Sagami Bay, central Japan. Limnology and Oceanography, 51(3):1432–1442, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Sakwinska:2000:TDT

[Sak00] Olga Sakwińska. Trimethylamine does not trigger antipredatory life history shifts in Daphnia. Limnology and Oceanography, 45 (4):988–990, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Salmaso:2005:ECF

[Sal05] Nico Salmaso. Effects of climatic fluctuations and vertical mixing on the interannual trophic variability of Lake Garda, Italy. Limnology and Oceanography, 50(4):553–565, March 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Schmidt:2006:PFS

[SAP^{+}06] Katrin Schmidt, Angus Atkinson, Klaus-Jürgen Petzke, Maren Voss, and David W. Pond. Protozoans as a food source for Antarctic krill, Euphausia superba: Complementary insights from stomach content, fatty acids, and stable isotopes. Limnology and Oceanography, 51(5):2409–2427, September 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Simo:2002:CDD

[SAPA^{+}02] Rafel Simó, Stephen D. Archer, Carlos Pedrós-Alió, Linda Gilpin, and Claire E. Stelfox-Widdicombe. Coupled dynamics of dimethylsulfoniopropionate and dimethylsulfide cycling and the microbial food web in surface waters of the North Atlantic. Limnology and Oceanography, 47(1):53–61, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:2007:ICM

[Sar07] Orlando Sarnelle. Initial conditions mediate the interaction between Daphnia and bloom-forming cyanobacteria. Limnology and Oceanography, 52(5):2120–2127, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:2003:TRA

[SAS^{+}03] Katrin Schmidt, Angus Atkinson, Dorothea Stübing, James W. McClelland, Joseph P. Montoya, and Maren Voss. Trophic relationships among Southern Ocean copepods and krill: Some uses and limitations of a stable isotope approach. Limnology and Oceanography, 48(1):277–289, January 2003. CODEN LI-OCAH. ISSN 0024-3590.

Sarma:2008:SVT

[SAS08] V. V. S. S. Sarma, O. Abe, and T. Saino. Spatial variations in time-integrated plankton metabolic rates in Sagami Bay using triple oxygen isotopes and O2 : Ar ratios. Limnology and Oceanography, 53(5):1776–1783, September 2008. CODEN LI-OCAH. ISSN 0024-3590.

Sellers:2003:PPL

[SB03] Tim Sellers and Paul A. Bukaveckas. Phytoplankton production in a large, regulated river: a modeling and mass balance assessment. Limnology and Oceanography, 48(4):1476–1487, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Solow:2000:CTV

[SBB00] Andrew R. Solow, Stephen M. Bollens, and Andrew Beet. Comparing two vertical plankton distributions. Limnology and Oceanography, 60(3):506–509, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Santos:2008:NBG

[SBC^{+}08] Isaac R. Santos Santos, William C. Burnett, Jeffrey Chanton, Benjamin Mwashote, I. G. N. A. Suryaputra, and Thorsten Dittmar. Nutrient biogeochemistry in a Gulf of Mexico subterranean estuary and groundwater-derived fluxes to the coastal ocean. Limnology and Oceanography, 53(2):705–718, March 2008. CODEN LIOCAH. ISSN 0024-3590.

See:2006:USD

[SBL06] Jason H. See, Deborah A. Bronk, and Alan J. Lewitus. Uptake of spartina-derived humic nitrogen by estuarine phytoplankton in nonaxenic and axenic culture. Limnology and Oceanography, 51(5):2290–2299, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Sutula:2004:ESS

[SBM04] Martha Sutula, Thomas S. Bianchi, and Brent A. McKee. Effect of seasonal sediment storage in the lower Mississippi River on the flux of reactive particulate phosphorus to the Gulf of Mexico. Limnology and Oceanography, 49(6):2223–2235, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Smith:2006:CEF

[SBR^{+}06] K. L,. Jr. Smith, R. J. Baldwin, H. A. Ruhl, M. Kahru, B. G. Mitchell, and R. S. Kaufmann. Climate effect on food supply to depths greater than 4,000 meters in the northeast Pacific. Limnology and Oceanography, 51(1):166–176, January 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Seidendorf:2007:ESR

[SBS07] Bernd Seidendorf, Maarten Boersma, and Klaus Schwenk. Evolutionary stoichiometry: The role of food quality for clonal differentiation and hybrid maintenance in a Daphnia species complex. Limnology and Oceanography, 52(1):385–394, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Stramski:2007:VOP

[SBW07] Dariusz Stramski, Marcel Babin, and Slawomir B. Woźniak. Variations in the optical properties of terrigenous mineralrich particulate matter suspended in seawater. Limnology and Oceanography, 52(6):2418–2433, November 2007. CODEN LI-OCAH. ISSN 0024-3590.

Striebel:2009:CSS

[SBZ^{+}09] Maren Striebel, Silvia Bartholmé, Rebekka Zernecke, Christina Steinlein, Florian Haupt, Sebastian Diehl, and Herwig Stibor. Carbon sequestration and stoichiometry of motile and nonmotile green algae. Limnology and Oceanography, 54(5):1746–1752, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Sanford:2000:MTV

[SC00] Lawrence P. Sanford and Sean M. Crawford. Mass transfer versus kinetic control of uptake across solid-water boundaries. Limnology and Oceanography, 45(5):1180–1186, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Saiz:2007:SFM

[SC07] Enric Saiz and Albert Calbet. Scaling of feeding in marine calanoid copepods. Limnology and Oceanography, 52(2):668– 675, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Stets:2008:IDO

[SC08] Edward G. Stets and James B. Cotner. The influence of dissolved organic carbon on bacterial phosphorus uptake and bacteriaphytoplankton dynamics in two Minnesota lakes. Limnology and Oceanography, 53(1):137–147, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Sanchez-Cabeza:2000:SCC

[SCARM00] J. A. Sanchez-Cabeza, I. Ani-Ragolta, and P. Masquè. Some considerations of the^{210}Pb constant rate of supply (CRS) dating model. Limnology and Oceanography, 45(4):990–995, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Saiz:2003:ESS

[SCB03] Enric Saiz, Albert Calbet, and Elisabetta Broglio. Effects of small-scale turbulence on copepods: The case of Oithona davisae. Limnology and Oceanography, 48(4):1304–1311, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Spivak:2007:TBC

[SCDR07] Amanda C. Spivak, Elizabeth A. Canuel, J. Emmett Duffy, and J. Paul Richardson. Top-down and bottom-up controls on sediment organic matter composition in an experimental seagrass ecosystem. Limnology and Oceanography, 52(6):2595– 2607, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Schell:2000:DCC

[Sch00] Donald M. Schell. Declining carrying capacity in the Bering Sea: Isotopic evidence from whale baleen. Limnology and Oceanography, 45(2):459–462, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Schell:2001:CIR

[Sch01] Donald M. Schell. Carbon isotope ratio variations in Bering Sea biota: The role of anthropogenic carbon dioxide. Limnology and Oceanography, 44(4):999–1000, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Schelske:2006:COF

[Sch06a] Claire L. Schelske. Comment on the origin of the “fluid mud layer” in Lake Apopka, Florida. Limnology and Oceanography, 51(5):2472–2480, September 2006. CODEN LIOCAH. ISSN 0024-3590. See [BHVC05].

Schindler:2006:RAU

[Sch06b] D. W. Schindler. Recent advances in the understanding and management of eutrophication. Limnology and Oceanography, 51(1 (part 2)):356–363, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Sakamoto:2007:PPA

[SCH07] Masaki Sakamoto, Kwang-Hyeon Chang, and Takayuki Hanazato. Plastic phenotypes of antennule shape in Bosmina longirostris controlled by physical stimuli from predators. Limnology and Oceanography, 52(5):2072–2078, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Schindler:2009:LSI

[Sch09] D. W. Schindler. Lakes as sentinels and integrators for the effects of climate change on watersheds, airsheds, and landscapes. Limnology and Oceanography, 54(6 (part 2)):2349–2358, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Smith:2008:EHI

[SCM08] Christopher G. Smith, Jaye E. Cable, and Jonathan B. Martin. Episodic high intensity mixing events in a subterranean estuary: Effects of tropical cyclones. Limnology and Oceanography, 53(2): 666–674, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Selander:2008:ENP

[SCP08] Erik Selander, Gunnar Cervin, and Henrik Pavia. Effects of nitrate and phosphate on grazer-induced toxin production in Alexandrium minutum. Limnology and Oceanography, 53(2): 523–530, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Soranno:2008:FDE

[SCS^{+}08] Patricia A. Soranno, Kendra Spence Cheruvelil, R. Jan Stevenson, Scott L. Rollins, Sarah W. Holden, Sylvia Heaton, and Eric Torng. A framework for developing ecosystem-specific nutrient criteria: Integrating biological thresholds with predictive modeling. Limnology and Oceanography, 53(2):773–787, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Strecker:2004:EEG

[SCV04] Angela L. Strecker, Tyler P. Cobb, and Rolf D. Vinebrooke. Effects of experimental greenhouse warming on phytoplankton and zooplankton communities in fishless alpine ponds. Limnology and Oceanography, 49(4):1182–1190, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Sakai:2000:EMD

[SCY00] Akiko Sakai, Kazuhisa Chikita, and Tomomi Yamada. Expansion of a moraine-dammed glacial lake, Tsho Rolpa, in Rolwaling Himal, Nepal Himalaya. Limnology and Oceanography, 45 (6):1401–1408, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Schenau:2000:NCM

[SD00] S. J. Schenau and G. J. De Lange. A novel chemical method to quantify fish debris in marine sediments. Limnology and Oceanography, 45(4):963–971, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Sakwinska:2005:LHS

[SD05] Olga Sakwińska and Piotr Dawidowicz. Life history strategy and depth selection behavior as alternative antipredator defenses among natural Daphnia hyalina populations. Limnology and Oceanography, 50(4):1284–1289, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Stief:2006:PMF

[SdB06] Peter Stief and Dirk de Beer. Probing the microenvironment of freshwater sediment macrofauna: Implications of depositfeeding and bioirrigation for nitrogen cycling. Limnology and Oceanography, 51(6):2538–2548, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Schlekat:2000:BPA

[SDC00] Christian E. Schlekat, Alan W. Decho, and G. Thomas Chandler. Bioavailability of particle-associated silver, cadmium, and zinc to the estuarine amphipod Leptocheirus plumulosus through dietary ingestion. Limnology and Oceanography, 45(1): 11–21, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Schmid:2007:SSM

[SDG^{+}07] M. Schmid, M. De Batist, N. G. Granin, V. A. Kapitanov, D. F. McGinnis, I. B. Mizandrontsev, A. I. Obzhirov, and A. Wüest. Sources and sinks of methane in Lake Baikal: a synthesis of measurements and modeling. Limnology and Oceanography, 52 (5):1824–1837, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Staahl-Delbanco:2002:EBR

[SDH02] Annika Ståhl-Delbanco and Lars-Anders Hansson. Effects of bioturbation on recruitment of algal cells from the “seed bank” of lake sediments. Limnology and Oceanography, 47(6):1836– 1843, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Sheibley:2003:INT

[SDJT03] Richard W. Sheibley, John H. Duff, Alan P. Jackman, and Frank J. Triska. Inorganic nitrogen transformations in the bed of the Shingobee River, Minnesota: Integrating hydrologic and biological processes using sediment perfusion cores. Limnology and Oceanography, 48(3):1129–1140, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Sobek:2009:OCB

[SDKZ^{+}09] Sebastian Sobek, Edith Durisch-Kaiser, Roland Zurbrügg, Nuttakan Wongfun, Martin Wessels, Natacha Pasche, and Bernhard Wehrli. Organic carbon burial efficiency in lake sediments controlled by oxygen exposure time and sediment source. Limnology and Oceanography, 54(6):2243–2254, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Selvendiran:2009:MDT

[SDM^{+}09] Pranesh Selvendiran, Charles T. Driscoll, Mario R. Montesdeoca, Hyun-Deok Choi, and Thomas M. Holsen. Mercury dynamics and transport in two Adirondack Lakes. Limnology and Oceanography, 54(2):413–427, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Struyf:2006:TMB

[SDV^{+}06] Eric Struyf, Armel Dausse, Stefan Van Damme, Kris Bal, Britta Gribsholt, Henricus T. S. Boschker, Jack J. Middelburg, and Patrick Meire. Tidal marshes and biogenic silica recycling at the land–sea interface. Limnology and Oceanography, 51(2):838–846, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Shiller:2006:POD

[SDvEB06] Alan M. Shiller, Shuiwang Duan, Peter van Erp, and Thomas S. Bianchi. Photo-oxidation of dissolved organic matter in river water and its effect on trace element speciation. Limnology and Oceanography, 52(2):1716–1728, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Schneider:2006:ECC

[SE06] Kenneth Schneider and Jonathan Erez. The effect of carbonate chemistry on calcification and photosynthesis in the hermatypic coral Acropora eurystoma. Limnology and Oceanography, 51(3): 1284–1293, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Schwarz:2008:RMM

[SEC08] Julia I. K. Schwarz, Werner Eckert, and Ralf Conrad. Response of the methanogenic microbial community of a profundal lake sediment (Lake Kinneret, Israel) to algal deposition. Limnology and Oceanography, 53(1):113–121, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Sukenik:2002:IGP

[SEL^{+}02] Assaf Sukenik, Rachel Eshkol, Alexander Livne, Ora Hadas, Meir Rom, Dani Tchernov, Assaf Vardi, and Aaron Kaplan. Inhibition of growth and photosynthesis of the dinoflagellate Peridinium gatunense by Microcystis sp. (cyanobacteria): a novel allelopathic mechanism. Limnology and Oceanography, 47(6): 1656–1663, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Stewart:2003:BPM

[SF03a] Gillian M. Stewart and Nicholas S. Fisher. Bioaccumulation of polonium-210 in marine copepods. Limnology and Oceanography, 49(4):2011–2019, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Stewart:2003:ESAa

[SF03b] Gillian M. Stewart and Nicholas S. Fisher. Experimental studies on the accumulation of polonium-210 by marine phytoplankton. Limnology and Oceanography, 48(3):1193–1201, May 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Stewart:2003:ESAb

[SF03c] Gillian M. Stewart and Nicholas S. Fisher. Experimental studies on the accumulation of polonium-210 by marine phytoplankton. Limnology and Oceanography, 48(4):1720, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Stone:2004:TDM

[SF04] Jeffery Robert Stone and Sherilyn C. Fritz. Three-dimensional modeling of lacustrine diatom habitat areas: Improving paleolimnological interpretation of planktic : benthic ratios. Limnology and Oceanography, 49(5):1540–1548, September 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Schwalbach:2005:WRA

[SF05] Michael S. Schwalbach and Jed A. Fuhrman. Wide-ranging abundances of aerobic anoxygenic phototrophic bacteria in the World Ocean revealed by epifluorescence microscopy and quantitative PCR. Limnology and Oceanography, 50(4):620–628, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Selbie:2009:ELC

[SFB^{+}09] Daniel T. Selbie, Bruce P. Finney, David Barto, Lynda Bunting, Guangjie Chen, Peter R. Leavitt, Erland A. MacIsaac, Daniel E. Schindler, Mark D. Shapley, and Irene Gregory-Eaves. Ecological, landscape, and climatic regulation of sediment geochemistry in North American sockeye salmon nursery lakes: Insights for paleoecological salmon investigations. Limnology and Oceanography, 54(5):1733–1745, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Statzner:2000:CGA

[SFC^{+}00] Bernhard Statzner, Eric Fièvet, Jean-Yves Champagne, Robert Morel, and Eric Herouin. Crayfish as geomorphic agents and ecosystem engineers: Biological behavior affects sand and gravel erosion in experimental streams. Limnology and Oceanography, 45(5):1030–1040, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Stevens:2005:TLB

[SFL05] Craig L. Stevens, Timothy S. R. Fisher, and Gregory A. Lawrence. Turbulent layering beneath the pycnocline in a strongly stratified pit lake. Limnology and Oceanography, 55(3): 197–206, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Six:2008:CPC

[SFR^{+}08] Christophe Six, Zoe V. Finkel, Francisco Rodriguez, Dominique Marie, Frédéric Partensky, and Douglas A. Campbell. Contrasting photoacclimation costs in ecotypes of the marine eukaryotic picoplankter Ostreococcus. Limnology and Oceanography, 53(1): 255–265, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Stachowicz:2002:BRB

[SG02] John J. Stachowicz and Edwin D. Grosholz. Book review: Bertness, Mark D., Steven D. Gaines, and Mark E. Hay [Eds.]. 2001. Marine community ecology. Sinauer Associates. 550 p. US$62.95. ISBN 0-87893-057-4. Limnology and Oceanography, 47(1):332, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Saito:2008:ZCC

[SG08] Mak A. Saito and Tyler J. Goepfert. Zinc-cobalt colimitation of Phaeocystis antarctica. Limnology and Oceanography, 53(1): 266–275, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Sansone:2004:MAW

[SGB04] Francis J. Sansone, Andrew W. Graham, and William M. Berelson. Methane along the western Mexican margin. Limnology and Oceanography, 49(6):2242–2255, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Suarez:2008:TLP

[SGG^{+}08] L. Velo Suárez, S. González Gil, P. Gentien, M. Lunven, C. Bechemin, L. Fernand, R. Raine, and B. Reguera. Thin layers of Pseudo-nitzschia spp. and the fate of Dinophysis acuminata during an upwelling-downwelling cycle in a Galician Ría. Limnology and Oceanography, 53(5):1816–1834, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Shcherbina:2008:MBH

[SGLT08] Andrey Y. Shcherbina, Glen G. Gawarkiewicz, Christopher A. Linder, and Simon R. Thorrold. Mapping bathymetric and hydrographic features of Glover’s Reef, Belize, with a REMUS autonomous underwater vehicle. Limnology and Oceanography, 53 (5 (part 2)):2264–2272, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Saurel:2007:SMF

[SGPK07] C. Saurel, J. C. Gascoigne, M. R. Palmer, and M. J. Kaiser. In situ mussel feeding behavior in relation to multiple environmental factors: Regulation through food concentration and tidal conditions. Limnology and Oceanography, 52(5):1919–1929, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Smith:2009:USS

[SGPM09] Geneviève K. Smith, Frédéric Guichard, Filip Petrović, and Christopher W. McKindsey. Using spatial statistics to infer scales of demographic connectivity between populations of the blue mussel, Mytilus spp. Limnology and Oceanography, 54(3): 970–977, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Saito:2008:STC

[SGR08] Mak A. Saito, Tyler J. Goepfert, and Jason T. Ritt. Some thoughts on the concept of colimitation: Three definitions and the importance of bioavailability. Limnology and Oceanography, 53(1):276–290, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Sieracki:2006:DPA

[SGT^{+}06] Michael E. Sieracki, Ilana C. Gilg, Edward C. Thier, Nicole J. Poulton, and Ralf Goericke. Distribution of planktonic aerobic anoxygenic photoheterotrophic bacteria in the northwest Atlantic. Limnology and Oceanography, 51(1):38–46, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Sunda:2000:EZM

[SH00] William G. Sunda and Susan A. Huntsman. Effect of Zn, Mn, and Fe on Cd accumulation in phytoplankton: Implications for oceanic Cd cycling. Limnology and Oceanography, 45(7):1501– 1516, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Skovgaard:2003:FUH

[SH03] Alf Skovgaard and Per Juel Hansen. Food uptake in the harmful alga Prymnesium parvum mediated by excreted toxins. Limnology and Oceanography, 48(3):1161–1166, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Sunda:2004:RAP

[SH04] William G. Sunda and Susan A. Huntsman. Relationships among photoperiod, carbon fixation, growth, chlorophyll a, and cellular iron and zinc in a coastal diatom. Limnology and Oceanography, 49(5):1742–1753, September 2004. CODEN LI-OCAH. ISSN 0024-3590.

Statham:2005:DIC

[SH05a] P. J. Statham and V. Hart. Dissolved iron in the Cretan Sea (eastern Mediterranean). Limnology and Oceanography, 50(4): 1142–1148, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Sunda:2005:ECS

[SH05b] William G. Sunda and Susan A. Huntsman. Effect of CO_{2} supply and demand on zinc uptake and growth limitation in a coastal diatom. Limnology and Oceanography, 50(4):1181–1192, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Smith:2006:WSN

[SH06] Stephen V. Smith and James T. Hollibaugh. Water, salt, and nutrient exchanges in San Francisco Bay. Limnology and Oceanography, 51(1 (part 2)):504–517, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Sunda:2007:AUG

[SH07] William G. Sunda and D. Ransom Hardison. Ammonium uptake and growth limitation in marine phytoplankton. Limnology and Oceanography, 52(6):2496–2506, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Sereda:2010:CEP

[SH10] Jeff M. Sereda and Jeff J. Hudson. Comparative estimate of P fluxes in lakes: a comment on “Fish decomposition in boreal lakes and biogeochemical implications” by Chidami and Amyot (2008). Limnology and Oceanography, 55(1):463–465, January 2010. CODEN LIOCAH. ISSN 0024-3590. See [CA08] and response [ACD10].

Steen:2008:DDC

[SHA08] Andrew D. Steen, Leila J. Hamdan, and Carol Arnosti. Dynamics of dissolved carbohydrates in the Chesapeake Bay: Insights from enzyme activities, concentrations, and microbial metabolism. Limnology and Oceanography, 53(3):936–947, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Stapel:2001:NRS

[SHBM01] Johan Stapel, Marten A. Hemminga, Cornelis G. Bogert, and Yvonne E. M. Maas. Nitrogen (^{15}N) retention in small Thalassia hemprichii seagrass plots in an offshore meadow in South Sulawesi, Indonesia. Limnology and Oceanography, 46(1):24–37, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Silsbe:2006:VCP

[SHGM06] G. M. Silsbe, R. E. Hecky, S. J. Guildford, and R. Mugidde. Variability of chlorophyll-a and photosynthetic parameters in a nutrient-saturated tropical great lake. Limnology and Oceanography, 51(5):2052–2063, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Shick:2004:CIU

[Shi04] J. Malcolm Shick. The continuity and intensity of ultraviolet irradiation affect the kinetics of biosynthesis, accumulation, and conversion of mycosporine-like amino acids (MAAs) in the coral Stylophora pistillata. Limnology and Oceanography, 49(2):442– 458, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Seitzinger:2005:MLC

[SHL^{+}05] S. P. Seitzinger, H. Hartnett, R. Lauck, M. Mazurek, T. Minegishi, G. Spyres, and R. Styles. Molecular-level chemical characterization and bioavailability of dissolved organic matter in stream water using electrospray-ionization mass spectrometry. Limnology and Oceanography, 50(1):1–12, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Stevens:2001:WMR

[SHS01] C. L. Stevens, C. L. Hurd, and M. J. Smith. Water motion relative to subtidal kelp fronds. Limnology and Oceanography, 46(6):668–678, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Stubing:2003:ULB

[SHS03a] Dorothea Stübing, Wilhelm Hagen, and Katrin Schmidt. On the use of lipid biomarkers in marine food web analyses: an experimental case study on the Antarctic krill, Euphausia superba. Limnology and Oceanography, 48(4):1685–1700, July 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Stumpf:2003:DWD

[SHS03b] Richard P. Stumpf, Kristine Holderied, and Mark Sinclair. Determination of water depth with high-resolution satellite imagery over variable bottom types. Limnology and Oceanography, 48(1 (part 2)):547–556, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Scott:2007:KIE

[SHSH^{+}07] K. M. Scott, M. Henn-Sax, T. L. Harmer, D. L. Longo, C. H. Frame, and C. M. Cavanaugh. Kinetic isotope effect and biochemical characterization of form IA RubisCO from the marine cyanobacterium Prochlorococcus marinus MIT9313. Limnology and Oceanography, 52(5):2199–2204, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Shull:2001:TMM

[Shu01] David H. Shull. Transition-matrix model of bioturbation and radionuclide diagenesis. Limnology and Oceanography, 46(4): 905–916, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Saggio:2001:MTF

[SI01] Angelo Saggio and Jörg Imberger. Mixing and turbulent fluxes in the metalimnion of a stratified lake. Limnology and Oceanography, 46(2):392–409, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Stocker:2003:HTD

[SI03] Roman Stocker and Jörg Imberger. Horizontal transport and dispersion in the surface layer of a medium-sized lake. Limnology and Oceanography, 48(3):971–982, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Shimizu:2008:EDB

[SI08] Kenji Shimizu and Jörg Imberger. Energetics and damping of basin-scale internal waves in a strongly stratified lake. Limnology and Oceanography, 53(5):1574–1588, July 2008. CODEN LIOCAH. ISSN 0024-3590.

SanMartin:2006:VTE

[SIH^{+}06] Elena San Martin, Xabier Irigoien, Roger P. Harris, Ângel López-Urrutia, Mikhail V. Zubkov, and Jane L. Heywood. Variation in the transfer of energy in marine plankton along a productivity gradient in the Atlantic Ocean. Limnology and Oceanography, 51(5):2084–2091, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Shimizu:2007:HSE

[SIK07] Kenji Shimizu, Jörg Imberger, and Michio Kumagai. Horizontal structure and excitation of primary motions in a strongly stratified lake. Limnology and Oceanography, 52(6):2641–2655, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Sharma:2005:MEP

[SIRO05] Kanika Sharma, Patrick W. Inglett, K. Ramesh Reddy, and Andrew V. Ogram. Microscopic examination of photoautotrophic and phosphatase-producing organisms in phosphorus-limited Everglades periphyton mats. Limnology and Oceanography, 50 (6):2057–2062, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Snucins:2000:IVT

[SJ00] Ed Snucins and Gunn John. Interannual variation in the thermal structure of clear and colored lakes. Limnology and Oceanography, 45(7):1639–1646, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Sato:2008:SVV

[SJ08] Mei Sato and Peter A. Jumars. Seasonal and vertical variations in emergence behaviors of Neomysis americana. Limnology and Oceanography, 53(5):1665–1677, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Straile:2003:CEW

[SJH03] Dietmar Straile, Klaus Jöhnk, and Rossknecht Henno. Complex effects of winter warming on the physicochemical characteristics of a deep lake. Limnology and Oceanography, 48(4):1432–1438, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Smith:2006:EFM

[SJH06] Val H. Smith, Samantha B. Joye, and Robert W. Howarth. Eutrophication of freshwater and marine ecosystems. Limnology and Oceanography, 51(1 (part 2)):351–355, January 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Schwenk:2004:EMG

[SJR^{+}04] Klaus Schwenk, Paavo Junttila, Milla Rautio, Finn Bastiansen, Jürgen Knapp, Outi Dove, Raquel Billiones, and Bruno Streit.

Ecological, morphological, and genetic differentiation of Daphnia (Hyalodaphnia) from the Finnish and Russian subarctic. Limnology and Oceanography, 49(2):532–539, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Sand-Jensen:2009:MRS

[SJRB09] Kaj Sand-Jensen, Ane Løvendahl Raun, and Jens Borum. Metabolism and resources of spherical colonies of Nostoc zetterstedtii. Limnology and Oceanography, 54(4):1282–1291, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Stanley:2003:EIM

[SJW03] Emily H. Stanley, Mark D. Johnson, and Amelia K. Ward. Evaluating the influence of macrophytes on algal and bacterial production in multiple habitats of a freshwater wetland. Limnology and Oceanography, 48(3):1101–1111, May 2003. CODEN LI-OCAH. ISSN 0024-3590.

Smith:2001:SSP

[SK01] Erik M. Smith and Michael W. Kemp. Size structure and the production/respiration balance in a coastal plankton community. Limnology and Oceanography, 46(3):473–485, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:2004:ZRA

[SK04] Orlando Sarnelle and Roland A. Knapp. Zooplankton recovery after fish removal: Limitations of the egg bank. Limnology and Oceanography, 49(4 (part 2)):1382–1392, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:2005:NRF

[SK05] Orlando Sarnelle and Roland A. Knapp. Nutrient recycling by fish versus zooplankton grazing as drivers of the trophic cascade in alpine lakes. Limnology and Oceanography, 50(6):2032–2042, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Smith:2001:PBC

[SKBC01] Kenneth, Jr. L. Smith, Ronald S. Kaufmann, Roberta J. Baldwin, and A. F. Carlucci. Pelagic–benthic coupling in the abyssal eastern North Pacific: an 8-year time-series study of food supply and demand. Limnology and Oceanography, 46(3):543–556, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Striegl:2001:CDP

[SKC^{+}01] Robert G. Striegl, Pirkko Kortelainen, Jeffrey P. Chanton, Kimberly P. Wickland, Glynnis C. Bugna, and Miitta Rantakari. Carbon dioxide partial pressure and ^{13}C content of north temperate and boreal lakes at spring ice melt. Limnology and Oceanography, 46(4):941–945, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Suggett:2001:APS

[SKH^{+}01] David Suggett, Gijsbert Kraay, Patrick Holligan, Margaret Davey, Jim Aiken, and Richard Geider. Assessment of photosynthesis in a spring cyanobacterial bloom by use of a fast repetition rate fluorometer. Limnology and Oceanography, 46 (4):802–810, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Shade:2007:IDP

[SKJ^{+}07] Ashley Shade, Angela D. Kent, Stuart E. Jones, Ryan J. Newton, Eric W. Triplett, and Katherine D. McMahon. Interannual dynamics and phenology of bacterial communities in a eutrophic lake. Limnology and Oceanography, 52(2):487–494, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Shaked:2005:GKM

[SKM05] Yeala Shaked, Adam B. Kustka, and François M. M. Morel. A general kinetic model for iron acquisition by eukaryotic phytoplankton. Limnology and Oceanography, 50(3):872–882, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Spiese:2009:RDD

[SKNK09] Christopher E. Spiese, David J. Kieber, Christopher T. Nomura, and Ronald P. Kiene. Reduction of dimethylsulfoxide to dimethylsulfide by marine phytoplankton. Limnology and Oceanography, 54(2):560–570, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Sellers:2001:FMW

[SKR01] Patricia Sellers, Carol A. Kelly, and John W. M. Rudd. Fluxes of methylmercury to the water column of a drainage lake: The relative importance of internal and external sources. Limnology and Oceanography, 46(6):623–631, May 2001. CODEN LI-OCAH. ISSN 0024-3590.

Shikano:2004:VHF

[SKTD04] Shuichi Shikano, Eisuke Kikuchi, Shigeto Takagi, and Hideyuki Doi. Volcanic heat flux and short-term holomixis during the summer stratification period in a crater lake. Limnology and Oceanography, 49(6):2287–2292, November 2004. CODEN LI-OCAH. ISSN 0024-3590.

Strauss:2000:RNA

[SL00] Eric A. Strauss and Gary A. Lamberti. Regulation of nitrification in aquatic sediments by organic carbon. Limnology and Oceanography, 45(8):1854–1859, December 2000. CODEN LI-OCAH. ISSN 0024-3590.

Stelzer:2001:EPR

[SL01] Robert S. Stelzer and Gary A. Lamberti. Effects of n: P ratio and total nutrient concentration on stream periphyton community structure, biomass, and elemental composition. Limnology and Oceanography, 46(2):356–367, March 2001. CODEN LI-OCAH. ISSN 0024-3590.

Sivan:2002:EDS

[SL02] Orit Sivan and Boaz Lazar. ^{14}C excess in deep-sea sediments porewater driven by diffusion — Southeast Mediterranean. Limnology and Oceanography, 47(2):565–570, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Shanks:2000:DOT

[SLB^{+}00] Alan L. Shanks, John Largier, Laura Brink, John Brubaker, and Rian Hooff. Demonstration of the onshore transport of larval invertebrates by the shoreward movement of an upwelling front. Limnology and Oceanography, 45(1):230–236, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Sondergaard:2000:NAF

[SlBWC^{+}00] Morten Søndergaard, Peter J. le B. Williams, Gustave Cauwet, Bo Riemann, Carol Robinson, Senka Terzic, E. Malcolm S. Woodward, and Jakob Worm. Net accumulation and flux of dissolved organic carbon and dissolved organic nitrogen in marine plankton communities. Limnology and Oceanography, 45 (5):1097–1111, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Stige:2009:CFZ

[SLC^{+}09] Leif Chr. Stige, Dmitry L. Lajus, Kung-Sik Chan, Padmini Dalpadado, Sünnje L. Basedow, Igor Berchenko, and Nils Chr.

Stenseth. Climatic forcing of zooplankton dynamics is stronger during low densities of planktivorous fish. Limnology and Oceanography, 54(4):1025–1036, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Silverman:2000:VSP

[SLD00] Harold Silverman, John W. Lynn, and Thomas H. Dietz. In vitro studies of particle capture and transport in suspension-feeding bivalves. Limnology and Oceanography, 45(5):1199–1203, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Schneider:2009:SDC

[SLDE09] Kenneth Schneider, Oren Levy, Zvy Dubinsky, and Jonathan Erez. In situ diel cycles of photosynthesis and calcification in hermatypic corals. Limnology and Oceanography, 54(6):1995– 2002, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Savage:2004:DRE

[SLE04] Candida Savage, Peter R. Leavitt, and Ragnar Elmgren. Distribution and retention of effluent nitrogen in surface sediments of a coastal bay. Limnology and Oceanography, 49(5):1503–1511, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Schladow:2002:OTA

[SLHK02] S. Geoffrey Schladow, Minhee Lee, Bernhard E. Hürzeler, and Peter B. Kelly. Oxygen transfer across the air–water interface by natural convection in lakes. Limnology and Oceanography, 47 (5):1394–1404, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Stepanauskas:2000:HDB

[SLJ00] Ramūnas Stepanauskas, Hjalmar Laudon, and Niels O. G. Jørgensen. High DON bioavailability in boreal streams during a spring flood. Limnology and Oceanography, 45(6):1298–1307, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Sponaugle:2005:FCF

[SLKP05] Su Sponaugle, Thomas Lee, Vassiliki Kourafalou, and Deanna Pinkard. Florida Current frontal eddies and the settlement of coral reef fishes. Limnology and Oceanography, 50(4):1033–1048, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Sundback:2004:BNF

[SLL^{+}04] Kristina Sundbäck, Frank Linares, Fredrik Larson, Angela Wulff, and Anna Engelsen. Benthic nitrogen fluxes along a depth gradient in a microtidal fjord: The role of denitrification and microphytobenthos. Limnology and Oceanography, 49 (4):1095–1107, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Stephens:2003:EMC

[SLRM03] F. Carol Stephens, Eric M. Louchard, R. Pamela Reid, and Robert A. Maffione. Effects of microalgal communities on reflectance spectra of carbonate sediments in subtidal optically shallow marine environments. Limnology and Oceanography, 48 (1 (part 2)):535–546, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Suzuki:2002:EWG

[SLS^{+}02] Koji Suzuki, Hongbin Liu, Toshiro Saino, Hajime Obata, Michiaki Takano, Kei Okamura, Yoshiki Sohrin, and Yukiyo Fujishima. East–west gradients in the photosynthetic potential of phytoplankton and iron concentration in the subarctic Pacific Ocean during early summer. Limnology and Oceanography, 47 (6):1581–1594, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Stedmon:2001:OCD

[SM01] C. A. Stedmon and S. Markager. The optics of chromophoric dissolved organic matter (CDOM) in the Greenland Sea: an algorithm for differentiation between marine and terrestrially derived organic matter. Limnology and Oceanography, 47(4): 2087–2093, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Shull:2002:DPR

[SM02] David H. Shull and Lawrence M. Mayer. Dissolution of particlereactive radionuclides in deposit-feeder digestive fluids. Limnology and Oceanography, 47(5):1530–1536, September 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Stedmon:2005:RVD

[SM05a] Colin A. Stedmon and Stiig Markager. Resolving the variability in dissolved organic matter fluorescence in a temperate estuary and its catchment using PARAFAC analysis. Limnology and Oceanography, 50(4):686–697, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Stedmon:2005:TPD

[SM05b] Colin A. Stedmon and Stiig Markager. Tracing the production and degradation of autochthonous fractions of dissolved organic matter by fluorescence analysis. Limnology and Oceanography, 50(5):1415–1426, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Steward:2000:GSD

[SMA00] Grieg F. Steward, Janine L. Montiel, and Farooq Azam. Genome size distributions indicate variability and similarities among marine viral assemblages from diverse environments. Limnology and Oceanography, 45(8):1697–1706, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Schoch:2006:FDS

[SMA^{+}06] G. Carl Schoch, Bruce A. Menge, Gary Allison, Maria Kavanaugh, Sarah Ann Thompson, and Spencer A. Wood. Fifteen degrees of separation: Latitudinal gradients of rocky intertidal biota along the California Current. Limnology and Oceanography, 51(6):2564–2585, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Sickman:2003:ENE

[SMC03] James O. Sickman, John M. Melack, and David W. Clow. Evidence for nutrient enrichment of high-elevation lakes in the Sierra Nevada, California. Limnology and Oceanography, 49(4): 1885–1892, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Sohm:2008:ARP

[SMC08] Jill A. Sohm, Claire Mahaffey, and Douglas G. Capone. Assessment of relative phosphorus limitation of Trichodesmium spp. in the North Pacific, North Atlantic, and the north coast of Australia. Limnology and Oceanography, 53(6):2495–2502, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Saito:2002:CLU

[SMCW02] Mak A. Saito, James W. Moffett, Sallie W. Chisholm, and John B. Waterbury. Cobalt limitation and uptake in Prochlorococcus. Limnology and Oceanography, 47(6):1629–1636, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Scully:2004:EDP

[SMD^{+}04] Norman M. Scully, Nagamitsu Maie, Susan K. Dailey, Joseph N. Boyer, Ronald D. Jones, and Rudolf Jaffé. Early diagenesis of plant-derived dissolved organic matter along a wetland, mangrove, estuary ecotone. Limnology and Oceanography, 49(5): 1667–1678, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Swadling:2004:LCP

[SMDG04] K. M. Swadling, A. D. McKinnon, G. De’ath, and J. A. E. Gibson. Life cycle plasticity and differential growth and development in marine and lacustrine populations of an Antarctic copepod. Limnology and Oceanography, 49(3):644–655, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Suggett:2006:GPL

[SMG06] David J. Suggett, Stephen C. Maberly, and Richard J. Geider. Gross photosynthesis and lake community metabolism during the spring phytoplankton bloom. Limnology and Oceanography, 51(5):2064–2076, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Shellenbarger:2006:ISG

[SMGP06] Gregory G. Shellenbarger, Stephen G. Monismith, Amatzia Genin, and Adina Paytan. The importance of submarine groundwater discharge to the near shore nutrient supply in the Gulf of Aqaba (Israel). Limnology and Oceanography, 53(2):1876–1886, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Suzumura:2002:CPM

[SMH^{+}02] Masahiro Suzumura, Toshihiro Miyajima, Hiroshi Hata, Yu Umezawa, Hajime Kayanne, and Isao Koike. Cycling of phosphorus maintains the production of microphytobenthic communities in carbonate sediments of a coral reef. Limnology and Oceanography, 47(3):771–781, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Soetaert:2006:LTC

[SMH^{+}06] Karline Soetaert, Jack J. Middelburg, Carlo Heip, Patrick Meire, Stefan Van Damme, and Tom Maris. Long-term change in dissolved inorganic nutrients in the heterotrophic Scheldt estuary (Belgium, The Netherlands). Limnology and Oceanography, 51 (1 (part 2)):409–423, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Smith:2000:RSE

[Smi00] Walker O. Smith, Jr. Ross Sea ecology: Italiantartide expeditions (1987–1995). Limnology and Oceanography, 45(8):1887, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Smith:2006:REC

[Smi06] Val H. Smith. Responses of estuarine and coastal marine phytoplankton to nitrogen and phosphorus enrichment. Limnology and Oceanography, 51(1 (part 2)):377–384, January 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Simons:2006:ZRE

[SMJ^{+}06] Rachel D. Simons, Stephen G. Monismith, Ladd E. Johnson, Gesche Winkler, and François J. Saucier. Zooplankton retention in the estuarine transition zone of the St. Lawrence Estuary. Limnology and Oceanography, 51(6):2621–2631, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Strom:2007:MGC

[SMO07] Suzanne L. Strom, Erin L. Macri, and M. Brady Olson. Microzooplankton grazing in the coastal Gulf of Alaska: Variations in top-down control of phytoplankton. Limnology and Oceanography, 52(4):1480–1494, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Sundareshwar:2001:OEI

[SMP^{+}01] P. V. Sundareshwar, J. T. Morris, P. J. Pellechia, H. J. Cohen, D. E. Porter, and B. C. Jones. Occurrence and ecological implications of pyrophosphate in estuaries. Limnology and Oceanography, 46(6):1570–1577, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Secord:2005:SDA

[SMP05] David Secord and Giséle Muller-Parker. Symbiont distribution along a light gradient within an intertidal cave. Limnology and Oceanography, 55(3):272–278, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Sharples:2001:PDS

[SMR^{+}01] Jonathan Sharples, Mark C. Moore, Tom P. Rippeth, Patrick M. Holligan, David J. Hydes, Neil R. Fisher, and John H. Simpson. Phytoplankton distribution and survival in the thermocline. Limnology and Oceanography, 46(3):486–496, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Stacey:2007:CDT

[SMS07] Mark T. Stacey, Margaret A. McManus, and Jonah V. Steinbuck. Convergences and divergences and thin layer formation and maintenance. Limnology and Oceanography, 52(4):1523– 1532, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Stacey:2009:RPL

[SMS09] Mark Stacey, Margaret McManus, and Jonah Steinbuck. Response to “Plankton layer profiles as determined by shearing, sinking, and swimming” by D. A. Birch, W. R. Young, and P. J. S. Franks. Limnology and Oceanography, 54(1):400, January 2009. CODEN LIOCAH. ISSN 0024-3590. See [BYF09].

Simon:2007:NDG

[SNFT07] Kevin S. Simon, Dev K. Niyogi, Russell D. Frew, and Colin R. Townsend. Nitrogen dynamics in grassland streams along a gradient of agricultural development. Limnology and Oceanography, 52(3):1246–1257, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Shikata:2008:FII

[SNM^{+}08] Tomoyuki Shikata, Sou Nagasoe, Tadashi Matsubara, Souta Yoshikawa, Yasuhiro Yamasaki, Yohei Shimasaki, Yuji Oshima, Ian R. Jenkinson, and Tsuneo Honjo. Factors influencing the initiation of blooms of the raphidophyte Heterosigma akashiwo and the diatom Skeletonema costatum in a port in Japan. Limnology and Oceanography, 53(6):2503–2518, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Sobrino:2009:ECI

[SNPK^{+}09] Cristina Sobrino, P. J. Neale, J. D. Phillips-Kress, R. E. Moeller, and J. A. Porter. Elevated CO_{2} increases sensitivity to ultraviolet radiation in lacustrine phytoplankton assemblages. Limnology and Oceanography, 54(6 (part 2)):2448–2459, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Saito:2007:CNF

[SO07] Hiroaki Saito and Kazufumi Osako. Confirmation of a new food chain utilizing geothermal energy: Unusual fatty acids of a deep-sea bivalve, Calyptogena phaseoliformis. Limnology and Oceanography, 52(5):1910–1918, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Sosik:2003:GRC

[SON^{+}03] Heidi M. Sosik, Robert J. Olson, Michael G. Neubert, Alexi Shalapyonok, and Andrew R. Solow. Growth rates of coastal phytoplankton from time-series measurements with a submersible flow cytometer. Limnology and Oceanography, 48(5):1756–1765, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Sosiak:2000:NRB

[Sos00] Al Sosiak. Nutrient reduction and biomanipulation as tools to improve water quality: The Lake Rinsjön story. Limnology and Oceanography, 45(7):1679, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Swart:2002:OSV

[SP02] P. K. Swart and R. Price. Origin of salinity variations in Florida Bay. Limnology and Oceanography, 47(4):1234–1241, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Smith:2004:BMG

[SP04] Erik M. Smith and Yves T. Prairie. Bacterial metabolism and growth efficiency in lakes: The importance of phosphorus availability. Limnology and Oceanography, 49(1):137–147, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Shanksa:2009:ART

[SP09] Alan L. Shanks and Catherine A. Pfister. Annual recruitment of three species of tide-pool fishes is driven by variation in springtime coastal hydrodynamics. Limnology and Oceanography, 54 (5):1481–1487, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Spero:2000:MF

[Spe00] Howard J. Spero. Modern Foraminifera. Limnology and Oceanography, 45(8):1890, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Simis:2005:RSC

[SPG05] Stefan G. H. Simis, Steef W. M. Peters, and Herman J. Gons. Remote sensing of the cyanobacterial pigment phycocyanin in turbid inland water. Limnology and Oceanography, 55(3):237– 245, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Schofield:2004:EIB

[SPM04] Kate A. Schofield, Catherine M. Pringle, and Judy L. Meyer. Effects of increased bedload on algal- and detrital-based stream food webs: Experimental manipulation of sediment and macroconsumers. Limnology and Oceanography, 49(4):900–909, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Socki:2002:SIS

[SPR02] Richard A. Socki, Eugene,, Jr. C. Perry, and Christopher S. Romanek. Stable isotope systematics of two cenotes from the northern Yucatan Peninsula, Mexico. Limnology and Oceanography, 47(6):1808–1818, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Singer:2005:FHE

[SPW^{+}05] Gabriel A. Singer, Michaela Panzenböck, Gabriele Weigelhofer, Christina Marchesani, Johann Waringer, Wolfgang Wanek, and Tom J. Battin. Flow history explains temporal and spatial variation of carbon fractionation in stream periphyton. Limnology and Oceanography, 50(4):706–712, March 2005. CODEN LI-OCAH. ISSN 0024-3590.

Scheffer:2003:FFW

[SPZ03] Marten Scheffer, Rob Portielje, and Luis Zambrano. Fish facilitate wave resuspension of sediment. Limnology and Oceanography, 49(4):1920–1926, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Sylvan:2007:EIP

[SQTA07] J. B. Sylvan, A. Quigg, S. Tozzi, and J. W. Ammerman. Eutrophication-induced phosphorus limitation in the Mississippi River plume: Evidence from fast repetition rate fluorometry. Limnology and Oceanography, 52(6):2679–2685, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Smith:2002:WCM

[SR02] Craig R. Smith and Christophe Rabouille. What controls the mixed-layer depth in deep-sea sediments? The importance of POC flux. Limnology and Oceanography, 47(2):418–426, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Smith:2004:RCB

[SR04] Craig R. Smith and Christophe Rabouille. Reply to comment by Boudreau on: What controls the mixed-layer depth in deepsea sediments? The importance of POC flux. Limnology and Oceanography, 49(2):623–624, March 2004. CODEN LIOCAH. ISSN 0024-3590. See [Bou04].

Simon:2007:DCD

[SR07] Meinhard Simon and Bernd Rosenstock. Different coupling of dissolved amino acid, protein, and carbohydrate turnover to heterotrophic picoplankton production in the Southern Ocean in austral summer and fall. Limnology and Oceanography, 52(1): 85–95, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Serret:2001:LVB

[SRF^{+}01] Pablo Serret, Carol Robinson, Emilio Fernández, Eva Teira, and Gavin Tilstone. Latitudinal variation of the balance between plankton photosynthesis and respiration in the eastern Atlantic Ocean. Limnology and Oceanography, 46(7):1642–1652, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Slawyk:2000:UMD

[SRG00] Gerd Slawyk, Patrick Raimbault, and Nicole Garcia. Use of ^{15}N to measure dissolved organic nitrogen release by marine phytoplankton (reply to comment by Bronk and Ward). Limnology and Oceanography, 45(8):1884–1886, December 2000. CODEN LIOCAH. ISSN 0024-3590. See [BW99].

Smith:2008:TAF

[SRKK08] K. L. Smith, Jr., H. A. Ruhl, R. S. Kaufmann, and M. Kahru. Tracing abyssal food supply back to upper-ocean processes over a 17-year time series in the northeast Pacific. Limnology and Oceanography, 53(6):2655–2667, November 2008. CODEN LI-OCAH. ISSN 0024-3590.

Statzner:2005:MGC

[SrL05] Bernhard Statzner, Sylvie Mé rigoux, and Maria Leichtfried. Mineral grains in caddisfly pupal cases and streambed sediments: Resource use and its limitation through conflicting resource requirements. Limnology and Oceanography, 50(4):713– 721, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Schloesser:2009:IVS

[SRL^{+}09] Ryan W. Schloesser, Jay R. Rooker, Patrick Louchuoarn, John D. Neilson, and David H. Secord. Interdecadal variation in seawater δ ^{13}C and δ ^{18}O recorded in fish otoliths. Limnology and Oceanography, 54(5):1665–1668, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Saito:2005:PCB

[SRM05] Mak A. Saito, Gabrielle Rocap, and James W. Moffett. Production of cobalt binding ligands in a Synechococcus feature at the Costa Rica upwelling dome. Limnology and Oceanography, 55 (3):279–290, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Salmon:2006:FMI

[SRNW06] Tim P. Salmon, Andrew L. Rose, Brett A. Neilan, and T. David Waite. The FeL model of iron acquisition: Nondissociative reduction of ferric complexes in the marine environment. Limnology and Oceanography, 52(2):1744–1754, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Scavia:2003:PRG

[SRT^{+}03] Donald Scavia, Nancy N. Rabalais, R. Eugene Turner, Dubravko Justić, and William J. Wiseman, Jr. Predicting the response of Gulf of Mexico hypoxia to variations in Mississippi River nitrogen load. Limnology and Oceanography, 48(3):951–956, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Simon:2004:CEM

[SRZ04] Meinhard Simon, Bernd Rosenstock, and Walter Zwisler. Coupling of epipelagic and mesopelagic heterotrophic picoplankton production to phytoplankton biomass in the Antarctic polar frontal region. Limnology and Oceanography, 49(4):1035–1043, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Sterner:2001:DIF

[SS01] Robert W. Sterner and Michael S. Schwalbach. Diel integration of food quality by Daphnia: Luxury consumption by a freshwater planktonic herbivore. Limnology and Oceanography, 46(2): 410–416, March 2001. CODEN LIOCAH. ISSN 0024-3590. See erratum [SS04].

Stelzer:2003:ISR

[SS03] C. P. Stelzer and T. W. Snell. Induction of sexual reproduction in Brachionus plicatilis (Monogononta, Rotifera) by a densitydependent chemical cue. Limnology and Oceanography, 48(2): 939–943, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Sterner:2004:EDI

[SS04] Robert W. Sterner and Michael S. Schwalbach. Erratum: Diel integration of food quality by Daphnia: Luxury consumption by a freshwater planktonic herbivore. Limnology and Oceanography, 49(1):314, January 2004. CODEN LIOCAH. ISSN 0024-3590. See [SS01].

Stelzer:2006:SCR

[SS06] Claus-Peter Stelzer and Terry W. Snell. Specificity of the crowding response in the Brachionus plicatilis species complex. Limnology and Oceanography, 51(1):125–130, January 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Schluter:2004:SDB

[SSA^{+}04] Michael Schlüter, Eberhard J. Sauter, Claus E. Andersen, Henning Dahlgaard, and Paul R. Dando. Spatial distribution and budget for submarine groundwater discharge in Eckernförde Bay (Western Baltic Sea). Limnology and Oceanography, 49(1):157– 167, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Stead:2005:SPS

[SSAH05] Tracey K. Stead, Jenny M. Schmid-Araya, and Alan G. Hildrew. Secondary production of a stream metazoan community: Does the meiofauna make a difference? Limnology and Oceanography, 55(3):398–403, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Shimeta:2001:IBB

[SSAZ01] Jeff Shimeta, Victoria R. Starczak, Oladipo M. Ashiru, and Cheryl Ann Zimmer. Influences of benthic boundary-layer flow on feeding rates of ciliates and flagellates at the sediment– water interface. Limnology and Oceanography, 46(7):1709–1719, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Stramski:2002:ETN

[SSC02] Dariusz Stramski, Antoine Sciandra, and Hervé Claustre. Effects of temperature, nitrogen, and light limitation on the optical properties of the marine diatom Thalassiosira pseudonana. Limnology and Oceanography, 47(2):392–403, March 2002. CO-DEN LIOCAH. ISSN 0024-3590.

Seebens:2007:PDF

[SSH^{+}07] Hanno Seebens, Dietmar Straile, Rainer Hoegg, Hans-Bernd Stich, and Ulrich Einsle. Population dynamics of a freshwater calanoid copepod: Complex responses to changes in trophic status and climate variability. Limnology and Oceanography, 52 (6):2364–2372, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Sampei:2009:SCP

[SSH^{+}09] Makoto Sampei, Hiroshi Sasaki, Hiroshi Hattori, Alexandre Forest, and Louis Fortier. Significant contribution of passively sinking copepods to downward export flux in Arctic waters. Limnology and Oceanography, 54(6):1894–1900, November 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Staehr:2007:TDR

[SSJ07] Peter A. Staehr and Kaj Sand-Jensen. Temporal dynamics and regulation of lake metabolism. Limnology and Oceanography, 52 (1):108–120, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Schelske:2006:HLL

[SSK06] Claire L. Schelske, Eugene F. Stoermer, and William F. Kenney. Historic low-level phosphorus enrichment in the Great Lakes inferred from biogenic silica accumulation in sediments. Limnology and Oceanography, 51(1 (part 2)):728–748, January 2006. CO-DEN LIOCAH. ISSN 0024-3590.

Swart:2005:TSV

[SSL05] Peter K. Swart, Amel Saied, and Kathryn Lamb. Temporal and spatial variation in the δ ^{15}N and δ ^{13}C of coral tissue and zooxanthellae in Montastraea faveolata collected from the Florida reef tract. Limnology and Oceanography, 50(4):1049–1058, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Scheuerell:2002:EAF

[SSLE02] Mark D. Scheuerell, Daniel E. Schindler, Arni H. Litt, and W. T. Edmondson. Environmental and algal forcing of Daphnia production dynamics. Limnology and Oceanography, 47(5):1477– 1485, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Sterner:2004:PTM

[SSM^{+}04] Robert W. Sterner, Tanya M. Smutka, R. Michael L. McKay, Qin Xiaoming, Erik T. Brown, and Robert M. Sherrell. Phosphorus and trace metal limitation of algae and bacteria in Lake Superior. Limnology and Oceanography, 49(2):495–507, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Steinbuck:2009:OTM

[SSM^{+}09] Jonah V. Steinbuck, Mark T. Stacey, Margaret A. McManus, Olivia M. Cheriton, and John P. Ryan. Observations of turbulent mixing in a phytoplankton thin layer: Implications for formation, maintenance, and breakdown. Limnology and Oceanography, 54(4):1353–1368, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Stramska:2000:EAB

[SSMM00] Malgorzata Stramska, Dariusz Stramski, B. Greg Mitchell, and Curtis D. Mobley. Estimation of the absorption and backscattering coefficients from in-water radiometric measurements. Limnology and Oceanography, 45(3):628–641, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Seitzinger:2002:BDN

[SSS02] Sybil P. Seitzinger, R. W. Sanders, and Renée Styles. Bioavailability of DON from natural and anthropogenic sources to estuarine plankton. Limnology and Oceanography, 47(2):353–366, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Steinke:2006:DST

[SSS06] Michael Steinke, Jacqueline Stefels, and Eize Stamhuis. Dimethyl sulfide triggers search behavior in copepods. Limnology and Oceanography, 53(2):1925–1930, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Striebel:2008:LIC

[SSS08] Maren Striebel, Gertrud Spörl, and Herwig Stibor. Lightinduced changes of plankton growth and stoichiometry: Experiments with natural phytoplankton communities. Limnology and Oceanography, 53(2):513–522, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Smith:2004:NGD

[SSV^{+}04] Jennifer E. Smith, Celia M. Smith, Peter S. Vroom, Kevin L. Beach, and Steven Miller. Nutrient and growth dynamics of Halimeda tuna on Conch Reef, Florida Keys: Possible influence of internal tides on nutrient status and physiology. Limnology and Oceanography, 49(6):1923–1936, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Scheffer:2001:CWC

[SSvNH01] Marten Scheffer, Dietmar Straile, Egbert H. van Nes, and Harry Hosper. Climatic warming causes regime shifts in lake food webs. Limnology and Oceanography, 46(7):1780–1783, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Scheffer:2003:CER

[SSvNH03] Marten Scheffer, Dietmar Straile, Egbert H. van Nes, and Harry Hosper. Climatic effects on regime shifts in lakes: a reply. Limnology and Oceanography, 48(4):1353–1356, May 2003. CODEN LIOCAH. ISSN 0024-3590. See [JSJ03].

Stead:2003:ESS

[ST03] Robert A. Stead and Raymond J. Thompson. The effect of the sinking spring diatom bloom on digestive processes of the coldwater protobranch Yoldia hyperborea. Limnology and Oceanography, 48(1):157–167, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Saijo:2004:CPP

[ST04] Sachiko Saijo and Eiichiro Tanoue. Characterization of particulate proteins in Pacific surface waters. Limnology and Oceanography, 49(4):953–963, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Stanford:2001:BRJ

[Sta01] Jack A. Stanford. Book review: Jones, J. B., and P. J. Mulholland [Eds.]. 1999. Streams and ground waters. Academic Press. xix + 425 p. $79.95. ISBN 0-12-389845-5. Limnology and Oceanography, 46(3):1582, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Scully:2003:PEI

[STC03] Norman M. Scully, Lars J. Tranvik, and William J. Cooper. Photochemical effects on the interaction of enzymes and dissolved organic matter in natural waters. Limnology and Oceanography, 48(5):1818–1824, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Sharples:2007:SNM

[STG^{+}07] Jonathan Sharples, Jacqueline F. Tweddle, J. A. Mattias Green, Matthew R. Palmer, Young-Nam Kim, Anna E. Hickman, Patrick M. Holligan, C. Mark Moore, Tom P. Rippeth, John H.

Simpson, and Vladimir Krivtsov. Spring-neap modulation of internal tide mixing and vertical nitrate fluxes at a shelf edge in summer. Limnology and Oceanography, 52(5):1735–1747, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Simis:2007:OSF

[STHG07] Stefan G. H. Simis, Marjolijn Tijdens, Hans L. Hoogveld, and Herman J. Gons. Optical signatures of the filamentous cyanobacterium Leptolyngbya boryana during mass viral lysis. Limnology and Oceanography, 52(1):184–197, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Suthers:2004:EAI

[STK^{+}04] Iain M. Suthers, C. T. Taggart, D. Kelley, D. Rissik, and J. H. Middleton. Entrainment and advection in an island’s tidal wake, as revealed by light attenuance, zooplankton, and ichthyoplankton. Limnology and Oceanography, 49(1):283–296, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Staal:2007:NFA

[StLHB^{+}07] Marc Staal, Sacco te Lintel Hekkert, Geert Jan Brummer, Marcel Veldhuis, Cor Sikkens, Stefan Persijn, and Lucas J. Stal. Nitrogen fixation along a north-south transect in the eastern Atlantic Ocean. Limnology and Oceanography, 52(4):1305–1316, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Sobek:2007:PRD

[STP^{+}07] Sebastian Sobek, Lars J. Tranvik, Yves T. Prairie, Pirkko Kortelainen, and Jonathan J. Cole. Patterns and regulation of dissolved organic carbon: an analysis of 7,500 widely distributed lakes. Limnology and Oceanography, 52(3):1208–1219, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Smyntek:2007:SPS

[STST07] Peter M. Smyntek, Mark A. Teece, Kimberly L. Schulz, and Stephen J. Thackeray. A standard protocol for stable isotope analysis of zooplankton in aquatic food web research using mass balance correction models. Limnology and Oceanography, 52(5): 2135–2146, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Steinberg:2008:BVZ

[SVB^{+}08] Deborah K. Steinberg, Benjamin A. S. Van Mooy, Ken O. Buesseler, Philip W. Boyd, Toru Kobari, and David M. Karl. Bacterial vs. zooplankton control of sinking particle flux in the ocean’s twilight zone. Limnology and Oceanography, 53(4):1327–1338, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Serra:2007:RSV

[SVC^{+}07] T. Serra, J. Vidal, X. Casamitjana, M. Soler, and J. Colomer. The role of surface vertical mixing in phytoplankton distribution in a stratified reservoir. Limnology and Oceanography, 52(2): 620–634, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Sell:2001:POC

[SvKM01] Anne F. Sell, Donna van Keuren, and Laurence P. Madin. Predation by omnivorous copepods on early developmental stages of Calanus finmarchicus and Pseudocalanus spp. Limnology and Oceanography, 46(4):953–959, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Vidal:2009:AME

[SVPK^{+}09] A. Sanchez-Vidal, C. Pasqual, P. Kerhervé, S. Heussner, A. Calafat, A. Palanques, X. Durrieu de Madron, M. Canals, and P. Puig. Across margin export of organic matter by cascading events traced by stable isotopes, northwestern Mediterranean Sea. Limnology and Oceanography, 54(5):1488–1500, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:2005:LAD

[SW05a] Orlando Sarnelle and Alan E. Wilson. Local adaptation of Daphnia pulicaria to toxic cyanobacteria. Limnology and Oceanography, 50(5):1565–1570, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Stramski:2005:RCP

[SW05b] Dariusz Stramski and Slawomir B. Woźniak. On the role of colloidal particles in light scattering in the ocean. Limnology and Oceanography, 50(5):1581–1591, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Schoemann:2001:EPA

[SWCL01] V. Schoemann, R. Wollast, L. Chou, and C. Lancelot. Effects of photosynthesis on the accumulation of Mn and Fe by Phaeocystis colonies. Limnology and Oceanography, 46(5):1065–1076, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Sahling:2005:PHS

[SWD^{+}05] Heiko Sahling, Klaus Wallmann, Anke Dählmann, Rolf Schmaljohann, and Sven Petersen. The physicochemical habitat of Sclerolinum sp. at Hook Ridge hydrothermal vent, Bransfield Strait, Antarctica. Limnology and Oceanography, 50(4):598–606, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Stramski:2004:OPA

[SWF04] Dariusz Stramski, Slawomir B. Woźniak, and Piotr J. Flatau. Optical properties of Asian mineral dust suspended in seawater. Limnology and Oceanography, 49(3):749–755, May 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Strom:2003:CDMa

[SWH^{+}03] Suzanne Strom, Gordon Wolfe, Jan Holmes, Hilmar Stecher, Carolyn Shimeneck, and Lambert Sarah. Chemical defense in the microplankton i: Feeding and growth rates of heterotrophic protists on the DMS-producing phytoplankter Emiliania huxleyi. Limnology and Oceanography, 48(1):217–229, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:2005:CIB

[SWH^{+}05] Orlando Sarnelle, Alan E. Wilson, Stephen K. Hamilton, Lesley B. Knoll, and David F. Raikow. Complex interactions between the zebra mussel, Dreissena polymorpha, and the harmful phytoplankter, Microcystis aeruginosa. Limnology and Oceanography, 50(3):896–904, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Southwell:2008:SFD

[SWML08] Melissa W. Southwell, Jeremy B. Weisz, Christopher S. Martens, and Niels Lindquist. In situ fluxes of dissolved inorganic nitrogen from the sponge community on Conch Reef, Key Largo, Florida. Limnology and Oceanography, 54(3):986–996, May 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Sobrino:2008:AEC

[SWN08] Cristina Sobrino, Mary Love Ward, and Patrick J. Neale. Acclimation to elevated carbon dioxide and ultraviolet radiation in the diatom Thalassiosira pseudonana: Effects on growth, photosynthesis, and spectral sensitivity of photoinhibition. Limnology and Oceanography, 53(2):494–505, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Strom:2003:CDMb

[SWS^{+}03] Suzanne Strom, Gordon Wolfe, Amy Slajer, Sarah Lambert, and Jennifer Clough. Chemical defense in the microplankton II: Inhibition of protist feeding by β-dimethylsulfoniopropionate (DMSP). Limnology and Oceanography, 48(1):230–237, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Shank:2006:PSC

[SWS^{+}06] G. Christopher Shank, Robert F. Whitehead, Michelle L. Smith, Stephen A. Skrabal, and Robert J. Kieber. Photodegradation of strong copper-complexing ligands in organic-rich estuarine waters. Limnology and Oceanography, 51(2):884–892, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Skoog:2002:MGU

[SWSJ02] Annelie Skoog, Kenia Whitehead, Frank Sperling, and Karen Junge. Microbial glucose uptake and growth along a horizontal nutrient gradient in the North Pacific. Limnology and Oceanography, 47(6):1676–1683, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Shaked:2006:ZAA

[SXLM06] Yeala Shaked, Yan Xu, Karine Leblanc, and François M. M. Morel. Zinc availability and alkaline phosphatase activity in Emiliania huxleyi: Implications for Zn-p co-limitation in the ocean. Limnology and Oceanography, 51(1):299–309, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Smith:2007:OBM

[SY07] S. Lan Smith and Yasuhiro Yamanaka. Optimization-based model of multinutrient uptake kinetics. Limnology and Oceanography, 52(4):1545–1558, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Szymczak-Zyla:2008:IMC

[SŻKL08] Malgorzata Szymczak-Żyla, Grażyna Kowalewska, and J. William Louda. The influence of microorganisms on chlorophyll-a degradation in the marine environment. Limnology and Oceanography, 57(3):851–862, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Tagliabue:2003:ALZ

[TA03] Alessandro Tagliabue and Kevin R. Arrigo. Anomalously low zooplankton abundance in the Ross Sea: an alternative explanation. Limnology and Oceanography, 48(2):686–699, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Trainer:2000:DAP

[TAB^{+}00] Vera L. Trainer, Nicolaus G. Adams, Brian D. Bill, Carla M. Stehr, John C. Wekell, Peter Moeller, Mark Busman, and Dana Woodruff. Domoic acid production near California coastal upwelling zones, June 1998. Limnology and Oceanography, 45(8): 1818–1833, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Teranes:2000:RNU

[TB00] Jane L. Teranes and Stefano M. Bernasconi. The record of nitrate utilization and productivity limitation provided by δ ^{15}N values in lake organic matter — a study of sediment trap and core sediments from baldeggersee, Switzerland. Limnology and Oceanography, 45(4):801–813, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Teranes:2005:FCV

[TB05] Jane L. Teranes and Stefano M. Bernasconi. Factors controlling δ ^{13}C values of sedimentary carbon in hypertrophic baldeggersee, Switzerland, and implications for interpreting isotope excursions in lake sedimentary records. Limnology and Oceanography, 50 (3):914–922, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Tremblay:2009:OMD

[TB09] Luc Tremblay and Ronald Benner. Organic matter diagenesis and bacterial contributions to detrital carbon and nitrogen in the Amazon River system. Limnology and Oceanography, 54(3): 681–691, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Twining:2004:ESI

[TBF04] Benjamin S. Twining, Stephen B. Baines, and Nicholas S. Fisher. Element stoichiometries of individual plankton cells collected during the Southern Ocean iron experiment (SOFeX). Limnology and Oceanography, 49(6):2115–2128, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Tessier:2001:GRI

[TBG01] Alan J. Tessier, Elena V. Bizina, and Kevin C. Geedey. Grazer– resource interactions in the plankton: Are all daphniids alike. Limnology and Oceanography, 46(7):1585–1595, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Tobias:2007:OIA

[TBH07] Craig R. Tobias, John Karl Böhlke, and Judson W. Harvey. The oxygen-18 isotope approach for measuring aquatic metabolism in high productivity waters. Limnology and Oceanography, 52 (4):1439–1453, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Torres:2001:RHF

[TBHS01] M. E. Torres, J. P. Barry, D. A. Hubbard, and E. Suess. Reconstructing the history of fluid flow at cold seep sites from Ba/ Ca ratios in vesicomyid clam shells. Limnology and Oceanography, 46(7):1701–1708, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Trees:2000:APV

[TCB^{+}00] Charles C. Trees, Dennis K. Clark, Robert R. Bidigare, Michael E. Ondrusek, and James L. Mueller. Accessory pigments versus chlorophyll-a concentrations within the euphotic zone: a ubiquitous relationship. Limnology and Oceanography, 45(5):1130–1143, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Tall:2006:GDS

[TCC06] Laure Tall, Louise Cloutier, and Antonella Cattaneo. Grazerdiatom size relationships in an epiphytic community. Limnology and Oceanography, 51(2):1211–1216, March 2006. CODEN LI-OCAH. ISSN 0024-3590.

Tsai:2005:SDV

[TCCG05] An-Yi Tsai, Kuo-Ping Chiang, Jeng Chang, and Gwo-Ching Gong. Seasonal diel variations of picoplankton and nanoplankton in a subtropical western Pacific coastal ecosystem. Limnology and Oceanography, 50(4):1221–1231, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Teubner:2003:EPA

[TCD^{+}03] Katrin Teubner, N. D. Crosbie, K. Donabaum, W. Kabas, A. K. T. Kirschner, G. Pfister, M. Salbrechter, and M. T. Dokulil. Enhanced phosphorus accumulation efficiency by the pelagic community at reduced phosphorus supply: a lake experiment from bacteria to metazoan zooplankton. Limnology and Oceanography, 48(3):1141–1149, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Thebault:2009:BMR

[TCL^{+}09] Julien Thébault, Laurent Chauvaud, Stéphane L’Helguen, Jacques Clavier, Aurélie Barats, Séverine Jacquet, Christophe Pécheyran, and David Amouroux. Barium and molybdenum records in bivalve shells: Geochemical proxies for phytoplankton dynamics in coastal environments? Limnology and Oceanography, 54(3):1002–1014, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Tobias:2003:PWD

[TCP^{+}03] Craig R. Tobias, Matthew Cieri, Bruce J. Peterson, Linda A. Deegan, Joseph Vallino, and Jeffrey Hughes. Processing watershed-derived nitrogen in a well-flushed New England estuary. Limnology and Oceanography, 48(5):1766–1778, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Tarling:2003:SMS

[TCR03] Geraint A. Tarling and Janine Cuzin-Roudy. Synchronization in the molting and spawning activity of northern krill (Meganyctiphanes norvegica) and its effect on recruitment. Limnology and Oceanography, 49(4):2020–2033, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Tranvik:2009:LRR

[TDC^{+}09] Lars J. Tranvik, John A. Downing, James B. Cotner, Steven A. Loiselle, Robert G. Striegl, Thomas J. Ballatore, Peter Dillon, Kerri Finlay, Kenneth Fortino, Lesley B. Knoll, Pirkko L. Kortelainen, Tiit Kutser, Soren. Larsen, Isabelle Laurion, Dina M. Leech, S. Leigh McCallister, Diane M. McKnight, John M. Melack, Erin Overholt, Jason A. Porter, Yves Prairie, William H. Renwick, Fabio Roland, Bradford S. Sherman, David W. Schindler, Sebastian Sobek, Alain Tremblay, Michael J. Vanni, Antonie M. Verschoor, Eddie von Wachenfeldt, and Gesa A. Weyhenmeyer. Lakes and reservoirs as regulators of carbon cycling and climate. Limnology and Oceanography, 54(6 (part 2)):2298–2314, November 2009. CODEN LI-OCAH. ISSN 0024-3590.

Thamdrup:2006:AAO

[TDJ^{+}06] Bo Thamdrup, Tage Dalsgaard, Marlene Mark Jensen, Osvaldo Ulloa, Laura Farías, and Ruben Escribano. Anaerobic ammonium oxidation in the oxygen-deficient waters off northern Chile. Limnology and Oceanography, 51(5):2145–2156, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Twiner:2001:TEH

[TDT01] Michael J. Twiner, S. Jeffrey Dixon, and Charles G. Trick. Toxic effects of Heterosigma akashiwo do not appear to be mediated by hydrogen peroxide. Limnology and Oceanography, 46(6):1400– 1405, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Town:2000:DME

[TF00] Raewyn M. Town and Montserrat Filella. Dispelling the myths: Is the existence of L1 and L2 ligands necessary to explain metal ion speciation in natural waters? Limnology and Oceanography, 45(6):1341–1357, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Twining:2004:TTT

[TF04] Benjamin S. Twining and Nicholas S. Fisher. Trophic transfer of trace metals from protozoa to mesozooplankton. Limnology and Oceanography, 49(1):28–39, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Troedsson:2007:MQD

[TFNT07] Christofer Troedsson, Marc E. Frischer, Jens C. Nejstgaard, and Eric M. Thompson. Molecular quantification of differential ingestion and particle trapping rates by the appendicularian Oikopleura dioica as a function of prey size and shape. Limnology and Oceanography, 52(1):416–427, January 2007. CODEN LI-OCAH. ISSN 0024-3590.

Turnewitsch:2003:VPS

[TG03] Robert Turnewitsch and Gerhard Graf. Variability of particulate seawater properties related to bottom mixed layer-associated internal waves in shallow water on a time scale of hours. Limnology and Oceanography, 48(4):1254–1264, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Talmage:2009:EEC

[TG09] Stephanie C. Talmage and Christopher J. Gobler. The effects of elevated carbon dioxide concentrations on the metamorphosis, size, and survival of larval hard clams (Mercenaria mercenaria), bay scallops (Argopecten irradians), and Eastern oysters (Crassostrea virginica). Limnology and Oceanography, 54(6):2072– 2080, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Timmermans:2001:GRL

[TGdB^{+}01] Klaas R. Timmermans, Loes J. A. Gerringa, Hein J. W. de Baar, Bas van der Wagt, Marcel J. W. Veldhuis, Jeroen T. M. de Jong, Peter L. Croot, and Marie Boye. Growth rates of large and small Southern Ocean diatoms in relation to availability o iron in natural seawater. Limnology and Oceanography, 46(2):260– 266, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Twining:2000:EHB

[TGF00] Benjamin S. Twining, John J. Gilbert, and Nicholas S. Fisher. Evidence of homing behavior in the coral reef mysid Mysidium gracile. Limnology and Oceanography, 45(8):1845–1849, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Treignier:2008:ELF

[TGFPT08] C. Treignier, R. Grover, C. Ferrier-Pagés, and I. Tolosa. Effect of light and feeding on the fatty acid and sterol composition of zooxanthellae and host tissue isolated from the scleractinian coral Turbinaria reniformis. Limnology and Oceanography, 53 (6):2702–2710, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Tobias:2001:QGD

[THA01] Craig R. Tobias, Judson W. Harvey, and Iris C. Anderson. Quantifying groundwater discharge through fringing wetlands to estuaries: Seasonal variability, methods comparison, and implications for wetland–estuary exchange. Limnology and Oceanography, 46(3):604–615, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Trainer:2002:BPD

[THH02] Vera L. Trainer, Barbara M. Hickey, and Rita A. Horner. Biological and physical dynamics of domoic acid production off the Washington coast. Limnology and Oceanography, 47(5):1438– 1446, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Thingstad:2000:ETM

[Thi00] T. Frede Thingstad. Elements of a theory for the mechanisms controlling abundance, diversity, and biogeochemical role of lytic bacterial viruses in aquatic systems. Limnology and Oceanography, 45(6):1320–1328, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Trimmer:2009:ERM

[THJ^{+}09] Mark Trimmer, Alan G. Hildrew, Michelle C. Jackson, James L. Pretty, and Jonathan Grey. Evidence for the role of methanederived carbon in a free-flowing, lowland river food web. Limnology and Oceanography, 54(5):1541–1547, September 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Trainer:2009:VPN

[THL^{+}09] Vera L. Trainer, Barbara M. Hickey, Evelyn J. Lessard, William P. Cochlan, Charles G. Trick, Mark L. Wells, Amoreena MacFadyen, and Stephanie K. Moore. Variability of pseudonitzschia and domoic acid in the Juan de Fuca eddy region and its adjacent shelves. Limnology and Oceanography, 54(1):289– 308, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Thornton:2009:SDD

[Tho09] Daniel C. O. Thornton. Spatiotemporal distribution of dissolved acid polysaccharides (dAPS) in a tidal estuary. Limnology and Oceanography, 54(5):1449–1460, September 2009. CODEN LI-OCAH. ISSN 0024-3590.

Tamburri:2000:CRF

[THR00] Mario N. Tamburri, Magdalena N. Halt, and Bruce H. Robison. Chemically regulated feeding by a midwater medusa. Limnology and Oceanography, 45(7):1661–1666, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Turner:2009:SME

[THT^{+}09] R. Eugene Turner, Brian L. Howes, John M. Teal, Charles S. Milan, Erick M. Swenson, and Dale D. Goehringer-Tonerb. Salt marshes and eutrophication: an unsustainable outcome. Limnology and Oceanography, 54(5):1634–1642, September 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Tang:2000:DBT

[THWS00] Degui Tang, Chin-Chang Hung, Kent W. Warnken, and Peter H. Santschi. The distribution of biogenic thiols in surface waters of Galveston Bay. Limnology and Oceanography, 45(6):1289–1297, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2001:CRT

[TIH^{+}01] Gordon T. Taylor, Maria Iabichella, Tung-Yuan Ho, Mary I. Scranton, Robert C. Thunell, Frank Muller-Karger, and Ramon Varela. Chemoautotrophy in the redox transition zone of the Cariaco Basin: a significant midwater source of organic carbon production. Limnology and Oceanography, 46(1):148–163, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2002:SHB

[TIH02] Derek J. Taylor, Christine R. Ishikane, and Robert A. Haney. The systematics of holarctic bosminids and a revision that reconciles molecular and morphological evolution. Limnology and Oceanography, 47(5):1486–1495, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Tang:2001:PGP

[TJV01] K. W. Tang, H. H. Jakobsen, and A. W. Visser. Phaeocystis globosa (Prymnesiophyceae) and the planktonic food web: Feeding, growth, and trophic interactions among grazers. Limnology and Oceanography, 46(8):1860–1870, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Truesdale:2003:RCI

[TKAW03] Victor W. Truesdale, Hilary Kennedy, Susana Agusti, and Tim J. Waite. On the relative constancy of iodate and totaliodine concentrations accompanying phytoplankton blooms initiated in mesocosm experiments in Antarctica. Limnology and Oceanography, 48(4):1569–1574, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Treude:2005:ECA

[TKBJ05] Tina Treude, Martin Krüger, Antje Boetius, and Bo Barker Jørgensen. Environmental control on anaerobic oxidation of methane in the gassy sediments of Eckernförde Bay (German Baltic). Limnology and Oceanography, 50(6):1771–1786, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Torgersen:2002:DVD

[TKK02] Thomas Torgersen, Egil Karlsbakk, and Stein Kaartvedt. Deviating vertical distribution and increased conspicuousness of parasitized Calanus. Limnology and Oceanography, 47(4):1187– 1191, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Toole:2003:PDD

[TKK^{+}03] Dierdre A. Toole, David J. Kieber, Ronald P. Kiene, David A. Siegel, and Norman B. Nelson. Photolysis and the dimethylsulfide (DMS) summer paradox in the Sargasso Sea. Limnology and Oceanography, 48(3):1088–1100, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Teschke:2007:SLR

[TKM07] Mathias Teschke, So Kawaguchi, and Bettina Meyer. Simulated light regimes affect feeding and metabolism of Antarctic krill, Euphausia superba. Limnology and Oceanography, 52(3):1046– 1054, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Turner:2004:SPC

[TLM04] Andrew Turner, Sophie M. Le Roux, and Geoffrey E. Millward. Speciation and partitioning of cadmium and zinc in two contrasting estuaries: The role of hydrophobic organic matter. Limnology and Oceanography, 49(1):11–19, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Trimborn:2007:EVC

[TLrR07] Scarlett Trimborn, Gerald Langer, and Bjö rn Rost. Effect of varying calcium concentrations and light intensities on calcification and photosynthesis in Emiliania huxleyi. Limnology and Oceanography, 52(5):2285–2293, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Tartarotti:2001:LVC

[TLS01] Barbara Tartarotti, Isabelle Laurion, and Ruben Sommaruga. Large variability in the concentration of mycosporine-like amino acids among zooplankton from lakes located across an altitude gradient. Limnology and Oceanography, 46(6):1546–1552, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Teira:2006:DAB

[TLvAH06] Eva Teira, Philippe Lebaron, Hendrik van Aken, and Gerhard J. Herndl. Distribution and activity of bacteria and archaea in the deep water masses of the North Atlantic. Limnology and Oceanography, 51(5):2131–2144, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Tortell:2002:SIC

[TM02] Philippe D. Tortell and François M. M. Morel. Sources of inorganic carbon for phytoplankton in the eastern subtropical and equatorial Pacific Ocean. Limnology and Oceanography, 47(4): 1012–1022, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Turner:2005:HRT

[TM05] Andrew Turner and Edward Mawji. Hydrophobicity and reactivity of trace metals in the low-salinity zone of a turbid estuary. Limnology and Oceanography, 50(3):1011–1019, May 2005. CO-DEN LIOCAH. ISSN 0024-3590.

Tobias:2001:TFH

[TMA^{+}01] Craig R. Tobias, Stephen A. Macko, Iris C. Anderson, Elizabeth A. Canuel, and Judson W. Harvey. Tracking the fate of a high concentration groundwater nitrate plume through a fringing marsh: a combined groundwater tracer and in situ isotope enrichment study. Limnology and Oceanography, 46(8):1977– 1989, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Tyler:2003:BAC

[TMA03] Anna Christina Tyler, Karen J. McGlathery, and Iris C. Anderson. Benthic algae control sediment–water column fluxes of organic and inorganic nitrogen compounds in a temperate lagoon. Limnology and Oceanography, 48(6):2125–2137, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Twining:2007:PRC

[TMB07] Benjamin S. Twining, Steven E. Mylon, and Gaboury Benoit. Potential role of copper availability in nitrous oxide accumulation in a temperate lake. Limnology and Oceanography, 52(4): 1354–1366, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Tortell:2006:ICU

[TMC06] Philippe D. Tortell, Cheryl L. Martin, and Miranda E. Corkum. Inorganic carbon uptake and intracellular assimilation by subarctic Pacific phytoplankton assemblages. Limnology and Oceanography, 51(5):2102–2110, September 2006. CODEN LI-OCAH. ISSN 0024-3590.

Torres:2003:PIM

[TMG03] Raymond Torres, Mwasi J. Mwamba, and Miguel A. Goñi. Properties of intertidal marsh sediment mobilized by rainfall. Limnology and Oceanography, 48(4):1245–1253, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Tremblay:2006:BDE

[TMH^{+}06] Jean-Éric Tremblay, Christine Michel, Keith A. Hobson, Michel Gosselin, and Neil M. Price. Bloom dynamics in early opening waters of the Arctic Ocean. Limnology and Oceanography, 51 (2):900–912, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Tanaka:2007:ICG

[TMK^{+}07] Yasuaki Tanaka, Toshihiro Miyajima, Isao Koike, Takeshi Hayashibara, and Hiroshi Ogawa. Imbalanced coral growth between organic tissue and carbonate skeleton caused by nutrient enrichment. Limnology and Oceanography, 52(3):1139–1146, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2002:RSM

[TMP02] Brad W. Taylor, Angus R. McIntosh, and Barbara L. Peckarsky. Reach-scale manipulations show invertebrate grazers depress algal resources in streams. Limnology and Oceanography, 47(3): 893–899, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Tank:2000:ANC

[TMS^{+}00] Jennifer L. Tank, Judy L. Meyer, Diane M. Sanzone, Patrick J. Mulholland, Jackson R. Webster, Bruce J. Peterson, Wilfred M. Wollheim, and Norman E. Leonard. Analysis of nitrogen cycling in a forest stream during autumn using a ^{15}N-tracer addition. Limnology and Oceanography, 45(5):1013–1029, July 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Tockner:2002:NOM

[TMUW02] Klement Tockner, Florian Malard, Urs Uehlinger, and J. V. Ward. Nutrients and organic matter in a glacial river–floodplain system (Val Roseg, Switzerland). Limnology and Oceanography, 47(1):266–277, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Takai:2002:CSD

[TMYH02] Noriyuki Takai, Yasufumi Mishima, Akemi Yorozu, and Akira Hoshika. Carbon sources for demersal fish in the western Seto Inland Sea, Japan, examined by δ ^{13}C and δ ^{15}N analyses. Limnology and Oceanography, 47(3):730–741, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Trimmer:2009:PNG

[TN09] Mark Trimmer and Joanna Claire Nicholls. Production of nitrogen gas via anammox and denitrification in intact sediment cores along a continental shelf to slope transect in the North Atlantic. Limnology and Oceanography, 54(2):577–589, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Thomas:2001:IPS

[TNM^{+}01] Steven A. Thomas, J. Denis Newbold, Michael T. Monaghan, G. Wayne Minshall, Theodore Georgian, and Colbert E. Cushing. The influence of particle size on seston deposition in streams. Limnology and Oceanography, 46(6):1415–1424, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Tzortziou:2008:TMS

[TNO^{+}08] Maria Tzortziou, Patrick J. Neale, Christopher L. Osburn, J. Patrick Megonigal, Nagamitsu Maie, and Rudolf Jaffé. Tidal marshes as a source of optically and chemically distinctive colored dissolved organic matter in the Chesapeake Bay. Limnology and Oceanography, 53(1):148–159, January 2008. CODEN LI-OCAH. ISSN 0024-3590.

Toolan:2001:CCB

[Too01] Tara Toolan. Coulometric carbon-based respiration rates and estimates of bacterioplankton growth efficiencies in Massachusetts Bay. Limnology and Oceanography, 46(6):1298–1308, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Tortell:2000:EEP

[Tor00] Philippe D. Tortell. Evolutionary and ecological perspectives on carbon acquisition in phytoplankton. Limnology and Oceanography, 45(3):744–750, May 2000. CODEN LIOCAH. ISSN 0024-3590.

Toledo:2003:SSF

[TP03] Gerardo Toledo and Brian Palenik. A Synechococcus serotype is found preferentially in surface marine waters. Limnology and Oceanography, 48(5):1744–1755, September 2003. CODEN LI-OCAH. ISSN 0024-3590.

Teal:2009:SDS

[TPFS09] Lorna R. Teal, Ruth Parker, Gary Fones, and Martin Solana. Simultaneous determination of in situ vertical transitions of color, pore-water metals, and visualization of infaunal activity in marine sediments. Limnology and Oceanography, 54(5):1801–1810, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Tortell:2008:ICU

[TPG^{+}08] Philippe D. Tortell, Chris Payne, Celine Gueguen, Robert F. Strzepek, Philip W. Boyd, and Björn Rost. Inorganic carbon uptake by Southern Ocean phytoplankton. Limnology and Oceanography, 53(4):1266–1278, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Takacs:2001:BDO

[TPM01] Cristina D. Takacs, John C. Priscu, and Diane M. McKnight. Bacterial dissolved organic carbon demand in McMurdo Dry Valley lakes, Antarctica. Limnology and Oceanography, 46(5): 1189–1194, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Teira:2001:DOC

[TPSF01] Eva Teira, María José Pazó, Pablo Serret, and Emilio Fernández. Dissolved organic carbon production by microbial populations in the Atlantic Ocean. Limnology and Oceanography, 46(6):1370– 1377, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Trudeau:2003:EWV

[TR03] Véronique Trudeau and Joseph B. Rasmussen. The effect of water velocity on stable carbon and nitrogen isotope signatures of periphyton. Limnology and Oceanography, 48(6):2194–2199, November 2003. CODEN LIOCAH. ISSN 0024-3590.

Turner:2006:PIW

[TRF^{+}06] R. E. Turner, N. N. Rabalais, B. Fry, N. Atilla, C. S. Milan, J. M. Lee, C. Normandeau, T. A. Oswald, E. M. Swenson, and D. A. Tomasko. Paleo-indicators and water quality change in the Charlotte Harbor Estuary (Florida). Limnology and Oceanography, 51(1 (part 2)):518–533, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Tortell:2000:ICA

[TRM00] Philippe D. Tortell, Greg H. Rau, and François M. M. Morel. Inorganic carbon acquisition in coastal Pacific phytoplankton communities. Limnology and Oceanography, 45(7):1485–1500, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Tanaka:2003:MPA

[TRT03] Tsuneo Tanaka, Fereidoun Rassoulzadegan, and T. Frede Thingstad. Measurements of phosphate affinity constants and phosphorus release rates from the microbial food web in Villefranche Bay, northwestern Mediterranean. Limnology and Oceanography, 48(3):1150–1160, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Tanaka:2004:OUH

[TRT04] Tsuneo Tanaka, Fereidoun Rassoulzadegan, and T. Frede Thingstad. Orthophosphate uptake by heterotrophic bacteria, cyanobacteria, and autotrophic nanoflagellates in Villefranche Bay, northwestern Mediterranean: Vertical, seasonal, and shortterm variations of the competitive relationship for phosphorus. Limnology and Oceanography, 49(4):1063–1072, July 2004. CO-DEN LIOCAH. ISSN 0024-3590.

Turley:2000:DRC

[TS00] C. M. Turley and E. D. Stutt. Depth-related cell-specific bacterial leucine incorporation rates on particles and its biogeochemical significance in the Northwest Mediterranean. Limnology and Oceanography, 45(2):419–425, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Thompson:2001:PBS

[TS01] Michelle L. Thompson and Linda C. Schaffner. Population biology and secondary production of the suspension feeding polychaete Chaetopterus cf. variopedatus: Implications for benthicpelagic coupling in lower Chesapeake Bay. Limnology and Oceanography, 46(8):1899–1907, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Tartarotti:2006:SOC

[TS06] Barbara Tartarotti and Ruben Sommaruga. Seasonal and ontogenetic changes of mycosporine-like amino acids in planktonic organisms from an alpine lake. Limnology and Oceanography, 51 (3):1530–1541, May 2006. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2008:VBC

[TS08] Gordon T. Taylor and Cornelius W. Sullivan. Vitamin B_{12} and cobalt cycling among diatoms and bacteria in Antarctic sea ice microbial communities. Limnology and Oceanography, 53(5): 1862–1877, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Tod:2009:MVM

[TSA09] S. P. Tod and J. M. Schmid-Araya. Meiofauna versus macrofauna: Secondary production of invertebrates in a lowland chalk stream. Limnology and Oceanography, 54(2):450–456, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Tovar-Sanchez:2009:RIR

[TSDK^{+}09] Antonio Tovar-Sanchez, Carlos M. Duarte, Adam B. Kustka, Susana Agustí, Jordi Dachs, David A. Hutchins, Douglas G. Capone, and Sergio A. Sañudo-Wilhelmy. Rejoinder to: Influence of river discharge in the tropical and subtropical North Atlantic Ocean. Limnology and Oceanography, 54(2):648–652, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Timmermans:2001:AEA

[TSG^{+}01] Klaas R. Timmermans, Josje Snoek, Loes J. A. Gerringa, Ingrid Zondervan, and Hein J. W. de Baar. Not all eukaryotic algae can replace zinc with cobalt: Chaetoceros calcitrans (Bacillariophyceae) versus Emiliania huxleyi (Prymnesiophyceae). Limnology and Oceanography, 46(6):699–703, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Tarling:2006:NGR

[TSH^{+}06] Geraint A. Tarling, Rachael S. Shreeve, Andrew G. Hirst, Angus Atkinson, Dave W. Pond, Eugene J. Murphy, and Jon L. Watkins. Natural growth rates in Antarctic krill (Euphausia superba): I. Improving methodology and predicting intermolt period. Limnology and Oceanography, 51(2):959–972, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Tang:2005:RNT

[TSKA05] Degui Tang, Martin M. Shafer, Dawn A. Karner, and David E. Armstrong. Response of nonprotein thiols to copper stress and extracellular release of glutathione in the diatom Thalassiosira weissflogii. Limnology and Oceanography, 50(4):516–525, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Townsend-Small:2005:CCN

[TSMB05] Amy Townsend-Small, Michael E. McClain, and Jay A. Brandes. Contributions of carbon and nitrogen from the Andes Mountains to the Amazon River: Evidence from an elevational gradient of soils, plants, and river material. Limnology and Oceanography, 50(4):672–685, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Turner:2007:HSS

[TSML07] R. Eugene Turner, Erick M. Swenson, Charles S. Milan, and James M. Lee. Hurricane signals in salt marsh sediments: Inorganic sources and soil volume. Limnology and Oceanography, 52 (3):1231–1238, May 2007. CODEN LIOCAH. ISSN 0024-3590.

Tovar-Sanchez:2006:EDD

[TSSWK^{+}06] Antonio Tovar-Sanchez, Sergio A. Sañudo-Wilhelmy, Adam B. Kustka, Susana Agustí, Jordi Dachs, David A. Hutchins, Douglas G. Capone, and Carlos M. Duarte. Effects of dust deposition and river discharges on trace metal composition of Trichodesmium spp. in the tropical and subtropical North Atlantic Ocean. Limnology and Oceanography, 52(2):1755–1761, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Tolli:2005:BCO

[TT05] John D. Tolli and C. D. Taylor. Biological CO oxidation in the Sargasso Sea and in Vineyard Sound, Massachusetts. Limnology and Oceanography, 50(4):1205–1212, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Treignier:2009:CIC

[TTG^{+}09] C. Treignier, I. Tolosa, R. Grover, S. Reynaud, and C. Ferrier-Pagé sa. Carbon isotope composition of fatty acids and sterols in the scleractinian coral Turbinaria reniformis: Effect of light and feeding. Limnology and Oceanography, 54(6):1933–1940, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Thomsen:2004:POC

[TTSC04] Uffe Thomsen, Bo Thamdrup, David A. Stahl, and Donald E. Canfield. Pathways of organic carbon oxidation in a deep lacustrine sediment, Lake Michigan. Limnology and Oceanography, 49 (6):2046–2057, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Tunnicliffe:2000:FSR

[Tun00] Verena Tunnicliffe. A fine-scale record of 130 years of organic carbon deposition in an anoxic fjord, Saanich Inlet, British Columbia. Limnology and Oceanography, 45(6):1380– 1387, September 2000. CODEN LIOCAH. ISSN 0024-3590.

Teira:2006:AUE

[TvAVH06] Eva Teira, Hendrik van Aken, Cornelis Veth, and Gerhard J. Herndl. Archaeal uptake of enantiomeric amino acids in the meso- and bathypelagic waters of the North Atlantic. Limnology and Oceanography, 51(1):60–69, January 2006. CODEN LI-OCAH. ISSN 0024-3590.

Timmermans:2004:GRH

[TvdWdB04] Klaas R. Timmermans, Bas van der Wagt, and Hein J. W. de Baar. Growth rates, half-saturation constants, and silicate, nitrate, and phosphate depletion in relation to iron availability of four large, open-ocean diatoms from the Southern Ocean. Limnology and Oceanography, 49(6):2141–2151, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Takesue:2002:NCD

[TvG02] Renee K. Takesue and Alexander van Geen. Nearshore circulation during upwelling inferred from the distribution of dissolved cadmium off the Oregon coast. Limnology and Oceanography, 47 (1):176–185, January 2002. CODEN LIOCAH. ISSN 0024-3590.

Thunell:2000:OCF

[TVL^{+}00] Robert C. Thunell, Ramon Varela, Martin Llano, James Collister, Frank Muller-Karger, and Richard Bohrer. Organic carbon fluxes, degradation, and accumulation in an anoxic basin: Sediment trap results from the Cariaco Basin. Limnology and Oceanography, 45(2):300–308, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Tian:2001:SDC

[TVSS01] Ru Cheng Tian, Alain F. Vézina, Michel Starr, and François Saucier. Seasonal dynamics of coastal ecosystems and export production at high latitudes: a modeling study. Limnology and Oceanography, 46(8):1845–1859, November 2001. CODEN LI-OCAH. ISSN 0024-3590.

Teasdale:2004:TBC

[TVT04] Michael Teasdale, Kay Vopel, and David Thistle. The timing of benthic copepod emergence. Limnology and Oceanography, 49 (3):884–889, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Tan:2009:RCD

[TW09] Qiao-Guo Tan and Wen-Xiong Wang. The regulation of calcium in Daphnia magna reared in different calcium environments. Limnology and Oceanography, 54(3):746–756, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Trainer:2009:ESM

[TWC^{+}09] Vera L. Trainer, Mark L. Wells, William P. Cochlan, Charles G. Trick, Brian D. Bill, Keri A. Baugh, Benjamin F. Beall, Julian Herndon, and Nina Lundholm. An ecological study of a massive bloom of toxigenic Pseudo-nitzschia cuspidata off the Washington State coast. Limnology and Oceanography, 54(5):1461–1474, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Tuit:2004:DVM

[TWR04] Caroline Tuit, John Waterbury, and Gregory Ravizza. Diel variation of molybdenum and iron in marine diazotrophic cyanobacteria. Limnology and Oceanography, 49(4):978–990, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Tang:2001:OCC

[TWS01] Degui Tang, Kent W. Warnken, and Peter H. Santschi. Organic complexation of copper in surface waters of Galveston Bay. Limnology and Oceanography, 46(2):321–330, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2003:PCC

[TWS03] Gordon T. Taylor, Juliette Way, and Mary I. Scranton. Planktonic carbon cycling and transport in surface waters of the highly urbanized Hudson River estuary. Limnology and Oceanography, 48(5):1779–1795, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Tank:2005:EUR

[TXH05] Suzanne E. Tank, Marguerite A. Xenopoulos, and Len L. Hendzel. Effect of ultraviolet radiation on alkaline phosphatase activity and planktonic phosphorus acquisition in Canadian boreal shield lakes. Limnology and Oceanography, 50(5):1345–1351, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Tanentzap:2008:CLW

[TYK^{+}08] Andrew J. Tanentzap, Norman D. Yan, Bill Keller, Robert Girard, Jocelyne Heneberry, John M. Gunn, David P. Hamilton, and Peter A. Taylor. Cooling lakes while the world warms: Effects of forest regrowth and increased dissolved organic matter on the thermal regime of a temperate, urban lake. Limnology and Oceanography, 53(1):404–410, January 2008. CODEN LI-OCAH. ISSN 0024-3590.

Upstill-Goddard:2001:UPR

[UGSMN01] Robert C. Upstill-Goddard, Jo M. Suijlen, Gillian Malin, and Philip D. Nightingale. The use of photolytic rhodamines WT and sulpho G as conservative tracers of dispersion in surface waters. Limnology and Oceanography, 46(4):927–934, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Urabe:2000:DCP

[UGY^{+}00] Jotaro Urabe, Tek Bahadur Gurung, Takehito Yoshida, Tatsuki Sekino, Masami Nakanishi, Masahiro Maruo, and Eiichiro Nakayama. Diel changes in phagotrophy by Cryptomonas in Lake Biwa. Limnology and Oceanography, 45(7):1558–1563, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Uitz:2008:RPP

[UHB^{+}08] Julia Uitz, Yannick Huot, Flavienne Bruyant, Marcel Babin, and Hervé Claustre. Relating phytoplankton photophysiological properties to community structure on large scales. Limnology and Oceanography, 53(2):614–630, March 2008. CODEN LI-OCAH. ISSN 0024-3590.

Umlauf:2005:IEM

[UL05] Lars Umlauf and Ulrich Lemmin. Interbasin exchange and mixing in the hypolimnion of a large lake: The role of long internal waves. Limnology and Oceanography, 50(5):1601–1611, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Ueyama:2005:WIM

[UM05] Rei Ueyama and Bruce C. Monger. Wind-induced modulation of seasonal phytoplankton blooms in the North Atlantic derived from satellite observations. Limnology and Oceanography, 50(6): 1820–1829, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Unrein:2007:SYR

[UMASG07] Fernando Unrein, Ramon Massana, Laura Alonso-Sáez, and Josep M. Gasol. Significant year-round effect of small mixotrophic flagellates on bacterioplankton in an oligotrophic coastal system. Limnology and Oceanography, 52(1):456–469, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Umezawa:2002:FSM

[UMY^{+}02] Yu Umezawa, Toshihiro Miyajima, Masumi Yamamuro, Hajime Kayanne, and Isao Koike. Fine-scale mapping of land-derived nitrogen in coral reefs by δ ^{15}N in macroalgae. Limnology and Oceanography, 47(5):1405–1416, September 2002. CODEN LI-OCAH. ISSN 0024-3590.

Underwood:2005:PMP

[UPC^{+}05] G. J. C. Underwood, R. G. Perkins, M. C. Consalvey, A. R. M. Hanlon, K. Oxborough, N. R. Baker, and D. M. Paterson. Patterns in microphytobenthic primary productivity: Speciesspecific variation in migratory rhythms and photosynthetic efficiency in mixed-species biofilms. Limnology and Oceanography, 50(3):755–767, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Urbach:2001:UBC

[UVY^{+}01] Ena Urbach, Kevin L. Vergin, Lei Young, Ariel Morse, Gary L. Larson, and Stephen J. Giovannoni. Unusual bacterioplankton community structure in ultra-oligotrophic Crater Lake. Limnology and Oceanography, 46(3):557–572, May 2001. CODEN LIOCAH. ISSN 0024-3590.

Ussher:2007:DRS

[UWA^{+}07] Simon J. Ussher, Paul J. Worsfold, Eric P. Achterberg, Agathe Laës, Stéphane Blain, Patrick Laan, and Hein J. W. de Baar. Distribution and redox speciation of dissolved iron on the European continental margin. Limnology and Oceanography, 52(6): 2530–2539, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Villar-Argaiz:2002:LHB

[VAS02] Manuel Villar-Argaiz and Robert W. Sterner. Life history bottlenecks in Diaptomus clavipes induced by phosphorus-limited algae. Limnology and Oceanography, 47(4):1229–1233, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Vaillancourt:2000:SDM

[VB00] Robert D. Vaillancourt and William M. Balch. Size distribution of marine submicron particles determined by flow field-flow fractionation. Limnology and Oceanography, 45(2):485–492, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Vollmer:2005:DWW

[VBH^{+}05] Martin K. Vollmer, Harvey A. Bootsma, Robert E. Hecky, Graeme Patterson, John D. Halfman, John M. Edmond, David H. Eccles, and Ray F. Weiss. Deep-water warming trend in Lake Malawi, East Africa. Limnology and Oceanography, 50 (4):727–732, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Ventura:2005:ROM

[VC05] Marc Ventura and Jordi Catalan. Reproduction as one of the main causes of temporal variability in the elemental composition of zooplankton. Limnology and Oceanography, 50(6):2043–2056, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Vidal:2005:IWF

[VCCS05] Javier Vidal, Xavier Casamitjana, Jordi Colomer, and Teresa Serra. The internal wave field in Sau reservoir: Observation and modeling of a third vertical mode. Limnology and Oceanography, 50(4):1326–1333, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Vila-Costa:2008:SVD

[VCKS08] Maria Vila-Costa, Ronald P. Kiene, and Rafel Simí. Seasonal variability of the dynamics of dimethylated sulfur compounds in a coastal northwest Mediterranean site. Limnology and Oceanography, 53(1):198–211, January 2008. CODEN LIOCAH. ISSN 0024-3590.

VanMooy:2008:ANL

[VD08] Benjamin A. S. Van Mooy and Allan H. Devol. Assessing nutrient limitation of prochlorococcus in the North Pacific subtropical gyre by using an RNA capture method. Limnology and Oceanography, 53(1):78–88, January 2008. CODEN LIOCAH. ISSN 0024-3590.

vonDassow:2005:BRD

[vDBL05] Peter von Dassow, Rachel N. Bearon, and Michael I. Latz. Bioluminescent response of the dinoflagellate Lingulodinium polyedrum to developing flow: Tuning of sensitivity and the role of desensitization in controlling a defensive behavior of a planktonic cell. Limnology and Oceanography, 50(4):607–619, March 2005. CODEN LIOCAH. ISSN 0024-3590.

vanderHeide:2009:PHS

[vdHPH^{+}09] T. van der Heide, E. T. H. M. Peeters, D. C. R. Hermus, M. M. van Katwijk, J. G. M. Roelofs, and A. J. P. Smolders. Predicting habitat suitability in temperate seagrass ecosystems. Limnology and Oceanography, 54(6):2018–2024, November 2009. CODEN LIOCAH. ISSN 0024-3590.

VanMooy:2004:RBB

[VDK04] Benjamin A. S. Van Mooy, Allan H. Devol, and Richard G. Keil. Relationship between bacterial community structure, light, and carbon cycling in the eastern subarctic North Pacific. Limnology and Oceanography, 49(4):1056–1062, July 2004. CODEN LIOCAH. ISSN 0024-3590.

vandePoll:2006:PME

[vdPAJ^{+}06] Willem H. van de Poll, Anne-Carlijn Alderkamp, Paul J. Janknegt, Jan Roggeveld, and Anita G. J. Buma. Photoacclimation modulates excessive photosynthetically active and ultraviolet radiation effects in a temperate and an Antarctic marine diatom. Limnology and Oceanography, 51(3):1239–1248, May 2006. CODEN LIOCAH. ISSN 0024-3590.

vandePoll:2007:ADI

[vdPVB07] Willem H. van de Poll, Ronald J. W. Visser, and Anita G. J. Buma. Acclimation to a dynamic irradiance regime changes excessive irradiance sensitivity of Emiliania huxleyi and Thalassiosira weissflogii. Limnology and Oceanography, 52(4):1430– 1438, July 2007. CODEN LIOCAH. ISSN 0024-3590.

derStaay:2000:ADP

[vdSvdSG^{+}00] Seung Yeo Moon van der Staay, Georg W. M. van der Staay, Laure Guillou, Daniel Vaulot, Herveé Claustre, and Linda K. Medlin. Abundance and diversity of prymnesiophytes in the picoplankton community from the equatorial Pacific Ocean inferred from ^{18}S rDNA sequences. Limnology and Oceanography, 45(1):98–109, January 2000. CODEN LIOCAH. ISSN 0024-3590.

vonElert:2002:DLP

[vE02] Eric von Elert. Determination of limiting polyunsaturated fatty acids in Daphnia galeata using a new method to enrich food algae with single fatty acids. Limnology and Oceanography, 47 (6):1764–1773, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Veuger:2007:NIR

[VEMM07] Bart Veuger, Bradley D. Eyre, Damien Maher, and Jack J. Middelburg. Nitrogen incorporation and retention by bacteria, algae, and fauna in a subtropical, intertidal sediment: an in situ ^{15}Nlabeling study. Limnology and Oceanography, 52(5):1930–1942, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Vaque:2008:SCP

[VGP^{+}08] Dolors Vaqué, Òscar Guadayol, Francesc Peters, Jordi Felipe, Laia Angel-Ripoll, Ramon Terrado, Connie Lovejoy, and Carlos Pedrós-Alioó. Seasonal changes in planktonic bacterivory rates under the ice-covered coastal Arctic Ocean. Limnology and Oceanography, 53(6):2427–2438, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Vopel:2006:PPB

[VH06] Kay Vopel and Ian Hawes. Photosynthetic performance of benthic microbial mats in Lake Hoare, Antarctica. Limnology and Oceanography, 51(5):1801–1812, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Verburga:2009:PWL

[VH09] Piet Verburg and Robert E. Hecky. The physics of the warming of Lake Tanganyika by climate change. Limnology and Oceanography, 54(6 (part 2)):2418–2430, November 2009. CODEN LI-OCAH. ISSN 0024-3590.

Verleyen:2004:QHS

[VHL^{+}04] Elie Verleyen, Dominic A. Hodgson, Peter R. Leavitt, Koen Sabbe, and Wim Vyverman. Quantifying habitat-specific diatom production: a critical assessment using morphological and biogeochemical markers in Antarctic marine and lake sediments. Limnology and Oceanography, 49(5):1528–1539, September 2004. CODEN LIOCAH. ISSN 0024-3590.

vanHengstum:2009:FTA

[vHRB^{+}09] Peter J. van Hengstum, Eduard G. Reinhardt, Patricia A. Beddows, Henry P. Schwarcz, and Jeremy J. Gabriel. Foraminifera and testate amoebae (thecamoebians) in an anchialine cave: Surface distributions from Aktun Ha (Carwash) cave system, Mexico. Limnology and Oceanography, 54(1):391–396, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Vincent:2000:EDP

[Vin00] Warwick F. Vincent. Ecosystem dynamics in a polar desert: The McMurdo Dry Valleys, Antarctica. Limnology and Oceanography, 45(8):1888, December 2000. CODEN LIOCAH. ISSN 0024-

Vahatalo:2007:PPB

[VJ07] Anssi V. Vähätalo and Marko Järvinen. Photochemically produced bioavailable nitrogen from biologically recalcitrant dissolved organic matter stimulates production of a nitrogenlimited microbial food web in the Baltic Sea. Limnology and Oceanography, 52(1):132–143, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Vadeboncoeur:2003:GGL

[VJV^{+}03] Yvonne Vadeboncoeur, Erik Jeppesen, M. Jake Vander Zanden, Hans-Henrik Schierup, Kirsten Christoffersen, and David M. Lodge. From Greenland to green lakes: Cultural eutrophication and the loss of benthic pathways in lakes. Limnology and Oceanography, 48(4):1408–1418, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Vestheim:2006:PCA

[VK06] Hege Vestheim and Stein Kaartvedt. Plasticity in coloration as an antipredator strategy among zooplankton. Limnology and Oceanography, 53(2):1931–1934, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Viitasalo:2007:BCP

[VKV07] Satu Viitasalo, Tarja Katajisto, and Markku Viitasalo. Bioturbation changes the patterns of benthic emergence in zooplankton. Limnology and Oceanography, 52(6):2325–2339, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Vijverberg:2005:EPF

[VKvD05] Jacobus Vijverberg, Hans Peter Koelewijn, and Wim L. T. van Densen. Effects of predation and food on the population dynamics of the raptorial cladoceran Leptodora kindtii. Limnology and Oceanography, 50(2):455–464, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Voss:2005:BRF

[VM05] Kenneth J. Voss and André Morel. Bidirectional reflectance function for oceanic waters with varying chlorophyll concentrations: Measurements versus predictions. Limnology and Oceanography, 50(4):698–705, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Vreca:2006:CAO

[VM06] Polona Vreca and Gregor Muri. Changes in accumulation of organic matter and stable carbon and nitrogen isotopes in sediments of two Slovenian mountain lakes (Lake Ledvica and Lake Planina), induced by eutrophication changes. Limnology and Oceanography, 51(1 (part 2)):781–790, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Vargas:2007:RIM

[VMC^{+}07] Cristian A. Vargas, Rodrigo A. Martínez, L. Antonio Cuevas, Marcelo A. Pavez, Carolina Cartes, Humberto E. González, Rubén Escribano, and Giovanni Daneri. The relative importance of microbial and classical food webs in a highly productive coastal upwelling area. Limnology and Oceanography, 52(4): 1495–1510, July 2007. CODEN LIOCAH. ISSN 0024-3590.

VanMooy:2001:ETC

[VMH^{+}01] Benjamin Van Mooy, Barbara MacGregor, David Hollander, Ken Nealson, and David Stahl. Evidence for tight coupling between active bacteria and particulate organic carbon during seasonal stratification of Lake Michigan. Limnology and Oceanography, 46(5):1202–1208, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Voparil:2008:EVM

[VMJ08] Ian M. Voparil, Lawrence M. Mayer, and Peter A. Jumars. Emulsions versus micelles in the digestion of lipids by benthic invertebrates. Limnology and Oceanography, 53(1):387–394, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Vesely:2003:ITD

[VMKN03] Josef Veselý, Vladimír Majer, Jiří Kopáček, and Stephen A. Norton. Increasing temperature decreases aluminum concentrations in Central European lakes recovering from acidification. Limnology and Oceanography, 48(6):2346–2354, November 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Voss:2003:SUR

[VMS^{+}03] Kenneth J. Voss, Curtis D. Mobley, Lydia K. Sundman, James E. Ivey, and Charles H. Mazel. The spectral upwelling radiance distribution in optically shallow waters. Limnology and Oceanography, 48(1 (part 2)):364–373, January 2003. CODEN LIOCAH. ISSN 0024-3590.

VandenMeersche:2004:CNC

[VMS^{+}04] Karel Van den Meersche, Jack J. Middelburg, Karline Soetaert, Pieter van Rijswijk, Henricus T. S. Boschker, and Carlo H. R. Heip. Carbon-nitrogen coupling and algal-bacterial interactions during an experimental bloom: Modeling a ^{13}C tracer experiment. Limnology and Oceanography, 49(3):862–878, May 2004. CODEN LIOCAH. ISSN 0024-3590.

Vuorio:2009:NIC

[VMS09] Kristiina Vuorio, Markus Meili, and Jouko Sarvala. Natural isotopic composition of carbon (δ ^{13}C) correlates with colony size in the planktonic cyanobacterium Gloeotrichia echinulata. Limnology and Oceanography, 54(3):925–929, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Vonk:2008:DON

[VMSB08] Jan Arie Vonk, Jack J. Middelburg, Johan Stapel, and Tjeerd J. Bouma. Dissolved organic nitrogen uptake by seagrasses. Limnology and Oceanography, 53(2):542–548, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Verweij:2008:SNC

[VNH^{+}08] Marieke C. Verweij, Ivan Nagelkerken, Ingmar Hans, Susanne M. Ruseler, and Paul R. D. Mason. Seagrass nurseries contribute to coral reef fish populations. Limnology and Oceanography, 53 (5):1540–1547, July 2008. CODEN LIOCAH. ISSN 0024-3590.

vanNugteren:2009:SDD

[vNHM^{+}09] Paul van Nugteren, Peter M. J. Herman, Leon Moodley, Jack J. Middelburg, Matthijs Vos, and Carlo H. R. Heip. Spatial distribution of detrital resources determines the outcome of competition between bacteria and a facultative detritivorous worm. Limnology and Oceanography, 54(5):1413–1419, September 2009. CODEN LIOCAH. ISSN 0024-3590.

vanOevelen:2009:CWC

[vODL^{+}09] Dick van Oevelen, Gerard Duineveld, Marc Lavaleye, Furu Mienis, Karline Soetaert, and Carlo H. R. Heip. The cold-water coral community as hotspot of carbon cycling on continental margins: a food-web analysis from Rockall Bank (northeast Atlantic). Limnology and Oceanography, 54(6):1829–1844, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Vieira:2008:RHE

[VOG^{+}08] Armando Augusto Henriques Vieira, Pedro Ivo Coelho Ortolano, Danilo Giroldo, Maria José Dellamano Oliveira, Thais Beraldo Bittar, Ana Teresa Lombardi, Antonio Luiz Sartori, and Berit Smestad Paulsen. Role of hydrophobic extracellular polysaccharide of Aulacoseira granulata (Bacillariophyceae) on aggregate formation in a turbulent and hypereutrophic reservoir. Limnology and Oceanography, 53(5):1887–1899, September 2008. CODEN LIOCAH. ISSN 0024-3590.

vanOevelen:2006:FBC

[vOMSM06a] Dick van Oevelen, Jack J. Middelburg, Karline Soetaert, and Leon Moodley. The fate of bacterial carbon in an intertidal sediment: Modeling an in situ isotope tracer experiment. Limnology and Oceanography, 51(3):1302–1314, May 2006. CODEN LIOCAH. ISSN 0024-3590.

vanOevelen:2006:TSB

[vOMSM06b] Dick van Oevelen, Leon Moodley, Karline Soetaert, and Jack J. Middelburg. The trophic significance of bacterial carbon in a marine intertidal sediment: Results of an in situ stable isotope labeling study. Limnology and Oceanography, 51(5):2349–2359, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Vrede:2002:IFQ

[VPA02] Tobias Vrede, Jonas Persson, and Grete Aronsen. The influence of food quality (P : C ratio) on RNA : DNA ratio and somatic growth rate of Daphnia. Limnology and Oceanography, 47(2): 487–494, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Volkenborn:2007:BBE

[VPH^{+}07] N. Volkenborn, L. Polerecky, S. I. C. Hedtkamp, J. E. E. van Beusekom, and D. de Beer. Bioturbation and bioirrigation extend the open exchange regions in permeable sediments. Limnology and Oceanography, 52(5):1898–1909, September 2007. CO-DEN LIOCAH. ISSN 0024-3590.

VanderZanden:2001:VTF

[VR01] M. Jake Vander Zanden and Joseph B. Rasmussen. Variation in δ ^{15}N and δ ^{13}C trophic fractionation: Implications for aquatic food web studies. Limnology and Oceanography, 47(4):2061– 2066, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Vidal:2007:SEH

[VRC07] Javier Vidal, Francisco J. Rueda, and Xavier Casamitjana. The seasonal evolution of high vertical-mode internal waves in a deep reservoir. Limnology and Oceanography, 52(6):2656–2667, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Vardaro:2009:CVC

[VRS09] Michael F. Vardaro, Henry A. Ruhl, and Kenneth, Jr. L. Smith. Climate variation, carbon flux, and bioturbation in the abyssal North Pacific. Limnology and Oceanography, 54(6):2081–2088, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Vahatalo:2000:SQY

[VSSTS00] Anssi V. Vähätalo, Mirja Salkinoja-Salonen, Petteri Taalas, and Kalevi Salonen. Spectrum of the quantum yield for photochemical mineralization of dissolved organic carbon in a humic lake. Limnology and Oceanography, 45(3):664–676, May 2000. CO-DEN LIOCAH. ISSN 0024-3590.

Vopel:2005:WIH

[VTO^{+}05] Kay Vopel, David Thistle, Jörg Ott, Monika Bright, and Hans Røy. Wave-induced H_{2}S flux sustains a chemoautotrophic symbiosis. Limnology and Oceanography, 50(1):128–133, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Vopel:2003:EBS

[VTR03] Kay Vopel, David Thistle, and Rutger Rosenberg. Effect of the brittle star Amphiura filiformis (Amphiuridae, Echinodermata) on oxygen flux into the sediment. Limnology and Oceanography, 49(4):2034–2045, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Vanormelingen:2009:LGA

[VVD^{+}09] Pieter Vanormelingen, Wim Vyverman, Dries De Bock, Katleen Van der Gucht, and Luc De Meester. Local genetic adaptation to grazing pressure of the green alga Desmodesmus armatus in a strongly connected pond system. Limnology and Oceanography, 54(2):503–511, March 2009. CODEN LIOCAH. ISSN 0024-3590.

Veuger:2006:FPI

[VvOBM06] Bart Veuger, Dick van Oevelen, Henricus T. S. Boschker, and Jack J. Middelburg. Fate of peptidoglycan in an intertidal sediment: an in situ ^{13}C-labeling study. Limnology and Oceanography, 51(4):1572–1580, July 2006. CODEN LIOCAH. ISSN 0024-3590.

VandenMeersche:2009:AAC

[VVSM09] Karel Van den Meersche, Pieter Van Rijswijk, Karline Soetaert, and Jack J. Middelburg. Autochthonous and allochthonous contributions to mesozooplankton diet in a tidal river and estuary: Integrating carbon isotope and fatty acid constraints. Limnology and Oceanography, 54(1):62–74, January 2009. CODEN LI-OCAH. ISSN 0024-3590.

VonElert:2001:SCF

[VW01] Eric Von Elert and Tania Wolffrom. Supplementation of cyanobacterial food with polyunsaturated fatty acids does not improve growth of Daphnia. Limnology and Oceanography, 46 (6):1552–1558, September 2001. CODEN LIOCAH. ISSN 0024-3590.

Vahatalo:2008:LTP

[VW08] Anssi V. Vähätalo and Robert G. Wetzel. Long-term photochemical and microbial decomposition of wetland-derived dissolved organic matter with alteration of ^{13}C : ^{12}C mass ratio. Limnology and Oceanography, 53(4):1387–1392, July 2008. CO-DEN LIOCAH. ISSN 0024-3590.

vonWachenfeldt:2009:MIF

[vWBT09] Eddie von Wachenfeldt, David Bastviken, and Lars J. Tranvik. Microbially induced flocculation of allochthonous dissolved organic carbon in lakes. Limnology and Oceanography, 54(5):1811– 1818, September 2009. CODEN LIOCAH. ISSN 0024-3590.

Valdes-Weaver:2006:LTT

[VWPP^{+}06] Lexia M. Valdes-Weaver, Michael F. Piehler, James L. Pinckney, Karin E. Howe, Karen Rossignol, and Hans W. Paerl. Longterm temporal and spatial trends in phytoplankton biomass and class-level taxonomic composition in the hydrologically variable Neuse–Pamlico estuarine continuum, North Carolina, U.S.A. Limnology and Oceanography, 51(3):1410–1420, May 2006. CO-DEN LIOCAH. ISSN 0024-3590.

vonWachenfeldt:2008:LAD

[vWSBT08] Eddie von Wachenfeldt, Sebastian Sobek, David Bastviken, and Lars J. Tranvik. Linking allochthonous dissolved organic matter and boreal lake sediment carbon sequestration: The role of light-mediated flocculation. Limnology and Oceanography, 53 (6):2416–2426, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Verschoor:2007:IMT

[VZM07] Antonie M. Verschoor, Yegor S. Zadereev, and Wolf M. Mooij. Infochemical-mediated trophic interactions between the rotifer Brachionus calyciflorus and its food algae. Limnology and Oceanography, 52(5):2109–2119, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Widdicombe:2001:IBP

[WA01] Stephen Widdicombe and Melanie C. Austen. The interaction between physical disturbance and organic enrichment: an important element in structuring benthic communities. Limnology and Oceanography, 46(7):1720–1733, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Wagner:2009:CDQ

[WA09] Carola Wagner and Rita Adrian. Cyanobacteria dominance: Quantifying the effects of climate change. Limnology and Oceanography, 54(6 (part 2)):2460–2468, November 2009. CO-DEN LIOCAH. ISSN 0024-3590.

Weeks:2008:IPC

[WAB^{+}08] S. J. Weeks, K. R. N. Anthony, A. Bakun, G. C. Feldman, and O. Hoegh-Guldberg. Improved predictions of coral bleaching using seasonal baselines and higher spatial resolution. Limnology and Oceanography, 53(4):1369–1375, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Weinbauer:2009:EVP

[WAGH09] Markus G. Weinbauer, Jesus-Maria Arrieta, Christian Griebler, and Gerhard J. Herndl. Enhanced viral production and infection of bacterioplankton during an iron-induced phytoplankton bloom in the Southern Ocean. Limnology and Oceanography, 54 (3):774–784, May 2009. CODEN LIOCAH. ISSN 0024-3590.

Wangersky:2002:BRG

[Wan02] Peter J. Wangersky. Book review: Gili, J. M., J. L. Pretus, and T. T. Packard [Eds.]. 2001. A marine science odyssey into the 21st century. Scientia Marina (suppl. 2) 65: 1–326. Institut de Ciències del Mar, Barcelona. 326 p. EURO90. ISBN ISSN 0214-8358. Limnology and Oceanography, 47(4):1270–1271, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Wieters:2009:BCS

[WBB09] Evie A. Wieters, B. R. Broitman, and G. M. Brancha. Benthic community structure and spatiotemporal thermal regimes in two upwelling ecosystems: Comparisons between South Africa and Chile. Limnology and Oceanography, 54(4):1060–1072, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Werne:2002:MIT

[WBD02] Josef P. Werne, M. Baas, and J. S. Sinninghe Damsté. Molecular isotopic tracing of carbon flow and trophic relationships in a methane-supported benthic microbial community. Limnology and Oceanography, 47(6):1694–1701, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Weinbauer:2003:LVI

[WBH03] Markus G. Weinbauer, Ingrid Brettar, and Manfred G. Höfle. Lysogeny and virus-induced mortality of bacterioplankton in surface, deep, and anoxic marine waters. Limnology and Oceanography, 48(4):1457–1465, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Werner:2006:STP

[WBP^{+}06] Ursula Werner, Markus Billerbeck, Lubos Polerecky, Ulrich Franke, Markus Huettel, Justus E. E. van Beusekom, and Dirk de Beer. Spatial and temporal patterns of mineralization rates and oxygen distribution in a permeable intertidal sand flat (Sylt, Germany). Limnology and Oceanography, 51(6):2549– 2563, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Webb:2009:PRM

[WBP09] Karen E. Webb, David K. A. Barnes, and Sverre Plankea. Pockmarks: Refuges for marine benthic biodiversity. Limnology and Oceanography, 54(5):1776–1788, September 2009. CODEN LI-OCAH. ISSN 0024-3590.

Winter:2003:EDE

[WBR^{+}03a] Thomas C. Winter, Donald C. Buso, Donald O. Rosenberry, Gene E. Likens, Alex, Jr. M. Sturrock, and David P. Mau. Evaporation determined by the energy-budget method for Mirror Lake, New Hampshire. Limnology and Oceanography, 48(3): 995–1009, May 2003. CODEN LIOCAH. ISSN 0024-3590.

Wissel:2003:EWC

[WBR03b] B. Wissel, W. J. Boeing, and C. W. Ramcharan. Effects of water color on predation regimes and zooplankton assemblages in freshwater lakes. Limnology and Oceanography, 49(4):1965– 1976, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Wacker:2002:FQE

[WBvE02] Alexander Wacker, Paul Becher, and Eric von Elert. Food quality effects of unsaturated fatty acids on larvae of the zebra mussel Dreissena polymorpha. Limnology and Oceanography, 47(4): 1242–1248, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Wang:2004:CDD

[WC04] Zhaohui Aleck Wang and Wei-Jun Cai. Carbon dioxide degassing and inorganic carbon export from a marsh-dominated estuary (the Duplin River): a marsh CO_{2} pump. Limnology and Oceanography, 49(2):341–354, March 2004. CODEN LIOCAH. ISSN 0024-3590.

Wells:2006:APR

[WCB^{+}06] Llyd E. Wells, Michael Cordray, Sarah Bowerman, Lisa A. Miller, Warwick F. Vincent, and Jody W. Deming. Archaea in particle-rich waters of the Beaufort Shelf and Franklin Bay, Canadian Arctic: Clues to an allochthonous origin? Limnology and Oceanography, 51(1):47–59, January 2006. CODEN LI-OCAH. ISSN 0024-3590.

Wirum:2007:SOA

[WCGBG07] Ferenc P. Wírum, Juan P. Carricart-Ganivet, Larry Benson, and Duncan Golicher. Simulation and observations of annual density banding in skeletons of Montastraea (Cnidaria: Scleractinia) growing under thermal stress associated with ocean warming. Limnology and Oceanography, 52(5):2317–2323, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Woulds:2007:OCS

[WCL^{+}07] Clare Woulds, Greg L. Cowie, Lisa A. Levin, Johan H. Andersson, Jack J. Middelburg, Sandra Vandewiele, Peter A. Lamont, Kate E. Larkin, Andrew J. Gooday, Stefanie Schumacher, Christine Whitcraft, Rachel M. Jeffreys, and Matthew Schwartz. Oxygen as a control on sea floor biological communities and their roles in sedimentary carbon cycling. Limnology and Oceanography, 52(4):1698–1709, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Wiedenmann:2009:CRA

[WCM09] John Wiedenmann, Katherine A. Cresswell, and Marc Mangel. Connecting recruitment of Antarctic krill and sea ice. Limnology and Oceanography, 54(3):799–811, May 2009. CODEN LI-OCAH. ISSN 0024-3590.

Wieland:2001:FSM

[WdBD^{+}01] Andrea Wieland, Dirk de Beer, Lars R. Damgaard, Michael Kühl, Dagmar van Dusschoten, and Henk Van As. Fine-scale measurement of diffusivity in a microbial mat with nuclear magnetic resonance imaging. Limnology and Oceanography, 46(2): 248–259, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Williamson:2007:DZC

[WDL07] Craig E. Williamson, Hendrika J. De Lange, and Dina M. Leech. Do zooplankton contribute to an ultraviolet clear-water phase in lakes? Limnology and Oceanography, 52(2):662–667, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Whitehead:2000:IUB

[WdMD^{+}00] Robert F. Whitehead, Stephen de Mora, Serge Demers, Michel Gosselin, Patrick Monfort, and Behzad Mostajir. Interactions of ultraviolet-B radiation, mixing, and biological activity on photobleaching of natural chromophoric dissolved organic matter: a mesocosm study. Limnology and Oceanography, 45(2):278–291, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Weitere:2008:DGM

[WDVA08] Markus Weitere, Johanna Dahlmann, Carsten Viergutz, and Hartmut Arndt. Differential grazer-mediated effects of high summer temperatures on pico- and nanoplankton communities. Limnology and Oceanography, 53(2):477–486, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Wang:2001:RZA

[WDX01] Wen-Xiong Wang, Robert C. H. Dei, and Yan Xu. Responses of Zn assimilation by coastal plankton to macronutrients. Limnology and Oceanography, 46(6):1524–1534, September 2001. CO-DEN LIOCAH. ISSN 0024-3590.

Weyhenmeyer:2004:SRB

[Wey04] Gesa A. Weyhenmeyer. Synchrony in relationships between the North Atlantic Oscillation and water chemistry among Sweden’s largest lakes. Limnology and Oceanography, 49(4):1191–1201, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Wang:2001:PCO

[WG01] Wen-Xiong Wang and Laodong Guo. Production of colloidal organic carbon and trace metals by phytoplankton decomposition. Limnology and Oceanography, 46(2):278–286, March 2001. CODEN LIOCAH. ISSN 0024-3590.

Wenzhofer:2004:SSS

[WG04] Frank Wenzhöfer and Ronnie N. Glud. Small-scale spatial and temporal variability in coastal benthic O_{2} dynamics: Effects of fauna activity. Limnology and Oceanography, 49(5):1471–1481, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Wishner:2000:LSE

[WGC00] Karen F. Wishner, Marcia M. Gowing, and Gelfman Celia. Living in suboxia: Ecology of an Arabian Sea oxygen minimum zone copepod. Limnology and Oceanography, 45(7):1576–1593, November 2000. CODEN LIOCAH. ISSN 0024-3590.

Williamson:2002:TDU

[WGD^{+}02] Craig E. Williamson, Gabriella Grad, Hendrika J. De Lange, Shawna Gilroy, and Danielle M. Karapelou. Temperaturedependent ultraviolet responses in zooplankton: Implications of climate change. Limnology and Oceanography, 47(6):1844–1848, November 2002. CODEN LIOCAH. ISSN 0024-3590.

Weider:2004:AAR

[WGKE04] Lawrence J. Weider, Karen L. Glenn, Marcia Kyle, and James J. Elser. Associations among ribosomal (r)DNA intergenic spacer length, growth rate, and C : N : P stoichiometry in the genus Daphnia. Limnology and Oceanography, 49(4 (part 2)):1417– 1423, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Waite:2005:LED

[WGLT05] Anya M. Waite, Örjan Gustafsson, Odd Lindahl, and Peter Tiselius. Linking ecosystem dynamics and biogeochemistry:

Sinking fractionation of organic carbon in a Swedish fjord. Limnology and Oceanography, 50(4):658–671, March 2005. CODEN LIOCAH. ISSN 0024-3590.

Wahl:2008:ERZ

[WGN^{+}08] David H. Wahl, Jodi Goodrich, Michael A. Nannini, John M. Dettmers, and Daniel A. Soluk. Exploring riverine zooplankton in three habitats of the Illinois River ecosystem: Where do they come from? Limnology and Oceanography, 53(6):2583–2593, November 2008. CODEN LIOCAH. ISSN 0024-3590.

Wilson:2007:DTC

[WH07] Alan E. Wilson and Mark E. Hay. A direct test of cyanobacterial chemical defense: Variable effects of microcystin-treated food on two Daphnia pulicaria clones. Limnology and Oceanography, 52 (4):1467–1479, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Wetz:2006:RIM

[WHC^{+}06] Michael S. Wetz, Burke Hales, Zanna Chase, Patricia A. Wheeler, and Michael M. Whitney. Riverine input of macronutrients, iron, and organic matter to the coastal ocean off Oregon, U.S.A., during the winter. Limnology and Oceanography, 51(5): 2221–2231, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Walseng:2006:MCL

[WHHS06] Bjørn Walseng, Dag O. Hessen, Gunnar Halvorsen, and Ann Kristin Schartau. Major contribution from littoral crustaceans to zooplankton species richness in lakes. Limnology and Oceanography, 51(6):2600–2606, November 2006. CODEN LI-OCAH. ISSN 0024-3590.

Walsh:2008:STT

[WIK08] Elizabeth M. Walsh, Anitra E. Ingalls, and Richard G. Keil. Sources and transport of terrestrial organic matter in Vancouver Island fjords and the Vancouver–Washington Margin: a multiproxy approach using δ ^{13}C_{org}, lignin phenols, and the ether lipid BIT index. Limnology and Oceanography, 54(3):1054–1063, May 2008. CODEN LIOCAH. ISSN 0024-3590.

Weyhenmeyer:2007:NDC

[WJA^{+}07] Gesa A. Weyhenmeyer, Erik Jeppesen, Rita Adrian, Lauri Arvola, Thorsten Blenckner, Thomas Jankowski, Eleanor Jennings, Peeter Nõges, Tiina Nõges, and Dietmar Straile. Nitratedepleted conditions on the increase in shallow northern European lakes. Limnology and Oceanography, 52(4):1346–1353, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Webb:2007:MAP

[WJMD07] Eric A. Webb, Rachel Wisniewski Jakuba, James W. Moffett, and Sonya T. Dyhrman. Molecular assessment of phosphorus and iron physiology in Trichodesmium populations from the western central and western South Atlantic. Limnology and Oceanography, 52(5):2221–2232, September 2007. CODEN LI-OCAH. ISSN 0024-3590.

Weyhenmeyer:2009:NRD

[WK09] Gesa A. Weyhenmeyer and Jan Karlsson. Nonlinear response of dissolved organic carbon concentrations in boreal lakes to increasing temperatures. Limnology and Oceanography, 54(6 (part 2)):2513–2519, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Wolinska:2004:DPL

[WKB^{+}04] Justyna Wolinska, Barbara Keller, Kerstin Bittner, Sandra Lass, and Piet Spaak. Do parasites lower Daphnia hybrid fitness? Limnology and Oceanography, 49(4 (part 2)):1401–1407, July 2004. CODEN LIOCAH. ISSN 0024-3590.

Wankel:2006:NSC

[WKFP06] Scott D. Wankel, Carol Kendall, Chris A. Francis, and Adina Paytan. Nitrogen sources and cycling in the San Francisco Bay Estuary: a nitrate dual isotopic composition approach. Limnology and Oceanography, 52(2):1654–1664, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Watanabe:2006:ASC

[WKH^{+}06] Atsushi Watanabe, Hajime Kayanne, Hiroshi Hata, Setsuko Kudo, Ken Nozaki, Ken Kato, Akira Negishi, Yutaka Ikeda, and Hiroya Yamano. Analysis of the seawater CO_{2} system in the barrier reef-lagoon system of Palau using total alkalinity-dissolved inorganic carbon diagrams. Limnology and Oceanography, 51(4): 1614–1628, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Willey:2008:RSF

[WKSM08] Joan D. Willey, Robert J. Kieber, Pamela J. Seaton, and Carrie Miller. Rainwater as a source of Fe(II)-stabilizing ligands to seawater. Limnology and Oceanography, 53(5):1678–1684, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Winter:2009:SCB

[WKW09] Christian Winter, Marie-Emmanuelle Kerros, and Markus G. Weinbauer. Seasonal changes of bacterial and archaeal communities in the dark ocean: Evidence from the Mediterranean Sea. Limnology and Oceanography, 54(1):160–170, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Worm:2006:EEG

[WL06] Boris Worm and Heike K. Lotze. Effects of eutrophication, grazing, and algal blooms on rocky shores. Limnology and Oceanography, 51(1 (part 2)):569–579, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Wing:2007:TSW

[WLP^{+}07] S. R. Wing, J. J. Leichter, C. Perrin, S. M. Rutger, M. H. Bowman, and C. D. Cornelisen. Topographic shading and wave exposure influence morphology and ecophysiology of Ecklonia radiata (C. Agardh 1817) in Fiordland, New Zealand. Limnology and Oceanography, 52(5):1853–1864, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Winters:2003:PSW

[WLRB03] Gidon Winters, Yossi Loya, Rüdiger Röttgers, and Sven Beer. Photoinhibition in shallow-water colonies of the coral Stylophora pistillata as measured in situ. Limnology and Oceanography, 48 (4):1388–1393, July 2003. CODEN LIOCAH. ISSN 0024-3590.

Warner:2006:EDC

[WLRT06] Mark E. Warner, Todd C. LaJeunesse, Jennifer D. Robison, and Rebecca M. Thur. The ecological distribution and comparative photobiology of symbiotic dinoflagellates from reef corals in Belize: Potential implications for coral bleaching. Limnology and Oceanography, 53(2):1887–1897, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Worm:2000:CFW

[WLS00] Boris Worm, Heike K. Lotze, and Ulrich Sommer. Coastal food web structure, carbon storage, and nitrogen retention regulated by consumer pressure and nutrient loading. Limnology and Oceanography, 45(2):339–349, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Wong:2003:AMM

[WLTF03] Wai Hing Wong, Jeffrey S. Levinton, Benjamin S. Twining, and Nicholas Fisher. Assimilation of micro- and mesozooplankton by zebra mussels: a demonstration of the food web link between zooplankton and benthic suspension feeders. Limnology and Oceanography, 48(1):308–312, January 2003. CODEN LI-OCAH. ISSN 0024-3590.

Willis:2006:RFC

[WM06] Theodore V. Willis and John J. Magnuson. Responses of fish communities in five north temperate lakes to exotic species and climate. Limnology and Oceanography, 52(6):2808–2820, November 2006. CODEN LIOCAH. ISSN 0024-3590.

Woodson:2007:FBC

[WM07] C. B. Woodson and M. A. McManus. Foraging behavior can influence dispersal of marine organisms. Limnology and Oceanography, 52(6):2701–2709, November 2007. CODEN LIOCAH. ISSN 0024-3590.

Watras:2006:MCL

[WMRK06] C. J. Watras, K. A. Morrison, O. Regnell, and T. K. Kratz. The methylmercury cycle in Little Rock Lake during experimental acidification and recovery. Limnology and Oceanography, 51(1): 257–270, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Wiltshire:2008:RNS

[WMW^{+}08] Karen Helen Wiltshire, Arne Michael Malzahn, Kai Wirtz, Wulf Greve, Silvia Janisch, Peter Mangelsdorf, Bryan F. J. Manly, and Maarten Boersma. Resilience of North Sea phytoplankton spring bloom dynamics: an analysis of long-term data at Helgoland Roads. Limnology and Oceanography, 53(4):1294–1302, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Waldbusser:2004:EIB

[WMWV04] George G. Waldbusser, Roberta L. Marinelli, Robert B. Whitlatch, and Pieter T. Visscher. The effects of infaunal biodiversity on biogeochemistry of coastal marine sediments. Limnology and Oceanography, 49(5):1482–1492, September 2004. CODEN LI-OCAH. ISSN 0024-3590.

Wey:2009:RDS

[WNAW09] Jennifer K. Wey, Helge Norf, Hartmut Arndt, and Markus Weitere. Role of dispersal in shaping communities of ciliates and heterotrophic flagellates within riverine biofilms. Limnology and Oceanography, 54(5):1615–1626, September 2009. CODEN LI-OCAH. ISSN 0024-3590.

Wai:2008:SFO

[WNL^{+}08] Tak-Cheung Wai, Jasmine S. S. Ng, Kenneth M. Y. Leung, David Dudgeon, and Gray A. Williams. The source and fate of organic matter and the significance of detrital pathways in a tropical coastal ecosystem. Limnology and Oceanography, 53(4): 1479–1492, July 2008. CODEN LIOCAH. ISSN 0024-3590.

Worden:2004:ADE

[WNP04] Alexandra Z. Worden, Jessica K. Nolan, and B. Palenik. Assessing the dynamics and ecology of marine picophytoplankton: The importance of the eukaryotic component. Limnology and Oceanography, 49(1):168–179, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Wollheim:2001:ISS

[WPD^{+}01] Wilfred M. Wollheim, Bruce J. Peterson, Linda A. Deegan, John E. Hobbie, Beth Hooker, William B. Bowden, Kenneth J. Edwardson, David B. Arscott, Anne E. Hershey, and Jacques Finlay. Influence of stream size on ammonium and suspended particulate nitrogen processing. Limnology and Oceanography, 46(1):1–13, January 2001. CODEN LIOCAH. ISSN 0024-3590.

Wuest:2000:TKE

[WPV00] Alfred Wüest, Gabriel Piepke, and David C. Van Senden. Turbulent kinetic energy balance as a tool for estimating vertical diffusivity in wind-forced stratified waters. Limnology and Oceanography, 45(6):1388–1400, September 2000. CODEN LIOCAH. ISSN 0024-3590.

West:2009:TBI

[WPW^{+}09] Elizabeth J. West, Kylie A. Pitt, David T. Welsh, Klaus Koop, and David Rissik. Top-down and bottom-up influences of jellyfish on primary productivity and planktonic assemblages. Limnology and Oceanography, 54(6):2058–2071, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Waiser:2000:CCR

[WR00] Marley J. Waiser and Richard D. Robarts. Changes in composition and reactivity of allochthonous DOM in a prairie saline lake. Limnology and Oceanography, 45(4):763–774, June 2000. CODEN LIOCAH. ISSN 0024-3590.

Werdell:2003:RAB

[WR03] P. Jeremy Werdell and Collin S. Roesler. Remote assessment of benthic substrate composition in shallow waters using multispectral reflectance. Limnology and Oceanography, 48(1 (part 2)):557–567, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Webster:2001:UBC

[WRD01] D. R. Webster, S. Rahman, and L. P. Dasi. On the usefulness of bilateral comparison to tracking turbulent chemical odor plumes. Limnology and Oceanography, 46(5):1048–1053, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Wuest:2005:CIL

[WRG^{+}05] Alfred Wüest, Thomas M. Ravens, Nikolai G. Granin, Otti Kocsis, Michael Schurter, and Michael Sturm. Cold intrusions in Lake Baikal: Direct observational evidence for deep-water renewal. Limnology and Oceanography, 55(3):184–196, January 2005. CODEN LIOCAH. ISSN 0024-3590.

Wroblewski:2001:BRH

[Wro01] Joe Wroblewski. Book review: Harrison, Paul J., and Timothy R. Parsons [Eds.]. 2000. Fisheries oceanography: an integrative approach to fisheries ecology and management. Blackwell Science. xii + 347 p. £30. ISBN 0-632-05566-9. Limnology and Oceanography, 46(7):1844, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Wells:2001:SPS

[WS01] Mathew Graeme Wells and Bradford Sherman. Stratification produced by surface cooling in lakes with significant shallow regions. Limnology and Oceanography, 46(7):1747–1759, November 2001. CODEN LIOCAH. ISSN 0024-3590.

Wiegner:2004:SBD

[WS04] Tracy N. Wiegner and Sybil P. Seitzinger. Seasonal bioavailability of dissolved organic carbon and nitrogen from pristine and polluted freshwater wetlands. Limnology and Oceanography, 49 (5):1703–1712, September 2004. CODEN LIOCAH. ISSN 0024-

Weisse:2006:EPG

[WS06] Thomas Weisse and Peter Stadler. Effect of pH on growth, cell volume, and production of freshwater ciliates, and implications for their distribution. Limnology and Oceanography, 52(2):1708– 1715, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Wolowicz:2006:EED

[WSBL06] Maciej Wołowicz, Adam Sokołowski, Abdullah Salem Bawazir, and Rafał Lasota. Effect of eutrophication on the distribution and ecophysiology of the mussel Mytilus trossulus (Bivalvia) in southern Baltic Sea (the Gulf of Gdańsk). Limnology and Oceanography, 51(1 (part 2)):580–590, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Warner:2005:NTE

[WSC^{+}05] Robert R. Warner, Stephen E. Swearer, Jennifer E. Caselle, Michael Sheehy, and Georges Paradis. Natal trace-elemental signatures in the otoliths of an open-coast fish. Limnology and Oceanography, 50(5):1529–1542, September 2005. CODEN LI-OCAH. ISSN 0024-3590.

Warner:2006:NTS

[WSC^{+}06] Robert R. Warner, Stephen E. Swearer, Jennifer E. Caselle, Michael Sheehy, and Georges Paradis. Natal trace=elemental signatures in the otoliths of an open-coast fish. Limnology and Oceanography, 51(1):349, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Webster:2008:EEN

[WSC^{+}08] Katherine E. Webster, Patricia A. Soranno, Kendra Spence Cheruvelil, Mary T. Bremigan, John A. Downing, Peter D. Vaux, Timothy R. Asplund, Linda C. Bacon, and Jody Connor. An empirical evaluation of the nutrient-color paradigm for lakes. Limnology and Oceanography, 54(3):1137–1148, May 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Winder:2009:DSC

[WSEL09] Monika Winder, Daniel E. Schindler, Timothy E. Essington, and Arni H. Litt. Disrupted seasonal clockwork in the population dynamics of a freshwater copepod by climate warming. Limnology and Oceanography, 54(6 (part 2)):2493–2505, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Waite:2000:MSP

[WSHN00] Anya M. Waite, Karl A. Safi, Julie A. Hall, and Scott D. Nodder. Mass sedimentation of picoplankton embedded in organic aggregates. Limnology and Oceanography, 45(1):87–97, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Winter:2005:LBR

[WSHW05] Christian Winter, Arjan Smit, Gerhard J. Herndl, and Markus G. Weinbauer. Linking bacterial richness with viral abundance and prokaryotic activity. Limnology and Oceanography, 50(3):968–977, May 2005. CODEN LIOCAH. ISSN 0024-3590.

White:2006:FES

[WSKL06] Angelicque E. White, Yvette H. Spitz, David M. Karl, and Ricardo M. Letelier. Flexible elemental stoichiometry in Trichodesmium spp. and its ecological implications. Limnology and Oceanography, 51(5):1777–1790, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Weisse:2002:IET

[WSL^{+}02] Thomas Weisse, Peter Stadler, Eva S. Lindström, Susan A. Kimmance, and David J. S. Montagnes. Interactive effect of temperature and food concentration on growth rate: a test case using the small freshwater ciliate Urotricha farcta. Limnology and Oceanography, 47(5):1447–1455, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Welker:2007:SSC

[WŠN^{+}07] Martin Welker, Lenka Šejnohová, Danka Némethová, Hanšvon Döhren, Jiří Jarkovský, and Blahoslav Maršálek. Seasonal shifts in chemotype composition of microcystis sp. communities in the pelagial and the sediment of a shallow reservoir. Limnology and Oceanography, 52(2):609–619, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Wand:2006:BMP

[WSNH06] Ulrich Wand, Vladimir A. Samarkin, Horst-Michael Nitzsche, and Hans-Wolfgang Hubberten. Biogeochemistry of methane in the permanently ice-covered Lake Untersee, central Dronning Maud Land, East Antarctica. Limnology and Oceanography, 51 (2):1180–1194, March 2006. CODEN LIOCAH. ISSN 0024-3590.

Ward:2000:UVO

[WSNM00] J. E. Ward, L. P. Sanford, R. I. E. Newell, and B. A. MacDonald. The utility of in vivo observations for describing particle capture processes in suspension-feeding bivalve molluscs. Limnology and Oceanography, 45(5):1203–1210, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Wilson:2006:ECT

[WST06] Alan E. Wilson, Orlando Sarnelle, and Angeline R. Tillmanns. Effects of cyanobacterial toxicity and morphology on the population growth of freshwater zooplankton: Meta-analyses of laboratory experiments. Limnology and Oceanography, 53(2):1915– 1924, July 2006. CODEN LIOCAH. ISSN 0024-3590.

Williamson:2009:LRS

[WSVS09] Craig E. Williamson, Jasmine E. Saros, Warwick F. Vincent, and John P. Smol. Lakes and reservoirs as sentinels, integrators, and regulators of climate change. Limnology and Oceanography, 54(6 (part 2)):2273–2282, November 2009. CODEN LIOCAH. ISSN 0024-3590.

Wells:2005:DAS

[WTC^{+}05] Mark L. Wells, Charles G. Trick, William P. Cochlan, Margaret P. Hughes, and Vera L. Trainer. Domoic acid: The synergy of iron, copper, and the toxicity of diatoms. Limnology and Oceanography, 50(6):1908–1917, November 2005. CODEN LIOCAH. ISSN 0024-3590.

Wiegner:2009:BED

[WTM09] Tracy N. Wiegner, Randee L. Tubal, and Richard A. MacKenzie. Bioavailability and export of dissolved organic matter from a tropical river during base- and stormflow conditions. Limnology and Oceanography, 54(4):1233–1242, July 2009. CODEN LIOCAH. ISSN 0024-3590.

Wurtsbaugh:2007:NCT

[Wur07] Wayne A. Wurtsbaugh. Nutrient cycling and transport by fish and terrestrial insect nutrient subsidies to lakes. Limnology and Oceanography, 52(6):2715–2718, November 2007. CODEN LI-OCAH. ISSN 0024-3590.

Whitehead:2000:IML

[WV00] Kenia Whitehead and Maria Vernet. Influence of mycosporinelike amino acids (MAAs) on UV absorption by particulate and dissolved organic matter in La Jolla Bay. Limnology and Oceanography, 45(8):1788–1796, December 2000. CODEN LI-OCAH. ISSN 0024-3590.

Wacker:2004:FQC

[WvE04] Alexander Wacker and Eric von Elert. Food quality controls egg quality of the zebra mussel Dreissena polymorpha: The role of fatty acids. Limnology and Oceanography, 49(5):1794–1801, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Wellnitz:2000:HIS

[WW00] Todd A. Wellnitz and J. V. Ward. Herbivory and irradiance shape periphytic architecture in a Swiss alpine stream. Limnology and Oceanography, 45(1):64–75, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Webster:2001:CGC

[WW01] D. R. Webster and M. J. Weissburg. Chemosensory guidance cues in a turbulent chemical odor plume. Limnology and Oceanography, 46(5):1034–1047, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Wood:2002:MST

[WW02] Rose Wood and John Widdows. A model of sediment transport over an intertidal transect, comparing the influences of biological and physical factors. Limnology and Oceanography, 47(3):848– 855, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Wetz:2003:PPO

[WW03] Michael S. Wetz and Patricia A. Wheeler. Production and partitioning of organic matter during simulated phytoplankton blooms. Limnology and Oceanography, 48(5):1808–1817, September 2003. CODEN LIOCAH. ISSN 0024-3590.

Wetz:2007:RDO

[WW07] Michael S. Wetz and Patricia A. Wheeler. Release of dissolved organic matter by coastal diatoms. Limnology and Oceanography, 52(2):798–807, March 2007. CODEN LIOCAH. ISSN 0024-3590.

Wilcox:2005:ELC

[WWMB05] Heidi S. Wilcox, J. Bruce Wallace, Judy L. Meyer, and Jonathan P. Benstead. Effects of labile carbon addition on a headwater stream food web. Limnology and Oceanography, 50 (4):1300–1312, July 2005. CODEN LIOCAH. ISSN 0024-3590.

Wulff:2000:EUR

[WWS^{+}00] Angela Wulff, Sten-Åke Wängberg, Kristina Sundbäck, Claes Nilsson, and Graham J. C. Underwood. Effects of UVB radiation on a marine microphytobenthic community growing on a sandsubstratum under different nutrient conditions. Limnology and Oceanography, 45(5):1144–1152, July 2000. CODEN LIOCAH. ISSN 0024-3590.

Woodson:2005:RCP

[WWWY05] C. B. Woodson, D. R. Webster, M. J. Weissburg, and J. Yen. Response of copepods to physical gradients associated with structure in the ocean. Limnology and Oceanography, 50(5):1552– 1564, September 2005. CODEN LIOCAH. ISSN 0024-3590.

Xie:2009:PCC

[XBD^{+}09] Huixiang Xie, Simon Bélanger, Serge Demers, Warwick F. Vincent, and Tim N. Papakyriakou. Photobiogeochemical cycling of carbon monoxide in the southeastern Beaufort Sea in spring and autumn. Limnology and Oceanography, 54(1):234–249, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Xenopoulos:2003:RCW

[XLF^{+}03] Marguerite A. Xenopoulos, David M. Lodge, Jason Frentress, Timothy A. Kreps, Scott D. Bridgham, Elizabeth Grossman, and Caryn J. Jackson. Regional comparisons of watershed determinants of dissolved organic carbon in temperate lakes from the Upper Great Lakes region and selected regions globally. Limnology and Oceanography, 48(6):2321–2334, November 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Xu:2007:ZCC

[XTSM07] Yan Xu, Degui Tang, Yeala Shaked, and François M. M. Morel. Zinc, cadmium, and cobalt interreplacement and relative use efficiencies in the coccolithophore Emiliania huxleyi. Limnology and Oceanography, 52(5):2294–2305, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Xu:2002:ADB

[XW02] Yan Xu and Wen-Xiong Wang. The assimilation of detritusbound metals by the marine copepod Acartia spinicauda. Limnology and Oceanography, 47(2):604–610, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Yager:2001:DBV

[YCM^{+}01] Patricia L. Yager, Tara L. Connelly, Behzad Mortazavi, K. Eric Wommack, Nasreen Bano, James E. Bauer, Stephen Opsahl, and James T. Hollibaugh. Dynamic bacterial and viral response to an algal bloom at subzero temperatures. Limnology and Oceanography, 46(4):790–801, June 2001. CODEN LIOCAH. ISSN 0024-3590.

Yamashita:2008:ADD

[YJMT08] Youhei Yamashita, Rudolf Jaffé, Nagamitsu Maie, and Eiichiro Tanoue. Assessing the dynamics of dissolved organic matter (DOM) in coastal environments by excitation emission matrix fluorescence and parallel factor analysis (EEM-PARAFAC). Limnology and Oceanography, 53(5):1900–1908, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Yund:2003:SUS

[YM03] Philip O. Yund and Susanne K. Meidel. Sea urchin spawning in benthic boundary layers: Are eggs fertilized before advecting away from females? Limnology and Oceanography, 48(2):795– 801, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Yokokawa:2004:GRM

[YNCK04] Taichi Yokokawa, Toshi Nagata, Matthew T. Cottrell, and David L. Kirchman. Growth rate of the major phylogenetic bacterial groups in the Delaware estuary. Limnology and Oceanography, 49(5):1620–1629, September 2004. CODEN LIOCAH. ISSN 0024-3590.

Yacobi:2007:UUP

[YPBL07] Yosef Z. Yacobi, Naomi Perel, Eugeni Barkan, and Boaz Luz. Unexpected underestimation of primary productivity by ^{18}O and ^{14}C methods in a lake: Implications for slow diffusion of isotope tracers in and out of cells. Limnology and Oceanography, 52(1):329–337, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Yoshiyama:2006:PRN

[YS06] Kohei Yoshiyama and Jonathan H. Sharp. Phytoplankton response to nutrient enrichment in an urbanized estuary: Apparent inhibition of primary production by overeutrophication.

Limnology and Oceanography, 51(1 (part 2)):424–434, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Yahel:2003:SFE

[YSM^{+}03] Gitai Yahel, Jonathan H. Sharp, Dominique Marie, Clivia Häse, and Amatzia Genin. In situ feeding and element removal in the symbiont-bearing sponge Theonella swinhoei: Bulk DOC is the major source for carbon. Limnology and Oceanography, 48(1): 141–149, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Yoder:2002:CPC

[YSO02] James A. Yoder, Stephanie E. Schollaert, and John E. O’Reilly. Climatological phytoplankton chlorophyll and sea surface temperature patterns in continental shelf and slope waters off the northeast U.S. coast. Limnology and Oceanography, 47(3):672– 682, May 2002. CODEN LIOCAH. ISSN 0024-3590.

Yamada:2003:DPC

[YT03] Namiha Yamada and Eiichiro Tanoue. Detection and partial characterization of dissolved glycoproteins in oceanic waters. Limnology and Oceanography, 48(3):1037–1048, May 2003. CO-DEN LIOCAH. ISSN 0024-3590.

Yamashita:2009:BSD

[YT09] Youhei Yamashita and Eiichiro Tanoue. Basin scale distribution of chromophoric dissolved organic matter in the Pacific Ocean. Limnology and Oceanography, 54(2):598–609, March 2009. CO-DEN LIOCAH. ISSN 0024-3590.

York:2007:SID

[YTVR07] Joanna K. York, Gabrielle Tomasky, Ivan Valiela, and Daniel J. Repeta. Stable isotopic detection of ammonium and nitrate assimilation by phytoplankton in the Waquoit Bay estuarine system. Limnology and Oceanography, 52(1):144–155, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Yu:2002:TMA

[YW02] Ri-Qing Yu and Wen-Xiong Wang. Trace metal assimilation and release budget in Daphnia magna. Limnology and Oceanography, 47(2):495–504, March 2002. CODEN LIOCAH. ISSN 0024-3590.

Yahel:2007:SFM

[YWR^{+}07] Gitai Yahel, Frank Whitney, Henry M. Reiswig, Dafne I. Eerkes-Medrano, and Sally P. Leys. In situ feeding and metabolism of glass sponges (Hexactinellida, Porifera) studied in a deep temperate fjord with a remotely operated submersible. Limnology and Oceanography, 52(1):428–440, January 2007. CODEN LI-OCAH. ISSN 0024-3590.

Yahel:2005:DPA

[YYB^{+}05] Ruthy Yahel, Gitai Yahel, Tal Berman, Jules S. Jaffe, and Amatzia Genin. Diel pattern with abrupt crepuscular changes of zooplankton over a coral reef. Limnology and Oceanography, 50 (3):930–944, May 2005. CODEN LIOCAH. ISSN 0024-3590.

Yahel:2002:DCS

[YYG02] Ruthy Yahel, Gitai Yahel, and Amatzia Genin. Daily cycles of suspended sand at coral reefs: a biological control. Limnology and Oceanography, 47(4):1071–1083, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Yasuhara:2007:ELT

[YYTH07] Moriaki Yasuhara, Hideo Yamazaki, Akira Tsujimoto, and Kotaro Hirose. The effect of long-term spatiotemporal variations in urbanization-induced eutrophication on a benthic ecosystem, Osaka Bay, Japan. Limnology and Oceanography, 52(4):1633– 1644, July 2007. CODEN LIOCAH. ISSN 0024-3590.

Ziegler:2000:ESR

[ZB00] Susan Ziegler and Ronald Benner. Effects of solar radiation on dissolved organic matter cycling in a subtropical seagrass meadow. Limnology and Oceanography, 45(2):257–266, March 2000. CODEN LIOCAH. ISSN 0024-3590.

Zibordi:2001:RBF

[ZB01] Giuseppe Zibordi and Jean-François Berthon. Relationships between Q-factor and seawater optical properties in a coastal region. Limnology and Oceanography, 46(5):1130–1140, July 2001. CODEN LIOCAH. ISSN 0024-3590.

Zaneveld:2003:IBM

[ZB03a] J. Ronald V. Zaneveld and Emmanuel Boss. The influence of bottom morphology on reflectance: Theory and two-dimensional geometry model. Limnology and Oceanography, 48(1 (part 2)): 374–379, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Zimmer:2003:BEA

[ZB03b] Martin Zimmer and Silvia Bartholmé. Bacterial endosymbionts in Asellus aquaticus (Isopoda) and Gammarus pulex (Amphipoda) and their contribution to digestion. Limnology and Oceanography, 48(6):2208–2213, November 2003. CODEN LI-OCAH. ISSN 0024-3590.

Zimmerman:2002:SGR

[ZC02] Andrew R. Zimmerman and Elizabeth A. Canuel. Sediment geochemical records of eutrophication in the mesohaline Chesapeake Bay. Limnology and Oceanography, 47(4):1084–1093, July 2002. CODEN LIOCAH. ISSN 0024-3590.

Zohary:2000:RND

[ZGP^{+}00] Tamar Zohary, Hans Güde, Utsa Pollingher, Bina Kaplan, Riki Pinkas, and Ora Hadas. The role of nutrients in decomposition of a thecate dinoflagellate. Limnology and Oceanography, 45(1): 123–130, January 2000. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2000:ETM

[Zha00] Jing Zhang. Evidence of trace metal limited photosynthesis in eutrophic estuarine and coastal waters. Limnology and Oceanography, 45(8):1871–1878, December 2000. CODEN LIOCAH. ISSN 0024-3590.

Zemmelink:2006:DMS

[ZHFD06] H. J. Zemmelink, L. Houghton, N. M. Frew, and J. W. H. Dacey. Dimethylsulfide and major sulfur compounds in a stratified coastal salt pond. Limnology and Oceanography, 51(1): 271–279, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Zimmer:2006:NEF

[ZHL06] Kyle D. Zimmer, Brian R. Herwig, and Leah M. Laurich. Nutrient excretion by fish in wetland ecosystems and its potential to support algal production. Limnology and Oceanography, 51(1): 197–207, January 2006. CODEN LIOCAH. ISSN 0024-3590.

Zollner:2009:EZM

[ZHSJ09] Eckart Zöllner, Hans-Georg Hoppe, Ulrich Sommer, and Klaus Jürgens. Effect of zooplankton-mediated trophic cascades on marine microbial food web components (bacteria, nanoflagellates, ciliates). Limnology and Oceanography, 54(1):262–275, January 2009. CODEN LIOCAH. ISSN 0024-3590.

Zimmerman:2003:BMI

[Zim03] Richard C. Zimmerman. A biooptical model of irradiance distribution and photosynthesis in seagrass canopies. Limnology and Oceanography, 48(1 (part 2)):568–585, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Zawada:2003:CFC

[ZJ03] David G. Zawada and Jules S. Jaffe. Changes in the fluorescence of the Caribbean coral Montastraea faveolata during heatinduced bleaching. Limnology and Oceanography, 48(1 (part 2)): 412–425, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Zinser:2007:ILT

[ZJC^{+}07] Erik R. Zinser, Zackary I. Johnson, Allison Coe, Erdem Karaca, Daniele Veneziano, and Sallie W. Chisholm. Influence of light and temperature on Prochlorococcus ecotype distributions in the Atlantic Ocean. Limnology and Oceanography, 52(5):2205–2220, September 2007. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2002:VSF

[ZLL^{+}02] Xiaodong Zhang, Marlon Lewis, Michael Lee, Bruce Johnson, and Gennady Korotaev. The volume scattering function of natural bubble populations. Limnology and Oceanography, 47(5): 1273–1282, September 2002. CODEN LIOCAH. ISSN 0024-3590.

Zajac:2003:RIP

[ZLP^{+}03] Roman N. Zajac, Ralph S. Lewis, Lawrence J. Poppe, David C. Twichell, Joseph Vozarik, and Mary L. DiGiacomo-Cohen. Responses of infaunal populations to benthoscape structure and the potential importance of transition zones. Limnology and Oceanography, 48(2):829–842, March 2003. CODEN LIOCAH. ISSN 0024-3590.

Zehr:2007:ELN

[ZMJ^{+}07] Jonathan P. Zehr, Joseph P. Montoya, Bethany D. Jenkins, Ian Hewson, Elizabeth Mondragon, Cindy M. Short, Matthew J. Church, Andrew Hansen, and David M. Karl. Experiments linking nitrogenase gene expression to nitrogen fixation in the North Pacific subtropical gyre. Limnology and Oceanography, 52(1): 169–183, January 2007. CODEN LIOCAH. ISSN 0024-3590.

Zugger:2008:OSP

[ZMK^{+}08] Michael E. Zugger, Alicia Messmer, Timothy J. Kane, Jennifer Prentice, Brian Concannon, Alan Laux, and Linda Mullen. Optical scattering properties of phytoplankton: Measurements and comparison of various species at scattering angles between 1^{o}and 170^{o}. Limnology and Oceanography, 53(1):381–386, January 2008. CODEN LIOCAH. ISSN 0024-3590.

Zeller:2006:CPG

[ZRL06] Martina Zeller, Thorsten B. H. Reusch, and Winfried Lampert. A comparative population genetic study on calanoid freshwater copepods: Investigation of isolation-by-distance in two Eudiaptomus species with a different potential for dispersal. Limnology and Oceanography, 51(1):117–124, January 2006. CODEN LI-OCAH. ISSN 0024-3590.

Zepp:2008:STV

[ZSS^{+}08] Richard G. Zepp, G. Christopher Shank, Erik Stabenau, Karen W. Patterson, Mike Cyterski, William Fisher, Erich Bartels, and Susan L. Anderson. Spatial and temporal variability of solar ultraviolet exposure of coral assemblages in the Florida Keys: Importance of colored dissolved organic matter. Limnology and Oceanography, 53(5):1909–1922, September 2008. CO-DEN LIOCAH. ISSN 0024-3590.

Zhang:2003:BRM

[ZVRL03] Hao Zhang, Kenneth J. Voss, R. Pamela Reid, and Eric M. Louchard. Bidirectional reflectance measurements of sediments in the vicinity of Lee Stocking Island, Bahamas. Limnology and Oceanography, 48(1 (part 2)):380–389, January 2003. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2006:SSM

[ZW06] Li Zhang and Wen-Xiong Wang. Significance of subcellular metal distribution in prey in influencing the trophic transfer of metals in a marine fish. Limnology and Oceanography, 51(5): 2008–2017, September 2006. CODEN LIOCAH. ISSN 0024-3590.

Zou:2004:BRF

[ZWC^{+}04] Li Zou, Xu-Chen Wang, Julie Callahan, Randolph A. Culp, Robert F. Chen, Mark A. Altabet, and Ming-Yi Sun. Bacterial roles in the formation of high-molecular-weight dissolved organic matter in estuarine and coastal waters: Evidence from lipids and the compound-specific isotopic ratios. Limnology and Oceanography, 49(1):297–302, January 2004. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2004:TWI

[ZWHP04] Jinzhong Zhang, Feiyue Wang, James D. House, and Bryan Page. Thiols in wetland interstitial waters and their role in mercury and methylmercury speciation. Limnology and Oceanography, 49(6):2276–2286, November 2004. CODEN LIOCAH. ISSN 0024-3590.

Zafiriou:2008:DCM

[ZXN^{+}08] Oliver C. Zafiriou, Huixiang Xie, Norman B. Nelson, Raymond G. Najjar, and Wei Wang. Diel carbon monoxide cycling in the upper Sargasso Sea near Bermuda at the onset of spring and in midsummer. Limnology and Oceanography, 53(2): 835–850, March 2008. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2001:MVP

[ZZPP01] Meng Zhou, Yiwu Zhu, Shawn Putnam, and Jay Peterson. Mesoscale variability of physical and biological fields in southeastern Lake Superior. Limnology and Oceanography, 46(6):679– 688, May 2001. CODEN LIOCAH. ISSN 0024-3590.